A Complete Bibliography of Lecture Notes in Mathematics (1980–1984)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

29 November 2023 Version 1.11

Title word cross-reference

&, ν, ⊃, λ, ∀, ϵ [SH84]. (0, ∞) [Ani83]. (1, 1) [MT84]. (1 − 1) [CB84a]. (2 + ϵ) [SC81j]. (2p + 1) [And80]. (4q + 2) [Kos80a]. (A − λB) [Kåg83]. (a_{2}, a_{3}) [Tam82e]. (dd^{c})^{n} [Bed82]. (e, M_{ω}^{∞}, p) [Die83f, Die83k]. (e, T_{ω}^{∞}, p) [Die83f, Die83j]. (G, σ) [KN80a]. (k; l) [Kwa83]. (m, n) [Laj81]. (n = 6, 7, 8) [Kim81]. (v, k, μ) [Ead80]. ∗ [Kan84a]. −ϵU′′ + U′ = F [Rei82b]. 0 [Miy81c, Obe82, Orb84, Ple83i]. 0 < p < ∞ [Sem83]. 0 < p ≤ 1 [Pop83b]. 0′ [EHK81]. 1 [Abd83, Arn82q, Cer81, Cho84e, CJ84, Gir83a, Mic83, Miy81c, Obe82, Ple83a, Ple83i, Sot82f]. 1 + k2^{N} [Sta81]. 1/2 [Gal80]. 10 [HB84, LTS82]. 10^{5} [CR84a]. 11 [MM83]. 15 [Her82a]. 1 < λ < ∞ [ESV83s]. 1_{p} [Leś84]. 2 [Arn82d, Bas80, BS84, Cha81c, Cos84, Hei84, Hem82a, Hen80c, Kea82, Küh80, Mar83g, Miy81d, Neu82b, Ple83h, Shi81a, ZVC80a]. 2 + 1 [Chr80b]. 20 [BS81]. 28 [Bro81e]. 2ème [Sch82n]. 2: 3 [Tam82j]. 2n [Bec82a]. 2n + 1 [Bui84]. 2 × 2 [Har80b, Küm84a]. 3 [Ber83b, Cer81, Esn81, Fic80, Gal83b, Kas82b, Roc84, Sch81b, Sie80a, Som82a, Spe81b, Sul81b, Zie81d]. 3 + 1 [Chr80b]. 4 [BM81b, Cat83b, FK83b, Kre84a, Lan80d, Lev80f]. 44 [GRW81]. 9 [Bro81e]. [0, 1]^{1+k} [Har80a]. ]0, ∞[ [CG83a]. ^{−} [Maj83]. eL^{2}(Γ/G, X) [Bru81n]. ^{∞} [Pro81]. ′ [CJ84]. ^{s}E^{∗} [Bau81]. _{1} [DeV84, Mek81]. 1F_{1}(1; c_{i}; z)}_{i=1}^{n} [dB84b]. _{2} [Dev81]. _{i} [Lev80c]. _{K}Ban_{1} →_{K} Vec [Hof82b]. _{sl} [Her84b]. A [Alb83a, Alb83b, Sot82b, Sot82g, Tri84]. A(t) = a(t)A [Gek84e]. A − λB [Kub83]. A^{∗}(I) [McA83b]. A^{∗}(I) − B^{∗}(I) [McA83a]. A^{∗} = Ā^{∗} [McA83l]. A^{ϵ} [Tri84]. A^{k}(D) [Jak80a]. A^{m}(D) [BO84]. A^{N} − 1 [Lün81]. a_{1} → −1/4 [JM84]. a_{3} + λa_{2} [Tam82h]. a_{∞} [Igu82]. A_{n} [BL82, Roc84]. A_{o} [Sch84h]. A_{p} [Jou83a]. ℵ [Dev81, Mek81]. α [Cho84a, Fri81, She82a]. B [CR83, Leg82c, Leg82a, SB80]. B(g, s) = (λg, λs) [Gru82a]. B(H) [Pas80]. B^{∗}(I) [McA83b]. Ā^{∗} [McA83h]. Ā^{∗}(I) [McA83c]. Ē^{∗} [Bau81]. ¯∂ [Lie80, Tal83]. ¯∂u = f [Las81a]. because [Bea82b]. ▽ [Maj84b]. BM(R^{3}) [PRW84]. BMO(R^{n}, dx) [Jou83b]. modm [Jos82]. Br(X) = Br′(X) [Hoo82]. C [Duv83a]. C(X) [Gie82r, Gie82a, Gie82u, Sem82c, Sem82d]. C(X; E) [Sch83a]. C∗ [CS82a, Cun83, Mun84, Sal82a]. C^{∗} [Ren80a, Ren80b]. C^{1,α} [Jos84b, Jos84a]. C^{2} [FY81, Sko80a]. C^{2,α} [Jos84j]. C^{∞} [CM81, DRR81c, Mal80c]. C^{∞}(ω) [Sag80]. C^{n} [Kan83l]. C^{O+α} [Bem80]. c_{1}^{2} = 3c_{2} [Ish83]. c_{o} [Hei83]. C_{P} (X; E) [Sch83b, Sch83c, Sch83f]. c_{x}^{p} [Ang81j]. cents^{n} [Hen80c]. CL(n) [Kim81]. Ct(k) [Nor80d]. Ct(n) [Nor80b]. P_{χo}Z [IS81b]. D [Lau83b]. D[0, 1] [Daf81]. D[0, ∞) [Kal83]. D_{λ}, K [Gre80d]. D_{n} [Van83a, Roc84]. Δ [Lev80c, Lan82b]. Δ_{∗} [Kre84g]. diff^{ω}(R^{n}) [Mas84a]. DiffΣ^{n} [RP80]. ẋ(t) = −f(x(t − 1)) [Pet81]. e′(P) [Sch83n, Sch83j, Sch83s]. E_{n} [Roc84]. ℓ_{1} [RU82]. Ext(A, Z) [SS81a]. EXT^{n} [Gui81]. F [Jun83, Mar83f, Jon82, Las81a]. F (1, 2) [KT84]. F (L^{P}) [Cór80b]. F_{σ} [Jun84]. \frac{∂}{∂x_{n}} [Van83a]. G [Bar83d, Bru81q, Goo83d, Gui80a, Poy83, Bre82b]. G/H [FJ81]. G/P [Aky82]. g = 1 [Nam80c]. G^{X} [Gui80a]. G_{∞} \ G [Bum84h]. G_{m}(w) [ESV83f, ESV83i]. G_{O}(Rπ) [Bas81]. Γ [DD83, Duv83b, Gie82a, Maj83, MS83c, dG84a, Bat81]. Γ ⋂ P/X [Sch83j]. ΓX [Sch83s]. ≥ 1 [Lau83c]. GL(2) [Bum84b, Car81b]. GL(3) [Bum84k, Fli82f]. GL(3, R) [Bum84c]. GL(N) [Ber83c, Jac83b]. GL(n, q) [Sri81]. GL_{n} [Gre80e, Gre81b, Hum80b, Hum80a, Sch80a, SP80, Vas81, Sil82]. GL_{n}(D) [Reh80]. GL_{n}(K) [Gre80f]. H′ [Jeu80f, BS80b, BBC80, BBCH81a, BBCH81c, BBCH81b, BBCH81e, BBCH81f, BBCH81g, BBCH81i, CN84a, HZ84a, Jan83b, Pop83a, SM81, Hau80b]. H^{−1} [Tho84b]. H^{1} [Bak84d, BHL81, Jon84, ND83, Ste83c, Tho84b]. H^{1}(T) [Smi83]. H^{∞} [Bou83b, Est83b, Jon84]. H^{P} [Cha82a, CMT^{+}84, Gun80a, Gun80b, Sem83]. H_{σ}^{τ} [Bru81m]. Ĵ [MR84a]. I [Hwa81]. ID_{ν} [Buc81b, Poh81b]. ∞ [Miy81b]. ∫ β × dβ [PRW84]. IP^{4} [Bea82b]. J_{m}(w) [ESV83g]. K [Bak84a, BJ80, Con81b, DW81, DeV84, Del81h, Den82a, Den82b, DK82c, FHM82, FJ81, FS81e, Gil81d, Gra82e, GWL81, Hil82b, HPR84, Igu82, Kas81, Kat83, Kna83a, Liu80, Mal84, Ore82, She82r, Sna82, Sna84, Sou81a, Sou82, Sus82b, Vog84, Wal82a, Wal84, Col80b, Kah82, Köh81, Sot82d, Sot82g, Sot82k, Sus82a, Wal83b, War82k, vOV81c]. K(a_{n}/1) [Gil82a, Gil82b, JM84, JR82, Rei82a]. K(a_{n}/1) [Jac82a]. K(π, n) [Leg82d]. k(x, σ, δ) [Cau80b]. k∗ [Kon82a]. K3 [Le 83b, Loo81a]. k[x, σ, δ] [Cau80b]. k[x, y] [DF82c]. K′ [She81d]. K_{∗}(ZZ/p^{2}) [Pri81]. K_{−i} [Ped84a]. K_{0}(Y)_{ϵ} [Cha83e]. K_{2} [Bak84b, Blo81a, Kol82b, Sti81, vdK81]. K_{2}(o) [Hur82c]. K_{2}F [Bro82c]. K_{3} [DI81b, Gra81b, Gui82b, Sti81]. K_{3}(Δ/I) → K_{2}(Δ, I) [Lod81]. K_{3}(Z[∈]) [Kas82b]. k_{4n,4n} [Lit82]. K_{6} [Mäk80d]. κ [Kec81c, Wal83a]. KK [Cun83]. L [Bum84a, Bum84g, Chr83e, Chr83d, Dev81, Don81, GPS83, Sch84k, Was81, MV80]. L(B_{t}, t) [Bar82b]. L(R) [MS83a, Ste83a]. L^{1} [Cel84, Don82d, Ros83b, Bad83]. l^{1}(ω_{n}) [Tho83]. L^{2} [Bro82a, Dem82b, Kni83, Sil83]. L^{2}(Γ/G, X) [Bru81a]. l^{∞} [Str83b]. L^{P} [MZ83, CD80, Che81a, Dom83c, Don82f, Gre83, Kan83a, Pra83, SC81h, SW83, Obe83]. L^{p}(R) [Dom83b]. L^{p}Z [SC81g]. L^{q} [Obe83]. L_{1} [AJ83, HHM83, BC82a]. L_{1}(μ) [AJ83]. L_{2} [Pit81]. L_{2}(ΓH, χ) [Hej83b]. L_{B}^{P}(R^{n}, dμ) [Jou83i]. L_{g} [Gru82u, Gru82c, Gru82d, Gru82q, Gru82f, Gru82u]. L_{∞_{ω_{1}}} [Kue81a]. L_{p} [Zaf82, Wat84b]. λ = 1 [PV81]. λ_{1} [Gal81]. ⟨X, X⟩ [Str82a]. lim a_{n} = 0 [Gil82a]. L log [JS82b]. M [Beh81, Gie82b, Gie82o, Hen83, Atk82, Kov83]. M(G) [Mél80]. M(x) [DM81e]. M_{0} [CdS81b, CdS81i]. M_{12} [Bue82]. M_{7} [SSW83]. M_{o} [CdS81c]. C [Sch80w]. C((t)) [Slo82]. C∗ [Som82b]. C0 [Meu82]. C^{2} [CB84a]. C^{N} [Ava82, AFHN84, BB84, Gav82, Gru80, Gru83, LT84c, Pen84, Tay83]. CP^{1} [Bur82]. CP^{n} [Bur82]. N [Ger84, Mic81a]. N × N [Led81a]. P^{2} [DM84, Mar83g]. P^{2}(R) [Knu82]. P^{n} [Mar83b]. P^{n}(C) [Dol83a]. P_{3} [Eis81]. P_{n} [Sch80n]. R [For84a, Hum80a, Dom83c]. R^{+} [Dom83c, Itô84b]. R^{3} [Koi84]. R^{7} [Mas84b]. R^{d} [HK80a]. R^{n} [BCJ^{+}81, Dom80, HYZ83]. Z [For84a]. L^{∞} [Bou81c]. L^{∞} [Bou81f]. L^{P} [Bou81d, Bou81g]. M [Mäk80h]. μ [Wal83a, Wal83b]. μ ∗ L^{1} [Mél80]. N [DM80c, Itô84b, Dem80c, Dyd81, KT83, Kla83a, Par83h, Sig83i]. N(t, k, v) [SM80]. ≠ 2 [Roc84]. Nϵ(T V) [Itô84b]. n ≥ 2 [Dom80]. N ∉ B [Itô84b]. ν [Lel82a]. O [Hop83k, Kis81b]. ω [Gie82b, Gie82s, Mur82b]. ω_{1} [Kue81a, EM83a, EM83b]. ω_{μ+1} [Buc81a]. ω_{n} [Tho83]. ωWh^{Diff}(∗) [Bök84]. P [Ric81d, Sch83n, Alm80, Alm83, Arn82n, BO82, Ber80c, BC81, Ber83c, Bes81, Blo82a, BF83f, Caj81d, Car80a, Car83a, CN84a, Dia84, Dix81, EH81b, Fon83, Gök83, Hub83a, Hub83b, Hum80b, Hür81, Kas82b, Lag81, LS83a, LMW80, Lin84a, LP81, MR83, Par83h, Ple83f, Pop83b, PR84g, PR84a, PR84b, PR84c, PR84f, PR84h, Pro81, Rod81, SS83, Sch82k, SC81b, SC81k, SC81l, SC81f, SC81m, SC81s, SC81t, She82l, Smi81, Sou81b, Suw83, Tah82, Was81, Wil81e]. P (ω) [She82i]. p = 2 [Gat84]. p > 0 [Del80b]. P^{2} [Dem80c, Dem80c]. P^{n} [Bui84]. P_{m}(C) [Kla83b]. P_{n+1}(C) [Kla83a]. ∂ ¯∂ [Lel80]. Π [Lev80g, Lev80d, Lev80q]. π(x) [LO84]. Π_{2}^{1} [Mar81h]. p ≠ 2 [Kas82b]. P ω [Tan81a]. P ω_{1}(λ) [Woo83a]. £ [ESV83i, ESV83a, ESV83r]. PSL(2, 13) [Meu82]. PSL(2, F_{q}[t]) [GvdP80g]. PSL(2, R) [Hej83c]. Q [Arn82b, Arn82i, Arn82l, Arn82p, Arn82n, Gro80h, Jos83a, KMS83, Kle84e, Kle84j, Kle84k, Wil82a, Erm80a, Lie80, Ple83h, SC81q, SC81s]. q = −p_{a} [Suw83]. Q_{m}(w) [ESV83f, ESV83i]. R [Dom83c, Lyu84, Mon80i, Pre83i, LR80b]. R^{−} [Duv83a]. R^{2} [Par83e]. R^{3} [DL82, Ioo80, Par83e]. R^{d} [Kos83a]. R^{G} [Mon80g, Mon80i]. R^{G} → R [Ale83b]. R^{∞} [Leś84]. R^{m} [RRM84]. R^{n} [Car83a]. ℜ(a_{3} + λa_{2}) [Tam82b]. S [BL80b, GL81, Joh81a, Kan83d, Lue81, Pre83d, Pre82, Que81, Web84b, Cha83i, CC82e, CS84b, Edm82, KJTB80, Kow80c, Kow80d, Kow80j, Kow80k, Kow80l]. S(b) [Tam82b]. S + ζV [Kan83c, Kan83d]. s = 0 [Was81]. s = 1 [Was81]. S^{1} [Kre84g, Sie81b, Sul81b]. S^{3} [Mor81b]. S^{3} × R [Sie80b]. S_{R} [Tam82c, Tam82g]. S_{R}(b) [Tam82h, Tam82i]. Σ [Bre83b, Ban81a, Chu82a, Grz82, Luk80b, Nic83, Whe80]. σ = 1/2 [Gon81]. Σ_{1}^{1} [Ste83b, Wil81d]. Σ_{n} [Kau81a]. Σ_{o} [Wil80a]. SK [Kre84c]. SK_{1} [Dra80e, Dra80c, DK80b, Dra80a, Dra80b, Oli81]. SL(2) [Del80a]. SL(2, IR) [CLM84]. SL(2, C) [Bar82c]. SL(2, C)^{∗} [AAS83]. SL(2; C) [EF81c]. SL(2; C) [Fal84]. SL(2; R) [EF81c]. SL(n, R) [Spe81a]. SL_{2} [SS83]. SL_{2}(Z) [Cas83b]. SL_{2}(q) [Ple83h]. SL_{3}(C[X]) [vdK82]. SL_{3}Q [Sch83k]. SL_{n} [Hum80b, Hum80a, Stu82]. \sqrt{|a_{n}|} + \sqrt{|a_{n−1}|} < 1 [Gil82b]. SU(1, 1) [EF82].

SU(N, 2) [KS83c]. ∑_{i=1}^{n} λ_{i}x_{i} \frac{∂u}{∂x_{i}} = F (x, u) [Cha83t]. t [Köh81]. τ

[BS80a, Mur82b]. Θ [Lau83a]. θ_{α} [Bru81j]. U(g)^{K} [FJ81]. U_{6} [Mäk80m]. U_{m}(w) [ESV83i]. U_{R} [Mäk80h]. \underline{lim (i)} [GJ81]. V [Jaj80]. V^{m} ⊂ P^{m(m+3)/2} [Gat84]. V_{h} [Saa84b]. V_{λ,K} [Gre80b]. φ.(ω, .) [Bis81n]. W^{k,1}(R^{n}) [Poo83]. Wh(y)^{ϵ} [Cha83f]. ξ_{A} [Mäk80g]. y^{(n)} + py = 0 [Jon81]. Z [Lev80w, Nel81, War82h]. z/pz [Alm83]. Z^{n} [GLP81]. Z_{p} [Köh82]. ζ [Szu83]. |G| [Mon80g]. |Q_{m}(w)| [ESV83c]. |U_{m}(w)| [ESV83c]. |w| = 1 [ESV83g].

* [Ima83]. *-semigroups [Ima83].

-a [KJTB80, Jos84a]. -a-priori [Jos84j]. -absolutely [LMW80, Kis81b]. -additive [Whe80]. -additivity [BS80a]. -adic [BF83f, Dia84, Hum80b, Lag81, LS83a, Ple83f, PR84a, PR84b, SS83, Sou81b, Was81]. -adically [PR84g, PR84c, PR84f]. -adique [Bes81]. -adiques [Fon83, LP81, Rod81, Tah82]. -algebra [Sot82d, Ren80a]. -algebras [Sch82k, Sot82g, Ren80b, Arn82b, Arn82i, Arn82l, Arn82p, Bad83, Cun83, Mun84, Sot82b, Sot82g]. -analytic [Kra80i]. -analytique [Sch80w]. -Basic [Pro81]. -block [Ple83h]. -body [Sig83i]. -bounded [Tri84, Web84b]. -Calculus [Lau83a]. -Capacity [Car80a, Car83a]. -Cobordism [Cha83i]. -Cobordisms [Hau80b]. -coefficient [Atk82]. -complète [Abd83]. -concave [Erm80a, Lie80]. -Cone [BBCH81e, Pop83a]. -Cones [BBCH81a, BBCH81c, BBCH81b, BBCH81f, BBCH81g, BBC80, BBCH81i, Jan83b]. -connectivity [Dyd81]. -continuous [Ban81a, Nic83, Nel81]. -convergence [DD83, dG84a]. -convex [Gie82u]. -convexity [SC81s]. -convolution [MZ83]. -coverings [SB80]. -cube [KT83]. -curve [Gro80h]. -D [Hem82a]. -decomposable [Jaj80]. -decomposition [MS83c]. -definability [Wil80a]. -demi-groupes [Bat81]. -Derham [Maj84b]. -designs [Köh81, Neu82b, SM81]. -Difference-cycles [Köh81]. -differentials [Sot82f]. -dimension [Caj81d]. -Dimensional [ZVC80a, Bas80, Chr80b, Fic80, Gal83b, Roc84]. -element [Ber83b]. -Elementarily [Kue81a]. -équivariantes [Sal82a]. -estimates [Jos84b]. -extensions [KN80a]. -Factorization [SC81q]. -fibres [Leg82a]. -fields [Grz82]. -finite [Cho84a, FJ81]. -folds [Som82a]. -Formes [Cer81]. -Fraction [Lan82b]. -function [Mur82b]. -functional [Mal84]. -functions [Bre82b, ESV83i, ESV83a, ESV83r, Tam82g, CM81, GPS83, ND83, Was81]. -Gauss [SC81k, SC81f]. -generic [Cho84e, CJ84]. -groupes [Que81]. -Groups [Hub83a, Smi81, GRW81, Hub83b, Kea82, Kle84e, Kle84j, Kle84k, Kre84c]. -halb- [BKK81]. -harmonic [Mar83f]. -hereditary [MV80]. -homology [Sna84]. -Ideal [Laj81]. -ideals [Hen83]. -inequalities [Che81a]. -integral [SC81l]. -invariant [Ber83c]. -invariants [Kah82, Ped84a]. -kernels [Kwa83]. -lifting [BS80b]. -local [CN84a]. -M [Hwa81]. -manifolds [Kos80a, Kow80c, Kow80j, Kow80l, Kre84a, Zie81d]. -matrices [CC82e, Par83h, Wil82a]. -meromorphic [Kov83]. -metrizable [Jun84]. -models [Chu82a]. -module [Car83a, Mon80i, Van83a]. -modules [BL82, Gie82r, Lau83b]. -Near [Pre82]. -numbers [Edm82]. -only [Lev80d]. -operators [Kan83l]. -pencils [Kåg83]. -Pietsch [SC81j, SC81m]. -Poincaré [Maj84b]. -Points [She82l]. -preduals [HHM83]. -primary [Lev80g, Lev80q]. -principal [Poy83]. -Products [Wal83b, Wal83a]. -Projective [Alb83a, Alb83b]. -projectivity [Jon82]. -Properness [She82a]. -radical [Con81b]. -Radonifying [SC81b]. -ranks [EH81b]. -RE [Fri81]. -regular [Maj83]. -Regularity [Kan84a]. -results [Mur82b]. -ring [Igu82]. -rings [Joh81a]. -Rule [Buc81a]. -semigroups [Bem80]. -semistable [LR80b]. -Separable [EM83a, Mek81, EM83b]. -sequence [Lyu84]. -series [Bum84a, Bum84g, Chr83d]. -simple [Arn82n, Arn82n]. -smoothness [SC81s]. -Souslin [Kec81c]. -space [BL80b]. -spaces [Gie82b, Gie82s, Jun83, Pop83b, Gre83, Bou81d, Bou81c, Bou81f, Bou81g, CR83, CN84a, Don82f, HZ84a]. -spheres [Sie80a]. -Spline [Gru82c, Gru82d]. -Splines [Gru82q, Gru82f, Gru82u, Gru82u]. -Stable [LMW80, Lin84a]. -stochastic [Kni83]. -structure [Beh81, Gie82o, Gre83]. -structures [Kow80d, Kow80k, Bre83b]. -submodules [Gie82a]. -summing [SC81t]. -systems [Lue81]. -th [Bec82a, FK83b]. -Théorie [GWL81, Kas81]. -theories [HPR84]. -Theory [Den82a, Den82b, FS81e, Bak84a, Sus82a, KMS83, She81d, BJ80, Cun83, DW81, DK82c, FHM82, Gie82b, Gil81d, Hil82b, Igu82, Kat83, Liu80, Ore82, She82r, Sna82, Sna84, Sou81a, Sou82, Sus82b, Vog84, Wal82a, Wal84]. -torsion [Kas82b, Mon80g]. -torsion-free [Lev80w]. -transformations [Bar83d]. -transforms [Sch84k]. -transitive [Küh80]. -trees [Dev81]. -type [Kna83a, SC81g]. -types [Del81h]. -uniformly [Hei84]. -uples [Kla83a]. -valent [Gök83]. -valuations [PR84h]. -valued [Don82d, Hei83]. -varieties [Bar82c]. -variety [vOV81c]. -vector [BS84].

10th [Gri84]. 1956 [Hop83d]. 1966/67 [AY81]. 1967/68 [Ano80, Ano81b, Ano81c]. 1976 [DK80b]. 1977 [DPT80, Men80a]. 1978 [Ano80, AY80, AY81, BLS80, Eve80b, Flo80j, Hen80d, JPW81, LS80, Mac81c, PPV80]. 1978/79 [Ano80, Ano80, AY80, AY81, LS80]. 1979 [Ano81b, AV80, BH81a, Ben80a, BFF80, BMR81, DG80d, DG80c, FKSW81, GPRS80, HN81a, Hen81a, Izé80, JP81, KM80, Köl80, KN80b, Ław80, Mar80b, NR80, Nor80a, PWW80, RY81, Rao80, Rau80, RSW80, Wat80, Wer80]. 1979/80 [Ano81b, AY81, Ano81b, BMR81, HN81a, RY81]. 1980 [AGP81, Ama81, Ano81a, Ano81c, AL81, AL82, AY82a, Bec81, CC82f, DAP82, DP82b, DLR81, Eel82, ES81, FF81, FS81e, Hen82c, Hin82, KO81a, KKS82, Kno81, KMS81, LW81, LST81, McA81, MTT81, Rao81, Ric81c, Rog81d, Wil81b, dBvR81, vO80]. 1980/81 [Ano81c, AY82b, LS82a, Ano81c, AY82a]. 1981 [AJ81, All82b, AD83b, ABGM82, BBC^{+}83, Ban82, BOWS82, BS83e, BS83f, BKS83, BKK82, Car82, CBJS83a, CBJS83b, CZ83a, Cer84a, CW82, CTCMR82, Con82b, Dol83b, EdJ82, FMSW83, Fle82c, Fuk82a, GP83a, GW81, HT82a, Hen82a, Hey82, HP84h, HJW83, JTW82, KPT82, KMS82, LR83a, LDS83, MS81a, MdJ82, MM82b, Pal83, PPS83, RW82, Tur82, Wat82, Win82a, dFH82, vOV82]. 1981/1983 [LDS83]. 1981/82 [AY83, HP84h]. 1982 [AFG84, AMT82, BP84a, Bak84a, BJ83a, BDM83, CV83, Cas83a, CCCN84, CGO83, ES82, EL83, FM84, FG83, Ghe83, Hen84a, IP83b, JV82, KR83a, KZ83a, KS83d, Ław83, MO84, MRW83, Nav84b, PR83, RS83a, Ves83, Vin84b, dV84]. 1982/83 [AY84]. 1983 [AGT84, CP84b, Eri84b, GLM83, GMSV84, GS84, Gri84, HLR83, HKLR83, HJLR84, Hey84, HN83, Jag84, KS84a, KN84, KY84, KMS84, Kom84b, LDS83, MP84b, Sal84a, Sal84b, SW84, TW84, WB84]. 1984 [Ken84b]. 1st [Cer84a, Ghe83, GS84, dFH82].

2nd [Sal84b]. 2nd1981 [Con82b].

30 [JW83]. 32ème [Mal80b]. 33ème [Mal81a]. 34ème [Mal82c]. 35ème [Mal83]. 3rd [MM82b, dV84].

4th [DP82b].

6th [KZ83a].

78 [GIR81d]. 79 [Ano80, AY80, AY81, KMM81, LS80].

80 [Ano81b, BMR81, HN81a, RY81]. 81 [Ano81c, AY82a, KMM83a]. 82 [HP84h]. 8th [Wat80].

9th [Wat82].

a-posteriori [Rei82b]. A-priori [Jos84b, Jos84c]. A.

[Cór80a, Gra83d, Mar81i, Mer80a, Mer80b, Pis80, SM83b, Web84a]. A.-D. [Mar81i]. Aachen [BOR^{+}84b, MR84b]. Aarhus [MO84]. Abbildung [phAD80a]. Abbildungen [Dit80, Küh83a, Küh83b, Woh83]. Abbildungsräume [Hau80a]. abbreviations [ŁK83b]. Abel [Bak82b]. Abel-type [Bak82b]. abelian [And84b, Băn83b, BM81b, Boi80, BR81c, BH82c, CQ81, CH84, Cor83c, Fel80, Fuc81a, Göb83a, Göb83b, Hau81, HW81c, Hub83b, LR81, Mek83b, Nam80a, Nyg83, Pas83, Pro81, Ran83b, Alb83a, Alb83b, Arn82f, BT82d, Cor83b, Del81b, Del81c, GW81, GLM83, Hau83a, Hau83b, HP84d, Hub83a, Mek83a, Met83, Ran83a, Rei81, Rei83a, Rei83b, Sch84j]. abélien [Del81f]. Abélienne [Tig81, Bea83a]. Abeliens [Bel83a, Bir83a, Bir83b, BBM82e]. Abelsche [Sch84b]. Abelscher [Sch84c, Sch84e, HP84d]. abgeschlossenen [DK82b]. Abhyankar [Orb84]. above [Kat80b]. absolue [BY80]. Absolute [Oku84a, Oku84b, Oku84c, Oku84e, Oku84d, Oku84f, Oku84g, Art80a, Fuk82b, Kus82, Oku84i, Oku84h, Ul’84]. Absolutely [Ruc82, Sch83a, Jak80b, Kis81b, LMW80]. Abstract [CP84a, Cuy84a, Da 84, Mos84a, Mun84, Ver84b, Ver84c, Ver84d, Car83b, Clé84, Cuy81, DGZ84, DKS84, HG82a, Hel80b, Kan83o, Ław82, Ver84i, Ver84j, Ver84k]. abstraites [Bro84]. abstraits [Dub82b]. Abundance [Hub83a, Hub83b]. Accélérabilité [Del84a]. Acceleration [Guz81, Kow81a, CA82, Gil82a, Jac82a, Lon81, Mag84]. according [Dre83, Gro81a, Mur82a]. accretive [HG82a]. accroissements [Jac84]. Accuracy [Ivn83, JN82, Ste84b]. accurate [Hem82a, Hem82b]. Acharya [Gil81c]. Achievement [EH84, Har83]. Acireale [GS84]. Acknowledgements [Gro80a]. acknowledgments [DRR81g]. acoustic [Col82c]. Actes [CV81, Rao80]. actin [AD83a]. Acting [Ivr84c, CRW^{+}80, Dic80e, Mon82, Van83a]. Action [BMR81, Bou82d, Chu82c, Hor80, Kon82a, Slo80d, Ull82]. Actions [Car82, Mor81b, Ben84e, BBŚ82, Bra84b, CS82a, Fis81, Fur81, Hes81, Kon82b, Kor82, Kos80b, Sco81, Slo80e, Som82b, Ung82]. acts [Kna83c, Lex83]. Acyclic [HH84]. AD [Woo83b, Woo83a]. adaptation [DKT84]. adapté [CMS83]. adaptés [Fey80b]. Adaptive [HR82, Alb80, Rei82b]. Adaptive-grid [HR82]. Addendum [Ogu81a, Ram82a]. Adding [She82a]. Additif [Bre81b]. Addition [Kre84b, Sto80a]. additional [Kre84f]. Additive [All82a, Arn82a, Arn82b, Bel83a, Nar84a, Sar84, All84, Bea83c, Erd81a, Fuk81, Hum80c, Mas82a, Nat84, Ric81a, Whe80]. additives [DFM82]. additivity [BS80a]. Adelaide [Cas83a]. adelic [Hum80b]. adequate [Gie82b]. ADI [Hor82a]. ADI-methods [Hor82a]. adiabatic [Bak83]. adic [BF83f, Dia84, Hum80b, Lag81, LS83a, Ple83f, PR84a, PR84b, SS83, Sou81b, Was81]. adically [PR84g, PR84c, PR84f]. adique [Bes81]. adiques [Fon83, LP81, Rod81, Tah82]. adjacency [SS81b]. Adjacents [Dra83i]. adjoint [BCS81, EG81, ELZ83, Fle82a, FK83b, Gir83b, Gre82a, Hön82, Mae80g, Mor80c, Mor82b, Slo80d, Slo80f]. adjointness [Dun82, EN83].

Adjoints [McD82, Fey84a]. adjugate [PV81]. adjunctions [MS82a]. Adleman [Len81]. Adler [Ver81a]. admettant [Jac84]. Admissible [Ana80, Cho84j, Bae84, ESV83a, KP84b]. admissibles [Car83c]. admitting [Ben82a, Köh82]. advanced [Kow80n]. Advances [Saz81c, LV84, Fle82c]. advantages [Bau84a]. Affine [Hil82b, Lep82, Mac81a, Pet82, RC81, Bar82c, Bre82a, Cal81, DW81, Éli84b, For84b, Fre82b, KS83a, Köh82, Kon82b, Kow80a, Kow80i, LP84a, Miy83b, Mon82, Sch80m, Ver82b, Win82b, Zel81d]. affine-ruledness [Miy83b]. affine-scheme [Ver82b]. after [Edw80, Len81, Orb84]. afterthought [Kin84b]. age [Gre84]. age-dependent [Gre84]. Aharoni [Mil84b]. Ahlfors [Sto83b]. aid [Col83, Tam82e]. aide [Jeu80b, Net82a]. aimed [AW83]. Aix [GPRS80]. Aix-en-Provence [GPRS80]. Aléatoire [Bis81k, Bis82]. Aléatoires [Fer81b, KMS81, BBP82, Éli82, Éli84b, Éme80c, Fer83a, Gui80d, Gui84a, Le 82b, Nob81, Web83a]. Alexander [Hil81a, Hil81i, Hil81l]. Algebra [Ama81, BV83a, Dra84, FG83, KPT82, PWW80, Sch82d, Zel81c, vOV81l, And81a, Aup83, Bet82, Bou83b, Cat84, Dav81a, DR80a, Dom83a, Enr83, Fos81, Gab81b, Gre80f, Gre82b, Han80, Jos83b, Lam84, Mor81c, Poy83, Ren80a, SSW83, Sot82b, Sot82d, Tho83, Ung82, Wil83b, Ama81]. Algebraic [Ama81, ABGM82, Brz81, Dol83b, FM84, Fom84a, Fra81, HJW83, Lev80e, LW81, Mat80, Miy81a, Nel82, PV83a, Rob81a, Sch80b, Sch80l, Shi83c, Slo80h, Sot82h, Tan81b, Tan81f, Tog81, Wal84, vOV81a, vOV81h, vOV82, Ben82a, BT82a, BBŚ82, Bjö81, Blo81a, Bor82b, BV83c, BV83g, BS84, CKK83, CV84a, Con83, DK82c, Duf84, FHM82, For84b, FL81, GP83c, GvdP80a, Gie82s, GM82d, Hes81, Hil82b, Igu82, Ila82, JW83, Kat83, Kon82b, Mar82b, Miy81b, Miy81c, Miy81d, Miy83b, Orb84, Osg84, PV83b, Pre84e, Ric81b, Rus84b, She82r, Sna84, Sou81a, Stu82, Sus82b, Suw83, Tan81j, Tau81, Uen82, Vog84, Wal82a, vOV81c, ABGM82, BP84a, Bak84a, CGO83, Con82b, Den82a, Den82b, Dol83b, FS81e, LW81, MO84, RS83a]. Algebraically [Dad80, Che84a]. algebraischen [End84]. algebraization [AT82e, Con84]. Algebras [Ama81, AL81, AL82, BBC^{+}83, Ben84c, BF83f, DG80d, HMT^{+}81b, MV80, MV81b, Ren80b, Rin84e, Rom83, Win82a, AS82a, All83a, And81a, AN81, Ani84, Arn82b, Arn82i, Arn82l, Arn82p, Bad83, BPV83, Bau84b, BN81, BO82, Ber83b, Blo81b, Blo82a, Boi80, BR81b, Bon81, Bon84, Bor84a, Bow82, BM83, Brz81, BF83b, CS80a, Can84, Cha80a, CC82e, Cun83, D’A84, Dal83, Dla83, DO82, EF81a, Est83a, Est83b, Fon84, Fra84, Fre83, Fre82b, FL84b, Gab81a, Gab80, GŁ84b, Gra83a, Gra82d, Gre81a, Grø83, Gur80, Gus82, HR81, Haz81a, Haz83, HMT81a, Hen84b, Hil82b, Hum80f, Hur82a, Ila82, IW80, Jaj84a, Jaj84b, JL80, Jøn80, Kac82, Kap82, Kau81c, Kha82h, Kop81, Kos81f, Kra82b, Kra80i, KS80b, Lam84, Lau83d, Lep82, LP84a, LM83, Lor81a, Loy83]. algebras [Mac81a, Mal80c, MS83b, Mis84b, Mun84, MSW^{+}84, Nel81, Oko81, Pet82, Pie83a, Pog80, Rei82c, Rin80b, Rin84b, Rin84g, RC82, Ros83a, Sal82b, Sch82k, Sel81b, Sel83, Slo80c, Sot82b, Sot82g, Sot82k, Sze82, Ton83, Van80, Wil81e, Wil81f, Wil82c, Win81, Zel81e]. Algèbre [Ioc84g, Tan83a, Lev82a, Mal80a, Mal82b, Mél80]. Algebren [Bon80, Stu80]. Algèbres [Cau80a, Gab81c, Ioc84a, Ioc84h, Tah82, Tan83e, Ver81a, Vig80a, Vig80c, Vig80b, Ale81b, Bou81a, BO84, Con81a, Ioc84b, KS83b, Lev80a, Tau80, Tig82, Yam80, Yam82, Zay82, Vig80e]. Algébrique [CTCMR82, GIR81d, GWL81, Tan83d, Ver81b]. algébriquement [Dur80, Vid80, Zay82]. Algébriques [Ang81b, GIR81d, Ang81e, Bre83e, Cos82b, Duv83a, Kah80, Ver81a, Ris82]. Algebroid [Cam80b, Cam80c, Cam80d, Cam80e, Cam80f, Cam80g, CC82b, Cas82a]. algorithm [CG83b, Cur80a, Eri83, Gol80, GN80, LS81b, Ove82b, Ruh83, Ste81b]. Algorithmes [Cor84a]. algorithmic [Can84]. Algorithmically [Bau84b]. algorithmische [Ewa80]. algorithms [Daw82, Geo81c, Hul82, Koj81, LO84, Len81, May80, Mos84a, Rin80a, Rin80b, Ste84b, ST82, Tod82, Van83b]. All-azimuth [Dra83h]. allgemeine [phAD80c, Spa83a, HP84a]. Allied [CP84b]. Allocution [Cho84l, Coe82]. Almost [BS80a, CS82a, Har81a, Pre83a, Tei81a, Tes83, BS80b, BB81a, BK81c, But81, DK82a, Dat83, Dev80, Duf84, GM80h, Jaj84a, Kan80a, Kar80a, Oku84b, Oku84c, Oku84f, Rei82c, Rob84a]. Almost-periodic [Har81a]. along [Bar83c]. already [Ara84]. alternating [Ale83a]. alternative [Eng84, Gie82c, Lau83a]. always [Joh81c]. am [HP84h]. amalgamation [Ber83a]. Amarts [GS83c, GS83d, Sch83g, Bel81b, CS81a]. ambient [Hoc81]. American [Car82, Kom84b, dFH82]. amicable [tR84]. Amorces [Sie80b]. Ampère [Ceg83, Ceg84, Kis84a, Lel83]. amphicheiral [Sie80a]. Ample [Som82a, Gol83a]. amplitude [Llo83]. Amsterdam [dBvR81]. Análise [Mac81c]. analog [AV83c]. analogs [Ul’84]. analogue [Duv83b, Fal84]. Analogues [Joh84, Gra83a]. Analyse [LS80, LS82a, Web83a, DG80b, AGT84]. Analysis [Ban82, Ben80a, BS83f, BGV84, BPK^{+}84, BFPS81, BKK82, CV81, CV83, CBJS83a, CBJS83b, CW82, CP84b, Eel82, FKSW81, Fuk82a, Goo83h, Gri84, HN81a, HHN84, Hen82a, Hur83, Ket82b, LW82, LG81d, Mac81c, Maj84c, MRW83, Nac82, Nel82, PPV80, Rei82b, RW82, Sal84a, Str84e, TW84, Tur82, Ver83b, Vin84b, Wat80, Wat82, Amb83, Att84, Bac83, Bak84a, Bey83, Ble80, Buc81b, BL81, Car80e, Ceg80, CC84b, CM80b, CK84, FS81a, FS81b, FJ84, Gil82b, Gur80, Gus83, HM83c, Hof83, Izé82a, Jer80, Jof83, KŁLS80, Kan83g, Kan83n, Kan83o, Kis81a, Kom84a, KW84, Lan82a, Lau83a, Liu84, LG81a, LG81b, LG81c, LS82b, Maj84d, McL82, Mil80a, Moe82, Mor84b, Osh81, Ped84b, Pen84, Per83b, Pis83, Poh81b, RS83c, SM83a, Ste80b, SH83, Str83b, ST82, gS83a]. analysis [TN81, Van83c, Ver83c, WSF^{+}84, War82d, Wei80]. Analytic [GvdP80a, Goo83a, Hen80c, Imm84a, Kus82, Abi80, AF84, BKS82, Boc82, Bru81b, Bum84a, CS80a, Car80b, Del81g, FR83, Has83i, Kle84b, Kom80, Kra80i, Lan82b, Las84, ŁK83a, Man82, Ram82a, Ram82b, Sch84k, Shi81b, SW80, Sig83f, Ste84a, SS84a, Ton83, Uen82, ZVC80d, JTW82, Kno81, Ław80, Ław83]. Analytical [Goo84, HJW83, Sch80v, DLR81]. analyticité [dB83]. Analyticity [Rot84l]. analytique [AT82a, Ava82, Ble82, DG83, Mil83a, Sch80s, Sch80w, Ten84]. Analytiques [AT82c, Sch80v, AT82f, Bar80d, Gav82, Gay80, Gru80, LJ80, Mil83a, Rab83, Sal82a, Sch80t]. and/or [Gea82]. Andrew [Sie80b]. angelic [Flo80g]. angle [Woh83]. angles [Wed83]. Anhang [Dra80d]. anharmonic [Vor82]. Anick [Lem81a]. Ann [Dol83b]. Anneau [Hür81, Avr82, Bea83a, Gui82b]. Anneaux [Lem81a, Lev81, Ber81c, Cha83a, CT82, Les80, Mau82, Mer80a, Que81, WK80b]. Année [Mal80b, Mal81a, Mal82c, Mal83]. Années [LS80, LS82a, LDS83]. annulation [Lel83]. Annullator [Sch84g]. Annullator-Kriterien [Sch84g]. Anomalous [FW80a, Beh82]. Anosov [BM81c, Dev80, FW80a, Goo83m]. ANR [Gau81]. Anscombe [RR80]. Anti [Clo81a, Mil81c, TV84]. Anti-basis [Clo81a]. anti-cupping [Mil81c]. anti-self-dual [TV84]. anticanoniques [Dem80e, Dem80f]. anticipating [All82d]. antiisolated [Fil84]. antinormal [SH81b]. Antwerp [vO80, vOV82]. Anwendung [Wei82a, Gru82a]. Anwendungen [Dra80f, phAD80d, Gru82g, Gru82u]. Anwendungsspektrum [Gru82b]. Aplicada [Pal83]. Apparition [Bar81a]. appearing [Wol82b]. Appendice [Del82a, Len80a, Mey82a]. Appendix [CGO84, Kle84c, vdA83]. applicable [KS83c]. Application [Die82b, Gui84b, HN80b, Joh83b, Kle84d, Kre84c, Mdz81, MLM82, Mok83, Not82c, Raf83, Vig80d, Adi80a, AP80, Ani83, Axe80, Bou82a, Bou82d, Brä82b, Bro81c, Bro83, Cam80a, Ceg80, Daw81, GHZ83, HPR80, Hey81, Kal83, MN82, Nam80a, Par82b, Ric81b, Sem83, SW83, SH83, Tau80, Bor83b, SC81m]. Applications [AFG84, BF80b, BF80a, BR80a, BDM83, Ber80a, BV83a, BV83b, BGV84, BLP83a, Cha83c, Cha83j, CC82f, Cuy84d, Dac82a, Dra83e, Dra84, EdJ82, EF81g, FM84, Flo80a, Fra83c, Huš82a, Jan83a, Jeu80a, JS84, Jos84d, KPT82, KS84a, Kom84b, Le 83a, LG81d, LS82b, Maj81b, Mer84, Nel82, PWW80, PPS83, Ple81, Ray81, Rog81d, RCC80, Sak82a, Sal84b, SJ82, Str84a, Tho81a, TW84, Wer81a, WB84, Wil80a, dBvR81, vH82, AHK82b, Att84, AS80, Aze80, BH84, Bas81, BL80a, BT82g, Bre84, Bro84, CG83b, CMZ80, CC84b, CC84a, CP81, CRSSC83, Dav81b, Day83, Dia84, DJK81, Dră83a, DH83, Eri83, FSY82, Fey84c, FL81, Gay80, Gie83b, Gil81d, Gre84, Gru83, Gru82u, Gul83, Gus81, Har81g, Har81n, Hej83d, Hof82a, HPR84, How82, Jof83, Kas84, Kom84a, ŁK83e, Laz84]. applications [Lem81b, LS81b, Mas84c, Mas81, MW84, McL82, Mil81a, Mil80b, Mil82, Mon80j, Mor80a, Mue84, Mus83g, Par83a, Pis83, Rog81b, RR80, Saa84b, SM83a, Šig83l, Str81d, Str83a, ST82, Sus82a, Wer81t, Wet83, Win81, Zel81a, Zie81d, dB83, dGdlV82, BF80b, Sou82]. Applied [Rau80, Ser84, Art84b, Cha80a, Est80, Gra84a]. Approach [Aul83, Kos81e, Mat80, Ren80b, Zel81c, Aus82, BC82a, Bel81a, Bro82d, Duf83, EG81, Eas83, Far82, Fer84, Fer81e, FN82, Fou81b, Gay84, Gir82, HS84, Hos82, Hür82b, KP84a, Kaw82, KW84, Kub83, KZ83b, Lau83a, Liu84, LMRW82a, LMRW82b, NT82, O’M82, Rog80, Ruy80, Sam80, Sch83g, Sel81a, Šig83l, Tak84]. Approaches [Kur84, Mar80b]. approximable [EHP83, Hin84]. approximant [Cuy81, GL81, Ort80]. Approximants [Cuy84d, Dra81, Dra83b, Dra83c, Dra84, Bos84, Bre81c, BvI84, Cuy81, Cuy84a, Cuy84c, Gel81a, KS81a, Kla81, Nut82, Sab84, vI84]. Approximate [Cor81a, DMPG83, Fro82a, HU83]. approximating [Ste84a]. Approximation [Bis81a, CP83b, Gek84a, GMSV84, HKD^{+}84, Kan80b, LO83, Lum80a, Mac81c, Mil80b, Sal84a, Saz81c, Sch80b, Sch80c, Sch80e, Sch80k, Sch80l, Str84g, Tho83, WB84, dB84a, dBvR81, Ani83, BD84, Bie81, Boi83, Bra84a, BK81b, BGM^{+}80, Che80a, CS84a, Cox82, DR82a, Dyn81, Gau83, God83, Gol80, GS83e, Hen83, LV84, Mas84c, O’F83, Osg84, Rav84, Riv84, Tho81a, Tog81, Ver84a, Wah82, Wal83d, Wat84b, Web81, Wuy81, dB82, dB84b, Rau80]. Approximation- [Mil80b]. approximation-theoretic [Bie81]. Approximations [Las84, Beu81, CC84b, CC84a, Cus81, Flo80b, Pol84, Rem82, Tho84l, Wil84]. approximative [Gau81]. April [Ama81, ES81, ES82, Fle82c, Hey84, KZ83a, Ław80, RW82, TW84]. Aproximacão [Mac81c]. Äquivalenz [Spa83a, Spa83b]. Äquivariante [Hau80a]. Arakelov [Fal83b]. arbitrarily [Ach84]. Arbitrary [Hop83p, Chu82c, Jos84a]. Arbor [Dol83b]. arc [Kea81, Rei84b]. Archimedean [Ber83c, Jou80, Len84]. Arcs [LJ80, LZ82, Ber84b, vS81b]. Area [Hop83f, PRW84]. arguments [FK83a]. arising [BCP83, Fuc81b, Hil82a, Noc82, Sau80]. Arithmetic [BMR81, BF83b, CC82f, Gro80b, Hum80d, PWW80, Pil81a, Pil81b, Pil81c, BEPS81, CC84a, Clo80, Clo81a, Gro80d, Gro80f, HP80, Mac80, Miy81e, Mor84b, New81, Par80, Par81, Sch81a, Wil80a, vdD80, vdK81, vdT81]. Arithmetical [NV82, Trn84, Goo84, Haf81]. Arithmetically [VV80, Sch83q, Sch83r]. arithmetically-defined [Sch83r]. Arithmetik [Sch84c]. Arithmétique [BMR81, Vig80f, Ric81a, Vig80e]. Arithmétiques [Jag84, Vig80d, Ber84f, Mor80a]. arithmetisch [Sch83k, Sch83q, Sch83r]. Arnoldi [Ruh83]. Aronszajn [She82p, Han81]. arrangements [Car81c]. array [Ket83b]. arrays [AD81, Sch83i, Tei81a]. article [Gal89, Sko80a]. artificial [vA82]. Artin [IW80, Rei82c, Smi82, Tan81c]. Artinian [DRS80]. artiniennes [Roq82]. artiniens [Cou81]. Asano [MR80]. Asian [KY84]. Aspect [GF84]. Aspects [Ano81a, BH81a, Ban82, BEY81, Bou83h, El 81, Hir84a, Tan81f, Ant80, AF84, Bet82, Gie82s, Haw84, Hem83, Jak82, STS82, Sim80, dM81a]. aspherical [Dav84]. Asplund [Bou83c, RU82]. Asscoli [Gra82c]. assertion [ESV83r]. assigned [GV83]. Associated [CdS81g, BR81a, BBCH81e, Bou83k, Bum84a, Che81b, CC82e, DM80c, ESV83i, EW83, KKM84c, KKM84b, KKM84f, Kos80b, Lic84a, NS81b, Ōsh82, Spi84, Tom82]. Associative [Mon80c, Bau84b]. associatives [Lic80b]. associativity [Tam82a]. associé [AR80a, AR80b]. associées [Air83, Com82h, Fer81c, Gér81a, Hig82a, KLM81, Lic80b, Sch80x]. Associés [Ioc84a, Ros84b, Bis81a, Bis81f, Bre83d, Éli84b, Mas83b, Sch80w]. Assorted [Par83a]. Assoziativen [Stu80]. assuming [EH81b]. assumptions [Sot82c]. Asymmetric [War82k]. Asymptotic [Aka83, Alb80, AG83b, Böh82, Doz81a, Dră83a, EG81, Eas82a, ES83, Fle81, lPBF80, Gra83c, Kin84c, Lon84, Maj84c, Maj84d, McA83d, McA83e, McA83f, McA83g, Min83a, Moe82, Obe82, SM83a, Sna82, Var84, Ver83c, Ver83b, Wal83c, Woy82, Ant83, Atk82, AF82, Brü81r, DH83, Eas83, Ell83, FR80, Gut83a, Har81h, HI83, Her83b, How82, IV80, Izé82a, Maj84e, Mil83b, PD82f, Ruy80, Wim80, Woy83, vdB83]. asymptotically [AG83b, BCF83]. Asymptotics [Dui82, Imm84a, Ivr84c, Eng83b, Ivr84a, Ivr84b, San83]. Asymptotique [Rao80, Aze82, BD80a, Cha83t, Com82d, Duv83b]. Asymptotiques [Com82e, Bar82a, Gui80d]. Atkinson [EL83]. atom [Die82b, Soa82]. atoms [Fro82a]. attachés [Sch80p]. attraction [Gin80, Gin83, HK80a, HK81, Man81a]. au-dessus [Leg82c]. Auflösungsverhalten [Nic82]. August [AL81, AL82, BP84a, Ban82, Bec81, BOWS82, Cas83a, CP84b, DG80d, DG80c, EdJ82, FM84, Fuk82a, GP83a, IP83b, KN84, KS83d, Ław83, LST81, Mac81c, MO84, Mar80b, MdJ82, McA81, Pal83, PPS83, Ric81c, RSW80, SW84, Tur82, dV84, vOV82]. Auslander [Bau80, Bün81, Gab80]. Auslander-Reiten [Gab80]. Australia [BOWS82, Cas83a, McA81, RSW80]. Australian [BOWS82, McA81, RSW80]. Austria [KS84a]. auto [Gab81c]. auto-injectives [Gab81c]. Autocorrelation [Gol82]. Autodual [BBC80]. automata [HL81, ON81, Röd84]. Automatic [Aup83, BBC^{+}83, Gea82, GG82]. automatique [Del81a]. automorphes [Gér81a]. Automorphic [Bru81c, Bru81f, Bum84b, Bum84c, Cas83b, GPS83, How81b, Bru81g, Bum84a, Bum84f, GvdP80i]. Automorphism [Mon80c, Gel83, GGHJG81, KT82b, Lan81a, She82i]. Automorphismes [Bas82]. Automorphisms [GvdP80b, Rog84b, ZVC80b, Bai81, FHV80, Gla81, Gup84, Mon80d, Mon80i]. Autonomes [DM81d]. Autonomous [Har81b, AG83b, CS83a, Cap83, Cho80b, Coo84, Cus83, Har81a, Har81c, Har81d, Par83g]. Autopolaire [Ioc84g, Ioc84i]. Autopolaires [Ioc84d, Ioc84c, Ioc84b]. autostability [Smi81]. Autour [BHL81, Cha83t]. Aux [BF80i, Vig80d, Bou84b, Bro84, Duv83a, Duv83c, Éli84b, Fer83b, Gay80, GJ81, Hel80a, Ioc84a, Las81b, Le 83a, LC80, Mey81g, Ray81, Ser80, Sou83, Com82h, Gér81a]. Auxiliary [Hey80a]. average [HM84b]. averages [ND83]. averaging [Raf83, RCC80]. avoidance [EH84]. axially [KB80]. Axiom [Kaw81, KSS81, KKMW81]. axiomatique [Las80b]. axiomatizable [Ber83a]. axiomatizations [Nic80]. Axiome [Smy84]. axioms [Gie82h]. azimuth [Dra83h]. Azimuts [Dra83h]. Azumaya [Gab81a, Sze82, Van80].

B [Dem80a, Hel80a, Ioc84a, Lei81, SM83b, Tam82c, Tam82g, Hej83f]. B. [Bar81a, Ioc84h, Mél80]. Bacharach [DM84]. Back [Hop83a, Kle82a, Meh83a, Sch80d, She82b]. Background [Bou83d]. Bäcklund [Chr80b, DM80b, DM80c, Jef80]. Bad [WB84, Bel83b]. Baer [Fuc81a]. Baire [CP83b, Kel80]. Bakaj [Gra83d]. balance [Gil81c]. Balanced [Mor80b, Ran83a, BR81c, KR83b, PV81, Ran83b]. balayage [BH80a, DJ80, Jan83b]. balayaged [Cor84b]. ball [Chr80a, Gie82h]. balls [God83, Kow81b]. Banach [BBC^{+}83, Bec81, BJ83a, CW82, All83a, Ant81, Aup83, Bea81, Beh81, BS83f, Bor82a, Bou83b, Bri83, Bur81, CS80a, Clé84, CRW^{+}80, CN84b, Det82, Det83, Det84, Dom83a, Don82h, Don82i, Ebe83, Est83a, Est83b, Fey80b, Gar82d, Gin80, Gin81, Gin83, GHZ83, Gra83a, GL82d, HH81a, Hei84, HM83b, HM83c, HJ82b, HT81, HTW81, Jam83, Kan83o, KP84a, Kir81, Kra82b, Kue81b, Kür83b, Kva80, Led81b, Led83, Led84a, Loy83, Mak80a, Man81a, MNP81, Mik82a, MSW^{+}84, Mus83a, Nic83, Pau80, Pau83, PR82, PPS83, Pis80, Pog80, Ros82, Ruc82, SC81e, SC81t, Sem82e, Ton83, Woy82, Woy83, Yas81, vDS83]. Banachic [AA83]. band [vL81]. bands [Sze81]. Bang [Val84a]. bank [Bea82a]. Barriers [Wer81b, Par82d]. bars [Ber84e, HHB84]. Base [Fli82f, She81b, Gie82g, Gie82k, Jen80, Lau83c, Sot82b]. based [Csö81, Cze84, EN83, GM84a, Gur80, MS83b, Rei82b, Ste81b, Tho84l]. Bases [BH80a, Sem82a, Sem82b, Sem82e, Daw82, HNP81, Kha82i, Nat84, Sem82c, Sem82d, TNV^{+}84]. Basic [Bou83e, Gie82d, ŁK83a, Pro81, Rog81a, Rot84a, She82d, Str84b, Bue81, BF83c, BF83d, KS83c, Kna83b, ŁK83e, Tan81j, Tan81k]. basis [Bro82a, Clo81a, Mar81d]. bass [Les83, Geo81a]. basse [Kas81]. Bauer [BR80b, BR80c, Loe80]. Bazilevic [DSS83]. be [Kis81b]. Beach [BBC^{+}83]. beam [FO82]. begin [Wil81c]. beginning [Wil81c]. Begriff [phAD80b]. behaved [Pre83k]. Behavior [Har81i, DK82a, Glu80, Har81h, Izé82a, Kör83, McG82, Ort80, Tei81a, Woy82, Woy83]. Behaviour [Eng84, Ant83, Atk80, Atk82, Dar82, Her83b, Kot81, Pre83g]. behavpour [Cow83]. being [Sot82b]. Beispiele [Sch82g]. Belgium [vO80, vOV82]. Beliebig [Dra80c]. below [CJ84, EHK81, Jos84c]. Bemerkungen [Dra80d, Küh83a, Lep80, Sch82c]. Bendixson [MT84]. Bergström [Mäk80b, Mäk80c]. Berlin [AJ81, FKSW81]. Berndtsson [Sko80a]. Bernoulli [Ebe83, Fei82, Fei84]. Bernstein [Bax83, Bjö81, BB80, Lac84a]. Besov [Mer84]. Bessel [Ars81, PY81, PY82, Yor84]. Best [Che80a, Cus81, Flo80b, God83]. Betreffend [Dra80f]. Between [Fef84, HO82, Jos84k, Ani84, Bet82, Car83a, CN82, Cun83, Dup80, FS81b, Gie82p, HTW84, Her84a, HW81c, Jon84, Jos84d, Jos84f, Jos84h, Ket82a, Kle84c, Lad81, Mar83a, Mue81, Nag82, Par82e, Par83d, Pis82, RU82, Sen83a, Ste83d, Wed83, Wil81d]. Beweis [Lün81]. Bewerteten [Dra80c, Dra80a, Dra80b, Kne80]. Bex [KO81a]. beyond [Ars81]. Bezeichnungen [Sch82i]. bi [Cor83a]. bi-excessive [Cor83a]. Bianchi [Ves83]. bias [Hem82a]. bibliography [GS83c]. bidisc [Gun80b, Jak83]. bidisque [Cha84d]. Bieberbach [Bae82, BK81c]. Bielefeld [Bak84a, BS83e, DK80b, Hin82, Men80a]. Biennial [Gri84, Wat80, Wat82]. Bierlein [Lem80]. biextension [Bre83d]. Biextensions [Bre83c]. bifurcating [IACT81, Van84]. Bifurcation [BB82a, Bus84b, CR80b, Ioo80, Ioo84, Izé80, AF81, Ant80, Bol80, CG83b, CMPY83, CR80c, Col81a, Fie83a, Guc81, Gui80c, Hal81, Hal84, Ize82b, Jos80, JB81, Kea81, Lou84, Mag80, Mit82b, RCC80, SH81a, Web81, BLS80, Sal84b]. Bifurcations [Dui84, Bro81a, Tak80, vS81a]. Big [Sho84, Tes83]. biharmoniques [Bou84c, Smy84]. biharmonisch [Gru82c]. Biholomorphic [Kró83, KŁLS80, Kra83]. bijections [Ket82a]. Bilaterally [Dom83b]. bilatère [Kas81]. bilinear [Ber80d]. Bilinearform [Gru82a, Gru82s]. Billingsley [Caj81a, Caj81b]. Bimeromorphic [Uen82]. binary [Chr83a, FR81, Hos82]. binding [WTL84]. biological [Hen81c]. biology [Gra83c]. Biomathematics [Bus84b]. bipartite [Ell83, EH84, Gan81, Jaw83, PM83]. bipartitioned [Gan81]. biplanes [Lan81a]. birational [dS82]. Birationality [Ver80b, vOV81b]. Birkhoff [Maj84g]. birthday [Sul81a]. bis [Mey82b]. Bishop [Goo81a]. Bismut [Mey81b]. Bisublinear [Don82a]. Bitopological [Iva84, Brü82b]. bivariant [And81a]. bivariate [AW83, Nie80, OR83, dB84a]. Błażejewko [Ław83, Wer80]. Blackwell [Sho84]. Blanchfield [Hil81e]. BLD [Fug82]. Blendingfunktionen [Gru82q]. block [Bai81, Bul84a, CC83, Dem83b, Kee83, Mor80b, PV81, Ple83h, Rem82]. Blocks [Ple83b, Ple83a, Mic81b, Sor83, Sri81, Tre84]. blocktheoretic [Ols81]. blow [Pav84]. blow-up [Pav84]. Blumenthal [Gir81b, Gir81a, Gir81c]. BMO [BHL81, DeV84, Blo83, Cha82a, CMT^{+}84, Gar80, Mil84a, ND83]. Bo [Sko80a]. board [Har83]. Bochner [BD83, Bou84e, Gre83, Hir84b, LT84b]. BOCS [Kle81, Roj80]. bodies [Dra80e, KSS80]. Body [Sig83h, Ens83, Sig83i, Tam82c, Tam82e, Tam82i]. Bogomolov [Ray81, Rou81b]. Bologna [Vin84b]. Boltzmann [Cer84a, Art84a, Elm84, Neu84b, Pal84, Spo84, Zwe84]. Bonn [JP81]. Book [HHN84]. Bookstores [Die83c]. Boolean [Bon84, Bur83c, DG80a, Fon84, Fra84, Kop81, Nam81, Pie83a]. bord [Kor80a, Sch80o]. Bordism [Kos81a, Kre84f, Kre84e, Kre84d, Kre84g]. Borel [Sch83l, Bia81, Cha83t, CP81, Gar82c, Grz84, Lou83a, Spa82a, Spa82e]. Borne [Mic81a]. bornée [Led81b, Sko80a]. bornés [CMS80, Vig83, Yan80]. bornological [CM81]. boule [Sko80a]. boundaries [Apa83, BN80, Cha83s, Hil84, Yor84]. Boundary [Beg83, Cha83g, CdS81g, Ivr84c, Krá80a, Nam80b, Ole83, Ran82, Sig83a, Sig83b, Ale81c, Bax81, Bax83, BH81b, BE83b, Böh81, Böh82, BN82, Bru82c, Col80a, Dol83a, Fad83, Fic80, FO82, FK83b, Gek84a, GM82f, Gus83, Hab80a, Hil82a, Hue80, IS81a, Izé82a, Kaw82, Kla81, Kre81c, Lod81, Loe80, Lor82, LS82b, Mae80f, MW84, Moe82, NSE82, O’M83, Shi81b, Skr84, Tak84, Val84c, War82b, War82e]. boundary-value [NSE82]. Bounded [Sch83b, Tam82f, AHS82, BB84, Ble83, Blo84b, BN82, Erk83, HG82a, Jan83a, Leu80, Loe80, MZ83, Pas80, Tri84, Web84b, Wer81q, vdK82]. Boundedness [AS83, Har81c, KS83e, Mor84a]. Bounding [Flo80c]. bounds [AS84a, Cus84, JN82, Jen83, MSP83, Neu82a, Osg84]. Bourbaki [Ano80, Ano81b, Ano81c]. Boussinesq [CD80, GLCS80]. Bowen [Kif80]. Boy [Lan84b]. braid [Joh84, Moi81]. branch [Bar83c, BS81, Moi81]. Branched [Whi83]. branches [Erd81b]. Branching [Sap84, Mon80a, Rog84a, vHSV81]. Brasil [Pal83]. Brauer [KO81a, Art82a, Blo81a, DF82c, Ele82, FS81c, Hür81, Hür82b, KO81b, Orz82, OV82, Rin80a, Tan81c, Van80, Ver82b, Ver82a, vOV82]. Brazil [Izé80, Mac81c, dFH82]. Brelot [Las80b, Saa84a, dlP82]. Bremen [AGP81, BH81a]. Brief [Sot82a, Ram82a, Ram82b]. Brill [EH83, Ghi83, Mal81b]. Brisbane [BOWS82]. British [Car82]. Brjuno [Mar81i]. Brocot [CLM84]. broken [Sem82a]. broken-line [Sem82a]. Brouwer [vdLT81]. Brownian [Bar81b, BP84b, Fuk84, Ken81, Ken84a, Kuo80, McG84, Pri84, Rog81b, Rog84a, Rus84a, SC81m, Yor84]. Brownien [JY81, GP81, Sch80o, Yor83, Gal82b]. Browniens [Fer81c]. Bruck [Jür83b, Jür83b]. Bruhat [Hof81a]. Brunswick [Win82a]. Brusselator [Rot84b]. Bucharest [BP84a, CBJS83a, CBJS83b, PPS83]. Buchsbaum [Sch82d, Sch82e, Sch82g]. Buchsbaum- [Sch82d]. Buchsbaum-Moduln [Sch82e]. Buchsbaum-Ringen [Sch82g]. Buffalo [Flo80j]. Bulgaria [DP82b]. Bulgarian [DP82b]. Bundle [Kos81e, And82, BH84, Eta84, Gie82h, Mar83g, Ung82]. bundled [vdB83]. Bundles [Gie82g, Gie82e, Gie82f, Nar82, BF83a, BT82a, BS84, Czy81, Eis81, Fed84, Gie82j, Gie82k, Gie82n, Gie82p, Gie82r, Gie82o, Gie82t, Gol83a, Jaw81, Kup80, KT84, Laz84, Mar82b, MR83, Meh83b, Mor82a, Sch80n]. Bures [CC82f]. Bures-sur [CC82f]. Burgers [Fit84]. Burkholder [Led81a]. Burnside [HN80b, Men80a]. Butler [BS83c, BS83d, Gio83a, Gio83b]. Butuzov [SM83b]. BVODEs [Rus82b].

C [Dem80a, Flo80b, Flo80f, Hus81a, Hej83g]. C.B.M.S [KKS82]. c.c [She82i]. C.I.M.E [Cer84a, Con82b, Ghe83, GS84, MM82b, Sal84b, dV84]. C.N.E.T [KMS81]. C.N.R.S [BMR81]. Cône [Ioc84g, Ioc84i, Ber80b, Hén80a]. Cônes [Cho83a, Ioc84d, Ioc84c, Ioc84a, Ioc84e, Fey84b, Ioc84b]. Cabal [KMM81, KMM83a]. cad [Le 84b, MS81b]. cad-lag [Le 84b, MS81b]. Calcul [BJ83b, Bis81c, Bis81b, Bis81d, Föl81a, Lel82a, Bis81g, Com82f, FLM81, Mél80]. calculation [Bas81, DO80b]. Calculations [Cha83d, Föl81a, Pow84, SJ82]. calculée [Zhe84]. Calculs [BBM82a, LP81]. Calculus [Lau83a, AHE81b, Ben84a, BF82, Bis81j, DD83, Kan83h, Kra80i, Mar83a, Mey81a, Str81d, Str83a, Ves81, Zaf82]. Calcutta [Rao81]. Calderón [CM80b, CS84b, Jou83f, Jou83c, Jou83e, Jou83d, Jou83j, Mey80c, Ric82]. calibration [WMW83]. California [BBC^{+}83]. Caltech [KMM81, KMM83a]. Campanato [JTW83]. Canada [BDM83, FF81]. cancellation [Arn82o]. Canonical [Abh82u, Fei82, Bou82d, Bou83k, DPV82, Kåg83, KK83, Mor82a, Roc84, Vog84, Wil82b]. canoniques [Bre83f, DI81a, Fra83a, Lel80]. Cantor [Mae81, MT84]. capacitaire [Kah84]. Capacitary [Fuk83]. capacité [Anc84b]. Capacités [Gal82b, DFM82, RP84]. Capacities [Del81g, HKD^{+}84, LT84c, Ceg80].

Capacity [Car80a, MS82e, Wer81c, Car83a]. capillary [Bem80, KWW83]. Cappell [Lan80d]. Caracas [PR83]. caractère [dlP82]. Caractères [Duf81, Mél80]. Caractérisation [Str84h, Zay82, Cho80a]. Caractérisations [Dem81]. Caractéristique [DI81b, Hür81, Vid80]. Caractéristiques [Las81d, Gui84a, Las81c, Las81b, Las81e, Led84b, Mis82]. Carbondale [Ama81]. cardinal [Hau81, Kak81]. cardinal-determined [Hau81]. cardinals [BM81a, Kec81c]. Carleman [Boi83, Gau83]. Carleson [Joh83b, JM82a]. Carleton [Ban82, DG80d, DG80c]. Carlos [Izé80]. Carlson [Shi83a]. carré [Bou84d, Kas81]. Cartan [Her83a]. Carter [Dix81, Gre80b]. cartesian [Ble80, BH82b, HW81c, McA80]. Cas [BD80a, Com82a, Com82b, Ang81h, BJ83b, Bou82a, Erm80a, HM82c]. case [Aub82, Ber83c, Bro81a, BK81c, CC82d, CC82c, DP83b, FK83b, Har81g, Kre84f, Kre84l, Kre84m, Mäk80c, Nam80c, Saz81a, Saz81b, SC81g, Sig83a, Sig83b, Spa83a]. cases [CdS81k, Har81l, Har81m, KS83c, Kna83b, Per84]. Casson [Sie80b]. Catalan [Fin82, Ket82a]. catalysts [Fie83a]. Catania [GS84, Sal84a]. Categorical [BH81a, Gir82, KP82, Ban82]. catégorie [Lau83b]. Categories [BV83c, DM81b, DM81c, Huš82b, Nel82, Năs80, AS82a, Arn82a, Gra82c, Her82d, Lad81, Lam82, MW82, Rin84d]. Category [Vin84a, BH82a, BW82, CP83b, HJ82a, Huš82a, Hyl81, Jos83a, LM84b, Wei82b, KPT82]. Cauchy [Beg83, BW82, Bré81a, DM82, Dwi83, Gra82b, Has83a, Iwa83, Jou83g, Jou83f, Las81c, Pal84, Ran82, SS84a, Ver84a]. Cauchy-completion [BW82]. Cayley [CQ83, CH84, DM84, Gre82b]. CBMS [CW82, KM80]. celebration [Sul81a]. cell [Wal81b]. cell-like [Wal81b]. cells [GM82a, Haz81b, Lus83]. Center [CCCN84, And81a]. centers [Doz81a, Hur82a, vDS83]. Central [BL80b, Gin81, Hab80b, Ros82, Hey81, Ila82, Led84a, Maj81d, NV82, Pau80, Roh81, Ryc84, Sla84, Wil82c]. centrales [Tah82]. centraliser [Joh81a]. Centre [DPT80, Eel82, MS81a]. Centro [Cer84a, Con82b, Ghe83, GS84, MM82b, Sal84b, dV84]. certain [AS84a, Ben82a, Bou81e, Bro81c, BS83h, Buc84, BF83l, Cat83b, CL84b, Ell82, EF81d, Fek83, Fom84a, Fom84b, Her80b, Hoh83, HRW84, JM81b, Kah82, Lad81, Laj81, Lor83, Los82, Mañ81b, MS83d, Onn82, OS80, Pap84b, Ruy80, Saw82, Sel81b, Shi83c, SV82, Ste82a, Tei81a, TW82b]. Certaines [Cau80a, Tal80, Ale81b, Cho83d, Cho84m, Fer81c, GV84, JY81, Kow81a, LLP80, Str84g, dB83]. Certains [Led83, BBP82, Mau82, PP81, Seb80a]. Cesari [Gra84a]. CF [Tre84]. CH [EH81b]. Chaînes [Die83a]. Chadam [Bar81a]. Chain [Mon80b, FN82, Kum81, Mar81d, Ped84a, vL80]. chain-complete [Mar81d]. chaines [Der84]. chains [BM83, DPV82, Die83a, Dol83a, GM84b, GGHJG81, MN82]. Chaired [ARMS82]. chaleur [Paq82]. challenging [FO82]. champ [AR80a, AR80b, Bou84d, GL82a, GR83b]. Champs [Bre83d, Kla83a, Mar81i]. Chandra [Del81h]. Change [Fli82f, Pet81, Eng84, She81b, Sot82b]. Changements [Com82c, Jac80b]. changes [AT82g, McG84]. Changing [She82c]. channel [Glo83, Sig83a]. channels [Ami80]. chaos [Cas84a, Pet81, Sie81b]. Chaotic [Sie81b]. chapitre [Car81c]. Chapter [BW81a, JM82b, JM82c, JM82d, JM82e, Jos84k, Min83b, Min83c, Min83d, Min83e, Min83f, Sot82c, Sto80b, Sto80c]. char [Suw83]. Character [Kin81, Ull82, SV82]. characterisation [Rog81b]. characterisations [Her80c]. characterised [Pra82a]. Characterising [dMPvS81]. Characteristic [Cam80b, Cam80c, Hop83g, Orb84, SS84a, Alm80, Alm83, And84b, BO82, Ber80c, BC81, Blo82a, BJ83c, Car80e, Che84b, Del80b, Dou82, FMW82, Gat84, Gil81b, Hab80b, Kap82, KS83e, MR83, Ost83, Roc84, Wil81e]. characteristics [BLP83d]. Characterization [Ach81, BH80b, Cho81, Das81, Gan81, HW82, Has83b, Hop83l, Iit83, KKM84a, KKM84c, KKM84b, Kna83c, Lan81a, Mad80a, Sch83c, BV83f, Cho80a, Czé83, Dav81b, DH80, Gun80b, HV81a, HPR80, Hoh83, KKM84f, KM83, Koś83c, Luk81, McA83h, Mdz81, Pan83a, Pra80, Sem83, Shi83b, Šig83l, Sor83, Wan81, Zah84]. Characterizations [BT81, yCV81, KKM84d, MS82b, OR83, Sot82b, DM84, JTW83, Kol81a, Laj81, Pet83, Sot82g]. Characterizing [Edw80]. Characters [Ros84d, Gre80h, Kle84h]. Charakterisierung [Ben80c, phAD80b]. Charakterisierungssätze [Gru82h]. Charakterisierungssätzen [Gru82o]. charakteristischen [phAD80b]. Charge [AR80a]. chargée [GL82a]. Chatteryji [SC81s]. Chebyshev [Del82b, Mas84c, vDS83]. check [Ver82a]. Chemical [Rot84c, DESS82, Ort80]. chemin [Nua81]. chemistry [Brä82b, WMW83]. Chen [Tan83c, Tan83f]. Cherlin [Ber81c]. Chern [Bea80]. Cherry [Gut83b]. Chevalley [Slo82]. Chicago [LW81]. Chinese [Day81]. chirurgie [Sie80b]. choice [Hul82, vdHM84]. Choix [Del81a]. Choquet [BF80i, Fey82a, Itô84d]. chow [Ang81c, Ang81i, Ang81b, Bea83a]. Christine [Gab81c]. Christoffels [Lei81]. chromatic [Wal84]. chronometric [Jak82]. CIM [JK83]. CIM-groupoids [JK83]. Circle [LW81, AF82, HL83a, Kom80, Kör83, Kos80b, ME80, TH84]. circles [IACT81]. Circulant [Ead80]. circular [DO80b, ESV83b, Jac82b]. circulations [GQ82]. circumference [ESV83g]. City [CCCN84, Kom84b]. Ck [Kan83o]. Ck-classification [Kan83o]. Clarke [DM81d]. Class [Kat83, Ark84, Arn82j, BCF82, Ber83a, Bjö81, BKS82, Böh81, Bou81d, Bou81c, Bré81a, Bus84a, CS83a, Cas84b, CL84c, CL84d, Doi84, Eas82a, Fek83, Gel81b, GL81, GH82, GT83, Haf81, HP80, HJ81, Hul80, IS81b, Ket82a, Kre84g, Kum81, Mae81, MM82a, Mäk80d, Mat82, Neu81, Orz82, Ove82a, Pot82, Rit81, Sam82b, Sch82b, Sco81, Sle84, Ull82]. classe [GH84, Las81b, Roq82]. Classes [Ang81a, Bou81g, Has83g, Mis82, Nar84g, Spa82b, Spa82a, VW83a, Zie81a, Ale81a, And84b, Bar80d, Bar80e, Bou81f, CL84b, Con81b, Djo81, Fie83b, Has83f, Her82a, HN81b, Jia81, KT82b, Kha82f, Kha82g, Laj81, Lin81, Los82, Mad80a, MW81, Mei80, NQ– D84, Oes80, Pic80b, Ros81a, Sch83o, Spi84, Thi81, VW83b, Zie81d, Zie81c, Zie81g]. classgroup [Frö81, Wil81g]. Classical [Ber82a, CCN82, Gal80, Hey80b, Hor80, Ver84l, Zel81c, AS81b, Cuy81, Djo81, Don82h, Dre83, Ell80, FS81b, Fef84, Hur82a, JvdS84j, LG81e, Man84, Mos81a, Net80, Sel81a, Sel81b, Wel83, Wer81u]. Classifiant [Tan83g]. Classification [Die83b, Esn81, Has83c, Kla83a, Lev80h, Lev80i, Lev80f, Lev80g, Lev80j, Mur82a, BH84, BY82, Ber83c, Blo82a, BL82, Car83c, Cha83b, DR80b, Est83a, Gro80c, HM83b, Jos83b, Kan83o, KS82, KS83c, Kos81a, Kos81b, Kow80a, Kow80b, Lor83, NiW84, Pra84, Ric81a, Wil81e]. Classifications [Adi80a]. Classifying [Gar82b, Ket83a, Kah82]. classique [BBM82b, Bis81o]. classiques [Let81, Net82a]. classless [Sch81a]. Clausthal [AD83b, DAP82]. clefs [Éli84b]. Cleveland [CW82]. Clifford [Sel83]. Clique [Pul83, CP83a]. clos [Dur80, Vid80, Zay82]. Close [KS83f]. Close-to-convex [KS83f]. Closed [Hop83b, Hop83k, Hop83p, Hop83r, Hop83s, AF83, Bou80, Che84a, Cur80a, Dic80a, Est83b, Gre82d, Ham82, JS82b, KS84c, Kre84a, Mok83, Poh81a, Pre81, PR84g, PR84c, PR84f, Pre84k, Sze81]. closedness [BH82b, Gre82e]. Closure [Ste81a, CV84a, Don82g, Gie82h]. Closures [Djo81, GM84a, Pre84e]. clue [Bai82a]. Clusters [SM83c]. CM [Sch84d, Sch84e]. CM-Typen [Sch84e]. CM-Varietäten [Sch84d]. Coarse [Don81]. coarseness [Mic83]. coatings [EV82]. coaxial [Par82e]. Cobordism [Cha83i, Szű80]. Cobordisms [Hus81b, Hau80b]. cocompact [FH84]. Cocoyoc [Hen82a]. cocycles [Fur81]. codes [BJ81, Her82a, Ple82, Rus82b, Tho82]. codimension [Bro81b, Ecc80, Guc81, Hen80c]. coding [Bet82]. coefficient [Atk82, Bre83k, Tam82c, Tam82e, Tam82i]. Coefficients [Bru81f, Bré81a, Bru81g, Cor84c, Fri83, Goo83e, Gru84, Kos83a, Seb80a, Des84]. Coercive [FW82b]. Coeur [Cau80b]. Cofinal [SS80b]. Cofinalities [She82c]. Cohen [Hus81a, Mer80a, Ore83, Zou83]. Coherent [Ban81b, LM84b, vOV81c, Ben82a, Bou82d]. Cohomological [Dic80b, Hür82b, Wal81b]. Cohomologie [Bes81, Fon83, KO81b, Ang81a, Jen80, Kum84b]. cohomologies [Lic84a]. cohomologiques [GJ81]. Cohomology [EF81g, Ogu81c, Ogu81d, Sch83o, Sch83r, Aky82, And84b, EF81a, EF81d, EF81c, EF82, Fal83a, FP81, Ill83, Kat82, Lyu84, Maj84b, Man82, Spe81a, Stu82, Suw83, Sch83q]. coincidence [CN84b, Gór81]. coinductive [Mos83]. cokernel [Van83a]. cold [Col84, Jam84]. cold-drawing [Jam84]. collapsantes [Tho81a]. collar [Gal83b]. collection [Daw84]. College [HLR83, HKLR83, HJLR84]. collision [Art84a, Die82b]. collision-induced [Die82b]. Collocation [Nør80j]. Colloque [CV81, KMS81, LS82a, MP84b, Ano81a, Cho84l]. Colloquium [AJ81, BFF80, BOR^{+}84b, FKSW81, KY84, MR84b]. color [Com83, Van82]. colour [ST84]. colour-representation [ST84]. coloured [Ell83]. colouring [Cam82b]. Columbia [Car82, CC82f]. columns [De 83]. combination [Pas80, Tam82h]. combinational [Erd80]. combinatoire [Car81c]. Combinatorial [BOWS82, Erd81b, Fie83b, Hsu80, KN84, McA81, RSW80, AD81, Cor83d, Jov84, JW83, Lep82, LS81b, Mil80a, Win81, ZVC80a, BOWS82, Cas83a, JV82, McA81, RSW80]. combinatoric [HKP83]. combinatorics [Bet82, Nam81, Obe82, Rao81]. combined [Raf83]. combustion [vH82]. comme [Yor83]. Comments [Krá80b, Cor80c, Maj81a]. common [Chu82c]. communtative [Est83b]. Commutant [Cau80a]. commutatifs [Cha83a, GG84, Gal89, Kla83a]. Commutative [All83a, CV81, CV83, Cat84, Dra84, Fle82c, Neu84a, Ton83, vOV81h, AW82, Est83a, Fos81, Jaj84b, Mal80a, Mal82b, Mon82, VVV81, Ver80b, Ver81b, Ver83d, Vog82, vOV81i]. Commutatives [Tan83e, Roq82]. commutativity [Kos82]. commutators [JP84b]. commuting [Chu82c]. comodules [Kle81]. Compact [Der81, Flo80j, Kit84, Lin82a, NR84, Rup84a, Sun82, AA83, BPV83, Băn83b, BD83, Boi80, Bra84b, BH82d, CS82a, Cor83a, CR80e, Din82, Din81, DP83b, Edg82, Flo80d, Gol83b, Gui80a, Haz82, Her83a, Hof82c, Hum80g, JvdS84d, Kan84a, Kau81b, Kin82, Kin84c, LR81, LM83, Lor83, Los82, Mañ81b, McC82, Nai84, Net82b, Ohn84, Ohs84, Ren80d, Rup83, Vek84, Wei80]. compacta [Fer81d, Geo81a, HI81a, O’F83]. compactes [Cer81]. Compactification [Nam80j, Has83h, Nam80a, Nam80f, Nam80i]. compactifications [Dyd81, GK81, Hof81c, Rup84e]. compactness [BD83, BR81a, CMZ80, Dac82b, Din82, EMS82, Flo80e, Flo80g, Fra81, HM81, SS81a, Sbo83]. Compacts [Bel83a, Cho83a, BC82b, Cam80a]. Comparaison [Bar80f, BBM82b, Lel80]. Comparison [Gil81d, GS82b, LT84c, AB83, Cuy81, Enr83, Ket82b, Rus82b, Tra82a]. Compensated [Dac82b]. Compensation [Éme80a]. competing [FN82]. competitions [Cla83]. Complement [She81c, DL81, Lag81, Del81f]. complementary [Das81, Gan81, Rei84b]. complements [BCJ^{+}81]. complet [Hir80]. Complete [BBŚ82, CC82f, DM84, DP83a, For84b, Hwa81, BP81a, Ber82a, Bor84a, DG84b, Fas80, Gat84, Hof81c, Kal83, Kat82, Kee83, Kop81, Mar81d, Miy81a, Miy81b, Miy81c, Miy81d, NiW84, Pic80b, Rob81c, Sch84j, SH81b, Val84b, Abd83, GS84]. Completely [Arn82c, AV83a, Lue81, SC81a, SC81b, Tho84c, Tho84d, VW83a, Arn81, Arn82h, Arn82m, AV83b, Chu82b, Chu82c, Clé84, Gér80, Gie82k, Hof81c, STS82, VW83b]. complètement [Kla83b]. Completeness [Ens83, SH84]. Completion [Jos83a, Lev80k, AS82a, Bal81b, BW82, But82, Ern81a, Hof82b, Lev80a, Mae80c, SC81d]. completion- [Ern81a]. Completions [But82, Fuc81b, GK81, Her82d, Mil81b]. complets [WK80b]. Complex [BHS81, Gro80b, HHN84, Hit82, LW82, Nac82, Nar82, Bed82, Ber84c, Bru82c, CS82a, Ceg80, Ceg83, Ceg84, CRW^{+}80, Con84, CB84b, CN84b, Dyn81, Ele82, Eva81, Gar83a, GM82d, Gre82b, Gro80i, HT82b, Ill83, Ivn83, Kai80b, Kau81b, Kis81a, MS82e, MS83d, Ohs84, Ran82, Riv84, Ros84a, Ros81a, SS80a, Sch80v, SS84a, Vla82, ZVC80d, CBJS83a, CBJS83b, Eel82, HN81a, Suc83]. complex-analytic [SS84a]. Complexe [AGT84, Cha84a, DG80b, GŁ82b, GR83b, Kis84a, Lel83, Sch80s, dB83]. Complexes [Nac82, Nor80a, OS81, Sch80v, Băn80, BH82c, Est84, Fos80, Fox80, HP84f, Lan80c, Ped84a, Sch80t, Sch80w, Suc83, Tho81b, ZVC80a]. complexité [Mic81a]. Complexity [Car81a, Bet82, DL84, Gur84, Mac81d, Man80, Wil80a, Yap84]. component [Chu82a, Chu82b, Gri80, Mor81c]. components [Bjö84, Fra83d, Nam80b]. Comportement [Bou84a, Ros84e]. composantes [HM82c]. composes [Jac81]. composite [SP83, Ste83d]. composite-media [Ste83d]. composition [Sot82b]. Compositions [vL81, Kis83, OR80]. Compressed [FL84b]. compressibilité [LB84]. compressible [LB84, Maj84a, PW84, Val84c]. Comptes [BMR81]. Computability [How81a, Nor80c]. computable [Nor80d]. Computation [Goo83b, Goo83c, GY82, SW81, Ame82, HN81b, IACT81, KN82, Kre84g, MB82a, Ore82, BOR^{+}84b]. Computational [CC83, Man80, Yap84, Bre83a, DL84, Gur84]. Computations [Alt81, BF83n, Gra81b]. computer [Bar84, Dav81a, Mäk80e, SB80]. computers [HN80b]. Computing [CW80, Gil81b, Gol80, Kåg83, Kla81, LO84, Mäk80g, Ste83f]. concave [Erm80a, Lie80]. Concentration [Hes81]. Concentric [KS84c]. Concept [Hop83f, EP84, Ern81a, EN83]. concepts [ŁK83a]. conceptual [Gue84]. concernant [Anc84b, Itô84d, Tal82a]. concerning [Gil81c, Kis84b, Pri81, Tau81, Wil81d]. Conclusion [GM80c]. concordances [HI81b]. concrete [Her82d]. condensing [AF81, Zvy84]. Condition [Apa83, Gir83a, Boi80, CCV82, Dem83b, Der84, Hau81, Pis83, Ric81d, RR80, Tri84, Vig80f, Wer81d]. conditional [McG84, NS81b]. conditioning [Cec84, Wuy81]. Conditions [GP84, GB84, MT80a, Air83, All81b, Aup83, Bau81, Bro82d, Cor80c, GM82f, Hey80a, IS81a, Kno80, KSS80, Las84, Mon80b, Tei82, Van82, vL80]. conduction [Sah83, SSTW82]. conductor [Ore82, Ple83c, vdB83]. Cone [BBCH81e, Don82b, Pop83a, Wer81d]. Cones [BBCH81a, BBCH81c, BBCH81b, BBCH81f, BBCH81g, BBC80, BBCH81i, DG84b, Gen83, Hof82a, Ioc84e, Itô82, Jan83b, Mas84b]. Conference [All82b, Ama81, ABGM82, AV80, AL81, AL82, BBC^{+}83, BP84a, BH81a, Ban82, Bec81, BDM83, BOWS82, BKS83, BKK82, CV83, Cas83a, CZ83a, CW82, CGO83, CP84b, Den82a, Den82b, DG80d, DAP82, Dol83b, DLR81, EdJ82, Eve80b, ES81, ES82, FM84, FF81, FG83, FS81e, Fuk82a, GW81, GLM83, GMSV84, Gri84, HT82a, Hey82, Hey84, HJW83, Izé80, JV82, JPW81, KR83a, KM80, KPT82, KS84a, Ken84b, KN84, KKS82, Kno81, Köl80, KMS82, KMS84, Ław80, Ław83, LW81, MS81a, Mar80b, MdJ82, MRW83, McA81, NR80, PWW80, PPV80, RS83a, RW82, Ric81c, RSW80, Rog81d, Sal84a, SW84, TW84, Vin84b, Wat80, Wat82, WB84, Win82a, dBvR81, Ano81a, Ara84, Bak84a, KS83d, MO84, JW83]. Conferences [GPRS80, MTT81]. Configuration [Blo80]. configurational [dF84]. configurations [Cae80, GS83e, KS80a]. Conformal [Jad80, Jos84e, Jos84k, Apa83, Car80a, Woh83]. conforme [Fer83b, Sch80w]. Conformes [Sch80v, Cră83, Fer83b, Sch80r, Sch80s, Sch80t]. Conforming [McA83i, Sch80v]. confronté [Sou83]. Congruence [Ber84d, Hop83c, Cas83b, Czé83, Hum80e, Reh82, SSW83, Spe81a, Van82]. congruences [SSW83, Str81a]. Conic [BF83a]. coniques [Mer80a].

Conjecture [Mor81b, Anc84b, Bae82, Car81b, EP81, FH84, Geo81a, Gil81c, HN80a, McA80, McG82, Par82a, Par83j, Pri81, Ray83, Ric81b, SC81p, Sna84, Tay81, Wil83a]. conjectures [Dem81, Hir83, Kos80b, Mac81b, Rin80a]. conjugacy [Dix81, Djo81]. conjugate [Axe80, Bri83, Eis83, Ket82b, Mus83a, Oku84d, Pow84]. Conjugates [MS81c, MS81d]. conjugation [Gam80]. Conley [Con80b]. Connected [Bil81, Has83g, Ach81, Ber84b, GL82c, Kun82, Mor81c, Röd84]. connectedness [lPBF80, Jaw83]. Connectednesses [Pre82]. Connecticut [BS83f, LSS81]. Connecting [Lum80b]. Connection [SS80a, CM82a, Kle84c, Koh83, YV81]. Connections [Ani84, Hes80, Maj84c, And82, Her82b, Jad80, JvdS84a, Kob81, Kol81b, Ste83d, vS81b]. Connectivity [FL81, Dyd81, Ell83, Har84]. connexes [HM82c, Mar81b]. connexion [GR83a]. Connexions [GH84, Bis81p, Kla83b]. conormal [Kle84a]. consecutive [Mue81]. Consequences [Tam82c, Bro81a, Maj81g]. conservation [BE81, Mel82, Par81]. conservative [GM80h, GM80i, GM80j]. conservativity [Háj84]. considerations [Jos84i, OW83]. Consistency [Mad80b, Šče81a, She82c, Woo83b]. constance [Str82a]. Constant [Hop83b, Hop83k, Hop83p, Hop83q, sC83, Gru84, MS82e, Nav84a, Riv84]. constants [BF83h, Gro81b, Seb80a, Tes83]. constituted [Itô82]. constrained [Gay84, GL82c, Lac84a, Spi84, Van83c]. constraint [GN80, Hön82]. Constraints [Śni83a, GP83d, Mar82a, Ver84l]. Constructibilité [Dem83a]. Constructing [Ell82, vdHM84, HD84]. Construction [AD81, AP81a, Bel83a, Bil82b, Cha83f, Cha83e, Fer81b, HR81, HP84i, Jak80b, Luk81, Mal82a, MS81e, MS81f, Rin84a, Sak82b, BY80, Bou82a, CP84a, CN84a, Daw81, Ers82, Gie83a, GV83, Gre82d, Jou80, Kim81, Köh81, Lev80o, Lor81b, Lou83b, Maj81i, Nie80, Sch82o, Zie81c]. Constructions [Kle84e, AT82d, Bur83c, Dev81, Gar80, Joc81, Rup84b]. Constructive [DM81e, Bee81, Bri81a, Cha81a, FS81b, Fef84, Gel81b, Gre81c, KW84, Sel81a, Sie81a, Ric81c]. construites [Bou81a]. Contact [Gui80c, Ber84d, GV84, Lic84a, Mit81, SS84b]. contacts [FKM83]. containing [MT84, RU82]. contains [GT83]. contents [Web84b]. context [GM84b]. contigual [BH82a]. contiguity [GM83b]. Contiguous [AS84b]. continua [Cha84b]. Continuation [Col80d, Kea83, All81b, AMPY83, Bru81b, Bum84a, Hac82b, Kea81, Kle84b]. continuations [Goo83a]. continue [BY80]. Continued [Chi84b, Ger84, JTW82, Bre81c, Fra83c, Gil82a, Gil82b, Jac82a, JW82, JM84, JW84, JS82c, JT82, JR82, Kan83k, Liu84, Mag84, Ove82a, Rei82a, TW82b, TW82a]. continues [AGY80, CG83a, Die82a, HYZ83, Str84g]. Continuité [Fer81b, Bou84b]. ’Continuity [Sch81c, BBC^{+}83, Cor84b, Dac82e, Thi84, Aup83, Fuk82b, GŁ84b, Kot81, Kus82, Lia84, Rup84d]. Continuous [Aul84a, BH81a, CS81a, EF81a, EF81b, EF82, GK81, Gie82j, Hof81c, MP82, Nel81, Sch83e, Sem82e, Sie82, Ban81a, BB81e, CC82a, Che80b, Čob84, Die82b, Ern81a, Fli82a, Flo80e, GH82, Hen81d, HM81, Jak80b, JS81, JJT80, Leu80, Lut82, MV81a, Mar84b, Mar81c, Nic83, Sch82a, Sch83d, Sha80, Ver80a, Zvy84, BH81a]. continuum [Bro82a]. contour [Ros84d]. contoured [Mis84c]. Contrôle [El 81, Lio83, DG80b, El 84]. contracted [Ple82]. Contraction [Mor81a, AD83a, MM82a, SC81f]. contraction-disassembly [AD83a]. Contractivity [Krá80c, Ise84b]. Contre [Tal80]. Contre-Exemples [Tal80]. contribution [Mur84]. contributions [Gra84a]. Control [MM82b, Bar80a, Bar83a, Ben82b, Dah82, Gal82a, Haz83, Hul82, Kno80, Mit82c, Par82b, PS84c]. Controllability [Mar81e]. Controlled [Cha83g, Cha83i, Cha83h, Cha83j, Cha83k, Ver84e, Aub84, Cha83n, Chi83]. controls [Ark84, Bar83b]. convection [Hem82b]. convenient [DK82c]. convention [Ver84g]. Conventions [BBM82c, DMOyS81d, Ogu81e, War82c, ESV83k, Gro80g]. Convergence [BC82b, BC83, Don82c, GB84, Gil82a, GO82, Har80a, HZ84b, Ing80, Jur81, Mil81d, MN82, Sib80, Art80a, AHS82, AW83, Att84, Axe82a, Bay81b, Blo84a, Bor83b, BS83g, Bra82a, Bus82, Col80b, CA82, DZM83, DD83, Ern81b, FW82b, Gaf83, GM82f, GM83b, Guz81, Hac82a, HYZ83, Hig82b, Iva83, Jac82a, Jac82b, JM84, Jer80, JR82, Kör83, Kow81a, Kur84, Kva80, Las84, Lon81, Naj84, Nut82, Oku84a, Pau80, Pot82, Roh81, Rüs80, Ruz82, RS84, Smi83, Wec81, Wes80, dB84b, dG84a]. Convergences [Jac81]. convergent [BJ80, HV81b, Kas82a]. converging [Gel81a]. converse [Das81]. Convex [Bou83h, Ale81a, Bas84, Bie81, Col80c, CM81, Din81, DM82, Eke83, Flo80h, Gen83, Gie82u, Had82, HU83, Itô82, Jan82, JK83, Kha82c, Kha82f, KS83f, Mar83a, Sch83a, Sch83c, Sic84, Thi84, BF80c]. convexe [Tho81a]. Convexes [BF80c, BF80d, BF80e, BF80j, Cho83a, Cră83, Cho83a]. Convexité [Bar80b, Fey80b]. Convexity [BBCH81f, CV84a, Ole84, SC81s]. Convolution [Bru81d, Dal83, Kan83a, Kin82, Lau80a, BŻ84, CM82b, Gra83a, Grø83, Hul80, Itô82, Itô84e, Itô84c, Itô84b, Itô84a, Lau80b, Mal80c, MZ83, McC82, Ruz82, Sta84]. convolutional [Mei80]. convolutions [CM80b, HJ81]. coordinates [Jos84j, dGdlV82]. coordonnées [Bis81b, Bis81d, DI81a]. Copenhagen [BFF80]. copy [RU82]. core [DJK81]. coreflective [BH82a]. corollaries [Sot82j]. Corollary [Rot84d, HMP82]. Corps [Tig81, Vig80a, Vig80c, Vig80b, Dur80, Oes80, Rol82, Tah82, Vid80, Yam80, Zay82]. correction [ABM81, Hac82c, Hem82b, Lin82a, Ste84b]. Corrections [Gin80, AS82b, Fle82b]. correctors [Cla81]. correlated [HM83a, PS84b]. correlation [Abr82, Düm84, MP82]. correspondence [BF83c, Fli82b, GM82e]. correspondences [BB84, BF83j, Kem83]. corrigé [Har80a]. Cortona [MRW83, MM82b]. CoSaks [CS81b]. coset [Lyn84]. cosh [Cor83d]. cosine [Kur82, Lut82]. cotangent [Bis81p]. Cotorsion [Min81, Oko81]. cotype [Pis82, SC81h, SC81t]. Cotypes [SC81o]. Couette [DO80b]. Countable [Nor80g, FS83, HP80, Mar83d, MN82, Nor80e, Ton83]. countable-many [Ton83]. Countably [Flo80d, Mus83b]. Counter [Mil81b]. Counter-examples [Mil81b]. counterexample [Gin83, Los80]. Counterexamples [Kha82a, Tri84, BT82a, CP84a, Vol82b]. counterparts [Gel81b]. Counting [KT82b, BFG82]. couple [Let81]. coupled [Die82b, Gar82a, Kir84, Sec82, dM83]. couples [CN84b]. coupling [Tes83]. coupures [Rol82]. Courant [AR80b, Mas83b]. Courants [Erm80b, Mas83b, CB84a, Dem82a, Dem84]. courbe [Del81f]. courbes [Eli84a, GP83e, Tei80a, Ver81a]. courbure [BB81c, Bis81a, DG83, Gav82, Lan80c, Lem81b]. course [Dud84]. covariance [Ali82, Ebe83, Pap84a]. Covariant [Hol84, Bor83a]. covariantes [Bis81b, Bis81d]. cover [Gab81b, Whe80, Zah84]. Covering [She82q, Bru81q, CP83a, Leh83a, SM80, Sta81]. coverings [GvdP80k, KS83a, Pul83, Rem82, SB80, SW81, Sta82b]. Coxeter [Dav84, Mac81b]. cr [Hen80c]. cr-functions [Hen80c]. crack [PMC83]. Cramér [Nie80]. creating [vS81a]. Crete [PPV80]. cristallin [BBM82e]. Cristalline [BBM82g, Fon83]. Cristaux [BBM82d, DI81a]. critère [Die82a]. Criteria [Art80a, BD83, Din82, KS83e, Top80]. criterion [DP83b, Sot82g, Sot82k, Wer81s]. Critical [AK82, DMT84, Fou81a, Sah83, Wie83, Eas82a]. critique [Bou82a]. critiques [Com82a, Com82g, Lan80c]. crochet [Str84g]. Crochets [Tan83b]. Croissance [Las81a, Lel80, Tit81]. croissants [Kah80, MMS81]. cross [Gra82b]. Crossed [BH82c, CV82, Gur80, HT82b, Hil83]. crossed-modules [Hil83]. crossing [BKQ83]. Crowell [Hil81b]. Cruces [Ric81c]. Crude [ESV83c]. cryptomorphic [Nic80]. crystal [Wis84]. Crystalline [Ogu81c, Ogu81d, BBM82g]. crystallographic [Ple81, Zie81g]. crystals [BBM82d, Eri84a]. Cube [AA84, Bre83g, Bre83i, Bre83j, KT83]. cubes [Hav83]. Cubic [HB84, CP83a, Ell82, Mäk80k]. cubique [Bea82b]. cubiques [Del80a]. cupping [FS81d, Mil81c]. current [Red83b]. Currents [HL83a, Mok83]. Curvatura [Hop83s]. Curvature [Hop83b, Hop83k, Hop83p, Hop83q, Hop83s, Bou81i, Bou82b, BK81b, sC83, Der81, GS83e, Ken81, Ken84a, Kit84, Tak84]. Curve [Tei80a, CC82d, ESV83b, ESV83h, GvdP80c, Gra82e, Gro80h, KK83, Mar83g]. Curved [Hay82]. Curves [Cam80b, DG84b, GvdP80j, Gro80b, BE83a, Böh81, Buc80, Cam80c, Cam80d, Cam80e, Cam80f, Cam80g, CC82b, Cas82a, Cat83b, CV84b, CL84b, CM80b, Eis81, FL84a, Ful83, GvdP80b, GvdP80a, GvdP80f, GvdP80i, GvdP80k, Gie83a, Gre82c, Hul83, Jou83g, Kur83a, LR83b, MP84a, Miy83c, Moi81, MM83, Poh81a, vOV81i]. Cuts [Pil81a, Kla81, Par80]. CW [Tho81b, HP84e, HP84f]. CW-complexes [Tho81b]. CW-Komplexe [HP84e, HP84f]. Cycle [Ang81b, Bar80b, Kas83]. Cycles [CH84, Del81b, Del81c, DMOyS81c, Hén80a, Hol83b, Ogu81c, Ogu81d, Als81, Bar80d, BT82a, Blo81a, Cam80a, Gér81a, Isc82, Köh81, Lau83c, Llo83, Oda83b, Rei84b, Rey80, Shi83c, Tei80b]. Cyclic [Hsu80, Mäk80f, Rei83a, GT83, Köh81, Lue81, Mäk80c, Mäk80k, Mäk80l, Ple83b, Rei83b, Wil81g, EV82]. cycliques [EV82, Tig82]. cyclotomic [Coa81, Coh83, Mil81a]. cyclotomiques [Oes80]. Cyclotomy [LL81]. Cylindric [HMT81a, HMT^{+}81b, AN81]. cylindric-relativized [AN81]. cylindrical [Che81b, SC81i]. cylindriques [Fey82b].

D [Dem80a, Sch80a, SP80, Hem82a]. D. [Mar81i, Mey82a]. d’Acceleration [GB84]. d’additivité [Bre83j]. d’adjonction [MT80a]. d’aire [Gun84, Sko80a]. d’Algèbre [Mal80b, Mal81a, Mal82c, Mal83]. d’algèbres [Lev80a, Mal82d, Roq82, Tig82]. d’algébricité [Cau80b]. d’algébrisation [AT82e]. d’Analyse [CV81, LDS83]. Daniell [Tho84a, Vol82a]. d’anneaux [Les83, Sch81b]. d’annulation [Abd83, Sko80b]. Dans [Led83, Ale81b, Anc84a, Ang81i, Ava82, Bar81a, Bea82b, Bea83a, Bea81, BC82b, BJ83b, Bis81p, Bis81o, Bou82a, BO84, Cau80b, Cha84a, CB84a, Cha84d, Cha83t, CR80b, Cor84a, DZM83, Del81a, Dem80c, Dub82a, Duv83a, Ehr83, Erm80a, Fer81a, FM83, GL82a, Geb83, GR83b, dMG83, Gru80, Gru83, Gue80, HH82, HYZ83, HZ84b, Hir84a, HM82c, JG81, Led81b, Lel82a, Lev80a, LR83d, Mau82, PP81, Pra83, Rea83, Rol82, Ros84e, Rou81a, Rou81b, Sch80t, Sch80x, Sch80w, Sch80u, Sko80a, Sko80c, Smy84, Spr83, Yam80, Yam82, Zhe84]. d’aplatissement [Bar80c]. d’applications [Bar80f, Gal82c]. d’approximation [Pis80]. d’après [Bar81a, Bea80, Ber81c, Ber81d, Del80a, Del81f, Dem81, DM81d, Éme81, Esn81, Fos80, Gab81c, Hel80a, Hus81a, Lan80d, Lem81a, Mar81i, Mél80, Mer80a, MT80a, Mer80b, Mey81b, Mor81b, Pis80, Ray81, Rou81b, Sie80b, Ver81a, dM80a]. d’arcs [Kah80]. d’Arithmétiques [JG81]. d’arrêt [FSY82]. Darstellungen [Ros84c]. Data [CP82, BC82a, CD80, Far82, GM84a, Tho84n]. d’auteurs [Ano80, Ano81b, Ano81c]. d’autres [Bar81a]. Davidenko [Geo81b]. Deakin [McA81]. Dec [BS83e]. decade [Bro82e]. decay [Kni80]. December [CZ83a, GLM83, GMSV84, JPW81, KKS82]. Decidable [Clo81b]. Decision [Che80b, Rob84b, MR81a, Man80, Wei84]. Decomposability [Ole84, Bea83c, For84a]. Decomposable [AV83a, Cor83b, Sch80m, VW83a, Arn81, Arn82c, Arn82h, Arn82m, AV83b, Cor83c, Jaj80, Jan82, VW83b, vHSV81]. Decomposition [All82c, BŻ84, CC83, Jak80a, LR81, MT82, Thi81, Bou84f, Rod81, BC82b, Bru81n, Cor83a, DMPG83, Doz81b, Hej83b, Hof81a, Hus81a, Itô84e, Jak83, Kim81, Kun81a, Lit82, MS83c, Mos81a, Nie80, Ple83i, Röd84, Sch83h, Sig83j, Ste83f, Wer81a, Wer81p, Hus81a, Itô84c, PY82]. Decompositions [Bir83a, Bir83b, Dic80c, Fos82, Goo83d, Arn82d, BM81b, Fin82, Tig82]. decoupling [O’M82, O’M83]. decreasing [Pet80]. Dectdable [Che84a]. Dedekind [Lad81, Lad83a, Lad83b, Lün82, MR81c, vdK81]. Defect [AS82b, Hac82c, Hem82b, Ple83b, Ste84b]. Defektrelationen [phAD80b]. deferred [Lin82a]. deficiencies [Aka83]. deficiency [EG81]. Definability [Bal81a, Che84b, Mot81, Wil80a]. definable [Bae84, Pil81c, Wil81a]. defined [Adi80b, Bul84b, CKL83, Kom80, LO83, Sch83q, Sch83r]. definierter [Sch83k, Sch83q, Sch83r]. definies [Ber84a]. définis [CMS83]. définissant [Ble82]. definite [BE83b, CR83, EF82, FT80, IS81a, Min83a]. definiteness [BB83a, Bur83a]. Definition [Dar82, Cai80, Gie82n, Koo80, Yas81, dG84a, Eli84a, Kis84a]. Definitions [BFPS81, Cha83l, Van83b, DRR81b, FS81a, Gie82d, Man83, Mar81h, Rot84a, Sie81a, Sla84, Bre83i]. definitizable [Lan82c]. Deformation [Slo80a, CC82d, Gre82c, Jac80a, Kal82, NS81a]. Deformations [Lev82b, Lic80a, Nar82, Slo80b, Kla83b, Sal82a, Băd84, Buc80, Gir84b, MP84a, Pie83b, Sar83, Slo82, Lic80b]. Degenerate [Str84f, Kos83a]. Degeneration [Cle82, BE83a]. degenerations [Sam80]. dégénérés [Com82a]. Degree [Bil80, CS84a, War82k, Abh82n, Abh82m, Bil81, Bui84, Cha81c, Das81, EH80, EH81a, Hur82c, LR83b, Lev80f, vdLT81, vdL84]. degree-pair [Das81]. Degrees [AS84b, Art84a, Bil82a, CJ84, EHK81, FS81d, Fri81, Kec81b, Mil81c, Odi81, Sho81]. Dehn [Goo83m]. d’Eichler [Vig80f]. d’Eilenberg [Tan83h]. Del [Dem80b, Dem80c, Dem80d, Dem80e, Dem80f]. d’Elasticité [Ruí83, CR80a]. delay [AHS82, dBC80, Coo84, DI83]. delays [Car80e]. deleted [McA80]. Deligne [Gre82c]. delivered [JW83]. demi [Bat81, Duv83a, Mer81]. demi-directs [Bat81]. demi-groupe [Mer81]. demi-modules [Bat81]. demi-plan [Duv83a]. Demonstration [Mol82a, Car81b, Web84a]. Denmark [MO84]. denombrables [GG84, Gal89]. dense [KS84c]. d’ensembles [Bel83a, Mil83a]. densité [BJ83b, Lel82a]. density [Acz81a, Cho84b, Gol81b, Nat81, Pau83]. d’entiers [Que81]. d’entrelacement [FLM81]. denumerable [Bil82a]. Deny [MP84b, Anc84a, Cho84l, Dav81b, Hir84a]. Departement [HP84h]. dépendance [Gui84a]. Dependence [Pap84a, Bar80e, Hen81d]. dependent [CR83, Gre84, GLCS80, HR82, Hej84, Ros82]. depending [Men83, Mus83c]. d’équations [Duv83c, Kum84b, Las81b, Las81d, Lum80a, Mas82c]. Derham [Maj84b]. Derivations [Eta84, Vid80, Sie81a]. derivative [Bré81a, Hem84, Kea81]. derivative-free [Kea81]. derivatives [HU83, Iwa83, Krá80h, Pre83i]. derived [Lev80h, Lev80l, Lev80o, Lev80p]. derivées [Ser80, Gav83, Lau83b, Mey81g]. d’Erreur [Cha84a]. Desarguesian [Gly83]. Descent [Sot82c]. descente [Bre83j]. described [Not82d, Not82e, Not82f]. Description [Bre82b, Sak81, Duv80, Eas82c, Gie82c, Gie82u, Tay82]. Descriptions [Per84]. design [DKT84, Fum81, PV81]. Designs [Hsu80, SS81b, AP81a, BJ81, BP81b, Bil82b, CC83, Ead80, FHV80, GH83, GHH83, Hug81, Kee83, Köh81, Mor80b, Neu82b, Rem82, Seb80b, SM81, Wol81b]. désingularisation [Mer80b]. Desingularization [Abh82u, Cos84, Gir84a]. d’espaces [Băn80, HM82a, Sal82a]. dessus [Leg82c]. Desuspension [Gra84b]. detailed [War82d]. d’Été [FMSW83, Hen80d, Hen81a, Hen82c, Hen84a]. detection [Gea82]. determinacy [KSS81, KKMW81]. determinant [Jos84c]. determinantal [BB83a, Bre84]. Determinante [Dra80f, Sch80a]. determinantielles [GM82b, GM82c]. determinants [Tay82]. Determinatal [Hil81c]. Determination [Mäk80f, Pan83b, Tam82e, Tam82d, Wer82]. determine [May80]. determined [Hau81, KS81a]. Determining [FR80]. deterministic [Gis84]. Deux [BF80j, Dra83b, Hei80, KMS81, dlP80, Bro81d, Gun80a, Lel80, MS81b, Mey81e, Zhe84]. developability [Maj84d]. Developing [Enr82]. Development [Hej83a, Bra82c, Bre83a, JvdS84b, Wis84]. Developments [HJW83, Hur83]. développée [LC80]. Développement [Com82d, Aze82].

Développements [Bar82a, Com82e]. deviating [FK83a]. deviations [Aze80, Bro84, Che83, And80, Aze80, BBP82]. device [CA82]. devil [Aub82]. d’évolution [Bar81a, Lum80a]. d’existence [AT82f, CR80c]. d’extensions [BBM82a]. DFN [Kra80i]. DGC [Kle81]. d’Hilbert [Gir81b, Gir81c]. d’homotopie [Hus81a]. d’hyperplans [Car81c]. diagonal [Per84]. diagonalize [Dem83b]. diagram [Bue81, Maj81g]. diagrammatic [Rog80]. diagrams [HPR80, Höp82, Kim81]. Diamonds [She82k]. diatomic [Die82b]. dichotomie [Éli82]. dictionary [GM82a]. d’idéaux [Oes80]. Dieudonné [Dra80f, BBM82b, BBM82d, BBM82g, Sch80a]. Dieudonné-Determinante [Dra80f, Sch80a]. diffeomorphism [DR80b, FY81, HJ82c, Kat80b, Kre84i, Kre84j]. difféomorphismes [Poé80]. Diffeomorphisms [DRR81a, Kre84e, Bas81, BM81c, Jos84g, Jos84f, Jos84c, Kre84f, Kre84d, Kun84, Sea82, dMPvS81, vS81b]. Difference [Aul84b, Imm84a, AK82, Bra80, Car80e, dBC80, Gru84, Imm84b, Imm84c, Imm84d, KY80, KB80, Lor82, Tan81a, Köh81]. difference-differential [dBC80]. differences [CP82, Duv83a, Duv83c]. different [Han81, Leś84, MS82b, Pre83g]. Différentbielle [Leg82b]. différentes [Dem82a, Dub84]. différentiabilité [RP84]. Differentiability [Rob81b, Rei84a]. Differentiable [Gra83b, Kow80c, Fle82b, Kre84a, Las84, Mil83a]. différentiables [Gui80b]. Differential [iA83, Amb83, AD83b, AV80, Aul84c, BR80a, BR80b, CdS81g, Est80, Eve80b, ES81, ES82, EL83, FKSW81, GP83c, Hin82, Hop83e, Hop83r, Hop83o, Izé80, Jou83f, Mar80b, MdJ82, Mat80, Min83h, Nac82, PD82a, Str84f, dFH82, Ahl81a, AHL81b, All82d, And82, AG83b, Atk82, AHS82, Bar80e, Bec80, BS83b, Bia81, Bra80, Brä82b, Bro81c, Bro83, Bum84d, BV83j, CMA83, dBC80, CC84b, Col83, CPD81, Col80d, Coo84, Cor84e, Da 84, Del82b, Die82b, Dră83a, Eas82a, Eas83, ES83, ESV83d, Eke83, Elb82, Elw81, Eng82, Eng83b, Erb83, Eve80a, EN83, EW83, FK83a, Gea82, GG82, GP83d, Gek84b, Gek84d, Gek84f, Gek84g, GBC82, Hab80b, Haz81a, HR82, Hel84a, IK80, IV80, JM81a, Jia81, Joh83b, Jou83e, Kar81, Kas83, Kat80a, Kau83, Kos83b, Kré81a, Kre81b]. differential [Kun81a, Kun82, Kun84, Kwo83, LM80, Las81d, Le 84a, LO83, Llo83, MW84, Mei80, Mey81a, MB82b, Not82d, Not82e, Not82f, Ole83, OS80, Per82b, RW84b, Rot84l, RCC80, Rus82a, Sag80, SS80a, Sch82l, Sib80, Sle80, Sot82l, SS84b, Str83a, Tal82b, Tam82j, Var84, Vin84a, Wil84, Wol82b, dCR83, vdH82, DAP82, GPRS80, Hop83d, Ken84b, Nav84b, dFH82]. Differential/algebraic [GP83c]. Differentialform [Sch83j]. Differentialgleichungen [phAD80a]. Differentialgleichungssysteme [Wei82a]. differentials [Bin83, Sot82f]. Differentiation [Akc83, PT82]. différentiel [Air83, Bis81c, Bis81b, Bis81g]. Différentielle [AY82b, Cal81, Éme82, Mey81b, Mey82b, Sch80u, Sch82n, Wei83]. Différentielles [Las81d, Ber80e, Éme80b, GR83b, Kum84b, Las81b, Le 83a, Sch82n, Tan83e, Yan83b]. différentiels [Bet83, BBP82, Dem83a, Las81e, Lev81, Seb80a, Sou80]. Difficulties [Lon81, Rus82a]. Diffusion [Ros84e, Rot84g, Sch80o, BB82a, Bey83, Bis81a, Bus84a, Car80d, Hem82b, Itô80, KW82, Lum84, Mas82b, Mas82c, TW81, Tom82]. diffusionen [Ros83c]. Diffusions [AHK82a, Bis81e, Pap83, ABHK84, BJ83b, Bis81f, Bis81m, Bis81p, BC83, Det84, Fuk81, Fuk82b, HP84i, KLM81, Mue82]. diffusive [KM82, MSP83, PW84]. Digital [JS82c, CS84a, JS84]. digraphs [Als81, Bri81b, Das81]. dihedral [Kea82, Wil81g]. dilatation [Erk83, Dit80]. Dilation [AC82, AC83, Sig83f]. Dilations [MS80b, FG84, Küm84a, Nie84]. Dilogarithm [Gra81b, Blo81b]. Dimension [Caj81a, Car81a, Kah84, MT80b, Wal81b, Bak82c, BF82, Ble83, Bor82c, Bro81b, Caj81b, Caj81d, Cer81, Cos84, DP82a, Dic80b, Esn81, Gir84a, Hoc81, Kas81, Koj81, Lan80d, Lau83c, LS82a, Lev82a, Mañ81b, Miy81b, Miy81c, Miy81d, Mum83, NS81a, Sch81b, Sie80b, Sul81a, Sul81b, WK80b]. dimension- [Miy81b]. Dimensional [ZVC80a, Arn82b, Arn82i, Arn82l, Arn82p, Bas80, BGMY80, BG84b, Bou81e, BF83k, Car80a, Car80c, sC83, Chr80b, Chu82b, CC82e, Chu82c, CC82c, Edw80, Fic80, Gal83b, Gar83a, GHH83, GM84b, GQ82, Gru82u, Guc80, Hal83, HM82d, Isc82, Jun84, Kac82, Kir82, Kol82a, Kre84l, Kre84m, Le 84a, Lin83, Loy83, MR81a, MW80, Miy83a, Oko81, ON81, Ore82, Pap84b, Par83d, PS84c, RC82, Roc84, Sal83, Saz81a, Saz81b, Seb80b, Vek84, Vog80, Wed83, Whi84, dO80a, KS84a]. Dimensions [GJ81, CP81, Fox80, Kör80b, Kow80a, Kow80b, Mar84a, Now82]. d’immersion [Sch81b]. d’indépendance [Dur80]. d’indices [LR83d]. D’Inéquations [BF80k]. d’intégrabilité [Air83]. d’intégrale [Cai80]. d’intégrales [Aze82]. d’intersection [Bis81n, Mal80a, Mal82b]. d’invariance [Bou81b]. Diophantine [CC84b, Cus81, Hür82b, McK80, Osg84, Sch84a, Sch80e]. Diperfect [Ber81a]. Dirac [Beh82]. Direct [Gek84b, Jos80, Arn82e, BM81b, DJW84, DZH81, EF81d, Har81q, Mc 81, Tom82]. Directed [Rin84b, Hav83]. Directes [Bir83a, Cam80a, Erm80a]. direction [Axe80]. directional [Hem82a, HU83]. Directionality [Red83a]. Directs [Bir83b, Bat81, Fey84a]. Dirichlet [AHK82b, ABHK84, All82c, BBCH81c, Bou84d, Bro82d, Föl81b, Fuk84, GŁ82b, How82, Jan83a, Jos84g, Kol84, Las80b, Mae80a, Mae80h, Net80, Net82b, Net82a, Ōsh82, Sto83a, Tom82]. Dirichlet-finite [Mae80h]. disassembly [AD83a]. disc [AF84, Bou83b]. disconnected [Onn82]. disconnectednesses [Pre82]. Disconnections [Koś83c]. Discontinuité [Lel83]. Discontinuités [MMS81, MS81b]. discontinuity [Ceg84]. Discontinuous [Ber84e, GvdP80d, Men84, ZVC80h, Ell81, GvdP80c, Iva83, Kos83a, Tho84p, ZVC80d, ZVC80f]. Discrete [ABM81, Aul84a, GM80e, Wac82, AV83c, Böh82, Dre83, FH84, Grø83, Kat82, Kin81, Sig83g, Tho84c, Tho84d, Wag82, vHSV81]. discretely [Wil81a]. Discretization [Gek84c, KM83, Die82b, Hen82b, Tho84p]. discretizations [FR83]. discretize [Tal82b]. discrets [GG84, Gal89, Mis82]. discriminant [DL81]. discriminator [Ber83a]. Discussion [ESV83e, Sot82a].

Disintegration [GM82e]. Disjoint [Fra83b]. disjunctive [Jür83b, Mos81a]. disjunktive [Jür83b]. disk [Dom83a, FY81, Hay83]. Diskret [Dra80a, Dra80b, Kne80]. disks [Par82e]. Disordered [HN83, MS83d, Ste83d, Ste83e]. displacement [Dou82, KWW83, Pin81, PW84]. dissipatifs [Dub82a]. dissipation [Daf83]. dissipative [Hal80b, Har81b, Har81a, Har81c, Har81h]. dissipativité [Dub84]. dissociation [Die82b]. Distance [Ale81c, Pet80, Cae80, Gar82b, Ove82b, Sei81]. distance-transitive [Gar82b]. Distances [Gen83]. Distinguished [Kow80d, MP84a, Bal81b]. Distribution [LR83a, Nar84g, Sto82a, Aka83, Ant83, AV83c, BT81, Brü81r, Fle81, HV81a, Hay83, Hoh83, KM83, Min83a, Mol83, MS83d, OR83, Pet83, Ric83, Ruy80, Shi83b, Wan81, Wer81g, Wer81r, Wol81a]. Distributional [Pol84, Bré81a]. Distributions [KKM84a, Ani83, Bec84, Ber83c, Brö82b, Buc84, Cae80, Dri82, JY81, KKM84d, KKM84b, KKM84f, Lia84, Pan83a, RS82b, RS84, vHSV81]. Distributive [Bai81, Ban81a, Ban83a, Hof81c, Rom83]. distributivity [Czé83]. d’Itô [Föl81a, Mas82c]. Divergence [Wal83d, Bro81b, BvS83, GY82]. divergence-free [GY82]. divergent [Ruc82]. divided [CP82, Taf82]. diviseur [Bea82b]. Diviseurs [Sko80a]. divisibilité [Mic81a]. Divisible [Itô82, AP81a, Ant83, Blo82b, EF81b, EF81f, EF82, LeP81, Len84, PY81, Sto82b, Wol81a]. divisibles [GG84, Gal89]. Division [BF83f, Dwi83, Gay80, Shi81b, BF83b, Sni83b, Tig82]. divisor [BKS82, Hag81, Ros81a]. Divisors [McA83e, BEPS81, Som82a]. d’Iwasawa [NQ– D84]. Dixmier [Tau80]. DLR [Pap84b]. do [Mac81c]. d’obstruction [FLM81]. d’obstructions [LP81]. Does [She82p, Des80b, vdK82]. Domain [Man81a, DMPG83, Dyn81, HK80a, HLSW80, Nam80f, Orz82, SS80a]. domaine [Bou81b, MMP80]. Domaines [Vig83, CB84a]. Domains [Lad83a, Sak82f, Aub84, BB84, Erk83, Fad83, Fle81, FW82a, lPBF80, Gin80, Gin83, HK81, Jak80a, Jos84a, Lad81, Lad83b, MR81c, OV82, Pen84, Sak82b, Sak82e, Tam82d]. dominated [Fer81d, Geo81a]. domination [Hig82a, Lau80a, Lau80b, Smy84]. dominée [Hig82b]. donnée [BY80]. dont [Éli84b]. d’opérateurs [Con81a, Lev81, Lum80a]. d’optimisation [Bis81o]. d’ordre [Ava82, Rea83, Seb80a, Tan83b]. d’Orlicz [Mey82c]. d’Ornstein [Bak82c, Mey82d]. d’Orsay [GIR81d]. Double [Ban81c, Krá80d, GP81, RS82a, Tei81a]. doubles [Del81f, Hau80b, Pin80a]. Doubly [Har80c, FHV80, GGHJG81]. downcrossing [Mai81]. drap [Yor83]. drawing [Col84, Jam84]. Driven [San83, Gal83a, SM83a]. droit [Bou82a]. droite [Der84, Ros84e, dV81]. Droites [HH82, Abd83]. DTr [HPR80]. DTr-periodic [HPR80]. Dual [Bou83f, Dac82c, Eke83, Le 82a, Des82, Loh83, RW83, Sch83f, Sou81a, TV84, Vog83]. Dualisation [Sal80]. Dualisierende [Sch82d]. Dualisierender [Sch82c, Sch82e]. Dualité [Ber81d, Spa82c, BBM82f, BW81b, BHL81]. dualities [HN82]. Duality [Con80c, Gie82f, Gie82i, Att84, Ban81a, Bou82c, FJ84, GS82a, Ill83, KWP84, Kür83b, Pis82]. Duals [Con80d, Mad80c, Sid83]. Dubreil [Mal80b, Mal81a, Mal82c, Mal83]. Dubrovnik [BKK82, MS81a]. due [DK80a, Nam80i]. Duffing [BS81]. Dugué [Wol83]. Duhamel [Paq82]. d’un [Avr82, Băn80, Bar80b, Bou81b, Bou82a, Bro84, CR80a, Dem82b, Erm80a, Fey84c, Gui82b, Gui84b, Hig82b, Hür81, Ioc84g, Jen80, Kas81, LB84, Le 84b, Leg82e, Les83, MS81b, Mer81, MMP80, Pie80, Ruí83, Spe81b, Tan83f, Tan83i, Vid80, dM80a, dlP82]. Dundee [Eve80b, ES81, ES82, EL83, Gri84, Wat80, Wat82]. d’une [BY80, Bea83a, Bel83a, BMR81, BJ83b, Cha84a, Del81f, DG83, Duv83b, GR83a, dMG83, GV84, Jeu80g, KS83b, Lev82a, Mey81b, MB82b, Rea83, Roq82, Ser80, Tan83h, Tan83g, Wei83]. d’univers [Sou83]. Dupin [Miy84]. dures [Ros84e]. Durham [Wil81b]. d’utilisation [Hei80]. dx [Jak80b]. dyadic [Soa82]. Dynamic [GM80e, Ort80, FS83, Hal84]. Dynamical [Glu80, Hen81b, NR80, SM83a, Ali84, Bar82d, Bra84b, Foi80, FG84, Hal83, Mel82, MS83d, PS84a, Tan81e, Tan81h, RY81]. Dynamics [Aul84a, Pal83, dV84, Fra80, GL84a, Gis84, NT82, Ver84l, BS83e]. dynamiques [Pri82]. Dynkin [HPR80].

E. [Esn81, Lei81]. E.D.S. [MB82b]. E.N.S.T [KMS81]. each [Gat84]. easy [Joc81]. Ebene [Ben80c]. Eberlein [Flo80d]. Éclater [Dem80c]. Eclidean [Wil82c]. École [DPT80, FMSW83, Hen80d, Hen81a, Hen82c, Hen84a]. economic [BFG82]. economical [MK83]. edge [CL84a, Neu81]. edge-regular [Neu81]. edges [Fic80]. Ee [phAD80b]. effect [Har81k]. Effective [Dav81a, Lou83b, Osg84, SY84, Enr82, Mor82a, Ros81a]. effectivity [Mor84b]. effects [Daf83, Red83a]. Efficiency [Fri80, HI83]. Efficient [GM81, SSTW82, EFG82, Wes82]. efficiently [Sta82a]. effilement [Anc84b]. eigen [Itô80]. Eigenfunction [CdS81a, FK81]. eigenfunctions [Atk80, Brü81r, FJ81, Goo83f, Var84]. Eigenschaften [HP84a, Dra80e]. Eigenvalue [BLS80, Kno83, Men83, AB83, AF82, BN80, Eva83, Gre82a, HS84, HLSW80, IS81a, Jen83, Mit82a, NO83, PS81, Ruh83, Saa83, Sch82p, gS83a]. eigenvalues [AF82, CdS81l, CW80, Edm82, Fle81, LO83, Min83a, Sch83i, Sig83d, Var84]. Eigenvectors [Nus81, CW80]. eight [GHMN80, Her80b]. Eighth [McA81]. ein- [NC80]. einem [DK82b, Küh83a]. einfach [HP84e]. Einfacher [Lün81]. Einfluss [Lei81]. Einheitskreisscheibe [phAD80a]. Einleitung [Gru82m, Sch84f, Sch83m]. Einschliessungssätze [Col81b]. Einstein [Bin83, Bou81i, Hit82, Loo81a, Mon80a]. Eisenstein [Bru81b, Bru81e, Bum84e, Chr83c, Chr83f, Sch83o, Sch83r]. Eisensteinkohomologie [Sch83k]. Eisensteinkohomologieklassen [Sch83o, Sch83n]. Eisensteinreihen [Sch83p, Sch83r, Sch83t]. Eisensteinseries [Chr83e]. Ekeland [DM81d]. elastic [HHB84]. elasticity [Ant80]. electrical [ŁK83e]. electromagnetic [Kre81c]. electron [Bar82d]. Electrostatics [Wer81e]. Element [Tho84e, Ber83b, BGM^{+}80, CG83b, DH82, DT80, DKT84, GLCS80, Hen82b, MM82a, Mor82b, Ove82a, Rei82a, Rei82b, Whi84]. élémentaire [Ceg81, Cha81b, Mey81e, Tig81]. élémentaires [Del84b]. Elementarily [Kue81a]. Elementary [Gla81, GGHJG81, HHM83, Hop83f, Hop83o, JTW83, LG81a, Mar82b, Rup84b, Ung82, Bai82a, Bak82a, BDL80, DM81e, Dui82, Kem83, Lev80i, Lev80n, Lev80t, Man80, Pan83a, TN81]. Elementen [Ros84c]. Elements [Bro82c, Est83a, Hop83s, Bac80, BMP80, Dri82, HT81, Itô80, Kan83p, Mon80h, Mon80g, Slo80c, Zou83]. Elimination [Ber80c, Len84, BMP80, Duf82a, Wei84]. Elliptic [Gro80b, Ivr84c, Str84f, Bar80e, BH81b, Böh82, Bur83b, CG82, DG84b, Duf82b, Fle81, FW82a, Gal82a, GS82b, Gre82a, Hab80a, Hem82a, HS84, Hul83, Kam82, Kaw82, Moe82, MP83b, Slo82, SSTW82, Tho84k, Tho84l, Whi84, dF82a, CC82d]. elliptical [Skr84]. elliptically [Mis84c]. elliptique [Ser80, dM80a]. elliptiques [Tal83]. embeddable [Per81a]. Embedded [Ion84, Orb84, Sig83d]. Embedding [Fai82a, Hav83, HI81a, Meu82, Ohs84, PR84d]. Embeddings [Ven81, Bro81e, CR81, Don82b, Fra84, Miy84, Nam80h, RF84, Ros83b, SS81b]. èmes [And80]. empirical [Bor83b, Csö81, Dud81, Dud84, Ing80, Rüs80]. empirique [Har80a]. empiriques [BD80a]. enclosing [CA82]. Encounter [BS83e]. end [Kau81a]. Endlich [HP84b]. Endliche [HP84c]. Endomorphism [Alb83a, Alb83b, Arn82d, EM83a, Göb83a, Göb83b, Hau83a, Rei83a, EM83b, Hau83b, Rei81, Rei83b]. endomorphisms [Mel83]. ends [Jac80a]. énergie [Anc84a, Bou84f]. Énergies [Lan84a]. Energy [Mae80b, Wer81f, GS82a, Gol82, Mae80h]. Energy-finite [Mae80b, Mae80h]. Enflo [SC81s]. engendré [Avr82]. Engineering [Kom84b, Bre83a, Hen81c, Kom84a, KI84, ŁK83e]. enlargment [Jeu80g]. Enleitung [Sch82f]. Enriched [Gra82d, Nel82]. Enright [Bou81a]. Enriques [Con82a, Con83]. ensemble [Ble82]. Ensembles [BF80j, Cos82b, Gru83, Leg82c, El 80, Gru80, LP84b, MR81b, Wat84a]. entières [Est84, Gay80]. entiers [Ceg81, Cha81b]. Entire [KLR^{+}84, Blo84a, Kis83, Ost83]. Entre [Tan83e, Del81a, DS80, Dem82a, Las81f, Mey82c, RP84]. entries [Wal83c]. Entropy [Edm82, BGMY80, BK83, Guc80, Kat80b, Man81c, Pis83]. entry [HL81]. enumerable [FS81d, Joc81, Mil81c]. enumerated [Ket83b]. enumerating [Ket83a]. Enumeration [FR81, Cam82a, Cam82b, FKM83]. enumerative [RS81]. envelopes [Shi81a]. enveloping [And81a, Cha80a, Jos83b]. enveloppante [Lev82a]. enveloppantes [Ale81b, Lev80a]. enveloppes [Cou81]. épaisseurs [Tal82a]. epi [Wet83, Att84]. EPI-convergence [Att84]. epi-limits [Wet83]. epidemic [GM83a]. epine [Gal82b]. Epstein [Chr83a]. Equadiff [KS83d]. equal [Del80b, Mic83]. Equalities [Buc84, Bak84c]. Equality [Fli82c, Car83a, Van83c]. Equation [BR80c, Cer84a, Ame82, BP83, Bas80, BS81, Beh82, Bin83, Bra82a, Bro82a, BHS81, BF83e, Car80e, Cas84b, dBC80, Chi84b, Chr80b, CCN82, Coh81, Cus83, DK82a, DH82, ESV83d, Elm84, Fer84, Fro82a, Geo81b, Gre84, Hal81, Har81h, Hem82b, Hol81, Hön82, Jer80, KP84a, Kre81c, MW84, MB82a, Men80b, Men82, MB82b, MP83a, Mol82b, Neu84b, Oli80, Pal84, PS83b, RW84b, Rus82a, San83, Sau80, Spo84, Sta84, Ste81b, Tam82j, Tes83, Tho84c, Tho84d, Tho84n, Wer81n, Wil82a, Zwe84, Car80d, Cha83t, LY83, Mey81b, Rea83, Ser80, Tro82a, Wei83]. equational [Jež83]. Equations [AGP81, AMT82, BR80a, BR80b, Eve80b, ES81, ES82, EL83, GR83b, Har81i, Hen81e, Hin82, Imm84a, Izé80, Mar80b, MdJ82, Mas84d, Mat80, Min83h, Par82b, Sch80i, Str84f, dFH82, AF81, ABM81, All82d, Amb83, Ami80, AHS82, Aul84c, Aul84b, Bak82b, Bar80a, Bar83a, Bar83b, Bar80e, Bec80, BS83b, BR81a, Bor84c, Bra80, Brä82b, BGM^{+}80, Bro80, Bro81c, BCP83, Bru84, Bum84a, Bum84d, BV83j, CMA83, CD80, CG82, Cat84, Cha83s, Cho80b, CC82d, CC84b, Clé84, Col83, CPD81, Col80d, Coo84, Cow83, Cuy84e, Da 84, DI83, DMT84, Del82b, DGZ84, Det84, DKS84, Die82b, DMPG83, DM80c, Dră83a, Duf83, EG81, Eas82a, Eas83, ES83, Eke83, Elb82, Eng82, Eng83b, Erb83, Fas80, FK83a, FR83, Foi80, Fre82b, Fri82, Gal83a, Gar82a, Gea82, GG82, GP83d]. equations [GBC82, GO82, Goo83a, Gra82b, GLCS80, Gri82, Gru84, Har81k, Har81j, HR82, Hel84a, Hel84b, Hen81c, Hen81h, Hey80b, Hit82, Hön80, Imm84c, Imm84d, IK80, IV80, JM81a, Jef80, Joh83b, Kan80b, Kar81, Kas83, Kas84, Kat80a, Kni80, Koj81, Kos83b, Kot81, KY80, KB80, Kre81b, Kun81a, Kun82, Kun84, LM80, Las81d, Le 84a, LO83, Llo83, Lon84, Lum80b, Maj84a, MNP81, Mas80, MW84, Mei80, Mik82a, Mil80b, Mil82, Miy83a, Mon80a, Naj84, Neu82a, Noc82, Not82d, Not82e, Not82f, Pan82, Per82b, PP82b, Rav84, Rob81b, Rot84l, RCC80, Sap84, SS80a, Sch84a, Sch82l, Sch82o, Sec82, Seg83, Sib80, Sle80, Sot82l, SS84b, Str83a, Tal82b, Tho84k, Tod82, Ver84a, Vin84a, Wac82, Wer81u, Wil84, Wol82b, Zeh83, dCR83, dM83, vdH82, Ber80e, CR80c, LB84]. équations [Le 83a, Mey81g, Sch82n, Yan83b, BV81, CR80b, CR80d, Duv83a, Éme80b, Lum84]. Equi [She82c]. Equi-Consistency [She82c]. equicontinuous [Don82g]. equidistribution [Kin84c, Shi83a]. èquilibre [Ros84e]. Equilibria [Aul84a, BF84, Dre83, Van84]. Equilibrium [KS80a, Dui84, Emc81, HD83, Hen81g, Hen81j, Wer81g, Wer81r]. equimultiplicity [HO82]. equioscillation [Tre84]. equipartition [Gol82]. equisingularity [Cam80e]. Equivalence [Ahl81a, CR84b, Dem83b, ES83, Gui80c, HHM83, Her82a, IV80, LP84b, Mag80, PD82f, Spa83a, Lic80b, Mil83a, Mil83a]. equivalences [FS81b, Ang81e, Sch80p]. Equivalent [SC81c, AHL81b, Kue81a, MV80, MV81b, SC81b, SC81h, SC81s]. Equivariant [FHM82, Liu80, Uhl82, BJ80, DR80b, DK82c, JS80, Kre84n]. équivariantes [Sal82a]. Erdos [ARMS82]. Erfahrungen [Wei82a]. Ergänzende [Dra80d]. Ergebnisse [Sch82i]. Ergodic [Ahl80, Bel81a, Bel83b, CP83b, De 82b, Fuk83, Fur81, Hen84b, Jaj84a, Jaj84b]. Ergodicity [KN80a, BE80]. ergodiques [Erm80b]. Errata [Min83g, Sak82c, Šče81b, Sch82m, Sch82m]. Erratum [Kin84a]. Error [BL81, Böh82, Dah82, Gil82b, Rei82b, Sha84a, Ste82b, Tho84n]. erste [phAD80b]. Erweiterungen [Jür83b, Sch84i]. erzeugte [HP84b]. Espace [Las81a, Tan83f, Bak84d, Bou84a, Cam80a, CMS83, Dem81, Dub82a, Geb83, Gue80, Gun80a, HH82, Kor80a, Lan80d, Leg82e, Lel82a, LBL81, Poé80, Saa84a, Tan83i, dlP82]. Espaces [BF80c, Bou84c, Ehr83, Fey80a, Jur83a, Las80b, Led83, Leg82f, Rou81a, Bar80f, Bea81, Cră83, Éli82, Éli84b, Éme80c, Fer81a, Fey80b, Hus81a, Led81b, Lem81a, LR83d, Mey82c, Pis80, Sch80x, Sch80w, Smy84]. essay [Hof82c]. Essays [Ale81a]. Essential [Bon82b, Dun82, Ada82, Atk80, Don82c, ELZ83]. Essentially [MS82a, Sot82d]. est [Del81f, Las81a]. estabilidad [Lew82]. Establishing [Hen80b]. estimate [Ceg83, Gaf83, Tay83, Wah82]. Estimates [ESV83f, ESV83g, Gul83, Kas84, Dem82b, DM81e, ESV83c, Jos84j, Jos84b, Jos84a, Jos84c, Lon84, Rei82b, Sam82a, SW83, Tho84f, Tho84i, Tho84n]. Estimation [Col80b, Ale83a, GP81, Hoh83, Ivn83, Kat80b, KMM83b, Sha84a, Ste82b]. Estimations [CB84a, Dem82b, Tal83]. estimator [CCV82]. estimators [Aka83]. Estivo [Cer84a, Con82b, Ghe83, GS84, MM82b, Sal84b, dV84]. Établissement [CR80a]. Etale [FP81, KS83a, Sot82a, Sot82b, Sou82, KO81b]. états [Lio83]. etc [Dem80a, Lem81a, vOV81j]. Etcheberry [Hes83]. etendu [Bou81b]. étendus [Hir84b]. ETH [HP84h]. Etherington [Ket83b]. étude [CMS83, Kas81, KS83b, Tro82a, Air83, BF80c, Bak84d, Cha83t, CR80b, CR80c, Duv83b, Har80a, Rol82, Roq82, Ruí83, Wei83]. Études [CC82f]. Euclidean [Aso84, Ben80a, Ber82a, Jur80, Kle84b, Kow81b, Ove82b, Str81c, vdL84]. Euler [Hop83g, Mar84a, Tul80]. evaluating [Rus82a]. evaluation [MM82a]. évanescents [Tei80b]. Evans [phAD80c]. Evanston [FS81e, NR80]. Evarist [Gin80]. Even [Kol82a, Gly83, Kre84l, Sel83]. even-dimensional [Kre84l]. éventail [Mer80b]. Eventails [Bry80]. éventuelle [Bar81a]. Every [SSW83, She82p, Sot82d, BS80b, BP84b, Wil81a]. everywhere [DK82a, Oku84b, Oku84c, Oku84f]. Evolution [Har81i, Bak83, Bin83, BCP83, Cas84b, DMT84, DM80c, Har81k, Hen82b, Hor82a, Jef80, KP84a, Kot81, Lum80b, Mar84a]. evolutions [Gol82]. Exact [Mat82, Sig83c, Bün81, Hil81b, Lev80l, MM82a, May80, Mdz81, Nat84, App83]. exactly [Pic80b]. example [AP81b, Kon82a, Per83a, Tan81c]. Examples [Arn82e, CM82b, Eel80, GvdP80e, Has83d, Hen81c, Kan83b, Küm84a, Nam80d, Ple83d, Bie81, Con83, Dar82, Gir84b, Hej83e, Hen81m, Ish83, KT84, Lou84, Mil81b, Mol83, Pan83b, Ren80e, Som82b, Sot82k]. exception [Hul82]. exceptional [BP84b, Kle84d]. exceptionnels [Gru83]. excessive [BB83b, Cor83a]. excessives [de 82a, dlP80]. Exchange [Bar83c, Arn82o]. Excision [GWL81]. excitations [Bai82b]. excluded [KT83]. Excursion [Gie82j, BP84b, McG84, Rog81b]. exemple [Léa83]. Exemples [Fer81b, Tal80, Éli84b, Hei80]. exercise [San83]. Existence [BJ83b, Ces83, Cor83b, ESV83h, Eng83b, Hen81d, Hig82a, Imm84b, JM81a, Jer80, Jos84h, Kos81b, Kos83a, Kow80e, Lem81b, Lic80b, Maj84e, Men80b, Mon80h, Mon80g, PD82b, Tal80, Wer81g, Art84a, BP83, Bax83, Ben84e, Cor83c, Har81f, Itô84b, Jos84g, Jos84f, Jos84a, KP84a, Kau81a, Lip83, Loh83, Mar83a, SC81j, Str82b]. existence-uniqueness [Bax83]. Existentially [Bou80]. existenzbeweis [Dit80]. Exit [Mue82, Lia84]. exotic [Kre84a]. exotique [Sie80b]. Expanded [ZVC80h]. expansion [Bul84b, Ste83c, Wim80]. Expansions [Abh82u, Alb80, Böh82, Bum84f, Cam80d, CdS81a, Dră83a, FK81, Goo83i, Lan82b, Rus84a, Wal83c]. expansive [CR80e]. experiment [GF84]. Experimental [Dup80]. experiments [GT84, Kea83]. Explicit [And81a, BS83b]. explizite [Gru82j, Gru82i]. Exploding [Dem80c]. expoential [SV82]. exponent [GM80b, GHMN80, Her80b, Iva83, New84]. Exponential [Atk80, AHS82, Bra84a, Dra81, DO82, Hoh83, KM83, OR83, Pic80a, Shi83b, Sle84]. Exponentialfunktion [End84]. exponentiality [Csö81]. exponentially [GH82]. exponentials [HNP81]. exponentiation [Mac81e]. exponentiel [Gay80]. Exponentielles [Com82e, Com82d]. exponents [Cam80c, MP82, Sah83]. Exposants [Gui84a, Com82g, Led84b]. exposé [Del82a, Éme82, Éme83, Len80a]. Exposed [Sid83]. Exposés [Ano80, Ano81b, Ano81c, AY81]. expresions [DH83]. Expressions [CdS81g, Min83h, And81a, Atk82, EN83, EW83, GV84]. Ext [EH81b, She82k, Lad81]. extend [BN80]. extended [Hes83]. Extending [BN81, BP81b, Gup84]. Extension [Ale83b, Ark84, Bou84b, Com83, Don82f, Don82e, Don82d, JS82b, LP81, Tig81, Bie82, Bor82c, Brü82b, Cai80, CKMN83, Čob84, Efr82, Ern81a, IW80, Kau83, Kov83, KS83f, Lau83a, Lem80, Mis84b, Mok83, SS80b, Wil81g, Yam80, Str80b]. Extensions [BBM82e, BT82c, Bre81b, CMH80, Gio83a, Gio83b, Goo80, Kru84b, Lin84a, Wan81, Whe80, BT82d, Bie80, BDR80, BD80b, Blo81b, Bré81a, BH82c, Bro83, BHS81, BF83l, CdS81b, CdS81c, CdS81i, Doi84, Eas83, Fey82b, Fre82a, Hod84, Hue82, Jür83b, Kau81a, KN80a, Kun81b, Mai81, ON81, Rin84f, Rui81, Sch80m, Wil82c, vW80]. extent [MS83a]. Exterior [Mag83]. external [Duv80]. Extraction [Bro82a]. extraordinary [HPR84]. extrapolation [Da 84, Del81a, Mak80a]. Extremal [KT83, Hue80, Tam82d, Tay83, vW80]. Extremalprobleme [Küh83b]. Extremalproblemen [Küh83a]. extremals [Fom84b]. Extreme [Bou83g, Jam83]. extremes [dHR81]. Extremum [Tam82f].

F [Dem80a, phAD80c, SM83b, Lel82a]. F. [Bar81a, Hus81a, Mél80]. F.V [EL83]. Faber [And84a]. Faces [BF80d, Fra83d, Ver84c, Ver84f]. Facettes [BF80d]. Facialement [Ioc84c, Ioc84g]. factor [Hor80]. Factorial [Rob84a]. Factorialité [Bay80, Bay81a]. Factoring [Hos82, Rus84b]. factorisation [Pri84]. factorizable [EF81b]. Factorization [Hof82a, Joh82, Kau83, McD82, Bro83, Coh82, Mal80c, SC81c, SC81q, SC81d, SC81n, SC81q]. factorizations [AC83, Rei82a]. factorized [CC82e]. factors [AA84, Kan84c, MT82, MV81b, Oku84e, Oku84d]. Facts [She82d, Rup84b, Tau81]. faiblement [Abd83, Dur80]. Faibles [Las81a, Yan83b]. fails [Lev80f]. faisant [Gui80b]. Faisceau [CR82]. faisceaux [BBM82a, BBM82e, Cha81b, Dem83a, Die82a]. faithful [Lin84b]. Fall [phAD80b, Spa83a, Spa83b]. Familie [Sch83j]. Families [Ket83b, MP81, Mus83c, CRW^{+}80, Daw82, KP80, Ket82a, Ket83a, Rin84a, Tei80a, Vol81]. Familles [Ang81b, Bar80d, Del84a, Tei80a]. family [Daw81, Kan83c, Kan83d, Kom80, Pet80]. Fano [Col82a, Col82b, Mur82a]. Far [Jef80]. fast [Mit82a, O’M83, Ove82b, SSTW82]. faux [Lan80d]. feasibility [LTS82]. feasible [GL82c, Rob83b]. features [Las81d]. Feb [AMT82]. February [BKS83, Car82, HN83, MTT81, Rao81]. Federal [AD83b, Bak84a, Mac81c]. Feedback [Tan81d, Bar83b, Glo83]. Fefferman [Cho84m]. Feller [Sto80d]. Fenchel [Zie81b]. Fermat [AGT84, Sch84d]. fermé [Mas83b, Mél80]. fermées [Cer81]. fermés [BO84, CB84a, Dem84]. Ferrero [Oes80]. ferromagnet [Abr82]. Feuilletages [Con81a]. few [Jaj84a, Sch82o]. Feyel [Tal82a]. Feynman [Aze82, CS80a, Man83, Sim80, SH83]. Fiber [Her82b, Kol81b, FH81]. fiber-preserving [FH81]. Fiberings [Ivr84c]. Fibration [Tan83h, Tan83g, RP80]. Fibrations [Kre84g, Cor81a]. Fibre [Jaw81, Bis81p, Jen80, Leg82d, Leg82e, Les83, Jen80]. fibrées [Sul81b]. Fibres [Leg82d, Le 83b, Ray81, Rou81b, Băn80, Bar80c, Bar82a, BB81d, Ghi83, Gol83a, Leg82a, Abd83, DS80, Dem82a, Dem82b, Sko80b, Sko80c]. Field [DP82b, Rui81, Ber82a, Bro81b, Des82, Duv80, Gra82e, Kat83, Kij80, KR83c, Lau83a, Mac81d, Mar82a, Mik82b, Pas83, Riz84, Rob83a, Roc84, SS83, Sot82k, Tau81, War82j, Wel83]. Fields [Car82, Hop83s, Kos81e, Mäk80f, PV83a, PR84b, Pre84a, Tra82b, Aky82, AHK82a, Ami82, AG80, Bak84b, Bec82b, Blo81a, Bon82a, Bou82b, Bou82c, Bro81a, BvS83, Cha83b, Che84a, Coa81, CL84d, CMH80, CB84b, Del80b, DP83b, Eas82c, FS81c, Gal80, Glu80, Gro82, Grz82, GY82, Hab80b, HM83a, Her82b, Hum80g, Hur82c, Jef80, Kat82, Kos80a, Lie82, Maj81i, Mäk80c, Mäk80k, Mäk80l, MP81, Man82, Mas84d, Mil81a, MJ82, Mon84, PV83b, Pre81, PR84g, PR84a, PR84c, PR84f, Pre84b, Pre84c, Pre84d, Pre84f, Pre84h, Pre84k, Ray83, Sus82a, VVV81, Ver84e, Wei84, Zel81d, vdL84]. Fifth [CBJS83a, CBJS83b, Eve80b, MdJ82, Kar81]. Figueroa [HS82]. films [VLA83]. filter [CS84a, Fra81]. filtered [WK80a]. Filtering [MM82b, Miy83a, DR82a, Fer84, Gri82, Haz81a, JS84, Kun82, Mit82c, Mit83, Par82b]. Filters [Bae84, Heb83, JS82c, Wei81b]. Filtrage [Yor81]. Filtration [Jeu80g, Bry82, Jac80b, Wal84]. filtrations [NS81b, Zhe84]. Filtrées [Cau80a]. filtres [WK80b, Lau83b]. final [OW83]. finding [Cur80a, Tam82i]. fine [Fro82c, lPBF80, RP84, RU82]. finement [Fug82]. Fini [Sak81, Avr82, Bas82, BO84]. Finie [Cau80a, Gab81c, LS82a, Zay82]. Finies [Die83k, Gal82c]. Finis [Mor81b]. finishing [vOV81d]. Finite [Arn82f, AV83a, Gra82e, Hod84, KB80, Lev80m, McK83, Mon80c, Mor82b, Ple83f, She81a, Tho84e, Zel81c, Zie81a, AHL81b, Arn81, Arn82b, Arn82i, Arn82l, Arn82p, AV83b, Bea82a, BR81b, Bou81e, Bou83k, BGM^{+}80, Bru80, Bur83b, CKMN83, CG83b, Cho84a, CC82e, Clo80, Cur80b, DP82a, DH82, DT80, DKT84, DRS80, FS84, Fer81d, Fis81, FJ81, Gab80, Gab81b, Gly83, GLCS80, HL81, Hap81, Hen82b, Hir83, HD84, Kol82a, Kum81, KS80b, Lev80c, Lev80r, Lev80u, Lor81a, Lor81b, Loy83, Mac81b, Mae80b, Mae80h, MM82a, MP81, Mar83b, MS83b, Oko81, Oli81, ON81, Ott81, Par83d, Pas80, Pra84, Rei82b, Rob81c, Rog84b, SSW83, Saz81a, Sch82b, She82i, Sot82d, TH84, Tha83, Tho81b, Tri84, Wed83, Whi84, Wil81d, Wil81g, Zel81d, Zie81d, Zie81c, vL81]. Finite-difference [KB80]. finite-dimensional [CC82e, Oko81]. Finitely [Fer81d, Hin84, Bau84b, Bea83c, Ber83a, Can84, Cze84, Geo81a, Kna83c, Mas82a, Rei81, HP84b]. finitely-additive [Mas82a]. Finitely-generated [HP84b]. Finiteness [GM80b, Ill83, Cha83h, Cha83n, Fer81e, Sig83g, YV81]. finitist [Mae81]. finitization [YV81]. finitude [Ang81g]. Finland [LR83a]. Finnish [CBJS83a, CBJS83b, Leh83b]. First [Dri82, EF81d, EF81c, Kan83e, KY84, Mat80, PPS83, Zel81a, Gek84d, Gek84f, Gek84g, Goo83k, GM80f, HU83, Not82d, Ohn84, ON81, Rav84]. first-order [HU83, ON81]. fit [Csö81]. fitted [GH82]. Fitting [Sot82e, BC82a, Win82b]. FitzHugh [Rot84e]. Fixed [Gau81, Geo81a, Had82, Jia81, JS81, Mon80c, vDS83, All81b, App83, Bel81a, BG84b, Bro81f, Cel84, FH81, Fou81b, Izé82a, Mon80h, Mon80g, Mon80i, Mon82, FF81]. fixed-point [App83]. fixes [Mas82b, Spa82e]. Fixmengen [Nic82]. Fixmengen-Systemen [Nic82]. Fixpunkte [Col81b]. Fixpunktprinzipien [Hof81b]. Fixpunktsätze [App83]. flag [KT84, LS83b]. flame [FN82]. flames [vH83b]. flares [HLSW80]. flat [BK81c, Kar80a, Lop84]. Flatness [Jon82, HO82]. Fleischer [SZ81a]. flight [Ste83d]. floating [Ste84b]. floating-point [Ste84b]. Florida [GMSV84]. Flot [Mey81b, Bis81n]. Flots [Bis81f, Bis81a, Léa83]. Flour [FMSW83, Hen80d, Hen81a, Hen82c, Hen84a, Ano81a]. flow [Ami80, Bro81b, DO80b, HM82d, KSS80, Maj84a, MP83a, Par82e, SJ82, Wac80, Wac82]. flows [Aso84, Bab80, Bem80, Bou83a, Chu82c, Fas80, FW80a, Goo83m, Gut83b, How82, KB80, Kun84, Man81c, dO80a]. ^{≪}mécanique [ZM84a]. fluid [Bab80, GL84a, KWW83, Maj84a, Ste83d, dV84]. fluid-state [Ste83d]. fluide [LB84]. fluids [BGM^{+}80, CN82, DMPG83, KS80a, Val84c]. Fluktuationen [Ros83c]. fluorescent [TAP84]. fluorescent-labelled [TAP84]. flux [GQ82, Gue80, Hey80a]. Fock [Col80c]. folds [Som82a]. foliation [Rob81b]. Foliations [EHP83, Sun82, Gar83b, Gir84b, Hol80, HK80b, KT82a, NR84]. following [Mun80]. Foncteur [Ang81c, Ang81j]. Foncteurs [Vus80, GJ81]. Fonction [Bel83a, Cha84a, Las81a, Saa84a, Duv83b, Hig82b]. fonctionnel [CMS83]. Fonctionnelles [Dra83d, Gay80, JY81]. fonctionnels [Fey80a]. Fonctions [Ber84a, Ble82, Bre83e, Bre83f, Bre83g, Dra83c, El 80, Fer81b, Fug82, GL82a, Le 84b, Nob81, Ava82, Bar82a, Ber84f, Cha84d, CR82, Dim83, Est84, Fer83a, Gav83, Gay80, Geb83, Gui80b, Lel82a, LS82a, Web83a, de 82a, dlP80, Nor80a]. fondamental [Del81f, Duv83a]. fondamentales [Bar80d]. forbidden [McK83]. forced [Har81h, Lev80b]. forces [Elm84]. forcibly [Cho81, Gan81]. Forcing [Mek83a, She82d, She82f, She82g, She82h, She82m, She82n, She82o, She82j, Mek83b]. Foreword [DK80c]. Form [Gru82l, Sch80i, Bou83k, Bro81a, Bum84a, Chr83a, De 83, Fuk84, Goo83k, Goo83l, Kåg83, Kri81, KK83, Sta81]. Formal [Bro81a, BDL80, BV83d, DRR81b, Jür83b, Mae81, Miy81e, Sib80, Dra83e]. formale [Jür83b]. formalism [Her80a, Mey81a]. formalisms [Kup80]. Formalizing [Bee81]. Formally [PR84a, PR84b, Pre84a, Bec82b, Pre84c, Pre84d]. formation [KJTB80]. Formations [NQ– D84]. forme [Bar80c, GV84]. Formel [phAD80c]. formelle [AT82a]. formellement [CT82]. formelles [Bay80, Bay81a]. Formels [Dra83e]. Formen [Sch83j]. Formes [Ang81d, Cer81, CTCMR82, Gér81a, Rea83, All82c, Ang81g, Des84, Oja82, Vog82, Cho83b]. Forms [Bru81f, Bum84c, CdS81b, Rin84e, AHK82b, ABHK84, Bor84a, Brz81, Bum84b, Bum84f, Cas83b, yCV81, CKLR82, CS82d, CS83b, Cus81, Enr83, Fre82b, GPS83, Gér80, GvdP80i, HG82b, HB84, How81b, HTW81, HJ82c, Kle82b, Mac81a, Nam80d, Pre84c, Pre84d, Raw80, Rin84c, Sar84, Tom82]. Formula [Fli82f, Goo83h, Hej83c, Art83, Ber82b, Bru81p, CM82a, Duf84, EN83, Fal84, Fit84, Fli82e, Fli82g, Gil81b, Gra81a, Gre82c, Hej83a, Hej83f, Hej83g, Kis84b, Kna83a, Kun81b, Mae80b, Maj81g, Mäk80b, Mäk80c, OW83, Pin84, Ple83c, Wet83]. formulae [AHE81a, AF82, Goo83k, Goo83l, Kos80b, Mos84b, OW83]. Formulas [Mäk80g, DO82, JW82, Kle82b, Lin82a, Mot81]. formulated [Bin83]. Formulation [Bis81g]. formulations [DKT84]. Formule [Aze82, Cha84a, Paq82, Ale81b]. fortement [DFM82, Erm80a]. Fortes [Las81a]. found [vdT81]. foundation [Bri81a, Mos84a]. Foundations [HL83b, Kol83, MSW80a, MSW80c, Gue84, LG81b, YV81]. four [Gir83a, Hur82c, LT84a, Sal83]. four-dimensional [Sal83]. Fourier [Art80a, Bea83a, Bru81g, Bru81f, Bum84f, CM80b, CP81, Des84, Goo83e, Goo83i, Her81, Joh83b, JM82a, Kuo80, McG82, Mha84, Oku84b, Oku84e, Oku84d, Oku84g, Oku84i, Oku84h, Osh81, SS83, Sam82a, Ste83c]. Fourth [FG83, IP83b, Mar80b, EG81, Eas82a, GS82b, vdL84]. fourth-order [Eas82a]. Foutz [Csö81]. FPF [Fai82b]. Fractal [SM83c, VLA83]. fractals [BD84]. Fraction [Lan82b, Chi84b, Ger84, JT82]. Fractional [Ble80, Ble83, Kal80]. Fractions [JTW82, Bre81c, Fra83c, Gil82a, Gil82b, Jac82a, JW82, JM84, JW84, JS82c, JR82, JS84, Mag84, MR84a, Ove82a, Rei82a, Sni83b, TW82b, TW82a, Vol81]. fragment [Sch84a]. fragments [Par81]. Framefields [Kos81c]. frames [Ban81b]. framework [KP82]. framings [Kre84f]. France [BLS80, BMR81, CC82f, Rao80, RS83a]. Franchis [Mar83c]. Frankel [Dem81]. Fréchet [AP81b, Sid83, Tra84]. Fredholm [Bor84c, Krá80e, Mar81g, Mil82, Sap84]. Free [Arn82f, Ban83a, Ber83b, HV81b, Lad83a, Ran83a, Val84c, ZVC80c, Alb83a, Alb83b, Art80b, BM81b, Ben84e, Bot83, BR81c, BN80, BN82, Bro81b, BvS83, Col80a, GY82, Hil84, Hil82a, Hof82c, JK83, Jon82, Jos82, Jou80, Kea81, Lad83b, LS81a, Lev80w, MZ83, Mat82, Met83, MP83a, Moe82, Pic80c, Pro81, Ran83b, Van83a, vdB83]. Freedman [LT84a, Sie80b]. freedom [Art84a]. freely [Ket83a]. freeness [Höp82]. Freie [AJ81, Hof81b]. freien [Ros84c]. French [A’C80, Abd83, Air83, Ale81b, Ale83b, All82c, Anc80, AT82a, AT82d, AT82e, AT82f, AT82g, Anc84b, Anc84a, And80, Ang81a, Ang81h, Ang81c, Ang81d, Ang81e, Ang81f, Ang81g, Ang81i, Ang81j, Ano80, Ano81a, Ano81b, Ano81c, AR80a, AR80b, Arq84, Ava82, Avr82, AGY80, Aze80, Aze82, BF80b, BF80a, BF80c, BF80d, BF80g, BF80e, BF80f, BF80h, BF80i, BF80j, BF80k, BM82a, BM82b, Bak82c, Bak84d, BD80a, Băn80, Bar81a, Bar80b, Bar80d, Bar80c, Bar82a, BY80, Bar80f, BS83a, Bas82, Bay80, Bay81a, Bea80, Bea82b, Bea83a, Bea83b, Bea81, Bel83a, BB81c, Ber80b, Ber84a, BV81, Ber81c, Ber84d, BC82b, Ber81d, BBM82b, BBM82d, BBM82a, BBM82c, BBM82e, BBM82g, BBM82f, Ber80e, Ber84f, Bes81, Bet83, BJ83b, BEY81, Bir83a, Bir83b, Bis81a, Bis81c, Bis81b, Bis81d]. French [Bis81e, Bis81f, Bis81g, Bis81i, Bis81h, Bis81k, Bis81l, Bis81n, Bis81m, Bis81p, Bis81o, Bis82, BW81b, Ble82, Bos80, Bou81a, Bou84a, Bou84b, Bou84c, Bou81b, Bou82a, Bou84f, Bou84d, Bou84e, BBP82, Bre81b, Bre83c, Bre83d, Bre83e, Bre83f, Bre83g, Bre83h, Bre83i, Bre83j, Bre83b, Bre83k, Bro84, Bro81d, BHL81, BO84, Bry80, Bry82, Cai80, Cal81, CG83a, Cam80a, Car83c, Car80d, Car81c, Car81b, Cau80a, Cau80b, Caz83, Ceg81, Cer81, Cha84a, Cha83a, CB84a, Cha84d, Cha83t, Cha81b, Cho83a, Cho83b, Cho83c, Cho84l, Cho80a, CMS80, Cho83d, Cho84m, CR80b, CR80a, CR80c, Coe82, CMS83, CT82, Col80b, Com82a, Com82c, Com82b, Com82d, Com82f, Com82g, Com82h, Con81a, Cor81b, Cor84a, CR82, Cou81, Cră83, DZM83, DG80b, Del81a, Del84a, Del80a, DI81a, Del81f, Del81d]. French [Del81e, DI81b, Del82a, DFM82, Del84b, DS80, Dem82a, Dem82b, Dem83a, DG83, Dem84, Dem80b, Dem80c, Dem80d, Dem80e, Dem80f, Dem81, Der84, Des84, DM81d, Die82a, Die83b, Die83a, Die83c, Die83d, Die83e, Die83f, Die83h, Die83k, Dim83, Dla80, Dor81, Dra83b, Dra83c, Dra83d, Dra83f, Dra83e, Dra83g, Dra83h, Dra83i, Dub82a, Dub82b, Dub84, Duf81, Dur80, Duv83b, Duv83a, Duv83c, Ehr83, Ehr84, El 81, El 84, El 80, Eli84a, Éli82, Éli84b, Éme80a, Éme80b, Éme80c, Éme81, Éme82, Éme83, Erm80b, Erm80a, Esn81, EV82, Est84, FSY82, Fer81c, Fer81b, Fer81a, FMSW83, Fer83a, Fer83b, Fey80a, Fey80c, Fey80b, Fey82b, Fey82a, Fey84a, Fey84c, Fey84b, Föl81a, Fon83, FM83, Fra83a, Fug82, FLM81, Gab81c, Gal82b, GG84, Gal89, Gal82c, Gal81, GL82a, Gav82, GŁ82b, Gav83]. French [Gay80, Geb83, GR83a, Gér81a, Gér81b, GB84, Ghi83, GIR81d, Gir81b, Gir81a, Gir81c, Gir83a, GM82b, GM82c, dMG83, GV84, GPV82, Gra82a, GH84, Gru80, Gru83, GJ81, GP83e, Gue80, Gui80a, Gui81, Gui83, Gui80b, GWL81, Gui82b, Gui80d, Gui84b, Gui84a, Gun80a, Gun84, GP81, Har80a, HH82, HYZ83, HZ84b, Hei80, Hel80a, Hén80a, Hen80d, Hen81a, Hen82c, Hen84a, HM82a, Hig82a, Hig82b, Hir80, Hir84a, Hir84b, HM82c, Hür81, Hus81a, Ioc84d, Ioc84a, Ioc84e, Ioc84f, Ioc84g, Ioc84i, Ioc84h, Itô84e, Itô84c, Itô84b, Itô84a, Itô84d, Jac81, Jac80b, Jac84, JG81, Jen80, Jeu80b, Jeu80c, Jeu80d, Jeu80e, Jeu80f, Jeu80g, JY81, Jur83a, Kah80, Kah84, Kas81, Kas82b, Kis82, Kis84a, Kla83a, Kla83b, KS83b, KO81b, KLM81, Kor80a, Kow81a, Kri82, Kum84b, LB84].

French [Lan80c, Lan84a, Lan80d, Lan84b, Las81c, Las81b, Las81e, Las81d, Las81f, Las81a, Las80a, Las80b, Lau83b, LT84b, Le 83a, LY83, Le 84b, LeL82b, Le 82b, Le 83b, Léa83, Led81a, Led81b, Led83, Led84b, Leg82b, Leg82c, Leg82d, Leg82e, Leg82f, LJ80, LJ82, Lel80, Lel82a, Lel83, LDS83, Lem81a, Lem81b, Len80a, Len80b, LLP80, Les83, Les80, LC80, Let81, Lev80a, Lev81, Lev82a, LBL81, Lic80b, Lie80, LP81, Lio83, LR83d, Loo81b, Lum80a, Lum84, Mal80a, Mal80b, Mal81a, Mal82b, Mal82c, Mal82d, Mal83, Mal81b, Mar81b, Mar81i, Mas81, Mas83b, Mas82b, Mas82c, MR80, Mau82, MMS81, MS81b, MR81b, Mél80, Mer80a, MT80a, Mer80b, Mer81, Mey80c, Mey80a, Mey80b, Mey81b, Mey81c, Mey81f, Mey81g, Mey81d, Mey81e, Mey82a, Mey82b, Mey82c, Mey82e]. French [Mey82d, MB82b, Mey84, Mic81a, MMP80, Mil83a, Mis82, Mor80a, Mor81b, Năs80, Net82b, Net82a, NQ– D84, Nob81, Nua81, Oes80, Oja82, Paq82, PP81, Pie80, Pin80a, Pin80b, Pin80c, Pin80d, Pis80, PY82, Poé80, Pra83, Pri82, Que81, Que84, Rab83, RP84, Ray81, Rea83, Ric81a, Ris82, Rod81, Rol82, Roq82, Ros84b, Ros84e, Rou81a, Rou81b, Ruí83, Saa84a, Saa84b, Sak81, Sal82a, Sal80, Sch81b, Sch80o, Sch80p, Sch80q, Sch80r, Sch80s, Sch80t, Sch80x, Sch80w, Sch80u, Sch80v, Sch82m, Sch82n, Seb80a, Ser80, Sie80b, Sko80a, Sko80b, Sko80c, Smy84, Sou80, Sou83, Spa82b, Spa82d, Spa82e, Spe81b, Spr83, Str80a, Str80b, Str82a, Str84g, Str84h, Suc83, Sul81b, Tah82, Tal80, Tal82a, Tal83, Tan83a, Tan83b, Tan83d, Tan83e, Tan83f, Tan83h, Tan83g]. French [Tan83i, Tau80, Tei80a, Tei80b, Tei81b, Tei82, Ten84, Tho81a, Tig81, Tig82, Tit81, Tro83, Tro82a, Ver81a, Ver83a, Ver81b, Vid80, Vig80d, Vig80a, Vig80c, Vig80b, Vig80f, Vig83, Vog82, Vus80, Wat84a, Web83a, Web84a, Wei83, WK80b, Yam80, Yam82, Yan80, Yan83b, Yan83a, Yor81, Yor83, Zay82, ZM84b, ZM84a, Zhe84, dB83, de 82a, dM80a, dV81, dlP80, dlP82]. Frenet [Ber84d]. FRG [DAP82, KS83d]. Fricke [Coh83]. Friedrichs [Kau83]. Frobenius [Gre84, GS83b, Rus84b]. Fröhlich [Rit81, Tay81, Ull82]. Front [DMOyS81a, Hop83h, Hop83i, Hop83j, Sch80f, She82e]. frontier [Lan82a]. frontière [Anc80, Bos80, Har80a]. frontières [Gav83, Mas82b]. Fuchs [GRW81]. Fuchs- [GRW81]. Fuchsian [BLP83a, BLP83b, BLP83f, Goo84, HL83a, Zie81b]. fuchsiens [Sul81b]. Fujita [Esn81]. Full [Duf82a, Gie82k]. Fulton [Del81f]. Funcional [Mac81c]. Function [BR80a, BR80c, Gie82c, GS83d, Hyl81, PR84c, AP81b, Atk82, Bie80, Bie81, Blo81a, Blo83, Bra84a, Bul84b, Din82, Ebe83, ESV83s, FS81c, Fra83c, Gar82d, Gil81a, Gon81, Gri80, Gun80b, HU83, HM84b, JM81b, Kal83, Kis83, Kno83, Mäk80h, Mas84d, Mer84, Mur82b, Pin84, Pop83b, Ray83, Sjö83, Ste83c, Ste84a, Sto80i, Tay83, VVV81, Wag82, Wer81i]. Functional [BF83e, Mae80c, Nel82, Amb83, AG83b, Bec80, BG84c, BPK^{+}84, BvI84, Bum84a, CMA83, dBC80, DeV84, Foi80, GBC82, Goo83a, IV80, Jos84c, KP84a, Kat80a, Kis81a, Kra80i, Mal84, Pas80, RCC80, BKK82, Fuk82a, Izé80, Mac81c]. Functionals [Dac82e, Kom84b, Nor80g, CS80a, CRSSC84, Coo84, Don82a, Fuk81, Kuo80, Kus82, Nor80e, Rus84a, RS84, Sch82a, Sch83b, TW81]. Functions [Ble83, Ław80, Ław83, MSW80c, Sch83e, Sem82e, Tam82f, Abr82, Ada82, All82a, All84, Ana80, AF84, Ars81, AW83, BD83, Bay81b, Bea83c, BT81, Blo84a, BBCH81b, BBCH81e, BBCH81g, BKS82, BE82, BT82i, Boi83, Bre83g, Bre82b, Bum84d, Bum84i, Bum84j, BM80, BR81d, Che80b, CR83, CC84a, CM81, Cór80a, Cor84b, Dia84, DSS83, Dui82, Düm84, Dwi83, Edg82, ESV83i, ESV83a, ESV83r, Efr82, EF82, Fek83, Fic80, FT80, Flo80e, lPBF80, Gau83, GPS83, Gel81a, Gho82, GL81, GWW81, Gök83, GL82d, Haf81, HPR80, Has83i, Hay83, Hen80c, HS84, Hue80, IP83a, Jan83a, JJT80, KLR^{+}84, KS81a, KT82b, Kol84, Kom80, Kov83, KS83f, Kur82, Lan82b, Lan82c, Leu80, Lic84b, Loe80, Lut82, Mae80b, Mae80h, Mae80i, MS80a, May80, MP82, MSW80b, MSW80d, Mil84a, Mon82]. functions [MS83d, Mor84a, Nar84a, Nar84d, Nar84e, ND83, Ore83, Osg84, Ost83, Pet82, Pre83d, Pre83h, Pre83k, Rom83, Sab84, Sch82a, Sch83d, Sem82a, Sem82b, Sha84b, Shi81b, Sic84, Ste84a, Stu82, Sul81a, Sze81, TW81, Tam82g, Tam82j, Tan81i, Ton83, Tra84, Vor82, Was81, Wer81q, Wil83b, Wim80, Wol81a, dBT80, dB84a, dB84b, vdHM84]. functor [HW81b]. Functorial [Cha83m, Aus82]. Functors [McD82, AR80c, BCS81, BB80, Fos80, Jos83a, Mar80a, Mar81f, Rin84f]. Fundamental [Dic80d, Kar81, Par83b, Str81a, ST82, Wer81h, Col82a, Col82b, DL81, DH83, HT82b, Ima83, JvdS84a]. fundamentals [Flo80h]. Funktion [phAD80b, phAD80c]. Funktionen [BKK81, phAD80d, phAD80e, NC80]. Further [BDL80, Cha83n, EF81e, JvdS84b, MSW80c, Par83c, Sem82c, Sta81, Nor80i]. future [Pee84].

G [EH81b, Mer80a]. G. [FL84a]. G.C.H. [She82q]. Gale [Par83b]. Galerkin [GLCS80, Mor80c, Raf83, Tho84e, Tho84o, Tho84p, Wah82]. Galerkin-averaging [Raf83]. Galios [Ull81]. Gallardo [Gal89]. Galois [BF83h, Chi84a, Jan84, Kat82, Que84, Ros83a]. galoisienne [Que81]. galoisiennes [Que84]. gambit [KS84b]. game [Har83, Mar83e]. Games [Sch80g, BJ84, Clo80, EH84, Mos81c]. gamma [Dia84, HV81a, Wan81]. gap [Eng83a, Har81o, Loh83]. Gaps [Mue81, Atk80, KB80]. Gardner [Vol84]. Garnett [Éme81]. Garnett-Jones [Éme81]. gases [Hil82a, Ytr84]. gauche [Les80, Mau82]. gauches [GP83e]. Gauge [Man82, MdJ82, Ne’82, Bai82b, Duv80, Fer80, Kai80a, Pas83, War82a]. Gauge-theory [Ne’82]. Gauss [Bra82a, Bro82a, BF83f, BF83h, BF83i, Del80a, Dra83g, Ehr83, Hop83b, Hop83q, Koi84, RF84, SC81k, SC81d, SC81f, SC81o]. Gauss-summing [SC81d]. Gaussian [Ano81a, Ant81, BŻ80, Che83, Duf82a, Her83b, IJ84, JM81b, Ław82, Pap84a, Pau83, Web83b]. Gaussiennes [Fer83a, Mey84]. gaussiens [Ano81a]. gazing [Bre82a]. GDR [PPS83]. Gebieten [NC80, phAD80e]. Gebuhrer [Gal89]. Geelong [McA81]. gehalten [HP84h]. gekoppelter [Ros83c]. Gelfand [BB80, Gui80a, BV83e, Let81]. gels [AD83a, Whi83]. ’Genaue’ [App83]. General [BLP83b, DMOyS81b, FM84, Gre82d, Har84, Har81d, Hop83b, Hop83k, Hop83r, JvdS84c, Kle84f, Lan80a, Maj84f, Nar84b, Ogu81b, SH81a, Sem82d, Bak82a, Bur80, BF83g, Cat83a, Chi84b, CS82d, Dra81, FK83a, GRW81, Hue82, Ish83, KP84a, KS80a, LR83b, PR84d, Sig83b, Skr84, Spa83a, Sri81, Tho84k, Tho84l, Wac80, Wil82b, Woy83, Zel81d, dGdlV82, Sig83a]. générale [AY81, HH82]. générales [BBM82c, Dem80d, Lum84]. Généralisation [Saa84b, Cor81b, Jac84]. généralisations [Cho84m]. généralisé [LR83d, Net82b]. Generalised [De 83, DM80c, GMJ84, HP84j, EN83, Fin82]. généralisée [Net82a]. Generalities [Har81e, vOV81e]. generalization [CS82b, Ers82, Gra81a, Lag81, Lau83a, Leu80, Mar81c, SZ81b]. Generalizations [BB80, Hop83b, Hop83k, Hop83p, Alt81, Arn82c, Arn82h, Cuy81, Net80, Nut82, Par83c]. Generalized [BB82b, Cla81, Cun83, Dau81, EH84, Fuc81b, Gil81a, GN80, HS81, Ise81, Kow80f, Kow80g, Kow80h, Lac81, Lan82a, LR80a, Mad80c, Min83h, Onn82, Ore83, Rem82, SH84, AB83, AC82, Axe80, Ben84a, Cha84b, Cor83d, DI83, Eis83, Eva83, GN82, Gol81b, HK80a, HK81, Jür83b, Kee83, Kol81a, Kos82, KW82, Kow80a, Kow80b, Kow80d, Kow80e, Kow80i, Lin81, Lip83, Rus84a, RR80, Sch82p, Ste83f, gS83a, Tam82g, Van83c, Wal83c, Zvy84, Dun82, Gru82t]. Generalizing [CCV82]. generate [Jou83j]. generated [Ban83a, CdS81b, CdS81c, Foi80, Her80b, HP84b, Ket83a, Kna83c, Lyu84, Rei81, Sor83, Sta84]. generates [Ket83b]. generateur [Bou81b]. Générateurs [Hir84b]. Generating [Nic80, Sta82a, Cae80]. Generation [Rey80, Gra82e, Rea81, Tri84]. generator [Itô80]. Generators [Are84, Ell80, All83a, And81a, Kre84k, Tri84]. généraux [Gér81b, Jeu80f]. Generic [Ami82, RF84, Sal82b, Cho84e, CJ84, Gat84]. généricité [Fra83a]. génériques [Bar80c]. genetics [Sau80]. Genova [CZ83a]. ’Genuine [EF81f]. Genus [Hop83k, Hop83p, VVV81, Arn82j, CKMN83, Cat83b, ESV83i, ESV83a, KK83, MM83]. geodesic [BB81d, Glu80, Kow81b, Man81c]. geodesics [AG80]. geodesiques [Fer83b, Gér81a]. Geometric [AS84a, Ant80, Bou83h, DAP82, Gie83b, Hen81e, Jos84i, Kra82a, MS81a, Sim80, Tan81f, Bla83, Cor81a, CK84, Fri82, Geo81a, Haw84, Her80a, Hes80, Joh82, Mey81a, Wel83, BF80c, KM80, MdJ82, Pal83]. Geometrical [Bou81e, Day83, GPRS80, Ali82, BT80, Top80, Mar80b]. Géométrie [AT82a, AY82b, Bea80, Cal81, Éme82, GIR81d, Mey81c, Mey82b, Sch82n, Ver83a, Car81c, Gru80, Lan84a, Ver81b, Lei81, CTCMR82]. Geometries [Bue82, Tha83, Per81a, Hal80a, Man84, AJ81]. Géometrique [BF80c, Erm80b, Ang81g, Bis81g, Bar80c]. Geometrische [Sch84g]. Geometry [ABGM82, AV80, BP84a, CGO83, DP82b, Dol83b, FKSW81, Hop83d, Hop83e, Hop83r, Hop83o, Ken84b, Kup80, LW81, Lic84a, Nav84b, RS83a, SC81e, Sot82h, Str81b, vOV81h, vOV82, Ale81c, iA83, Bac80, Bak84a, BT82b, Brz81, Bue81, Bur81, Bur83b, Czy81, Ele82, Elw81, Erd80, Est80, FL81, GvdP80f, Gin82, GT80b, Hir83, HJ82b, Hul83, Hür82b, Kal82, Kir82, McL82, Mok83, Mor84b, Ott81, RS81, Ros84a, Sal83, Seg82b, SS84b, Tan81i, Tan81j, Uen82, Wac80, AV80, LST81, Ves83]. geordneten [Kun81c]. geordneter [BKK81]. George [Fle82c]. Gérard [Coe82]. German [App83, BKK81, Ben80c, Bon80, Col81b, DK82b, phAD80b, phAD80c, phAD80d, phAD80a, phAD80e, Dit80, Dra80d, Dra80e, Dra80f, Dra80c, Dra80a, DK80c, Dra80b, End84, Ewa80, Gru82g, Gru82b, Gru82a, Gru82h, Gru82m, Gru82j, Gru82i, Gru82k, Gru82n, Gru82e, Gru82o, Gru82p, Gru82c, Gru82d, Gru82q, Gru82f, Gru82r, Gru82s, Gru82l, Gru82t, Gru82u, Har81p, Hau80a, HP84a, HP84b, HP84c, HP84d, HP84e, HP84f, HP84g, Hof81b, IvW82, Jür83b, Kne80, Küh83a, Küh83b, Kun81c, Lei81, Lep80, Lja81, Löf80, Lün81, NC80, Nic82, Ost80, Reh80, Ros84c, Sch82c, Sch82e, Sch82f, Sch82g, Sch82h, Sch82i, Sch82j, Sch80a, Sch84b, Sch84d, Sch84f, Sch84g, Sch84h, Sch84e, Sch84i, SP80, Sch83k, Sch83m, Sch83o, Sch83n, Sch83j, Sch83p, Sch83q, Sch83r, Sch83l, Sch83s, Sch83u, Sch83t, Spa83a]. German [Spa83b, Stu80, Tho84a, War82i, Wei82a, Woh83, ZVC80h]. Germany [AD83b, AMT82, Bak84a, BH81a, BJ83a, BS83e, DLR81, Hey82, HJW83, JPW81, Köl80, KMS82, Men80a, Rau80, Rog81d, WB84]. germes [Gal82c, Mil83a, Sal82a]. Germs [DRR81a, Din81]. gersten [Par82a]. Getting [Pre83b]. Gevrey [LR83d]. gewisser [End84]. Gherardelli [FL84a]. ghost [Ne’82]. ghost-gauge [Ne’82]. ghosts [Ne’82]. Gibbsian [DP83b]. Gierer [Rot84f]. Gierer-Meinhardt [Rot84f]. Gine [Gin80]. Girsanov [AHE81a]. given [Böh81, Cer84a, DP82b, Dol83a, GS84, Kan84c, Sal84b, Ves83, dV84]. Glassey [Bar81a]. Gleason [Kru84b]. gleichgewicht [Ros83c]. Global [AF81, BGV84, Bus82, Dre83, FKSW81, Fie83a, Gro80d, Har81i, Par83d, Par83e, Par83f, Rot84g, Rus82b, Sha84a, Sho81, Ste82b, Tan81e, Vig80b, AMPY83, Ant80, Axe82a, Bak84b, BKS82, Čer84b, Dah82, DP82a, FS81c, GL82c, Har81f, LMGC82, Par83b, Par83g, Sch80q, Seg82b, SW80, Sto80i, War82e, FKSW81, NR80]. globale [Gru80]. globales [GV84, Sko80c]. globalization [BT80]. glued [DW81]. GNS [Bot83]. Golay [BJ81]. gold [VLA83]. Goldie [Jos81, Ros81b]. goodness [Csö81]. goodness-of-fit [Csö81]. Gordon [Gar82a, BHS81, Chr80b, Hol81]. Gorenstein [Bor84a]. Göttingen [DK80b]. governed [Gal82a]. graceful [Ach84]. Grade [Mal80a, Mal82b, McA83d]. Graded [Van80, vOV81f, CV82, Czy81, Gre81a, RC82, VV80]. Gradient [Mae80d, Eis83, Fey80c, Ket82b, KI84, Pow84]. graduable [Ple83e]. Graduate [CCCN84]. Graduated [Ple83e, Ple83c]. Graduées [Tan83e]. Graham [San81]. Graiev [Gui80a]. Grandes [Aze80, BBP82, Bro84]. Graph [Erd81b, Kan84c, KY84, Mut83a, Rao81, Ber84b, Bil81, Bil82a, CP84a, DR80a, Gil81b, KR83b, Mut83b, PM83, Rea81, RW84c, Yap84, BKS83, KY84]. graphic [Cha81c, Cho81, Tam82a]. graphical [Buc84, ST84]. graphics [Bar84]. Graphs [Ber81a, GM80a, Sei81, Als81, BP81a, Ber80a, BMKQ83, Bri81b, BMS81, CP83a, CQ81, CQ83, CH84, CL84a, Dic80e, Ell82, Ell83, EH84, Gar82b, Gel83, Hav83, Hol83b, IP83a, KS84c, Kri81, Lyn84, Mag83, Maj83, MS82b, Mic83, Neu81, Pra80, Pra82a, Pra84, Pul83, Qui84, Tay84, WTL84, ZVC80c]. gravitating [Ver80a]. Gravitation [Tra82a]. gravitational [Bou82b, Bou82c, Gal80]. gravity [Des80b]. Greece [PPV80]. Green [Has83e, Eas83, EN83, Mae80b, Mae80g, Saa84a, Wer81i]. Greensche [phAD80c]. grid [DKT84, Hac82a, Hac82c, Hac82b, HR82, Mit82b, SJ82, Wes80, vA82]. Griffiths [Bea80, DS80, Shi83a]. Grilliot [Mos81b]. Gromov [BK81c, Kar80a, Mun80]. gross [Ehr84]. Grössencharaktere [Sch84h]. Grössenklassen [Kun81c]. Grossissement [Jeu80b, Jeu80c, Jeu80d, Jeu80g, Mey80a]. Grothendieck [Arn82g, But81, CS83b, Flo80d, Kai83, Năs80, Nie84, Que81, SC81q, SC81d]. Group [Ama81, Ben84c, BT82d, BT82c, GW81, GLM83, Gre81a, HLR83, HKLR83, HJLR84, Jak82, KN84, Ple83f, San81, STS82, She82r, War82k, Web84b, AP81a, Adi80a, And81a, Bac80, Bak82a, Bal81b, BD83, Ben84e, BT82g, Boi80, Bra84b, Bru81q, Bue82, BF83l, BH82d, CQ81, Col82a, Col82b, De 83, DF82c, Din82, DL81, EP84, FH84, Fis81, GPS83, GvdP80c, GvdP80g, Gre80g, Gre82b, HL83a, Han80, Hel80b, HW81c, Hue82, Hul80, Hum80c, Hum80h, Hür82b, Itô80, Iva83, Joh84, KT82b, Kle84b, Köh82, Kor82, Kos81f, Kre84k, Lor81a, MZ83, Mon80j, Mor81c, MP83b, Nam80g, Oda83a, Oli81, Ple83d, Rit81, Rog81c, Rog84b, Röh84, SZ81b, ST84, Sie82, Slo80e, Sna82, Sni83b, Sri81, Ste82a, Sul81a, Tan81c, Tay82, Tri84, Ull82, Ver82b, Wag82, War82j, Wil81g]. Group [Car82, MS81e, MS81f]. Group- [Web84b]. Group-graded [Gre81a]. group-theoretic [BT82g]. Group-theoretical [STS82]. Groupe [Bes81, Bre81b, Hür81, KO81b, Avr82, Bou82a, Del81f, Éli84b, Gui82b, Kas82b, Las80a, Mer81, Per81b, KO81a]. Groupes [Bel83a, Bir83a, Bir83b, Del81d, Del81e, Jen80, Mor81b, Sou80, Tit81, Vig80d, Bas82, Bat81, BC82b, BW81b, Bou84b, Duf81, FLM81, Gal82b, Gér81b, Gui81, Hus81a, JG81, Leg82d, Leg82e, LP81, Mis82, PP81, Que81, Rod81, Rol82, Spa82a, Spa82e, Sul81b]. Groupoid [Ren80b, HT82b, Ren80a]. groupoids [JK83, Ren80d]. Groups [AJ81, Alb83a, Arn82f, AV83a, BS83c, CV81, CV83, Cor83b, Dic80e, Dic80f, EM83a, Gal89, GP83a, GvdP80e, GvdP80j, Gio83a, GWW81, GHMN80, Hau83a, Hey82, Hey84, Hub83a, Hum80d, Mek83a, Men80a, Met83, Mon80c, Ple83f, Ran83a, Rei83a, Slo80h, Smi81, VW83a, Zel81c, ZVC80h, Zie81a, vOV82, AAS83, Adi80b, Ahl80, AHK82a, Alb83b, Arn81, Arn82b, Arn82c, Arn82e, Arn82d, Arn82g, Arn82h, Arn82m, Arn82n, Arn82q, AV83b, Bak82a, Băn83b, Bau84b, BM81b, BB81b, Ber80d, BJ81, BS83d, Blo81a, BKS82, Boi80, BR81c, Buj83, Bur83b, But81, BŻ84, CR84a, Can84, CL84c, CL84d, Coh83, Con81b, Cor83a, Cor83c, Cor80c, Cor84c, Cor84e, CK84, Cur80b, Dav84, Del81h, Dic80d, Dix81, Djo81, DG84c, Duf84, DO82, EM83b, Ele82, Ell80, EF81a, EF81d, EF81e]. groups [FS81c, Fel80, FP81, Fuc81a, Fur81, Gel83, GvdP80d, Gio83b, GGHJG81, GRW81, Göb83a, Göb83b, Goo84, Gra82e, GM80b, Gui82a, Gup84, Gur80, Hab80b, HTW84, Hau81, Hau83b, Haz81b, Haz82, Haz84b, Hel84a, Her83a, Her80b, Hes81, Hil81d, Hil82c, HP84b, HP84d, HP84g, Hod84, Hof82c, Hol83a, HRW84, HJ82c, HD84, Hub83b, Hue82, Hul80, Hum80b, Hum80g, Hus81a, Jac80a, Jan84, Joh81a, Kac81, Kan84a, Kar83, Kea82, Kee83, Kir82, Kle84d, Kle84e, Kle84j, Kle84k, KS81b, Kol82a, Kon82b, Kra83, Kre84c, Kre84d, Kre84n, Kró83, LR81, Lan81a, Len84, Lor81a, Lus83, Mac81b, Mag81, Mal80c, Mar81e, Mas83a, Mat82, McC82, Mek81, Mek83b, Mel83, Men84, Mon82, New84, Onn82, Ore82, OS81, Orz82, OV82, Par82a, Ple81, Ple83b, Pra84, Pro81, Ran83b]. groups [Rei81, Rei83b, Rhe81, Ric82, Ric81b, Rob81c, Rob84b, Rog84b, Ros81b, Rot84l, Rup84e, Ruz82, Sch84j, Sch80m, Sch83q, Sch83r, Sco81, SH81b, She81b, Slo82, Slo83, TH84, Van80, Ver82a, VW83b, Wal83c, Wei80, Zel81d, ZVC80b, ZVC80d, ZVC80a, ZVC80c, ZVC80f, Zie81d, Zie81b, Zie81c, Zie81g, KN84]. Growth [Goo81b, Sul81a, Gri80, Guc80, Kis81a, Kis83, Mac83, Ost83, PMC83, Wag82]. Grundformel [phAD80d]. Gruppe [Dra80f]. Gruppen [HP84h, Sch83r, HP84a, HP84b, HP84c, HP84d, HP84g, Sch83q]. Gruppenalgebren [Lep80]. GSP [Bou83c]. Guide [Bra82c, LG81d]. Guillemin [dM80a]. Gummersbach [KPT82].

H [CGO84, DeV84, Fle82c, LDS83, Sie80b]. H. [Geo81a]. Hadamard [De 83, Hop83l, Seb80b]. Hahn [FM83, Don82i, Mag84, Yas81]. Haifa [AV80]. Hakone [MTT81]. halb [BKK81]. Halbgruppe [Kun81c]. Halbgruppen [Jür83b, Lja81]. Halbsystem [Sch84e]. half [Art84a, Raw80, Ytr84]. half-forms [Raw80]. half-space [Ytr84]. Hamburger [Cam80d, MR84a]. Hamilton [BT80]. Hamiltonian [BCF82, BCF83, Bor83c, Bot83, Bro80, CS83a, Cap83, CQ81, CQ83, CC82d, Chu82c, Chu82d, CZ83b, Dui84, Fom84a, Kup80, PS84a, Ver83c, Zeh83]. Hamiltonians [CC82e]. hamiltonienne [Bis81o]. hamiltoniennes [Bis81m]. hamiltoniens [Com82h, Ver81a, DM81d]. handles [Kre84b]. handling [Hul82]. Hanges [BLP83e]. Hankel [AAK^{+}84]. hard [She81a]. Hardy [Cha82b, Has83f, Ale81a, BH83, Fek83, Has83g, Jou83h, Sjö82]. Harish [Del81h]. Harish-Chandra [Del81h]. Harmonic [Ben80a, BS83f, Bou82b, Bur82, CV81, CV83, CW82, CK84, FJ84, Gur80, JP84a, Jos84j, Jos84k, Kal82, KT82a, Leu80, Mae80a, Mae80e, Pen84, Sea82, AS84a, Bac83, BE81, Ble80, BH80b, DL82, Der81, Eel80, Gau83, Gra82b, Hue80, Jan83a, Jos84d, Jos84g, Jos84f, Jos84h, Jos84c, Ken81, Leu82, Loe80, Mae80b, Mae80g, Mae80h, Mar83f, Nag82, Pet80, Pis83, Sam82b, Sul81a, Uhl82, Wer81q, KKS82, MRW83, PPV80, RW82]. Harmonique [CV81, Gue80, Kor80a, dlP82]. harmoniques [Ava82, Gav83, Lem81b, dB83]. harmonisch [Gru82c, Gru82d]. harmonische [phAD80d, phAD80d]. harmonizable [PS84b]. Harnack [Anc80, Hue80]. Harrington [Mos81b, vdT81]. Harris [Bou82a]. Hartree [Men80b]. Hartshorne [Dem81]. Hauptsatz [phAD80b]. Hausdorf [Sul81a]. Hausdorff [Hof81c, Hof82a, JvdS84d, Kah84, Lem82, NR84, Sun82, Tan81a]. Hautes [CC82f]. hauteur [Gir83a]. Having [McD82]. Havsbad [KR83a]. Hawaii [GLM83]. Hayman [ESV83a]. heat [SSTW82]. heavy [Aub84]. Hecke [Bum84g, Fre82b, Mos84b, Sco81]. height [Acz81a]. Heisenberg [CK84, DG84c]. Held [AJ81, All82b, ABGM82, AV80, AL81, AL82, BBC^{+}83, BH81a, Ban82, Bec81, BJ83a, Ben80a, BFF80, BOWS82, BS83f, BKK82, CV83, Car82, CZ83a, Cer84a, CW82, CC82f, CGO83, Con82b, Den82a, Den82b, DAP82, DP82b, DLR81, EdJ82, Eel82, Eve80b, ES81, ES82, FF81, FKSW81, Fle82c, Flo80j, FG83, FS81e, Fuk82a, GP83a, GPRS80, Ghe83, HT82a, Hen82a, Hey82, Hin82, HJW83, Izé80, JP81, JV82, JPW81, KR83a, KM80, KZ83a, KPT82, KKS82, Kno81, Köl80, KMS82, KN80b, LR83a, Ław80, LST81, MS81a, MRW83, McA81, Men80a, MM82b, MTT81, NR80, PWW80, PR83, PPV80, PPS83, RY81, Rao81, Rau80, RW82, Ric81c, RSW80, Wat80, Win82a, dBvR81, dFH82, AFG84, AGT84, AD83b, AMT82, BP84a, Bak84a, BDM83, BS83e, BOR^{+}84b, BKS83]. held [Cas83a, CBJS83a, CBJS83b, CCCN84, CP84b, Dol83b, Eri84b, EL83, FM84, GLM83, GMSV84, GS84, Gri84, HLR83, HKLR83, HJLR84, Hey84, HN83, IP83b, Jag84, JTW82, KS84a, Ken84b, KN84, KS83d, KY84, KMS84, Kom84b, Ław83, MO84, MP84b, MR84b, Nav84b, Pal83, RS83a, Rog81d, Sal84a, Sal84b, SW84, TW84, Vin84b, Wat82, WB84, dV84]. Helmholtz [Ste81b]. hemigroups [Sie82]. Henryk [Edw80]. Hensel [Del80b]. Henselian [Pre84f]. Henselsch [Dra80c]. Hereditary [DRS80, Jøn80, Kha82b, Dla83, MV80, MV81b, Min81, Oko81]. Hermite [Dor81, Rus84a]. Hermitian [Fal83a, Frö81, Kan80a, Lan82a, Pan82]. Herz [Onn82]. heuristic [Cor80c, EFG82]. Heuristics [CL84c, CL84d]. Hewitt [Fuc81b]. Heyting [TT81]. HFD [BJ84]. Hierarchies [EHK81, DJW84]. hierarchy [Bal81a, Kec81c, Lou83a, Par80, Spo84, Tan81a]. High [Mil81c, Bas80, How82, Ste84b]. high-accuracy [Ste84b]. Higher [Bor82a, JP84b, Kno80, Seb80b, Sel83, Ars81, DK82c, EG81, GHH83, Sou81b, Ver84a]. Highest [Hum80f, Hem84, RC82]. Highly [Sic84, GG82]. Higman [San81]. Hilbert [Maj84g, A’C80, Ang81i, Bec80, BM80, Cas84b, CdS81c, DKS84, Dub82a, Gir81b, Gir81a, Goo81b, Gra82a, GL82d, HG82a, Har81a, Hop83b, Jaj80, Kah80, Kot81, Las81a, Oda83b, Ore83, Pap84a, PT82, SC81a, Sto80g, dM80a]. Hilbertian [Har81g]. Hironaka [CGO84, Oda83a]. Hirzebruch [Ish83]. History [Maj81a, Bre81c, Gus81, Gus82]. HNN [HD84]. HNN-constructing [HD84]. Hochheben [Sch83j]. Hochschild [DI82]. Hodge [Bry82, Cas83b, Del81b, Del81c, DMOyS81c, LBL81, Miy83c, Ogu81c, Ogu81d]. holoidal [BS83h]. Holomorfia [Mac81c]. holomorph [Gru82d]. holomorphe [Dem82b, GR83a, Jen80]. holomorphes [Bar80f, Dim83, Duv83a, GL82a, Gav83, Geb83, Gru83, Kla83a, Mar81i, Seb80a]. Holomorphic [BB84, Din81, Gil83, Kai80a, Lie82, Nar82, Sch80n, Sun82, Bay81b, BY82, CKK83, Cha83b, Dol83a, Eel80, Gir84b, Gri80, Hol80, Kob81, Kra80i, Dem82b].

Holomorphieaussagen [Sch83t, Sch83p]. Holomorphy [Col80c, Mac81c, Bec82a, Bia81, Bur83a, Kau81c, MN81]. holonômes [Bry82, Dem83a]. holonomic [Kas83, MJ82]. Holonorphic [Ohs84]. hom [Lad81]. homeomorphic [Pra82b]. homeomorphism [DR80b, Trn84]. Homeomorphismes [Cră83]. homeomorphisms [Bou83a, Dat83, GMW84, Par83d]. homoclinic [dCR83]. Homogène [Ioc84g]. homogeneity [Cha84b]. Homogeneous [Arn82h, Kec81a, Lev80n, MR83, Mot81, Bar82c, CS82a, Elm84, GM84b, Gil83, GGHJG81, HG82b, Hel84b, Kit84, KP84b, Lop84, Pen84, Tho84c, Tho84n, TV84, Vor82]. Homogènes [Ioc84c, Éli82, Éli84b, LJ80]. Homogenization [Pin81, SP83]. Homographique [Cor84a]. Homological [Fox80, WK80a]. Homologie [BW81b, Sch81b, Ris82]. homologique [Kas81]. Homology [Cur80b, Bau81, Ben82c, DI82, Fra80, Kra83, Kró83, Mdz81, Neu80, Sie80a, Sna84]. homomorphism [Bon84, Cha83q]. Homomorphisms [HW81c, Cun83, Del81h, Don82c, Sot82a, Sot82b, Sot82c]. homotopic [BH84, Mdz81]. homotopical [HPR84]. Homotopie [Leg82e, Leg82f, Tan83c]. Homotopielineare [Löf80]. homotopies [LR83c]. Homotopique [Tan83a, Bre83h]. Homotopy [Cha83j, HK80b, All81a, Ber80d, Bök84, Bou82d, Far84, Fer81e, Gra82d, HJ82a, HH81b, HT82b, HJ82c, Hus81a, JS80, Kos81b, Liu80, Tod82, Wal84, Leg82f]. homtervals [Pre83b]. honnêtes [Jeu80b]. Honnef [WB84]. honneur [Coe82, LS82a]. Honolulu [GLM83]. Hopf [Bak84c, Fie83a, LMRW82a, LMRW82b, RW84b, Wil82a, Zel81c]. Hopkins [Năs80]. Horizontal [Jer80]. Hörmander [Bis81j]. Horn [Fel80]. Horn-theories [Fel80]. horocycles [Hof81a]. Horrocks [CS82b]. horseshoes [vS81a]. Host [Mél80]. Hua [BV81]. Hudson [Fle82c]. Hukuhara [DM82]. hull [BH82a]. hulls [Ple83c]. Hunt [JM82a, Itô82, Itô84e, Itô84c]. Hurwitz [Gon81]. hybrid [Eng82]. hydrodynamic [Kan80b, MSP83]. hydrodynamical [Dup80]. hydrodynamics [Ioo84]. Hydrodynamik [Ros83c]. Hyper [BB81b]. Hyper-indecomposable [BB81b]. hyperanalytic [Gil81a]. Hyperbolic [Bou83a, Gro81a, Apa83, Aul83, BLP83a, Cho80b, Gal83b, Ghi80, Hey81, KS83e, Mar81g, Mun80, Rav84, Wil84]. hyperbolicity [CS80b, Fri80]. hyperbolique [Las81c]. hyperboliques [Gér81a, Las81f, Sul81b]. hyperbolischen [phAD80b]. hyperbolischer [Wei82a]. hyperdiameter [MS82e]. Hypererlang [Ani83]. Hyperfinite [Hel83, Lin83]. hypergeometric [Pan83a, dB84b]. hypergraphic [Bil82a]. hypergraphs [Bil80, Leh83a]. Hypergravities [Des80a]. hypergroupes [GG84, Gal89]. hypergroups [Blo82b]. Hyperhypersimple [Cho84c]. hypermaps [Fed84]. Hyperplane [Băd84, Som81, Con82a, Con83]. hyperprojective [Kec81c]. hyperspaces [Cha84c]. hyperstonean [Zah84]. hypersurface [Bea82b, Hoc81, Rea83, Shi83c]. hypersurfaces [Cal81, yCV81, Dwi83, Gav82, Miy84, Ohs84, Tak84]. hypothèse [Jeu80f]. Hypothesis [She82p, Gol81b].

I- [DRR81e]. I. [Jøn80, SZ81a]. i.i.d [Goo81b, Hei84]. I.S. [Yan80, Zhe84]. iced [vdB83]. Ideal [JK83, Lau83d, McA83j, Est83b, GK81, Kak81, Mau82, Sen83a, Wil81a, Ble82, Kas81, Laj81]. Ideal-free [JK83]. Ideals [Hil81a, Art82b, Bru82a, Fis81, For83a, Hen83, Hil81i, Hil81l, Jos83b, Kna83c, Lor81a, Lyu84, McA83g, Ruc82, Sot82e, Tau81]. Idéaux [Ber81c, BO84, Eli84a, Lev80a, Yam82, Mer81, Yam80]. idempotent [vL81]. idempotents [GM84b, HH81b]. identical [Blo80, Cae80]. Identification [GP81, Fou81a]. identifying [EFG82]. identities [Bre84, CC82c, HKP83, Ila80, Lep82, Mon80e, Par82c]. identity [Don82h, EN83, Tam82i]. ideology [Chi83]. ihre [Gru82u]. II [DG80d, HKLR83, Spa83b, Wer80, AV83a, AV83b, BF80c, BM82b, Bay81a, Beh82, Bel83b, BBM82g, BD80b, BOR^{+}84b, BS84, Cho83a, Col82b, CR84b, Cus81, Dem80c, Den82b, DRR81e, EP84, Ern81b, FZ80b, Fro82c, GR83b, Gir81a, Goo83c, Hof82a, Joh82, JM82c, Kno83, LeP81, LMRW82b, Pin80c, Ram82b, Rin80b, Tam82f, Tay84, Tei80b, Tei82, Ver83b, Wol81a, Woy82]. II-singularities [DRR81e]. III [Bec81, HJLR84, SW84, CL84a, Cho83b, Dem80d, DRR81f, Gir81c, JM82d, Oli81]. III-singularities [DRR81f]. IIMAS [Hen82a]. Iitaka [Esn81]. Ikeda [Kis84b]. ILL [SC81d]. Illinois [Ama81, LW81, NR80]. im [Gru82j, Gru82k, Gru82n, Gru82c, Gru82d, Küh83b, Ros83c, Sch83s]. IMA [Eri84b, HN83]. Images [Cam80a, Erm80a, Kru84a]. imbedding [NSE82]. immersions [BH84, Ber80d, Bin83, Ecc80, GS83e, Lan84b, Ohn84]. Immeuble [Bes81]. Immeubles [Gér81b]. immiscible [Dup80]. impaire [Hus81a]. Implementation [Eri83, But80]. implications [Bou81e, Sen84]. implicit [But80, Eng82, GP83d, HW82, KY80, PW84]. implies [CS80b, Day81]. Imply [She82p, GP84, Sie81b]. imprimitivity [Fre82a, Hol84]. Improper [She82g]. improvements [Wer81u]. Impulses [PD82a]. Impulsive [PD82c]. Inada [Par83b]. Incidence [DO82, Bou83k, Han80]. including [FO82, Ste83d]. Inclusion [Col83, Ole83]. incoming [BL81]. incommensurate [Aub82]. incomplete [PV81]. incompleteness [vdT81]. incompressible [Bab80, BGM^{+}80, DMPG83, GLCS80, How82, KSS80, Maj84a, Wac82]. increasing [Gho82, Wal81a]. increments [Det83]. indecomposable [Arn82e, BB81b, Jež83]. indéfiniment [Mil83a]. indefinite [HG82b, HS84, MLM82]. indépendante [Nua81]. indépendants [Jac84]. Independence [LS81a, TN81, Acz81b, Clo80, Clo81a, D’A84, Deh80, Deh81, Gie82m, NS81b, Šče81a]. independent [Det83, HTW81, Woy83]. Indeterminate [Abh82a, Abh82b, Abh82d, Abh82c, Abh82n, Abh82m]. index [AMPY83, Bro81f, Bro82d, CMPY83, CZ83b, Fou81b, Izé82a, Kra82b, Mil81a, Mil82, Sil83]. indexed [Bou82d, Ryc84]. indexées [CG83a, Led81a]. Indexing [Abh82e]. India [All82b]. Indian [Rao81]. indice [Bet83, GP81, Per81b]. Indices [KMS81, Bro81d, EG81, Gun80a, MS81b, Mey81e]. Induced [Kle81, Die82b, Röh84]. Induction [BF83h, Mae81]. Inductive [BFPS81, FS81a, Lev80o, Lev80p, Sie81a, BF83d, BF83g, Mar81h, Yas81]. industry [AB83]. inégalité [Cho84m, Ehr84, Itô84d, Len80b, Mey82e]. Inégalités [Fey80b, Gun80a, Led81a, Len80a, BM82a, BM82b, Bro81d, Cho83d, LLP80, Rab83]. Inequalities [Hop83n, Ols81, Str84f, AHS82, Bro83, Che81a, CS82d, CS83b, EW83, Fuk83, GL81, Hen82d, Hor82a, Lac84a, Moe82, Pas80, Pit81, Saw82, dF82b]. Inequality [Hop83n, All83b, DSS83, Dru83, Jac83a, Kai83, Kwo83, Nie84, Sch82b, SC81g, Tam82g, Tam82i, dAKL83]. inertia [Kre84k]. Inf [AA83]. Inf-compact [AA83]. inference [iA83]. inférieure [Har80a]. infini [Bou84a, Seb80a]. infinie [Bak82c, LS82a]. infiniment [GG84, Gal89]. infinis [Lum84]. infinitary [EH80, Wil81d]. Infinite [BS83c, Car80c, Chu82a, Chu82b, Hal83, Kir82, Mad80d, Ruz82, AHL81b, BS83d, BG84b, Car80a, CPD81, Die82b, Edw80, Elm84, Gal82a, Gar83a, GM84b, Hue82, Kač80, Kac82, MW80, Miy83a, PS84c, Rob84b, RC82, Saz81b, SH81b, Vog80, KS84a]. infinite-dimensional [BG84b, GM84b, KS84a]. infinitedimensional [HI83]. Infinitely [Blo82b, BvS83, Has83g, Ant83, CP81, EF81b, EF81f, EF82, LeP81, MT82, PY81, Sto82b, Wol81a]. Infinitesimal [CS80b, Str83b, Bou81b, Itô80, Ves81]. infinitesimale [Web83a]. infinitesimales [Băn80]. infinity [CS83a, IV80, YV81]. Informal [JP81]. infra [Sén83b]. Inherently [MS83b]. inhomogeneous [MP82, Tho84d, Wol82b]. Initial [Gek84c, AS82a, Bjö84, BV83d, CD80, Dah82, Enr82, Gek84a, GH82, Gus83, Jeu80c, KS83e, Las84, Mae81, Ste82b, Jeu80c]. initial-boundary [Gus83]. injectifs [Sal80]. Injective [Joh81b, VW83a, Fai82b, Jon82, Kol82b, VW83b]. injectives [Cou81, Gab81c]. injury [Cho84a]. inmiscible [CN82]. Inner [BS83h, Mon80d, Mon80i, OW83, War82k, ND83, Shi81b, Wed83]. inneren [Ewa80]. input [PS84c]. insoluble [Bau84b]. inspired [FG84]. Instabilité [Rou81a, Rou81b]. Instability [Sig83d, DT80, Dup80]. Instables [Ray81]. instanton [Chu82a]. instantons [Bou82c]. Institut [CC82f]. Institute [Rao81]. Instituto [Pal83]. Integer [Sch80h, Ach81, BEPS81, GWW81, OR80]. integer-pair [Ach81]. integer-valuated [GWW81]. Integers [Nar84g, Ple83f, All82a, Jos82, Lic84b, Sou81a, Sta81, Ull81, vdL84]. Integra [Hop83s]. intégrabilité [AGY80]. Integrability [CC82f, Ber82a, BŻ80, Cor80c, LR80b]. Integrable [Maj84c, BD83, CS80a, Chu82b, Chu82c, DM80c, Fom84a, Gér80, Mik82b, STS82, MMS81]. intégrables [Éme80a, Kla83b]. Integral [AHS82, Bak83, BV83f, Bou83i, Gin82, Goo83f, Jou83f, Krá80f, Min83h, MP83a, Rin84c, Rin84e, Sor83, Tho81b, Bak82b, Bak84c, Bar83a, Ben84a, Bru84, CMZ80, Del82b, Det84, DK80a, Edg83, EW83, Fek83, Gal83a, Gul83, Gus81, Hön80, Hön82, Hor82b, Itô80, JP84b, KS84b, Kre81c, Lon84, Maj81f, Man83, Mei80, Mon84, Oka84, Pav84, PRW84, Rog81c, Rog84b, RCC80, RS84, Saw82, SC81l, Sen84, Sim80, SH83, Tau81, Rog81d]. Intégrale [Cho83c, GŁ82b, Duv83b, Gun84, Lan84a, de 82a, dlP80]. Intégrales [Bis81h, DFM82, Gav82, Jac80b, ZM84b, CMS80, CMS83, Com82d, Fer81c, Lan80c, Mas82c, Mey80c, Sch80x, Sch80w, Com82e]. Integralidentitäten [phAD80d]. Integrals [Mae80a, Maj81b, Wil81b, And81c, BF84, Bis81h, Bra80, CMT^{+}84, Goo83b, Goo83c, Her83b, Kot81, Maj81c, Mar83a, Per83a, Rog81a, RS82a, Ros84d, SS83, Str81c, Tar82, Vol82a, Wim80, ZM84b, Tho84a]. Integrands [Sal84a]. Integration [Gru82a, Hin82, JvdS84d, JvdS84e, JvdS84f, JvdS84g, AG83b, Cec84, Det83, Doz81b, Ell81, Geo81b, HW81a, JvdS84b, Kal80, Lin83, Wol82b, Zak81, vdH82, Bar82a, Gru82l]. integro [Eng83b, Mei80]. integro-differential [Eng83b, Mei80]. integrodifferential [DGZ84]. Intensively [KKM84a, KKM84e]. Inter [MS81a]. Inter-University [MS81a]. interacting [Gal80, HS81]. interaction [CRSSC83, Men82, vH83b]. interactive [Bar84]. Interdiffusion [TAP84]. interesting [Lit82]. interface [CN82, TAP84]. interfaces [Dup80]. intermèdiaire [Bea82b]. intermédiaires [Mas83b]. Intermediate [CdS81d, CdS81j]. intermittancies [SM83c]. International [AFG84, ABGM82, AL81, AL82, BP84a, Ban82, Bec81, CV83, CZ83a, DG80d, DAP82, Eel82, FM84, FG83, Fuk82a, KZ83a, KPT82, Nav84b, NR80, Pal83, PR83, Rau80, TW84, Vin84b, Bak84a, KS83d]. Internationale [GS84, Sal84b]. Internazionale [Cer84a, Con82b, Ghe83, MM82b, dV84]. interpoint [Cae80]. interpolants [GMS84]. interpolating [BT82b, Blo84a, Gru82u]. Interpolation [Cha84a, GMSV84, KWP84, Mey82c, TNV^{+}84, Ber84c, CRW^{+}80, CS84b, Cor81b, CN84b, GM81, GMJ84, Gul83, Her84a, HV81b, JP84a, Jon84, KP82, LL84, Mer84, Mil84a, Mit84, Mun84, PP82a, Pee84, Per84, Sag82, Sem83, Wol82a, CP84b]. interpolierende [Gru82u]. interpretability [Lin84b]. Interprétation [Ang81g, Bre83h, GS82d]. Interpretations [Man84, JG81]. intersection [DG84b, Fal83b, FL84a, Gat84, Hoc81]. Intersections [Ang81e, GS84, Bor84a, DM84, For84b, God83, NiW84, Tak80, Val84b]. Intertwining [Bru81h, KS81b]. Interval [ND83, Pre83e, Bon84, Jak80b, JS82b, SC81f, SC81o]. Intervalle [DK82b]. intervalles [Str82a]. intervals [Lan81b, Har81p]. intervenir [Gui80b]. intracellular [AD83a]. Intrinsic [Rup84a]. Introduction [AT82b, Ang81f, Aul84d, Ben80b, BT82e, Bis81i, BBCH81d, BLP83c, Bru81i, BF83i, Caj81c, Cha83o, Cho84d, CR80c, CdS81e, Con80e, Dac82d, DMOyS81b, DRR81b, ESV83j, Gie82l, Goo83g, Gra83d, Gre80c, Gro80e, Gui82a, GM80d, Hal84, Hen81f, Hop83m, Ioc84f, Ize82b, Jos84l, Kan83f, KMS83, Kle84g, Kos81d, Kre84h, LW82, LC80, Mad80e, Mäk80i, MJ82, Not82a, Ogu81b, Per82a, Ple83g, PR84e, Pre83c, Ren80c, Rot84h, Rup84c, Sak82d, SM83b, Seg82a, Shi83a, Sig83e, SK80, Sto82a, Str84c, War82f, Zel81b, vH83a, vOV81g, vOV81h, AS80, Bai82b, Bak82b, BE82, Fri82, Gut83a, HTW84, HP83, JvdS84c, Koj81, Maj84f, Neu84b, RW84a, SH81a, Tod82, Top83, Ang81f, Bis81i, Ioc84f, Sch82f]. introductive [Cho84l]. Introductory [Krá80g, Cho84l]. Intuitionistic [Ues81, Šče81a, SH84, TT81, Ves81].

Invariance [Ang81g, Tan81f, BHS84, Bor83b, BS83g, Cor84a, Rév81, Sot82b]. Invariant [Bor82c, Brü81r, Dat83, Fos81, Hen81g, Tan81g, AF84, BP81a, Ber83c, Bre82b, CMPY83, Cor84e, Dom83b, Dom83c, Ern81a, Fer83b, Gie83b, GQ82, HL83a, HN80a, Has83k, Hen81k, Ioo80, IACT81, Jak80b, Kac83, Kri81, Los84, Mañ81b, Mik82b, NSE82, Neu80, Pop83b, Sie80a, Str82b, Tak80, Tan81e, Tan81h, Var84, Ghe83, Tho84a]. Invariante [Bel83a, phAD80a]. Invariants [Alm83, Avr82, Bum84h, Gal82c, Hür81, Bak83, Bea81, Cam80e, GF84, Hap81, Haz83, Hil81c, Hun83, HK80b, Ion84, Kah82, KŁLS80, Kim81, Kir82, Kra83, Kró83, Lev80s, Mon84, Ols81, Ped84a, Spr83, Ver83d, Vus80, War82l]. Inventaire [Tan83d]. Inverse [Mar81a, Abr82, Col82c, GMJ84, HG82a, Jef80, LV84, Mik82b, NO83, PdS80]. inverses [Imm84b]. Inversion [AHL83, Her81, Ste81b]. Invertible [Har80b]. investigation [KSS80, Skr84]. Involution [Raţ80, Tig81, Vog84]. Involutionen [Löf80]. involutions [Her80b]. involutive [BLP83d]. involutory [Bac80]. involved [JS82a]. Involving [PD82a, AF81, All83b, BP83, Buc84, GM83a, HN82, JW82, Le 84a, Mil82, PS83b]. Iokalen [Sch82d]. Ionescu [Ers82]. ionized [Hil82a]. Ioue [Dub82b]. Iraklion [PPV80]. Irrational [Sch80c]. irrationnelles [Lem81a]. Irreducible [Kow80i, RS82b, Bil82b, Bur80, FMW82, Fre83, KS82, Ohn84, RC82, Wol80]. irreductibles [Bay80, Bay81a]. irredutibles [Car83c]. irregulîeres [Duv83c]. Irregular [Maj84c, Lev82b, Sib80]. irregularities [Bec84]. irrégulières [Kla83b]. Irreversibility [Cas84a]. Irreversible [AFG84, Gis84]. Ising [Abr82, HM83a, MP82]. Iskovskih [Mur82a]. Ismagilov [Gui80a]. iso [Kla83b]. iso-irrégulières [Kla83b]. Isobars [Abh82f]. Isoclinism [BT82f]. Isolated [Fil84, AMPY83, BY82, CKK83, HN80a, MP84a]. isolées [Com82b, GM82b, GM82c]. isometric [Kre84i]. Isometries [Gre83]. isomorphic [Bau81, Isl81]. isomorphism [Arn82a, Arn82j, Kop81]. Isomorphisms [Ara83, Rog84b, Röh84]. Isoperimetric [Hop83n]. isospectral [CC82d]. Israel [AV80]. issues [CC83]. Itô [Maj81g, AHE81a, Bis81a, Bis81d, Bis81g, Föl81a, Kun81b, Maj81c, Maj81b, Maj81f, Mey81a]. Italy [AFG84, Cer84a, Con82b, Ghe83, GS84, MRW83, Sal84a, Sal84b, Vin84b, dV84]. Iterated [BFPS81, DZH81, Gaf83, Kue81b, Leś84, dAKL83]. Iterates [Pre83e, Pre83d]. Iteration [She82f, CA82, GM84a, Hem82b, Kas82a, Sch82p]. iteration-based [GM84a]. iterations [AS82b, Hac82c]. Iterative [BGM^{+}80, Axe82a, Eva83, GO82, Guz81, Wer81u, AMT82]. iterativen [Gru82k]. itéré [Led81b]. Itpack [KO83]. IUTAM [Rau80]. IV [BH81a, BJ83a, Nor80a, Čer84b, Dem80e, JM82e, OW83]. Ivasev [Kör80b]. IX [BOWS82, Hen81a].

J [Del80a, ZVC80h]. J. [Bar81a, Ber81d, Hel80a, Hus81a, Ioc84a, Ioc84h, Web84a]. Jacobi [BT80, GN82, Mag84]. Jacobian [GvdP80c, GvdP80h, Sot82g, Sot82k]. Jacobians [Par83k]. jacobienne [Bea82b]. Jacobischen [Sch84d]. Jacobson [Mon80g]. Jacques [MP84b, Anc84a, Hir84a]. Jacquet [Bum84i]. James [Edg80, Flo80b, Flo80f]. Jan [AMT82]. Janeiro [Mac81c, Pal83]. January [ABGM82, Car82, GW81, GLM83, Hen82a, Ken84b, MS81a]. Japan [Fuk82a, IP83b, MTT81, RS83a]. Jenkins [DSS83]. Jensen [Gam80]. Jersey [Win82a]. jet [Kup80]. jets [Tro82b]. Jeulin [Mey80a]. Joensuu [LR83a]. Johannson [Jac80a]. John [Bar81a]. join [Jež83]. join-indecomposable [Jež83]. Joint [Rup84d, FJ81]. Jones [Éme81]. Jordan [Ioc84b, Ioc84g, Kau81c, Sch83h, Win82b]. Jørgensen [Gro81a]. Josephson [Cus83, San83]. Journées [AGT84, CTCMR82, Jag84, Rao80]. Juin [Nor80a, Rao80, Ano81a, CV81]. Julius [Lan80d]. July [BBC^{+}83, Bak84a, BJ83a, BOR^{+}84b, CBJS83a, CBJS83b, CW82, DAP82, Eel82, EL83, Gri84, Hey82, Izé80, Jag84, KPT82, Köl80, KMS84, MRW83, Men80a, MM82b, MR84b, Pal83, PPV80, Sal84b, Tur82, Wil81b, dFH82]. Jump [CRSSC83, AHE81b, Cho84e]. jumping [Mar83g]. jumps [DH80]. June [All82b, Ano81a, BDM83, CV83, CBJS83a, CBJS83b, Cer84a, CGO83, Con82b, Den82a, Den82b, DLR81, FF81, Ghe83, GS84, Gri84, Hey82, JP81, KS84a, KMS82, KMS84, KN80b, LR83a, Men80a, MP84b, NR80, PR83, Rog81d, Sal84a, Sal84b, Wat80, Wat82, dFH82]. Jung [Gir83a].

K. [ARMS82, Esn81, MR80]. Kahane [SC81f, SC81g]. Kähler [Kal82, Loo81a, Mok83, Sot82f, Riz84]. kählérienne [Abd83]. Kan [Fre82a]. Kannan [Mas81]. Kantorovitch [Fer81a]. Kapazitätsfunktion.Eine [phAD80c]. Kaplansky [Wil81f]. Karl [BR80b]. Kármán [CR80b, CR80a, CR80c, CR80d]. Karpacz [PWW80]. Kashiware [EP81]. Katata [Fuk82a]. Kawamata [Esn81]. Kazamaki [Yan83a]. KdV [Chu82b, Mol82b]. KdV-equation [Mol82b]. Kelley [Cha84c]. Kellogg [Wer81j]. Kelly [SC81s]. Kelvin [Wer81k]. Kennzeichnung [Ewa80]. Kern [Gru82e]. Kernel [Che81b, BT82d, Her84b, Jou83g, Kis84b, Lau80b, MW80, Sjö83]. kernels [Ale83a, AC82, AA83, Bru84, Clé84, Ers82, HNP81, Itô82, Kwa83, Lau80a, Ric82]. Kharchenko [GPV82]. Khintchin [Fal84]. Kieby [MR81b]. Killing [CB84b]. kind [Bar82d, Han81, Nam80f]. kinds [Han81]. kinematic [Her80c]. kinematics [ADT83]. kinetic [Ame82, Ytr84, Cer84a]. kinetics [DESS82]. Kingdom [GMSV84]. KK [vDS83]. Klainerman [Bar81a]. Klassische [phAD80d]. klassischen [IvW82]. Klein [Buj83, Gar82a, Pin80b, Pin80c]. Kleinian [GP83a, Slo83, Sul81a]. Kloosterman [Goo83e]. Kneading [Pre83f]. Kneser [DM82]. Knight [Mc 81]. Knot [Lev80e, Böh81, Hay82]. Knots [Far84, Fra80]. Kobayashi [Gar83a]. Kodaira [Miy81b, Miy81c, Miy81d, Mum83]. Kohomologie [Sch83l, Sch83q, Sch83r, Sch83s]. Kohomologiemoduln [Sch82j]. Kolmogorov [BD83, Kah80]. Kolmogorovian [GF84]. Köln [HT82a]. Kombinatorik [Sch82h]. Kommutatoren [Reh80]. Kompaktifizierung [Sch83l]. Komplex [Sch82e]. Komplexe [Sch82d, Sch82h, Spa83b, HP84e, HP84f, Sch82c]. komplexer [Dit80, Sch84c, Sch84b]. konfigurationsräume [Hau80a]. konformen [phAD80a]. Konstruktion [Gru82k, Gru82n, Sch82g]. Korea [KN84]. Korovkin [Bro81c, Don82e, Don82g, Don82h]. Körper [DK82b, Dra80e]. Körpern [Dra80c, Dra80a, Dra80b, Kne80]. Korteweg [Fre82b]. Kostant [Jos81]. Köthe [Mad80c]. Kouchnirenko [Mer80a]. Kozubnik [Ław80]. Krein [KW82, Lan82c, Nus81]. Kriterien [Sch84g]. Kronecker [GH83, GHH83, Kåg83, Las80a]. Krull [Car81a, Cha83a, Lev82a, Orz82, OV82]. Kuiper [Sul81a]. Kulikov [Bru80]. Kunita [Mey81b]. Künneth [Ill83]. Kupka [FY81, Rob81a]. Kürzungssätze [BKK81]. Kutta [But80, Gek84g, HW82]. Kutzko [Car81b]. Kyoto [Fuk82a, Ken84b, RS83a]. Kyoungju [KN84].

L [CdS81f, Lan80d, JS82b]. L. [Gal89, Mey80b]. labelled [Ell83, TAP84]. Labelling [vdLT81]. Laboratory [Mol82a]. lacets [Hus81a, Lem81a]. lack [Shi83b]. lag [Le 84b, MS81b]. Łagów [BKS83]. Lagrange [EN83, Tul80]. Lagrangian [Kup80]. Lagrangiennes [Com82e, Com82h]. Laguerre [Mag84]. Lamberto [Gra84a]. Lancaster [Tur82]. Lanczos [AB83, CG83b, CW80, Eri83]. land [Ars81]. Lang [Ray83]. langages [Ric81a]. Langlands [Sch83t, Car81b, MS81e, MS81f, Wil83a]. Langlands-Quotienten [Sch83t]. language [Sot82i]. languages [Jür83b]. Laplace [BH83, Bra80]. Laplacian [AHL83, BK81a, BB81d, Brü81r, Now82]. Large [Hin82, Hop83d, Noc82, Che83, CW80, Daf81, Fum81, Hei84, Hia84, HT81, Hue82, LR83b, O’M82, Saa83, Sch82p, tR84, Aze80]. largenes [EP84]. largest [AS82a, Hag81, Mar83d]. larmes [Mey81c]. Lasry [DM81d]. last [Bro82e, Lia84]. Latin [dFH82, Bou83k]. Lattice [FG83, AL80, Bal81b, CKL83, Con81b, Don82c, Hol83a, Kop81, Kos82, Mar81c, Rog80, SSW83]. Lattices [BH81a, Aub82, AS81b, Ban81a, Ban83a, CN84b, Day83, Det82, Don82b, Don82h, Ern81a, Fra83d, Gie82j, Hof81c, HM81, Kür83b, Leu80, McK83, Nic83, Rom83, Sch81c]. Laurent [Bul84a, Bul84b]. law [BE81, Gaf83, HV81a, Hei84, IJ84, Kue81b, Leś84, Rog81b, dAKL83]. lawless [vdHM84]. laws [Ant83, Art84b, Daf81, DG84c, HK80a, Hia84, HT81, Jan82, KMM83b, LR80a, Mac81e, Mel82, Obe82, PY81, SC81k]. Lawson [Str81a]. layer [Krá80d]. layered [MP82]. LCTVS [LR80b]. leading [Gek84e]. Least [Bar83d, Kea83, Van83c]. Lebesgue [Gal82b, Gro81b, Top80]. Lecture [Gal89, LM88]. Lectures [Ben82b, Cer84a, CGO84, DP82b, Flo80j, GS84, Haz83, Hop83d, JB81, Mil84b, Mit82c, Pre84a, Sal84b, Ves83, dV84]. Leffler [ESV83s]. Lefschetz [Gór81, Mos84b]. Left [Art82b, Lus83, McD82, BE83b, Cor84e, IS81a]. left-definite [BE83b, IS81a]. Legendre [Eve80a, Kac81]. Lelong [Coe82, Kis82, LeL82b, LS82a]. Lemma [Nav84a, She82q, Gal83b, Hes83, Kif80, Lin84a, Maj84b, vdLT81]. lemmas [Gru82r, Maj84e]. Lemmata [Gru82o, Gru82r]. lemme [Lel82a]. Lemmes [Mor80a]. length [Bri81b, Her82a, May80, Miy81e, vdK82]. lengths [Lan80b]. Leningrad [FM84, HN81a]. Lenstra [Bas81]. Leontief [Par83k]. less [CR84a]. leur [Ava82]. leurs [Cou81, Gav83, Sch80r]. level [Axe82b, How81a, MM82a, Wet83]. Levels [BP84b]. Levi [Bay81b, Rea83]. Levin [Oli80]. Levine [Web84a]. Lévy [MS83d, Bou84b, Fal84, Kan82, Kis84b, Mai81, Sch82b, Sen83a]. Liaison [LJ82, Tan83e, Hun83]. Liapunov [dBC80, Coo84, Hen81b]. Liberal [Gre81c]. Librairies [Die83b, Die83c, Die83f, Die83g, Die83j, Die83k]. libre [Spe81b]. libres [Dla80, KS83b]. Lichtenbaum [Sna84]. Lie [Ama81, CV81, CV83, Bau84b, And81a, BO82, Bes81, BW81b, Blo81b, Blo82a, Bou81a, Bow82, Can84, CC82e, Cor80c, Cor84c, Cor84e, Del81h, Djo81, Duf81, Duf84, Enr83, EF81a, EF81e, FH84, Fer83b, Fre82b, FLM81, Fur81, Gal82b, Gui81, Gui82a, Haz81a, Haz83, HLR83, HKLR83, Her83a, HJLR84, Hil82b, HL83b, Hul80, Hum80f, Hur82a, Jos83b, Kac81, Kac82, Kap82, Kir82, KS83b, Kum84b, Lep82, LP84a, Lev80a, Lev82a, Mac81a, Mal80c, Mal82d, Mar81e, Mis84b, Pet82, Ric82, RC82, Rot84l, Sch82k, Sel81b, Sie82, Slo80c, Slo83, Tan83e, Tau80, Ung82, Ver81a, Wil81e, Wil81f, Wil82c, Win81, Win82a]. Lien [RP84]. Liénard [Bar80a, DI83]. Lifting [AAS83, BS80b, BHS84, Kaz83, Los80]. Liftings [Vol82a, And82, BS80a, Los82, Los84]. Liftung [Dra80f]. Ligure [CZ83a]. like [Bus82, Dev81, HN80b, Kre84f, Lin81, Nar84e, Nav84a, Pot82, Wal81b]. Lille [Coe82]. Limit [Art84b, GHZ83, Maj81b, AF82, Att84, Ber81b, Bjo82, DH80, Dri82, FK83b, Gin81, Gne83a, Hei83, Hey81, HM84b, JM84, Led83, Led84a, Llo83, Maj81e, Maj81d, Pau80, Rey80, Roh81, Ros82, RR80, Ryc84, Sul81a, dHR81]. limit- [FK83b]. limite [Gui84b, Yor83]. Limites [Bak81, HM82a, LB84, Led83, Hel80a, Las81b, Le 82b]. limits [Bou82d, DJW84, Des82, Düm84, Gra82d, HS81, Mar81a, Wet83]. Line [Hop83s, Dal83, Jer80, Kan82, MS82b, RU82, Sem82a, vdB83]. Linéaire [Ruí83, Air83, CR80a, Ser80]. Linéaires [BF80k, Cho83b, Dra83d, Bar81a, Ber80e, Dla80, Gér81b, Gui80b, Hir84a, Kla83b, Mey81g, Yor83]. Linear [Bar83a, Dac82e, DP82b, Doy81, Gek84e, Gek84d, HHN84, Hen81h, Hin82, Imm84c, Jun83, Kré81a, Nar84c, Aup83, Bak82a, Bar83d, Bar80e, Ber84c, Bor83b, Bor84c, BvI84, Bro80, CM80a, Car80e, CC84b, CL84b, Cus81, Die82b, Djo81, Dră83a, DH83, Duf83, Eas82c, Elb82, Eng84, GL84a, Gie82m, Gli84, GO82, Gri82, Har81g, Hej84, HLSW80, Hön82, HTW81, Hue82, JN82, Jur81, KKM84c, KKM84b, Kun82, KZ83b, Ler81, Lin83, LS81b, Mañ81b, MW80, Mik82a, Miy83a, Nus81, O’M82, O’M83, Par82b, Pas80, Pav84, PS84b, PS84c, Rav84, Sag80, Sau80, Sch82a, SS80a, Sch83b, Sch82o, Sri81, SS84b, Tam82h, Tan81e, Tan81h, Tod82, Wer81t, Wol82b, Zel81d, dM83, AD83b]. linear-quadratic [Dră83a]. Linearization [Kor82]. Linearized [Pas83]. linearly [Gay84, Gla81, HW82]. lines [GLP81, GS83e, Has83e, Hay82, Iit83, Mar83g, NS81b]. linguistic [BKQ83]. link [Fou81a, Hil81d]. Linkage [LR83b]. Linked [Dev80, BE80]. linking [Poh81a]. Links [Hil81a, Hil81d, Hil81j]. Linksideale [Lep80]. LINPACK [CCV82]. Liouville [App83, Eas83, Min83a]. Liouville-Probleme [App83]. Lipschitz [Bou83a, CM80b, Gol83b, Jou83g, Onn82, Sjö82, Soa83]. lipschitziennes [Nob81]. liquid [Eri84a, Wis84]. liquids [Dup80]. Lisses [Led83]. list [Ver82a]. Littlewood [Jou83d, Jou83h, McG82, Mey81d]. living [Rup84f]. LMS [Wil81b]. Local [Art82c, Dyd81, Fli82d, Gie82m, Goo83h, Gro80f, HM81, Kan83g, Kle84h, Kot81, Kür83b, McC82, Oku84i, Oku84h, Per82b, Rév81, Rob83b, Sha80, Sto80e, Sto80f, Tan81h, Vig80c, dHR81, Avr82, Bar81b, BP83, BK83, BF83b, CV82, Cha83b, CN84a, D’A84, Fuk81, Gui84b, Har81e, Jac83a, Kan83o, Kat81, Kol84, Le 83a, Le 84a, Lum80b, Lyu84, Poh81b, Pra82a, PS83b, Rog84a, Ros84e, Sch80q, Sil82, Sot82d, Sto80a, Vid80, Wei83]. locale [Ang81a, Car81b]. Localement [Bel83a, Cho83a, BC82b, Cră83, Dub82a, Éme80a, Kla83b]. locales [CG83a, Dim83, Geb83, Hyl81, Roq82]. Localic [BV83a]. Localisation [Kol84, Sch80p, Sch80q, Vog82, Yam80, Pop83a]. localisations [Sch80r]. Localization [Hil83, Bie81, Wal84, vOV81d, HP84d, HP84f, HP84g]. Localizations [BV83g, Bor82b, Mal82a]. Localizing [Hil81e]. Locally [Gie82n, Hum80g, Kha82c, Kir81, Kow80j, Ren80d, Ale81a, Băn83b, BD83, Bie81, Boi80, BH82d, Col80c, CM81, Cor83a, Din82, Din81, DM82, Fal83a, Flo80h, Gie82g, Had82, Haz82, Jan82, JvdS84d, Kan84a, Kha82f, Kin82, Kin84c, LR81, LM83, Mat82, McC82, Sch83c, Var84, Wei80]. Locating [BCJ^{+}81]. location [CC82a, KMM83b, Ove82b]. locaux [Dub82a, Gun84, Lem81a, Les83, Lum80a, Sch81b, Yor83]. Loen [JTW82]. Loewy [Lan80b]. Log [IJ84]. logarithm [Gaf83, Kue81b, Leś84, dAKL83]. logarithme [Led81b]. logarithmic [Har81j, Miy81b, Miy81c, Miy81d, Tay81, Tay82]. logarithmique [Ehr84, Itô84a, Kow81a]. logarithmiques [BM82a, BM82b]. logarithmischen [phAD80d]. Logic [BOR^{+}84b, KPT82, KMM81, KMM83a, MR84b, PWW80, Bea83c, Gur84, Len84, LSS81, MTT81]. Logical [DL84, Kol83, LG81b, Röd84]. logics [Cze84]. loi [Har80a, Led81b, Yor83]. Lois [GG84, Gal89, Fer81c, Mey84]. łojasiewicz [Rab83]. lokalen [Kun81c]. Lokaler [Sch82j]. Lokalisierung [HP84d, HP84e, HP84f, HP84g]. Long [BBC^{+}83, Als81, Bre81c, Edg80]. Longitudes [Hil81f]. Longueur [Cau80a]. loop [Hus81a]. Loops [Str81b]. Lorentz [Gre82b, Mer84]. Lorenz [Rob81b]. lorsque [LB84]. Losange [Cor84a]. lost [And81c]. Lösung [Ewa80, Gru82j, Gru82i, Gru82k, Gru82n, Wei82a]. lot [BFG82]. Lotka [Rot84k]. Low [EF81g, Gir84a, Her82c, How81b, Kow80a, Kow80b]. Lowell [KM80]. Lower [Cus84, Dac82e, JN82]. Löwner [Tam82i]. Löwner-identity [Tam82i]. Lublin [SW84]. Luigi [Ves83]. Lukacs [Wan81]. Luminy [CV83, CV81]. lumped [Tho84g]. Lund [CP84b]. Lusin [Gun84, McG82]. Lusztig [Gre80b]. Lyapunov [PD82d].

M [CdS81f, Hwa81]. M. [BD83, Kar80a, Loo81b, Ver81a]. même [Zhe84]. Maass [Sar84]. Macaulay [Mer80a, Ore83]. Machines [Ros84b]. Machlup [TW81]. Mackey [Str83b]. MacQueen [Mos81b]. macromolecules [TAP84]. macroscopique [Ros84e, Tro82a]. magnetic [Beh82]. magnétique [GL82a]. magnetostatique [AR80b]. magnification [Jeu80c, Jeu80d]. Mai [CTCMR82, LS82a]. Main [Nam80e, Beg83]. Major [Cho84f]. Majorantes [Fer81b, Hei80]. Majoration [Bar80c, DG83, Fer81b, Ber84a]. Majorations [Pra83]. Majorized [CS82d]. Majorizing [Hei83]. make [Ber84b]. Malliavin [BF82, BJ83b, Bis81j, Mey81b, Str81d]. manifold [BK81c, FHY83, Hen81k, Lic84a, SS84a]. Manifolds [Aul84a, Ivr84c, LW82, Not82c, Ver84f, Ale81c, ADT83, Apa83, Aul83, Bak83, Bjö81, BG84b, CR80e, Dar82, Dav84, Der81, Edw80, FP83, FW80b, GN80, Gro81a, Gut83b, Ham82, Hen81g, Hes80, Hit82, KŁLS80, Kan80a, Kar80a, Kos80a, Kow80c, Kow80j, Kow80l, Kow81b, Kre84f, Kre84a, Lac81, LS83a, LS83b, LT84a, Mon84, Nav84a, Not82b, Ohs84, Pet80, Pie83b, Riz84, Rup84f, Sto83b, Tak80, Tar82, Tog81, Zie81d, dO80a]. manifolds-with-boundary [Ale81c]. Many [Sig83h, Böh81, CP81, DH80, Kör80b, MT82, Ton83]. Many-Body [Sig83h]. many’moduli [BvS83]. map [BH84, Koi84, Lod81, Wil81g]. Mapping [Zie81a, Aup83, CKK83, Don82h, Gel83, Jaw83, Kre84j, Obe83, Par83a, Woh83, Zie81d, Zie81c, Zie81g]. Mappings [Jos84k, ŁK83d, Ver84h, AS84a, AHK82a, AF81, All81a, Car83b, Čob84, Dev80, Fru83, HRW84, IACT81, IK80, Jos84e, Kee83, Kir81, Kis83, ŁK83a, Mar83f, Mil82, Par83k, Pet80, RS83b, Ric83, Sie81b, Ver84d, Ver84i, Ver84k, Wal81b, Zvy84, vDS83]. Maps [Jos84k, KKS82, Pre83e, Sto82a, War82k, BE81, BY82, Ber80d, BGMY80, Bur82, BE80, Čer84b, CR80e, Dar82, Eel80, FH81, GM82d, Gra82b, Guc80, HH84, HTW84, HZ84a, Hus81b, Jak80b, Jaw81, Jos84d, Jos84h, Jos84j, Kal82, Ken81, Leu82, Mañ81b, MS82c, Nag82, RF84, Sam82b, SC81a, SC81b, SC81d, SC81l, SC81t, Sto83b, Szű80, Uhl82]. March [AV80, Eri84b, Eve80b, ES81, ES82, EL83, FS81e, KR83a, KM80, MTT81, WB84]. marches [Éli82, Éli84b, Gui84b]. Marcinkiewicz [CS83b, HT81]. marge [Éme82]. Margherita [CZ83a]. Marie [Mal80b, Mal81a, Mal82c, Mal83]. Marked [Zie81b]. Markov [Fuk82a, Ark84, Bas84, BBCH81e, Bou81b, DJ80, Der84, Die83a, FS83, FG84, GM84b, GQ82, Gui84b, HM83a, Küm84a, Lac84a, Ōsh82, Sto80f, Sto83a, Str82b]. Markovian [Düm84, Le 82a]. markovienne [KLM81, Gui84a]. marriage [Mil84b]. marriages [Hwa81]. Marseille [CV81, CV83]. Marshall [AV83c, OR83]. Martin [Anc80, Bos80, Has83h]. Martingale [Bur81, Chi83, SC81h, BY80, Fuk81, Kos83a, LMRW82a, LMRW82b, Sch80p, Sch80x, Sch80w, Wal82b, CW82]. Martingales [Bel81b, Bis81j, Dar82, Gho82, HZ84b, Jeu80g, Nua81, RU82, Sch80r, Sch80s, Sch80t, Sch80v, AGY80, Bak81, Bis81d, Bro81d, CG83a, Cha82a, Cha82b, Che81a, Cho83d, DZM83, Ell81, Gal83a, Geb83, Gun80a, Gut83a, GP81, HW81a, HYZ83, Jac83a, Led81a, Len80a, LLP80, MMS81, Sch80p, Sch80q, Sch80x, Sch80w, Sch80u, Sch82n, Sha80, Woy82, Woy83, Zak81]. Maryland [Ben80a, HLR83, HKLR83, HJLR84]. Maslov [GH84, Per81b].

Massachusetts [KM80]. masses [Tho84g]. Massey [And80, Tan83b]. Massless [Des82]. matching [She81b]. Matématica [Pal83]. Matematico [Cer84a, Con82b, Ghe83, GS84, MM82b, Sal84b, dV84]. material [Chr83b]. Materialized [Fum81]. materials [MS83d]. Mathematical [DAP82, DP82b, GPRS80, IP83b, KM80, Kom84b, MSW80c, Rob83a, Vla82, Emc81, Gra83c, KKM84f, Kin84b, Leh83b, Par82d, Rab84, vdT81, CZ83a]. mathematicians [Est80]. Mathematics [BOWS82, BS83e, Cas83a, Erd81b, HN83, LR83a, LM88, McA81, Ric81c, RSW80, Bee81, Dav81a, Fef84, Goo81a, Gra84a, Sel81a, Ser84, YV81]. Mathematik [HP84h]. Mathematik-Departement [HP84h]. mathématique [Caz83]. Mathématiques [DPT80]. Mathieu [BJ81, Bue82, MSW80b, MSW80c, Tes83]. Matiyassévitch [Mar81b]. matrice [dV81]. Matrices [Mad80d, Bou83k, CC82e, Coh82, CW80, DG80a, De 83, Fie83b, GM84b, Gir83b, Gui80d, Gui84a, Har80b, Haz81b, Hos82, Küm84a, Lan82a, Le 82b, PM83, Par83h, PdS80, Seb80b, Wil82a]. Matrix [Cor84c, Kwo83, Roj80, Ara83, Bro82a, dBC80, Dem83b, Dră83a, DH83, Dro80, Duf82a, Duf82b, GP83c, Gek84e, Gil81b, Ila80, Mad80a, Mal82a, Mol82a, Tho82, Wal83c, dM81a, KR83a]. Matroid [Daw82, Daw84]. matroids [Bor83c, Daw81, Nic80]. Matscience [All82b]. Matter [DMOyS81a, Hop83a, Hop83h, Hop83i, Hop83j, Kle82a, Meh83a, Sch80d, Sch80f, She82b, She82e]. Maurey [SC81q]. Maximal [Arn82i, Bea82a, Cha80a, Cho84g, Cór80a, Gun80b, Art82b, Art82c, Blo83, CdS81f, Fuk83, Jou83h, Mun80, Nor80f, Pit81, Saw82, Sch83n, Sjö83, War82g, Wil81g]. maximality [Rot84d]. Maximaux [MR80]. Maximum [Tho84f, Wer81l, Arq84, Dub82b, EFG82, Hir80, Jer80, MR80]. Maximum-Norm [Tho84f]. Maximumprinzip [phAD80d]. Maxwell [Gar82a]. May [AJ81, AGT84, BFF80, CGO83, HJW83, JV82, JW83, Kno81, KY84, Kom84b, LW81, Vin84b, Win82a, vO80]. Mazur [Coa81]. McMullen [Vol84]. Mean [Hop83k, Hop83p, BEPS81, sC83, Kit84, Leu80, Tak84]. meaning [BT80, Geo81a]. means [Bre82b, Mad80g]. measurability [Per83a]. Measurable [BM81a, Gra80, Sch82a, Bie80, Edg82, Han84, Iof80, Kak81, Los80, Mah84, Mar84b, Thi84, Wag80]. Measure [BS80b, Bie80, Con80c, Fro82c, KS84b, LM83, Ada82, BPV83, Bel81a, Bie82, Che81b, Csö81, DL82, Ers82, Ghi80, Gil81c, GM82e, JvdS84e, Lem80, Mar83f, Pap84a, Sch83g, Sen84, Sul81a, Vol82b, Whe80, vW80, BDM83, Köl80, KMS82, KMS84]. Measure-fine [Fro82c]. measurements [Hol84]. Measures [Gal89, Hey82, Hey84, Sch80g, Ali82, All83b, AG83a, BR81a, Blo82b, Blo84b, Bor83b, BK81b, But82, BH82d, BŻ84, Che83, Con80d, Con80h, CS81b, Cor83a, Dat83, Det82, Die83d, DU83, Fil84, Gam80, Gar82c, Gar82d, GQ82, GM83b, Hei83, Hul80, JM81b, Jaj80, Jak80b, Jan82, Joh83b, Jov84, Jur80, Jur81, Kar83, KKMW81, Kin82, LR81, Ław82, Lem82, LMW80, Lin84a, Loe80, LR80b, Lou83b, Mae80d, Maj81h, Mas82a, Mis84c, MN82, Pap84b, Pra82b, PT82, RRM84, RS82a, Ruz82, Sch83h, SC81i, Sie82, Str82b, Woo83a, dG84a].

Mecanique [Bis82, Bis81k]. Mechanical [Sig83h, Dri82]. Mechanics [Rau80, AHK82b, Att84, Bis81k, Bri81a, Car80b, Col84, Emc81, Gli84, Gro82, Gue84, HD83, LP84b, Noc82, Str81c, dF84]. mechanism [Cuy84b, Dup80]. Medford [Bec81]. Media [HN83, Dou82, Pap83, PW84, SP83, Ste83d, Ste83e]. medium [Jer80]. Meeks [Mor81b]. Meeting [AMT82, Kom84b]. Mehler [Sjö83]. mehrfach [phAD80e, NC80]. Meinhardt [Rot84f]. mélangeant [BD80a]. Mellin [Bum84j]. Melrose [Hel80a]. membranes [DL82]. memory [Shi83b]. ménage [Fos81]. Mennicke [Sus82a]. Merit [BT82i]. meromorpher [phAD80e, NC80]. méromorphes [Duv83a]. Meromorphic [Gér80, Sto82a, Boi83, Gri80, JJT80, KLR^{+}84, Kov83, Sto83b]. merotopic [BH82a]. Mesurables [Tal80]. Mesure [dMG83]. Mesures [Cho83b, Die83d, Fer81b, FM83, Hei80]. Metabelian [New84]. metaplectique [Per81b]. Method [Hop83b, KKM84a, Abr82, Axe80, BT80, BL80a, Bra82a, CC82a, Cha83s, Cho84a, CRW^{+}80, CS84b, CA82, DH82, DT80, Dou82, ESV83l, Eis83, FJ84, GL82c, GLCS80, Hem82a, Hey81, Izé82a, Jac82a, JJT80, Kea81, LO83, MLM82, Mik82b, Now82, PD82d, Par82c, Poh81b, Pow84, Raf83, RCC80, SP83, Ste83f, Tam82e, Tho84g, Tho84h, Tho84o, Tho84p, Wer81m, Wes80, Wes82, WMW83, vA82]. Methode [phAD80d, Col80b, Éme80b]. Méthodes [DG80b, Lel80, MR81b]. Methods [Ben84b, DAP82, DLR81, GPRS80, Gek84c, HT82a, Hub83a, Hub83b, KM80, ŁK83d, PR83, Rau80, Tho84e, AB83, All84, ABM81, All81a, Axe82a, Axe82b, Bjö84, Böh82, BGM^{+}80, Bus82, But80, Cas82b, Cat84, Cho84i, CN84b, DESS82, Del82b, DR82a, DMPG83, Eke83, Elw81, Eng82, Enr82, Eva83, GG82, Gek84b, Gek84e, Gek84d, Gek84f, Gek84g, Ger84, Goo84, GH82, Gra83c, HK81, HW82, HR82, Hel80b, Hem84, Her84a, Hor82a, JN82, JT82, Kan84b, Kas82a, Kea83, Ket82b, Kla81, Koc82, KN82, KI84, Kra82a, Kre81c, LM80, ŁK83c, ŁK83e, MM82a, MW84, Mit82b, Mor80c, Mor82b, NO83, Nør80j, Pot82, Rab84, Rei82b, Rup83, Saa83, SJ82, Sch82p, Skr84, SSTW82, Ste84a, ST82, Tay81, Tho84b, Tho84l, Van83c, Wer81u, Whi84]. Metivier [Len80b, Pra83, Éme80b]. Metric [Hop83s, Bay81b, BCJ^{+}81, CO81, Gar83a, Kal83, Pis83, Sch82b]. metrics [Bou81i, Rac83, Sen83a, Szu83]. métrique [dlP82]. Metrisabilité [Éme80c]. metrizable [CPD81, Jun84, Kar83]. Mexico [AL81, AL82, FG83, GP83a, Hen82a, Ric81c]. Meyer [Éme82, Wei81a]. MGOO [FW82a]. Michael [Est84, Sie80b]. Michigan [Dol83b]. microlocale [Ver83a]. microlocales [Las81c, Las81b]. Microscopic [PMC83]. middle [BB81a, Mic83]. midrank [Ruy80]. Midwest [Dol83b, LW81]. mild [Kot81]. Mills [Aso84, Bou82b, Bou82c, Fri82, Her82b, Tra82a, Tra82b]. Mills-equations [Fri82]. Milne [Ber81d]. Milnor [BY82]. Milutin [Hes83]. min [Dix81]. min- [Dix81]. minded [Kai83]. minima [Mar83a]. Minimal [CJ84, Cho84h, CdS81f, Hof82b, Kna83a, Rac83, Sha84b, AF84, Bro82d, CO81, Del81h, Hil84, Koi84, Mas84b, Nie84, Ohn84, dC81]. minimale [LJ80]. minimality [Hen80b]. Minimally [Wil80b]. minimax [Cas82b, Mil83b]. Minimisatin [HL81]. minimization [Pow84]. minimizing [Sbo83]. minimum [AK82, Ber84b, RV84]. Minkowski [Art80b]. Minneapolis [Eri84b, HN83, RW82]. Minnesota [Eri84b, HN83, RW82]. minorants [Bas84]. Minorations [Gal81]. Miquel [Sam80]. Miscellaneous [Kre84k, KDB^{+}84, McA83k, Ivr84a]. miscible [Dou82, PW84]. mistake [LM88]. Mittag [ESV83s]. Mittag-Leffler [ESV83s]. Mittel [Küh83a, Küh83b]. Mixed [Hem82b, AL80, FP83, HN82, Mek81, Tho84h]. mixte [Bos80]. Mixtes [Die83b, Die83g, Die83k, Dub82b, Jur83a]. Miyata [Wil81g]. mobility [Pic80c, RW84c]. Möbius [Ahl80, AF84, Iva83]. Möbius-invariant [AF84]. modal [FK81]. Model [Clo81b, FZ80c, JL80, JP81, Kol81a, PR84f, Sch84j, AD83a, Art84a, Bar83c, Bey83, Bru81b, DJK81, Gal80, Gis84, HM83a, MM82a, MV81a, Man84, Mil81b, Mun84, PM83, Par82d, Pre84i, Rob81b, Rot84f, Rot84i, Rot84k, ST82, vdD80, BMR81, PWW80]. Model- [Sch84j]. model-theory [Mun84]. Modèle [Tan83f, Tan83h, Tan83g, CR80a, Sou83]. Modèles [BMR81, Dem80f, Tan83i, Cha81b, Ric81a, Tan83c]. Modeling [HN83, O’M82]. Modelling [Cla83, Cre84, Rab84]. Models [KZ83a, Pil81a, Pil81b, Ste83d, BM81a, Bou80, Bru81c, Bru81e, Bru81g, Bru81o, Chu82a, For83b, GM83a, HP80, Hel83, KR83c, Kue81a, Luk80b, MP82, MK83, Mik82b, Mos81c, Par80, Sch81a, Wie83, Wil80b, vdD80, MR84b]. modern [RS81]. Modification [JW84]. Modifications [AT82c, BS83a, TW82a, AT82f, AT82g]. Modified [BF83j, Dou82]. modifiée [Duv83b]. modul [LL81]. modulaires [Des84]. Modular [Ben84b, Mus83d, Mus83f, Ste80b, Tra84, Bur83b, Day83, Fre83, Fre82b, Kos81f, Mac81a, Mit84, Mum83, Mus83g, Oda83b, Röd84]. modulared [Mus83b]. modulars [Mus83c]. modulated [DR80a]. module [BV83b, Car83a, Mon80i, Rin84d, Van83a, Wei81a]. Modules [Ben84c, Ben84d, Bry82, Cau80a, Cha83a, Dic80f, Hau83a, Lad83a, Lev80e, AS80, Bat81, Bjö81, Bor82a, Bou80, BL82, Car83c, Coh82, Cos82a, Cou81, Dad80, Dau82, DZH81, Fai82a, Fox80, GM82a, Gie82c, Gie82r, Gie82u, Gre80b, Gre80d, HPR80, Hau81, Hau83b, Hil81c, Hil83, Hum80f, Jan84, Jon82, Kle81, Knu82, Lad81, Lad83b, Lan80b, Lau83b, LP84a, Lev80i, Lev80g, Lev80m, Lev80c, Lev80n, Lev80r, Lev80q, Lev80t, Lev80u, Lev80w, Lin82b, Lyu84, Min81, Mon80j, NQ– D84, Oko81, Pog80, RC82, Sal80, Sta82a, Tay81, Ull81, WK80b, vOV81c]. Moduli [Cas82a, Cat83a, Tak80, vS81b, Cat83b, Gie83b, GŁ84b, Kat81, MM83, Nam80a, RV84, Tan81e, Tan81h]. Modulikörper [Sch84i]. Moduln [Sch82e]. Modulo [Wal83b, Wal83a]. modulus [Ghi80, KP80, MP81, dMPvS81, SC81s]. Moerbeke [Ver81a]. molecular [Elm84]. molecule [Die82b]. molecules [Wol82b]. Molien [Ale81b]. moment [AC82, Beh82, BL81, JT82, MR84a]. Momente [IvW82]. Moments [All84, GS82c, DHH^{+}84, Hey81, Kir82]. monômiales [BS83a]. monadic [Lam82, MS82a, ON81]. Monadicity [Fon84]. Mondragone [AFG84]. Monge [Ceg83, Ceg84, Kis84a, Lel83]. Monochromatic [GLP81, Rei84b]. monodromies [Moi81]. monodromy [Por83]. monoid [Ros81a, vL83a]. monoidal [Gre82d, Gre82e]. monoïde [Spe81b]. Monoïdes [Sak81, PP81, Spe81b]. monoids [Kna83c]. monomial [Buc80, BS83a]. monomiales [Eli84a]. monomials [Lyu84]. monomorphisms [BH84]. Monopole [Cor82]. Monorientable [Hil81g]. Monotone [KKM84a, Har81l, Har81m, Har81n, KKM84e, Mar81h, PP82a, Pre83h, Pre83k]. monotones [ZM84b]. monotonically [Kas82a]. Monotonicity [Col80a, Col83]. monotonously [CA82]. monsieur [Coe82]. Montecatini [Cer84a, Ghe83, Sal84b]. Montessus [GMS84]. Moore [Hus81a, PdS80, Tan83h]. Morass [Dev81]. Morass-like [Dev81]. morasses [Don81]. Morava [HPR84]. Mordell [And81c, Ray83]. Mori [Dem81]. morphism [Dem82b, Rus84b]. morphisme [Ang81i, Băn80, Cam80a, Dem82b, Erm80a]. Morphisms [Gie82p, Kos81e, Sot82h, Joh82, MT82, Sot82i, dS82]. Morrey [JTW83]. Morse [AG80, Bas81, Sha84b]. Moschovakis [Fey84c]. Moscow [KZ83a]. Moser [DM81d]. most [Bui84]. Mostow [Mun80]. Motifs [Del81d, Del81e]. motion [Bas80, Fuk84, Gur80, Ken81, Ken84a, KSS80, McG84, SC81m, TAP84, Yor84]. motivation [DRR81g, Mar81c, PR84e]. Motives [DMOyS81c]. mountain [Rab82]. mouvement [Gal82b, JY81, Sch80o]. mouvements [Fer81c]. moving [Cha83s]. moyennables [Bou84b]. moyennes [Des84, Gay80]. moyennes- [Gay80]. moyens [Bro84]. Muckenhoupt [Bru82c]. Multi [Hac82a, Hac82b, Mas81, Mit82b, Düm84, Hac82c, Koh83, Mar81e, SJ82, SJ84, Tay84, vA82]. Multi-applications [Mas81]. multi-graphs [Tay84]. Multi-grid [Hac82a, Hac82b, Mit82b, Hac82c, SJ82]. multi-parameter [SJ84]. multi-point [Koh83]. multi-trajectories [Mar81e]. multiaffine [Sem82b]. multiderivative [Gek84b, Gek84f]. Multidimensional [LeP81, Fer84, Fuk82b, Jer80]. multidimensionnel [BD80a, Har80a]. multifacility [CC82a]. multifrontal [Duf83]. Multifunctions [Spi84, Art84b, Iof80, Mar84b, Thi84, Sal84a]. multigraphic [Bil81, EH81a]. multigraphs [BJ83c, SS81b]. Multigrid [FW82a, Hem84, ST82, AS82b, Axe82b, Bra82a, Bra82c, DH82, Ket82b, MLM82, Now82, SSTW82, Wes82, HT82a]. multigroupe [Mas82b]. Multilinéaires [Ang81d, Ang81g]. multilinear [CM80b]. multimeasures [GT80a]. multinomial [Pan83a]. Multiparameter [Bro82e, Cha80b, Sle80, Rüs80]. Multiple [Bol80, Ecc80, Maj81c, Maj81b, Amb83, BLP83d, CG82, HL81, HW81a, Kle82b, Wes80]. multiple-point [Kle82b]. multiplers [TNV^{+}84]. multiples [Lio83]. Multiplication [Gro80b, Ceg81, Cha81b, Cla83, Gro80i, Shi81b]. Multiplicative [Chi84a, Has83i, Nar84d, Erd81a, Hum80h]. Multiplicator [BT82c, BT82g]. multiplicité [Las81c, Las81b, Las81e]. Multiplicités [Las81d, Tei82]. multiplicities [CKK83, Las81d]. multiplicity [Mic81b, Ple83a]. multiplier [Joh83b]. Multipliers [Bad83, BPV83, Cór80b, Zaf82, MS80b]. Multiplikation [Sch84c, Sch84b]. Multiplikative [Dra80f]. Multipoint [Gel81a, Wal83d]. Multipoints [Cha84a, vI84]. multisets [Bil82a]. Multistep [Gek84f, Enr82, Gek84b, Gek84e, Gek84d, JN82]. multivalued [AF81, Hia84]. multivariable [Kri81]. Multivariate [Cuy84c, Deh81, LL84, Cuy84b, KKM84d, Pan83a, Sab84, WMW83, dB82]. Mumford [GvdP80b, GvdP80f, GvdP80i, GvdP80j, Nam80i, Tan81c]. muni [Mic81a]. Mustov [Kör80b]. mutuel [Fey80c]. my [Ram82a]. Mysore [All82b].

N [Bax83, Mer80b]. N. [BLP83e, DK80a, Kis84b]. N.B [Ara84]. N.S.F [KKS82]. N.S.F.-C.B.M.S [KKS82]. Nachbin [Fuc81b]. Nachdiplomvorlesung [HP84h]. Nagumo [Rot84e]. Nakano [DS80]. Nash [BF84, BKS82, BE82, CR82, Efr82, Mil84b]. Nash-Williams [Mil84b]. Natural [Fri81, Pop83a, Isl81, Ket82a, Zel81f]. naturelles [Com82f]. naturels [Ceg81, Cha81b]. Navier [Rau80, Ami80, Bas80, BGM^{+}80, DMPG83, Fas80, Foi80, GLCS80, Hey80a, Hey80b, Kni80, KY80, KB80, Mar84a, Mas80]. n’ayant [Del81f]. Nazarova [Rin80a]. Near [Arn82j, Pre82, Arn82a, Aul84a, Bro81b, Dui84, ESV83g, Hen81g, Kow81b]. near-isomorphism [Arn82a]. nearaffine [MT80b]. nearness [Ben82c, Gre82d]. NEC [Buj83]. necessarily [Had82, Kol84]. necessary [All82d, Kno80]. necking [ZC84]. need [Fer81d]. needed [Kis81b]. Negative [Hop83q, Tho84i, Bis81a, Ken81, Ken84a, Lic84b, Pre83i]. negatives [Ber84a]. Neighborhood [Hen81j, Hen81i, Hen81k]. neighbourhood [IV80, vdHM84]. neighbourhoods [GM80a]. neighbours [Bie80]. Neisendorfer [Hus81a]. Neofields [Hsu80, Kee83]. n’est [Vig80f]. Netherlands [Jag84, LST81, Mar80b, MdJ82, dBvR81]. Nets [Abh82h, Abh82g, Abh82a, Abh82b, Abh82j, Abh82m, Abh82o, Don82g, Mun84]. Nettings [Vek84]. network [Fou81a]. networks [Dre83, Lum80b, Pol84]. neuer [Dit80]. Neumann [Bro83, Dru83, GŁ84b, Hen84b, Jaj84a, Jaj84b, Krá80c, Krá80e, Tal83]. neutral [CMA83, GBC82, IV80]. Neutralisés [Tig81]. Nevanlinna [Cha84d, phAD80c, Leh83b, Osg84, Top83]. Nevanlinnas [phAD80b]. Nevanlinnasche [phAD80b]. Nevanlinnaschen [phAD80b]. Newcastle [RSW80]. Newton [Böh82, Bus82, JJT80, Mer80b, Pot82]. Newton-like [Bus82, Pot82]. Newton-method [JJT80]. Newtonian [ABHK84, Car80c, Gau83]. Newtoniennes [RP84]. next [Coh81]. nice [Ver84k]. nicht [Bon80, Sch83t]. Nicht-Unitarisierbarkeit [Sch83t]. nichtlineare [App83]. nichtlinearen [Nic82]. nichtlinearer [Wei82a]. Nico [Sul81a]. Nielsen [Fos80, Zie81c, Zie81f, Zie81e]. Nijenhuis [Riz84]. Nikaido [Par83b]. Nikodým [AS83, Bou83h, FM83, JvdS84j, Mus83a, Tra84]. nilmanifolds [Cor84d]. Nilpotence [VL83b]. Nilpotent [Slo80c, AP80, BMP80, Cor84c, Cor84e, Haz81b, Hel84a, Hil82c, HP84b, HP84f, HP84g, HRW84, Mon80h, Ric82, Rot84l, TH84]. nilpotente [HP84b, HP84c, HP84h, HP84h]. nilpotenter [HP84a, HP84f, HP84g]. nilpotentes [Lev80a]. nilpotents [Gal82b]. Ninth [BOWS82]. niveau [Bou84f]. NK [Nav84a, Wei81a]. NK-manifolds [Nav84a]. No [HM80, HM82b, HM84a, FY81, Han81, Mon80h, Mon80g, She82i]. nodal [Ful83]. nodaux [Wat84a]. Noether [Cam80d, EH83, Eng83a, Ghi83, Mal81b]. Noetherian [Bor82c, Gol81a, Min81, She81d, Sot82c]. noetheriennes [Yam80]. Nohel [Oli80]. noise [SM83a, SH83]. Nombre [dM80a, Kis82, Lel82a, Oes80]. nombres [Ten84]. Nomenclature [Gru82p]. Nomenklatur [Gru82p]. noms [Ano80, Ano81b, Ano81c]. Nomura [HJ82a]. Non [Ber83c, Bro80, CV81, Dac82e, Deh80, DP82b, Dra84, ELZ83, FK81, Fle82c, Jaj84b, Kow80k, LG81c, Maj81d, Mal80a, Miy81a, Miy81b, Miy81c, Miy81d, Raw80, SC81j, Sig83f, Tak80, Tal80, War82g, Wer81t, dO80a, vOV81h, ABHK84, Ale81a, All82d, And84b, Art84a, AW82, Băn83b, Bar81a, BP83, BF83a, Ben84e, Ber84c, Ber80c, BC82b, Boi80, Bor84c, BLP83d, BH82c, Brü82b, Cer81, Cha83a, Cho83a, CMS80, CR80a, CPD81, Com82a, Coo84, Der81, DP83b, Dur80, Eke83, Elb82, Emc81, Éme80a, Fer83a, Fle82a, FJ81, Fle82b, GM84b, Gli84, Gol83a, GL82c, GT83, Gre82a, Gri82, Gru84, Gui80b, GM80j, HW82, HG82b, Han84, Har81m, HN80a, Hei84, HLSW80, Hig82b, Isl81, KKM84b, Kun82, LL81, Lic84b, Liu84, LG81a, LG81b, LG81e, Mal82b]. non [Mañ81b, Mar83a, Mey81g, Min83a, Miy83a, Mon82, Mor80c, Mor82b, Par82b, Pas83, Pav84, Ped84b, Rao80, Roq82, Ruí83, Sau80, SC81o, Sch82o, Ser80, She82i, Ste83c, Tho84n, VVV81, Ver80b, Ver81b, Ver84l, Vog82, Wal82b, Yan80, ZM84b, dF84, vOV81i, AD83b, CV81, CV83, Jos84a]. non-abelian [And84b, Băn83b, BH82c, Pas83]. non-ample [Gol83a]. non-anticipating [All82d]. Non-Archimedean [Ber83c]. non-autonomous [Coo84]. Non-central [Maj81d]. non-classical [Ver84l]. Non-Commutative [Dra84, Fle82c, vOV81h, Jaj84b, AW82, Mon82, VVV81, Ver80b, Ver81b, vOV81i]. non-compact [DP83b]. Non-complete [Miy81a, Miy81b, Miy81c, Miy81d]. non-connected [GL82c]. non-conservative [GM80j]. non-constant [Gru84]. non-convex [Eke83]. non-definite [Min83a]. non-dégénérés [Com82a]. non-differentiable [Fle82b]. Non-dilation-analytic [Sig83f]. non-equilibrium [Emc81]. Non-Existence [Tal80, SC81j, Art84a, Ben84e]. non-homogeneous [GM84b, HG82b]. non-involutive [BLP83d]. non-isolated [HN80a]. non-isomorphic [Isl81]. non-Linéaire [Ruí83]. Non-Linear [Dac82e, DP82b, Bro80, Wer81t, Ber84c, Bor84c, Elb82, Gli84, KKM84b, Kun82, Mañ81b, Miy83a, Par82b, Pav84, Sch82o, AD83b]. non-linearly [HW82]. Non-maximal [War82g]. non-metrizable [CPD81]. Non-modal [FK81]. non-negative [Lic84b]. Non-parallel [Kow80k, Der81]. Non-positive [Raw80]. non-rational [BF83a]. non-Riemannian [FJ81]. Non-self-adjoint [ELZ83, Fle82a, Gre82a, Mor80c, Mor82b]. non-selfadjoint [Ped84b]. non-separable [Han84]. Non-singular [dO80a]. non-smooth [Tho84n]. non-squarefree [LL81]. non-standard [ABHK84, LG81b, LG81e]. non-strictly [GT83]. non-strictly-convex [Mar83a]. non-summable [Ste83c]. non-symétrique [Hig82b]. Non-transversal [Tak80]. non-trivial [SC81o, She82i]. non-unique [Brü82b]. non-uniqueness [BP83]. Non-variational [Jos84a]. nonadiabatic [Wol82b]. nonarchimedean [BN81]. nonautonomous [Hen81h]. Noncocommutative [Taf82]. Noncommutative [Cec84, Mun84, And81b, Sta82a, vOV81l]. noncompact [Por83]. nonconservative [CM80a]. noncontinuable [Sic84]. noncontractions [MV81a]. Nonconvex [Pow84, Mar83a]. nonexpansive [Kir81, vDS83]. nonfinitely [MS83b, Sie80a]. nonhomogeneous [MN82]. nonintegrability [Chu82d]. nonisotropic [Soa83]. Nonlinear [AGP81, AMT82, BLS80, Bax81, Bax83, Ces84, For83b, GBC82, Har81i, Imm84d, Lor82, Seg83, Amb83, Ame82, Ant80, App83, BC82a, BH81b, Ber82a, BT82i, BN80, Bus84a, CG82, Cha83s, Cho80b, CCN82, Cuy84e, DR82a, ES83, Eng84, Erb83, Fer84, FR83, FO82, Gek84d, Har81h, Har81k, Har81l, Har81m, Haz81a, HM83b, Hen81c, Her80a, Hor82a, Jef80, Jef81, Kan84b, Kaw82, Kea83, Kre81b, LR83c, MNP81, MW80, MLM82, MB82a, Mil80b, Mil82, Mit82a, Mit82c, Mit83, Neu84b, Noc82, Nus81, Ort80, Pan82, Pan83b, PS81, PP82b, Rei84a, Rob83b, Sib80, Skr84, SSTW82, SJ84, Tho84j, Vin84a, Wer81u, Ytr84, dM83, MM82b, Vin84b]. nonlinearities [CG82]. nonlinearity [Har81j]. nonlinearly [Men83]. nonlocal [Men82]. Nonmonotonic [ZM84b]. nonnegative [GM84b, Mil82]. nonnormality [Bar82c]. Nonoscillation [FK83a, Elb82, Erb83]. nonparametric [Ale83a]. nonself [Gir83b]. nonself-adjoint [Gir83b]. nonsingular [KKMW81]. nonsmooth [Jof83, Las84]. Nonstandard [HM83c, Hur83, LG81d, Kaw81, Lau83a, LS82b, Mur84, Per83b, RS83c]. nonstationary [Mas80]. nonsymmetric [Ruh83]. nonzero [Che84b]. Noordwijkerhout [Jag84]. Nordic [LR83a]. Nörlund [Mad80g, Oku84b, Oku84c, Oku84e, Oku84d, Oku84h]. Norm [CR83, Sch80i, Tho84f, Bak84b, Gri80, KN82, Loy83, Mun80, ND83, Pau83, Pic80a, Saw82, Tho84i, dF82b]. Norm-dependent [CR83]. normable [AP81b]. Normal [Sag80, Saz81c, Bro81a, DG84b, Eis81, Gol83a, HO82, Krá80h, Kre84f, Mas82a, Sie81a, Vas81]. Normale [Ves83, KS83b]. normales [Tho81a]. Normalisation [Mar81i]. normalisé [Har80a]. Normalité [BD80a]. normality [Cos82a, DHH^{+}84]. normalized [Bak82a, RV84]. Normalizer [Bow82]. Normalizing [BDR80, BD80b]. norme [RP84]. normed [Fru83]. Normen [Stu80]. normes [Bro81d, Dub84, Oja82]. norming [HK81]. norms [BN81, CS82d, Wil83b, vDS83]. North [Car82]. Northwestern [FS81e, NR80]. Norway [JTW82]. not-autonomous [Cap83]. Notation [Abh82i, Arn82k, Gro80g, JvdS84h, Kle84i, ŁK83b, Mad80f, Rin84d, Rin84f, Sto80b, Str84d, Ver84g, Cha83l, Bis81l]. Notational [ESV83k, Gie82q]. Notationen [Sch83u]. Notations [BF80g, Bis81l, DMOyS81d, Ogu81e, Spa82d, BBCH81j, Rot84a]. Note [CD81, Éme83, GN82, Mey82d, Sto80g, All82d, Ben82c, BJ83c, CS83a, Clo81b, DK82a, Gil81c, HJ82c, Hur82a, Kec81b, Lop84, Pit81, Wil82a, Wol82a]. notebook [And81c, BW81a]. Notes [AL82, Com82g, Gal89, Kan82, LM88, She82g, She82h, She82o, Acz81b]. Notion [Gue80, Háj84, Isl81, Mar81c]. notions [DS80, Dem82a, Pre84i]. nouveal [CMS83]. nouveaux [Pis80]. nouvelles [Mey80c]. Nov [BS83e]. November [BH81a, BKK82, CZ83a, Dol83b, FKSW81, HT82a]. Novikov [FH84]. noyau [BV81, Bou84a, Col80b, Hig82a, Hig82b]. noyau-fonction [Hig82b]. Noyaux [Éli84b, Del84b, Itô84e, Itô84c, Itô84b, Itô84a, Saa84b, Fey82b]. NSF [CW82, KM80]. NSF-CBMS [CW82]. nuclear [Bjo82, Kis81b, Rot84i, SC81a]. nuclearity [Joh83a]. nuclei [Spa83a]. nucleus [CL84a]. nuisance [HI83]. nul [Kas81]. null [But82, Grz82]. null-completion [But82]. Nulle [DI81b, Vid80]. nullsets [Top80]. Nullstellensatz [CD81, CT82]. Nullstellensätze [McK80]. Number [Kno81, PV83a, AS84a, AD81, Bak84a, Bas80, BB81a, BY82, Ber84b, Beu81, CL84d, Dem83b, Erd81a, Erd82, FR80, Gne83a, Hag81, How82, Isl81, JS82a, KR83b, Mae81, MM82a, Mäk80d, Man80, Mic83, PV83b, Pin84, Sar84, Sch84a, SW81, Tau81, WTL84, vdD80, All82b, CCCN84, Jag84]. number-theoretic [SW81]. Numbers [Sch80b, Sch80c, Sch80l, BB82b, BKQ83, CP83a, Daf81, Edm82, Fin82, Hei84, Hia84, HT81, Lün82, MS82d, Ple83i, Ram82a, Ram82b, Sou81a, SM80]. Numerical [Bey83, Cuy81, Fas80, Geo81b, GT84, Hem83, KSS80, LM80, Mäk80j, MB82a, NSE82, NO83, SJ84, Tur82, Wol82b, Bak82b, Bas80, Bru84, DESS82, ESV83e, Enr82, Hem82a, Hul82, KY80, Lan82a, McL82, Mit81, PW84, Ste80b, Wac80, Wat84b, Web81, Wil84, dM83, AGP81, Gri84, Hen82a, Hin82, PR83, Wat80, Wat82]. numerically [Geo81c, Mor82a]. Numérique [Ruí83]. numerischer [Wei82a].

O. [Gal89]. Oaxtepec [GP83a]. Oberwolfach [AMT82, BJ83a, Den82a, Den82b, DLR81, EdJ82, GW81, Hey82, Hey84, HJW83, JPW81, JW83, Köl80, KMS82, KMS84, Rog81d]. objects [Ban83a, Ket83a, Shi81a]. Obrechkoff [DK80a]. observables [dF84]. observation [Par82b, Wil81d]. observations [Eis83, Sou83]. observed [Dup80]. obstacles [Bem80]. Obstruction [GWL81, Cha83h, Cha83n, DI82, Fer81e, HTW84]. obstructions [Ham82]. obtenues [Bar82a]. Occurrence [Pre83g]. octagon [War82d, War82h]. October [BLS80, Nav84b, RS83a, dBvR81, She82h]. Octobre [Nor80a]. Odd [Mil81a, CG82, FMW82, Gra84b, Hag81, Hus81a, Kre84m, Ple83h, Pra80]. odd-dimensional [Kre84m]. ODE [Tho82]. ODE-codes [Tho82]. ODEs [Sha82, Sha84a]. oeuvre [Anc84a, Caz83]. off [Per84]. off-diagonal [Per84]. offshore [AB83]. Ohio [CW82]. Old [Sta82b]. Olkin [AV83c, OR83]. Olomorphic [Tra82b]. One [BF83k, Le 84a, MS80a, Alb80, BO82, BF82, BGMY80, Bro81b, Bün81, Chu82c, CC82c, Dic80b, Dyd81, Ecc80, Gol83b, GQ82, Guc80, Isc82, Jan82, Kom80, LP84a, Mic81b, Ore82, Pap84b, Sie81b, Šig83l, VVV81, Wil81f, dMPvS81]. One-dimensional [BF83k, Le 84a, Chu82c, GQ82, Isc82, Ore82, Pap84b]. one-parameter [Bün81, Kom80]. one-point [Dyd81]. one-sample [Alb80]. One-to-one [MS80a]. ones [EHP83]. only [Lev80d]. Onsager [TW81]. ont [Dur80]. onto [AJ83, Aup83]. Open [CGO83, Kru84a, Boc82, CKMN83, Cha81a, Düm84, GP83b, GM80c, HM83b, JK83, SC81f, Sto83a]. opérant [Avr82]. Opérateur [Bou84d, Kis84a, Lel83, Lie80, dM80a, dV81]. Opérateurs [FLM81, Arq84, Bet83, CMS83, Las81e, Las81f, Seb80a]. operation [Gam80]. Operational [Kan83h, Zaf82]. Operationen [Lja81]. Operations [Kow80l, Sou82, Wal82a, Del84a]. Operator [KS84a, Mad80g, SSM^{+}84, ABM81, And84a, Bed82, Ceg83, Ceg84, Cho84e, CCN82, Cuy84a, Cuy84e, Gal82a, GL84a, HK81, Jou83h, Kis81b, Krá80c, Krá80e, Lac84b, Lut82, Mil80b, Mil82, MS80b, Par82c, Pas80, Ruc82, Van83a]. Operators [AD83b, BR80a, BR80b, BLP83d, BLP83f, Cuy84d, Don82e, EL83, HP84j, Ivr84c, Jou83f, KKM84a, Kal80, Kan83o, Krá80f, Mad80d, Nac82, Sto80f, AAK^{+}84, Ahl81a, AHL81b, And81a, And82, And84b, BG84a, Bia81, BB83a, Bor84c, Bré81a, Bro82d, Bro83, Bru81d, Bru81h, Bru81j, Bum84g, CMZ80, CV84a, CRW^{+}80, Col80c, Cor84e, CM82b, Don82f, Don82d, Don82g, Eva81, ELZ83, Eve80a, Fle81, Fle82a, Gie82e, Hab80b, HG82a, Har81n, HJ82a, Haz81a, Imm84b, Ivr84b, JP84b, Jou83c, Jou83e, Jou83d, Jou83j, Jun83, KKM84e, Kan83a, Kan83l, Kau83, KS81b, Kom80, Kré81a, Lan80a, Lan82c, LMW80, Lum80b, MZ83, Mar81g, MW80, Mos84b, Nic83, Nus81, PS84c, Sag80, Saw82, SH84, SY84, Sto80a, Sto80e, Tho84k, Var84, Dun82]. Opératurs [Dub82a]. opposite [Daw82]. Optimal [Bar83b, DKT84, Gal82a, Not82b, Not82c, Sch82l, Ces83, DG80b, Kno80, Lio83, PS84c, Rem82]. Optimalité [Die83e]. optimality [Ead81, EMS82]. Optimization [Ame82, CZ83a, Sal84a, Vin84b, Art84b, Ces84, Dră83a, Eke83, Fle82b, Gay84, GL82c, GT84, HHB84, Jof83, KI84, LMGC82, May80, Spi84]. Optional [Wal81a]. optionnelle [Str80a]. oracle [She82i]. oracle-c.c [She82i]. orbit [Aso84, CMPY83, HT82b, Tan81g]. orbital [SS83]. Orbites [DM81d]. Orbits [Cam82a, Cam82b, AMPY83, AF83, Cur80a, Hap81, JB81, Kir82, Kon82b, Mis84b]. Order [BBCH81f, EF81g, Ise84b, Ise84a, Jos82, Mat80, Nic83, Abh82n, Abh82m, Atk82, Bax83, BE83b, Bor82a, Brä82b, Bro81e, Bro82c, CC82a, CR84a, CQ83, CPD81, DM80c, EG81, Eas82a, Eas83, ESV83s, Enr82, Erb83, Eri84a, EF81d, EF81c, EW83, Fle82b, FK83b, Gal82a, Gek84b, Gek84d, Gek84f, Gek84g, Gly83, GS82b, Gri80, GM80f, GM80g, HU83, HN81b, HO80, Ise81, Ivr84b, JP84b, Kar81, Kno80, Kwo83, KZ83b, LTS82, MW84, Nat84, Not82d, Not82f, O’M82, ON81, Rav84, Sot82l, Ver84a, Zel81f, dB84a, dCR83]. Order-free [Jos82]. orderable [Huš82b]. Ordered [Abh82j, Abh82k, Ama81, Kha82d, Bal81b, Bün81, Con81b, Ern81b, Gla81, Han80, Hol83a, JM83, Kha82e, Len84, Rhe81, Rom83, Sch84j, Wec81, Wil81a]. ordering [Jon81]. Orderings [Pre84b, HP80, Ler81, MR81a, Pre84j]. Orders [Arn82l, And81b, Arn82i, Art82b, Art82c, AS81b, Brz81, BR81d, Cha80a, Jou80, MR80, Ple81, Ple83c, Ple83e, Ple83i, Rog80, War82g]. ordinaires [Bet83, DI81a, Del81f]. Ordinal [Buc81b, Mos81c, Acz81a, How81a]. ordinals [Bae84, KP84b, Sie81a]. Ordinary [CdS81g, Min83h, Die82b, Elb82, Eng82, Gea82, GG82, GP83d, Kau83, LM80, LO83, Not82e, OS80, SS80a, Sch82l, Ste82b, Wol82b, Eve80b, ES81, ES82, EL83]. ordonnéas [Rol82]. ordre [Sch82n]. Ordres [MR80]. organisé [LS82a]. Orientable [Ioc84e]. Orientables [Ioc84e]. Orientation [TAP84, Wis84, Kre84f, Kre84d, vdLT81]. Orienting [Eri84b]. origin [JvdS84j, Pee84]. original [Maj81f]. origine [Cha83t]. origins [PMC83]. Orleans [KKS82]. Orlicz [Blo84b, Gar82d, GT80b, Gos84, IJ84, Ław82, Mus83e, Mus83f]. Ornstein [HS81]. Orsay [MP84b]. Orthogonal [Oku84g, AD81, Dra81, Dra83e, GH83, GHH83, Mag84, Oku84a, Oku84c, Oku84f, Seb80b, Dra83f]. Orthogonalgeometrie [Ben80c]. orthogonally [Har80c, Nie84]. Orthogonaux [Dra83f, Dra83e, Dra83i]. orthonormal [Gro81b]. Oscillation [AS81a, Elb82, Erb83, Kre81b, Jon81, Kwo83, KZ83b, Leu80]. oscillations [Lev80b]. oscillators [Vor82]. oscillatory [Gea82, GG82]. Oseen [MP83a]. Osterwalder [Kle84b]. Other [BT82g, Cam80e, Kos81e, Cuy81, Erd81b, Mar83d, Ul’84]. Ottawa [Ban82, DG80d, DG80c]. our [ESV83e, ESV83t]. outer [Mas82a, Mon80d, Mon80i]. outerplanar [Cho81]. outgoing [BL81]. Outline [ESV83l]. output [Ivn83, PS84c, Ryd83]. Ovaloids [Hop83c, Hop83l]. ovals [Gly83, LTS82, LZ82]. Overrelaxation [Kas82a]. overview [Sec82].

P [ARMS82, Gie82a, Gro80h]. P. [Bea80, DS80, Hus81a, Jøn80, Mor81b, Ver81a]. P.D.E. [BLP83a]. P.Dolbeault [LDS83]. P.I. [Ver80b]. P.Lelong [LDS83]. Pade [Beu81, Bos84, Bre81c, BvI84, Bul84a, CC84b, CC84a, Cuy81, Cuy84a, Cuy84b, Cuy84c, Cuy84d, Dor81, Dra83b, Dra84, Gel81a, GL81, Ise84a, KS81a, Kla81, LV84, Mag82, Nut82, Ort80, Sab84, Tre84, Wal83d, WB84, Wil84, Wuy81, dBvR81]. Padé-approximant [Cuy81]. Padé-approximants [Cuy81]. Pade-approximations [Beu81]. Padé-type [BvI84, Sab84]. padéons [LM84a]. Paderborn [Rau80]. Painlevé [CM82a]. pair [Ach81, CdS81f, Das81]. pairing [Hil81e]. Pairs [CdS81g, Ban81c, Cho84h, HJ82a, tR84]. Palaiseau [DPT80]. Paley [Blo84b, Jou83d, Mey81d]. Panelled [Mas83a]. paper [Gra83d, Ram82a, SZ81a, SM83b]. papers [JW83, vH83a]. parabol [Dub82b]. Parabolic [Hen81e, Tho84e, Tho84k, Bar83b, Bar80e, BV83j, Cha83s, Col80d, Eng83b, Hal81, Hen82b, Hen81c, Hes84, Kir84, Naj84, Nam80b, Sto83b, Wah82, Wol80, vdH82]. parabolich [Sch83n]. parabolique [Pie80]. paraboliques [Dub84]. parabolischen [phAD80b]. paracompact [Gie82g]. Paracompactness [Fro84]. Paradoxales [Bir83a, Bir83b]. parallel [Der81, FS84, Gli84, Kit84, Kow80k, Pic80b, Pin81].

Parallelepipeds [Sch80h]. parallelism [Kol81b]. parallelisms [Her80c]. Parameter [Bar80e, AHE81a, AHE81b, AF81, BB81e, Bün81, Che81a, CS81a, Doz81b, HI83, Hem84, IS81a, Kom80, Men83, Mer84, Mus83c, SJ84, Tan81d]. parametered [HW81a]. Parameterization [Bas80]. Parameterizations [Rei82a]. parameterized [Spi84]. parameters [Col81a, Eri84a, Fit84, KMM83b, KK83]. paramètres [Del81a]. parametric [CKK83, Deh80]. Parametrical [ŁK83d, ŁK83c]. Paramétrices [Las81c, Las81b, Sig83c]. Paramétrique [Rao80]. parametrix [Mal80c]. Parametrizations [Cam80f]. parametrized [FR83, Kea83]. Parfait [Bre81b]. Paris [BLS80, BMR81, KMS81, Mal80b, Mal81a, Mal82c, Mal83, Acz81b, HM80, HM82b, HM84a, MR81b, vdT81]. Park [HLR83, HKLR83, HJLR84]. Parreau [Has83b, Has83d, Has83j, Mél80]. Part [BOR^{+}84b, CBJS83a, CBJS83b, Den82a, Den82b, MR84b, Beg83, Eng84, Lau80b, Van83a, FZ80a, FZ80b, LeP81, Rob83b]. Partial [AD83b, BR80a, BR80b, Eve80b, ES81, ES82, GM84a, Hel84a, Pic80b, Str84f, And82, BS83b, Bia81, Bré81a, Bro81c, Col80d, Gin83, Háj84, HR82, Joh83b, Kos83b, Kun82, MR81a, Not82d, Not82f, Par82b, Ros83a, Rot84l, Sag80, Str83a, Zel81f]. Partially [Ama81, Bün81, Ern81b, GT84, Han80, Rom83, Wec81]. Particle [Rav84, Seg82b, NT82, Per82a]. Particles [Bon82a, Bai82a, Blo80, Duv80, Gal80]. particular [Doz81a]. particule [GL82a]. Particulières [Dra83d]. partie [A’C80]. partielle [phAD80a, Lja81]. partiellen [Gru82a, Gru82l]. partielles [Mey81g, Ser80]. partir [CR80a]. Partition [Pil81c, Clo80, DPV82, KKMW81]. Partitions [BFG82, KP84b, Van82, CS82c, Ers80, GLP81, Mag83]. parts [Bur80]. pass [Rab82]. passage [Sch80q]. past [Bem80, Kni83]. Path [Lia84, Tar82, All82d, Bri81b, Hor82b]. path-length [Bri81b]. path-wise [All82d]. Pathological [Kar83, Kon82a]. paths [Als81, Fra83b, lPBF80, Wal81a]. pathwise [Gal83a, Mit83, Per82b]. Pattern [KJTB80, Her82c]. patterns [Fie83b, Kau80, Tam82a]. Patterson [Del80a]. Paul [Mal80b, Mal81a, Mal82c, Mal83, Bou84b]. Paule [Mal80b, Mal81a, Mal82c, Mal83]. Paulo [dFH82]. PDE [Duf82b]. Peano [Ber81c, Clo80, Clo81a, HP80, Pil81b, Pil81c, Ric81a, Sch81a, vdT81]. Peixoto [CKMN83]. Pellaumail [Éme80b, Len80b, Pra83]. Pellian [Coh81]. penalization [BN82]. penalty [May80]. Pencils [KR83a, Dem83b, GP83c, Kåg83]. pendulum [MW84]. Peñíscola [Nav84b]. Penrose [Eas82b, PdS80]. Percolation [HD83, HN83, KWW83, VLA83, HM83a, Red83a, Red83b, Sah83, Whi83, Wie83]. Perfect [Cha81c, MS82d, AK82, Hag81, Sta82b, Tam82e, HMP82]. perforated [MSP83]. Periodic [Bec80, BCF82, BGMY80, CMA83, Cho80b, Cus83, DI83, Fri83, Ver80a, Zeh83, Adi80a, Adi80b, AMPY83, Bar83c, BB82a, Bro81b, CMPY83, CZ83b, Dui84, HPR80, Har81a, Har81d, Hen81i, Hes84, Hol81, Jac82b, JM84, JB81, Kos82, MB82a, Sle80]. periodic-parabolic [Hes84].

Periodically [Lev80b, PS84b]. périodicité [Tro83]. périodiques [DM81d, Gay80]. periods [Sie81b]. permanent [Gol80]. permanental [BJ83c]. permeabilities [MSP83]. permutable [VL83b]. Permutation [Die83h, Hal80a, Küh80]. permutations [Die83a]. Perron [Gre84, Wer81m]. perspective [Cor81a]. Perturbation [Bou82a, gS83a, Axe80, Con80a, DH80, FN82, Gre82a, Hem82a, Hem83, Kam82, Kaw82, LG81c, LS82b, vH82]. Perturbations [BB83b, EdJ82, BH80b, Eng84, FW82b, GM83a, Hab80a, Hol81, HO80, HM82d, MW80, Mil82, Naj84, O’M82, Pri82, Tri84, Zvy84]. perturbed [BV83j, Dră83a, DH83, FO82, Hil82a, How82, Mik82a, Moe82, Nør80j]. Perturbée [CR80b]. Peschl [BR80c]. Pesin [FHY83]. Peterson [Lin81]. petites [Pri82]. Petrov [Mor80c]. Pettis [AW82, Gul83, Sen84]. Pezzo [Dem80b, Dem80c, Dem80d, Dem80e, Dem80f]. Pfaff [GR83b, KS83b]. Pfaffian [Gér80]. Ph. [Car81b]. Phase [Gar82a, Hor80, LM80, NT82, Rey80]. phase-plane [LM80]. Phases [Com82f, Com82c]. phenomena [Bus84b, Maj84g, Sna82]. phenomenological [Col84]. Philadelphia [Kno81]. philosophy [Goo81a]. phylogenetic [FR80, FR81, Hen80b]. physical [Ano81a, Bai82b, BMKQ83, Hen81c]. physico [Ort80]. physico-chemical [Ort80]. Physics [BS83e, DAP82, Eel82, GPRS80, HN83, KM80, Bai82a, BCP83, Hil82a, JS82a, Vla82, Zuc84]. physiques [Ano81a]. Picard [Bre83d, Sco81, Ste82a]. Pick [Ise84b]. Pick-type [Ise84b]. picking [BJ84]. Pictures [CS82c]. Piecewise [LR83c, Pre83h, GH82, Pre83k, Tod82, Wah82]. piecewise-linear [Tod82]. Pierre [Coe82, LeL82b]. Pietsch [SC81j, SC81n, SC81o, SC81p, SC81m]. pinched [BM81c]. Pisier [Kai83, SC81h]. Pite [KR83a]. places [Bec82b, Pre84g]. plan [Duv83a]. Planar [Abr82, ZVC80f, ZVC80h, Bou83k, EFG82, Lit82, Mah84, Sul81a, ZVC80b, ZVC80d]. Plancherel [Duf84, Her81, Her83a]. Plane [DRR81a, Kle82b, ŁK83d, Ami80, BH81b, Cam80c, Cam80e, Cas82a, CS82c, Del81f, Fra83b, Has83c, Iit83, Jou80, LTS82, LM80, MP84a, Mon84, Nam80g, O’F83, Par83g, RS83b, Rey80]. planes [Art80b, Bou83k, Bre82a, Bro81e, DR81, FMW82, GM82b, GM82c, Gly83, HS82, Lor81b, Pic80b, Tei82]. planets [ABHK84]. Plans [KO81a, Web84a]. Plans-sur-Bex [KO81a]. Plate [Ber81d]. plateau [Böh81]. Platitude [Szp80, Sal80]. Platonic [Slo83]. plats [Sal80]. Plattsburgh [Fle82c]. playful [BM81a, Mos81c]. pleinement [LC80]. PLS [WMW83]. plumbed [Neu80]. pluriharmonic [dBT80]. plurisousharmoniques [Ble82, El 80, Lel82a, LS82a]. plurisubharmonic [Kis83, Tay83]. plus [Col80b, Esn81, FS81d, Vig80f]. plus-cupping [FS81d]. Plusieurs [Nor80a, Ber84f, Est84, Mor80a]. Poincaré [BW81b, Bor84a, Bru81l, Goo83i, HH84, Hej83b, Hop83s, Kle84b, Lem81a, Maj84b, Spr83]. Point [BJ84, FF81, App83, Bai82b, Bax81, Bec84, BF84, Bro81f, Cel84, Dyd81, FH81, Fou81b, Gau81, GS83b, Had82, Hen81g, Hen81j, Izé82a, Jia81, JS81, Kle82b, Koh83, LG81e, Mag82, McA80, MMP80, NSE82, Sib80, Ste84b, Wol81b].

Point-picking [BJ84]. Pointclasses [JM83, Ste81a]. pointed [BL82]. Points [Dra83b, Maj84c, Sch80h, She82l, Spa82e, All81b, BS81, Bel81a, Ber80e, BGMY80, BG84b, Bou83g, Bro81e, BS83h, Col80b, Com82a, DMT84, Del81f, Dem80c, Dui84, Ecc80, Fri83, Geo81a, HM84b, Jam83, Kam82, Lan80c, Loh83, Lor82, Luk80a, Mar83b, Pin80a, Sid83, dCR83, vDS83]. Pointwise [Cow83, Flo80e, Col83, Edg82, HH81a, Kow80f]. pointwise-compact [Edg82]. Poisson [BV81, Bra82a, Det82, HV81a, Her84b, Kas84, Kos82, Ryd83, Wer81n]. Pol [Tro82a]. polaire [KS83b]. polaires [El 80, HM82a, Tei82]. Poland [BKS83, Ław80, Ław83, PWW80, SW84, Wer80]. Polar [Web83b, Kan82, dGdlV82]. polarization [Pan83b]. polarizations [Raw80]. polarized [Nam80a]. pole [Yor84]. pole-seeking [Yor84]. poles [KS81a]. Police [Die83e]. policies [FS83]. Polish [Car82]. polonais [Fer81a]. Pólya [Šig83l]. Polycyclic [Rhe81, Lor81a, Ros81b]. polycyclic-by-finite [Lor81a]. Polyèdre [Mer80b]. Polyèdres [BF80b, BF80c, BF80e, BF80h]. polygonal [lPBF80]. polygons [Rey80]. polygroups [Com83]. polyhedra [BF80b, BF80c, JS81]. polyhedral [BK81b]. polymer [TAP84, Whi83]. polymer-polymer [TAP84]. Polymers [Eri84b, ABHK84, Jam84, Whi83, Wis84, ZC84]. polynôme [dM80a]. Polynômes [Dra83f, Dra83i, Hür81, Dra83e]. Polynômiale [Las81a, Tit81]. polynômiales [Duv83c]. polynômiaux [Vus80]. Polynomial [Gre80e, Gre80f, Gre81b, Mon80e, Nar84e, Nar84f, Ste84a, Daw84, Dyn81, Gil81b, Har80b, Kri81, Lin82b, Llo83, Mas84c, PdS80, Sti81, Wah82, Wer82]. Polynomial-like [Nar84e]. polynomially [Sic84]. polynomials [Abh82f, Blo84a, BJ83c, Bro82a, CKL83, Dic80a, Dra81, Dra83f, Dra83e, ES83, Eve80a, GN82, Ila80, Ila82, Kau80, Kin81, Lac84a, Mag84, Maj81j, RV84, Sle84, War82i]. polyominoes [Ach84]. Polytechnique [DPT80]. polytopaler [Ewa80]. Polytopes [BF80c, BF80f, Tei81b]. ponctuelle [Col80b]. ponctuels [Gra82a, Jac81, Kri82]. Ponomarev [BB80]. Pontrjagin [Ani84]. ponts [PY82]. pooled [Aka83]. population [Gre84, Sau80]. porous [DL82, Dou82, Fie83a, Jer80, PW84]. portée [AR80a]. portion [ESV83b]. Porz [HT82a]. posedness [Tri84, Vol84]. Poset [Naz81, Cla83]. posets [Hof81c, Mar81d]. positif [Arq84, Mas83b]. positifs [Abd83, CB84a, Dem82a, Dem82b, Dem84, Sko80b, Sko80c]. position [HH82]. Positions [Dem80d]. Positive [Bur83a, Cam80b, Die83f, Die83g, Don82e, Itô80, PS81, dF82a, Are84, Bou82a, Bro83, CKLR82, CR83, Clé84, CdS81c, Dem82b, Don82f, Don82d, Don82g, EF82, FT80, Hab80b, HG82b, Hes84, Kle84b, Laz84, Mok83, Nam80d, Nus81, Raw80, Sul81a]. Positively [CR80e]. Positives [Cho83b]. positivités [DS80]. possible [And81b]. posteriori [Rei82b]. Postgraduate [MS81a]. Postulation [GP83e]. Potential [BBCH81f, BBCH81g, Krá80f, Ōsh82, Sto82b, Wer81o, BK81a, DH82, Eva81, Fic80, Haw84, Wil82a, BFF80]. potential-theoretic [Wil82a]. potentially [Ach81, Das81]. Potentials [Băn83b, AA83, Car80c, FW80b, Krá80d, Krá80h, Mae80g, Mae80i, Sig83f, Wer81l]. Potentiel [Fey84c, HM80, HM82b, HM84a, MP84b, Anc84a, Bou84a, Bou84f, El 84, Hir84a, Pie80]. Potentiels [Fey84a, Lel80, Éli84b, Fey80c, Fey84b, Gav82, Hig82b]. Pour [Ber81d, Ver81b, Abd83, BJ83b, BBP82, Car80d, Car81b, DS80, Der84, Dub84, Éli82, Fey80b, FM83, Ghi83, Gru83, Gun80a, Har80a, Hel80a, Kas82b, Kow81a, Las81c, Las81b, Las81e, Las81f, Le 82b, Led81a, Lel80, Lel82a, Mey84, Mil83a, Mis82, Năs80, Net82b, Pra83, Sko80b, Tau80, dV81]. POV [Ali82]. POV-measures [Ali82]. Power [ESV83u, She82k, Alb80, All83a, BDL80, Che84b, Gra83a, Guz81, Kör83, Kue81a, LR80a, New84, Sib80, Tam82g, Tam82i, WK80a]. power-inequality [Tam82g, Tam82i]. power-series [All83a]. powers [AHL83, Bec82a, HW81c, Kin82, Ple83h, Taf82, Wil83b]. PP [dB84a]. PQ [CQ83]. Practical [Cox82, LG81d, CG83b]. practice [LG81a]. Präsentation [Sch80a]. precedent [Éme83, Del82a, Len80a]. Prechips [Abh82l, Abh82f]. Precipitousness [Kak81]. precis [Pie80]. Precise [Ivr84c, Koo80]. Precision [Hul82]. Preconditioned [Eva83, Ket82b]. predicativity [Poh81b]. prediction [Kni83]. preduals [HHM83]. Preface [GPR80a, GPR80b, JW83]. preferences [Hwa81]. preimage [Lem82]. Préliminaires [Com82g, Jeu80e]. Preliminaries [Bou81h, BLP83e, Con80f, Goo83j, Hen81l, Hil81h, JvdS84i, Jou83i, Mon80f, PD82e, Par83i, Arn82k, Zie81f, Jeu80e]. Preliminary [Art84a, Str84e]. première [A’C80]. premiers [Les80, Lev80a, Yam80]. premisses [LS81a]. premonic [Abh82f]. Preparational [Chr83b]. prepared [Ara84]. Preprojective [AS80, AS81b, DR80a]. Prequantisation [Hor82b]. prequantization [Hor80]. prequantum [Duv80]. Prerequisites [Kha82e]. preschemes [Sot82i]. Prescribed [Por83, Tak84]. presence [HI83]. Present [DG80c]. Présentation [LLP80]. presentations [ZVC80a, Spe81b]. presented [Bau84b, Can84]. Preservation [She82l, Mun84]. preserves [SS80b]. preserving [Bro81a, FH81, Jaw81, Kre84d]. presque [Dem80d, Dim83, Tro83]. presque-holomorphes [Dim83]. presque-périodicité [Tro83]. pressure [Hey80a, Wac82]. presymplectic [GN80]. previous [Sot82i]. prévisibles [CMS80]. prewellordering [KSS81]. Primality [Len81]. primariness [Pop83b]. Primary [Cor84d, EM83a, EM83b, Lev80g, Lev80q]. Prime [DG80a, Lor81a, McA83e, Bou80, Gra84b, Hag81, Hof81c, Mon80g, New84, Pin84, Ple83h, RS82b]. prime-power [New84]. Primeiler [Lün81]. primes [Sta81]. primitifs [Yam82]. primitive [Jos83b, Lün81]. Primorsko [DP82b]. principal [GS83e, LS83a, Mau82, Ple83h, Poy83, SSW83, Wil81a]. principale [Rod81]. principality [Hil81f]. principaux [Leg82a, Les80]. Principe [Anc80, Ehr83, Hir80, Arq84, Dub82b, Hig82a]. principle [AS83, Bor83b, BS83g, Lau80a, Lau80b, McG82, Mil83b, Mor81a, Mue84, Pre84k, SC81f, Van84, Wer81l]. principles [GS82b, Hue80, Jer80]. priori [Jos84j, Jos84b, Jos84a, Jos84c]. probabiliste [Bak84d]. Probabilistes [El 81, Bou84e]. Probabilistic [MIS^{+}84a, GS82d, Kif80]. probabilité [GG84, Let81]. Probabilités [AY80, AY81, AY82a, AY82b, AY83, AY84, FMSW83, Hen80d, Hen81a, Hen82c, Hen84a, Föl81a, Gal89]. probabilities [Föl81a, Haz82, Haz84a, Haz84b, Wie83]. Probability [AFG84, Bec81, BS83f, Caj81a, DLR81, Gal89, Gue84, HJ82b, Kol83, SC81e, Acc84, Ani83, Bet82, BH82d, BŻ84, Cha81a, Csö81, GHZ83, Gir82, GQ82, GM83b, Jur80, Kin82, LR81, LR80b, Poh81a, RRM84, Sie82, TW81, Bec81, BJ83a, Hey82, Hey84, IP83b, SW84, Wer80]. Problem [ARMS82, CC82f, phAD80a, HHN84, AB83, Adi80b, Alb80, Art84b, Axe80, BHS84, Bay81b, Ben83, BFG82, Blo82a, BMP80, Böh81, Bol80, Bré81a, BL81, CC82a, CD80, Col82c, Cór80a, Cor83d, DH80, Dom83c, Est83b, Eva83, Fad83, Fai82a, FO82, FK83b, GL82c, Gre81a, GM80d, HH81a, Har81f, Hem82a, HLSW80, Hil82a, HM82d, Hum80e, IS81a, Jos81, Jos84g, Kah82, KP84a, Koh83, Kos83a, Kub83, Kur83a, Loy83, MS80a, Maj84g, Maj81e, Man81a, MR84a, MLM82, Nam80e, Nat81, Net80, Ove82b, Pal84, Par83i, Pic80c, PW84, PS84c, Reh82, SSTW82, SS84a, Sto83a, ST82, gS83a, Tak84, Tho84j, Tho84l, Tho84m, TN81, Vol84, Wil81e, WMW83, Wol83, Wuy81, Ytr84, Zie81f, Zie81e, dF84]. Problème [Bos80, Est84, Las81a, Ruí83, A’C80, Ghi83, Gui80b, Itô84d, Kah80, Las81c, Mél80, Net82b, Net82a, Paq82, Tal83, Tho81a, App83]. problèmen [Gal82b]. Problèmes [Bis81m, Dub82b, Kor80a, Let81, Ros84b, Anc80, Bar81a, Bis81o, Hel80a, Las81b]. Problems [BLS80, CGO83, CdS81g, DP82b, Gek84c, HHN84, Hop83b, KZ83a, Lor83, New80, Rau80, Tam82f, Tho84e, Acc84, ARMS82, App83, Ara84, AC82, Art83, AF82, Att84, Aub84, Bau84b, Bax81, Bax83, BH81b, Beg83, BE83b, Ber81b, Bjö84, Boc82, BT82i, Böh82, Brä82b, BN80, BN82, Bro82e, Bru82c, Bus84a, Can84, Car80b, CG83b, Cha81a, Cha84b, Che80b, Chr80a, Col80a, Col81a, Dac82c, Dah82, Dav81b, DM80c, Dră83a, Dro80, DM82, Emc81, Eng84, Enr82, Erd80, Erd81b, Erd81a, Erd82, FO82, FW82a, Fom84b, Gek84a, Gek84e, Gie83b, GH82, GS82b, Gre82a, GT84, GM80c, Gus83, Hab80a, Hac82b, HM83b, Hem83, Hem84, Hen82b, Hen81c, HS84, Hes84, Hey80a, How82, Ivr84a, Izé82a, JS82a, Jen83, Joh83b, JT82, Kac82, KKM84c, KKM84b, KKM84f, Kam82, Kan84b]. problems [KT83, Kaw82, KM83, Kno83, Krá80a, KDB^{+}84, KS83e, LV84, LM80, LMRW82a, LMRW82b, Lor82, LG81c, LS82b, MM82a, Maj81a, MR81a, Man80, Mar83b, Mar81g, Mau80, MW84, MIS^{+}84a, Men83, Min83a, Mit81, Mit82a, Mit82b, Miy83a, Mor84b, Mor80c, Mor82b, NSE82, NO83, Now82, O’M83, Pav84, Ped84b, PS81, Rin80a, Rin80b, Rob84b, Röd84, Roj80, Ruh83, Saa83, Sbo83, SS80a, Sch82p, Šig83l, Skr84, SJ84, Spi84, Ste80a, Ste82b, Tan81g, Val84c, Wac80, Wac82, Wah82, Wat84b, Web81, Wer81t, Whi84, Wil80a, dF82a, dM81a]. procedure [BH84, Dou82, GO82, PW84, RW83]. Procedures [AD83b, Wei84]. Proceedings [AFG84, AJ81, All82b, AGP81, AGT84, Ama81, AD83b, AMT82, ABGM82, AV80, AL81, BBC^{+}83, BP84a, BH81a, Ban82, BLS80, Bec81, BJ83a, BDM83, Ben80a, BFF80, BOWS82, BS83e, BS83f, BOR^{+}84b, BKS83, BKK82, CV83, Car82, Cas83a, CBJS83a, CBJS83b, CZ83a, CW82, CC82f, CGO83, Con82b, CP84b, Den82a, Den82b, DG80d, DG80c, DAP82, Dol83b, DLR81, EdJ82, Eel82, Eri84b, Eve80b, ES81, ES82, EL83, FM84, FF81, FKSW81, Fle82c, FG83, FS81e, Fuk82a, GP83a, GPRS80, Ghe83, GW81, GLM83, GMSV84, Gri84, HT82a, Hen82a, HLR83, HKLR83, HJLR84, Hey82, Hey84, Hin82, HJW83, HN83, IP83b, Izé80, Jag84, JP81, JTW82, JV82, JPW81, KR83a, KM80, KZ83a, KPT82, KS84a, KMM81, KMM83a, Ken84b, KN84, KKS82, KS83d, Kno81, KY84, Köl80]. Proceedings [KMS82, KMS84, Kom84b, KN80b, LR83a, Ław80, Ław83, LW81, LST81, Mac81c, MO84, Mal80b, Mal81a, Mal82c, Mal83, MS81a, Mar80b, MdJ82, MRW83, McA81, Men80a, MM82b, MP84b, MTT81, MR84b, Nav84b, NR80, PWW80, Pal83, PR83, PPV80, PPS83, RY81, Rao81, Rau80, RS83a, RW82, Ric81c, RSW80, Rog81d, Sal84a, SW84, TW84, Tur82, Vin84b, Wat80, Wat82, WB84, Wer80, Wil81b, Win82a, dBvR81, dFH82, vO80, vOV82, Bak84a, AY83, AY84]. process [AB83, AHE81b, BB81e, BF83g, Cuy84b, Fei82, Guz81, Izé82a, Kan80b, Kni83, Kus82, Le 84a, Leś84, Par82d, Rüs80]. Processes [AFG84, BS83e, Fuk82a, GS83d, Not82c, Not82d, Not82e, Not82f, Sto80f, Akc83, AHE81a, Ano81a, BSW84, Bas84, BBCH81e, Bre82b, Cha80b, Chi83, CRSSC83, DJ80, Det83, Doz81b, Dud81, Dud84, Ebe83, Föl81b, GQ82, GM82f, Hal80b, HKP83, HS81, HP83, Ing80, Kan82, Kle84c, KW82, Le 82a, LeP81, Mak80a, MS83d, Nie80, Not82b, Ōsh82, Per83b, PY81, PS84b, Rog84a, SC81i, SM83c, Sto82b, TW81, Tom82, Web83b, Yor84, vHSV81]. processing [JS84]. Processus [Kri82, Mas82b, Ano81a, Bak82c, BD80a, Bou81b, Bou82a, Bou84d, Bro84, CMS80, Éme80a, Éme80c, Fey80a, Har80a, Jac81, Jac84, Le 84b, MMS81, MS81b, Mey81f, Mey81e, Mey82d, Sch80p, Yan80, KMS81]. proches [Col80b]. Product [Lev80r, Lev80q, RS82a, War82k, All83b, Bro81f, Bru81l, Ceg80, Che80a, Fal83b, GBC82, Gur80, Har81q, JvdS84f, Kan80a, Kel80, Lad81, Lev80h, Lev80j, Mäk80b, Mäk80c, Oka84, OW83, Sto80a, Sto80g, Tom82, Trn84, WTL84, Wed83, Spa83a]. production [Ben83]. Products [Wal83b, vOV81j, Beh81, Ble80, CV82, DZH81, EF81b, EF81d, EF82, GH83, GHH83, Gra82b, Ila80, Joh83a, McA80, Pie83a, Pri84, Wal83a]. Produit [Las80a, Les83, Tig82]. Produits [And80, Bat81, Gui80d, Gui84a, Le 82b, Mal82d, Tah82, Tan83b]. Produkte [BKK81]. Produkten [Ros84c]. Produktzerlegung [Spa83a, Spa83b]. Prof. [ARMS82]. Professeur [Coe82]. program [AP80, Mäk80e]. Programmée [BMR81]. programming [BT82i, FS83, LS81b, Rob83b]. programs [FS84, RS83c, RW83]. Progress [DU83, Red83b, vOV81n]. progressif [Jeu80d]. progressions [New81]. progressive [Per83a, Jeu80d]. prohomotopy [LM84b]. Prohorov [Sen83a]. projectif [Ang81h, Dem81, HH82, Lan80d, LBL81]. projectifs [Sal80]. Projection [Saa83, Str80a, Gat84]. Projections [AJ83, CC82b, Cor84d]. Projective [Alb83a, Ber80d, Dic80f, GP83b, Ham82, Hof81d, Hul83, VW83a, VW83b, Alb83b, BE83a, BT82h, Bou83k, Bro81e, Cos82a, Dol82, Gre80a, HS82, Iit83, Ion84, Kec81a, Kol82a, Kon82b, LTS82, Lan80b, Lin82b, Lor81b, Lue81, MS82e, Per81a, Rob84a, Ros84a, Ste82a, Tha83, WK80b, Whe80]. projectives [Fai82a, HM82c]. projectivity [Hau81, Jon82]. projectors [Pas80]. Prolongement [Gue80, Seb80a, Str80b, Ava82]. promotor [Leh83b]. prononcée [Coe82]. Proof [BOR^{+}84b, BFPS81, DRR81c, DRR81d, ESV83b, ESV83p, ESV83m, ESV83n, ESV83o, ESV83q, FS81b, Flo80f, Kre84l, Kre84m, Poh81b, SC81p, Sot82g, Cil83, ESV83r, EP81, FHY83, Kai83, Lev80k, Liu84, Mae81, Maj81g, Mc 81, Mun80, Rog81b, Wil81f]. proof-theoretic [Liu84]. Proof-Theoretical [BFPS81, Poh81b]. proofs [GM80b, Imm84b, Jos84a, Miy81e, Rin84d, Rin84f]. Propaedeutic [Mar81d]. propagates [Mar83e]. Propagation [BLP83f, Hel80a, Las81e, Las81d, Dem84, DGZ84, Jef81, KJTB80, Las81f, Las81d]. Proper [Mek83a, Mek83b, She82h, She82m, She82n, She82o, She82j]. Properness [She82a]. Properties [Bos84, ESV83r, ESV83s, Fal83b, Kon82b, Met83, RRM84, Sak82e, AD81, Ahl80, AG83a, Arn82o, Att84, Bab80, Bot83, Bou83b, Bri81b, Bro82a, Brü82b, Čer84b, Cha83m, Cha83n, CV84a, CV84b, Erb83, Fie83b, Fom84b, Fuk84, Gar83b, GM80f, GM80g, Hen83, Jov84, KKM84c, KKM84e, Kan84c, Kar83, Kar81, KKMW81, Kha82b, Kna83c, Kni80, Kre81b, Lor82, Mas84c, Mun84, NS81a, Obe83, OS80, Pil81c, Pra82a, Roc84, Sch83c, SC81b, SC81c, SC81h, Sch81c, Sem82c, Sem82d, SM81, Ste81a, Sto80i, Trn84, Van83b, Var84, VLA83, WK80a, Yap84, Zwe84, Dra80e, SC81s]. Property [Bou83h, Pre83i, Ali82, Bie81, Cha84c, DM82, Gos84, Jak80a, KSS81, Kec81c, Mil81c, Mus83a, NS81b, Oku84i, Oku84h, Sjö82, Smi82, Sot82b]. propos [PP81, Sko80b]. proposed [ARMS82]. propre [Băn80]. propres [Erm80a, Gui80b, dM80a]. Proprétés [Bis81n]. propriété [Dur80, Hig82b, KLM81, Pis80]. Propriétés [Ber84f, Bou81b, Dim83, Fra83a, Bre83j, Col80b, Gui80d, Led84b, Mer81, Com82e]. proscribed [CH84]. prospects [Pee84]. protein [Hen80b]. Provence [GPRS80]. proving [Chu82d]. Pseudo [Jou83f, Las81d, Pre81, Avr82, Che84a, Fon84, Jou83e, Las81b, Las81e, Mor81a, de 82a]. pseudo-algebraically [Che84a]. pseudo-Boolean [Fon84]. Pseudo-Differential [Jou83f, Jou83e, Las81d]. pseudo-différentielles [Las81b]. pseudo-différentiels [Las81e]. pseudo-excessives [de 82a]. pseudo-reflexions [Avr82]. pseudo-uniform [Mor81a]. pseudoconvex [Fad83, Jak80a]. Pseudodifferential [And84b]. Pseudographic [EH80]. pseudogroupes [Kum84b]. pseudoisotopy [DI82]. pseudopotentials [Mol82b]. Pseudovariétés [Suc83]. pseudovarieties [Suc83]. PSH [Zel81e]. PSH-algebras [Zel81e]. Puebla [AL81, AL82, FG83]. punctured [RS83b]. Puppe [HJ82a]. Pura [Pal83]. Pure [Arn81, Arn82m, AV83a, AV83b, BV83h]. puzzles [Red83b]. pyramidal [Sem82b].

QE [BC81]. quadrantale [Bak81]. Quadratic [Pre84c, Pre84d, Rin84e, BCF83, Brz81, Chr83a, Dră83a, EW83, HG82b, Nam80d, PS84c, Rin84c, RW83, Wil82a]. Quadratiques [CTCMR82, Oja82, Vog82]. Quadrature [Sak82f, Bro82a, Sak82b, Sak82e]. Quadratures [Dra83g]. quadrics [Bur80]. Quadruple [Köh82, Har80c, Har81q]. Qualitative [O’F83, OS80, Con80a]. quand [Vig80f]. quandles [Neu84a]. Quantenrauschens [IvW82]. Quantifier [Wei84, BMP80, Mar83e]. quantifiers [Ber80c, Len84, Mac80, SH84]. Quantization [AD83b, KR83c, Bla83, Hes80, Lic80a, Pan83b, Śni83a]. Quantum [AFG84, ADT83, BSW84, CC82e, DP82b, Sig83h, Zuc84, Acc84, AHK82a, AHK82b, Ali84, Bea83c, Bon82a, Brä82b, Bri81a, Cas84a, CCN82, Dri82, Düm84, Ens83, FG84, Gis84, Gue84, HP84i, Kai80b, Kle84c, KR83c, MJ82, Pas83, Rob83a, Str81c, AFG84]. quartic [Vor82]. Quasi [AL80, Bro81b, AP81b, Arn82a, Bar80e, Bar82c, Bot83, Cou81, Cră83, Har81a, Har81c, Har81d, Jan83a, LJ80, Neu82b, Sha82, Sul81b, Ver82b]. quasi-autonomous [Har81a, Har81c, Har81d]. quasi-conformes [Cră83]. quasi-free [Bot83]. quasi-fuchsiens [Sul81b]. quasi-homogeneous [Bar82c]. quasi-injectives [Cou81]. quasi-isomorphism [Arn82a]. quasi-linear [Bar80e]. quasi-normable [AP81b]. quasi-symmetric [Neu82b]. Quasi-units [AL80]. Quasibound [Sig83g, Sig83d]. quasibound-state [Sig83d]. quasicoherent [vOV81c]. Quasiconformal [Car83b, ŁK83d, Mar83f, RS83b, KS83f, ŁK83a, Pie83b, Sar83]. quasiconformality [Fru83]. Quasiharmonic [Sto80h]. quasikonforme [Dit80]. quasikonformen [Küh83a, Küh83b]. quasilinear [Kar81]. quasimeromorphic [Gök83]. Quasimultipliers [Est83b]. quasioptimal [Wah82]. quasiregular [IK80, Ric83]. quasiresonance [Mey82f]. quasireversible [Pol84]. quasivarieties [vOV81b]. quaternion [Brz81, Sze82]. Quaternionic [Knu82, Luk80b]. quaternionique [LBL81]. Quaternions [Vig80e, Vig80a, Vig80c, Vig80b, GvdP80e, Vig80f]. quatre [Sie80b]. que [Del81f]. Québec [BDM83, FF81]. Queensland [BOWS82]. quelconque [Ava82]. Quelques [AT82d, Bea83a, Ber80b, BEY81, Bou84e, Dor81, Éli84b, Fey84c, Gui80d, Led84b, Mer81, ZM84a, Col80b, Des84, Éme80c]. question [McG82, Mey81f]. questions [Cau80b, HN80b, MK83, Wil82b]. queueing [Ryd83]. queues [Pol84]. quick [Has83l]. Quillen [Sna84, Tan83c, Tan83h, Tan83g, Tan83i]. Quivers [Rin84d, Bau80, Bün81, Hap81, Kač80, Kac83]. quotient [AP80, Dyd81, Heb83, NiW84, PM83, Sch82p, Slo80d, Slo80f, Wei81b]. Quotienten [Sch83t]. Quotients [Gie82r, Bac83, BBŚ82, Fra81, Lue83].

R [phAD80c, Her84b]. R. [Flo80b, Flo80f, Hel80a, HS82]. R.E. [AS84b]. Rábida [ABGM82]. Rabinowitz [DM81d]. Rademacher [SC81o]. radiating [Bar82d]. radical [Bac83, Con81b, Est83a, Lau83d, Loy83, Mon80g, Tho83, vL83a, BBC^{+}83]. Radicals [vL80, Bea82a, MW82]. radiciels [Gir83a]. radius [Bay81b, Krá80e]. Radon [Bou83h, GP84, HH81a, JvdS84j, Mus83a, Obe83]. Radonification [SC81r]. Radonifying [SC81b]. Rainwater [Flo80i]. Ramachandra [ARMS82]. Ramanujan [Alt81, And81c, BW81a, Bum84k, Mur82b]. ramifies [Szp80]. Ramsey [EH84, Len84, Mac80, She81a, SK80]. Rand [Sch83s]. Randes [Sch83l].

Randgebiet [Gru82c, Gru82d, Gru82k]. Randgebietes [Gru82i]. Random [BMKQ83, BB81e, GMW84, HN83, HP83, Maj81h, Qui84, Ant81, BR81a, BG84c, Bre82b, BŻ80, CLM84, Fra83d, Gar82d, Ger84, GHZ83, Gir83b, Gne83a, Gol82, Goo81b, GM82f, Gro82, Hei83, Hia84, HM84b, HT81, HTW81, IJ84, Jaw83, KR83b, Kue81b, Kva80, Lac84b, Pap83, Poh81a, Rac83, Ros82, RS82a, RR80, Ryc84, Ste83d, Woy83, Bis81k]. random-flight [Ste83d]. randomly [MSP83, Ryc84]. randomnesses [Ivn83]. Randoperatoren [Gru82e]. Randwertaufgaben [Hof81b]. rang [BD80a, Spe81b]. Range [MW80, Arn82o, Elm84]. Rangig [Dra80c]. Rank [Arn82f, AV83a, BS83c, Kra82b, Alb80, Arn81, Arn82d, Arn82q, AV83b, BM81b, BO82, BS83d, Eng84, Gol81a, Gol83b, How81b, Jos81, Mäk80m, Mar83g, Ros81b, Tri84, Wil81f]. rank- [Arn82d, Arn82q]. ranks [EH81b]. Rapisarda [GF84]. rappels [Spa82d]. rapport [GP81, Jac80b, MMP80, Ten84]. Rate [Bor83b, Las84, BS83g, Bra82a, Dyn81, Gaf83, Jer80, Naj84, Roh81, RS84, Wes80]. rates [Goo81b, Pau80]. rather [Sch81a]. Rational [BD84, Kle84j, Lev80s, Sou81a, BF83a, BKS82, Bök84, Bra84a, Bul84b, CC84a, CS84a, Cla81, Cuy81, DG84b, FS81c, GM81, GMJ84, GMS84, GV83, Mas84c, MP83b, O’F83, Ste84a, Tan81i, Vol81, dB84b, vI84, GMSV84]. rationaler [phAD80b]. rationality [Cat83b, Lie82]. Rationals [Sch80c]. Rationnelle [Cha84a, Tan83c, Cor81b]. rationnelles [Bea83b, Pin80d]. rationnels [Pin80a]. Ratios [Ull81]. Rauischholzhausen [JV82]. Räume [BKK81, HP84h]. Raumkeimen [Spa83a, Spa83b]. Raumwinkelkonforme [Woh83]. Ravello [CGO83]. ray [Hel81, Kni84, Le 84b, Mc 81]. Rayleigh [DT80, Sch82p]. Rayleigh-quotient [Sch82p]. RE [Fri81]. reachability [Rei84b]. reacting [KM82]. Reaction [Rot84g, BB82a, Bey83]. reaction-diffusion [BB82a]. reactions [FN82, Rot84c]. reactor [Rot84i]. Real [Brö82b, Bru82a, Mäk80f, Mäk80k, Mäk80l, Pre84a, Pre84e, Pre84f, Pre84g, Abi80, And81a, Bec82a, Bec82b, Ben82a, BT82a, BKS82, Boc82, CN84b, Dal83, Del81h, Dic80a, Djo81, Duf84, Enr83, Hur82c, Isc82, Iwa83, Kis81a, Mäk80c, Mar83e, Ohs84, Pre81, Pre84c, Pre84d, Pre84k, Rac83, She81b, Wal83c]. realisable [Bil80]. realisations [Bil81, Bil82a]. Realization [Lev80t, Lev80u, MS82c, Nam80f, Cha83p, CS84a, Lev80v, Tan81b, Zie81b, Zie81f, Zie81e, Zie81g]. realizations [EH80, EH81a, Ila80]. Realizing [PV83a, PV83b]. Reals [She82a]. rearrangement [Pop83b]. Rearrangements [Mil84a]. reasonable [Dra80e]. recents [Ten84, Ber80e]. Rechteck [Gru82c, Gru82d, Gru82j, Gru82n]. Reciprocity [Alm80, GS83b]. recognize [Kis81a]. Reconstruction [Bis81o, KKM84b, Koc82, Kum81, Vol84]. recourse [RW83]. recovery [Lev80p]. rectangle [Mag83]. rectangles [Mag83]. Rectangular [Kar80b, FW82a]. Rectification [Gal89]. Recurrence [Die83g, Die83f, Der84, Gil81b, Gui84b]. Recurrences [Die83i, GM80e, GM80h, GM80i, GM80f, GM80g, GM80j]. recurrent [Nar84c, Ōsh82, Sto82b]. Recursion [Cho84j, Nor80g, How81a, Kol81a, Mos84a, Nor80c]. recursive [CP84a, CR84b, KW84, Ler81, Miy81e, SS80b]. Recursively [Sch81a, FS81d, Joc81, Mil81c]. red [Sén83b]. Reduced [Hil81i, Bea82a, Gol81a, Han80, O’M82, Van84]. reduced-order [O’M82]. reducibility [Tan81a]. Reducing [Van83b]. réductifs [Rod81]. Reduction [Ber81b, Cas82b, FW82b, Bur80, Kan83e, Kan83i, Nam80d, SSTW82, Rea83]. Reductions [Pre83j, GvdP80a, Gil83, GS82c]. Reductive [Kow80m, FP81, Kor82, Wal83c]. réduit [Ang81h]. reduites [Cor84b, Oja82]. réel [CT82, CR82, Lan80d]. réeles [Sch80x]. reell [DK82b]. Réelle [CTCMR82, Mil83a, Rea83, dV81, Ros84e]. reellen [Ben80c]. réelles [HM82c, Ris82]. réels [CT82]. Rees [Smi82]. references [Krá80b]. Reflection [Mar80a, BB80, Col80d, Kre81c, Mar81f, Mey82f, OS81]. Reflections [Goo81a]. reflexions [Avr82, Sko80a]. reflexive [SC81s]. reflexivity [SC81h, SC81s]. Régime [Ros84b]. region [Gay84]. Regional [KKS82]. regions [GL82c, Has83c, Jac82b, JR82, Ove82a, Rei82a]. regression [Col80b]. Regular [CdS81g, Hal80a, Mac83, Mel83, Neu82b, Slo80e, Tro82b, Ben82c, BKS82, Cat84, Cha82a, Cha82b, DPV82, Dla83, Ell83, Gér80, Gie82k, GM80a, Goo80, Hay83, IS81a, Ima83, Kåg83, Kan83m, Kow80j, Lin82b, Loe80, Maj83, Men80b, Neu81, Pap84b, Qui84, Sot82d]. Regularité [Bro81d, Fer83a, dMG83, Hig82b, Mey81g]. Regularity [Bea83c, Ead81, Kan84a, Rob83b, Boi80, Gar82c, HG82a, Mas80, Mil81d, Ran82]. Regularization [BN82, Mas82a]. regulators [Cus84, Sou81b]. regulières [Bry82]. Reilly [Jür83b, Jür83b]. Reinhold [Fuc81a]. Reiten [Bau80, Bün81, Gab80]. Related [GP83a, Hop83g, Hop83n, Kre84e, Met83, Win82a, Bel81b, Car80b, Cat83b, Daw82, Daw84, DMPG83, Dom83a, Flo80i, HD83, HMT81a, Hen83, JT82, Kla81, Kra83]. relatif [Hir80]. relation [Clo81a, Gre81a, Her84a, Kre84k, Las81f, Mar83a, Mic81a, Sen83a]. relational [FS84]. Relations [DS80, Dem82a, Dra83h, Lad81, Adi80b, Clo80, Ell80, Leh83b, McA83i]. Relationship [IV80, Wil81d]. Relative [Hap81, Hil82c, Ver83d, Bak82a, Kas81, Kim81, Kró83]. relatively [AMPY83]. relatives [Bar80d]. Relativistic [Kai80b, Mar82a, NT82, Pan82]. Relativistically [Mik82b]. relativity [KS80a]. relativized [AN81]. Relaxation [Kan84b, KN82, Bra82a, Ket82b, Lev80b]. relaxed [Dac82c]. release [FW82a]. Relévement [Bis81p, DI81b, Sko80c, Tal80]. reliability [KKM84f]. reliable [GM81]. reliées [GJ81]. remain [Wol80]. remainder [Day81, Pin84]. Remark [Col81a, Mak80a, vW80, ABHK84, Alt81, BH82b, CS84b, Krá80g, Lau80b, Led84a, Miy83c]. Remarks [BT82a, BG84c, CO81, Coh83, Cze84, Dru83, Gök83, Gre80a, Grz84, Gus81, Haz84a, Hop83b, Hop83r, Isl81, MW82, Mar83c, Mit83, Sau80, vL83a, AHK82b, BDL80, Fra84, FG84, GS82a, Gie82q, Ima83, Jaj84a, Kač80, Kos81f, Lan80b, Los84, Man83, Mar84b, Mar81f, Men82, Mis84c, OV82, Ott81, Par83c, Sch82o, Šig83l, Zak81, dCR83, vdT81]. remarque [Yan83b, Zhe84].

Remarques [Bak82c, Fey82a, Fey84b, Gui80b, HZ84b, Pri82, Sko80b, Yan80, Bea83a, Ber80b, dlP80]. remember [Hen82d]. removal [Sot82c]. Rendus [BMR81]. renewal [Ket83b]. Rennes [CTCMR82]. répartition [Erm80b]. repeated [Jos80]. repères [Ber84d]. Réponse [LeL82b]. Report [Eis81, Kar80a, Kre84n, Rin80a, Fai82a]. Représentabilité [Ang81j, Die82a]. representable [Maj83]. representables [GP81]. Representant [Pie80]. Representation [Ben84b, BR81b, DG80c, Gre80g, Iof80, Pas80, vOV81k, Arq84, Cho83c, Aus82, BCS81, BV83h, Bru80, Chu82c, CRSSC84, DRS80, Fei82, Fos81, Fuk81, Gab80, Gab81b, Haf81, Han84, Hen80c, HN81b, Kle84b, Kra82a, Kri81, KS80b, Lip83, Mah84, Ott81, Rei82c, ST84, Wel83, Cha81b, Gab81c, Net82a, Zay82, de 82a, dlP80, DG80d, DG80c]. Representation-finite [BR81b, Gab80, Gab81b]. Representations [AL81, AL82, BO82, BT82c, Bru81k, BF83l, Con80g, DG80d, Fre82b, Gre80e, HLR83, HKLR83, HJLR84, Kle84k, Per81b, Rog81d, Sch82k, Sel81b, Tho81b, Wol80, Zel81c, Zel81d, BF80a, BF80h, Gui80a, Gui83, Alm80, Ber83c, BT82h, Bor83a, Bou83i, Bro82a, Bur83b, BF83k, Com83, Cur80b, Dla83, Edg83, Enr83, EF81b, EF81f, Est83b, Fek83, Goo83f, Gre80f, Gre81b, Gui82a, Gui83, Gus81, Gus82, Haz81b, Haz83, Itô80, JL80, Kač80, Kac82, Kac83, Kin81, Kir82, Kle81, Kle84j, KS82, KS83c, Kna83b, MZ83, Naz81, Roj80, She82r, Sil83, Spe81a, Tau81, Vog80, Wal83c, Bou81a, Fon83, Gui81, Las80a, LP81, Mis82, AL81, AL82]. représentées [Oja82]. representing [Ami80, Loe80]. reproducing [HNP81]. Republic [AD83b, Bak84a]. Research [HHN84, vH83a]. réseaux [Lum84]. Résidu [GR83a, LJ82]. residual [AG83a, Jac83b]. Residue [Nar84g]. residuel [Vid80]. résidus [Ang81a]. resolubles [Tau80, Duf81, FLM81, Gui81]. Resolution [CGO84, Mas82c, Pin80a, Tei80a, Tei80b, LJ80, Oda83a, Orb84, Slo80f, Tei80a, Tul80, Lel80]. Resolutions [RC82, Gre80a, Mar81a, Roc84]. resolvable [BP81b, FHV80, Har80c]. resolvantes [Hig82a]. Resolvent [CC82c, Sig83a, Sig83b]. Resolvents [BBCH81h, CdS81h, Lon84, Sto80d, Sto83a]. resonance [CS83a]. respect [BK81b, GM82e, HS84, Jak80b, Zak81, vL81]. response [Ber84e]. Restricted [Abh82n, Abh82m, Bor82c, Yas81]. restriction [Kis82, Meh83b, Sto83a]. restrictions [Abh82b, Abh82d, Abh82j, Abh82k, Abh82o, Abh82p, Van82]. result [Acz81b, Ber84c, BB83a, JS82b]. résultat [Fey82a, Loo81b, Mey80b, Mey82a]. Résultats [CR80c, Jeu80f, Bou84e, Dor81, Mar81b, Mey80a, Tal82a, ZM84a]. resulting [Pan83b]. Results [GM84b, She82c, Sig83h, Str84b, Art84a, Bec84, Bie82, Bou83e, BLP83e, BF83m, Clo80, Clo81a, Dud81, DRR81b, DRR81c, DRR81d, ESV83e, ESV83t, Erd81b, Erd81a, Erd82, EF81e, FP83, GHZ83, GT80a, Hal80b, HG82b, Har81d, Har81o, Hil84, Ivn83, Kle84f, Kow80n, Kre84k, Lou83a, Mäk80j, Mar83b, Mas84a, Mur82b, Nam80e, Nar84b, Nor80i, Pal84, Par81, Par83a, Par83b, Par83e, Ram82a, Ram82b, Rin84d, Rin84f, Rob81c, Rup83, Šče81a, Sho81, SW80, SM80, Sta81, Sta82b, Ste83d, Tha83, Woo83b, dB84b, SC81h].

Retarded [DKS84, dBC80, RCC80]. Retour [BF80i, Mey81d]. retournement [MMP80]. Retzhof [KS84a]. revêtements [Del80a, Gir83a]. reversals [Rei84b]. reverse [Ber84b, CS82c]. reversible [Wal82b]. reversing [Kre84f]. Revétements [EV82, Szp80]. Review [DP82b, Rot84j, BLP83e, Has83l, ŁK83e]. Revised [ZVC80h]. revisited [Fis81, Fos82, vOV81k]. Reynolds [Bas80, How82]. Rham [Ill83, LBL81, Fon83, Suw83]. Ribbon [Hil81j]. Riccati [Chi84b, Dră83a, Kwo83, Mag84]. Ricci [Bou81i, Der81]. Richardson [Del81a, GO82]. riches [FSY82]. rid [Pre83b]. Rident [Lag81]. Riedtmann [Gab81c]. Riemann [CC82f, Ben84a, Dwi83, Gér81a, Gra82b, Has83g, Has83j, Has83l, Iwa83, Kno83, Maj84g, Pin84, Ran82, VVV81, Ver84a, ZVC80d, Zie81b, vOV81i]. Riemannian [Ale81c, Beg83, BB81d, CO81, Der81, FJ81, Gli84, Hop83s, Kow80b, Kow80g, Kow81b, Mon84, Pet80, Sal83]. riemannienne [Bis81a]. riemanniennes [BB81c, DZM83, Gal81]. Riesz [Bak84d, BD83, BC82b, Cor83a, FW80b, GS82d, Mey84, Oku84f, Oku84i, Wer81a, Wer81p]. right [Heb83, Imm84b, Kna83c, Rhe81, Wei81b]. right-ordered [Rhe81]. Rigid [Gre82e, Sch84k, Dad80]. Rigid-analytic [Sch84k]. Rigidity [Bac80, Fur81, Tay84, Apa83, Liu80, Mun80]. Rigorous [Sig83h]. rim [Ili84]. rim-type [Ili84]. Ring [BV83a, Fle82c, Kre84k, vOV82, Bec82a, Igu82, SZ81b, Sil82, Sni83b, Sot82d, Sou81a, Vas81, War82m, Wil82b, vO80]. Ringe [Sch82d]. Ringen [Sch82g]. Rings [Alb83a, Arn82f, Ben84d, Che80b, EM83a, Hau83a, Lad83a, Mon80c, Mon80h, Mon80g, Ple83f, Rei83a, San81, Abh82h, Abh82s, Alb83b, Arn82d, Ban83a, Bea82a, BDL80, Ber80c, BC81, BKS82, BMP80, BV83e, BV83h, Bor82c, Bou83k, Bru82a, CV82, Che84b, DP82a, Dau81, Dic80a, DRS80, EM83b, Fai82b, Fis81, Göb83a, Göb83b, Gol81a, Goo80, Har80b, Hau83b, Ila80, Joh81a, Lad83b, Lin82b, Min81, Mon80d, Mon80j, Mon80i, Mon82, NV82, Oli81, Ore82, Ple83d, PdS80, Rei81, Rei83b, Rog81c, Rog84b, Röh84, Sna82, Sta82a, Sti81, Stu82, Sze82, TH84, Tay82, Ull81, VV80, Ver80b, Ver83d, WK80a, Wil81a, vOV81f, vdK81, vdL84, Kun81c]. Rio [Mac81c, Pal83]. Ritt [Cha83t]. robust [Fer84, Ket82b, Wes82]. Robuste [BEY81]. Roch [VVV81, vOV81i]. Rohlin [Sie80a]. Rojter [Rin80a]. role [KS83c, Leh83b]. Rolf [Leh83b]. Romania [BP84a, PPS83]. Romanian [CBJS83a, CBJS83b, PPS83, Leh83b]. Room [Sti83]. Roos [Lem81a]. Root [Kac83, Bra84a, Kač80, Mac81b, Yor84]. roots [Kov83, Wer82]. Rosenthal [SC81h]. rotatable [AP81a]. rotating [Par82e]. Rotations [ME80]. Roth [Lag81, Sch80j]. Rouen [Rao80]. row [Tei81a]. rows [AD81, De 83]. Royden [Mae80f]. Rubinstein [Fer81a]. Ruelle [Kif80]. Rule [Buc81a]. ruled [BS84, Hay82]. ruledness [Miy83b]. rules [SH84, vdLT81]. Rumely [Len81]. Runge [But80, Gek84g, HW82]. Ruscheweyh [BR80c]. Rutman [Nus81].

S [Bea80, Ber81d, CZ83a, Gra83d]. S. [Bar81a, Bax83, Bea80, Del80a, DS80, Esn81, Kis84b]. s.c [Tri84]. S.E.R.C [Tur82]. S.I.P. [Kar83]. S.M.F [CTCMR82]. s.p.c [Ohs84]. S.U.N.Y [Flo80j]. Saddle [vS81b, AW83]. sailing [Her82c]. Saint [FMSW83, Hen80d, Hen81a, Hen82c, Hen84a, Ano81a]. Saint-Flour [FMSW83, Hen80d, Hen81a, Hen82c, Hen84a, Ano81a]. Saks [FM83]. Salamanca [GPRS80]. same [Aka83]. Sample [GŁ84b, Alb80, Ing80, dHR81]. samples [Aka83]. sampling [BB81e]. Samuel [dM80a]. Sans [Bir83b, Bir83a, Föl81a, Mey81c, Pis80, Sch80o]. São [Izé80, dFH82]. SAP [Pre84h]. SAP-fields [Pre84h]. Satake [Nam80f]. satisfactory [DESS82]. satisfied [Bum84a, Bum84d]. satisfying [Bro82d, Lau80a, Lau80b]. saturated [Caj81b, For83a, Sch81a, Wil80b]. saturation [Ric81a, SS80b]. Satz [Lün81]. sauts [BJ83b]. scalaire [BB81c]. scalaires [Yam80]. Scalar [Bru81l, Atk82, Car80e]. scale [Bas80, Kec81c, KMM83b, O’M82, Wim80]. Scales [Mos83, Ste83a, Ste83b, Kec81a, MS83a, Mar83e]. Scarf [Par83j]. scattered [Nie84]. Scattering [Ali84, Sig83i, Sig83h, Abr82, Col82c, DM80c, Ens83, Jef80, Mik82b, Pan83b, Sec82, Sig83k, SW83, dV81]. Schauder [Sem82e]. scheduling [BFG82]. Schéma [Ang81b, Ang81i]. schémas [Bas82, Gra82a]. scheme [Kle84a, Ver82b, dM83]. schemes [BL81, But80, Com83, Con84, DW81, Ket82b, KY80, Lor82, Oda83a, Rob84a, She81d, Sna84, Tho84c, Tho84d, vOV81m]. Scheveningen [Mar80b, MdJ82]. Schiefkörper [Kne80]. Schiefkorpern [Dra80c, DK80b]. Schiefkörpers [Dra80f]. Schiffer [Tam82j]. Schloss [JV82]. Schmidt [SC81a]. School [CCCN84, DP82b, Eel82, LR83a, Tur82, dFH82]. Schottky [GvdP80j]. Schrader [Kle84b]. Schröder [Ket83b]. Schrödinger [Bro82a, Car80d, CCN82, Cow83, DK82a, Dun82, Eva81, Fle82a, Fro82a, Har81j, Lac84b, Par82c, Per81b, dV81]. Schubert [GM82a, Haz81b, RS81]. Schur [Ber84a, BB82b, BT82c, BT82g, BT82h, Gre80f, JS84, Nav84a]. Schur-like [Nav84a]. Schwartz [AP81b, Bie81, Mey80b]. Schwarz [Lel82a, Mor80a]. Schwarzian [Pre83i]. Sciences [Kom84b]. scientific [Ser84]. Scientifiques [CC82f]. Scindage [Dem82b]. Scotland [Eve80b, ES81, ES82, EL83, Gri84, Wat80, Wat82]. Scott [Ern81b, Joh81c, Mar81c, Wec81]. Scuola [Ves83]. search [Als81, BP81a, LTS82, SB80]. Second [CM82a, DG80d, Fle82b, Kan83i, Atk82, Bax83, BE83b, BW81a, Brä82b, CC82a, DI82, Eas83, Enr82, Erb83, EW83, Gek84b, Goo83l, GM80g, Hag81, HU83, HN81b, HO80, Ivr84b, Kwo83, KZ83b, Mae81, MW84, Not82f, PD82d, Sot82l, Tam82c, Tam82i, dCR83]. second-order [Atk82, BE83b, CC82a, Eas83, EW83, HU83, Ivr84b]. secondary [Web81]. secteurs [Cha83t]. section [Leg82f]. Sections [Bau80, ESV83u, Leg82f, Nor80h, Shi81a, Tal80, Băd84, Con82a, Con83, Die82a, Gie82c, Gie82u, GM82b, GM82c, Leg82e, MP83b, Sko80c, Som81, Tei82]. seeking [Yor84]. segment [Mae81]. segments [BV83d, Hay82]. Segrè [Bur80]. Seidel [Bra82a]. Seidenberg [Ric81d]. Seifert [HN80a]. Seite [Sch83s]. seizième [A’C80]. séjours [Bro84]. Sekiguchi [Yan83a]. Selberg [Chr83c, Chr83e, Chr83d, Goo83h, Hej83c, Hej83f, Hej83g, OW83].

Selbergsche [phAD80c]. Selected [Bou83j, Hej83d, Hop83o]. selection [Ers80, Han84, Mau80, Wag80]. selections [Gra80, Los80]. selectors [GT80a]. Self [Bou82c, For84a, FK83b, Mae80g, TV84, vHSV81, Das81, Dun82, EG81, ELZ83, Fai82b, Fle82a, Gan81, Gre82a, HZ84a, Jan82, Maj81i, Mor80c, Mor82b, Ver80a]. Self-adjoint [FK83b, Mae80g, EG81]. self-adjointness [Dun82]. self-complementary [Das81, Gan81]. self-converse [Das81]. Self-decomposability [For84a]. Self-decomposable [vHSV81, Jan82]. Self-dual [TV84]. Self-duality [Bou82c]. self-gravitating [Ver80a]. self-injective [Fai82b]. self-maps [HZ84a]. self-similar [Maj81i]. Selfadjoint [CdS81i, Bro83, CdS81a, CdS81b, CdS81c, Ped84b]. selfdecomposability [Jur81]. Selfdual [Ple83i, Ple83c]. selfmaps [JS81]. Selick [Hus81a]. selon [Gui80a]. semantics [FS84, GM84a]. Semi [Cos82a, Jeu80g, Mae81, Mar81g, Sch80u, Sch80v, She82j, Abd83, Aup83, Ber80c, Bis81d, BMP80, Bou81a, Bou82a, Cos82b, Czé83, Dem82b, EF81d, EF81e, Fur81, GP81, Hab80b, Har80a, Haz84a, Hul80, Itô80, Len80a, Lev82a, Mar84b, Sch80p, Sch80q, Sch80t, Sch80x, Sch80w, Sch82n, Sil82, Sko80b, Sko80c, Wal82b, Yam80, Jeu80g, Sch80v]. semi- [Haz84a]. semi-algébriques [Cos82b]. semi-continuous [Mar84b]. semi-corrigé [Har80a]. semi-direct [EF81d]. semi-distributivity [Czé83]. Semi-formal [Mae81]. Semi-Fredholm [Mar81g]. semi-group [Itô80]. semi-groupe [Bou82a]. semi-groups [Hul80]. semi-local [Sil82]. semi-martingale [Sch80p, Sch80x, Sch80w, Wal82b]. Semi-Martingales [Jeu80g, Sch80v, Sch80u, Bis81d, GP81, Sch80p, Sch80q, Sch80t, Sch82n, Jeu80g, Sch80v]. Semi-normality [Cos82a]. semi-positifs [Abd83, Dem82b, Sko80b, Sko80c]. semi-positive [Dem82b]. semi-premiers [Yam80]. Semi-Proper [She82j]. semi-simple [Aup83, Ber80c, BMP80, EF81e, Fur81, Hab80b, Lev82a]. semi-simples [Bou81a]. semialgebraischen [DK82b]. Semibounded [CdS81j]. Semicontinuity [Dac82e]. semicrystalline [ZC84]. semidefinite [CKLR82]. Semidiscrete [Tho84l]. semidynamical [Izé82a]. semifinite [Jaj84b]. Semigroup [Abh82o, Abh82p, Rup84e, Abh82j, Abh82k, BH82d, Dau81, Mor84a, Thi81]. semigroup-valued [Mor84a]. Semigroupe [FM83, dMG83]. Semigroupes [Tro83]. Semigroups [Abh82q, CKL83, HJW83, KS84a, Kur82, Rup84a, Rup84f, Sta84, Ali84, Are84, BG84a, Bem80, Bra84b, BS83h, BŻ84, FG84, GM84b, Gra83b, Heb83, HL83b, Hof83, Ima83, JW83, Jür83b, Kan83m, Kin82, Kin84c, Laj81, Las84, Le 82a, Lue83, MW81, McC82, Mel83, MS83c, MN82, Pav84, Rei84a, Rup83, Sto80g, Str82b, Wei81b, Win82b, JPW81]. semilattices [BPV83, LM83, Str81a]. Semilinear [Hen81e, Bar83b, Bec80, Hes84, HO80, dF82a]. semimartingale [Bar82b, Jac80b]. semimartingales [Cho80a, Gri82, HYZ83, Jac84, Pra83, Str80a, Str80b, Str84g, Str84h]. Séminaire [GIR81d, KO81a, Nor80a, Ano80, Ano81b, Ano81c, AY80, AY81, AY82a, AY82b, AY83, AY84, DPT80, GR83b, HM80, HM82b, HM84a, LS80, LS82a, LDS83, Mal80b, Mal81a, Mal82c, Mal83]. Seminar [BJ83a, Car82, CBJS83a, CBJS83b, CC82f, CCCN84, DK80b, HN81a, Hop83d, JTW82, KZ83a, KMM81, KMM83a, Lad83a, PPS83, Ves83, Lad83b]. Seminário [Mac81c]. Seminars [Ben80a]. seminorm [LR80b]. Seminormality [Day81, DR81]. semiorderings [Pre84b, Pre84j]. Semiprime [Fis81, Mon80d]. semiregular [Luk80a]. semirings [Zel81f]. semisimple [Bow82, Del81h, Enr83, FJ81, FJ84, Her83a, Hum80f, Jos83b, KS81b, Lop84, MW81, Osh81, Ric81b, Zuc84]. semistable [Haz84b, Jan82, LR80b, Meh83b, RRM84]. Semitopological [Rup84a, Rup83]. semitopologiques [Tro83]. semiuniverselles [Sal82a]. Sens [MR80, Bou84e, Hir84b]. sense [ESV83a, MR80]. Sensitivity [Kom84a, Ped84b, Rab84, Kom84b]. Separable [EM83a, Hub83a, Mek81, Ran83a, BR81c, EM83b, GT84, Han84, Hub83b, Ran83b, Sid83]. séparablement [Dur80]. separably [Gre83]. Separating [Rin83, Bjö84, Rin84a]. Séparation [BF80b, BF80j, HM82c, BF80b, Gie82h, Koo80, Spa83a]. Separatrices [HN80a]. separatrix [Rey80]. September [AFG84, CP84b, DP82b, GPRS80, PWW80, PPS83, Rau80, Wer80, dV84, She82o]. sequence [BB81a, Bil81, BFG82, BJ80, Bro82a, Cha83r, DW81, DR80b, EH80, Hil81b, Ket83b, Lau83d, Lyu84]. Sequenceable [Kee83]. Sequences [Nar84g, New81, Ach81, Bel83b, Bil80, But81, CLM84, Cha81c, Cho81, Das81, EH81a, Gab80, Gan81, Gil81d, Gly83, Hen80b, Jac82b, Lev80l, Lev80o, Lev80p, Maj83, Mat82, McA83f, McA83g, Nar84c, Nar84f, Pre83f, Rei82c, Ryc84, Sbo83, Taf82, Wei81a, vdHM84]. Sequential [Flo80g, Ste84b, Ul’84]. Serge [Ray83]. serie [Les83, Rod81]. Series [Dra83c, ESV83u, JM82a, Oku84g, Spr83, All83a, Art80a, BDL80, Bos84, Bru81l, Bum84a, Bum84e, Bum84g, Che84b, Chr83c, Chr83f, Chr83d, CL84b, Goo83i, Gra83a, Hej83b, HV81b, Isl81, Kin81, Kör83, Kva80, Oku84a, Oku84b, Oku84c, Oku84e, Oku84d, Oku84f, Oku84i, Oku84h, Poy83, Rin83, Ruc82, Sab84, Sch83r, Sib80, WK80a, Bay80, Bay81a, Lem81a]. Serre [Kat81, Ric81b, Sch83l]. Serre-Kompaktifizierung [Sch83l]. Serre-tate [Kat81]. ses [Tal83]. Session [Cer84a, Con82b, Ghe83, GS84, Kom84b, MM82b, Sal84b, dV84, ARMS82]. Set [Cho84i, Čob84, Ekl83a, Ekl83b, GS83d, HW81a, HMT^{+}81b, Hub83a, Kum81, Ada82, AG83a, AN81, Bor84a, Bro82a, DG84b, Gat84, Gre81c, Han80, Hay81, HMT81a, Hub83b, Kaw81, Luk80a, Mor84a, Šče81a, Sto83a, TT81, TN81, Val84b, dB84b, JP81]. Set-parametered [HW81a]. Set-reconstruction [Kum81]. Set-theoretic [Cho84i, Gat84, Val84b]. Set-valued [Čob84]. Sets [Bou83h, Flo80j, MR84b, AK82, All82a, AMPY83, Boc82, Bün81, Caj81b, Cho84c, Cho84e, CJ84, Cho84g, Cre84, Daw81, Daw84, Del81g, Die82b, Don82c, Edg82, EHK81, Ern81b, Ers80, Flo80c, Flo80d, Gau83, GHZ83, Gla81, Grz82, HN80a, Hin84, JK83, JS81, Joc81, Kak81, Kan82, Kec81c, Kel80, KP84b, Küh80, Lou83a, Mañ81b, Mar83b, Mos83, Nat81, Neu82b, Rob83b, Rom83, Sei81, Sho84, SW80, Sic84, Ste83b, Sul81a, Sze81, Web83b, Wec81, Wet83]. setting [DK82c, SY84]. Seventh [ES82, RSW80]. Several [Šig83l, Vla82, AF81, dBC80, Col81a, Fei84, GL82c, Ivn83, Iwa83, Kam82, Sem82b]. Severi [Art82a]. Sextic [Mäk80f, Mäk80c, Mäk80l]. Shaneson [Lan80d]. Shape [HHB84, Bau81, Bau84a, Cat81, HI81a, HI81b, LM84b, Mdz81, She81c, Ven81, MS81a]. shaped [Tho83]. sharply [Küh80]. sheaves [Bor82a, BV83c, BV83g, Dem82b, DR82b, Kas83, Man82, PR82, vOV81c, vOV81m]. sheets [MSP83]. Shelah [Mil84b]. Sherbrooke [BDM83, FF81]. Shift [Has83k, Ruz82]. shift-convergence [Ruz82]. Shift-invariant [Has83k]. shifts [GT83]. Shigefumi [Dem81]. Shih [Leg82b]. Shimura [DMOyS81c, MS81c, MS81d]. ship [Her82c]. shipdesign [AB83]. sided [Ruh83]. Sidon [Blo84b]. Siegel [Lag81, Chr83f, Mum83, Nam80f, Nam80g, Nam80j]. Siegen [KN80b]. sieve [All84]. sigma [For83b]. Sign [Ros83b, Fie83b]. Sign-embeddings [Ros83b]. sign-patterns [Fie83b]. signal [JS84]. signals [Glo83]. signature [Mäk80m]. signatures [Brö82b]. signe [Jac81]. Silva [Bia81, MN81]. Silva-holomorphy [Bia81, MN81]. similar [Maj81i]. Similarity [Kan83j, Kan83k, Kan83o]. Simion [Caz83, Leh83b]. Simple [Cha83j, EH81a, Hop83p, IACT81, Kac81, Lex83, Mol82b, Mon80i, Neu82a, Slo80h, Slo80g, Sot82h, Sot82i, Arn82i, Arn82l, Arn82n, Aup83, Ber80c, BF82, Blo82a, BMP80, CR84a, EF81e, Fer81e, FL84a, Fur81, Goo80, Hab80b, Kai83, Kap82, Lev82a, Mit82b, Pog80, Pra84, RW83, Slo80b, Slo80c, Sot82b, Sot82c, Sot82d, Sot82g, Sot82k, Wil81e, Wil81f]. simple-homotopy [Fer81e]. simple-minded [Kai83]. simples [Bou81a, Tah82]. simplest [Szű80]. Simplexes [BF80i]. simplical [Fou81b]. Simplices [Mun80]. simplicial [Geo81c]. simpliciaux [Leg82c]. simplified [EG81, KP84a]. Simpliziale [Sch82h]. simply [Slo82]. Simulation [Dou82, HN83, Bas80, DT80, Fas80, Ivn83]. simultanée [Pin80a, Tei80a, Tei80b]. Simultaneous [Sch80k, Sch80l, Sch82p, HV81a, Lem82, Wer82, dB84b, Tei80a]. sinc [Ste84a]. sine [BHS81, Chr80b, Hol81]. sine-Gordon [BHS81, Chr80b, Hol81]. Singapore [KY84]. single [Bar83c, Rei84b, Sem82a, Sig83a]. single-channel [Sig83a]. Singular [CMT^{+}84, EdJ82, FT80, GP83d, GM83a, Gre82a, Hab80a, HO80, HM82d, Kan83l, Maj84c, Mik82a, Axe80, Bra80, Bré81a, Bru84, Con80a, DH80, Eve80a, FN82, Fle82a, FW82b, Har81q, Hem82a, Hem83, JP84b, Kam82, Kaw82, Lau80b, LS82b, Naj84, NSE82, O’M82, PP82b, Sib80, Ste83f, gS83a, Tho84m, Wer81t, dO80a, vH82, Gui83]. Singular-singularly [Mik82a]. singularité [KS83b]. Singularités [DPT80, GM82b, GM82c, Gra82a, Lan80c, Las81d, Pin80b, Pin80c, Pin80d, Bar81a, Bea82b, Bry82, Com82b, Dem84, Hel80a, Las81e, Las81f, LJ80]. Singularities [BLP83f, CGO84, Fic80, Hop83q, Hop83s, KS81a, Mar83g, Slo80h, Wer81q, Whi84, AHL81b, BY82, Bey83, BvS83, DRR81f, DRR81e, Gau83, Gér80, GV83, Hue80, Las81d, MP84a, NiW84, Oda83a, Ort80, Roc84, Sha80, Slo80b, Slo80i, Slo80g, Slo82, Slo83, Szű80]. Singularity [Kos81e, Bro81b]. Singularly [BV83j, FO82, Dră83a, DH83, How82, Mik82a, Moe82]. singulières [Cer81, CMS83, GH84, Gui83, Mey80c]. singuliers [Ber80e]. sinks [FY81]. Site [Bre81b, BBM82e]. Six [SC81b]. Sixth [ES81, Hey82]. sixtieth [Sul81a]. size [Mor80b, Vog80]. sizes [Kum81]. Sjöstrand [Hel80a]. sketch [Hop83e]. Skew [Mon80j, WK80a, CS82c, CMH80, Har80b, Kur83a, PdS80]. Skolem [Sle84]. Skorohod [Har80a]. slender [GRW81]. Slepian [Lin84a]. slightly [PW84]. Slow [Aub84, O’M83, Bak83]. Slow/fast [O’M83]. Smale [AG80, Bas81, FY81, Rob81a]. Small [Hop83e, Llo83, Bro82c, Chr80a, Hem84, Ion84, SM83a, Tes83, Wol82b]. SMF [AGT84]. Smith [Mor81b]. Smoller [Con80b]. Smooth [Aub84, BK81b, Frö82b, Lan81b, Maj84a, PP82a, Tho84n, All81a, ADT83, Hei84, Led83, Sap84, Shi81b, Ste82a, dB84a]. Smoothability [Gut83b]. smoother [EHP83]. Smoothing [Har81k, LT84a, Ben83, CP82, Far82]. smoothness [Mha84, SC81s]. sober [Joh81c, Hof81d]. Sobolev [BM82a, BM82b, Ehr84, Gos84, Hed80, Hen82d, Mer84, Poo83]. Sobre [Lew82]. social [Gil81c]. Society [Kom84b]. software [DESS82, Duf82b, FW82a, Rus82a, Ste80b]. Solar [HLSW80]. solenoidal [GY82]. solenoids [Kru84a]. Solid [Woh83]. solids [Slo83]. Solitary [LM84a]. solitons [Cor80c, Cor82, KR83c, LM84a]. soluble [Rob84b]. Solution [AGP81, AMT82, GLCS80, Wil83a, Bar83d, BR81a, BGM^{+}80, Bru84, Chi84b, Cuy84e, DH80, DMPG83, Dom83c, Enr82, Fas80, FW82a, Fro82a, Hac82b, Hem82a, Hem82b, Hen81i, How82, MP83a, Mit81, NSE82, Net82a, Pan82, RW83, SSTW82, Wac80, Wat84b, Wer81u, Wil84]. Solutions [Duv83c, Gru84, Har81i, Las81d, Las81a, PHM^{+}84, Rot84g, AF81, All82d, Amb83, Ami80, Art84a, AHS82, Aul83, Bar83c, Bar80e, Bec80, BCF82, Ber82a, BS83b, Böh81, Bol80, BHS81, CMA83, CG82, Ces83, Cho80b, Chr80b, Chu82a, CC84b, Col83, Col80d, CZ83b, Cow83, Cus83, DK82a, DI83, Dem83a, Dui84, Duv83a, ES83, Eng84, Eng83b, Gal83a, Gér80, Har81h, Hes84, Hey80b, JM81a, Kar81, Kas84, Kni80, Kos83a, Kos83b, Kot81, KB80, Kun81a, Lum80a, Mac83, Maj84a, Maj84e, Mar82a, Mas80, MB82a, Men80b, Mey81g, MB82b, Mon80a, Net82a, Ole83, Pav84, PS81, Pet81, PP82b, Rot84l, Sap84, Sch82l, Sch82o, Seb80a, Ser80, Sib80, Ver84a, Ver80a, Wer81h, Yan83b, Zeh83, dF82a]. Solvability [Cor84e, Mil82, Bor84c, CPD81, MNP81, Mil80b, Sag80]. solvable [Bau84b, Kow80e]. solve [Jos84g]. solved [WMW83]. solver [Mit82a]. Solving [Noc82, Sha82, CC82a, Cha83s, Duf83, GO82, GL82c, JT82, Koj81, KY80, Kub83, NO83, Rin80a, Rin80b, Saa83, Sec82, Tod82]. Some [Acc84, AHL81b, AHK82b, Ara84, AF84, AF82, Bac83, Bak84c, BL80a, Bet82, Bie82, Bjö84, Boc82, Bre84, Bri81b, Bro81e, BCP83, But80, CS80a, Car80b, Cha84b, CV84b, CdS81k, Cor80c, Daf81, DJK81, Dud81, Eis83, Emc81, Eng82, Erd80, Erd81b, Erd81a, Erd82, Ers80, FR83, Fin82, FP83, Flo80h, Fra84, Gab81a, GS82a, Gie82s, Gir84b, GT80a, GM80f, GM80g, Hal80b, HG82b, Har81l, Har81m, Har81o, Haw84, HM83b, Hej83e, Hel80b, Hil84, Hop83r, Ima83, Jac82b, JW82, Jan82, Jef81, Jou83j, Jov84, Kač80, Kac82, Kan83m, Kap82, KS81b, Kni80, Koc82, Kos81f, Kos80b, Kow80n, Kre84a, Kun81b, KT84, Lan80b, Laz84, Los84, Lou84, Lou83a, LS81b, LG81e, Lyn84, Mac81b, MR81a, Man83, Mar84b, Mar81f, Mar83b, Mas84a, Mas84c, Mau80, McK80]. Some [McL82, Mei80, Men82, MK83, Mis84c, Mol83, Mon84, Mor84b, Mur82b, Mus83g, Nor80i, NS81a, OV82, Ott81, Par81, Pav84, Pis83, PP82b, Pre84i, Ran80, Ren80e, Roc84, Röd84, Ryc84, Saz81c, SH81b, Sen84, SW80, Sla84, Som82b, Sot82j, SM80, SM81, Ste80a, Ste83d, Str83a, Tan81j, Tan81k, Tau81, Val84a, Wer81u, Wil82b, Woo83b, Zak81, Zie81d, dB84b, dCR83, vOV81l, vdD80, vdT81, AC82, Art83, AS80, BPV83, Bie81, Bil82b, BKS82, Brä82b, Bro81a, Bro83, Bru82c, Can84, Cha81a, CL84b, Con84, Daw82, DM82, DRR81g, Dup80, Eas83, Elb82, Eng83b, Fie83b, FG84, Fru83, Gne83b, Goo83b, Goo83c, GS82b, Grz84, Gul83, GM80c, Har81k, Hen83, Hen84b, Jak80b, KKM84f, KŁLS80, Ket83a, Kim81, Kom84a, Led83, Lon84, LMRW82a, LMRW82b, MM82a]. some [Maj81g, Mas84b, Meh83b, Mil80b, Ore82, Pas80, Per84, Pog80, Ram82a, Ram82b, Rin84d, Rin84f, Sot82g, SW81, Thi81, Van83c, Vol82b, Web83b, Wer81t, Wet83, WK80a]. Sommes [Del80a, Kah80]. sont [Mer80a]. sound [vH83b]. sources [FY81, Lor81a]. Sous [Bea81, Sch80x, Sch80w, Spa82a, DFM82, Del84b, Kis82, Rab83, Ric81a, Spa82e, Spe81b, Tal83, Sch80x, Sch80w]. sous-additives [DFM82]. sous-analytiques [Rab83]. sous-elliptiques [Tal83]. Sous-espaces [Bea81, Sch80x, Sch80w, Sch80x, Sch80w]. Sous-groupes [Spa82a, Spa82e]. sous-langages [Ric81a]. sous-monoïdes [Spe81b]. sous-noyaux [Del84b]. sous-variété [Kis82]. Souslin [Kec81c, She82p]. Southeast [KY84]. Southern [Ama81]. Space [Hol81, Kan83o, NT82, PPS83, Abi80, AP81b, Ant81, AF84, Art84a, Aso84, BD83, Bec80, Ber80d, Bin83, BL80b, Bou83b, Bre84, Bru81m, BM80, CC82b, Cas84b, yCV81, sC83, Det83, Ebe83, Edg80, Elm84, FJ81, Fuc81b, Fuk84, Gar82a, GP84, Gie82k, Gie83a, GHZ83, Goo81b, GL82d, HG82a, Har81a, Hel81, HT82b, HV81b, HJ82c, Jam83, Jaw81, Kal83, KS83a, KP84a, Kot81, Kue81b, Loe80, Mak80a, MS82e, MM83, Mun80, Nor80b, Pap84a, Pre84j, Rac83, Rin80a, Rin80b, Ros84a, SC81t, Sig83k, Soa82, Sor83, Spa83a, Tan81g, Wed83, Ytr84]. Space- [Hol81]. space-homogeneous [Elm84]. Space-time [NT82]. Spaces [Bec81, BJ83a, BS83f, Caj81a, CW82, CP84b, Gie82f, Gre83, Kha82a, Kow80h, Mae80a, Mae80h, Mar82a, Mus83f, Nam80j, Sch83d, Sch83e, SC81e, Sem82e, SW84, War82k, Wer80, And81b, Ara83, BS80b, BF80c, Bay81b, Bed82, Beh81, Ben82a, BH83, BH82a, Ben82c, Bie81, Bjo82, BH80a, Blo80, BBCH81c, Bor83b, BG84c, BH80b, Bou81d, Bou81c, Bou81f, Bou81g, BPK^{+}84, Bou83c, Bou83f, Bri83, BM80, Bur81, Car80a, Car83b, Cat83b, Cha82b, Che80a, CR83, Clé84, CdS81c, CN84a, CRW^{+}80, Col80c, CM81, CPD81, Con80d, CS81b, Da 84, DJ80, Det84, DKS84, Din82, Din81, Doi84, Don82f, Doy81, Doz81a, DM82, Dyd81, Fal83a, FJ84, Flo80e, Flo80g, Flo80h, For83b, For84b, Fro82c, Fro84, Fru83, Fuc81b, Gar83a, Gar82d, Gie82c, Gie82b, Gie82g, Gie82s]. spaces [Gie82u, GBC82, Gil83, Gin80, Gin81, Gin83, Gol83b, GT80b, Gos84, Gre82d, Gre82e, Gul83, Had82, HH81a, HH84, Han84, HZ84a, Haz84b, Hed80, Hei84, HM83b, Hel84b, Hen83, HM83c, Her82b, Her80c, Hey81, HJ82b, Hof81d, Hof82a, HT81, HTW81, Huš82b, Hus81a, Hyl81, Igu82, Ili84, IJ84, Iva84, JM81b, Jaj80, JTW83, Jan82, Jan83b, JvdS84d, JvdS84e, JvdS84f, Jun83, Jun84, Jur80, Jur81, Kah82, Kar80b, Kel80, Kha82c, Kha82d, Kha82e, Kha82f, Kha82g, Kha82j, Kir81, Kol84, Kow80a, Kow80b, Kow80d, Kow80e, Kow80f, Kow80g, Kow80i, Kow80k, Kow80m, Kra80i, Kré81a, Kva80, Lan82c, Ław82, Led83, Led84a, Leg82f, Lem82, Lin83, LM84b, Lop84, Lor83, Los82, Mae80e, Mae80g, Mal84, Man81a, Mas82a, MNP81, Mer84, Mik82a, MT80b, Mit84, Mor81a]. spaces [Mus83a, Mus83b, Mus83d, Mus83e, Mus83g, Nag82, Nai84, Ohn84, Onn82, Osh81, Ōsh82, Par83d, Pau80, Pau83, Pee84, PR82, Per81a, Per84, Poo83, Pop83b, PT82, RU82, Rob84a, Ros82, Sch82a, Sch82b, Sch83c, SC81j, SC81m, SC81s, Shi81b, Sid83, Sjö82, Soa83, Ste82a, Sto80a, Sto80g, Sto80h, Tha83, Tog81, Trn84, Var84, Vek84, Vog84, Wal82a, Wal84, Wol82a, Woy82, Woy83, Zuc84, vDS83]. spacetime [Kai80b]. Spain [ABGM82, Nav84b]. Spanish [Lew82]. spanning [Hil81g]. Sparse [Duf82b, Duf82a, Duf83, Saa83, Sch82p, Tho82]. spatial [Cae80, Sar83]. Special [BS83f, HLR83, HKLR83, HJLR84, Jos84k, Kha82f, Kha82g, Kom84b, She82p, All82a, Ars81, BR81c, Bré81a, Cho80a, CdS81k, Doi84, Gie83a, GN80, Soa82, Wei81b]. spéciales [Cho80a]. specie [Bar83c]. specification [Dat83]. Spector [HN81b]. Spectra [CdS81l, Ber83b, Brö82b, Eva81, ELZ83, Gra82d, Hof81c, MT84]. Spectral [Bot83, Bru81n, DO80b, Gir83b, HN81a, Hed80, HN82, Ivr84c, Kan83n, Kan83o, Lan82c, Ros84a, Sig83j, WSF^{+}84, Atk82, BJ80, DW81, Dom80, Eri83, Gil81c, Gil81d, Gru82t, Har81o, Hej83b, KN82, KW82, Kub83, Maj81h, Sjö82, Ste80a, Wei80]. Spectrale [Tan83h, Leg82b]. spectrality [Kan83o]. Spectre [LBL81, CR82]. Spectrum [BV83i, Atk80, BM81c, Fli82a, HW81b, Jac83b, MV81a, Sig83g, Str82b]. speech [Cho84l]. speed [Her82c]. Spektralsatz [Gru82t]. Sperner [vdLT81]. Spezielle [Gru82q]. speziellen [Küh83a, Sch83p]. Sphären [Ewa80, Löf80]. sphere [AHL83, Jaw81, Tho81a]. spheres [Ben84e, Ber80d, Loh83, Neu80, Sie80a, Uhl82, Bak84d, Ros84e]. Spherical [IP83a, CR81, Dui82, Gui83, Her83b, HJ82c, KB80]. spherically [Bre82b, HK80a]. sphériques [Gui83]. Spheroidal [MSW80c, MSW80d]. Spin [Kre84f, Gal80, Hel83, JS82a, MP82, Pap84b]. Spin- [Kre84f]. spin-systems [JS82a]. spinning [Duv80]. Spinor [GP82]. spiralés [Cha83t]. spirit [Fro84]. Spitzenformen [Sch83n]. Spline [Gru82c, Gru82d, Cox82, PP82a, Sch82l]. Splines [Bar84, Gru82q, Gru82f, Gru82u, Wer81t, Gru82b]. split [BB81a, But81, GvdP80k, Hab80b, Rei82c]. Splitting [HH81b, Lad83a, Sze82, Ami82, Cha83q, Cha83r, Dem82b, Lad83b, Maj84e]. Sprachen [Jür83b]. Spring [Flo80j, Hin82, Kom84b]. Springer [Gol83a]. Spuren [Stu80]. spurious [PS81]. Square [Tre84, Bra84a, Tam82e, Yor84]. square-root [Yor84]. squarefree [LL81]. squares [Bar83d, Bou83k, Kea83, Sti83, Van83c]. Stabilité [Kis82, Ros84b]. Stability [CM80a, Coo84, DRR81f, DRR81e, Gus83, Hal81, HTW81, IK80, Kat80a, Nag82, PD82f, Sar83, Sus82b, Van84, dC81, vdK81, Bar83c, BvS83, CKMN83, CN82, DO80b, Gne83b, Hed80, Hen81b, Hoh83, Hol80, JM81a, Jur80, Jur81, KP84a, KM83, Koi84, Kol82b, Leu82, Lew82, MV81a, MP81, MK83, OR83, Par83g, Pet83, Shi83b, dMPvS81, vS81b, KZ83a]. stabilization [Tan81d]. Stabilizing [Daf83]. Stable [Arn82o, DP82a, DG84c, Gar82d, Gek84c, Haz82, Haz84b, LMW80, Moi81, Bal81a, Ber80d, Eng82, Far84, FHY83, Geo81c, GS83e, HK80a, HW82, Haz84a, Hwa81, KMM83b, Lin84a, Mas84b, MS83d, Rin84a, Rob81b, Sot82l, Wal84, Wol81b, dO80a]. stables [Gal82c, Sch80x, Sch80w]. stably [MV80, MV81b]. stage [BH84, GO82]. stair [Aub82]. Standard [BBCH81i, Bru81o, DJ80, Jan83b, LP84a, ABHK84, BH80a, BBCH81e, BBCH81g, Isl81, Liu84, LG81a, LG81b, LG81c, LG81e, Ohn84, Pop83a, Rot84j, Tho84o]. Standing [KM82]. Stanford [Hop83d]. Star [Bre82a, FS84, Ise81, Tho83]. star-finite [FS84]. star-shaped [Tho83]. stars [Ise84b, Ise84a]. start [AK82]. State [BBC^{+}83, CW82, Ric81c, Bot83, FS83, Fuc81b, SH81a, Sha82, Sig83d, Ste83d]. state-space [Fuc81b]. Statement [Lev80v, DRR81b, GM80d, Par83i]. Statements [ESV83t, Mil80a]. States [GMSV84, Dre83, MN82, Sig83g]. Stationary [Die83c, FS83, Aul83, FG84, Kre81c, Mak80a, Nie80, Die83d]. Stationnaires [Die83c, Die83d]. stations [JvdS84a]. Statistical [D’A84, Gar83b, Rao81, AHK82b, Ale83a, iA83, Car80b, Emc81, Gro82, GF84, HD83, LP84b, Mil83b, Ano81a]. Statistics [Ant81, IP83b, KKM84c, KKM84b, Ruy80]. Statistique [BEY81, DG83, Kri82, Rao80]. Statistiques [Rao80, Ano81a, LC80]. Status [KS82]. Steady [Ami80, Bab80, Dre83, SH81a, Sha82]. Steenrod [Bau81, Kah82, Lam84]. Stein [For84b, Jen80, JS82b, Kra80i, LT84b, Lin81, Mey82e]. Steinberg [Hue82]. Steiner [Doy81, LZ82]. Steinitzproblems [Ewa80]. Stellen [phAD80b]. Stem [Hue82]. step [BF83d, May80]. Stephan [BR80c]. steps [Mal80c]. stetige [phAD80d]. Stieltjes [Bar83a, Ber80b, Bos84, Gel81a, Hön80, Hön82, Min83h]. Stieltjes-integral [Hön80, Hön82]. stiff [Bjö84, Dah82, Gea82, Sha84a]. stiffly [Eng82]. Stillwell [ZVC80h]. Stochastic [AHE81b, CRSSC84, Det84, Ell81, Elw81, Gri82, HKP83, KZ83a, Kos83b, Kun82, Kun84, Lin83, MM82b, Par82d, Per83a, Per83b, Rog81a, Sal84a, All82d, BSW84, Ben82b, BB81e, Chi83, Det83, Doz81b, Fer84, Gal83a, Gue84, HW81a, JM81a, Kle84c, Kni83, Kot81, Kun81a, Le 84a, MB82b, MK83, Mit82c, MN82, PMC83, Par82b, Per82b, Pin81, RW83, RS82a, SC81i, SM83c, Str81c, Str83a, Tal82b, ZM84b, dF84, Bis81h, TW84, Wil81b]. Stochasticity [GM80h, GM80i, GM80j]. Stochastique [El 81, Aze82, Bis81c, Bis81b, Bis81o, Cai80, El 84, Éme82, LY83, Mey81b, Mey81c, Mey82b, Wei83, AY82b]. stochastique^{≫} [ZM84a]. stochastiques [Bis81h, CMS80, Éme80b, Jac81, Jac80b, Le 83a, Léa83, Mas82c, Sch80x, Sch80w, Sch82n, Yan83b, ZM84b]. Stocks [Poy83]. Stoilow [Caz83, Leh83b]. Stokes [Rau80, Ami80, Bas80, BGM^{+}80, DMPG83, Fas80, Foi80, GLCS80, Hey80a, Hey80b, Kni80, KY80, KB80, Maj84g, Mar84a, Mas80, Not82c]. Stone [Gie82t]. stopping [Bel81b, EMS82, NS81b]. stratifiable [Jun84]. stratifications [Tro82b, Ver84b, Ver84c, Ver84j]. Stratified [Ver84h]. stratipications [RF84]. strictly [GT83, Jak80a, KR83b, Mar83a]. Strings [Abh82s, Abh82r, Abh82d, Abh82c, Abh82e, Abh82k, Abh82n, Abh82p, KW82]. strip [Lac84b]. Strong [BP83, Cat81, Hia84, Los82, She82q, Wol81b, All82d, BS80a, Bau84a, BvS83, CS83a, Daf81, ES83, Hei84, JM81a, KKMW81, LM84b, Los80, MR84a, Mdz81, Mil81d, Riv84, Smi83]. Strongly [Lut82, Ran83a, Ben82a, Ber84b, BP81b, CQ81, GM80j, Jun84, Ran83b, Sho84]. Structural [Zel81e, CKMN83, CR82, Gil81c, Kle84f, Rup83, dO80a]. Structurally [Sot82l]. Structure [Bre83i, Bre83j, Dla80, Ekl83a, Kle84k, Lev80e, Rog81c, Sch83f, Ver84h, vOV81m, Art82c, Beh81, Ben82a, Bis81o, Bul84a, Chi84a, CC82d, Dor81, DR82b, Ekl83b, Fom84a, GM84b, Gie82o, Gre83, HPR84, Ise84a, Jan84, Kij80, Kop81, Kre84i, Kre84k, Kre84a, Kup80, Lau83d, Lev80h, Lev80j, Lev80r, Lev80q, Mag82, Mäk80h, Mis84b, Miy83c, Nam80b, Nor80d, Nor80f, Pet81, Que81, Ric81a, Rob83a, Rob83b, SS81a, Ver84i, Win82b]. Structures [Ama81, Die83b, Die83h, Die83i, Nar82, Que84, AL80, Bai81, BKQ83, BB83b, Bou81e, Bre83b, CB84b, Frö82b, GP83b, GP82, Gre82d, Grz84, HMT81a, HN82, Huš82a, Jad80, Kan80a, Kow80d, Kow80k, Kre84f, Pan82, Sch81c, Sla84, TV84, Wei81a]. Struktur [SP80]. Studies [BGV84, BFPS81, MS81a, MSW80c, Šig83l]. study [BF80c, CN82, Cle82, DRR81b, Dup80, How82, JS82a, LTS82, SP83, Van83a]. Sturm [App83, App83, Bro82a, Min83a]. Styria [KS84a]. sub [Sto83a]. sub-Markov [Sto83a]. subadditive [HPR80, Jaj84b]. subbilinear [Don82a]. subcanonical [CV84b]. subcategories [PR82]. Subdirectly [Fre83]. subfield [Mah84]. subfields [MS80a]. subgeneric [Hap81]. subgraph [EFG82]. subgraphs [Bil81, KR83b, KT83, KS84c, Lit82, McA80, Qui84]. subgrid [Bas80]. subgrid-scale [Bas80]. subgroup [Hum80e, Reh82, Wil81g]. Subgroups [AV83a, Bak82a, Ran83a, Arn81, Arn82m, AV83b, Cas83b, Dix81, FH84, Her83a, IS81b, Mil81a, Nam80b, Pro81, Ran83b, SH81b, Spe81a, Vas81, Wol80]. subharmonic [lPBF80, KLR^{+}84]. subharmonicity [Gam80]. subharmonische [phAD80d]. Sublinks [Hil81k]. submanifold [BHS81, CO81, KP80]. Submanifolds [Ken84b, Ban81c, Hen80c, Kas84, Kit84, Nai84, Ros84a, dC81]. submartingales [Mil81d]. submersions [BB81d]. submodules [Gie82a]. Subordinate [DR82b]. Subordinated [Maj81i]. subordinateurs [Bou84d]. subordination [Bou84e, Hir84b, PS84b]. subquadratic [Cap83]. Subrecursive [DJW84]. Subregular [Slo80i]. Subrings [Arn82p, Fai82b, Arn82b, DF82c]. subsets [Bae84, Cel84, Cho84f, Din81, KS84c, Nat84, Sch83a, Sle84, Thi81]. subsolutions [BL80a]. subspace [Dom83c, Rin84d]. subspaces [AJ83, CdS81a, CdS81d, CdS81f, CdS81j, Dom83b, Has83k, MM82a, SC81h, Van83b, Wed83]. subsquares [Sti83]. substitute [Ole84]. substitution [Arn82o]. Substitutions [Abh82t]. substructure [Sch84j]. Subsystems [BFPS81, FS81a]. subtheories [MR81a]. subtraction [Kre84b]. subtree [Han81]. subvarieties [vOV81j]. successor [ON81]. sufficiency [Tro82b]. suffisante [Der84]. Suffisantes [GB84]. suggested [Art83]. Suisse [KO81a]. Suite [Tan83h, Leg82b]. Suites [Del84a, Del81a, Kow81a]. Sullivan [Tan83c, Tan83i]. Sum [Bru81p, Goo83k, Goo83l]. Summability [Oku84g, Bri83, Oku84b, Oku84c, Oku84e, Oku84d, Oku84f, Oku84i, Oku84h]. summable [Ste83c]. Summary [DRR81g, Hop83e, Ram82a, Ram82b]. summatory [Haf81]. Summer [DP82b, Eel82, LR83a, Tur82]. summing [Kis81b, LMW80, SC81a, SC81b, SC81c, SC81d, SC81t]. Sums [BF83f, BF83i, Dau82, Arn82e, Bec82a, Bum84k, BF83h, Daf81, DZH81, Goo83e, Goo81b, Pic80a, SV82, Tei81a, Woy83]. sup [Flo80a]. sup-theorem [Flo80a]. Super [SC81s, BL80a, SC81h]. Super-reflexive [SC81s]. super-reflexivity [SC81h, SC81s]. superadditive [Akc83]. superatomic [Bon84]. supercentre [SZ81b]. supercompact [CL84a, Woo83a]. supercomputers [KO83]. superconvergence [Tho84i]. superficiel [AR80b]. Supergravity [Fer80, Des80b]. Superharmonic [Mae80i, Ana80, Kol84]. supérieur [Rea83, Tan83b]. supérieure [Mic81a]. Superinfinitesimals [Ben84a]. Superiore [Ves83]. superposition [Kah80]. superproperty [SC81o]. Superrenormalizable [Sén83b]. supersingular [Nyg83]. supersymmetric [Gal80, Luk80b]. supersymmetry [Fer80]. Supertopological [Doi84]. supgroups [BR81c]. Supplément [AY82b]. Support [LR80b, BH82d, Las81f]. supporting [Loe80]. supports [Erm80a, Kis81a]. Surface [CGO84, Hop83s, AAS83, AR80a, Col82a, Col82b, Ken84a, MP83a, MP83b, Rus84b]. Surfaces [Ber81d, DI81b, Dem80b, Dem80c, Dem80d, Dem80e, Dem80f, DPT80, Gir81b, Gir81a, Gir81c, Has83g, Hop83b, Hop83e, Hop83k, Hop83p, Hop83r, Hop83q, Jos84k, Miy81a, Ray81, ZVC80g, ZVC80h, Zie81a, Art82c, BF83a, Bem80, Ber80a, BK81b, BS84, Bui84, Buj83, Bur83b, CKMN83, CS82a, Cat83a, Cat84, sC83, Con82a, Con83, DF82c, FL84a, GP83b, Gér81a, GV83, Has83b, Has83d, Has83j, Has83l, Hay82, Hil84, Hil81g, Ish83, Jos84d, Jos84f, Jos84h, Kau81b, Koi84, Le 83b, LJ80, Loo81a, Miy81b, Miy81c, Miy81d, Oda83b, Orb84, Pin80d, Poé80, Por83, Ran80, Ris82, Sea82, Ste82a, Suw83, ZVC80d, ZVC80e, Zie81b, GIR81d]. surgery [DR80b, Goo83m, Ham82, HTW84, Kol82a]. surharmonioue [Gue80]. surharmoniques [Fug82, dlP80]. Survey [JT82, LV84, Wag80, Woy83, All81a, Ars81, Chr80a, DR80b, Foi80, Gin81, Gla81, HK81, Hol83a, Man81a, Pal84, Par82e, Pra82b, Pul83, Rea81, Reh82, SS81b, TW82a]. Surveys [Ver83b]. SVD [Van83c]. Swan [Tay81]. Swansea [TW84]. Sweden [CP84b, KR83a]. swirling [Par82e]. Sylvain [Lan80d]. Symbol [HRW84, Kac81]. Symbolic [Fra80]. symbols [AD81, Sus82a]. symétrique [Bou84a, Bre83d, Hig82b, Las80a]. symétriques [Bre83c, Vig83]. Symétrisation [Ehr83]. Symmetric [Grz82, Kow80h, KS80b, Nai84, Pra80, Pra84, Bro83, CW80, DP83a, EW83, Fal83a, FJ81, FJ84, For83b, Fuk81, Gol83b, Gre80g, HK80a, Haz81b, Hel81, IP83a, Jen83, Kow80a, Kow80b, Kow80d, Kow80e, Kow80f, Kow80g, Kow80i, Kow80k, KB80, Mal84, Nag82, Neu82b, Ohn84, Osh81, Ōsh82, Pra82a, SH81b, Sto82b, Var84, Zuc84]. symmetrical [PV81]. Symmetries [Bor83a, Mel82, PS84a]. symmetrizable [Fuk82b]. Symmetry [Bai82a, LS83b, Pog80, BW82, EN83, KT82b, Win81]. symplectic [GS83b, Hes80, Kij80, Nam80g, Wei82b]. symplectique [Lic80b]. symplectiques [Bis81e]. Symposia [MTT81]. Symposium [EL83, Fle82c, IP83b, JP81, KN80b, LST81, Nav84b, Pal83, RY81, Rao81, Rau80, Wil81b, Cho84l]. syntactiques [PP81]. syntax [DL84]. Synthesis [BF83m, Dom80, Hed80, Sjö82, Ste80a, WSF^{+}84]. System [Tan81f, BB82a, Bar82d, BH81b, Chu82c, Chu82d, CZ83b, Fit84, Foi80, Gil81c, Gro81b, KJTB80, Miy81e, Par83g, Rot84e, Sib80, Tan81k, dGdlV82]. Système [Duv83a, Air83]. Systemen [Nic82]. Systèmes [BF80k, Dem80e, Dra83i, Dub84, KS83b, BBP82, Com82h, Dem83a, DM81d, GR83b, Kum84b, Lio83, Mas82c, Pri82, Sal80, Ver81a]. Systems [AMT82, Bur80, DP82b, GM80e, Hin82, KS84a, NR80, PD82a, RY81, Rot84g, Sig83h, AK82, Ali82, AG83b, Bak83, Bar83d, Beg83, BCF82, BCF83, Bos84, BLP83b, Bro80, Cam80e, CS83a, Cap83, CM80a, Cas84a, Chu82b, CC84a, Doy81, DH83, Dui84, Düm84, Enr82, Fei84, FO82, Fom84a, Fre82a, Gal82a, GH83, GHH83, GP83c, Gek84b, Gek84d, Gek84f, Gek84g, Gér80, GO82, Hal83, HD83, Har81b, Har81a, Har81c, Har81d, Har80c, Har81q, Haz81b, Haz83, Hen81b, Hol84, HS81, HO80, HP83, Hwa81, Ivn83, Izé82a, JS82a, Kač80, Kac83, Kas83, KM82, Kaw81, Kir84, Kle84c, Köh82, Koj81, Kwo83, LZ82, Lue81, Mac81b, Mel82, Neu82a, Noc82, O’M82, OS80, Ort80, PD82c, Pap84b, PS84a, PS84c, RC81, RW84a, Ryd83, Sch82l, SM83a, STS82]. systems [Sig83i, Tan81d, Tan81e, Tan81h, Tod82, Ung82, Ver80a, Ver83c, vdD80]. Systolic [Sch83i]. Szankowski [Pis80]. Szego [ESV83b, ESV83h].

T. [Bou81a, Esn81]. table [Ano80, Ano81b, Ano81c, AY81, Bul84a, Mag82]. tableaux [Ise84a]. Tables [BMS81, Spa82f, Dor81, Tre84]. tactical [And81b]. Tail [GM84b]. tailored [Gur84]. tails [Ant83, JW82]. Talagrand [Loo81b]. Tamano [Fro84]. Tamarkin [BD83]. Tame [Dro80, Rin80b, Rin84e, Dla83, Hap81, Oko81]. Tampa [GMSV84]. tangency [Ban81c]. tangent [Eta84, Hén80a, Jac84]. tangents [HM82a]. Taniyama [Del81d, Del81e, MS81e, MS81f]. Tannakian [DM81b, DM81c]. tate [Kat81]. Tau [LO83]. Tauberian [Mad80h]. Taut [Miy84]. Taylor [Aze82, DT80, HV81b]. Tbilisi [IP83b]. Tchebycheff [MS82e]. teach [Des80b]. Technical [DAP82, FKSW81, Gru82r]. technique [Kea81]. Techniques [Cho84j, Kea83, MdJ82, Bel81b, Cle82, Duf82a, FR83, Jou83j, Rea81, Sec82, SJ84, Tho82, War82a, vH82]. Technische [Gru82r]. Teichmüller [Abi80, Gar83a, Poé80]. television [Glo83]. Temple [Kno81]. Temps [FSY82, Gun84, Bro84, Le 83a, Wei83, Yor83]. tend [LB84]. tending [Mac83]. tension [ZC84]. Tensor [Fos80, Joh83a, Pie83a, Beh81, Che80a, Der81, EF81b, EF82, Lad81, Oka84, Riz84]. tensoriel [Tig82]. tensorproducts [Ben84e]. Teoria [Mac81c]. teratology [Mur84]. term [Pin84, Sze81]. termes [Cha81b]. Terminologie [BF80g, Sch83u]. Terminology [BBCH81j, Mad80f]. terms [BB81a, Sot82b]. ternary [Bou83k, Brz81, Cus81, Mor80b]. test [Csö81]. testing [Csö81, Len81]. tests [Alb80, Deh80, Deh81]. th [Bec82a, FK83b, DM80c]. thêta [Ber84f, Bre83e, Bre83f]. Their [Gie82f, Hau83a, MSW80c, Rei83a, Adi80a, AHK82b, Bai81, Bou83k, Bro81e, Brü81r, Bur80, CC84a, Clo81a, Com83, Czy81, Dia84, Dra84, ELZ83, FR83, GM84b, GT80a, Goo83i, Gru82u, Gui82a, Gus82, Hau83b, Itô80, JvdS84a, Kac82, KKM84e, Kas84, Kir82, Kna83c, Kör83, KSS80, Kre84g, Lan81a, Mag83, Nut82, Ore82, Pra82a, Rei81, Rei83b, Rus82b, Sus82a, Tra84, Whi84]. Thématique [BMR81]. Theme [Rab82]. Theorem [Cha83g, Cha83i, ESV83b, ESV83r, ESV83p, ESV83m, ESV83n, ESV83o, ESV83q, Hop83c, Hop83s, Kre84l, Kre84m, Not82c, Sch80j, SC81h, SC81r, Sot82j, AS83, AW82, Bak84b, BG84a, Bax83, Bis81j, Blo84b, BV83f, Bre83g, Bro81c, Bro81f, Bru80, BK81c, CKMN83, CS82b, CMZ80, Cel84, Cha83p, Cho84b, Cil83, CM80b, Con80a, Con80b, Dav81b, Day81, DPV82, Dix81, Don82i, Efr82, EH83, Eng83a, FHY83, FS81d, Flo80a, Flo80b, Flo80d, Flo80f, Flo80i, Fro84, Fuk83, GŁ84b, Gór81, GMS84, Gra82c, Gre84, Gru82t, GS83b, GS83e, HG82a, Han84, Her81, Her83a, Hey81, HMP82, Ise84b, Jac80a, JvdS84j, Kru84b, Lag81, Led84a, Leh83a, Lem80, Lev80v, Loo81a, Mai81, Maj84b, Mar83c, Mc 81, McG84, Mil84b, Mit84, Mok83, Mos81a, Mos81b, Mun80, Nus81, Pau80, Pop84, PR84d, Pre83a]. theorem [Rab82, Rin80a, Roh81, Rot84d, Sag82, Sam80, Sar83, SC81q, SC81d, SC81n, SC81s, She81b, Šig83l, Smi83, VVV81, Wer81j, Wer81p, Wil81f, Yas81, Zie81b, dHR81, dS82, vOV81i, JM82a, SC81s, Wil81g]. Théorème [Ber81d, Bre83k, LT84b, Mal80a, Năs80, Rab83, Abd83, Bar80c, Bou84b, Bre83g, Cha83t, DZM83, Éli82, Éme81, Fer81a, Fey84c, Gui84b, Kah80, Mal82b, Mal81b, Mar81b, Mau82, Yan83a]. Théorèmes [AT82e, AT82f, BBM82f, FM83, Le 82b, Led83, LR83d, Sko80b, Web84a]. Theorems [Don82e, Ekl83a, KS83c, Maj81b, Par83f, She82l, Alm80, AS81a, App83, Ban81c, Bro81a, BC83, CP83b, Chr80a, Clo81a, Col83, Don82h, Ekl83b, Elb82, Ers80, FK83a, FL84a, Gab81a, Gau81, Gel81b, Gie82t, Gin81, GHZ83, Gne83a, Gne83b, Had82, Has83a, HJ81, Hei83, Hen84b, Ill83, Iof80, Jaj84a, Jaj84b, Jak83, Joh82, Jos84h, Jos84m, ŁK83a, Led83, LR80b, Luk81, Mad80b, Mad80h, Maj84d, Maj84e, Maj81d, Mau80, Meh83b, Mot81, Pet80, Raţ80, Ros82, Rot84j, RR80, Ryc84, She81c, Smi81, Sot82i, Val84a, Wag80, AT82e]. Theoretic [BR80a, Bel81a, BT82g, Bie81, Bor84a, Cho84i, FS81b, Gat84, Kol81a, Liu84, Sch83g, SZ81b, Sen84, SW81, Val84b, Wil82a]. Theoretical [BFPS81, Eel82, Rau80, Gil81b, JS82a, Jak82, Poh81b, STS82]. theoretically [DG84b]. Theorie [DK82b, Lja81, Sch82c, AT82g, BF80a, BMR81, BBM82g, Car80d, Ceg81, El 84, GWL81, Hür81, Mey81e, Tal80, Yor81, Anc84a, BBM82b, Cha81b, Cho83d, Fey84c, Hir84a, Kas81, Léa83, LLP80, MR81b, Mey81f, Mey81d, Mic81a, Spr83, Ten84, Vus80, HM80, HM82b, HM84a, MP84b]. Theories [CZ83a, Cer84a, MdJ82, Bai82b, Bal81a, BV83f, Che84a, Des82, FS81b, Fel80, Hay81, HN81b, HPR84, Jež83, JW83, Lue83, Mar82a, Mdz81, Mos81a, Ne’82, Pas83, Sch84j, Sén83b, Ste83d, Ues81, War82a, Wel83, LC80]. Theory [AFG84, All82b, Ama81, BR80c, BDM83, Ben84b, BFF80, BMR81, BS83f, BBCH81f, BV83a, BOR^{+}84b, BKS83, CW82, Cha83j, Cho84j, CCCN84, Con80c, Cuy84d, Den82a, Den82b, DG80d, DG80c, DP82b, DLR81, Eas82b, Ekl83a, Erd81b, FF81, Fle82c, FZ80c, FG83, FS81e, Ghe83, GW81, GLM83, Har81n, Has83l, HN81a, Hen81e, HJW83, Hub83a, IP83b, Jag84, Jam84, JvdS84g, JP81, JTW82, JV82, KPT82, Kan83o, KMS83, KN84, Kno81, KY84, Kol83, Köl80, KMS82, KMS84, Krá80f, LR83a, Mac81c, MS81a, Nam80h, NR80, PWW80, PPS83, Rao81, Rup84a, Sal84b, She81d, Sig83h, Sto82a, SW84, Tan81f, Wer81o, Wer80, vO80, vOV82, Abi80, Adi80a, Ali84, Att84, Aus82, Bak84a, Bak84c, Bau84a, BCS81, Bec84, BDL80, Bec82b, Bel81a, Bel83b, BT82a, Ber80a, Ber82a, BBM82g, Bet82, Beu81]. theory [BT82h, BBCH81g, BJ80, Bon82a, BV83b, BV83i, BG84b, Bos84, Brä82b, BC83, CP84a, Cas83b, Cat81, Cha81a, Chi83, CP83b, Clo81b, Col84, CZ83b, Cun83, Cuy84a, DW81, De 82b, DK82c, Dri82, EG81, Eas82a, Eas83, EP84, Ekl83b, Enr83, Erd81a, Erd82, EH84, FH81, Fer80, FHM82, FN82, Fli82d, Fos81, Fra83c, GM84a, Gie82b, Gie83b, Gil81a, Gil81d, Gir83b, Gir82, GGHJG81, GRW81, Goo84, Gre80g, Gre81c, Gro80i, Gui80c, Gut83a, Hac82a, Har81e, Har81g, Haw84, Haz83, Hej84, Hel80b, HM83c, Her80a, Hil82b, Hof83, Hub83b, Huš82a, Igu82, Ise81, Ize82b, Jak82, JS80, JvdS84b, JL80, Jou83d, Kac83, KP82, KWP84, Kai80a, Kai80b, KKM84f, Kal82, Kat83, Kaw81, Kem83, Kij80, Kno80, Koc82, Kol81a, KW82, Kra82a, KR83c, KZ83b, LV84, LT84a, ŁK83a]. theory [Laz84, Len84, Liu80, LS82b, Mac81d, Mag80, Man80, Man84, Mar84a, Mar82b, Mdz81, MK83, Mik82b, Mil83b, Mol83, MS83d, Mos84a, Mun84, Nam80a, Nam80d, Ne’82, ON81, Ore82, Osg84, Ott81, Pee84, Per82a, Ple81, PP82b, PR84f, Pre84i, Ram82a, Ram82b, Rei82c, Rob81a, Rob83a, Röd84, Rog81a, RW84b, Ros83a, Rup83, Ruy80, Sar84, Šče81a, Sch84a, Sec82, Seg82b, She81a, She81c, She82r, Shi83a, Sig83i, Sig83k, SSM^{+}84, Slo80a, Sna82, Sna84, Sou81a, Sou82, SK80, Sto83b, Sto82b, SW83, Sus82a, Sus82b, TT81, Tam82a, Tan81b, Tan81g, Tan81k, Tho81b, Top83, Tra82a, TN81, Ven81, Vog84, Vol82b, Wal82a, Wal84, Whi83, Wil80a, Win81, Wol82b, Ytr84, Zel81e, ZVC80d, ZVC80e, Zou83, Zvy84, dB82, vH82, vL83a, vOV81k, vdD80]. theory [vdB83, AT82g, EdJ82]. thermal [CRSSC84]. thermodynamic [Eri84a]. Thermodynamics [Gro82, Dre83]. thermoelastodynamics [dM83]. these [Ara84]. Thesis [San81]. theta [Bre83g, Pet82, Bre83g]. theta-functions [Pet82]. thick [Vol81]. Thin [Top80, Vol81, Kak81, VLA83]. Third [All82b, AL81, AL82, Bec81, BS83e, Dol83b, Hen82a, Gie82u, Nam80f]. Thom [Ver84d, Ver84i, Ver84k]. those [HN80b]. thought [Ser84]. Thrall [Rin80a]. Three [CGO84, Joc81, Art84a, Bro81b, sC83, Ens83, FL84a, Her80b, Hur82c, KK83, Lac81, Mor80b, Now82, Sie81b, Whi84, dO80a]. three-dimensional [sC83, Whi84, dO80a]. three-manifolds [Lac81]. Threefolds [Con82b, Mor82a, Cle82, Con82a, Con83, Lev82b, Mur82a, Uen82, Wil82b]. threshold [GM83a]. thresholds [Sah83]. Thue [Lag81, Mue84]. Thurston [Gro81a, Mor81b, Mun80, Poé80, Sul81b]. Tight [Ada82, CR81]. tightness [McC82]. Tilted [Bon81, HR81]. Tilting [Rin84f]. Time [Hej84, McG84, Tho84p, Bar81b, BP83, Bel81b, Düm84, GLCS80, HR82, Hol81, Jac83a, NT82, Per82b, PS83b, Rév81, Tan81e, Tan81h, vdH82]. Time-dependent [Hej84, GLCS80, HR82]. time-invariant [Tan81e, Tan81h]. time-periodic [Hol81]. times [EMS82, Le 84a, Mue82, Rog84a, Sha80]. tissus [Bea80]. Titchmarsh [Atk82]. TLC [SC81d]. Toda [Kos82]. Toeplitz [AAK^{+}84, AC82, CS82d, CS83b, Mad80c]. together [Leh83b]. Tokyo [RS83a]. tomography [Ste81b]. tool [Est80, LG81c]. Topics [CP84b, GP83a, Hen82b, Hop83g, Hop83o, Kre84e, LMGC82, Sal83, Win82a, dB82, Bou83j, Jef81, JT82, LG81e, Nor80i, PP82b, Tur82]. topologica [Lew82]. Topological [All81b, Bai82b, BH81a, BG84b, FM84, FZ80c, Gie82f, HJW83, Kha82a, Kha82h, Kha82i, Kha82j, Mag80, Skr84, Sto80i, Zah84, dM81a, Ben82a, BGMY80, Brü82b, Doi84, DM82, DRR81d, Fon84, Gra82c, Gre82d, Guc80, Had82, Hel80b, Hof82b, Kat80b, Kha82d, Kha82g, Lew82, LM84b, Man81c, Mas82a, Nor80b, Ruz82, Sch81c, SY84, ZVC80e]. Topologically [BS84, Edw80]. topologie [Har80a]. topologies [Bar80f, Kum81, Tra84]. topologique [Kla83a, Lel82a]. topologiques [Gal82c]. topologischer [BKK81]. Topology [Ban82, Boc82, FM84, KPT82, MO84, MS81a, Ern81b, Ers80, Flo80c, Flo80i, GM82d, Hof83, Kau81b, Kni84, Kur83a, Ran80, Sot82k, Str83b, Wec81, FM84, KN80b]. Topos [BV83a, LS81a, Mak81]. Toposes [Lam82, Hay81, Joh81b, Ues81]. Torelli [Cil83, Loo81a]. tores [Cam80a]. Tori [Zie81g, Fri80, Ioo80]. Toriques [Tei81b, Bry80]. Toroidal [Nam80i, Nam80j, Har83]. Torsion [Alb83a, Arn82f, Bir83a, Bir83b, Blo81a, Coh82, Lad83a, Lue83, Met83, Ran83a, Alb83b, Bea82a, BM81b, BR81c, Cha83k, HHB84, Hus81a, Jon82, Kas82b, Lad83b, Lev80d, Lev80w, Mon80g, Pro81, Ran83b, Van83a]. Torsion-Free [Ran83a, Alb83a, Met83, Alb83b, BR81c, Jon82, Ran83b, Van83a]. Toruńczyk [Edw80]. torus [BBŚ82, Kre84j, Nam80h, Oli80, Ros81a]. Total [Hop83s, Mic83, Wat84b, Jür83b]. Totally [Bré81a, GvdP80k, BB81d, Hur82c, Onn82]. totient [MS82d]. touch [vOV81d]. Toulouse [AGT84]. tourists [Bae82]. tournaments [Rei84b].

Tournantes [Ros84b]. Tous [Dra83h]. towers [Bow82]. Trace [BF83n, Fli82f, Goo83h, Hej83c, Mon82, Par82c, Art83, Ber82b, CC82c, Fli82e, Fli82g, Hej83a, Hej83f, Hej83g, OW83]. Traces [Ioc84i, Ioc84h, dBT80, Fli82c, Mon80h]. traditional [AB83]. Trägheitsschiefkörper [Ost80]. trajectoires [Le 84b]. trajectories [Aub84, Har81c, Mar81e]. Transcendent [CM82a]. Transcendental [Oda83b]. transfer [Pre84k, Tan81i, Wil81g]. transfer-principle [Pre84k]. transfinite [Mae81, MS82e]. Transform [Eas82b, BH83, CP81, GL82d, HH81a, Hel81, Kuo80, McG82, Obe83, SS83, Sam82a, Wer81k]. transformation [Aub82, Bea83a, DK80a, Eri83, Jef80, Kni83, PS84b, TW82b]. Transformations [Mey84, Ahl80, Bar83d, Ber81b, BvI84, Chr80b, Dem83b, DM80b, DM80c, Fra83a, Lan81b, Mar82b]. transformées [Bak84d]. transforms [Bia81, Bum84j, Bur81, GS82d, Kis81a, McA83j, Mha84, Sch84k]. Transience [Die83j, Die83k, Gui84b]. Transientes [Die83i]. transition [Ers82, Ioo84, Sto80i]. transitive [Als81, Gar82b, Küh80]. Translated [ZVC80h]. Translation [Blo84b, ZVC80h, BL80b, Dom83b, Dom83c, FMW82, Gli84, HH81a, Sot82i]. translation-invariant [Dom83b, Dom83c]. Transmission [Glo83, vdB83]. transonic [DH82, SJ82]. transport [GL84a, Hej84, Kan80b]. Transposition [Die83e]. Transversal [CKLR82, Tak80]. transversalité [HM82a]. Transversally [Sun82]. Transversals [Joh81a]. transversely [Gir84b]. Travaux [Ber80e, Oes80, Poé80, Sul81b, GPV82, Hir84a, Mél80, Ten84]. travelling [BB82a]. treatment [Bak82b, Gea82, Whi84]. Tree [BM83, She82p, BR81b, Gup84, Han81, Koś83c, Mil80a, SC81s]. Trees [Cho84k, Dic80f, Odi81, DPV82, Dev81, Ead81, FR80, FR81, Ger84, Hen80b, Kec81a, Top80]. treillis [Dla80, JG81]. treizième [Kah80]. Trends [Ben84b, DG80c, Ver83b, Acc84, Hof83]. Trennungsaxiom [Ben80c]. triangles [RS81]. Triangulability [Ver84h]. Triangulation [Ver84j, Ver84k]. tribus [Mey80a]. Tribute [EL83]. tridiagonalization [CW80]. tridimensionnelle [CR80a]. Trieste [Eel82]. trigonometric [BL81, CRSSC83]. triple [LZ82, RC81]. Triplets [Ser80]. Trivial [IW80, BS84, Fed84, Gie82n, SC81o, She82i]. trois [Fos81]. tronqué [Har80a]. tronqués [BD80a]. truncated [Sti81]. Truncation [Gil82b]. trust [Gay84]. trust-region [Gay84]. Tsetlin [Die83k]. Tsirelson [LY83]. tube [Gra81a]. Tubular [Rin84g, Rin83, Rin84a, Rin84f]. Tufts [Bec81]. Tulane [KKS82]. Tulcea [Ers82]. tunnelling [GL82c]. Turbulence [RY81, Ioo84, Jos80, LC80]. turbulent [Wac80]. turning [Kam82, Lor82]. Tutte [Daw84]. TV [Itô84d]. TV-inégalité [Itô84d]. Twelfth [Fle82c]. Twelve [SC81h]. twin [JR82]. twin-convergence [JR82]. twist [BE80, Dev80]. Twisted [Cam80g, CM82b, Fli82g, Fri83, Hau80b]. Twistor [Eas82b, War82a, Eas82c, Per82a, Wel83, DP82b]. twistor-geometric [Wel83]. twistors [Gin82]. Two [Acz81b, Con83, Edg83, FL84a, Gar80, Gly83, HP80, Hen81m, Saw82, Sig83k, Smi81, AHE81a, AHE81b, ARMS82, AC83, Axe82b, BH84, Bax81, BMS81, Car80e, CN82, Che81a, Chu82b, Chu82c, Doz81b, Enr83, GO82, Guc81, Gut83b, Her84a, Hin82, HN81b, How81a, HM82d, Hur82c, Iit83, Ing80, JvdS84f, Mag82, MM82a, Mal80c, MR81a, Mar84a, McA80, NSE82, ON81, Pic80b, Ruh83, Sei81, Sie81b, Sul81a, vdB83, Gru82u]. two-conductor [vdB83]. two-dimensional [Chu82b, Chu82c, HM82d, MR81a, ON81, Gru82u]. two-distance [Sei81]. two-graphs [BMS81]. two-level [Axe82b, MM82a]. two-manifolds [Gut83b]. two-parameter [Che81a, Doz81b]. two-point [Bax81, Mag82, McA80, NSE82]. two-sample [Ing80]. two-sided [Ruh83]. Two-space [Sig83k]. two-stage [GO82]. Typ [Sch84h]. Type [Met83, Mut83a, Mut83b, Sak81, SC81t, Axe82b, Bak82b, Bas82, BD83, Bök84, Bos80, BvI84, BO84, Bru80, BV83j, Cat83a, Cha84d, Chu82b, CS83b, CK84, DRS80, Dra81, DSS83, DRR81f, DRR81e, Fer81d, Fle82a, Fre82b, Gay80, Gel81a, Ghi83, Glo83, Hap81, Has83b, Has83d, Has83j, How81a, Hur82a, Ili84, Ise84b, Ish83, Itô84a, Kle84d, Kna83a, Kow80e, KS80b, Leg82d, LP84a, LR83d, Mas81, Mei80, Mek81, Men80b, Nav84a, Nor80f, Par83k, Pis82, Pra83, Rog80, RS84, Sab84, Sch82b, SC81f, SC81g, SC81h, Shi81a, Šig83l, Sot82d, Szu83, Wil82b, vDS83, vdK81, SC81h]. Type-Related [Met83]. Typen [Sch84e]. Types [Arn82q, Bou80, Ami80, Bia81, Bon84, Del81h, GM84a, Gla81, HN81b, Ivn83, Jon81, Jur81, Kop81, Mac81d, Mak81, MN81, Pil81c, Pre83g, Trn84]. typical [Gre84].

U.I.A [vO80, vOV82]. U.S.A [CC82f]. Überlagerungsflächen [phAD80a]. UCLA [KMM81, KMM83a]. Ueno [Esn81]. Uhlenbeck [Bak82c, HS81, Mey82d]. Ultra [Mal82d, Tah82]. Ultra-produits [Mal82d, Tah82]. ultrafilters [Bae84]. Ultrapowers [SC81j]. ultraproducts [Kür83b]. ultrasonic [Ste81b]. Unabhängigkeit [End84]. unbounded [Fle81, Gau83, HLSW80, Kan83d, Pen84, PS84c]. uncertainty [Tan81d]. Unconditional [HNP81, HT81, Kva80]. uncoupling [dM83]. Undecidability [CR84b, ON81]. Understanding [Vog83]. undirected [Hav83]. unfoldings [BvS83]. unicellular [GT83]. unicité [Itô84c, Net82b]. unidimensionnel [BJ83b]. unidimensionnelles [Le 83a, Yan83b]. unified [Far82, KW84, Ruy80, Sam80]. unifiée [LLP80]. Uniform [CS81b, DH83, JR82, Nar84g, Ver84a, Wim80, AS83, CS81a, Dau82, Fro82c, Fro84, Gre82d, Huš82a, Jaj84a, Mor81a, Mor84a]. Uniforme [FM83, AGY80, GL82a]. uniformization [Pop84]. Uniformly [MZ83, Hei84, HV81b, Sch82a]. unimodal [BT81]. unimodality [Wol81a]. Union [Rau80]. unions [DR81, JM83]. unipotent [DO82, Spa82b]. Unipotentes [Spa82a, Spa82b]. unipotents [LP81]. unique [Brü82b, Pri84]. uniquely [Bil80, Per81a]. Uniqueness [DHH^{+}84, Jos84m, KI84, Rot84d, Wer81r, All82d, BP83, Bax83, CL84b, DP83b, ESV83h, Gal83a, HJ81, Hen81d, Loy83, Mar83a, PD82b, Pap84b, Per82b, Riv84, Woo83a]. unirationality [Tan81c]. unirationnelles [Bea83b]. uniruledness [MM83]. Uniserial [AR80c, Hau83a, Hau83b]. unit [AHL83, AF84, Gie82h, Goo80, Hay83, Pet83]. unit-regular [Goo80]. unitals [Bro81e]. unitarily [AHL81b]. Unitarisierbarkeit [Sch83t]. Unitary [Enr83, Kna83b, Sna84, Spe81a, Ara83, Ber83c, GPS83, KS82, KS83c, Kre84f, Sil83, Vog83]. United [GMSV84]. Unités [Oja82]. Units [Mäk80f, San81, AL80, Mag81, Mil81a]. unity [Kov83]. Univalence [Par83f, Par83k, Par83a, Par83b]. Univalent [Tam82f, Par83e]. Universal [Her82d, Kra80i, Bel83b, Bru81q, Gab81b, GvdP80k, Grz82, Mal82a, Wil82c, FG83]. universel [Bre83k]. universellen [phAD80a]. universelles [Yam82]. universes [TT81]. Universidade [Mac81c]. Universität [AJ81]. Université [BLS80, Coe82, CTCMR82]. University [AV80, BBC^{+}83, BH81a, Ban82, Bec81, Ben80a, BOWS82, BS83f, Car82, CW82, CC82f, CCCN84, DG80d, DG80c, DAP82, Dol83b, Eri84b, FKSW81, FS81e, GLM83, HLR83, HKLR83, HJLR84, Hin82, Hop83d, HN83, KM80, KKS82, Kno81, KN80b, LSS81, LW81, LST81, MS81a, McA81, Men80a, NR80, PPV80, RY81, Rau80, RW82, Ric81c, RSW80, vO80, vOV82]. unknown [Le 84a]. unoriented [Kre84f]. unprovable [Mil80a]. unramified [Sot82a]. Unremovable [AF83]. unsolvability [Sho81]. Unstable [Lam84]. unsteady [Fas80, KSS80]. unsymmetric [PM83]. Untergruppen [Ros84c, Sch83k]. Untersuchung [Gru82s]. unusual [Jam83]. unverzweigte [Sch84i]. update [Geo81c, Wag80]. uples [Kla83a]. upper [CS84b]. upperhalf [Nam80g]. USA [Bec81, Dol83b, Fle82c, GLM83, LSS81]. Use [Now82, Ale83a, CG83b, May80, Rei82c, Sec82, SSTW82, Tho82]. useful [JW82]. Uses [Mag81, Lyn84, Ple82]. using [Com83, CW80, GLCS80, HPR80, Woo83b]. USSR [IP83b, KZ83a]. Utilisation [Cor84a]. Utrecht [LST81].

V [Dem80f, Sto80c]. V. [GPV82, SM83b, dM80a]. V.F. [Éme80a]. vacuum [Bin83]. vague [BC82b]. val [MT80a]. valent [Gök83]. valeur [BY80]. Valeurs [FM83, Gav83, dMG83, Dub82a, Geb83, Gui80b, Sch80r, Sch80t, Sch80x, Sch80w, Sch80u, dM80a]. valid [Gel81b]. validity [Par83j]. valuated [GWW81]. valuation [Bru82a, Kat82]. Valuations [Bec82b, CMH80, PR84h]. Value [CdS81g, Gek84c, Hay83, Ric83, Sto82a, Bax81, Bax83, BH81b, Beg83, BE83b, Bjö84, Böh82, CD80, Col80a, Dah82, Enr82, Fad83, FO82, FK83b, Gek84a, GH82, Gus83, Hab80a, Izé82a, Kaw82, Krá80a, KS83e, Lor82, LS82b, MW84, Mol83, NSE82, O’M83, Ove82a, Skr84, Ste82b, Ste83f, gS83a, Tak84, LR83a]. Valued [Sch83e, Ant81, BM80, Cha80b, Čob84, CS83b, Det83, Don82d, Don82i, Ebe83, Goo81b, GMJ84, GMS84, Gre83, GL82d, Hei83, Kot81, Kue81b, Leś84, Mak80a, Mor84a, Nam81, Sch83d, TT81, Web84b, Wei84, dF82b]. Values [Lic84b, AW83, Fle82a, HG82b, HM84b, Kla81, LS83a, Sig83a, Sig83b]. Vanishing [Kas83, Lau83c, Hil81l, Maj84d, Sie80a, dS82]. Varchenko [Mer80b]. Varenna [Con82b, dV84]. Variable [Cha84a, AD81, Bru82c, Koj81, Las81c, Las81b, Las81e, Las81d, Sem82a, VVV81]. Variables [Las81d, Nor80a, Ant81, Ber84f, Est84, Fei84, Gne83a, Goo81b, HB84, Hei83, Hia84, HTW81, IJ84, Iwa83, Jeu80b, Koo80, Kue81b, LeP81, Mas82b, Mor80a, Rac83, Rea83, Sem82b, Ton83, Vla82]. Variantes [CT82, Die83k]. Variation [MB82b, MMP80, Ble83, Fit84, KP80, MMS81, Miy83c, Nua81, MB82b]. Variational [Att84, DH80, MSP83, MW84, RW84a, BF84, Eke83, Fom84b, Her80a, Hor82a, Jos84a, ŁK83e, Moe82, Sbo83, Tak84]. variationnel [Bis81d]. variationnels [Bis81m]. variations [Bon82b, Com82f, DD83, Mar83a, Rab82]. Variationsrechnung [Gru82f]. Varietäten [Sch84b, Sch84d, Sch84c]. variété [Abd83, Bea83a, Bis81a, DG83, GŁ82b, HYZ83, Kis82, LT84b, Lic80b, Lie80, Sch80o, Sch80s, Sch80x, Sch80w, Sch80u, Sch82n, Spa82e]. Variétés [Bea83b, Com82h, HZ84b, Sch80v, Tei81b, Tei82, BB81c, Bis81e, Bry80, Cer81, DZM83, Esn81, Gal81, GM82b, GM82c, HM82a, HM82c, Lem81b, Sch80t, Sul81b]. Varieties [Del81b, Del81c, DMOyS81c, MS81c, MS81d, Art82a, Bar82c, BT82b, DP83a, Dol82, For84b, Fre83, GvdP80h, Hol83a, Ion84, Miy83b, Nam80a, Nyg83, Ore83, Pan82, Sch80v, VL83b, Ver80b, vOV81a]. variety [Ban83a, Ber83a, Bur80, DL81, Gat84, GvdP80c, KT84, Mum83, vOV81c]. Various [Han81, Ami80, Mai81, Man83]. Varopoulos [Éme81]. varying [KSS80]. Vasil’ev [Her82a]. Vasil’eva [SM83b]. vecteurs [Kla83a, Mar81i]. Vector [Aky82, And82, Car82, Con80h, CS83b, Gie82f, Gra82b, Kha82a, Kos80a, Kos81e, Nar82, Sch83e, SW84, Wer80, AG83a, BF80c, BH84, Bay81b, BT82a, Bjo82, Bre84, BS84, Bro81a, Bro81b, BvS83, But82, Cha83b, Cha80b, Col83, CM81, CB84b, Dem82b, DU83, Don82b, Don82c, Don82i, DM82, GMJ84, GMS84, GY82, Had82, Haz84b, Ila80, Ivn83, Jaw81, Kha82d, Kha82e, Kha82g, Kha82j, Kré81a, KT84, Laz84, Leu80, Lie82, MP81, Mar82b, Mar83g, MW84, Oka84, Pau83, Rin80a, Rin80b, Sch83d, Sch83h, Sch80n, Ver84e, Web84b, dF82b]. vector-space [Rin80a]. Vector-Valued [Sch83e, Cha80b, Don82i, Sch83d, Web84b]. Vectorbundles [Tra82b, Haz81b]. vectorfields [Haz81a, Tak80]. vectorial [AC82, Gar82a]. vectoriel [Lel82a]. vectorielle [Jac80b, Sch80p]. vectorielles [Sch80p, Sch80r]. Vectoriels [BF80c, Ray81, Rou81a, Rou81b, DS80, Dem82b, Ghi83, Le 83b]. vectors [BŻ80, Ros82, Ryc84, Woy83]. Vekua [Lan80a]. verallgemeinerte [Gru82l, Jür83b]. verallgemeinerten [Gru82a]. Verallgemeinerter [Gru82t]. verbände [BKK81]. Verchik [Gui80a]. Verfahren [Wei82a]. Verfeinerungs [BKK81]. Verfeinerungs- [BKK81]. Vergne [EP81]. verifiant [Hig82a, Arq84, Itô84b]. vérifiée [Vig80f]. Verma [GM82a]. Vernünftige [Dra80e]. Veronese [Gat84]. vers [LB84]. Verschwinden [Sch82j]. Version [Hej83f, Hej83g, BF82, Hej83a, Kif80, Pre83a, vOV81i]. versions [Leś84]. versus [NT82]. vertex [Als81, CH84, Koś83c]. vertex-transitive [Als81]. Vertices [Lor81a, Fic80]. Very [Cor83b, Cor83c, Ebe83, tR84]. via [Cas82b, DJW84, GvdP80i, Gup84, Mil81b]. viability [Aub84]. viable [Aub84]. vibration [CG83b]. vibrations [vdB83]. Viehweg [Esn81]. Vietoris [Wei81a]. view [Bai82b, Eri84a, GS83b, LG81e]. viewpoint [CK84, Dav81a, Hor82b, Jia81, RS81]. VII [Gal89, Hey84, RSW80]. VIII [Hen80d, McA81]. Villa [AFG84]. Villetaneuse [BLS80]. Vinberg [HPR80]. virtually [Ros81b]. viscoelastic [Ber84e]. viscosity [vA82]. viscous [Bab80, BGM^{+}80, CN82, DMPG83, HM82d, KSS80, Val84c]. Visibility [Hof81a]. Vitali [FM83]. Vlasov [Neu84b]. voisinage [Bar80b, Har80a]. vol [Ano80, Ano81b, Ano81c]. Vollständig [Dra80a, Dra80b, Kne80]. Volterra [Bak82b, Bar83a, Clé84, Hön80, Hön82, Kan83o, Kan83p, Min83h, Rot84k, Wil83b]. Volume [Hej83c, Hop83f, Bar80c, Bro81a, Kif80, Kow81b, Mun80, Pet80]. vom [Sch84h]. Vorbereitende [Sch82i]. vorgegebener [Dit80]. Voronoi [Nam80a]. Vortex [Mar84a]. Vortrag [Dra80d]. Vorwort [DK80c]. Vries [Fre82b]. vs [Fri80, Nor80c, SC81c, SC81o].

W. [Del81f, Ioc84a]. Wadge [Kec81b, Lou83a, Tan81a]. Waldspurger [PS83a]. Walk [HN83]. walks [CLM84, Ger84, Pap83]. Walsh [Tre84]. Wandering [Har81p]. waring [Ang81g, Nat81, Ang81d]. Warwick [RY81]. Washington [Oes80]. Watanabe [Kis84b]. Wave [DGZ84, Hel84b, Her82c, Mey82f, Bro80, Har81h, Har81j, Jef81, KJTB80, MB82a, Men82, Pan82, Seg83]. Waves [Kir84, Ste83e, BB82a, Col82c, Her80a, KM82, LM84a]. Weak [BR81a, Clo80, Dac82e, Din82, JM81a, Krá80h, Rüs80, SS81a, She82k, BD83, Ber81b, Daf81, Dat83, Ebe83, Flo80c, Flo80e, Gra80, GM82f, GM83b, Hof82a, HT81, Kur84, Mus83a, Sjö82, Zak81, vdD80]. weakest [SC81o]. weakly [Bru84, Kir84, Kre81b, Riz84, Flo80j]. weakly-coupled [Kir84]. weaver [Hos82]. Weaving [Kau80]. web [Mas83a]. wedding [MS83d]. Wege [DK82b]. Weierstraß [Gie82t]. weight [EFG82, HS84, Hum80f, KT82b, RC82, Saw82]. Weighted [Abh82u, BH83, Dol82, Gra83a, Grø83, Str84f, dF82b, AC82, Bad83, Bar83d, Bie81, Blo83, CS82d, CN84b, Daf81, Dom83b, Dom83c, Rob84a, Rüs80, Ste82a]. weighting [Van83c]. Weights [De 82b, Gre80h, Bru82c, Jou83a]. Weil [Hir83, Lip83]. Weingarten [Hop83b]. Weinstein [DM81d]. Weissen [IvW82]. Well [Pre83k, Vol84, JM83, Tri84]. Well-behaved [Pre83k]. well-ordered [JM83]. Well-posedness [Vol84, Tri84]. were [Ara84]. Werte [End84]. Werten [Sch83p]. Wertverteilung [NC80]. Wertverteilungsprobleme [phAD80e]. Wertverteilungstheorie [phAD80b]. Weyl [Atk82, Brä82b, Gra81a, HP84j, Kle84d, Lus83, Sto83b]. where [Gil82b, JM84]. Which [Hau83a, AP81b, AG83b, BS80b, BP84b, BN80, CMPY83, DG84b, Fed84, GT83, Hau83b, Ila80, Jos84g, Kow81b, NiW84, Sot82d, Wil81a]. while [GM84a]. White [SH83, SM83a]. Whitehead [Cha83k, Höp82, Kea82, Mag81, Tan83b]. Whitehead-diagrams [Höp82]. Whitney [Tei82]. Whittaker [Bru81o, Bum84d, Bum84i, Bum84j, GvdP80e]. whose [Con82a, Con83, De 83, Mor82a]. Wick [Maj81j]. Widom [Has83b, Has83d, Has83j]. Wiener [Bak84c, BG84a, Car83b, Fuk84, Kuo80, Kus82, Ław82, Leś84, LMRW82a, LMRW82b, Maj81c, Maj81b, Maj81f, Par82d, RW84b, Wer81s, Wil82a]. Wiener-Itô [Maj81c]. wild [Dro80]. Wiles [Coa81]. Williams [Len81, Mey82a, Mil84b, Rog81b]. Wimereux [LS82a]. wise [All82d]. within [Kre84g]. without [BMP80, Ers80, Föl81a, Hem82a, Her83a, Loh83, New81]. Witt [Ill83, Kre84n, Par82a, War82k, War82j, War82l, War82m]. WKB [Par82c]. Wold [Nie80]. word [Adi80b, vdK82]. wordchain [Tam82a]. words [Adi80a]. Work [PS83a, vOV81n, Coa81, Fuc81a, LT84a, Leh83b]. Workshop [AFG84, AL82, BH81a, DG80c, Eri84b, Fuk82a, GP83a, Hen82a, HN83, JTW82, Men80a, PR83, AD83b, FG84]. Workshops [Hin82]. world [BMKQ83]. Wu [Hor80, Lin81]. Wu-like [Lin81]. Würzburg [KS83d].

X [Cas83a, Hen82c, Lel82a, Hel81]. X-ray [Hel81]. XI [FMSW83]. XII [Hen84a]. XIII [BLS80]. XIV [AY80]. XV [AY81, EH84]. XVI [AY82a, AY82b]. XVII [AY83]. XVIII [AY84].

Y. [Esn81]. Yang [Aso84, Bou82b, Bou82c, Fri82, Her82b, Hor80, Tra82a, Tra82b]. Yau [Mor81b]. Year [BS83f, HLR83, HKLR83, HJLR84, LSS81]. Yoneda [Ani84]. York [CC82f, CCCN84, Hop83d, Kom84b]. Young [Kim81]. Yugoslavia [BKK82, MS81a]. Yvette [CC82f].

Zariski [Hén80a, NV82, Pop84, Ran80, Sot82k]. Zeitgeordnete [IvW82]. zero [CKK83, ESV83i, ESV83a, Gre81a, Jan82, Jun84, Mac83, MW81, Nat81, Sul81a, Vek84, LB84]. zero-dimensional [Jun84, Vek84]. zero-one [Jan82]. zero-tending [Mac83]. Zeros [Bul84b, ESV83u, CKLR82, Gon81, Mue81, Neu82a, Pin84, Cha84d, Rab83]. Zeta [BR81d, Chr83e, Vor82, Chr83d, Gon81, Kno83, Lic84b, Mue81, Pin84]. Zeta- [Chr83e, Chr83d]. zeta-function [Gon81, Pin84]. Zeta-functions [BR81d, Lic84b]. zetafunctions [Chr83a]. ZFC [Woo83b]. Zsigmondy [Lün81]. zügellos [Bon80]. zugeordnet [Sch83n]. Zum [Dra80d, Sch82j]. zur [Sch82c, Wei82a, DK82b, End84]. Zürich [HP84h]. zusammenhängenden [phAD80e, NC80]. zusammenhängender [HP84e]. Zweidimensionale [Gru82u]. Zweifach [Dra80b]. zweite [phAD80b]. Zygmund [Cór80a, CS83b, HT81, Jou83f, Jou83c, Jou83e, Jou83d, Jou83j, Ric82]. Zykellose [Bon80].

References

Atteia:1983:ICP

[AA83] M. Atteia and J. Audounet. Inf-compact potentials and Banachic kernels. Lecture Notes in Mathematics, 991:7–27, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061554/.

Akiyama:1984:CF

[AA84] Jin Akiyama and David Avis. Cube factors. Lecture Notes in Mathematics, 1073:93–99, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073108/.

Adamyan:1984:HTO

[AAK^{+}84] V. M. Adamyan, D. Z. Arov, M. G. Krein, S. C. Power, Sheldon Axler, Douglas N. Clark, V. V. Peller, R. G. Douglas, Carl Sundberg, N. Ya. Krupnik, I.È Verbitsky, Yu. I. Karlovich, I. M. Spitkovskii, L. A. Coburn, M. G. Krein, et al. Hankel and Toeplitz operators. Lecture Notes in Mathematics, 1043:251–307, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072188/.

Abikoff:1983:LSG

[AAS83] W. Abikoff, K. Appel, and P. Schupp. Lifting surface groups to SL(2, C)^{∗}. Lecture Notes in Mathematics, 971:1–5, 1983. CODEN LNMAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067067/.

Aasland:1983:GEP

[AB83] Liv Aasland and Petter Bjørstad. The generalized eigenvalue problem in shipdesign and offshore industry — a comparison of traditional methods with the Lanczos process. Lecture Notes in Mathematics, 973:146–155, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062099/.

Abdelkader:1983:TAP

[Abd83] Osama Abdelkader. Un théorème d’annulation pour les fibrés en droites semi-positifs sur une variété kählérienne faiblement 1-complète. (French) []. Lecture Notes in Mathematics, 1014: 1–16, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072066/.

Aroca:1982:AGP

[ABGM82] José Manuel Aroca, Ragnar Buchweitz, Marc Giusti, and Michel Merle, editors. Algebraic Geometry: Proceedings of the International Conference on Algebraic Geometry Held at La Rábida, Spain, January 7–15, 1981, volume 961 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 961. URL http:

//link.springer.com/book/10.1007/BFb0071272;   http:/
/www.springerlink.com/content/978-3-540-39367-2.

Abhyankar:1982:IN

[Abh82a] Shreeram S. Abhyankar. Indeterminate nets. Lecture Notes in Mathematics, 910:39–45, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093075/.

Abhyankar:1982:INR

[Abh82b] Shreeram S. Abhyankar. Indeterminate nets with restrictions. Lecture Notes in Mathematics, 910:46–52, 1982. CODEN LN-MAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093076/.

Abhyankar:1982:ISa

[Abh82c] Shreeram S. Abhyankar. Indeterminate strings. Lecture Notes in Mathematics, 910:14–20, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093067/.

Abhyankar:1982:ISR

[Abh82d] Shreeram S. Abhyankar. Indeterminate strings with restrictions. Lecture Notes in Mathematics, 910:21–25, 1982. CODEN LN-MAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093068/.

Abhyankar:1982:ISb

[Abh82e] Shreeram S. Abhyankar. Indexing strings. Lecture Notes in Mathematics, 910:29–30, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093070.

Abhyankar:1982:IPP

[Abh82f] Shreeram S. Abhyankar. Isobars for prechips and premonic polynomials. Lecture Notes in Mathematics, 910:59–66, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093079/.

Abhyankar:1982:Nb

[Abh82g] Shreeram S. Abhyankar. Nets. Lecture Notes in Mathematics, 910:31–32, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093071.

Abhyankar:1982:NR

[Abh82h] Shreeram S. Abhyankar. Nets on rings. Lecture Notes in Mathematics, 910:37–38, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093074.

Abhyankar:1982:Na

[Abh82i] Shreeram S. Abhyankar. Notation. Lecture Notes in Mathematics, 910:3, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093061.

Abhyankar:1982:OSN

[Abh82j] Shreeram S. Abhyankar. Ordered semigroup nets with restrictions. Lecture Notes in Mathematics, 910:36, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093073.

Abhyankar:1982:OSS

[Abh82k] Shreeram S. Abhyankar. Ordered semigroup strings with restrictions. Lecture Notes in Mathematics, 910:10, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093065.

Abhyankar:1982:P

[Abh82l] Shreeram S. Abhyankar. Prechips. Lecture Notes in Mathematics, 910:57–58, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093078.

Abhyankar:1982:RDOb

[Abh82m] Shreeram S. Abhyankar. Restricted degree and order for indeterminate nets. Lecture Notes in Mathematics, 910:53–56, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093077/.

Abhyankar:1982:RDOa

[Abh82n] Shreeram S. Abhyankar. Restricted degree and order for indeterminate strings. Lecture Notes in Mathematics, 910:26–28, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093069/.

Abhyankar:1982:SNR

[Abh82o] Shreeram S. Abhyankar. Semigroup nets with restrictions. Lecture Notes in Mathematics, 910:33–35, 1982. CODEN LN-MAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093072/.

Abhyankar:1982:SSR

[Abh82p] Shreeram S. Abhyankar. Semigroup strings with restrictions. Lecture Notes in Mathematics, 910:7–9, 1982. CODEN LN-MAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093064/.

Abhyankar:1982:Sa

[Abh82q] Shreeram S. Abhyankar. Semigroups. Lecture Notes in Mathematics, 910:4, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093062.

Abhyankar:1982:Sb

[Abh82r] Shreeram S. Abhyankar. Strings. Lecture Notes in Mathematics, 910:5–6, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093063.

Abhyankar:1982:SR

[Abh82s] Shreeram S. Abhyankar. Strings on rings. Lecture Notes in Mathematics, 910:11–13, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093066/.

Abhyankar:1982:Sc

[Abh82t] Shreeram S. Abhyankar. Substitutions. Lecture Notes in Mathematics, 910:67–72, 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093080/.

Abhyankar:1982:WEC

[Abh82u] Shreeram S. Abhyankar. Weighted Expansions for Canonical Desingularization, volume 910 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11195-6 (print), 3-540-38992-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 236 pp. LCCN QA3 .L28 no. 910.

URL http://link.springer.com/book/10.1007/BFb0093060;

 http://www.springerlink.com/content/978-3-540-38992-
7.

Albeverio:1984:NDP

[ABHK84] S. Albeverio, Ph. Blanchard, and R. Høegh-Krohn. Newtonian diffusions and planets, with a remark on non-standard Dirichlet forms and polymers. Lecture Notes in Mathematics, 1095: 1–24, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099118/.

Abikoff:1980:RAT

[Abi80] William Abikoff. The real analytic theory of Teichmüller space, volume 820 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 0-387-10237-X, 3-540-10237-X. ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 144 pp. LCCN QA3 .L28 no. 820; QA3 .L28 no. 820; QA331.

Allgower:1981:DCM

[ABM81] E. L. Allgower, K. Böhmer, and S. Mc Cormick. Discrete correction methods for operator equations. Lecture Notes in Mathematics, 878:30–97, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090676/.

Abraham:1982:PIF

[Abr82] D. B. Abraham. Planar Ising ferromagnet: correlation functions and the inverse scattering method. Lecture Notes in Mathematics, 925:37–56, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093498/.

ACampo:1980:PPD

[A’C80] Norbert A’Campo. Sur la première partie du seizième problème de Hilbert. (French) []. Lecture Notes in Mathematics, 770:208– 227, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096244/.

Arocena:1982:DGT

[AC82] R. Arocena and Prof. M. Cotlar. Dilation of generalized Toeplitz kernels and some vectorial moment and weighted problems. Lecture Notes in Mathematics, 908:169–188, 1982. CODEN LN-MAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093286/.

Arsene:1983:DTF

[AC83] Gr. Arsene and Zoia Ceauşescu. Dilation of two factorizations. Lecture Notes in Mathematics, 1014:256–264, 1983. CODEN LN-MAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072084/.

Accardi:1984:STP

[Acc84] Luigi Accardi. Some trends and problems in quantum probability. Lecture Notes in Mathematics, 1055:1–19, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071706/.

Achuthan:1981:CPC

[Ach81] Nirmala Achuthan. Characterization of potentially connected integer-pair sequences. Lecture Notes in Mathematics, 885:153– 164, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092259/.

Acharya:1984:APA

[Ach84] B. Devadas Acharya. Are all polyominoes arbitrarily graceful? Lecture Notes in Mathematics, 1073:205–211, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073118/.

Aczel:1981:OHD

[Acz81a] Peter Aczel. The ordinal height of a density. Lecture Notes in Mathematics, 890:25–31, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095655/.

Aczel:1981:TNP

[Acz81b] Peter Aczel. Two notes on the Paris independence result. Lecture Notes in Mathematics, 890:21–24, 1981. CODEN LN-MAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095654/.

Adhikary:1981:CCP

[AD81] Basudeb Adhikary and Premadhis Das. Construction and combinatorial properties of orthogonal arrays with variable number of symbols in rows. Lecture Notes in Mathematics, 885:165– 170, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092260/.

Alt:1983:CDM

[AD83a] Wolfgang Alt and Micah Dembo. A contraction-disassembly model for intracellular actin gels. Lecture Notes in Mathematics, 1017:1–9, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103231/.

Andersson:1983:NLP

[AD83b] Stig I. Andersson and Heinz-Dietrich Doebner, editors. Nonlinear Partial Differential Operators and Quantization Procedures: Proceedings of a workshop held at Clausthal Federal Republic of Germany, 1981, volume 1037 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1037. URL http:/

/link.springer.com/book/10.1007/BFb0073167;   http://
www.springerlink.com/content/978-3-540-38695-7.

Adamski:1982:TSF

[Ada82] Wolfgang Adamski. Tight set functions and essential measure. Lecture Notes in Mathematics, 945:1–14, 1982. CODEN LN-MAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096659/.

Adian:1980:CPW

[Adi80a] S. I. Adian. Classifications of periodic words and their application in group theory. Lecture Notes in Mathematics, 806:1–40, 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091266/.

Adian:1980:WPG

[Adi80b] S. I. Adian. On the word problem for groups defined by periodic relations. Lecture Notes in Mathematics, 806:41–46, 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091267/.

Angermann:1983:QKS

[ADT83] B. Angermann, H.-D. Doebner, and J. Tolar. Quantum kinematics on smooth manifolds. Lecture Notes in Mathematics, 1037: 171–208, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073173/.

Alexander:1981:GBS

[AF81] J. C. Alexander and P. M. Fitzpatrick. Global bifurcation for solutions of equations involving several parameter multivalued condensing mappings. Lecture Notes in Mathematics, 886:1–19, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092172/.

Atkinson:1982:SLC

[AF82] F. V. Atkinson and C. T. Fulton. Some limit circle eigenvalue problems and asymptotic formulae for eigenvalues. Lecture Notes in Mathematics, 964:28–55, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064986/.

Asimov:1983:UCO

[AF83] D. Asimov and J. Franks. Unremovable closed orbits. Lecture Notes in Mathematics, 1007:22–29, 1983. CODEN LN-MAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061407/.

Arazy:1984:SAM

[AF84] Jonathan Arazy and Stephen David Fisher. Some aspects of the minimal, Möbius-invariant space of analytic functions on the unit disc. Lecture Notes in Mathematics, 1070:24–44, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099090/.

Accardi:1984:QPA

[AFG84] Luigi Accardi, Alberto Frigerio, and Vittorio Gorini, editors. Quantum Probability and Applications to the Quantum Theory of Irreversible Processes: Proceedings of the International Workshop held at Villa Mondragone, Italy, September 6–11, 1982, volume 1055 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1055. URL http:/

/link.springer.com/book/10.1007/BFb0071705;   http://
www.springerlink.com/content/978-3-540-38798-5.

Alexander:1984:C

[AFHN84] H. Alexander, Frank Forelli, G. M. Henkin, and R. G. Novikov. C^{n}. Lecture Notes in Mathematics, 1043:619–630, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072195/.

Asimov:1980:MSF

[AG80] Daniel Asimov and Herman Gluck. Morse–Smale fields of geodesics. Lecture Notes in Mathematics, 819:1–17, 1980. CO-DEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086976/.

Anantharaman:1983:PRS

[AG83a] R. Anantharaman and K. M. Garg. The properties of a residual set of vector measures. Lecture Notes in Mathematics, 1033: 12–35, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099842/.

Arino:1983:AIF

[AG83b] O. Arino and I. Györi. Asymptotic integration of functional differential systems which are asymptotically autonomous. Lecture Notes in Mathematics, 1017:43–55, 1983. CODEN LN-MAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103234/.

Allgower:1981:NSN

[AGP81] Eugene L. Allgower, Klaus Glashoff, and Heinz-Otto Peitgen, editors. Numerical Solution of Nonlinear Equations: Proceedings, Bremen 1980, volume 878 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 0-387-10871-8, 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 878; QA3 .L471 no.878; QA3.L471. URL http:/

/link.springer.com/book/10.1007/BFb0090674;   http://
www.springerlink.com/content/978-3-540-38781-7.

Amar:1984:ACP

[AGT84] Eric Amar, Roger Gay, and Nguyen Van Thanh, editors. Analyse Complexe: Proceedings of the Journées Fermat — Journées SMF, held at Toulouse, May 24–27, 1983, volume 1094 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1094.

URL http://link.springer.com/book/10.1007/BFb0099148;
 http://www.springerlink.com/content/978-3-540-39096-
1.

Azema:1980:IUM

[AGY80] J. Azéma, R. F. Gundy, and M. Yor. Sur l’intégrabilité uniforme des martingales continues. (French) []. Lecture Notes in Mathematics, 784:53–61, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089469/.

Al-Hussaini:1981:IGF

[AHE81a] Ata Al-Hussaini and Robert J. Elliott. Itô and Girsanov formulae for two parameter processes. Lecture Notes in Mathematics, 851:464–469, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088736/.

Al-Hussaini:1981:SCT

[AHE81b] Ata Al-Hussaini and Robert J. Elliott. Stochastic calculus for a two parameter jump process. Lecture Notes in Mathematics, 863: 233–244, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091103/.

Albeverio:1982:DQF

[AHK82a] Sergio Albeverio and Raphael Høegh-Krohn. Diffusions, quantum fields and groups of mappings. Lecture Notes in Mathematics, 923:133–145, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093039/.

Albeverio:1982:SRD

[AHK82b] Sergio Albeverio and Raphael Høegh-Krohn. Some remarks on Dirichlet forms and their applications to quantum mechanics and statistical mechanics. Lecture Notes in Mathematics, 923:120– 132, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093038/.

Ahlfors:1980:EPG

[Ahl80] Lars V. Ahlfors. Ergodic properties of groups of Möbius transformations. Lecture Notes in Mathematics, 798:1–9, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097254/.

Ahlbrandt:1981:EDO

[Ahl81a] Calvin D. Ahlbrandt. Equivalence of differential operators. Lecture Notes in Mathematics, 846:1–16, 1981. CODEN LN-MAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089820/.

Ahlbrandt:1981:SUE

[AHL81b] Calvin D. Ahlbrandt, Don B. Hinton, and Roger T. Lewis. Some unitarily equivalent differential operators with finite and infinite singularities. Lecture Notes in Mathematics, 846:17–25, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089821/.

Ahlbrandt:1983:IUS

[AHL83] Calvin D. Ahlbrandt, Don B. Hinton, and Roger T. Lewis. Inversion in the unit sphere for powers of the Laplacian. Lecture Notes in Mathematics, 1032:1–8, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076789/.

Atkinson:1982:IIE

[AHS82] F. V. Atkinson, J. R. Haddock, and O. J. Staffans. Integral inequalities and exponential convergence of solutions of differential equations with bounded delay. Lecture Notes in Mathematics, 964:56–68, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064987/.

Airault:1983:ECI

[Air83] H. Airault. Étude des conditions d’intégrabilité associées à un système différentiel linéaire. (French) []. Lecture Notes in Mathematics, 1028:1–26, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071675/.

Aigner:1981:GGP

[AJ81] Martin Aigner and Dieter Jungnickel, editors. Geometries and Groups: Proceedings of a Colloquium Held at the Freie Universität Berlin, May 1981, volume 893 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 893. URL http:

//link.springer.com/book/10.1007/BFb0091009;   http:/
/www.springerlink.com/content/978-3-540-38639-1.

Alspach:1983:PS

[AJ83] Dale E. Alspach and William B. Johnson. Projections onto L_{1} subspaces of L_{1}(μ). Lecture Notes in Mathematics, 995:1–7, 1983. CODEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061884/.

Abrham:1982:CPS

[AK82] Jaromir Abrham and Anton Kotzig. Critical perfect systems of difference sets with the minimum start. Lecture Notes in Mathematics, 969:1–11, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062983/.

Akahira:1983:ADE

[Aka83] Masafumi Akahira. Asymptotic deficiencies of estimators for pooled samples from the same distribution. Lecture Notes in Mathematics, 1021:6–14, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072898/.

Akcoglu:1983:DSP

[Akc83] M. A. Akcoglu. Differentiation of superadditive processes. Lecture Notes in Mathematics, 1033:1–11, 1983. CODEN LN-MAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099841/.

Akyildiz:1982:VFC

[Aky82] Ersan Akyıldız. Vector fields and cohomology of G/P. Lecture Notes in Mathematics, 956:1–9, 1982. CODEN LNMAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101504/.

Arsove:1980:QUM

[AL80] Maynard Arsove and Heinz Leutwiler. Quasi-units in mixed lattice structures. Lecture Notes in Mathematics, 787:35–54, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086323/.

Auslander:1981:RAP

[AL81] Maurice Auslander and Emilo Lluis, editors. Representations of Algebras: Proceedings of the Third International Conference on Representations of Algebras Held in Puebla, Mexico, August 4–8 1980, volume 903 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 903. URL http:/

/link.springer.com/book/10.1007/BFb0092979;   http://
www.springerlink.com/content/978-3-540-38963-7.

Auslander:1982:RAW

[AL82] Maurice Auslander and Emilio Lluis, editors. Representations of Algebras: Workshop Notes of the Third International Conference on Representations of Algebras Held in Puebla, Mexico, August 4–8, 1980, volume 944 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11577-3 (print), 3-540-39313-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 944. URL http:

//link.springer.com/book/10.1007/BFb0094055;   http:/
/www.springerlink.com/content/978-3-540-39313-9.

Albers:1980:AEP

[Alb80] Willem Albers. Asymptotic expansions for the power of adaptive rank tests in the one-sample problem. Lecture Notes in Mathematics, 821:108–158, 1980. CODEN LNMAA2. ISBN 3-540-10239-6 (print), 3-540-38318-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097427/.

Albrecht:1983:ERPa

[Alb83a] Ulrich Albrecht. Endomorphism rings and A-projective torsionfree Abelian groups. Lecture Notes in Mathematics, 1006:209– 227, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21560-9_9/.

Albrecht:1983:ERPb

[Alb83b] Ulrich Albrecht. Endomorphism rings and A-projective torsionfree Abelian groups. Lecture Notes in Mathematics, 1006:209– 227, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103704/.

Aleksandrov:1981:ENL

[Ale81a] A. B. Aleksandrov. Essays on non locally convex Hardy classes. Lecture Notes in Mathematics, 864:1–89, 1981. CODEN LN-MAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096996/.

Alev:1981:FMC

[Ale81b] J. Alev. Sur la formule de Molien dans certaines algèbres enveloppantes. (French) []. Lecture Notes in Mathematics, 867:351– 364, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090392/.

Alexander:1981:DGR

[Ale81c] S. Alexander. Distance geometry in Riemannian manifolds-with-boundary. Lecture Notes in Mathematics, 838:12–18, 1981. CO-DEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088837/.

Alekseev:1983:UAK

[Ale83a] V. G. Alekseev. On the use of alternating kernels in nonparametric statistical estimation. Lecture Notes in Mathematics, 1021: 15–25, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072899/.

Alev:1983:EF

[Ale83b] J. Alev. Sur l’extension R^{G} → R. (French) []. Lecture Notes in Mathematics, 1029:267–282, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098935/.

Ali:1982:GPP

[Ali82] S. Twareque Ali. A geometrical property of POV-measures and systems of covariance. Lecture Notes in Mathematics, 905:207– 228, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092439/.

Alicki:1984:STQ

[Ali84] Robert Alicki. Scattering theory for quantum dynamical semigroups. Lecture Notes in Mathematics, 1055:20–31, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071707/.

Allgower:1981:SHM

[All81a] E. L. Allgower. A survey of homotopy methods for smooth mappings. Lecture Notes in Mathematics, 878:1–29, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090675/.

Alligood:1981:TCC

[All81b] Kathleen T. Alligood. Topological conditions for the continuation of fixed points. Lecture Notes in Mathematics, 886:20–32, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092173/.

Alladi:1982:AFS

[All82a] Krishnaswami Alladi. Additive functions and special sets of integers. Lecture Notes in Mathematics, 938:1–49, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097172/.

Alladi:1982:NTP

[All82b] Krishnaswami Alladi, editor. Number Theory: Proceedings of the Third Matscience Conference Held at Mysore, India, June 3–6, 1981, volume 938 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 938. URL http:/

/link.springer.com/book/10.1007/BFb0097171;   http://
www.springerlink.com/content/978-3-540-39279-8.

Allain:1982:DFD

[All82c] Guy Allain. Decomposition des formes de Dirichlet. (French) []. Lecture Notes in Mathematics, 906:1–26, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093259/.

Allinger:1982:NSN

[All82d] Deborah Allinger. A note on strong, non-anticipating solutions for stochastic differential equations: When is path-wise uniqueness necessary? Lecture Notes in Mathematics, 939:1–5, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096253/.

Allan:1983:CBA

[All83a] G. R. Allan. Commutative Banach algebras with power-series generators. Lecture Notes in Mathematics, 975:290–294, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064560/.

Allan:1983:IIP

[All83b] G. R. Allan. An inequality involving product measures. Lecture Notes in Mathematics, 975:277–279, 1983. CODEN LN-MAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064557/.

Alladi:1984:MAF

[All84] Krishnaswami Alladi. Moments of additive functions and sieve methods. Lecture Notes in Mathematics, 1052:1–25, 1984. CO-DEN LNMAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071538/.

Almkvist:1980:RTR

[Alm80] Gert Almkvist. Reciprocity theorems for representations in characteristic p. Lecture Notes in Mathematics, 795:1–9, 1980. CO-DEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090111/.

Almkvist:1983:IC

[Alm83] Gert Almkvist. Invariants of z/pz in characteristic p. Lecture Notes in Mathematics, 996:109–117, 1983. CODEN LN-MAA2. ISBN 3-540-12319-9 (print), 3-540-40043-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063237/.

Alspach:1981:SLP

[Als81] Brian Alspach. The search for long paths and cycles in vertextransitive graphs and digraphs. Lecture Notes in Mathematics, 884:14–22, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091804/.

Alter:1981:CGR

[Alt81] Ronald Alter. Computations and generalizations on a remark of Ramanujan. Lecture Notes in Mathematics, 899:183–196, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096461/.

Amayo:1981:ACL

[Ama81] Ralph K. Amayo, editor. Algebra Carbondale 1980: Lie Algebras, Group Theory, and Partially Ordered Algebraic Structures Proceedings of the Southern Illinois Algebra Conference, Carbondale, April 18 and 19, 1980, volume 848 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 848. URL http:

//link.springer.com/book/10.1007/BFb0090552;   http:/
/www.springerlink.com/content/978-3-540-38549-3.

Ambrosetti:1983:DEM

[Amb83] Antonio Ambrosetti. Differential equations with multiple solutions and nonlinear functional analysis. Lecture Notes in Mathematics, 1017:10–37, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103232/.

Ames:1982:ONK

[Ame82] W. F. Ames. Optimization of nonlinear kinetic equation computation. Lecture Notes in Mathematics, 968:165–189, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064887/.

Amick:1980:SSN

[Ami80] C. J. Amick. Steady solutions of the Navier–Stokes equations representing plane flow in channels of various types. Lecture Notes in Mathematics, 771:1–11, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086898/.

Amitsur:1982:GSF

[Ami82] S. A. Amitsur. Generic splitting fields. Lecture Notes in Mathematics, 917:1–24, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092225/.

Alligood:1983:IGC

[AMPY83] Kathleen T. Alligood, John Mallet-Paret, and James A. Yorke. An index for the global continuation of relatively isolated sets of periodic orbits. Lecture Notes in Mathematics, 1007:1–21, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061406/.

Ansorge:1982:ISN

[AMT82] Rainer Ansorge, Theodor Meis, and Willi Törnig, editors. Iterative Solution of Nonlinear Systems of Equations: Proceedings of a Meeting held at Oberwolfach, Germany, Jan. 31–Feb. 5, 1982, volume 953 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 953. URL http:/

/link.springer.com/book/10.1007/BFb0069370;   http://
www.springerlink.com/content/978-3-540-39379-5.

Andreka:1981:CRS

[AN81] H. Andréka and I. Németi. On cylindric-relativized set algebras. Lecture Notes in Mathematics, 883:131–315, 1981. CODEN LNMAA2. ISBN 3-540-10881-5 (print), 3-540-38790-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095614/.

Anandam:1980:ASF

[Ana80] V. Anandam. Admissible superharmonic functions. Lecture Notes in Mathematics, 787:1–8, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086320/.

Ancona:1980:PHF

[Anc80] Alano Ancona. Principe de Harnack à la frontière et problèmes de frontière de Martin. (French) []. Lecture Notes in Mathematics, 787:9–28, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086321/.

Ancona:1984:LTD

[Anc84a] Alano Ancona. L’énergie et la théorie du potentiel dans l’oeuvre de Jacques Deny. (French) []. Lecture Notes in Mathematics, 1096: 10–21, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100104/.

Ancona:1984:CCC

[Anc84b] Alano Ancona. Sur une conjecture concernant la capacité et l’effilement. (French) []. Lecture Notes in Mathematics, 1096: 34–68, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100106/.

Andre:1980:PME

[And80] Michel André. Produits de Massey et (2p + 1)èmes déviations. (French) []. Lecture Notes in Mathematics, 795:341–359, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090127/.

Anderson:1981:EEG

[And81a] Mary J. Anderson. Explicit expressions for the generators of the center of the enveloping algebra of real Lie algebras and for the algebra of bivariant operators on the group. Lecture Notes in Mathematics, 848:84–109, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090558/.

Andre:1981:PON

[And81b] Johannes André. On possible orders of noncommutative tactical spaces. Lecture Notes in Mathematics, 893:149–156, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091015/.

Andrews:1981:MIR

[And81c] George E. Andrews. Mordell integrals and Ramanujan’s “lost” notebook. Lecture Notes in Mathematics, 899:10–48, 1981. CO-DEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096452/.

Andersson:1982:VBC

[And82] Stig I. Andersson. Vector bundle connections and liftings of partial differential operators. Lecture Notes in Mathematics, 905: 119–132, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092432/.

Anderson:1984:FO

[And84a] J. M. Anderson. The Faber operator. Lecture Notes in Mathematics, 1105:1–10, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072396/.

Andersson:1984:POC

[And84b] Stig I. Andersson. Pseudodifferential operators and characteristic classes for non-abelian cohomology. Lecture Notes in Mathematics, 1045:1–10, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072160/.

Angeniol:1981:CCL

[Ang81a] Bernard Angéniol. Classes de cohomologie locale et résidus. (French) []. Lecture Notes in Mathematics, 896:33–56, 1981. CO-DEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091699/.

Angeniol:1981:FCA

[Ang81b] Bernard Angéniol. Familles de Cycle Algébriques — Schéma de Chow, volume 896 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 140 pp. LCCN QA3 .L28 no. 896. URL http:

//link.springer.com/book/10.1007/BFb0091696;   http:/
/www.springerlink.com/content/978-3-540-38644-5.

Angeniol:1981:FCF

[Ang81c] Bernard Angéniol. Foncteur de chow. (French) []. Lecture Notes in Mathematics, 896:76–82, 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091701/.

Angeniol:1981:FMW

[Ang81d] Bernard Angéniol. Formes multilinéaires de Wäring. (French) []. Lecture Notes in Mathematics, 896:9–32, 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091698/.

Angeniol:1981:IEA

[Ang81e] Bernard Angéniol. Intersections — Équivalences algébriques. (French) []. Lecture Notes in Mathematics, 896:131–137, 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091705/.

Angeniol:1981:IFI

[Ang81f] Bernard Angéniol. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 896:1–8, 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091697/.

Angeniol:1981:IFM

[Ang81g] Bernard Angéniol. Invariance des formes multilinéaires de waring: Interprétation géométrique, finitude. (French) []. Lecture Notes in Mathematics, 896:57–75, 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091700/.

Angeniol:1981:CPC

[Ang81h] Bernard Angéniol. Le cas projectif — le cas réduit. (French) []. Lecture Notes in Mathematics, 896:96–107, 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091703/.

Angeniol:1981:MDS

[Ang81i] Bernard Angéniol. Le morphisme du schéma de Hilbert dans le schéma de chow. (French) []. Lecture Notes in Mathematics, 896: 108–130, 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091704/.

Angeniol:1981:RDF

[Ang81j] Bernard Angéniol. Représentabilité du foncteur c_{x}^{p}. (French) []. Lecture Notes in Mathematics, 896:83–95, 1981. CODEN LNMAA2. ISBN 3-540-11169-7 (print), 3-540-38644-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091702/.

Anichkin:1983:HAP

[Ani83] S. A. Anichkin. Hypererlang approximation of probability distributions on (0, ∞) and its application. Lecture Notes in Mathematics, 982:1–16, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082056/.

Anick:1984:CBY

[Ani84] David J. Anick. Connections between Yoneda and Pontrjagin algebras. Lecture Notes in Mathematics, 1051:331–350, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075575/.

Anonymous:1980:SBV

[Ano80] Anonymous, editor. Séminaire Bourbaki vol. 1978/79 Exposés 525– 542: Avec table par noms d’auteurs de 1967/68 à 1978/ 79. (French) [], volume 770 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 770. URL http:/

/link.springer.com/book/10.1007/BFb0096231;   http://
www.springerlink.com/content/978-3-540-38547-9.

Anonymous:1981:ASA

[Ano81a] Anonymous, editor. Aspects statistiques et aspects physiques des processus gaussiens: [colloque], Saint-Flour, 22–29 juin 1980. (French) [Statistical aspects and physical aspects of Gaussian processes: [conference] Saint-Flour, 22–29 June 1980], volume 307 of Colloques internationaux du Centre national de la recherche scientifique. Éditions du Centre national de la recherche scientifique, Paris, France, 1981. ISBN 2-222-02946-5. LCCN QA3 .L28 no. 307; QA274.4 .A87 1981.

Anonymous:1981:SBVa

[Ano81b] Anonymous, editor. Séminaire Bourbaki vol. 1979/80 Exposés 543– 560: Avec table par noms d’auteurs de 1967/68 à 1979/ 80. (French) [], volume 842 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 842. URL http:/

/link.springer.com/book/10.1007/BFb0089923;   http://
www.springerlink.com/content/978-3-540-38437-3.

Anonymous:1981:SBVb

[Ano81c] Anonymous, editor. Séminaire Bourbaki vol. 1980/81 Exposés 561–578: Avec table par noms d’auteurs de 1967/68 à 1980/ 81. (French) [], volume 901 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 901. URL http:/

/link.springer.com/book/10.1007/BFb0097185;   http://
www.springerlink.com/content/978-3-540-38956-9.

Antman:1980:GAG

[Ant80] Stuart S. Antman. Geometric aspects of global bifurcation in nonlinear elasticity. Lecture Notes in Mathematics, 775:1–29, 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092018/.

Antoniadis:1981:SBS

[Ant81] A. Antoniadis. Statistics on Banach space valued Gaussian random variables. Lecture Notes in Mathematics, 860:1–8, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090604/.

Antonov:1983:ABT

[Ant83] S. N. Antonov. On the asymptotic behaviour of the tails of infinitely divisible distribution laws. Lecture Notes in Mathematics, 1021:41–44, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072901/.

Alford:1980:ANQ

[AP80] W. A. Alford and Bodo Pietsch. An application of the nilpotent quotient program. Lecture Notes in Mathematics, 806:47– 48, 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091268.

Adhikary:1981:CGD

[AP81a] Basudeb Adhikary and Rajendranath Panda. Construction of group divisible rotatable designs. Lecture Notes in Mathematics, 885:171–184, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092261/.

Ansemil:1981:EQN

[AP81b] J. M. Ansemil and S. Ponte. An example of a quasi-normable Fréchet function space which is not a Schwartz space. Lecture Notes in Mathematics, 843:1–8, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089266/.

Apanasov:1983:CCR

[Apa83] Boris Nikolaevič Apanasov. Condition of conformal rigidity of hyperbolic manifolds with boundaries. Lecture Notes in Mathematics, 1039:1–8, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073354/.

Appell:1983:GFN

[App83] Jürgen Appell. ’Genaue’ Fixpunktsätze und nichtlineare Sturm– Liouville-Probleme. (German) [’Exact’ fixed-point theorems and nonlinear Sturm–Liouville problems]. Lecture Notes in Mathematics, 1017:38–42, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103233/.

Arques:1980:CPP

[AR80a] Didier Arquès and Jean-Pierre Roth. Charge portée par une surface et champ associé. (French) []. Lecture Notes in Mathematics, 814:1–19, 1980. CODEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094145/.

Arques:1980:CSC

[AR80b] Didier Arquès and Jean-Pierre Roth. Courant superficiel et champ magnetostatique associé. (French) []. Lecture Notes in Mathematics, 787:29–34, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086322/.

Auslander:1980:UF

[AR80c] Maurice Auslander and Idun Reiten. Uniserial functors. Lecture Notes in Mathematics, 832:1–47, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088457/.

Arazy:1983:IUM

[Ara83] Jonathan Arazy. Isomorphisms of unitary matrix spaces. Lecture Notes in Mathematics, 991:1–6, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061553/.

Arazy:1984:BSP

[Ara84] J. Arazy. N.b. — some of these problems were prepared already for the 1982 conference. Lecture Notes in Mathematics, 1070: 232–239, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099105/.

Arendt:1984:GPS

[Are84] W. Arendt. Generators of positive semigroups. Lecture Notes in Mathematics, 1076:1–15, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072760/.

Arkin:1984:ECM

[Ark84] V. I. Arkin. Extension of the class of Markov controls. Lecture Notes in Mathematics, 1091:55–65, 1984. CODEN LN-MAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098801/.

Alladi:1982:PSC

[ARMS82] Krishnaswami Alladi, K. Ramachandra, M. Ram Murthy, and R. Sivaramakrishnan. Problem session (chaired by Prof. P. Erdős) two problems proposed by K. Ramachandra. Lecture Notes in Mathematics, 938:170–177, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097183/.

Arnold:1981:PSF

[Arn81] David M. Arnold. Pure subgroups of finite rank completely decomposable groups. Lecture Notes in Mathematics, 874:1– 31, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090520/.

Arnold:1982:ACQ

[Arn82a] David M. Arnold. Additive categories, quasi-isomorphism and near-isomorphism. Lecture Notes in Mathematics, 931:69–86, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094253/.

Arnold:1982:AGS

[Arn82b] David M. Arnold. Additive groups of subrings of finite dimensional Q-algebras. Lecture Notes in Mathematics, 931:160– 166, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094260/.

Arnold:1982:CDG

[Arn82c] David M. Arnold. Completely decomposable groups and generalizations. Lecture Notes in Mathematics, 931:58–68, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094252/.

Arnold:1982:ERD

[Arn82d] David M. Arnold. Endomorphism rings and decompositions of rank-2 groups. Lecture Notes in Mathematics, 931:36–39, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094249/.

Arnold:1982:EIG

[Arn82e] David M. Arnold. Examples of indecomposable groups and direct sums. Lecture Notes in Mathematics, 931:21–35, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094248/.

Arnold:1982:FRT

[Arn82f] David M. Arnold. Finite Rank Torsion Free Abelian Groups and Rings, volume 931 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 191 pp. LCCN QA3 .L28 no. 931. URL http:

//link.springer.com/book/10.1007/BFb0094245;   http:/
/www.springerlink.com/content/978-3-540-39234-7.

Arnold:1982:GG

[Arn82g] David M. Arnold. Grothendieck groups. Lecture Notes in Mathematics, 931:148–159, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094259/.

Arnold:1982:HCD

[Arn82h] David M. Arnold. Homogeneous completely decomposable groups and generalizations. Lecture Notes in Mathematics, 931:47– 57, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094251/.

Arnold:1982:MOF

[Arn82i] David M. Arnold. Maximal orders in finite dimensional simple Q-algebras. Lecture Notes in Mathematics, 931:129–138, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094257/.

Arnold:1982:NIG

[Arn82j] David M. Arnold. Near isomorphism and genus class. Lecture Notes in Mathematics, 931:139–147, 1982. CODEN LN-MAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094258/.

Arnold:1982:NP

[Arn82k] David M. Arnold. Notation and preliminaries. Lecture Notes in Mathematics, 931:1–4, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094246/.

Arnold:1982:OFD

[Arn82l] David M. Arnold. Orders in finite dimensional simple Q-algebras. Lecture Notes in Mathematics, 931:119–128, 1982. CODEN LN-MAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094256/.

Arnold:1982:PSC

[Arn82m] David M. Arnold. Pure subgroups of completely decomposable groups. Lecture Notes in Mathematics, 931:40–46, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094250/.

Arnold:1982:SSG

[Arn82n] David M. Arnold. Q-simple and p-simple groups. Lecture Notes in Mathematics, 931:167–177, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094261/.

Arnold:1982:SRS

[Arn82o] David M. Arnold. Stable range, substitution, cancellation, and exchange properties. Lecture Notes in Mathematics, 931:87– 107, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094254/.

Arnold:1982:SFD

[Arn82p] David M. Arnold. Subrings of finite dimensional Q-algebras. Lecture Notes in Mathematics, 931:108–118, 1982. CODEN LN-MAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094255/.

Arnold:1982:TRG

[Arn82q] David M. Arnold. Types and rank-1 groups. Lecture Notes in Mathematics, 931:5–20, 1982. CODEN LNMAA2. ISBN 3-540-11557-9 (print), 3-540-39234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094247/.

Arques:1984:ROV

[Arq84] Didier Arquès. Représentation des opérateurs vérifiant le principe du maximum positif. (French) []. Lecture Notes in Mathematics, 1061:1–13, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099014/.

Arscott:1981:LBB

[Ars81] F. M. Arscott. The land beyond Bessel: a survey of higher special functions. Lecture Notes in Mathematics, 846:26–45, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089822/.

Artemiades:1980:CAC

[Art80a] Nicolas K. Artémiades. Criteria for absolute convergence of Fourier series. Lecture Notes in Mathematics, 781:1–7, 1980.

CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097641/.

Artzy:1980:FMP

[Art80b] R. Artzy. On free Minkowski planes. Lecture Notes in Mathematics, 792:1–7, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088653/.

Artin:1982:BSV

[Art82a] M. Artin. Brauer–Severi varieties. Lecture Notes in Mathematics, 917:194–210, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092235/.

Artin:1982:LIM

[Art82b] M. Artin. Left ideals in maximal orders. Lecture Notes in Mathematics, 917:182–193, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092234/.

Artin:1982:LSM

[Art82c] M. Artin. Local structure of maximal orders on surfaces. Lecture Notes in Mathematics, 917:146–181, 1982. CODEN LN-MAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092233/.

Arthur:1983:SPS

[Art83] James Arthur. On some problems suggested by the trace formula. Lecture Notes in Mathematics, 1041:1–49, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073144/.

Arthur:1984:PRN

[Art84a] M. D. Arthur. Preliminary results on the non-existence of solutions for a half space Boltzmann collision model with three degrees of freedom. Lecture Notes in Mathematics, 1048:176–191, 1984.

CODEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071880/.

Artstein:1984:LLM

[Art84b] Zvi Artstein. Limit laws for multifunctions applied to an optimization problem. Lecture Notes in Mathematics, 1091:66–79, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098802/.

Auslander:1980:PMI

[AS80] M. Auslander and S. O. Smalø. Preprojective modules: an introduction and some applications. Lecture Notes in Mathematics, 832:48–73, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088458/.

Anderson:1981:OT

[AS81a] Robert J. Anderson and H. M. Stark. Oscillation theorems. Lecture Notes in Mathematics, 899:79–106, 1981. CODEN LN-MAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096454/.

Auslander:1981:PLC

[AS81b] M. Auslander and S. O. Smalø. Preprojective lattices over classical orders. Lecture Notes in Mathematics, 882:326–344, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092503/.

Adamek:1982:LIC

[AS82a] J. Adámek and G. E. Strecker. On the largest initial completion of categories of algebras. Lecture Notes in Mathematics, 915: 1–15, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092867/.

Auzinger:1982:DCM

[AS82b] W. Auzinger and H. J. Stetter. Defect corrections and multigrid iterations. Lecture Notes in Mathematics, 960:327–351, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069932/.

Antosik:1983:NBT

[AS83] Piotr Antosik and Charles Swartz. The nikodým boundedness theorem and the uniform boundedness principle. Lecture Notes in Mathematics, 1033:36–42, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099843/.

Adachi:1984:GBN

[AS84a] T. Adachi and T. Sunada. Geometric bounds for the number of certain harmonic mappings. Lecture Notes in Mathematics, 1090: 24–36, 1984. CODEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101564/.

Ambos-Spies:1984:CRD

[AS84b] Klaus Ambos-Spies. Contiguous R.E. degrees. Lecture Notes in Mathematics, 1104:1–37, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099477/.

Asorey:1984:EYM

[Aso84] M. Asorey. Euclidean Yang–Mills flows in the orbit space. Lecture Notes in Mathematics, 1045:11–20, 1984. CODEN LN-MAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072161/.

Ancona:1982:GAF

[AT82a] Vincenzo Ancona and Giuseppe Tomassini. Géométrie analytique formelle. (French) []. Lecture Notes in Mathematics, 943: 5–20, 1982. CODEN LNMAA2. ISBN 3-540-11570-6 (print), 3-540-39287-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095884/.

Ancona:1982:I

[AT82b] Vincenzo Ancona and Giuseppe Tomassini. Introduction. Lecture Notes in Mathematics, 943:1–4, 1982. CODEN LNMAA2. ISBN 3-540-11570-6 (print), 3-540-39287-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095883/.

Ancona:1982:MA

[AT82c] Vincenzo Ancona and Giuseppe Tomassini. Modifications Analytiques, volume 943 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11570-6 (print), 3-540-39287-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 120 pp. LCCN QA3 .L28 no. 943. URL http:

//link.springer.com/book/10.1007/BFb0095882;   http:/
/www.springerlink.com/content/978-3-540-39287-3.

Ancona:1982:QCF

[AT82d] Vincenzo Ancona and Giuseppe Tomassini. Quelques constructions. (French) []. Lecture Notes in Mathematics, 943:21–33, 1982. CODEN LNMAA2. ISBN 3-540-11570-6 (print), 3-540-39287-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095885/.

Ancona:1982:TAF

[AT82e] Vincenzo Ancona and Giuseppe Tomassini. Théorèmes d’algébrisation. (French) [Theorems of algebraization]. Lecture Notes in Mathematics, 943:82–109, 1982. CODEN LNMAA2. ISBN 3-540-11570-6 (print), 3-540-39287-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095888/.

Ancona:1982:TEM

[AT82f] Vincenzo Ancona and Giuseppe Tomassini. Théorèmes d’existence des modifications analytiques. (French) []. Lecture Notes in Mathematics, 943:54–81, 1982. CODEN LNMAA2. ISBN 3-540-11570-6 (print), 3-540-39287-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095887/.

Ancona:1982:TMF

[AT82g] Vincenzo Ancona and Giuseppe Tomassini. Théorie des modifications. (French) [Theory of changes]. Lecture Notes in Mathematics, 943:34–53, 1982. CODEN LNMAA2. ISBN 3-540-11570-6 (print), 3-540-39287-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095886/.

Atkinson:1980:EBE

[Atk80] F. V. Atkinson. Exponential behaviour of eigenfunctions and gaps in the essential spectrum. Lecture Notes in Mathematics, 827:1–24, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091372/.

Atkinson:1982:ABT

[Atk82] F. V. Atkinson. On the asymptotic behaviour of the Titchmarsh– Weyl m-coefficient and the spectral function for scalar secondorder differential expressions. Lecture Notes in Mathematics, 964: 1–27, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064985/.

Attouch:1984:VPE

[Att84] H. Attouch. Variational properties of EPI-convergence, applications to limit analysis problems in mechanics and duality theory. Lecture Notes in Mathematics, 1091:80–104, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098803/.

Aubry:1982:DSC

[Aub82] Serge Aubry. The devil’s stair case transformation in incommensurate lattices. Lecture Notes in Mathematics, 925:221–245, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093512/.

Aubin:1984:SHV

[Aub84] Jean-Pierre Aubin. Slow and heavy viable trajectories of controlled problems. smooth viability domains. Lecture Notes in Mathematics, 1091:105–116, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098804/.

Aulbach:1983:AHM

[Aul83] Bernd Aulbach. Approach to hyperbolic manifolds of stationary solutions. Lecture Notes in Mathematics, 1017:56–66, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103235/.

Aulbach:1984:CDD

[Aul84a] Bernd Aulbach. Continuous and Discrete Dynamics near Manifolds of Equilibria, volume 1058 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13329-1 (print), 3-540-38853-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 142 pp. LCCN QA3 .L28 no. 1058.

URL http://link.springer.com/book/10.1007/BFb0071569;
 http://www.springerlink.com/content/978-3-540-38853-
1.

Aulbach:1984:DEb

[Aul84b] Bernd Aulbach. Difference equations. Lecture Notes in Mathematics, 1058:59–99, 1984. CODEN LNMAA2. ISBN 3-540-13329-1 (print), 3-540-38853-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071572/.

Aulbach:1984:DEa

[Aul84c] Bernd Aulbach. Differential equations. Lecture Notes in Mathematics, 1058:9–58, 1984. CODEN LNMAA2. ISBN 3-540-13329-1 (print), 3-540-38853-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071571/.

Aulbach:1984:I

[Aul84d] Bernd Aulbach. Introduction. Lecture Notes in Mathematics, 1058:1–8, 1984. CODEN LNMAA2. ISBN 3-540-13329-1 (print), 3-540-38853-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071570/.

Aupetit:1983:ACC

[Aup83] Bernard Aupetit. Automatic continuity conditions for a linear mapping from a Banach algebra onto a semi-simple Banach algebra. Lecture Notes in Mathematics, 975:313–316, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064564/.

Auslander:1982:FAR

[Aus82] Maurice Auslander. A functorial approach to representation theory. Lecture Notes in Mathematics, 944:105–179, 1982. CODEN LNMAA2. ISBN 3-540-11577-3 (print), 3-540-39313-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094058/.

Artzy:1980:GDG

[AV80] Rafael Artzy and Izu Vaisman, editors. Geometry and Differential Geometry: Proceedings of a Conference Held at the University of Haifa, Israel, March 18–23, 1979, volume 792 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 792.

URL http://link.springer.com/book/10.1007/BFb0088652;
 http://www.springerlink.com/content/978-3-540-39214-
9.

Arnold:1983:PSFa

[AV83a] D. Arnold and C. Vinsonhaler. Pure subgroups of finite rank completely decomposable groups II. Lecture Notes in Mathematics, 1006:97–143, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21560-9_3/.

Arnold:1983:PSFb

[AV83b] D. Arnold and C. Vinsonhaler. Pure subgroups of finite rank completely decomposable groups II. Lecture Notes in Mathematics, 1006:97–143, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103698/.

Azlarov:1983:DAM

[AV83c] T. A. Azlarov and N. A. Volodin. On the discrete analog of Marshall–Olkin’s distribution. Lecture Notes in Mathematics, 982:17–23, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082057/.

Avanissian:1982:FHO

[Ava82] V. Avanissian. Sur les fonctions harmoniques d’ordre quelconque et leur prolongement analytique dans C^{N}. (French) []. Lecture Notes in Mathematics, 919:192–281, 1982. CODEN LN-MAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097052/.

Avramov:1982:IGF

[Avr82] Luchézar L. Avramov. Invariants d’un groupe fini engendré par des pseudo-reflexions, opérant sur un anneau local. (French) []. Lecture Notes in Mathematics, 924:375–383, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092941/.

Azarnia:1982:NCP

[AW82] N. Azarnia and J. D. Maitland Wright. A non-commutative Pettis theorem. Lecture Notes in Mathematics, 945:208–212, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096676/.

Attouch:1983:CBF

[AW83] Hédy Attouch and Roger J.-B. Wets. A convergence for bivariate functions aimed at the convergence of saddle values. Lecture Notes in Mathematics, 979:1–42, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066247/.

Axelsson:1980:GCD

[Axe80] O. Axelsson. A generalized conjugate direction method and its application on a singular perturbation problem. Lecture Notes in Mathematics, 773:1–11, 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094159/.

Axelsson:1982:GCI

[Axe82a] O. Axelsson. On global convergence of iterative methods. Lecture Notes in Mathematics, 953:1–19, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069371/.

Axelsson:1982:MMT

[Axe82b] O. Axelsson. On multigrid methods of the two-level type. Lecture Notes in Mathematics, 960:352–367, 1982. CODEN LN-MAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069933/.

Azema:1980:SPX

[AY80] Jacques Azéma and Marc Yor, editors. Séminaire de Probabilités XIV 1978/79, volume 784 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 784. URL http:

//link.springer.com/book/10.1007/BFb0089463;   http:/
/www.springerlink.com/content/978-3-540-38642-1.

Azema:1981:SPX

[AY81] Jacques Azéma and Marc Yor, editors. Séminaire de Probabilités XV 1979/80: Avec table générale des exposés de 1966/ 67 à 1978/79, volume 850 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 850. URL http:/

/link.springer.com/book/10.1007/BFb0088355;   http://
www.springerlink.com/content/978-3-540-38610-0.

Azema:1982:SPXa

[AY82a] Jacques Azéma and Marc Yor, editors. Séminaire de Probabilités XVI 1980/81, volume 920 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 920. URL http:

//link.springer.com/book/10.1007/BFb0092765;   http:/
/www.springerlink.com/content/978-3-540-39158-6.

Azema:1982:SPXb

[AY82b] Jacques Azéma and Marc Yor, editors. Séminaire de Probabilités XVI, 1980/81 Supplément: Géométrie Différentielle Stochastique, volume 921 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11486-6 (print), 3-540-39167-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 921. URL http:/

/link.springer.com/book/10.1007/BFb0092646;   http://
www.springerlink.com/content/978-3-540-39167-8.

Azema:1983:SPX

[AY83] Jacques Azéma and Marc Yor, editors. Séminaire de Probabilités XVII 1981/82: Proceedings, volume 986 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 986. URL http:

//link.springer.com/book/10.1007/BFb0068294;   http:/
/www.springerlink.com/content/978-3-540-39614-7.

Azema:1984:SPX

[AY84] J. Azéma and M. Yor, editors. Séminaire de Probabilités XVIII 1982/83: Proceedings, volume 1059 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1059. URL http:/

/link.springer.com/book/10.1007/BFb0100027;   http://
www.springerlink.com/content/978-3-540-38858-6.

Azencott:1980:GDA

[Aze80] R. Azencott. Grandes déviations et applications. (French) [Large deviations and applications]. Lecture Notes in Mathematics, 774: 1–176, 1980. CODEN LNMAA2. ISBN 3-540-09741-4 (print), 3-540-38567-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/

BFb0089623/. English translation available in Large deviations at Saint-Flour, Springer (2012), ISBN 3-642-33199-8..

Azencott:1982:FTS

[Aze82] Robert Azencott. Formule de Taylor stochastique et développement asymptotique d’intégrales de Feynman. (French) []. Lecture Notes in Mathematics, 921:237–285, 1982. CODEN LNMAA2. ISBN 3-540-11486-6 (print), 3-540-39167-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092653/.

Babenko:1980:PSV

[Bab80] K. I. Babenko. On properties of steady viscous incompressible fluid flows. Lecture Notes in Mathematics, 771:12–42, 1980. CO-DEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086899/.

Bachmann:1980:RGI

[Bac80] F. Bachmann. Rigidity in the geometry of involutory elements of a group. Lecture Notes in Mathematics, 792:8–13, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088654/.

Bachelis:1983:SRQ

[Bac83] Gregory F. Bachelis. Some radical quotients in harmonic analysis. Lecture Notes in Mathematics, 975:301–308, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064562/.

Bade:1983:MWA

[Bad83] William G. Badé. Multipliers of weighted l^{1}-algebras. Lecture Notes in Mathematics, 975:227–247, 1983. CODEN LN-MAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064553/.

Badescu:1984:HSD

[Băd84] Lucian Bădescu. Hyperplane sections and deformations. Lecture Notes in Mathematics, 1056:1–33, 1984. CODEN LN-MAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071767/.

Baernstein:1982:BCT

[Bae82] Albert Baernstein II. Bieberbach’s conjecture for tourists. Lecture Notes in Mathematics, 908:48–73, 1982. CODEN LN-MAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093280/.

Baeten:1984:FUD

[Bae84] Jos Baeten. Filters and ultrafilters over definable subsets of admissible ordinals. Lecture Notes in Mathematics, 1103:1–8, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099378/.

Bailey:1981:DBS

[Bai81] R. A. Bailey. Distributive block structures and their automorphisms. Lecture Notes in Mathematics, 884:115–124, 1981. CO-DEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091813/.

Bais:1982:SCP

[Bai82a] F. A. Bais. Symmetry as a clue to the physics of elementary particles. Lecture Notes in Mathematics, 926:85–106, 1982. CODEN LNMAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092658/.

Bais:1982:TEG

[Bai82b] F. A. Bais. Topological excitations in gauge theories; an introduction from the physical point of view. Lecture Notes in Mathematics, 926:107–142, 1982. CODEN LNMAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092659/.

Bakry:1981:LQM

[Bak81] D. Bakry. Limites ‘quadrantales’ des martingales. Lecture Notes in Mathematics, 863:40–49, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091091/.

Bak:1982:SGL

[Bak82a] Anthony Bak. Subgroups of the general linear group normalized by relative elementary groups. Lecture Notes in Mathematics, 967:1–22, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061897/.

Baker:1982:INT

[Bak82b] Christopher T. H. Baker. An introduction to the numerical treatment of Volterra and Abel-type integral equations. Lecture Notes in Mathematics, 965:1–38, 1982. CODEN LNMAA2. ISBN 3-540-11967-1 (print), 3-540-39558-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063199/.

Bakry:1982:RPO

[Bak82c] D. Bakry. Remarques sur le processus d’Ornstein Uhlenbeck en dimension infinie. (French) []. Lecture Notes in Mathematics, 920: 134–137, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092774/.

Bakaj:1983:IMA

[Bak83] A. S. Bakaj. Integral manifolds and adiabatic invariants of systems in slow evolution. Lecture Notes in Mathematics, 985:99– 136, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062365/.

Bak:1984:AEN

[Bak84a] Anthony Bak, editor. Algebraic K-theory, number theory, geometry, and analysis: proceedings of the international conference held at Bielefeld, Federal Republic of Germany, July 26–30, 1982, volume 1046 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 0-387-12891-3, 3-540-12891-3. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1046; QA3 .L28 no. 1046; QA612.33. DM55.00 (Germany).

Bak:1984:NTG

[Bak84b] Anthony Bak. A norm theorem for K_{2} of global fields. Lecture Notes in Mathematics, 1051:1–7, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075560/.

Baker:1984:SIE

[Bak84c] Neil Baker. Some integral equalities in Wiener–Hopf theory. Lecture Notes in Mathematics, 1095:169–186, 1984. CODEN LN-MAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099129/.

Bakry:1984:EPT

[Bak84d] D. Bakry. Étude probabiliste des transformées de Riesz et de l’espace H^{1} sur les sphères. (French) []. Lecture Notes in Mathematics, 1059:197–218, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100045/.

Baldwin:1981:DHS

[Bal81a] John T. Baldwin. Definability and the hierarchy of stable theories. Lecture Notes in Mathematics, 859:1–15, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090935/.

Ball:1981:DCL

[Bal81b] Richard N. Ball. The distinguished completion of a lattice ordered group. Lecture Notes in Mathematics, 848:208–217, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090567/.

Banica:1980:FIM

[Băn80] Constantin Bănică. Sur les fibres infinitesimales d’un morphisme propre d’espaces complexes. (French) []. Lecture Notes in Mathematics, 807:122–144, 1980. CODEN LNMAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090877/.

Banaschewski:1981:DDC

[Ban81a] B. Banaschewski. The duality of distributive σ-continuous lattices. Lecture Notes in Mathematics, 871:12–19, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089901/.

Banaschewski:1981:CF

[Ban81b] Bernhard Banaschewski. Coherent frames. Lecture Notes in Mathematics, 871:1–11, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089900/.

Banchoff:1981:DTT

[Ban81c] Thomas F. Banchoff. Double tangency theorems for pairs of submanifolds. Lecture Notes in Mathematics, 894:26–48, 1981. CODEN LNMAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096223/.

Banaschewski:1982:CAT

[Ban82] B. Banaschewski, editor. Categorical Aspects of Topology and Analysis: Proceedings of an International Conference Held at Carleton University, Ottawa, August 11–15, 1981, volume 915 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 915. URL http:/

/link.springer.com/book/10.1007/BFb0092866;   http://
www.springerlink.com/content/978-3-540-39041-1.

Bandelt:1983:FOV

[Ban83a] Hans-J. Bandelt. Free objects in the variety generated by rings and distributive lattices. Lecture Notes in Mathematics, 998:255– 260, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062034/.

Banulescu:1983:PLC

[Băn83b] Martha Bănulescu. Potentials on locally compact non-abelian groups. Lecture Notes in Mathematics, 1014:129–154, 1983. CO-DEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072076/.

Barbanti:1980:LEC

[Bar80a] Luciano Barbanti. Liénard equations and control. Lecture Notes in Mathematics, 799:1–22, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089306/.

Barlet:1980:CAV

[Bar80b] D. Barlet. Convexité au voisinage d’un cycle. (French) []. Lecture Notes in Mathematics, 807:102–121, 1980. CODEN LN-MAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090876/.

Barlet:1980:MDV

[Bar80c] D. Barlet. Majoration du volume des fibres génériques et forme géomètrique du théorème d’aplatissement. (French) []. Lecture Notes in Mathematics, 822:1–17, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097745/.

Barlet:1980:FAC

[Bar80d] Daniel Barlet. Familles analytiques de cycles et classes fondamentales relatives. (French) []. Lecture Notes in Mathematics, 807:1–24, 1980. CODEN LNMAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090873/.

Barnsley:1980:PDS

[Bar80e] M. F. Barnsley. Parameter dependence of solutions of classes of quasi-linear elliptic and parabolic differential equations. Lecture Notes in Mathematics, 782:1–15, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090425/.

Barroso:1980:CTE

[Bar80f] Jorge Alberto Barroso. Comparaison de topologies sur des espaces d’applications holomorphes. (French) []. Lecture Notes in Mathematics, 822:18–32, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097746/.

Bardos:1981:AES

[Bar81a] Claude Bardos. Apparition éventuelle de singularités dans des problèmes d’évolution non linéaires [d’après S. Klainerman, B. Glassey, J. Chadam, F. John (et d’autres)]. (French) []. Lecture Notes in Mathematics, 842:215–224, 1981. CODEN LN-MAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089936/.

Barlow:1981:BLT

[Bar81b] M. T. Barlow. On Brownian local time. Lecture Notes in Mathematics, 850:189–190, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088367.

Barlet:1982:DAF

[Bar82a] D. Barlet. Développements asymptotiques des fonctions obtenues par intégration sur les fibres. (French) []. Lecture Notes in Mathematics, 919:282–293, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097053/.

Barlow:1982:S

[Bar82b] M. T. Barlow. L(B_{t}, t) is not a semimartingale. Lecture Notes in Mathematics, 920:209–211, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092782/.

Bartels:1982:NAQ

[Bar82c] Dina Bartels. On nonnormality of affine quasi-homogeneous SL(2, C)-varieties. Lecture Notes in Mathematics, 924:384–399, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092942/.

Barut:1982:WKD

[Bar82d] A. O. Barut. What kind of a dynamical system is the radiating electron? Lecture Notes in Mathematics, 905:90–98, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092429/.

Barbanti:1983:LVS

[Bar83a] L. Barbanti. Linear Volterra–Stieltjes integral equations and control. Lecture Notes in Mathematics, 1017:67–72, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103236/.

Barbu:1983:OFC

[Bar83b] Viorel Barbu. Optimal feedback controls for semilinear parabolic equations. Lecture Notes in Mathematics, 979:43–70, 1983. CO-DEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066248/.

Bardi:1983:ESA

[Bar83c] Martino Bardi. Exchange of stability along a branch of periodic solutions of a single specie model. Lecture Notes in Mathematics, 1017:73–82, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103237/.

Bareiss:1983:LSS

[Bar83d] Erwin H. Bareiss. Least squares solution of weighted linear systems by G-transformations. Lecture Notes in Mathematics, 1005: 1–15, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112522/.

Bartels:1984:SIC

[Bar84] Richard H. Bartels. Splines in interactive computer graphics. Lecture Notes in Mathematics, 1066:1–29, 1984. CODEN LN-MAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099515/.

Basdevant:1980:PSS

[Bas80] C. Basdevant. Parameterization of subgrid-scale motion in numerical simulation of 2-dimensional Navier–Stokes equation at high Reynolds number. Lecture Notes in Mathematics, 771:43– 58, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086900/.

Bass:1981:LCA

[Bas81] Hyman Bass. Lenstra’s calculation of G_{O}(Rπ), and applications to Morse–Smale diffeomorphisms. Lecture Notes in Mathematics, 882:287–318, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092501/.

Bass:1982:ASG

[Bas82] Hyman Bass. Automorphismes de schémas et de groupes de type fini. (French) []. Lecture Notes in Mathematics, 924:310–321, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092938/.

Bass:1984:MPC

[Bas84] Richard F. Bass. Markov processes and convex minorants. Lecture Notes in Mathematics, 1059:29–41, 1984. CODEN LN-MAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100029/.

Batbedat:1981:DGD

[Bat81] A. Batbedat. γ-demi-groupes, demi-modules, produits demidirects. Lecture Notes in Mathematics, 855:1–18, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089804/.

Bautista:1980:SAR

[Bau80] R. Bautista. Sections in Auslander–Reiten quivers. Lecture Notes in Mathematics, 832:74–96, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088459/.

Bauer:1981:UWC

[Bau81] Friedrich W. Bauer. Under what conditions are shape homology Ē^{∗} and Steenrod homology ^{s}E^{∗} isomorphic? Lecture Notes in Mathematics, 870:186–214, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089716/.

Bauer:1984:ASS

[Bau84a] Friedrich W. Bauer. The advantages of strong shape theory. Lecture Notes in Mathematics, 1060:119–128, 1984. CODEN LN-MAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099928/.

Baumslag:1984:AIP

[Bau84b] Gilbert Baumslag. Algorithmically insoluble problems about finitely presented solvable groups, lie, and associative algebras. Lecture Notes in Mathematics, 1098:1–14, 1984. CODEN LN-MAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099656/.

Baxley:1981:NTP

[Bax81] John V. Baxley. Nonlinear two-point boundary value problems. Lecture Notes in Mathematics, 846:46–54, 1981. CODEN LN-MAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089823/.

Baxley:1983:NSO

[Bax83] John V. Baxley. Nonlinear second order boundary value problems: an existence-uniqueness theorem of S. N. Bernstein. Lecture Notes in Mathematics, 1032:9–16, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076790/.

Bayart:1980:FSF

[Bay80] Marc Bayart. Factorialité et séries formelles irreductibles. I. (French) []. Lecture Notes in Mathematics, 795:402–433, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090130/.

Bayart:1981:FSF

[Bay81a] Marc Bayart. Factorialité et séries formelles irreductibles II. (French) []. Lecture Notes in Mathematics, 867:174–213, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090387/.

Bayoumi:1981:LPR

[Bay81b] Aboubakr Bayoumi. The Levi problem and the radius of convergence of holomorphic functions on metric vector spaces. Lecture Notes in Mathematics, 843:9–32, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089267/.

Brenner:1980:GBG

[BB80] Sheila Brenner and M. C. R. Butler. Generalizations of the Bernstein–Gelfand–Ponomarev reflection functors. Lecture Notes in Mathematics, 832:103–169, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088461/.

Bautista:1981:NTM

[BB81a] Raymundo Bautista and Sheila Brenner. On the number of terms in the middle of an almost split sequence. Lecture Notes in Mathematics, 903:1–8, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092980/.

Benabdallah:1981:HIG

[BB81b] K. Benabdallah and A. Birtz. Hyper-indecomposable groups. Lecture Notes in Mathematics, 874:70–75, 1981. CODEN LN-MAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090525/.

Berard-Bergery:1981:CSV

[BB81c] Lionel Bérard-Bergery. La courbure scalaire des variétés riemanniennes. (French) []. Lecture Notes in Mathematics, 842:225– 245, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089937/.

Bergery:1981:LRS

[BB81d] L. Bérard Bergery and J.-P. Bourguignon. Laplacian and Riemannian submersions with totally geodesic fibres. Lecture Notes in Mathematics, 838:30–35, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088839/.

Blum:1981:RSC

[BB81e] Julius R. Blum and Russell A. Boyles. Random sampling from a continuous parameter stochastic process. Lecture Notes in Mathematics, 861:15–24, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097308/.

Barrow:1982:BPT

[BB82a] D. L. Barrow and P. W. Bates. Bifurcation of periodic travelling waves for a reaction-diffusion system. Lecture Notes in Mathematics, 964:69–76, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064988/.

Beutelspacher:1982:GSN

[BB82b] Albrecht Beutelspacher and Walter Brestovansky. Generalized Schur numbers. Lecture Notes in Mathematics, 969:30–38, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062985/.

Binding:1983:DRD

[BB83a] Paul Binding and Patrick J. Browne. A definiteness result for determinantal operators. Lecture Notes in Mathematics, 1032: 17–30, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076791/.

Boboc:1983:PES

[BB83b] N. Boboc and Gh. Bucur. Perturbations in excessive structures. Lecture Notes in Mathematics, 1014:155–187, 1983. CODEN LN-MAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072077/.

Bedford:1984:HCB

[BB84] Eric Bedford and Steve Bell. Holomorphic correspondences of bounded domains in C^{n}. Lecture Notes in Mathematics, 1094: 1–14, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099149/.

Boboc:1980:AC

[BBC80] N. Boboc, Gh. Bucur, and A. Cornea. Autodual H-cones. Lecture Notes in Mathematics, 787:64–77, 1980. CODEN LN-MAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086325/.

Bachar:1983:RBA

[BBC^{+}83] John M. Bachar, William G. Bade, Philip C. Curtis, H. Garth Dales, and Marc P. Thomas, editors. Radical Banach Algebras and Automatic Continuity: Proceedings of a Conference Held at California State University, Long Beach, July 17–31, 1981, volume 975 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 975. URL http:/

/link.springer.com/book/10.1007/BFb0064545;   http://
www.springerlink.com/content/978-3-540-39454-9.

Boboc:1981:C

[BBCH81a] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. H-cones. Lecture Notes in Mathematics, 853:35–67, 1981. CO-DEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090451/.

Boboc:1981:CF

[BBCH81b] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. H-cones of functions. Lecture Notes in Mathematics, 853:68– 95, 1981. CODEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090452/.

Boboc:1981:CDS

[BBCH81c] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. H-cones on Dirichlet spaces. Lecture Notes in Mathematics, 853: 245–277, 1981. CODEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090456/.

Boboc:1981:I

[BBCH81d] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. Introduction. Lecture Notes in Mathematics, 853:1–5, 1981. CO-DEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090448/.

Boboc:1981:MPA

[BBCH81e] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. Markov processes associated with a standard H-cone of functions. Lecture Notes in Mathematics, 853:230–244, 1981. CODEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090455/.

Boboc:1981:OCP

[BBCH81f] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. Order and Convexity in Potential Theory: H-Cones, volume 853 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 286 pp. LCCN QA3 .L28 no. 853. URL http:

//link.springer.com/book/10.1007/BFb0090447;   http:/
/www.springerlink.com/content/978-3-540-38618-6.

Boboc:1981:PTS

[BBCH81g] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. Potential theory on standard H-cones of functions. Lecture Notes in Mathematics, 853:152–229, 1981. CODEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090454/.

Boboc:1981:R

[BBCH81h] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. Resolvents. Lecture Notes in Mathematics, 853:7–34, 1981. CO-DEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090450/.

Boboc:1981:SC

[BBCH81i] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. Standard H-cones. Lecture Notes in Mathematics, 853:96–151, 1981. CODEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090453/.

Boboc:1981:TN

[BBCH81j] Nicu Boboc, Gheorghe Bucur, Aurel Cornea, and Herbert Höllein. Terminology and notations. Lecture Notes in Mathematics, 853: 6, 1981. CODEN LNMAA2. ISBN 3-540-10692-8 (print), 3-540-38618-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090449.

Berthelot:1982:CFE

[BBM82a] Pierre Berthelot, Lawrence Breen, and William Messing. Calculs de faisceaux d’extensions. (French) []. Lecture Notes in Mathematics, 930:53–109, 1982. CODEN LNMAA2. ISBN 3-540-11556-0 (print), 3-540-39227-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093028/.

Berthelot:1982:CAT

[BBM82b] Pierre Berthelot, Lawrence Breen, and William Messing. Comparaison avec la théorie de Dieudonné classique. (French) []. Lecture Notes in Mathematics, 930:151–204, 1982. CODEN LN-MAA2. ISBN 3-540-11556-0 (print), 3-540-39227-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093030/.

Berthelot:1982:CGF

[BBM82c] Pierre Berthelot, Lawrence Breen, and William Messing. Conventions générales. (French) []. Lecture Notes in Mathematics, 930:1–7, 1982. CODEN LNMAA2. ISBN 3-540-11556-0 (print), 3-540-39227-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093026/.

Berthelot:1982:CDF

[BBM82d] Pierre Berthelot, Lawrence Breen, and William Messing. Cristaux de Dieudonné. (French) [Dieudonné crystals]. Lecture Notes in Mathematics, 930:110–150, 1982. CODEN LNMAA2. ISBN 3-540-11556-0 (print), 3-540-39227-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093029/.

Berthelot:1982:EFA

[BBM82e] Pierre Berthelot, Lawrence Breen, and William Messing. Extensions de faisceaux abeliens sur le site cristallin. (French) []. Lecture Notes in Mathematics, 930:8–52, 1982. CODEN LNMAA2. ISBN 3-540-11556-0 (print), 3-540-39227-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093027/.

Berthelot:1982:TDF

[BBM82f] Pierre Berthelot, Lawrence Breen, and William Messing. Théorèmes de dualité. (French) []. Lecture Notes in Mathematics, 930:205–257, 1982. CODEN LNMAA2. ISBN 3-540-11556-0 (print), 3-540-39227-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093031/.

Berthelot:1982:TDC

[BBM82g] Pierre Berthelot, Lawrence Breen, and William Messing. Théorie de Dieudonné Cristalline II. (French) [Crystalline Dieudonné theory], volume 930 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11556-0 (print), 3-540-39227-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 261 pp. LCCN QA3 .L28 no. 930. URL http:/

/link.springer.com/book/10.1007/BFb0093025;   http://
www.springerlink.com/content/978-3-540-39227-9.

Brancovan:1982:GDP

[BBP82] M. Brancovan, F. Bronner, and P. Priouret. Grandes déviations pour certains systèmes différentiels aléatoires. (French) []. Lecture Notes in Mathematics, 920:159–183, 1982. CODEN LN-MAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092779/.

Bialynicki-Birula:1982:CQA

[BBŚ82] Andrzej Białynicki-Birula and Joanna Świecicka. Complete quotients by algebraic torus actions. Lecture Notes in Mathematics, 956:10–22, 1982. CODEN LNMAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101505/.

Berline:1981:QRC

[BC81] C. Berline and G. Cherlin. QE rings in characteristic p. Lecture Notes in Mathematics, 859:16–31, 1981. CODEN LN-MAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090936/.

Bartels:1982:AND

[BC82a] Richard H. Bartels and Andrew R. Conn. An approach to nonlinear l_{1} data fitting. Lecture Notes in Mathematics, 909:48–58, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092959/.

Berruyer:1982:CVD

[BC82b] Jacques Berruyer and René Carmona. Convergence vague et decomposition de Riesz dans des groupes non localement compacts. (French) []. Lecture Notes in Mathematics, 906:27–52, 1982. CO-DEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093260/.

Brooks:1983:CTT

[BC83] J. K. Brooks and R. V. Chacon. Convergence theorems in the theory of diffusions. Lecture Notes in Mathematics, 1033:79–93, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099848/.

Benci:1982:PSC

[BCF82] V. Benci, A. Capozzi, and D. Fortunato. Periodic solutions of a class of Hamiltonian systems. Lecture Notes in Mathematics, 964:86–94, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064990/.

Benci:1983:AQH

[BCF83] V. Benci, A. Capozzi, and D. Fortunato. On asymptotically quadratic Hamiltonian systems. Lecture Notes in Mathematics, 1017:83–92, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103238/.

Bridges:1981:LMC

[BCJ^{+}81] Douglas S. Bridges, Allan Calder, William Julian, Ray Mines, and Fred Richman. Locating metric complements in R^{n}. Lecture Notes in Mathematics, 873:241–249, 1981. CODEN LN-MAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090737/.

Brugarino:1983:SEE

[BCP83] T. Brugarino, A. Canino, and P. Pantano. Some evolution equations arising in physics. Lecture Notes in Mathematics, 1017: 107–114, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103241/.

Bautista:1981:AFR

[BCS81] R. Bautista, L. Colavita, and L. Salmerón. On adjoint functors in representation theory. Lecture Notes in Mathematics, 903:9– 25, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092981/.

Balacheff:1980:NAP

[BD80a] Serge Balacheff and Ghislain Dupont. Normalité asymptotique des processus empiriques tronqués et des processus de rang (cas multidimensionnel mélangeant). (French) []. Lecture Notes in Mathematics, 821:19–45, 1980. CODEN LNMAA2. ISBN 3-540-10239-6 (print), 3-540-38318-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097423/.

Bit-David:1980:NEI

[BD80b] J. Bit-David. Normalizing extensions II. Lecture Notes in Mathematics, 825:6–9, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089116/.

Batt:1983:WCC

[BD83] Jürgen Batt and Nicolae Dinculeanu. On the weak compactness criteria of Kolmogorov–Tamarkin and M. Riesz type in the space of Bochner integrable functions over a locally compact group. Lecture Notes in Mathematics, 1033:43–58, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099844/.

Barnsley:1984:RAF

[BD84] Michael F. Barnsley and Stephen G. Demko. Rational approximation of fractals. Lecture Notes in Mathematics, 1105:73–88, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072400/.

Becker:1980:FRE

[BDL80] J. Becker, J. Denef, and L. Lipshitz. Further remarks on the elementary theory of formal power series rings. Lecture Notes in Mathematics, 834:1–9, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090156/.

Belley:1983:MTA

[BDM83] Jean-Marc Belley, Jacques Dubois, and Pedro Morales, editors. Measure Theory and its Applications: Proceedings of a Conference held at Sherbrooke, Québec, Canada, June 7–18, 1982, volume 1033 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1033. URL http:/

/link.springer.com/book/10.1007/BFb0099840;   http://
www.springerlink.com/content/978-3-540-38690-2.

Bit-David:1980:NE

[BDR80] J. Bit-David and J. C. Robson. Normalizing extensions I. Lecture Notes in Mathematics, 825:1–5, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089115/.

Burton:1980:ELT

[BE80] Robert Burton and Robert W. Easton. Ergodicity of linked twist maps. Lecture Notes in Mathematics, 819:35–49, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086978/.

Baird:1981:CLH

[BE81] P. Baird and J. Eells. A conservation law for harmonic maps. Lecture Notes in Mathematics, 894:1–25, 1981. CODEN LN-MAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096222/.

Bochnak:1982:INF

[BE82] J. Bochnak and Gustave Efroymson. An introduction to Nash functions. Lecture Notes in Mathematics, 959:41–54, 1982. CO-DEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062249/.

Ballico:1983:DPC

[BE83a] E. Ballico and Ph. Ellia. On degeneration of projective curves. Lecture Notes in Mathematics, 997:1–15, 1983. CODEN LN-MAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061634/.

Bennewitz:1983:SOL

[BE83b] C. Bennewitz and W. N. Everitt. On second-order left-definite boundary value problems. Lecture Notes in Mathematics, 1032: 31–67, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076792/.

Beauville:1980:GTD

[Bea80] Arnaud Beauville. Géométrie des tissus [d’après S. S. Chern et P. A. Griffiths]. (French) []. Lecture Notes in Mathematics, 770: 103–119, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096238/.

Beauzamy:1981:SEI

[Bea81] B. Beauzamy. Sous-espaces invariants dans les espaces de Banach. (French) []. Lecture Notes in Mathematics, 842:95–111, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089930/.

Beachy:1982:MTR

[Bea82a] John A. Beachy. Maximal torsion radicals over rings with finite reduced bank. Lecture Notes in Mathematics, 951:41–46, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067323/.

Beauville:1982:SDD

[Bea82b] Arnaud Beauville. Les singularités du diviseur because de la jacobienne intermèdiaire de l’hypersurface cubique dans IP^{4}. (French) []. Lecture Notes in Mathematics, 947:190–208, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093588/.

Beauville:1983:QRT

[Bea83a] A. Beauville. Quelques remarques sur la transformation de Fourier dans l’anneau de Chow d’une variété abélienne. (French) []. Lecture Notes in Mathematics, 1016:238–260, 1983. CODEN LN-MAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099965/.

Beauville:1983:VRU

[Bea83b] A. Beauville. Variétés rationnelles et unirationnelles. (French) []. Lecture Notes in Mathematics, 997:16–33, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061635/.

Beaver:1983:RDF

[Bea83c] Olga R. Beaver. Regularity and decomposability of finitely additive functions on a quantum logic. Lecture Notes in Mathematics, 1033:59–67, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099845/.

Becker:1980:PSS

[Bec80] Ronald I. Becker. Periodic solutions of semilinear functional differential equations in a Hilbert space. Lecture Notes in Mathematics, 799:23–44, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089307/.

Beck:1981:PBS

[Bec81] Anatole Beck, editor. Probability in Banach Spaces III: Proceedings of the Third International Conference on Probability in Banach Spaces Held at Tufts University, Medford, USA, August 4–16, 1980, volume 860 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 0-387-10822-X, 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 860.

URL http://link.springer.com/book/10.1007/BFb0090603;
 http://www.springerlink.com/content/978-3-540-38710-
7.

Becker:1982:RHR

[Bec82a] Eberhard Becker. The real holomorphy ring and sums of 2nth powers. Lecture Notes in Mathematics, 959:139–181, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062253/.

Becker:1982:VRP

[Bec82b] Eberhard Becker. Valuations and real places in the theory of formally real fields. Lecture Notes in Mathematics, 959:1–40, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062248/.

Beck:1984:NRT

[Bec84] J. Beck. New results in the theory of irregularities of point distributions. Lecture Notes in Mathematics, 1068:1–16, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099437/.

Bedford:1982:OCS

[Bed82] Eric Bedford. The operator (dd^{c})^{n} on complex spaces. Lecture Notes in Mathematics, 919:294–323, 1982. CODEN LN-MAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097054/.

Beeson:1981:FCM

[Bee81] M. J. Beeson. Formalizing constructive mathematics: Why and how? Lecture Notes in Mathematics, 873:146–190, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090733/.

Begehr:1983:BVP

[Beg83] Heinrich Begehr. Boundary value problems for systems with Cauchy–Riemannian main part. Lecture Notes in Mathematics, 1014:265–279, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072085/.

Behrends:1981:STP

[Beh81] Ehrhard Behrends. M-structure in tensor products of Banach spaces. Lecture Notes in Mathematics, 843:41–54, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089269/.

Behncke:1982:DEA

[Beh82] Horst Behncke. The Dirac equation with an anomalous magnetic moment II. Lecture Notes in Mathematics, 964:77–85, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064989/.

Belley:1981:MTA

[Bel81a] J.-M. Belley. A measure theoretic approach to fixed points in ergodic theory. Lecture Notes in Mathematics, 886:33–43, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092174/.

Bellow:1981:MAR

[Bel81b] A. Bellow. Martingales, amarts and related stopping time techniques. Lecture Notes in Mathematics, 860:9–24, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090605/.

Belley:1983:CFE

[Bel83a] Jean-Marc Belley. Construction d’une fonction d’ensembles additive et invariante sur les groupes localement compacts abeliens. (French) []. Lecture Notes in Mathematics, 1033:68–73, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099846/.

Bellow:1983:BUS

[Bel83b] Alexandra Bellow. On ‘bad universal’ sequences in ergodic theory (II). Lecture Notes in Mathematics, 1033:74–78, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099847/.

Bemelmans:1980:SFP

[Bem80] J. Bemelmans. C^{O+α}-semigroups for flows past obstacles and for flows with capillary surfaces. Lecture Notes in Mathematics, 771:59–77, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086901/.

Benedetto:1980:EHA

[Ben80a] John J. Benedetto, editor. Euclidean Harmonic Analysis: Proceedings of Seminars Held at the University of Maryland, 1979, volume 779 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 779. URL http:/

/link.springer.com/book/10.1007/BFb0087664;   http://
www.springerlink.com/content/978-3-540-38602-5.

Benedetto:1980:I

[Ben80b] John J. Benedetto. Introduction. Lecture Notes in Mathematics, 779:1–4, 1980. CODEN LNMAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087665/.

Benz:1980:TOC

[Ben80c] Walter Benz. Ein Trennungsaxiom in der Orthogonalgeometrie und eine Charakterisierung der reellen Ebene. (German) []. Lecture Notes in Mathematics, 792:14–19, 1980. CODEN LN-MAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088655/.

Benedetti:1982:CTS

[Ben82a] Riccardo Benedetti. On certain topological spaces admitting strongly coherent real algebraic structure. Lecture Notes in Mathematics, 959:182–197, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062254/.

Bensoussan:1982:LSC

[Ben82b] A. Bensoussan. Lectures on stochastic control. Lecture Notes in Mathematics, 972:1–62, 1982. CODEN LNMAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064859/.

Bentley:1982:NHR

[Ben82c] H. L. Bentley. A note on the homology of regular nearness spaces. Lecture Notes in Mathematics, 962:1–7, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066879/.

Bensoussan:1983:PSP

[Ben83] A. Bensoussan. On the production smoothing problem. Lecture Notes in Mathematics, 979:71–87, 1983. CODEN LN-MAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066249/.

Benninghofen:1984:SCG

[Ben84a] Benjamin Benninghofen. Superinfinitesimals and the calculus of the generalized Riemann integral. Lecture Notes in Mathematics, 1103:9–52, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099379/.

Benson:1984:MRT

[Ben84b] David J. Benson. Modular Representation Theory: New Trends and Methods, volume 1081 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13389-5 (print), 3-540-38940-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 231 pp. LCCN QA3 .L28 no.

1081. URL http://link.springer.com/book/10.1007/3-540-
38940-7;   http://www.springerlink.com/content/978-3-
540-38940-8.

Benson:1984:MGA

[Ben84c] David J. Benson. Modules for group algebras. Lecture Notes in Mathematics, 1081:23–172, 1984. CODEN LNMAA2. ISBN 3-540-13389-5 (print), 3-540-38940-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-38940-7_2/.

Benson:1984:RM

[Ben84d] David J. Benson. Rings and modules. Lecture Notes in Mathematics, 1081:1–22, 1984. CODEN LNMAA2. ISBN 3-540-13389-5 (print), 3-540-38940-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-38940-7_1/.

Bentzen:1984:NET

[Ben84e] Steffen Bentzen. The non-existence of tensorproducts for free group actions on spheres. Lecture Notes in Mathematics, 1051: 8–24, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075561/.

Bateman:1981:AMD

[BEPS81] Paul T. Bateman, Paul Erdős, Carl Pomerance, and E. G. Straus. The arithmetic mean of the divisors of an integer. Lecture Notes in Mathematics, 899:197–220, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096462/.

Beresina:1980:ATS

[Ber80a] L. Ja. Beresina. Applications of the theory of surfaces to the theory of graphs. Lecture Notes in Mathematics, 792:20–23, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088656/.

Berg:1980:QRC

[Ber80b] Christian Berg. Quelques remarques sur le cône de Stieltjes. (French) []. Lecture Notes in Mathematics, 814:70–79, 1980. CO-DEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094147/.

Berline:1980:EQN

[Ber80c] Chantal Berline. Elimination of quantifiers for non semi-simple rings of characteristic p. Lecture Notes in Mathematics, 834:10– 19, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090157/.

Berrick:1980:PSI

[Ber80d] A. J. Berrick. Projective space immersions, bilinear maps and stable homotopy groups of spheres. Lecture Notes in Mathematics, 788:1–22, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099235/.

Bertrand:1980:TRP

[Ber80e] Daniel Bertrand. Travaux récents sur les points singuliers des équations différentielles linéaires. (French) []. Lecture Notes in Mathematics, 770:228–243, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096245/.

Berge:1981:DG

[Ber81a] Claude Berge. Diperfect graphs. Lecture Notes in Mathematics, 885:1–8, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092250/.

Bergstrom:1981:RWL

[Ber81b] Harald Bergström. Reduction of weak limit problems by transformations. Lecture Notes in Mathematics, 861:1–9, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097306/.

Berline:1981:IAP

[Ber81c] Chantal Berline. Idéaux des anneaux de Peano (d’après Cherlin). (French) []. Lecture Notes in Mathematics, 890:32–43, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095656/.

Berthelot:1981:TDP

[Ber81d] P. Berthelot. Le théorème de dualité plate pour les surfaces (d’après J. S. Milne). (French) []. Lecture Notes in Mathematics, 868:203–237, 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090649/.

Berger:1982:CSN

[Ber82a] Mel S. Berger. Classical solutions in nonlinear Euclidean field theory and complete integrability. Lecture Notes in Mathematics, 925:123–133, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093504/.

Bers:1982:TF

[Ber82b] Lipman Bers. On trace formula. Lecture Notes in Mathematics, 925:209–214, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093510/.

Bergman:1983:ACD

[Ber83a] Clifford Bergman. The amalgamation class of a discriminator variety is finitely axiomatizable. Lecture Notes in Mathematics, 1004:1–9, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063427/.

Berman:1983:FSE

[Ber83b] Joel Berman. Free spectra of 3-element algebras. Lecture Notes in Mathematics, 1004:10–53, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063428/.

Bernstein:1983:IDC

[Ber83c] Joseph N. Bernstein. p-invariant distributions on GL(N) and the classification of unitary representations of GL(N) (non-Archimedean case). Lecture Notes in Mathematics, 1041:50– 102, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073145/.

Berg:1984:FDN

[Ber84a] Christian Berg. Fonctions definies negatives et majoration de Schur. (French) []. Lecture Notes in Mathematics, 1096:69– 89, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100107/.

Berge:1984:MNA

[Ber84b] C. Berge. On the minimum number of arcs to reverse to make a graph strongly connected. Lecture Notes in Mathematics, 1073: 31–34, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073103/.

Bergh:1984:NLC

[Ber84c] Jöran Bergh. A non-linear complex interpolation result. Lecture Notes in Mathematics, 1070:45–47, 1984. CODEN LN-MAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099091/.

Bernard:1984:CCR

[Ber84d] D. Bernard. Congruence, contact et repères de Frenet. (French) []. Lecture Notes in Mathematics, 1045:21–35, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072162/.

Bernstein:1984:DRV

[Ber84e] Barry Bernstein. Discontinuous response in viscoelastic bars. Lecture Notes in Mathematics, 1063:64–75, 1984. CODEN LN-MAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072153/.

Bertrand:1984:PAF

[Ber84f] Daniel Bertrand. Propriétés arithmétiques de fonctions thêta à plusieurs variables. (French) []. Lecture Notes in Mathematics, 1068:17–22, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099438/.

Besson:1981:GLA

[Bes81] G. Besson. Groupe de Lie p-adique, immeuble et cohomologie. (French) []. Lecture Notes in Mathematics, 867:130–140, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090383/.

Beth:1982:SAC

[Bet82] Thomas Beth. Some aspects of coding theory between probability, algebra, combinatorics and complexity theory. Lecture Notes in Mathematics, 969:12–29, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062984/.

Betina:1983:IOD

[Bet83] Kamel Betina. Sur l’indice des opérateurs différentiels ordinaires. (French) []. Lecture Notes in Mathematics, 1015:1–49, 1983. CO-DEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071348/.

Beukers:1981:PAN

[Beu81] F. Beukers. Pade-approximations in number theory. Lecture Notes in Mathematics, 888:90–99, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095578/.

Bickel:1981:QAS

[BEY81] Peter J. Bickel, Nicole El Karoui, and Marc Yor. Quelques aspects de la statistique robuste. (French) []. Lecture Notes in Mathematics, 876:xi–72, 1981. CODEN LNMAA2. ISBN 3-540-10860-2 (print), 3-540-38774-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097498/.

Beyn:1983:NAS

[Bey83] W.-J. Beyn. Numerical analysis of singularities in a diffusion reaction model. Lecture Notes in Mathematics, 1017:93–100, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103239/.

Bair:1980:ATR

[BF80a] Jacques Bair and René Fourneau. Applications de la théorie des représentations. (French) []. Lecture Notes in Mathematics, 802: 161–191, 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089015/.

Bair:1980:APS

[BF80b] Jacques Bair and René Fourneau. Applications des polyèdres à la séparation. (French) [Applications of polyhedra to separation]. Lecture Notes in Mathematics, 802:192–212, 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089016/.

Bair:1980:EGE

[BF80c] Jacques Bair and René Fourneau. Étude Géometrique des Espaces Vectoriels II. Polyèdres et Polytopes Convexes. (French)

[Geometric study of vector spaces II. Convex polyhedra and polytopes], volume 802 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 283 pp. LCCN QA3 .L28 no. 802. URL http:

//link.springer.com/book/10.1007/BFb0089008;   http:/
/www.springerlink.com/content/978-3-540-39293-4.

Bair:1980:FFC

[BF80d] Jacques Bair and René Fourneau. Faces et facettes des convexes. (French) []. Lecture Notes in Mathematics, 802:22–51, 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089011/.

Bair:1980:PCF

[BF80e] Jacques Bair and René Fourneau. Les polyèdres convexes. (French) []. Lecture Notes in Mathematics, 802:52–98, 1980. CO-DEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089012/.

Bair:1980:PF

[BF80f] Jacques Bair and René Fourneau. Les polytopes. (French) []. Lecture Notes in Mathematics, 802:99–133, 1980. CODEN LN-MAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089013/.

Bair:1980:NTF

[BF80g] Jacques Bair and René Fourneau. Notations et terminologie. (French) []. Lecture Notes in Mathematics, 802:1, 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089009.

Bair:1980:RPF

[BF80h] Jacques Bair and René Fourneau. Représentations de polyèdres. (French) []. Lecture Notes in Mathematics, 802:134–160, 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089014/.

Bair:1980:RSC

[BF80i] Jacques Bair and René Fourneau. Retour aux simplexes de Choquet. (French) []. Lecture Notes in Mathematics, 802:213– 235, 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089017/.

Bair:1980:SDE

[BF80j] Jacques Bair and René Fourneau. Séparation de deux ensembles convexes. (French) []. Lecture Notes in Mathematics, 802: 2–21, 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089010/.

Bair:1980:SIL

[BF80k] Jacques Bair and René Fourneau. Systèmes d’inéquations linéaires. (French) []. Lecture Notes in Mathematics, 802:236– 250, 1980. CODEN LNMAA2. ISBN 3-540-09993-X (print), 3-540-39293-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089018/.

Bichteler:1982:SVM

[BF82] Klaus Bichteler and David Fonken. A simple version of the Malliavin calculus in dimension one. Lecture Notes in Mathematics, 939:6–12, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096254/.

Beltrametti:1983:CBN

[BF83a] M. Beltrametti and P. Francia. Conic bundles on non-rational surfaces. Lecture Notes in Mathematics, 997:34–89, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061636/.

Bushnell:1983:ALD

[BF83b] Colin J. Bushnell and Albrecht Fröhlich. Arithmetic of local division algebras. Lecture Notes in Mathematics, 987:1–7, 1983.

CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066414/.

Bushnell:1983:BC

[BF83c] Colin J. Bushnell and Albrecht Fröhlich. The basic correspondence. Lecture Notes in Mathematics, 987:53–67, 1983. CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066418/.

Bushnell:1983:BIS

[BF83d] Colin J. Bushnell and Albrecht Fröhlich. The basic inductive step. Lecture Notes in Mathematics, 987:68–86, 1983. CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066419/.

Bushnell:1983:FE

[BF83e] Colin J. Bushnell and Albrecht Fröhlich. Functional equation. Lecture Notes in Mathematics, 987:34–41, 1983. CODEN LN-MAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066416/.

Bushnell:1983:GSA

[BF83f] Colin J. Bushnell and Albrecht Fröhlich. Gauss Sums and p-adic Division Algebras, volume 987 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 187 pp. LCCN QA3 .L28 no. 987.

URL http://link.springer.com/book/10.1007/BFb0066413;
 http://www.springerlink.com/content/978-3-540-39618-
5.

Bushnell:1983:GIP

[BF83g] Colin J. Bushnell and Albrecht Fröhlich. The general inductive process. Lecture Notes in Mathematics, 987:87–98, 1983. CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066420/.

Bushnell:1983:ICG

[BF83h] Colin J. Bushnell and Albrecht Fröhlich. Induction constants for Galois Gauss sums. Lecture Notes in Mathematics, 987:146– 155, 1983. CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066423/.

Bushnell:1983:IGS

[BF83i] Colin J. Bushnell and Albrecht Fröhlich. Introduction to Gauss sums. Lecture Notes in Mathematics, 987:8–33, 1983. CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066415/.

Bushnell:1983:MC

[BF83j] Colin J. Bushnell and Albrecht Fröhlich. Modified correspondences. Lecture Notes in Mathematics, 987:166–182, 1983. CO-DEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066425/.

Bushnell:1983:ODR

[BF83k] Colin J. Bushnell and Albrecht Fröhlich. One-dimensional representations. Lecture Notes in Mathematics, 987:42–52, 1983. CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066417/.

Bushnell:1983:RCG

[BF83l] Colin J. Bushnell and Albrecht Fröhlich. Representations of certain group extensions. Lecture Notes in Mathematics, 987:99– 121, 1983. CODEN LNMAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066421/.

Bushnell:1983:SR

[BF83m] Colin J. Bushnell and Albrecht Fröhlich. Synthesis of results. Lecture Notes in Mathematics, 987:156–165, 1983. CODEN LN-MAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066424/.

Bushnell:1983:TC

[BF83n] Colin J. Bushnell and Albrecht Fröhlich. Trace computations. Lecture Notes in Mathematics, 987:122–145, 1983. CODEN LN-MAA2. ISBN 3-540-12290-7 (print), 3-540-39618-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066422/.

Bensoussan:1984:NPE

[BF84] A. Bensoussan and J. Frehse. Nash point equilibria for variational integrals. Lecture Notes in Mathematics, 1107:28–62, 1984. CODEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101492/.

Berg:1980:PTC

[BFF80] Christian Berg, Gunnar Forst, and Bent Fuglede, editors. Potential Theory Copenhagen 1979: Proceedings of a Colloquium Held in Copenhagen, May 14–18, 1979, volume 787 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 787.

URL http://link.springer.com/book/10.1007/BFb0086319;
 http://www.springerlink.com/content/978-3-540-39183-
8.

Billington:1982:PSC

[BFG82] Elizabeth J. Billington and James M. Fitz-Gerald. Partitions and the sequence counting problem in economic lot scheduling. Lecture Notes in Mathematics, 952:197–224, 1982. CODEN LN-MAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061980/.

Buchholz:1981:IID

[BFPS81] Wilfried Buchholz, Solomon Feferman, Wolfram Pohlers, and Wilfried Sieg. Iterated Inductive Definitions and Subsystems of Analysis: Recent Proof-Theoretical Studies, volume 897 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11170-0 (print), 3-540-38649-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 383 pp. LCCN QA3 .L28 no. 897. URL http:/

/link.springer.com/book/10.1007/BFb0091894;   http://
www.springerlink.com/content/978-3-540-38649-0.

Ballotti:1984:WTS

[BG84a] M. E. Ballotti and J. A. Goldstein. Wiener’s theorem and semigroups of operators. Lecture Notes in Mathematics, 1076:16– 22, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072761/.

Borisovich:1984:TTF

[BG84b] Yu. G. Borisovich and Yu. E. Gliklikh. Topological theory of fixed points on infinite-dimensional manifolds. Lecture Notes in Mathematics, 1108:1–23, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099549/.

Borowski:1984:RRF

[BG84c] D. Borowski and T. Gołebiewski. Remarks on random functional spaces. Lecture Notes in Mathematics, 1080:1–8, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099782/.

Bristeau:1980:FEA

[BGM^{+}80] M. O. Bristeau, R. Glowinski, B. Mantel, J. Periaux, P. Perrier, and O. Pironneau. A finite element approximation of Navier– Stokes equations for incompressible viscous fluids. iterative methods of solution. Lecture Notes in Mathematics, 771:78–128, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086902/.

Block:1980:PPT

[BGMY80] Louis Block, John Guckenheimer, Michał Misiurewicz, and Lai Sang Young. Periodic points and topological entropy of one dimensional maps. Lecture Notes in Mathematics, 819:18–34, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086977/.

Borisovich:1984:GAS

[BGV84] Yurii G. Borisovich, Yurii E. Gliklikh, and A. M. Vershik. Global Analysis — Studies and Applications I, volume 1108 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 301 pp. LCCN QA3 .L28 no. 1108. URL http:/

/link.springer.com/book/10.1007/BFb0099548;   http://
www.springerlink.com/content/978-3-540-39132-6.

Bliedtner:1980:BSB

[BH80a] J. Bliedtner and W. Hansen. Bases in standard balayage spaces. Lecture Notes in Mathematics, 787:55–63, 1980. CODEN LN-MAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086324/.

Boukricha:1980:CPH

[BH80b] A. Boukricha and W. Hansen. Characterization of perturbations of harmonic spaces. Lecture Notes in Mathematics, 787: 78–83, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086326/.

Banaschewski:1981:CLP

[BH81a] Bernhard Banaschewski and Rudolf-Eberhard Hoffmann, editors. Continuous Lattices: Proceedings of the Conference on Topological and Categorical Aspects of Continuous Lattices (Workshop IV) Held at the University of Bremen, Germany, November 9–11, 1979, volume 871 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 871. URL http:/

/link.springer.com/book/10.1007/BFb0089899;   http://
www.springerlink.com/content/978-3-540-38755-8.

Begehr:1981:NBV

[BH81b] Heinrich Begehr and George C. Hsiao. On nonlinear boundary value problems for an elliptic system in the plane. Lecture Notes in Mathematics, 846:55–63, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089824/.

Bentley:1982:CHC

[BH82a] H. L. Bentley and H. Herrlich. The coreflective hull of the contigual spaces in the category of merotopic spaces. Lecture Notes in Mathematics, 915:16–26, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092868/.

Brandenburg:1982:RCC

[BH82b] H. Brandenburg and M. Hušek. A remark on cartesian closedness. Lecture Notes in Mathematics, 962:33–38, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066883/.

Brown:1982:CCN

[BH82c] Ronald Brown and Philip J. Higgins. Crossed complexes and non-abelian extensions. Lecture Notes in Mathematics, 962:39– 50, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066884/.

Byczkowska:1982:SMS

[BH82d] H. Byczkowska and A. Hulanicki. On the support of the measures in a semigroup of probability measures on a locally compact group. Lecture Notes in Mathematics, 939:13–17, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096255/.

Benedetto:1983:WHS

[BH83] J. J. Benedetto and H. P. Heinig. Weighted Hardy spaces and the Laplace transform. Lecture Notes in Mathematics, 992:240– 277, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069163/.

Banghe:1984:TSP

[BH84] Li Banghe and Nathan Habegger. A two stage procedure for the classification of vector bundle monomorphisms with applications to the classification of immersions homotopic to a map. Lecture Notes in Mathematics, 1051:293–314, 1984. CODEN LN-MAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075573/.

Bru:1981:ADB

[BHL81] B. Bru, H. Heinich, and J.-C. Lootgieter. Autour de la dualité (H^{1}, BMO). (French) []. Lecture Notes in Mathematics, 850:259– 275, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088373/.

Bryan:1981:CES

[BHS81] Alan C. Bryan, Christopher R. Haines, and Allan E. G. Stuart. Complex extensions of a submanifold of solutions of the sine-Gordon equation. Lecture Notes in Mathematics, 846:74–81, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089826/.

Babiker:1984:LIP

[BHS84] A. G. A. G. Babiker, G. Heller, and W. Strauss. On a lifting invariance problem. Lecture Notes in Mathematics, 1089:79–85, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072604/.

Bianchini:1981:SHT

[Bia81] Mauro Bianchini. Silva-holomorphy types, Borel transforms and partial differential operators. Lecture Notes in Mathematics, 843: 55–92, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089270/.

Bierlein:1980:MEM

[Bie80] Dieter Bierlein. Measure extensions and measurable neighbours of a function. Lecture Notes in Mathematics, 794:1–23, 1980.

CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088207/.

Bierstedt:1981:ATL

[Bie81] Klaus-D. Bierstedt. The approximation-theoretic localization of Schwartz’s approximation property for weighted locally convex function spaces and some examples. Lecture Notes in Mathematics, 843:93–149, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089271/.

Bierlein:1982:SNR

[Bie82] Dieter Bierlein. Some new results on measure extension. Lecture Notes in Mathematics, 945:15–23, 1982. CODEN LN-MAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096660/.

Billington:1980:DSU

[Bil80] David Billington. Degree sequences uniquely realisable by hypergraphs. Lecture Notes in Mathematics, 829:59–68, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088900/.

Billington:1981:CSG

[Bil81] David Billington. Connected subgraphs of the graph of multigraphic realisations of a degree sequence. Lecture Notes in Mathematics, 884:125–135, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091814/.

Billington:1982:GHR

[Bil82a] David Billington. The graph of hypergraphic realisations of denumerable multisets of degrees. Lecture Notes in Mathematics, 952:150–181, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061978/.

Billington:1982:CSI

[Bil82b] Elizabeth J. Billington. Construction of some irreducible designs. Lecture Notes in Mathematics, 952:182–196, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061979/.

Binz:1983:EEE

[Bin83] E. Binz. Einstein’s evolution equation for the vacuum formulated on a space of differentials of immersions. Lecture Notes in Mathematics, 1037:2–37, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073168/.

Birtz:1983:DDPa

[Bir83a] Alain Birtz. Sur des decompositions directes paradoxales de groupes abeliens sans torsion. (French) []. Lecture Notes in Mathematics, 1006:358–361, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21560-9_19/.

Birtz:1983:DDPb

[Bir83b] Alain Birtz. Sur des decompositions directs paradoxales de groupes abeliens sans torsion. (French) []. Lecture Notes in Mathematics, 1006:358–361, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103714/.

Bismut:1981:AFA

[Bis81a] Jean-Michel Bismut. Approximation des flots associés à une diffusion de Itô sur une variété riemannienne à courbure negative. (French) []. Lecture Notes in Mathematics, 866:474–505, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088604/.

Bismut:1981:CDSb

[Bis81b] Jean-Michel Bismut. Calcul différentiel stochastique en coordonnées covariantes. (French) []. Lecture Notes in Mathematics, 866:424–438, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088602/.

Bismut:1981:CDSa

[Bis81c] Jean-Michel Bismut. Calcul différentiel stochastique. (French) []. Lecture Notes in Mathematics, 866:157–220, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088597/.

Bismut:1981:CVC

[Bis81d] Jean-Michel Bismut. Calcul variationnel en coordonnées covariantes sur des semi-martingales de Itô. (French) []. Lecture Notes in Mathematics, 866:506–553, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088605/.

Bismut:1981:DVS

[Bis81e] Jean-Michel Bismut. Diffusions sur les variétés symplectiques. (French) []. Lecture Notes in Mathematics, 866:221–241, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088598/.

Bismut:1981:FAD

[Bis81f] Jean-Michel Bismut. Flots associés à des diffusions. (French) []. Lecture Notes in Mathematics, 866:36–102, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088594/.

Bismut:1981:FGD

[Bis81g]  Jean-Michel Bismut.  Formulation géométrique du calcul

différentiel de Itô. (French) []. Lecture Notes in Mathematics, 866: 380–423, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088601/.

Bismut:1981:ISF

[Bis81h] Jean-Michel Bismut. Intégrales stochastiques. (French) [Stochastic integrals]. Lecture Notes in Mathematics, 866:123–156, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088596/.

Bismut:1981:IFI

[Bis81i] Jean-Michel Bismut. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 866:3–35, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0088593/.

Bismut:1981:MMC

[Bis81j] Jean-Michel Bismut. Martingales, the Malliavin calculus and Hörmander’s theorem. Lecture Notes in Mathematics, 851:85– 109, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088724/.

Bismut:1981:MAF

[Bis81k] Jean-Michel Bismut. Mécanique Aléatoire. (French) [Random mechanics], volume 866 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 563 pp. LCCN QA3 .L28 no. 866. URL http:/

/link.springer.com/book/10.1007/BFb0088591;   http://
www.springerlink.com/content/978-3-540-38734-3.

Bismut:1981:NFN

[Bis81l] Jean-Michel Bismut. Notations. (French) [Notation]. Lecture Notes in Mathematics, 866:1–2, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088592.

Bismut:1981:PVD

[Bis81m] Jean-Michel Bismut. Problèmes variationnels et diffusions hamiltoniennes. (French) []. Lecture Notes in Mathematics, 866:242– 362, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088599/.

Bismut:1981:PID

[Bis81n] Jean-Michel Bismut. Proprétés d’intersection du flot φ.(ω, .). (French) []. Lecture Notes in Mathematics, 866:103–122, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088595/.

Bismut:1981:RSH

[Bis81o] Jean-Michel Bismut. Reconstruction de la structure hamiltonienne dans les problèmes d’optimisation stochastique classique. (French) []. Lecture Notes in Mathematics, 866:363–379, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088600/.

Bismut:1981:RCF

[Bis81p] Jean-Michel Bismut. Relévement de connexions dans un fibre cotangent et relévement de diffusions. (French) []. Lecture Notes in Mathematics, 866:439–473, 1981. CODEN LNMAA2. ISBN 3-540-10840-8 (print), 3-540-38734-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088603/.

Bismut:1982:MAF

[Bis82] J.-M. Bismut. Mecanique aléatoire. (French) []. Lecture Notes in Mathematics, 929:1–100, 1982. CODEN LNMAA2. ISBN 3-540-11547-1 (print), 3-540-39220-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095618/.

Bojanowska:1980:SSC

[BJ80] Agnieszka Bojanowska and Stefan Jackowski. A spectral sequence convergent to equivariant K-theory. Lecture Notes in Mathematics, 788:245–256, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099248/.

Beth:1981:MGD

[BJ81] Thomas Beth and Dieter Jungnickel. Mathieu groups, with designs, and Golay codes. Lecture Notes in Mathematics, 893:157– 179, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091016/.

Beck:1983:PBS

[BJ83a] Anatole Beck and Konrad Jacobs, editors. Probability in Banach Spaces IV: Proceedings of the Seminar Held in Oberwolfach, Germany, July 1982, volume 990 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 0-387-12295-8, 3-540-12295-8 (print), 3-540-39870-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 990; QA3 .L28 no. 990 QA273.43. URL http:/

/link.springer.com/book/10.1007/BFb0064259;   http://
www.springerlink.com/content/978-3-540-39870-7.

Bichteler:1983:CMP

[BJ83b] Klaus Bichteler and Jean Jacod. Calcul de Malliavin pour les diffusions avec sauts: Existence d’une densité dans le cas unidimensionnel. (French) []. Lecture Notes in Mathematics, 986: 132–157, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068309/.

Borowiecki:1983:NCP

[BJ83c] Mieczysław Borowiecki and Tadeusz Jóźwiak. A note on characteristic and permanental polynomials of multigraphs. Lecture Notes in Mathematics, 1018:75–78, 1983. CODEN LN-MAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071615/.

Berner:1984:PPG

[BJ84] Andrew J. Berner and István Juhász. Point-picking games and HFD’s. Lecture Notes in Mathematics, 1103:53–66, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099380/.

Bjork:1981:BCM

[Bjö81] Jan-Erik Björk. The Bernstein class of modules on algebraic manifolds. Lecture Notes in Mathematics, 867:148–156, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090385/.

Bjon:1982:NLV

[Bjo82] Sten Bjon. On nuclear limit vector spaces. Lecture Notes in Mathematics, 915:27–39, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092869/.

Bjorck:1984:SMS

[Bjö84] Å. Björck. Some methods for separating stiff components in initial value problems. Lecture Notes in Mathematics, 1066:30–43, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099516/.

Barthel:1981:LP

[BK81a] Dominique Barthel and Reinhard Kümritz. Laplacian with a potential. Lecture Notes in Mathematics, 838:19–29, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088838/.

Brehm:1981:SAP

[BK81b] U. Brehm and W. Kühnel. Smooth approximation of polyhedral surfaces with respect to curvature measures. Lecture Notes in Mathematics, 838:64–68, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088842/.

Buser:1981:BCG

[BK81c] Peter Buser and Hermann Karcher. The Bieberbach case in Gromov’s almost flat manifold theorem. Lecture Notes in Mathematics, 838:82–93, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088844/.

Brin:1983:LE

[BK83] M. Brin and A. Katok. On local entropy. Lecture Notes in Mathematics, 1007:30–38, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061408/.

Bauer:1981:VKP

[BKK81] Heiko Bauer, Klaus Keimel, and Roland Köhler. Verfeinerungsund Kürzungssätze für Produkte geordneter topologischer Räume und für Funktionen (-halb-) verbände. (German) []. Lecture Notes in Mathematics, 871:20–44, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089902/.

Butkovic:1982:FAP

[BKK82] Davor Butković, Hrvoje Kraljević, and Svetozar Kurepa, editors. Functional Analysis: Proceedings of a Conference Held at Dubrovnik, Yugoslavia, November 2–14, 1981, volume 948 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 948. URL http:/

/link.springer.com/book/10.1007/BFb0069839;   http://
www.springerlink.com/content/978-3-540-39356-6.

Bloom:1983:CNL

[BKQ83] Gary S. Bloom, John W. Kennedy, and Louis V. Quintas. On crossing numbers and linguistic structures. Lecture Notes in Mathematics, 1018:14–22, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071606/.

Bochnak:1982:DCG

[BKS82] Jacek Bochnak, Wojciech Kucharz, and Masahiro Shiota. The divisor class groups of some rings of global real analytic, Nash or rational regular functions. Lecture Notes in Mathematics, 959: 218–248, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062257/.

Borowiecki:1983:GTP

[BKS83] M. Borowiecki, John W. Kennedy, and Maciej M. Sysło, editors. Graph Theory: Proceedings of a Conference held in Łagów, Poland, February 10–13, 1981, volume 1018 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1018.

URL http://link.springer.com/book/10.1007/BFb0071604;
 http://www.springerlink.com/content/978-3-540-38679-
7.

Berestycki:1980:SAM

[BL80a] H. Berestycki and P.-L. Lions. Some applications of the method of super and subsolutions. Lecture Notes in Mathematics, 782: 16–41, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090426/.

Bonetti:1980:CTS

[BL80b] Flavio Bonetti and Guglielmo Lunardon. Central translation Sspace. Lecture Notes in Mathematics, 792:24–29, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088657/.

Bultheel:1981:EAI

[BL81] A. Bultheel and K. U. Leuven. Error analysis of incoming and outgoing schemes for the trigonometric moment problem. Lecture Notes in Mathematics, 888:100–109, 1981. CODEN LN-MAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095579/.

Britten:1982:CPM

[BL82] D. J. Britten and F. W. Lemire. A classification of pointed A_{n}- modules. Lecture Notes in Mathematics, 933:63–70, 1982. CO-DEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093353/.

Blattner:1983:GQ

[Bla83] R. J. Blattner. On geometric quantization. Lecture Notes in Mathematics, 1037:209–241, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073174/.

Blei:1980:FCP

[Ble80] Ron C. Blei. Fractional cartesian products in harmonic analysis. Lecture Notes in Mathematics, 781:8–15, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097642/.

Blel:1982:FPI

[Ble82] Mongi Blel. Fonctions plurisousharmoniques et idéal définissant un ensemble analytique. (French) []. Lecture Notes in Mathematics, 919:26–55, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097042/.

Blei:1983:FBV

[Ble83] Ron C. Blei. Functions of bounded variation and fractional dimension. Lecture Notes in Mathematics, 992:309–322, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069165/.

Bloore:1980:CSI

[Blo80] F. J. Bloore. Configuration spaces of identical particles. Lecture Notes in Mathematics, 836:1–8, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089723/.

Bloch:1981:TAC

[Blo81a] S. Bloch. Torsion algebraic cycles, K_{2}, and Brauer groups of function fields. Lecture Notes in Mathematics, 844:75–102, 1981. CODEN LNMAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090478/.

Bloch:1981:DEL

[Blo81b] Spencer Bloch. The dilogarithm and extensions of Lie algebras. Lecture Notes in Mathematics, 854:1–23, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089515/.

Block:1982:CPS

[Blo82a] Richard E. Block. The classification problem for simple Lie algebras of characteristic p. Lecture Notes in Mathematics, 933: 38–56, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093351/.

Bloom:1982:IDM

[Blo82b] Walter R. Bloom. Infinitely divisible measures on hypergroups. Lecture Notes in Mathematics, 928:1–15, 1982. CODEN LN-MAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093216/.

Bloom:1983:MFW

[Blo83] Steven Bloom. The maximal function on weighted BMO. Lecture Notes in Mathematics, 992:227–239, 1983. CODEN LN-MAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069162/.

Bloom:1984:CIP

[Blo84a] Thomas Bloom. On the convergence of interpolating polynomials for entire functions. Lecture Notes in Mathematics, 1094: 15–19, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099150/.

Bloom:1984:TBM

[Blo84b] Walter R. Bloom. Translation bounded measures and the Orlicz– Paley–Sidon theorem. Lecture Notes in Mathematics, 1064:1– 9, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073629/.

Bove:1983:AFH

[BLP83a] Antonio Bove, Jeff E. Lewis, and Cesare Parenti. Applications to Fuchsian hyperbolic P.D.E. Lecture Notes in Mathematics, 984: 96–132, 1983. CODEN LNMAA2. ISBN 3-540-12285-0 (print), 3-540-39609-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065387/.

Bove:1983:GFS

[BLP83b] Antonio Bove, Jeff E. Lewis, and Cesare Parenti. General Fuchsian systems. Lecture Notes in Mathematics, 984:24–95, 1983. CODEN LNMAA2. ISBN 3-540-12285-0 (print), 3-540-39609-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065386/.

Bove:1983:I

[BLP83c] Antonio Bove, Jeff E. Lewis, and Cesare Parenti. Introduction. Lecture Notes in Mathematics, 984:1–9, 1983. CODEN LNMAA2. ISBN 3-540-12285-0 (print), 3-540-39609-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065384/.

Bove:1983:OMN

[BLP83d] Antonio Bove, Jeff E. Lewis, and Cesare Parenti. Operators with multiple non-involutive characteristics. Lecture Notes in Mathematics, 984:133–157, 1983. CODEN LNMAA2. ISBN 3-540-12285-0 (print), 3-540-39609-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065388/.

Bove:1983:PRR

[BLP83e] Antonio Bove, Jeff E. Lewis, and Cesare Parenti. Preliminaries and review of results of N. Hanges. Lecture Notes in Mathematics, 984:10–23, 1983. CODEN LNMAA2. ISBN 3-540-12285-0 (print), 3-540-39609-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065385/.

Bove:1983:PSF

[BLP83f] Antonio Bove, Jeff E. Lewis, and Cesare Parenti. Propagation of Singularities for Fuchsian Operators, volume 984 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12285-0 (print), 3-540-39609-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 161 pp. LCCN QA3 .L28 no. 984. URL http:/

/link.springer.com/book/10.1007/BFb0065383;   http://
www.springerlink.com/content/978-3-540-39609-3.

Bardos:1980:BNE

[BLS80] C. Bardos, J. M. Lasry, and M. Schatzman, editors. Bifurcation and Nonlinear Eigenvalue Problems: Proceedings, Université de Paris XIII, Villetaneuse, France, October 2–4, 1978, volume 782 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 782. URL http:/

/link.springer.com/book/10.1007/BFb0090424;   http://
www.springerlink.com/content/978-3-540-38637-7.

Burbea:1980:HSH

[BM80] J. Burbea and P. Masani. Hilbert spaces of Hilbert space valued functions. Lecture Notes in Mathematics, 828:1–20, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097390/.

Becker:1981:MCP

[BM81a] Howard S. Becker and Yiannis N. Moschovakis. Measurable cardinals in playful models. Lecture Notes in Mathematics, 839: 203–214, 1981. CODEN LNMAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090242/.

Benabdallah:1981:DDT

[BM81b] Khalid Benabdallah and Otto Mutzbauer. On direct decompositions of torsion free abelian groups of rank 4. Lecture Notes in Mathematics, 874:62–69, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090524/.

Brin:1981:ADP

[BM81c] M. Brin and A. Manning. Anosov diffeomorphisms with pinched spectrum. Lecture Notes in Mathematics, 898:48–53, 1981. CO-DEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091906/.

Bakry:1982:ISLa

[BM82a] D. Bakry and P.-A. Meyer. Sur les inégalités de Sobolev logarithmiques. I. (French) []. Lecture Notes in Mathematics, 920: 138–145, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092775/.

Bakry:1982:ISLb

[BM82b] D. Bakry and P.-A. Meyer. Sur les inégalités de Sobolev logarithmiques. II. (French) []. Lecture Notes in Mathematics, 920: 146–150, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092776/.

Brenner:1983:TAC

[BM83] Gary Brenner and Donald Monk. Tree algebras and chains. Lecture Notes in Mathematics, 1004:54–66, 1983. CODEN LN-MAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063429/.

Bloom:1983:RGP

[BMKQ83] Gary S. Bloom, Małgorzata T. Mandziuk, John W. Kennedy, and Louis V. Quintas. Random graphs and the physical world. Lecture Notes in Mathematics, 1018:94–110, 1983. CODEN LN-MAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071618/.

Boffa:1980:QEP

[BMP80] M. Boffa, A. Macintyre, and F. Point. The quantifier elimination problem for rings without nilpotent elements and for semi-simple rings. Lecture Notes in Mathematics, 834:20–30, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090158/.

Berline:1981:MTA

[BMR81] Chantal Berline, Kenneth McAloon, and Jean-Pierre Ressayre, editors. Model Theory and Arithmetic: Comptes Rendus d’une Action Thématique Programmée du C.N.R.S. sur la Théorie des Modèles et l’Arithmétique, Paris, France, 1979/ 80, volume 890 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 890. URL http:/

/link.springer.com/book/10.1007/BFb0095651;   http://
www.springerlink.com/content/978-3-540-38629-2.

Bussemaker:1981:TTG

[BMS81] F. C. Bussemaker, R. A. Mathon, and J. J. Seidel. Tables of twographs. Lecture Notes in Mathematics, 885:70–112, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092256/.

Brauner:1980:NEP

[BN80] C.-M. Brauner and B. Nicolaenko. On nonlinear eigenvalue problems which extend into free boundaries problems. Lecture Notes in Mathematics, 782:61–100, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090428/.

Beckenstein:1981:ENN

[BN81] Edward Beckenstein and Lawrence Narici. Extending nonarchimedean norms on algebras. Lecture Notes in Mathematics, 843:33–40, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089268/.

Brauner:1982:RBP

[BN82] C.-M. Brauner and B. Nicolaenko. Regularization and bounded penalization in free boundary problems. Lecture Notes in Mathematics, 942:19–42, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094737/.

Benkart:1982:RRO

[BO82] Georgia M. Benkart and J. Marshall Osborn. Representations of rank one Lie algebras of characteristic p. Lecture Notes in Mathematics, 933:1–37, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093350/.

Bruna:1984:IFT

[BO84] Joaquim Bruna and Joaquín M. Ortega. Idéaux fermés de type fini dans les algèbres A^{m}(D). (French) []. Lecture Notes in Mathematics, 1094:20–28, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099151/.

Bochnak:1982:TRA

[Boc82] J. Bochnak. Topology of real analytic sets — some open problems. Lecture Notes in Mathematics, 959:212–217, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062256/.

Bohme:1981:PPM

[Böh81] R. Böhme. A plateau problem with many solutions for boundary curves in a given knot class. Lecture Notes in Mathematics, 838:36–41, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088840/.

Bohmer:1982:AEE

[Böh82] Klaus Böhmer. Asymptotic error expansions and discrete Newton methods for elliptic boundary value problems. Lecture Notes in Mathematics, 968:292–300, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064896/.

Boidol:1980:RCG

[Boi80] J. Boidol. On a regularity condition for group algebras of non abelian locally compact groups. Lecture Notes in Mathematics, 781:16–21, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097643/.

Boivin:1983:CAM

[Boi83] André Boivin. On Carleman approximation by meromorphic functions. Lecture Notes in Mathematics, 1039:9–15, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073355/.

Bokstedt:1984:RHT

[Bök84] Marcel Bökstedt. The rational homotopy type of ωWh^{Diff}(∗). Lecture Notes in Mathematics, 1051:25–37, 1984. CODEN LN-MAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075562/.

Bolley:1980:MSB

[Bol80] C. Bolley. Multiple solutions of a bifurcation problem. Lecture Notes in Mathematics, 782:42–60, 1980. CODEN LN-MAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090427/.

Bongartz:1980:ZAZ

[Bon80] Klaus Bongartz. Zykellose Algebren sind nicht zügellos. (German) []. Lecture Notes in Mathematics, 832:97–102, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088460/.

Bongartz:1981:TA

[Bon81] Klaus Bongartz. Tilted algebras. Lecture Notes in Mathematics, 903:26–38, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092982/.

Bongaarts:1982:PFQ

[Bon82a] P. J. M. Bongaarts. Particles, fields and quantum theory. Lecture Notes in Mathematics, 926:143–159, 1982. CODEN LN-MAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092660/.

Bongiorno:1982:EV

[Bon82b] Benedetto Bongiorno. Essential variations. Lecture Notes in Mathematics, 945:187–193, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096674/.

Bonnet:1984:HTS

[Bon84] Robert Bonnet. On homomorphism types of superatomic interval Boolean algebras. Lecture Notes in Mathematics, 1103:67– 81, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099381/.

Borceux:1982:HOS

[Bor82a] Francis Borceux. Higher order sheaves and Banach modules. Lecture Notes in Mathematics, 915:40–49, 1982. CODEN LN-MAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092870/.

Borceux:1982:AL

[Bor82b] Francis Borceux. On algebraic localizations. Lecture Notes in Mathematics, 962:13–22, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066881/.

Borho:1982:IDR

[Bor82c] Walter Borho. Invariant dimension and restricted extension of Noetherian rings. Lecture Notes in Mathematics, 924:51–71, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092925/.

Borchers:1983:SCR

[Bor83a] H.-J. Borchers. Symmetries and covariant representations. Lecture Notes in Mathematics, 1031:1–15, 1983. CODEN LN-MAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072107/.

Borisov:1983:RCI

[Bor83b] I. S. Borisov. Rate of convergence in invariance principle in linear spaces. Application to empirical measures. Lecture Notes in Mathematics, 1021:45–58, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072902/.

Borowiecki:1983:HM

[Bor83c] Mieczysław Borowiecki. On Hamiltonian matroids. Lecture Notes in Mathematics, 1018:23–27, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071607/.

Boratynski:1984:PFG

[Bor84a] Maksymilian Boratyński. Poincaré forms, Gorenstein algebras and set theoretic complete intersections. Lecture Notes in Mathematics, 1092:270–290, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099369/.

Borger:1984:CPT

[BOR^{+}84b] Egon Börger, Walter Oberschelp, Michael M. Richter, Brigitta Schinzel, and Wolfgang Thomas, editors. Computation and Proof Theory: Proceedings of the Logic Colloquium held in Aachen, July 18–23, 1983 Part II, volume 1104 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1104. URL http:/

/link.springer.com/book/10.1007/BFb0099476;   http://
www.springerlink.com/content/978-3-540-39119-7.

Borisovich:1984:SNL

[Bor84c] Yu. G. Borisovich. On solvability of non-linear equations with Fredholm operators. Lecture Notes in Mathematics, 1108:194– 208, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099558/.

Bosgiraud:1980:PTM

[Bos80] Jacques Bosgiraud. Problème de type mixte sur la frontière de Martin. (French) []. Lecture Notes in Mathematics, 814:20– 69, 1980. CODEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094146/.

Bose:1984:PPA

[Bos84] N. K. Bose. Properties of Padé approximants to Stieltjes series and systems theory. Lecture Notes in Mathematics, 1105: 182–188, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072411/.

Botvich:1983:SPG

[Bot83] D. D. Botvich. Spectral properties of GNS–Hamiltonian in quasifree state. Lecture Notes in Mathematics, 1021:67–72, 1983. CO-DEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072904/.

Bouscaren:1980:ECM

[Bou80] Elisabeth Bouscaren. Existentially closed modules: Types and prime models. Lecture Notes in Mathematics, 834:31–43, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090159/.

Bouaziz:1981:RAL

[Bou81a] A. Bouaziz. Sur les représentations des algèbres de Lie semisimples construites par T. Enright. (French) []. Lecture Notes in Mathematics, 880:57–68, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090403/.

Bouleau:1981:PID

[Bou81b] N. Bouleau. Propriétés d’invariance du domaine du generateur infinitesimal etendu d’un processus de Markov. (French) []. Lecture Notes in Mathematics, 850:167–188, 1981. CODEN LN-MAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088366/.

Bourgain:1981:CSb

[Bou81c] Jean Bourgain. A class of L^{∞}-spaces. Lecture Notes in Mathematics, 889:84–101, 1981. CODEN LNMAA2. ISBN 3-540-11156-5 (print), 3-540-38611-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096363/.

Bourgain:1981:CSa

[Bou81d] Jean Bourgain. A class of L^{P}-spaces. Lecture Notes in Mathematics, 889:48–83, 1981. CODEN LNMAA2. ISBN 3-540-11156-5 (print), 3-540-38611-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096362/.

Bourgain:1981:GIC

[Bou81e] Jean Bourgain. Geometrical implications of certain finite dimensional structures. Lecture Notes in Mathematics, 889:14–24, 1981. CODEN LNMAA2. ISBN 3-540-11156-5 (print), 3-540-38611-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096360/.

Bourgain:1981:NCSa

[Bou81f] Jean Bourgain. New classes of L^{∞}-spaces. Lecture Notes in Mathematics, 889:25–47, 1981. CODEN LNMAA2. ISBN 3-540-11156-5 (print), 3-540-38611-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096361/.

Bourgain:1981:NCSb

[Bou81g] Jean Bourgain. New Classes of L^{P}-spaces, volume 889 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11156-5 (print), 3-540-38611-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 143 pp. LCCN QA3 .L28 no. 889. URL http:/

/link.springer.com/book/10.1007/BFb0096358;   http://
www.springerlink.com/content/978-3-540-38611-7.

Bourgain:1981:P

[Bou81h] Jean Bourgain. Preliminaries. Lecture Notes in Mathematics, 889:2–13, 1981. CODEN LNMAA2. ISBN 3-540-11156-5 (print), 3-540-38611-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096359/.

Bourguignon:1981:RCE

[Bou81i] Jean-Pierre Bourguignon. Ricci curvature and Einstein metrics. Lecture Notes in Mathematics, 838:42–63, 1981. CODEN LN-MAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088841/.

Bouleau:1982:PPS

[Bou82a] N. Bouleau. Perturbation positive d’un semi-groupe droit dans le cas critique, application à la construction de processus de Harris. (French) []. Lecture Notes in Mathematics, 906:53–87, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093261/.

Bourguignon:1982:HCG

[Bou82b] Jean-Pierre Bourguignon. Harmonic curvature for gravitational and Yang–Mills fields. Lecture Notes in Mathematics, 949:35– 47, 1982. CODEN LNMAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069754/.

Bourguignon:1982:SDY

[Bou82c] Jean-Pierre Bourguignon. Self-duality of Yang–Mills fields and of gravitational instantons. Lecture Notes in Mathematics, 925: 95–102, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093502/.

Bourn:1982:CAI

[Bou82d] Dominique Bourn. A canonical action on indexed limits an application to coherent homotopy. Lecture Notes in Mathematics, 962:23–32, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066882/.

Boudourides:1983:HLH

[Bou83a] M. A. Boudourides. Hyperbolic Lipschitz homeomorphisms and flows. Lecture Notes in Mathematics, 1017:101–106, 1983. CO-DEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103240/.

Bourgain:1983:NBS

[Bou83b] J. Bourgain. New Banach space properties of the disc algebra and H^{∞}. Lecture Notes in Mathematics, 995:8–13, 1983. CODEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061885/.

Bourgin:1983:ASG

[Bou83c] Richard D. Bourgin. Asplund spaces and the GSP. Lecture Notes in Mathematics, 993:116–172, 1983. CODEN LNMAA2. ISBN 3-540-12296-6 (print), 3-540-39873-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069326/.

Bourgin:1983:B

[Bou83d] Richard D. Bourgin. Background. Lecture Notes in Mathematics, 993:1–14, 1983. CODEN LNMAA2. ISBN 3-540-12296-6 (print), 3-540-39873-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069322/.

Bourgin:1983:BR

[Bou83e] Richard D. Bourgin. Basic results. Lecture Notes in Mathematics, 993:15–37, 1983. CODEN LNMAA2. ISBN 3-540-12296-6 (print), 3-540-39873-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069323/.

Bourgin:1983:DS

[Bou83f] Richard D. Bourgin. Dual spaces. Lecture Notes in Mathematics, 993:71–115, 1983. CODEN LNMAA2. ISBN 3-540-12296-6 (print), 3-540-39873-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069325/.

Bourgin:1983:EP

[Bou83g] Richard D. Bourgin. Extreme points. Lecture Notes in Mathematics, 993:38–70, 1983. CODEN LNMAA2. ISBN 3-540-12296-6 (print), 3-540-39873-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069324/.

Bourgin:1983:GAC

[Bou83h] Richard D. Bourgin. Geometric Aspects of Convex Sets with the Radon–Nikodým Property, volume 993 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12296-6 (print), 3-540-39873-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 448 (est.) pp. LCCN QA3 .L28 no. 993.

URL http://link.springer.com/book/10.1007/BFb0069321;
 http://www.springerlink.com/content/978-3-540-39873-
8.

Bourgin:1983:IR

[Bou83i] Richard D. Bourgin. Integral representations. Lecture Notes in Mathematics, 993:173–239, 1983. CODEN LNMAA2. ISBN 3-540-12296-6 (print), 3-540-39873-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069327/.

Bourgin:1983:ST

[Bou83j] Richard D. Bourgin. Selected topics. Lecture Notes in Mathematics, 993:240–448, 1983. CODEN LNMAA2. ISBN 3-540-12296-6 (print), 3-540-39873-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069328/.

Bourn:1983:CFI

[Bou83k] Stephen Bourn. A canonical form for incidence matrices of finite projective planes and their associated latin squares and planar ternary rings. Lecture Notes in Mathematics, 1036:111–120, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071512/.

Bougerol:1984:CID

[Bou84a] Philippe Bougerol. Comportement à l’infini du noyau potentiel sur un espace symétrique. (French) []. Lecture Notes in Mathematics, 1096:90–115, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100108/.

Bougerol:1984:EDT

[Bou84b] Philippe Bougerol. Extension du théorème de continuité de Paul Lévy aux groupes moyennables. (French) []. Lecture Notes in Mathematics, 1064:10–22, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073630/.

Boukricha:1984:EBF

[Bou84c] Abderrahman Boukricha. Espaces biharmoniques. (French) []. Lecture Notes in Mathematics, 1096:116–148, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100109/.

Bouleau:1984:OCD

[Bou84d] N. Bouleau. Opérateur carré du champ, subordinateurs et processus de Dirichlet. (French) []. Lecture Notes in Mathematics, 1061: 82–91, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099018/.

Bouleau:1984:QRP

[Bou84e] N. Bouleau. Quelques résultats probabilistes sur la subordination au sens de Bochner. (French) []. Lecture Notes in Mathematics, 1061:54–81, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099017/.

Bouleau:1984:DEP

[Bou84f] Nicolas Bouleau. Décomposition de l’énergie par niveau de potentiel. (French) []. Lecture Notes in Mathematics, 1096:149– 172, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100110/.

Bowser:1982:NTS

[Bow82] S. E. Bowser. Normalizer towers in semisimple Lie algebras. Lecture Notes in Mathematics, 933:57–62, 1982. CODEN LN-MAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093352/.

Billington:1982:CMI

[BOWS82] Elizabeth J. Billington, Sheila Oates-Williams, and Anne Penfold Street, editors. Combinatorial Mathematics IX: Proceedings of the Ninth Australian Conference on Combinatorial Mathematics Held at the University of Queensland, Brisbane, Australia, August 24–28, 1981, volume 952 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 952. URL http:

//link.springer.com/book/10.1007/BFb0061970;   http:/
/www.springerlink.com/content/978-3-540-39375-7.

Balasubramanian:1981:SCI

[BP81a] K. Balasubramanian and K. R. Parthasarathy. In search of a complete invariant for graphs. Lecture Notes in Mathematics, 885: 42–59, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092254/.

Beutelspacher:1981:ESR

[BP81b] Albrecht Beutelspacher and Ursula Porta. Extending strongly resolvable designs. Lecture Notes in Mathematics, 893:180–182, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091017/.

Barlow:1983:SEU

[BP83] M. T. Barlow and E. Perkins. Strong existence, uniqueness and non-uniqueness in an equation involving local time. Lecture Notes in Mathematics, 986:32–61, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068297/.

Badescu:1984:AGB

[BP84a] Lucian Bădescu and Dorin Popescu, editors. Algebraic Geometry Bucharest 1982: Proceedings of the International Conference held in Bucharest, Romania, August 2–7, 1982, volume 1056 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1056.

URL http://link.springer.com/book/10.1007/BFb0071766;
 http://www.springerlink.com/content/978-3-540-38811-
1.

Barlow:1984:LWE

[BP84b] M. T. Barlow and E. Perkins. Levels at which every Brownian excursion is exceptional. Lecture Notes in Mathematics, 1059: 1–28, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100028/.

Bourgain:1984:AFS

[BPK^{+}84] J. Bourgain, A. Pełczyński, I. A. Komarchev, B. M. Makarov, P. Wojtaszczyk, P. G. Casazza, E. M. Semënov, P. Wojtaszczyk, Peter W. Jones, E. M. Semënov, V. P. Zahariuta, O. S. Semiguk, N. I. Skiba, V. P. Zahariuta, F. Haslinger, et al. Analysis in functional spaces. Lecture Notes in Mathematics, 1043:2–47, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072184/.

Baker:1983:MSM

[BPV83] J. W. Baker, J. S. Pym, and H. L. Vasudeva. Multipliers for some measure algebras on compact semilattices. Lecture Notes in Mathematics, 998:8–30, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062025/.

Bauer:1980:DOP

[BR80a] Karl Wilhelm Bauer and Stephan Ruscheweyh. Differential Operators for Partial Differential Equations and Function Theoretic Applications, volume 791 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09975-1 (print), 3-540-39211-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 258 pp. LCCN QA3 .L28 no. 791.

URL http://link.springer.com/book/10.1007/BFb0103468;
 http://www.springerlink.com/content/978-3-540-39211-
8.

Bauer:1980:KWB

[BR80b] Karl Wilhelm Bauer and Stephan Ruscheweyh. Karl Wilhelm Bauer differential operators for partial differential equations. Lecture Notes in Mathematics, 791:1–187, 1980. CODEN LN-MAA2. ISBN 3-540-09975-1 (print), 3-540-39211-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103469/.

Bauer:1980:SRF

[BR80c] Karl Wilhelm Bauer and Stephan Ruscheweyh. Stephan Ruscheweyh on the function theory of the Peschl–Bauer equation. Lecture Notes in Mathematics, 791:191–253, 1980. CODEN LNMAA2. ISBN 3-540-09975-1 (print), 3-540-39211-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103470/.

Bharucha-Reid:1981:WCS

[BR81a] A. T. Bharucha-Reid. Weak compactness of solution measures associated with random equations. Lecture Notes in Mathematics, 860:25–34, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090606/.

Bongartz:1981:RFT

[BR81b] Klaus Bongartz and Claus Michael Ringel. Representation-finite tree algebras. Lecture Notes in Mathematics, 903:39–54, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092983/.

Bowman:1981:SBS

[BR81c] H. Bowman and K. M. Rangaswamy. On special balanced supgroups of torsion-free separable abelian groups. Lecture Notes in Mathematics, 874:32–40, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090521/.

Bushnell:1981:ZFO

[BR81d] Colin J. Bushnell and Irving Reiner. Zeta-functions of orders. Lecture Notes in Mathematics, 882:159–173, 1981. CODEN LN-MAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092491/.

Braaksma:1980:LIS

[Bra80] B. L. J. Braaksma. Laplace integrals in singular differential and difference equations. Lecture Notes in Mathematics, 827: 25–53, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091373/.

Braess:1982:CRM

[Bra82a] Dietrich Braess. The convergence rate of a multigrid method with Gauss–Seidel relaxation for the Poisson equation. Lecture Notes in Mathematics, 960:368–386, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069934/.

Brandas:1982:WTS

[Brä82b] Erkki Brändas. Weyl’s theory for second order differential equations and its application to some problems in quantum chemistry. Lecture Notes in Mathematics, 968:22–39, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064879/.

Brandt:1982:GMD

[Bra82c] Achi Brandt. Guide to multigrid development. Lecture Notes in Mathematics, 960:220–312, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069930/.

Braess:1984:RAE

[Bra84a] Dietrich Braess. On rational approximation of the exponential and the square root function. Lecture Notes in Mathematics, 1105: 89–99, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072401/.

Bratteli:1984:DSC

[Bra84b] Ola Bratteli. On dynamical semigroups and compact group actions. Lecture Notes in Mathematics, 1055:46–61, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071709/.

Bredimas:1981:TSE

[Bré81a] Antoine Pind. Brédimas. Totally singular extensions of the Cauchy problem for a special class of partial derivative operators with distributional coefficients. Lecture Notes in Mathematics, 846:64–73, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089825/.

Breen:1981:EDG

[Bre81b] Lawrence Breen. Extensions du groupe additif sur le site parfait. (French) []. Lecture Notes in Mathematics, 868:238–262, 1981.

CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090650/.

Brezinski:1981:LHC

[Bre81c] Claude Brezinski. The long history of continued fractions and Padé approximants. Lecture Notes in Mathematics, 888:1–27, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095574/.

Breach:1982:SGA

[Bre82a] D. R. Breach. Star gazing in affine planes. Lecture Notes in Mathematics, 952:1–33, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061971/.

Brehm:1982:DSI

[Bre82b] H. Brehm. Description of spherically invariant random processes by means of g-functions. Lecture Notes in Mathematics, 969: 39–73, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062986/.

Brebbia:1983:NDC

[Bre83a] C. A. Brebbia. A new development in computational engineering. Lecture Notes in Mathematics, 1005:16–36, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112523/.

Breen:1983:SFS

[Bre83b] Lawrence Breen. Σ-structures. (French) [Σ-structures]. Lecture Notes in Mathematics, 980:55–66, 1983. CODEN LN-MAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065688/.

Breen:1983:BSF

[Bre83c] Lawrence Breen. Biextensions symétriques. (French) []. Lecture Notes in Mathematics, 980:1–11, 1983. CODEN LNMAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065684/.

Breen:1983:CPA

[Bre83d] Lawrence Breen. Champs de Picard associés à une biextension symétrique. (French) []. Lecture Notes in Mathematics, 980:72– 84, 1983. CODEN LNMAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065690/.

Breen:1983:FTA

[Bre83e] Lawrence Breen. Fonctions thêta algébriques. (French) []. Lecture Notes in Mathematics, 980:45–54, 1983. CODEN LN-MAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065687/.

Breen:1983:FTC

[Bre83f] Lawrence Breen. Fonctions thêta canoniques. (French) []. Lecture Notes in Mathematics, 980:67–71, 1983. CODEN LN-MAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065689/.

Breen:1983:FTT

[Bre83g] Lawrence Breen. Fonctions theta et théorème du cube. (French) [Theta functions and the cube theorem], volume 980 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 115 pp. LCCN QA3 .L28 no. 980. URL http:/

/link.springer.com/book/10.1007/BFb0065683;   http://
www.springerlink.com/content/978-3-540-39478-5.

Breen:1983:IHF

[Bre83h] Lawrence Breen. Interprétation homotopique. (French) []. Lecture Notes in Mathematics, 980:85–101, 1983. CODEN LN-MAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065691/.

Breen:1983:SDCa

[Bre83i] Lawrence Breen. Structure du cube: définitions. (French) []. Lecture Notes in Mathematics, 980:12–28, 1983. CODEN LN-MAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065685/.

Breen:1983:SDCb

[Bre83j] Lawrence Breen. Structure du cube: propriétés d’additivité et de descente. (French) []. Lecture Notes in Mathematics, 980:29– 44, 1983. CODEN LNMAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065686/.

Breen:1983:TDC

[Bre83k] Lawrence Breen. Théorème du coefficient universel. (French) []. Lecture Notes in Mathematics, 980:102–109, 1983. CODEN LNMAA2. ISBN 3-540-12002-5 (print), 3-540-39478-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065692/.

Brezinski:1984:SDI

[Bre84] Claude Brezinski. Some determinantal identities in a vector space, with applications. Lecture Notes in Mathematics, 1071:1–11, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099606/.

Bridges:1981:TCF

[Bri81a] Douglas S. Bridges. Towards a constructive foundation for quantum mechanics. Lecture Notes in Mathematics, 873:260–273, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090739/.

Bridges:1981:SPL

[Bri81b] W. G. Bridges. Some path-length properties of graphs and digraphs. Lecture Notes in Mathematics, 885:185–194, 1981. CO-DEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092262/.

Brigola:1983:SCB

[Bri83] R. Brigola. On summability in conjugate Banach spaces. Lecture Notes in Mathematics, 991:28–34, 1983. CODEN LN-MAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061555/.

Broer:1980:NLW

[Bro80] L. J. F. Broer. Non-linear wave equations as Hamiltonian systems. Lecture Notes in Mathematics, 810:143–158, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089979/.

Broer:1981:FNF

[Bro81a] Henk Broer. Formal normal form theorems for vector fields and some consequences for bifurcations in the volume preserving case. Lecture Notes in Mathematics, 898:54–74, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091907/.

Broer:1981:QPF

[Bro81b] Henk Broer. Quasi periodic flow near a codimension one singularity of a divergence free vector field in dimension three. Lecture Notes in Mathematics, 898:75–89, 1981. CODEN LN-MAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091908/.

Brosowski:1981:AKT

[Bro81c] Bruno Brosowski. An application of Korovkin’s theorem to certain partial differential equations. Lecture Notes in Mathematics, 843:150–162, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089272/.

Brossard:1981:RMD

[Bro81d] Jean Brossard. Regularité des martingales à deux indices et inégalités de normes. (French) []. Lecture Notes in Mathematics, 863:91–121, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091095/.

Brouwer:1981:SUP

[Bro81e] A. E. Brouwer. Some unitals on 28 points and their embeddings in projective planes of order 9. Lecture Notes in Mathematics, 893: 183–188, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091018/.

Brown:1981:PTF

[Bro81f] Robert F. Brown. On the product theorem for the fixed point index. Lecture Notes in Mathematics, 886:44–48, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092175/.

Broad:1982:ECP

[Bro82a] John T. Broad. Extraction of continuum properties from L^{2} basis set matrix representations of the Schrödinger equation: the Sturm sequence polynomials and Gauss quadrature. Lecture Notes in Mathematics, 968:53–70, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064881/.

Brocker:1982:RSD

[Brö82b] Ludwig Bröcker. Real spectra and distributions of signatures. Lecture Notes in Mathematics, 959:249–272, 1982. CODEN LN-MAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062258/.

Browkin:1982:ESO

[Bro82c] Jerzy Browkin. Elements of small order in K_{2}F. Lecture Notes in Mathematics, 966:1–6, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062164/.

Brown:1982:ADI

[Bro82d] Richard C. Brown. An approach to the Dirichlet index for operators satisfying minimal conditions. Lecture Notes in Mathematics, 964:110–117, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064992/.

Browne:1982:MPL

[Bro82e] Patrick J. Browne. Multiparameter problems: The last decade. Lecture Notes in Mathematics, 964:95–109, 1982. CODEN LN-MAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064991/.

Brown:1983:NFS

[Bro83] Richard C. Brown. A von Neumann factorization of some selfadjoint extensions of positive symmetric differential operators and its application to inequalities. Lecture Notes in Mathematics, 1032:68–92, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076793/.

Bronner:1984:GDA

[Bro84] F. Bronner. Sur les grandes deviations abstraites applications aux temps de séjours moyens d’un processus. (French) []. Lecture Notes in Mathematics, 1059:82–90, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100034/.

Brune:1980:FRT

[Bru80] Hermann Brune. On finite representation type and a theorem of Kulikov. Lecture Notes in Mathematics, 832:170–176, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088462/.

Bruggeman:1981:LGG

[Bru81a] Roelof W. Bruggeman. L^{2}(Γ/G, X). Lecture Notes in Mathematics, 865:93–102, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089866/.

Bruggeman:1981:ACE

[Bru81b] Roelof W. Bruggeman. The analytic continuation of the Eisenstein model. Lecture Notes in Mathematics, 865:103–124, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089867/.

Bruggeman:1981:AM

[Bru81c] Roelof W. Bruggeman. Automorphic models. Lecture Notes in Mathematics, 865:33–40, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089860/.

Bruggeman:1981:CO

[Bru81d] Roelof W. Bruggeman. Convolution operators. Lecture Notes in Mathematics, 865:148–165, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089870/.

Bruggeman:1981:EM

[Bru81e] Roelof W. Bruggeman. Eisenstein models. Lecture Notes in Mathematics, 865:84–92, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089865/.

Bruggeman:1981:FCAb

[Bru81f] Roelof W. Bruggeman. Fourier Coefficients of Automorphic Forms, volume 865 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 201 pp. LCCN QA3 .L28 no. 865. URL http:/

/link.springer.com/book/10.1007/BFb0089856;   http://
www.springerlink.com/content/978-3-540-38726-8.

Bruggeman:1981:FCAa

[Bru81g] Roelof W. Bruggeman. Fourier coefficients of automorphic models. Lecture Notes in Mathematics, 865:61–71, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089863/.

Bruggeman:1981:IO

[Bru81h] Roelof W. Bruggeman. Intertwining operators. Lecture Notes in Mathematics, 865:49–60, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089862/.

Bruggeman:1981:I

[Bru81i] Roelof W. Bruggeman. Introduction. Lecture Notes in Mathematics, 865:1–18, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089857/.

Bruggeman:1981:O

[Bru81j] Roelof W. Bruggeman. The operators θ_{α}. Lecture Notes in Mathematics, 865:72–83, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089864/.

Bruggeman:1981:R

[Bru81k] Roelof W. Bruggeman. Representations. Lecture Notes in Mathematics, 865:24–32, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089859/.

Bruggeman:1981:SPP

[Bru81l] Roelof W. Bruggeman. Scalar product of Poincaré series. Lecture Notes in Mathematics, 865:166–176, 1981. CODEN LN-MAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089871/.

Bruggeman:1981:S

[Bru81m] Roelof W. Bruggeman. The space H_{σ}^{τ}. Lecture Notes in Mathematics, 865:133–147, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089869/.

Bruggeman:1981:SD

[Bru81n] Roelof W. Bruggeman. Spectral decomposition of ^{e}L^{2}(Γ/G, X). Lecture Notes in Mathematics, 865:125–132, 1981. CODEN LN-MAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089868/.

Bruggeman:1981:SWM

[Bru81o] Roelof W. Bruggeman. Standard and Whittaker models. Lecture Notes in Mathematics, 865:41–48, 1981. CODEN LN-MAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089861/.

Bruggeman:1981:SF

[Bru81p] Roelof W. Bruggeman. Sum formula. Lecture Notes in Mathematics, 865:177–195, 1981. CODEN LNMAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089872/.

Bruggeman:1981:UCG

[Bru81q] Roelof W. Bruggeman. The universal covering group G. Lecture Notes in Mathematics, 865:19–23, 1981. CODEN LN-MAA2. ISBN 3-540-10839-4 (print), 3-540-38726-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089858/.

Bruning:1981:IEL

[Brü81r] Jochen Brüning. Invariant eigenfunctions of the Laplacian and their asymptotic distribution. Lecture Notes in Mathematics, 838: 69–81, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088843/.

Brumfiel:1982:RVR

[Bru82a] G. W. Brumfiel. Real valuation rings and ideals. Lecture Notes in Mathematics, 959:55–97, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062250/.

Brummer:1982:NUE

[Brü82b] G. C. L. Brümmer. On the non-unique extension of topological to bitopological properties. Lecture Notes in Mathematics, 915: 50–67, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092871/.

Bruna:1982:MWS

[Bru82c] Joaquim Bruna. Muckenhoupt’s weights in some boundary problems of a complex variable. Lecture Notes in Mathematics, 908: 74–85, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093281/.

Brunner:1984:NSI

[Bru84] Hermann Brunner. The numerical solution of integral equations with weakly singular kernels. Lecture Notes in Mathematics, 1066: 50–71, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099518/.

Brylinski:1980:EVT

[Bry80] J. L. Brylinski. Eventails et variétés toriques. (French) []. Lecture Notes in Mathematics, 777:247–288, 1980. CODEN LN-MAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085887/.

Brylinski:1982:MHS

[Bry82] J.-L. Brylinski. Modules holonômes à singularités regulières et filtration de Hodge. (French) []. Lecture Notes in Mathematics, 961:1–21, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071273/.

Brzezinski:1981:AGT

[Brz81] Juliusz Brzeziński. Algebraic geometry of ternary quadratic forms and orders in quaternion algebras. Lecture Notes in Mathematics, 882:319–325, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092502/.

Babiker:1980:ASL

[BS80a] A. G. A. G. Babiker and W. Strauß. Almost strong liftings and τ-additivity. Lecture Notes in Mathematics, 794:220–227, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088225/.

Babiker:1980:MSW

[BS80b] A. G. A. G. Babiker and W. Strauß. Measure spaces in which every lifting is an almost H-lifting. Lecture Notes in Mathematics, 794:228–232, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088226/.

Becker:1981:DEM

[BS81] K.-H. Becker and R. Seydel. A Duffing equation with more than 20 branch points. Lecture Notes in Mathematics, 878:98–107, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090677/.

Bartijn:1983:MMF

[BS83a] Jaap Bartijn and Jan R. Strooker. Modifications monômiales. (French) [Monomial modifications]. Lecture Notes in Mathematics, 1029:192–217, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098932/.

Berger:1983:ESP

[BS83b] M. Berger and A. Sloan. Explicit solutions of partial differential equations. Lecture Notes in Mathematics, 983:1–14, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065332/.

Bican:1983:BGIa

[BS83c] L. Bican and L. Salce. Butler groups of infinite rank. Lecture Notes in Mathematics, 1006:171–189, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21560-9_6/.

Bican:1983:BGIb

[BS83d] L. Bican and L. Salce. Butler groups of infinite rank. Lecture Notes in Mathematics, 1006:171–189, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103701/.

Blanchard:1983:DPP

[BS83e] Ph. Blanchard and L. Streit, editors. Dynamics and Processes: Proceedings of the Third Encounter in Mathematics and Physics, held in Bielefeld, Germany Nov. 30–Dec. 4, 1981, volume 1031 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1031. URL http:/

/link.springer.com/book/10.1007/BFb0072106;   http://
www.springerlink.com/content/978-3-540-38691-9.

Blei:1983:BSH

[BS83f] Ron C. Blei and Stuart J. Sidney, editors. Banach Spaces, Harmonic Analysis, and Probability Theory: Proceedings of the Special Year in Analysis, Held at the University of Connecticut 1980–1981, volume 995 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 995. URL http:/

/link.springer.com/book/10.1007/BFb0061883;   http://
www.springerlink.com/content/978-3-540-40036-3.

Borovkov:1983:RCI

[BS83g] A. A. Borovkov and A. I. Sahanenko. On the rate of convergence in invariance principle. Lecture Notes in Mathematics, 1021:59– 66, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072903/.

Brown:1983:IPC

[BS83h] D. R. Brown and J. W. Stepp. Inner points in certain holoidal semigroups. Lecture Notes in Mathematics, 998:31–49, 1983. CO-DEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062026/.

Brinzanescu:1984:TTA

[BS84] Vasile Brînzănescu and Manuela Stoia. Topologically trivial algebraic 2-vector bundles on ruled surfaces. II. Lecture Notes in Mathematics, 1056:34–46, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071768/.

Barnett:1984:QSP

[BSW84] C. Barnett, R. F. Streater, and I. F. Wilde. Quantum stochastic processes. Lecture Notes in Mathematics, 1055:32–45, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071708/.

Benenti:1980:GMG

[BT80] S. Benenti and W. M. Tulczyjew. The geometrical meaning and globalization of the Hamilton–Jacobi method. Lecture Notes in Mathematics, 836:9–21, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089724/.

Bertin:1981:CUD

[BT81] E. M. J. Bertin and R. Theodorescu. Characterizations of unimodal distribution functions. Lecture Notes in Mathematics, 861: 10–14, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097307/.

Benedetti:1982:RCT

[BT82a] R. Benedetti and A. Tognoli. Remarks and counterexamples in the theory of real algebraic vector bundles and cycles. Lecture Notes in Mathematics, 959:198–211, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062255/.

Berenstein:1982:GIV

[BT82b] Carlos A. Berenstein and B. A. Taylor. On the geometry of interpolating varieties. Lecture Notes in Mathematics, 919:1–25, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097041/.

Beyl:1982:GER

[BT82c] F. Rudolf Beyl and Jürgen Tappe. Group Extensions, Representations, and the Schur Multiplicator, volume 958 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CO-DEN LNMAA2. ISBN 3-540-11954-X (print), 3-540-39542-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 278 pp. LCCN QA3 .L28 no. 958. URL http:/

/link.springer.com/book/10.1007/BFb0067022;   http://
www.springerlink.com/content/978-3-540-39542-3.

Beyl:1982:GEA

[BT82d] F. Rudolf Beyl and Jürgen Tappe. Group extensions with Abelian kernel. Lecture Notes in Mathematics, 958:5–66, 1982. CODEN LNMAA2. ISBN 3-540-11954-X (print), 3-540-39542-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067024/.

Beyl:1982:Ia

[BT82e] F. Rudolf Beyl and Jürgen Tappe. Introduction. Lecture Notes in Mathematics, 958:1–4, 1982. CODEN LNMAA2. ISBN 3-540-11954-X (print), 3-540-39542-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067023/.

Beyl:1982:Ib

[BT82f] F. Rudolf Beyl and Jürgen Tappe. Isoclinism. Lecture Notes in Mathematics, 958:123–178, 1982. CODEN LNMAA2. ISBN 3-540-11954-X (print), 3-540-39542-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067026/.

Beyl:1982:OGT

[BT82g] F. Rudolf Beyl and Jürgen Tappe. Other group-theoretic applications of the Schur multiplicator. Lecture Notes in Mathematics, 958:179–260, 1982. CODEN LNMAA2. ISBN 3-540-11954-X (print), 3-540-39542-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067027/.

Beyl:1982:STP

[BT82h] F. Rudolf Beyl and Jürgen Tappe. Schur’s theory of projective representations. Lecture Notes in Mathematics, 958:67–122, 1982. CODEN LNMAA2. ISBN 3-540-11954-X (print), 3-540-39542-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067025/.

Boggs:1982:MFN

[BT82i] Paul T. Boggs and Jon W. Tolle. Merit functions for nonlinear programming problems. Lecture Notes in Mathematics, 909: 1–10, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092955/.

Buchweitz:1980:DMC

[Buc80] R. O. Buchweitz. On deformations of monomial curves. Lecture Notes in Mathematics, 777:205–220, 1980. CODEN LN-MAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085884/.

Buchholz:1981:R

[Buc81a] W. Buchholz. The ω_{μ+1}-rule. Lecture Notes in Mathematics, 897: 188–233, 1981. CODEN LNMAA2. ISBN 3-540-11170-0 (print), 3-540-38649-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091898/.

Buchholz:1981:OA

[Buc81b] W. Buchholz. Ordinal analysis of ID_{ν}. Lecture Notes in Mathematics, 897:234–260, 1981. CODEN LNMAA2. ISBN 3-540-11170-0 (print), 3-540-38649-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091899/.

Buckley:1984:EIC

[Buc84] Fred Buckley. Equalities involving certain graphical distributions. Lecture Notes in Mathematics, 1073:179–192, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073116/.

Buekenhout:1981:BDG

[Bue81] Francis Buekenhout. The basic diagram of a geometry. Lecture Notes in Mathematics, 893:1–29, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091010/.

Buekenhout:1982:GMG

[Bue82] F. Buekenhout. Geometries for the Mathieu group M_{12}. Lecture Notes in Mathematics, 969:74–85, 1982. CODEN LN-MAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062987/.

Buium:1984:SDM

[Bui84] Alexandru Buium. On surfaces of degree at most 2n + 1 in P^{n}. Lecture Notes in Mathematics, 1056:47–67, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071769/.

Bujalance:1983:NGK

[Buj83] Emilio Bujalance. NEC groups and Klein surfaces. Lecture Notes in Mathematics, 971:6–14, 1983. CODEN LNMAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067068/.

Bultheel:1984:BSL

[Bul84a] A. Bultheel. On the block structure of the Laurent–Padé table. Lecture Notes in Mathematics, 1105:160–169, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072409/.

Bultheel:1984:ZRF

[Bul84b] Adhemar Bultheel. Zeros of a rational function defined by its Laurent expansion. Lecture Notes in Mathematics, 1071:34–48, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099608/.

Bump:1984:ACF

[Bum84a] Daniel Bump. The analytic continuation and functional equations satisfied by the L-series associated with an automorphic form. Lecture Notes in Mathematics, 1083:133–147, 1984. CO-DEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100155/.

Bump:1984:AFa

[Bum84b] Daniel Bump. Automorphic forms on GL(2). Lecture Notes in Mathematics, 1083:1–16, 1984. CODEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100148/.

Bump:1984:AFb

[Bum84c] Daniel Bump. Automorphic Forms on GL(3, R), volume 1083 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 184 pp. LCCN QA3 .L28 no. 1083. URL http:/

/link.springer.com/book/10.1007/BFb0100147;   http://
www.springerlink.com/content/978-3-540-39055-8.

Bump:1984:DES

[Bum84d] Daniel Bump. The differential equations satisfied by Whittaker functions. Lecture Notes in Mathematics, 1083:17–39, 1984. CO-DEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100149/.

Bump:1984:ES

[Bum84e] Daniel Bump. Eisenstein series. Lecture Notes in Mathematics, 1083:99–132, 1984. CODEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100154/.

Bump:1984:FEA

[Bum84f] Daniel Bump. Fourier expansions of automorphic forms. Lecture Notes in Mathematics, 1083:63–72, 1984. CODEN LN-MAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100151/.

Bump:1984:HOS

[Bum84g] Daniel Bump. Hecke operators and L-series. Lecture Notes in Mathematics, 1083:148–160, 1984. CODEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100156/.

Bump:1984:I

[Bum84h] Daniel Bump. Invariants of G_{∞} \ G. Lecture Notes in Mathematics, 1083:73–80, 1984. CODEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100152/.

Bump:1984:JWF

[Bum84i] Daniel Bump. Jacquet’s Whittaker functions. Lecture Notes in Mathematics, 1083:40–62, 1984. CODEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100150/.

Bump:1984:MTW

[Bum84j] Daniel Bump. The Mellin transforms of the Whittaker functions. Lecture Notes in Mathematics, 1083:161–180, 1984. CODEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100157/.

Bump:1984:RS

[Bum84k] Daniel Bump. Ramanujan sums on GL(3). Lecture Notes in Mathematics, 1083:81–98, 1984. CODEN LNMAA2. ISBN 3-540-13864-1 (print), 3-540-39055-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100153/.

Bunermann:1981:ARQ

[Bün81] Dieter Bünermann. Auslander–Reiten quivers of exact oneparameter partially ordered sets. Lecture Notes in Mathematics, 903:55–61, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092984/.

Burau:1980:SQT

[Bur80] W. Burau. Systems of quadrics through a general variety of Segrè and their reduction to irreducible parts [2]. Lecture Notes in Mathematics, 792:30–39, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088658/.

Burkholder:1981:MTG

[Bur81] D. L. Burkholder. Martingale transforms and the geometry of Banach spaces. Lecture Notes in Mathematics, 860:35–50, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090607/.

Burns:1982:HM

[Bur82] D. Burns. Harmonic maps from CP^{1} to CP^{n}. Lecture Notes in Mathematics, 949:48–56, 1982. CODEN LNMAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069755/.

Burbea:1983:PDH

[Bur83a] Jacob Burbea. Positive definiteness and holomorphy. Lecture Notes in Mathematics, 1039:16–31, 1983. CODEN LN-MAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073356/.

Burns:1983:GEM

[Bur83b] D. Burns. On the geometry of elliptic modular surfaces and representations of finite groups. Lecture Notes in Mathematics, 1008: 1–29, 1983. CODEN LNMAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065696/.

Burris:1983:BC

[Bur83c] Stanley Burris. Boolean constructions. Lecture Notes in Mathematics, 1004:67–90, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063430/.

Bus:1982:GCN

[Bus82] Jacques C. P. Bus. Global convergence of Newton-like methods. Lecture Notes in Mathematics, 909:11–17, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092956/.

Busenberg:1984:CND

[Bus84a] S. N. Busenberg. A class of nonlinear diffusion problems. Lecture Notes in Mathematics, 1076:23–31, 1984. CODEN LN-MAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072762/.

Busenberg:1984:BPB

[Bus84b] Stavros N. Busenberg. Bifurcation phenomena in biomathematics. Lecture Notes in Mathematics, 1057:1–56, 1984. CO-DEN LNMAA2. ISBN 3-540-12931-6 (print), 3-540-38812-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098593/;
 http://link.springer.com/content/pdf/bfm:978-3-540-38812-
8/1.pdf.

Butcher:1980:SIS

[But80] J. C. Butcher. Some implementation schemes for implicit Runge– Kutta methods. Lecture Notes in Mathematics, 773:12–24, 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094160/.

Butler:1981:GGA

[But81] M. C. R. Butler. Grothendieck groups and almost split sequences. Lecture Notes in Mathematics, 882:357–368, 1981. CODEN LN-MAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092505/.

Butkovic:1982:CNC

[But82] Davor Butković. Completions and the null-completion of vector measures. Lecture Notes in Mathematics, 948:230–234, 1982. CODEN LNMAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069847/.

Berline:1981:EHN

[BV81] Nicole Berline and Michèle Vergne. Équations de Hua et noyau de Poisson. (French) []. Lecture Notes in Mathematics, 880:1– 51, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090401/.

Borceux:1983:ALT

[BV83a] Francis Borceux and Gilberte Van den Bossche. Algebra in a Localic Topos with Applications to Ring Theory, volume 1038 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 240 pp. LCCN QA3 .L28 no. 1038. URL http:/

/link.springer.com/book/10.1007/BFb0073030;   http://
www.springerlink.com/content/978-3-540-38696-4.

Borceux:1983:AMT

[BV83b] Francis Borceux and Gilberte Van den Bossche. Applications to module theory. Lecture Notes in Mathematics, 1038:128– 138, 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073036/.

Borceux:1983:CAS

[BV83c] Francis Borceux and Gilberte Van den Bossche. Categories of algebraic sheaves. Lecture Notes in Mathematics, 1038:1–12, 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073031/.

Borceux:1983:FIS

[BV83d] Francis Borceux and Gilberte Van den Bossche. The formal initial segments. Lecture Notes in Mathematics, 1038:13–51, 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073032/.

Borceux:1983:GR

[BV83e] Francis Borceux and Gilberte Van den Bossche. Gelfand rings. Lecture Notes in Mathematics, 1038:194–229, 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073038/.

Borceux:1983:ITC

[BV83f] Francis Borceux and Gilberte Van den Bossche. Integral theories and characterization theorem. Lecture Notes in Mathematics, 1038:108–120, 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073034/.

Borceux:1983:LAS

[BV83g] Francis Borceux and Gilberte Van den Bossche. Localizations and algebraic sheaves. Lecture Notes in Mathematics, 1038:52– 107, 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073033/.

Borceux:1983:PRR

[BV83h] Francis Borceux and Gilberte Van den Bossche. Pure representation of rings. Lecture Notes in Mathematics, 1038:139–193, 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073037/.

Borceux:1983:ST

[BV83i] Francis Borceux and Gilberte Van den Bossche. Spectrum of a theory. Lecture Notes in Mathematics, 1038:121–127, 1983. CODEN LNMAA2. ISBN 3-540-12711-9 (print), 3-540-38696-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073035/.

Butuzov:1983:SPD

[BV83j] V. F. Butuzov and A. B. Vasil’eva. Singularly perturbed differential equations of parabolic type. Lecture Notes in Mathematics, 985:38–75, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062362/.

Brezinski:1984:PTA

[BvI84] Claude Brezinski and Jeannette van Iseghem. Padé-type approximants and linear functional transformations. Lecture Notes in Mathematics, 1105:100–108, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072402/.

Broer:1983:IMS

[BvS83] Henk Broer and Sebastian van Strien. Infinitely many’moduli of strong stability in divergence free unfoldings of singularities of vector fields. Lecture Notes in Mathematics, 1007:39–59, 1983.

CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061409/.

Berndt:1981:CRS

[BW81a] Bruce C. Berndt and B. M. Wilson. Chapter 5. Ramanujan’s second notebook. Lecture Notes in Mathematics, 899:49–78, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096453/.

Blanc:1981:HGL

[BW81b] Philippe Blanc and David Wigner. Homologie des groupes de Lie et dualité de Poincaré. (French) []. Lecture Notes in Mathematics, 880:52–56, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090402/.

Betti:1982:SCC

[BW82] R. Betti and R. F. C. Walters. The symmetry of the Cauchycompletion of a category. Lecture Notes in Mathematics, 962: 8–12, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066880/.

Barlow:1980:CMC

[BY80] M. T. Barlow and M. Yor. Sur la construction d’une martingale continue, de valeur absolue donnée. (French) []. Lecture Notes in Mathematics, 784:62–75, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089470/.

Bennett:1982:MNC

[BY82] Bruce Bennett and Stephen S.-T. Yau. Milnor number and classification of isolated singularities of holomorphic maps. Lecture Notes in Mathematics, 949:1–34, 1982. CODEN LNMAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069753/.

Byczkowski:1980:IGR

[BŻ80] T. Byczkowski and T. Żak. On the integrability of Gaussian random vectors. Lecture Notes in Mathematics, 828:21–29, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097391/.

Byczkowski:1984:DCS

[BŻ84] T. Byczkowski and T. Żak. Decomposition of convolution semigroups of probability measures on groups. Lecture Notes in Mathematics, 1064:23–35, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073631/.

Cornelius:1982:DAC

[CA82] H. Cornelius and G. Alefeld. A device for the acceleration of convergence of a monotonously enclosing iteration method. Lecture Notes in Mathematics, 953:68–79, 1982. CODEN LN-MAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069374/.

Caelli:1980:GSC

[Cae80] Terry Caelli. On generating spatial configurations with identical interpoint distance distributions. Lecture Notes in Mathematics, 829:69–75, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088901/.

Cairoli:1980:EDI

[Cai80] R. Cairoli. Sur l’extension de la definition d’intégrale stochastique. (French) []. Lecture Notes in Mathematics, 784:18–25, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089466/.

Cajar:1981:BDP

[Caj81a] Helmut Cajar. Billingsley Dimension in Probability Spaces, volume 892 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11164-6 (print), 3-540-38638-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 106 pp. LCCN QA3 .L28 no. 892. URL http:/

/link.springer.com/book/10.1007/BFb0091622;   http://
www.springerlink.com/content/978-3-540-38638-4.

Cajar:1981:BDS

[Caj81b] Helmut Cajar. The Billingsley dimension of saturated sets. Lecture Notes in Mathematics, 892:52–101, 1981. CODEN LN-MAA2. ISBN 3-540-11164-6 (print), 3-540-38638-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091625/.

Cajar:1981:I

[Caj81c] Helmut Cajar. Introduction. Lecture Notes in Mathematics, 892: 1–6, 1981. CODEN LNMAA2. ISBN 3-540-11164-6 (print), 3-540-38638-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091623/.

Cajar:1981:D

[Caj81d] Helmut Cajar. p-dimension. Lecture Notes in Mathematics, 892: 7–51, 1981. CODEN LNMAA2. ISBN 3-540-11164-6 (print), 3-540-38638-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091624/.

Calabi:1981:GDA

[Cal81] E. Calabi. Géométrie différentielle affine des hypersurfaces. (French) []. Lecture Notes in Mathematics, 901:189–204, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097198/.

Campana:1980:IDC

[Cam80a] Frédéric Campana. Images directes de cycles compacts par un morphisme et application à l’espace des cycles des tores. (French) []. Lecture Notes in Mathematics, 807:25–65, 1980. CODEN LNMAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090874/.

Campillo:1980:ACP

[Cam80b] Antonio Campillo. Algebroid Curves in Positive Characteristic, volume 813 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10022-9 (print), 3-540-38178-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 168 pp. LCCN QA3 .L28 no. 813. URL http:/

/link.springer.com/book/10.1007/BFb0090823;   http://
www.springerlink.com/content/978-3-540-38178-5.

Campillo:1980:CEP

[Cam80c] Antonio Campillo. Characteristic exponents of plane algebroid curves. Lecture Notes in Mathematics, 813:69–116, 1980. CODEN LNMAA2. ISBN 3-540-10022-9 (print), 3-540-38178-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090826/.

Campillo:1980:HNE

[Cam80d] Antonio Campillo. Hamburger–Noether expansions of algebroid curves. Lecture Notes in Mathematics, 813:45–68, 1980. CODEN LNMAA2. ISBN 3-540-10022-9 (print), 3-540-38178-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090825/.

Campillo:1980:OSI

[Cam80e] Antonio Campillo. Other systems of invariants for the equisingularity of plane algebroid curves. Lecture Notes in Mathematics, 813:117–140, 1980. CODEN LNMAA2. ISBN 3-540-10022-9 (print), 3-540-38178-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090827/.

Campillo:1980:PAC

[Cam80f] Antonio Campillo. Parametrizations of algebroid curves. Lecture Notes in Mathematics, 813:1–44, 1980. CODEN LNMAA2. ISBN 3-540-10022-9 (print), 3-540-38178-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090824/.

Campillo:1980:TAC

[Cam80g] Antonio Campillo. Twisted algebroid curves. Lecture Notes in Mathematics, 813:141–163, 1980. CODEN LNMAA2. ISBN 3-540-10022-9 (print), 3-540-38178-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090828/.

Cameron:1982:OE

[Cam82a] Peter J. Cameron. Orbits and enumeration. Lecture Notes in Mathematics, 969:86–99, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062988/.

Cameron:1982:OEC

[Cam82b] Peter J. Cameron. Orbits, enumeration and colouring. Lecture Notes in Mathematics, 952:34–66, 1982. CODEN LN-MAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061972/.

Cannonito:1984:SAP

[Can84] Frank B. Cannonito. On some algorithmic problems for finitely presented groups and Lie algebras. Lecture Notes in Mathematics, 1098:21–28, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099658/.

Capozzi:1983:SAH

[Cap83] A. Capozzi. On subquadratic not-autonomous Hamiltonian systems. Lecture Notes in Mathematics, 1017:122–131, 1983. CO-DEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103243/.

Caraman:1980:CCC

[Car80a] Petru Caraman. p-capacity and conformal capacity in infinite dimensional spaces. Lecture Notes in Mathematics, 798:68– 108, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097257/.

Carleson:1980:SAP

[Car80b] Lennart Carleson. Some analytic problems related to statistical mechanics. Lecture Notes in Mathematics, 779:5–45, 1980. CODEN LNMAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087666/.

Carmona:1980:IDN

[Car80c] René Carmona. Infinite dimensional Newtonian potentials. Lecture Notes in Mathematics, 828:30–43, 1980. CODEN LN-MAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097392/.

Cartier:1980:TDP

[Car80d] Pierre Cartier. Théorie de la diffusion pour l’équation de Schrödinger. (French) []. Lecture Notes in Mathematics, 770:132– 150, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096240/.

Carvalho:1980:ACE

[Car80e] L. A. V. Carvalho. An analysis of the characteristic equation of the scalar linear difference equation with two delays. Lecture Notes in Mathematics, 799:69–81, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089309/.

Carlson:1981:CKD

[Car81a] Jon F. Carlson. Complexity and Krull dimension. Lecture Notes in Mathematics, 903:62–67, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092985/.

Cartier:1981:CLL

[Car81b] Pierre Cartier. La conjecture locale de Langlands pour GL(2) et la démonstration de Ph. Kutzko. (French) []. Lecture Notes in Mathematics, 842:112–138, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089931/.

Cartier:1981:AHC

[Car81c] Pierre Cartier. Les arrangements d’hyperplans: Un chapitre de géométrie combinatoire. (French) []. Lecture Notes in Mathematics, 901:1–22, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097186/.

Carrell:1982:GAV

[Car82] James B. Carrell, editor. Group Actions and Vector Fields: Proceedings of a Polish–North American Seminar Held at the University of British Columbia January 15–February 15, 1981, volume 956 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 956. URL http:/

/link.springer.com/book/10.1007/BFb0101503;   http://
www.springerlink.com/content/978-3-540-39528-7.

Caraman:1983:AEB

[Car83a] Petru Caraman. About the equality between the p-module and the p-capacity in R^{n}. Lecture Notes in Mathematics, 1039:32– 83, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073357/.

Caraman:1983:QMA

[Car83b] Petru Caraman. Quasiconformal mappings in abstract Wiener spaces. Lecture Notes in Mathematics, 1013:23–49, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066516/.

Carmona:1983:CMA

[Car83c] Jacques Carmona. Sur la classification des modules admissibles irredutibles. (French) []. Lecture Notes in Mathematics, 1020: 11–34, 1983. CODEN LNMAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071495/.

Casas:1982:MAP

[Cas82a] E. Casas. Moduli of algebroid plane curves. Lecture Notes in Mathematics, 961:32–83, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071275/.

CastroB:1982:RMM

[Cas82b] Alfonso Castro B. Reduction methods via minimax. Lecture Notes in Mathematics, 957:1–20, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066231/.

Casse:1983:CMX

[Cas83a] Louis Reynolds Antoine Casse, editor. Combinatorial Mathematics X: Proceedings of the Conference held in Adelaide, Australia, August 23–27, 1982, volume 1036 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1036. URL http:/

/link.springer.com/book/10.1007/BFb0071504;   http://
www.springerlink.com/content/978-3-540-38694-0.

Casselman:1983:AFH

[Cas83b] William Casselman. Automorphic forms and a Hodge theory for congruence subgroups of SL_{2}(Z). Lecture Notes in Mathematics, 1041:103–140, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073146/.

Casati:1984:ICQ

[Cas84a] Giulio Casati. Irreversibility and chaos in quantum systems. Lecture Notes in Mathematics, 1055:62–75, 1984. CODEN LN-MAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071710/.

Castaing:1984:NCE

[Cas84b] C. Castaing. A new class of evolution equation in a Hilbert space. Lecture Notes in Mathematics, 1091:117–128, 1984. CODEN LN-MAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098805/.

Cathey:1981:SST

[Cat81] Fritz Cathey. Strong shape theory. Lecture Notes in Mathematics, 870:215–238, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089717/.

Catanese:1983:MSG

[Cat83a] F. Catanese. Moduli of surfaces of general type. Lecture Notes in Mathematics, 997:90–112, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061637/.

Catanese:1983:RCM

[Cat83b] F. Catanese. On the rationality of certain moduli spaces related to curves of genus 4. Lecture Notes in Mathematics, 1008:30– 50, 1983. CODEN LNMAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065697/.

Catanese:1984:CAM

[Cat84] Fabrizio Catanese. Commutative algebra methods and equations of regular surfaces. Lecture Notes in Mathematics, 1056: 68–111, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071770/.

Cauchon:1980:CML

[Cau80a] G. Cauchon. Commutant des modules de longueur finie sur certaines algèbres filtrées. (French) []. Lecture Notes in Mathematics, 825:10–18, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089117/.

Cauchon:1980:CQA

[Cau80b] Gérard Cauchon. Coeur de k[x, σ, δ] et questions d’algébricité dans k(x, σ, δ). (French) []. Lecture Notes in Mathematics, 795: 179–190, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090119/.

Cazacu:1983:OMS

[Caz83] Cabiria Andreian Cazacu. Sur l’oeuvre mathématique de Simion Stoilow. (French) []. Lecture Notes in Mathematics, 1013:8– 21, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066515/.

Charpentier:1984:ECP

[CB84a] Philippe Charpentier and Aline Bonami. Estimations des (1 − 1) courants positifs fermés dans les domaines de C^{2}. (French) []. Lecture Notes in Mathematics, 1094:44–52, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099154/.

Curras-Bosch:1984:KVF

[CB84b] C. Currás-Bosch. Killing vector fields and complex structures. Lecture Notes in Mathematics, 1045:36–42, 1984. CODEN LN-MAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072163/.

Cazacu:1983:CAFa

[CBJS83a] Cabiria Andreian Cazacu, Nicu Boboc, Martin Jurchescu, and Ion Suciu, editors. Complex Analysis — Fifth Romanian–Finnish Seminar: Part 1 Proceedings of the Seminar held in Bucharest, June 28–July 3, 1981, volume 1013 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1013. URL http:/

/link.springer.com/book/10.1007/BFb0066513;   http://
www.springerlink.com/content/978-3-540-38671-1.

Cazacu:1983:CAFb

[CBJS83b] Cabiria Andreian Cazacu, Nicu Boboc, Martin Jurchescu, and Ion Suciu, editors. Complex Analysis — Fifth Romanian–Finnish Seminar: Part 2 Proceedings of the Seminar held in Bucharest, June 28–July 3, 1981, volume 1014 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1014. URL http:/

/link.springer.com/book/10.1007/BFb0072065;   http://
www.springerlink.com/content/978-3-540-38672-8.

Calamai:1982:SOM

[CC82a] P. H. Calamai and A. R. Conn. A second-order method for solving the continuous multifacility location problem. Lecture Notes in Mathematics, 912:1–25, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093145/.

Campillo:1982:PSA

[CC82b] A. Campillo and J. Castellanos. On projections of space algebroid curves. Lecture Notes in Mathematics, 961:22–31, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071274/.

Chudnovsky:1982:RTI

[CC82c] D. Chudnovsky and G. Chudnovsky. Resolvent and trace identities in the one dimensional case. Lecture Notes in Mathematics, 925:215–220, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093511/.

Chudnovsky:1982:HSI

[CC82d] D. V. Chudnovsky and G. V. Chudnovsky. Hamiltonian structure of isospectral deformation equations. Elliptic curve case. Lecture Notes in Mathematics, 925:134–146, 1982. CODEN LN-MAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093505/.

Chudnovsky:1982:QHA

[CC82e] D. V. Chudnovsky and G. V. Chudnovsky. Quantum Hamiltonians associated with finite-dimensional Lie algebras and factorized s-matrices. Lecture Notes in Mathematics, 925:147–156, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093506/.

Chudnovsky:1982:RPC

[CC82f] David V. Chudnovsky and Gregory V. Chudnovsky, editors. The Riemann Problem, Complete Integrability and Arithmetic Applications: Proceedings of a Seminar Held at the Institut des Hautes Études Scientifiques, Bures-sur Yvette, France, and at Columbia University, New York, U.S.A., 1979–1980, volume 925 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 925.

URL http://link.springer.com/book/10.1007/BFb0093494;
 http://www.springerlink.com/content/978-3-540-39152-
4.

Colbourn:1983:DBD

[CC83] Charles J. Colbourn and Marlene J. Colbourn. Decomposition of block designs: Computational issues. Lecture Notes in Mathematics, 1036:141–146, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071515/.

Chudnovsky:1984:PRA

[CC84a] D. V. Chudnovsky and G. V. Chudnovsky. Padé and rational approximations to systems of functions and their arithmetic applications. Lecture Notes in Mathematics, 1052:37–84, 1984. CODEN LNMAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071540/.

Chudnovsky:1984:PAS

[CC84b] D. V. Chudnovsky and G. V. Chudnovsky. Padé approximations to solutions of linear differential equations and applications to Diophantine analysis. Lecture Notes in Mathematics, 1052:85– 167, 1984. CODEN LNMAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071541/.

Chudnovsky:1984:NTS

[CCCN84] David V. Chudnovsky, Gregory V. Chudnovsky, Harvey Cohn, and Melvin B. Nathanson, editors. Number Theory: A Seminar held at the Graduate School and University Center of the City University of New York 1982, volume 1052 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1052.

URL http://link.springer.com/book/10.1007/BFb0071537;
 http://www.springerlink.com/content/978-3-540-38788-
6.

Chudnovsky:1982:CQO

[CCN82] D. V. Chudnovsky, G. V. Chudnovsky, and A. Neveu. Classical and quantum operator nonlinear Schrödinger equation. I. Lecture Notes in Mathematics, 925:157–177, 1982. CODEN LN-MAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093507/.

Cline:1982:GLC

[CCV82] Alan K. Cline, Andrew R. Conn, and Charles F. Van Loan. Generalizing the LINPACK condition estimator. Lecture Notes in Mathematics, 909:73–83, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092961/.

Cannon:1980:IVP

[CD80] J. R. Cannon and Emmanuele DiBenedetto. The initial value problem for the Boussinesq equations with data in L^{p}. Lecture Notes in Mathematics, 771:129–144, 1980. CODEN LN-MAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086903/.

Cherlin:1981:NN

[CD81] Gregory L. Cherlin and Max A. Dikmann. Note on a Nullstellensatz. Lecture Notes in Mathematics, 890:111–114, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095659/.

Coddington:1981:EES

[CdS81a] Earl A. Coddington and Hendrik S. V. de Snoo. Eigenfunction expansions for selfadjoint subspaces. Lecture Notes in Mathematics, 858:168–182, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089066/.

Coddington:1981:FGS

[CdS81b] Earl A. Coddington and Hendrik S. V. de Snoo. Forms generated by selfadjoint extensions of M_{0}. Lecture Notes in Mathematics, 858:36–53, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089060/.

Coddington:1981:HSG

[CdS81c] Earl A. Coddington and Hendrik S. V. de Snoo. Hilbert spaces generated by positive selfadjoint extensions of M_{o}. Lecture Notes in Mathematics, 858:54–63, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089061/.

Coddington:1981:IS

[CdS81d] Earl A. Coddington and Hendrik S. V. de Snoo. Intermediate subspaces. Lecture Notes in Mathematics, 858:70–105, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089063/.

Coddington:1981:I

[CdS81e] Earl A. Coddington and Hendrik S. V. de Snoo. Introduction. Lecture Notes in Mathematics, 858:1–20, 1981. CODEN LN-MAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089058/.

Coddington:1981:MMS

[CdS81f] Earl A. Coddington and Hendrik S. V. de Snoo. Minimal and maximal subspaces for the pair l,m. Lecture Notes in Mathematics, 858:64–69, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089062/.

Coddington:1981:RBV

[CdS81g] Earl A. Coddington and Hendrik S. V. de Snoo. Regular Boundary Value Problems Associated with Pairs of Ordinary Differential Expressions, volume 858 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 219 (est.) pp. LCCN QA3 .L28 no. 858. URL http:

//link.springer.com/book/10.1007/BFb0089057;   http:/
/www.springerlink.com/content/978-3-540-38670-4.

Coddington:1981:R

[CdS81h] Earl A. Coddington and Hendrik S. V. de Snoo. Resolvents. Lecture Notes in Mathematics, 858:138–167, 1981. CODEN LN-MAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089065/.

Coddington:1981:SEa

[CdS81i] Earl A. Coddington and Hendrik S. V. de Snoo. Selfadjoint extensions of M_{0}. Lecture Notes in Mathematics, 858:21–35, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089059/.

Coddington:1981:SIS

[CdS81j] Earl A. Coddington and Hendrik S. V. de Snoo. Semibounded intermediate subspaces. Lecture Notes in Mathematics, 858:183– 204, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089067/.

Coddington:1981:SSC

[CdS81k] Earl A. Coddington and Hendrik S. V. de Snoo. Some special cases. Lecture Notes in Mathematics, 858:205–219, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089068/.

Coddington:1981:SEb

[CdS81l] Earl A. Coddington and Hendrik S. V. de Snoo. Spectra and eigenvalues. Lecture Notes in Mathematics, 858:106–137, 1981. CODEN LNMAA2. ISBN 3-540-10706-1 (print), 3-540-38670-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089064/.

Cecchini:1984:NIC

[Cec84] C. Cecchini. Noncommutative integration and conditioning. Lecture Notes in Mathematics, 1055:76–85, 1984. CODEN LN-MAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071711/.

Cegrell:1980:PCA

[Ceg80] Urban Cegrell. On product capacities with application to complex analysis. Lecture Notes in Mathematics, 822:33–45, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097747/.

Cegielski:1981:TEM

[Ceg81] Patrick Cegielski. Théorie élémentaire de la multiplication des entiers naturels. (French) []. Lecture Notes in Mathematics, 890: 44–89, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095657/.

Cegrell:1983:ECM

[Ceg83] Urban Cegrell. An estimate of the complex Monge–Ampère operator. Lecture Notes in Mathematics, 1039:84–87, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073358/.

Cegrell:1984:DCM

[Ceg84] Urban Cegrell. On the discontinuity of the complex Monge– Ampère operator. Lecture Notes in Mathematics, 1094:29–31, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099152/.

Cellina:1984:FPT

[Cel84] Arrigo Cellina. A fixed point theorem for subsets of L^{1}. Lecture Notes in Mathematics, 1091:129–137, 1984. CODEN LN-MAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098806/.

Cerf:1981:FFN

[Cer81] Jean Cerf. 1-formes fermées non singulières sur les variétés compactes de dimension 3. (French) []. Lecture Notes in Mathematics, 901:205–219, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097199/.

Cercignani:1984:KTB

[Cer84a] Carlo Cercignani, editor. Kinetic Theories and the Boltzmann Equation: Lectures given at the 1st 1981 Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) Held at Montecatini, Italy, June 10–18, 1981, volume 1048 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1048. URL http:/

/link.springer.com/book/10.1007/BFb0071876;   http://
www.springerlink.com/content/978-3-540-38777-0.

Cerin:1984:GPM

[Čer84b] Zvonko Čerin. On global properties of maps IV. Lecture Notes in Mathematics, 1060:129–139, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099929/.

Cesari:1983:ESE

[Ces83] Lamberto Cesari. Existence of solutions and existence of optimal solutions. Lecture Notes in Mathematics, 979:88–107, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066250/.

Cesari:1984:NO

[Ces84] L. Cesari. Nonlinear optimization. Lecture Notes in Mathematics, 1107:63–81, 1984. CODEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101493/.

CastroB:1982:MSN

[CG82] Alfonso Castro B. and J. V. A. Gonçalves. On multiple solutions of nonlinear elliptic equations with odd nonlinearities. Lecture Notes in Mathematics, 957:21–33, 1982. CODEN LN-MAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066232/.

Calais:1983:MLC

[CG83a] J. Y. Calais and M. Genin. Sur les martingales locales continues indexées par ]0, ∞[. (French) []. Lecture Notes in Mathematics, 986:162–178, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068311/.

Carnoy:1983:PUL

[CG83b] E. G. Carnoy and M. Geradin. On the practical use of the Lanczos algorithm in finite element applications to vibration and bifurcation problems. Lecture Notes in Mathematics, 973:156–176, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062100/.

Ciliberto:1983:AGO

[CGO83] Ciro Ciliberto, Franco Ghione, and Ferruccio Orecchia, editors. Algebraic Geometry — Open Problems: Proceedings of the Conference Held in Ravello, May 31–June 5, 1982, volume 997 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 997. URL http:/

/link.springer.com/book/10.1007/BFb0061633;   http://
www.springerlink.com/content/978-3-540-40047-9.

Cossart:1984:RSS

[CGO84] Vincent Cossart, Jean Giraud, and Ulrich Orbanz. Resolution of Surface Singularities: Three Lectures with an Appendix by H. Hironaka, volume 1101 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13904-4 (print), 3-540-39125-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 132 pp. LCCN QA3 .L28 no. 1101. URL http:/

/link.springer.com/book/10.1007/BFb0072258;   http://
www.springerlink.com/content/978-3-540-39125-8.

Chen:1984:CAC

[CH84] C. C. Chen and D. A. Holton. Cycles in abelian Cayley graphs with a proscribed vertex. Lecture Notes in Mathematics, 1073: 22–30, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073102/.

Chamarie:1980:MOA

[Cha80a] M. Chamarie. Maximal orders applied to enveloping algebras. Lecture Notes in Mathematics, 825:19–27, 1980. CODEN LN-MAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089118/.

Chatterji:1980:MPV

[Cha80b] S. D. Chatterji. Multiparameter processes and vector-valued processes. Lecture Notes in Mathematics, 828:44–46, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097393/.

Chan:1981:SOP

[Cha81a] Y. K. Chan. On some open problems in constructive probability theory. Lecture Notes in Mathematics, 873:44–53, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090726/.

Chatzidakis:1981:RTF

[Cha81b] Zoè Chatzidakis. La représentation en termes de faisceaux des modèles de la théorie élémentaire de la multiplication des entiers naturels. (French) []. Lecture Notes in Mathematics, 890: 91–110, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095658/.

Chawathe:1981:PGD

[Cha81c] P. D. Chawathe. 2-2 perfect graphic degree sequences. Lecture Notes in Mathematics, 885:195–202, 1981. CODEN LN-MAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092263/.

Chao:1982:BRM

[Cha82a] J.-A. Chao. H^{P} and BMO regular martingales. Lecture Notes in Mathematics, 908:274–284, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093294/.

Chao:1982:HSR

[Cha82b] J.-A. Chao. Hardy spaces on regular martingales. Lecture Notes in Mathematics, 939:18–28, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096256/.

Chamarie:1983:MAK

[Cha83a] Marc Chamarie. Modules sur les anneaux de Krull non commutatifs. (French) []. Lecture Notes in Mathematics, 1029:283– 310, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098936/.

Chaperon:1983:LCH

[Cha83b] Marc Chaperon. On the local classification of holomorphic vector fields. Lecture Notes in Mathematics, 1007:96–103, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061412/.

Chapman:1983:A

[Cha83c] T. A. Chapman. Applications. Lecture Notes in Mathematics, 1009:5–11, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066068/.

Chapman:1983:Cc

[Cha83d] T. A. Chapman. Calculations. Lecture Notes in Mathematics, 1009:70–79, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066079/.

Chapman:1983:Cb

[Cha83e] T. A. Chapman. Construction of K_{0}(Y)_{ϵ}. Lecture Notes in Mathematics, 1009:26–28, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066073/.

Chapman:1983:Ca

[Cha83f] T. A. Chapman. Construction of Wh(y)^{ϵ}. Lecture Notes in Mathematics, 1009:13–17, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066070/.

Chapman:1983:CBT

[Cha83g] T. A. Chapman. The controlled boundary theorem. Lecture Notes in Mathematics, 1009:79–84, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066080/.

Chapman:1983:CFO

[Cha83h] T. A. Chapman. Controlled finiteness obstruction. Lecture Notes in Mathematics, 1009:28–33, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066074/.

Chapman:1983:CCT

[Cha83i] T. A. Chapman. The controlled s-Cobordism theorem. Lecture Notes in Mathematics, 1009:84–89, 1983. CODEN LN-MAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066081/.

Chapman:1983:CSH

[Cha83j] T. A. Chapman. Controlled Simple Homotopy Theory and Applications, volume 1009 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 94 pp. LCCN QA3 .L28 no. 1009. URL http:/

/link.springer.com/book/10.1007/BFb0066066;   http://
www.springerlink.com/content/978-3-540-40973-1.

Chapman:1983:CWT

[Cha83k] T. A. Chapman. Controlled Whitehead torsion. Lecture Notes in Mathematics, 1009:19–26, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066072/.

Chapman:1983:DN

[Cha83l] T. A. Chapman. Definitions and notation. Lecture Notes in Mathematics, 1009:11–13, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066069/.

Chapman:1983:FP

[Cha83m] T. A. Chapman. Functorial properties. Lecture Notes in Mathematics, 1009:17–19, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066071/.

Chapman:1983:FPC

[Cha83n] T. A. Chapman. Further properties of the controlled finiteness obstruction. Lecture Notes in Mathematics, 1009:33–41, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066075/.

Chapman:1983:I

[Cha83o] T. A. Chapman. Introduction. Lecture Notes in Mathematics, 1009:1–5, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066067/.

Chapman:1983:RT

[Cha83p] T. A. Chapman. The realization theorem. Lecture Notes in Mathematics, 1009:60–70, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066078/.

Chapman:1983:SH

[Cha83q] T. A. Chapman. The splitting homomorphism. Lecture Notes in Mathematics, 1009:41–50, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066076/.

Chapman:1983:SS

[Cha83r] T. A. Chapman. The splitting sequence. Lecture Notes in Mathematics, 1009:50–59, 1983. CODEN LNMAA2. ISBN 3-540-12338-5 (print), 3-540-40973-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066077/.

Charpentier:1983:MSN

[Cha83s] Benito Chen Charpentier. A method for solving nonlinear parabolic equations with moving boundaries. Lecture Notes in Mathematics, 1005:37–45, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0112524/.

Charriere:1983:EAS

[Cha83t] Hélène Charrière. Étude asymptotique dans des secteurs spi-

ralés autour de l’origine de l’équation ∑_{i=1}^{n} λ_{i}x_{i} \frac{∂u}{∂x_{i}} = F (x, u) et

théorème de Borel–Ritt. (French) []. Lecture Notes in Mathematics, 1028:69–82, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071678/.

Chaffy:1984:FEI

[Cha84a] C. Chaffy. Formule d’erreur dans l’interpolation rationnelle multipoints d’une fonction de la variable complexe. (French) []. Lecture Notes in Mathematics, 1071:49–61, 1984. CODEN LN-MAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099609/.

Charatonik:1984:SPG

[Cha84b] J. J. Charatonik. Some problems on generalized homogeneity of continua. Lecture Notes in Mathematics, 1060:1–6, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099914/.

Charatonik:1984:PKH

[Cha84c] Włodzimierz J. Charatonik. On the property of Kelley in hyperspaces. Lecture Notes in Mathematics, 1060:7–10, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099915/.

Charpentier:1984:ZFT

[Cha84d] Philippe Charpentier. Sur les zéros des fonctions de type Nevanlinna dans le bidisque. (French) []. Lecture Notes in Mathematics, 1094:32–43, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099153/.

Cheney:1980:BAT

[Che80a] E. W. Cheney. Best approximation in tensor product spaces. Lecture Notes in Mathematics, 773:25–32, 1980. CODEN LN-MAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094161/.

Cherlin:1980:RCF

[Che80b] Gregory Cherlin. Rings of continuous functions: Decision problems. Lecture Notes in Mathematics, 834:44–91, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090160/.

Chevalier:1981:ITP

[Che81a] L. Chevalier. L^{p}-inequalities for two-parameter martingales. Lecture Notes in Mathematics, 851:470–475, 1981. CODEN LN-MAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088737/.

Chevet:1981:KAC

[Che81b] Simone Chevet. Kernel associated with a cylindrical measure. Lecture Notes in Mathematics, 860:51–84, 1981. CODEN LN-MAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090608/.

Chevet:1983:GML

[Che83] Simone Chevet. Gaussian measures and large deviations. Lecture Notes in Mathematics, 990:30–46, 1983. CODEN LN-MAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064261/.

Cherlin:1984:DTP

[Che84a] Gregory L. Cherlin. Dectdable theories of pseudo-algebraically closed fields. Lecture Notes in Mathematics, 1103:83–101, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099382/.

Cherlin:1984:DPS

[Che84b] Gregory L. Cherlin. Definability in power series rings of nonzero characteristic. Lecture Notes in Mathematics, 1103:102–112, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099383/.

Chitashvili:1983:MIT

[Chi83] R. J. Chitashvili. Martingale ideology in the theory of controlled stochastic processes. Lecture Notes in Mathematics, 1021:73– 92, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072905/.

Chinburg:1984:MGS

[Chi84a] Ted Chinburg. Multiplicative Galois structure. Lecture Notes in Mathematics, 1068:23–32, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099439/.

Chisholm:1984:CFS

[Chi84b] J. S. R. Chisholm. Continued fraction solution of the general Riccati equation. Lecture Notes in Mathematics, 1105:109–116, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072403/.

Chou:1980:CSS

[Cho80a] Ching-Sung Chou. Une caractérisation des semimartingales spéciales. (French) [A characterization of special semimartingales]. Lecture Notes in Mathematics, 784:116–117, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089475.

Chow:1980:PSN

[Cho80b] Shui-Nee Chow. Periodic solutions of nonlinear autonomous hyperbolic equations. Lecture Notes in Mathematics, 799:126– 139, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089312/.

Choudum:1981:CFO

[Cho81] S. A. Choudum. Characterization of forcibly outerplanar graphic sequences. Lecture Notes in Mathematics, 885:203–211, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092264/.

Choquet:1983:ICC

[Cho83a] Gustave Choquet. II. Convexes et cônes convexes non localement compacts. (French) []. Lecture Notes in Mathematics, 1033: 125–135, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099851/.

Choquet:1983:IFL

[Cho83b] Gustave Choquet. III. Formes linéaires positives et mesures. (French) []. Lecture Notes in Mathematics, 1033:135–143, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099852/.

Choquet:1983:RIF

[Cho83c] Gustave Choquet. Représentation intégrale. (French) []. Lecture Notes in Mathematics, 1033:114–124, 1983. CODEN LN-MAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099850/.

Chou:1983:CIT

[Cho83d] C. S. Chou. Sur certaines inégalités de théorie des martingales. (French) []. Lecture Notes in Mathematics, 986:117–120, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068305/.

Chong:1984:FIM

[Cho84a] Chi-Tat Chong. The α-finite injury method. Lecture Notes in Mathematics, 1106:49–71, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099167/.

Chong:1984:DT

[Cho84b] Chi-Tat Chong. The density theorem. Lecture Notes in Mathematics, 1106:138–164, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099172/.

Chong:1984:HS

[Cho84c] Chi-Tat Chong. Hyperhypersimple sets. Lecture Notes in Mathematics, 1106:114–126, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099170/.

Chong:1984:I

[Cho84d] Chi-Tat Chong. Introduction. Lecture Notes in Mathematics, 1106:1–36, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099165/.

Chong:1984:JOG

[Cho84e] Chi-Tat Chong. The jump operator and 1-generic sets. Lecture Notes in Mathematics, 1106:37–48, 1984. CODEN LN-MAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099166/.

Chong:1984:MSb

[Cho84f] Chi-Tat Chong. Major subsets. Lecture Notes in Mathematics, 1106:90–113, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099169/.

Chong:1984:MSa

[Cho84g] Chi-Tat Chong. Maximal sets. Lecture Notes in Mathematics, 1106:72–89, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099168/.

Chong:1984:MP

[Cho84h] Chi-Tat Chong. Minimal pairs. Lecture Notes in Mathematics, 1106:127–137, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099171/.

Chong:1984:STM

[Cho84i] Chi-Tat Chong. Set-theoretic methods. Lecture Notes in Mathematics, 1106:181–195, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099174/.

Chong:1984:TAR

[Cho84j] Chi-Tat Chong. Techniques of Admissible Recursion Theory, volume 1106 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 214 pp. LCCN QA3 .L28 no. 1106. URL http:/

/link.springer.com/book/10.1007/BFb0099164;   http://
www.springerlink.com/content/978-3-540-39121-0.

Chong:1984:T

[Cho84k] Chi-Tat Chong. Trees. Lecture Notes in Mathematics, 1106:165– 180, 1984. CODEN LNMAA2. ISBN 3-540-13902-8 (print), 3-540-39121-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099173/.

Choquet:1984:AIA

[Cho84l] Gustave Choquet. Allocution introductive au colloque Deny. (French) [Introductory speech at the Deny symposium]. Lecture Notes in Mathematics, 1096:1–9, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100103/.

Chou:1984:CGI

[Cho84m] Ching-Sung Chou. Sur certaines généralisations de l’inégalité de Fefferman. (French) []. Lecture Notes in Mathematics, 1059:219– 222, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100046/.

Christensen:1980:SSB

[Chr80a] Jens Peter Reus Christensen. A survey of small ball theorems and problems. Lecture Notes in Mathematics, 794:24–30, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088208/.

Christiansen:1980:BTS

[Chr80b] P. L. Christiansen. On Bäcklund transformations and solutions to the 2 + 1 and 3 + 1-dimensional sine-Gordon equation. Lecture Notes in Mathematics, 810:43–62, 1980. CODEN LN-MAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089973/.

Christian:1983:EZB

[Chr83a] Ulrich Christian. Epstein’s zetafunctions of a binary quadratic form. Lecture Notes in Mathematics, 1030:1–28, 1983. CODEN LNMAA2. ISBN 3-540-12701-1 (print), 3-540-38688-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071380/.

Christian:1983:PM

[Chr83b] Ulrich Christian. Preparational material. Lecture Notes in Mathematics, 1030:28–88, 1983. CODEN LNMAA2. ISBN 3-540-12701-1 (print), 3-540-38688-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071381/.

Christian:1983:SESa

[Chr83c] Ulrich Christian. Selberg’s Eisenstein series. Lecture Notes in Mathematics, 1030:149–180, 1983. CODEN LNMAA2. ISBN 3-540-12701-1 (print), 3-540-38688-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071383/.

Christian:1983:SZS

[Chr83d] Ulrich Christian. Selberg’s zeta- and L-series. Lecture Notes in Mathematics, 1030:89–148, 1983. CODEN LNMAA2. ISBN 3-540-12701-1 (print), 3-540-38688-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071382/.

Christian:1983:SZE

[Chr83e] Ulrich Christian. Selberg’s Zeta-, L-, and Eisensteinseries, volume 1030 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12701-1 (print), 3-540-38688-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 196 pp. LCCN QA3 .L28 no. 1030. URL http:/

/link.springer.com/book/10.1007/BFb0071379;   http://
www.springerlink.com/content/978-3-540-38688-9.

Christian:1983:SESb

[Chr83f] Ulrich Christian. Siegel’s Eisenstein series. Lecture Notes in Mathematics, 1030:181–190, 1983. CODEN LNMAA2. ISBN 3-540-12701-1 (print), 3-540-38688-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071384/.

Chudnovsky:1982:ICM

[Chu82a] D. V. Chudnovsky. Infinite component σ-models and instanton solutions. Lecture Notes in Mathematics, 925:57–70, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093499/.

Chudnovsky:1982:ICT

[Chu82b] D. V. Chudnovsky. Infinite component two-dimensional completely integrable systems of KdV type. Lecture Notes in Mathematics, 925:71–84, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093500/.

Chudnovsky:1982:RAT

[Chu82c] D. V. Chudnovsky. The representation of an arbitrary, twodimensional completely integrable system as the common action of two commuting one-dimensional Hamiltonian flows. Lecture Notes in Mathematics, 925:85–94, 1982. CODEN LN-MAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093501/.

Churchill:1982:PNH

[Chu82d] Richard C. Churchill. On proving the nonintegrability of a Hamiltonian system. Lecture Notes in Mathematics, 925:103–122, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093503/.

Ciliberto:1983:PTT

[Cil83] Ciro Ciliberto. On a proof of Torelli’s theorem. Lecture Notes in Mathematics, 997:113–123, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061638/.

Chong:1984:MDG

[CJ84] C. T. Chong and C. G. Jockusch. Minimal degrees and 1-generic sets below 0′. Lecture Notes in Mathematics, 1104:63–77, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099479/.

Cowling:1984:HAH

[CK84] Michael Cowling and Adam Korányi. Harmonic analysis on Heisenberg type groups from a geometric viewpoint. Lecture Notes in Mathematics, 1077:60–100, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072337/.

Chadzynski:1983:PAM

[CKK83] Jacek Chadzyński, Tadeusz Krasiński, and Wojciech Kryszewski. On the parametric and algebraic multiplicities of an isolated zero of a holomorphic mapping. Lecture Notes in Mathematics, 1039: 88–101, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073359/.

Cohen:1983:SDL

[CKL83] H. Cohen, R. J. Koch, and J. D. Lawson. Semigroups defined by lattice polynomials. Lecture Notes in Mathematics, 998:50– 56, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062027/.

Choi:1982:TZP

[CKLR82] Man-Duen Choi, Manfred Knebusch, Tsit-Yuen Lam, and Bruce Reznick. Transversal zeros and positive semidefinite forms. Lecture Notes in Mathematics, 959:273–298, 1982. CODEN LN-MAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062259/.

Camacho:1983:EPS

[CKMN83] C. Camacho, M. Krych, R. Mañé, and Z. Nitecki. An extension of Peixoto’s structural stability theorem to open surfaces with finite genus. Lecture Notes in Mathematics, 1007:60–87, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061410/.

Chia:1984:SGI

[CL84a] Gek-Ling Chia and Chong-Keang Lim. On supercompact graphs III: The edge nucleus. Lecture Notes in Mathematics, 1073:1– 10, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073100/.

Ciliberto:1984:UCL

[CL84b] Ciro Ciliberto and Robert Lazarsfeld. On the uniqueness of certain linear series on some classes of curves. Lecture Notes in Mathematics, 1092:198–213, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099363/.

Cohen:1984:HCGa

[CL84c] H. Cohen and H. W. Lenstra, Jr. Heuristics on class groups. Lecture Notes in Mathematics, 1052:26–36, 1984. CODEN LN-MAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071539/.

Cohen:1984:HCGb

[CL84d] H. Cohen and H. W. Lenstra, Jr. Heuristics on class groups of number fields. Lecture Notes in Mathematics, 1068:33–62, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099440/.

Clarysse:1981:GRC

[Cla81] T. H. Clarysse. Generalized rational correctors. Lecture Notes in Mathematics, 888:110–123, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095580/.

Clarke:1983:MCP

[Cla83] R. J. Clarke. Modelling competitions by poset multiplication. Lecture Notes in Mathematics, 1036:128–140, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071514/.

Clemens:1982:DTS

[Cle82] C. H. Clemens. Degeneration techniques in the study of threefolds. Lecture Notes in Mathematics, 947:93–154, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093586/.

Clement:1984:AVE

[Clé84] Ph. Clément. On abstract Volterra equations in Banach spaces with completely positive kernels. Lecture Notes in Mathematics, 1076:32–40, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072763/.

Chassaing:1984:BSR

[CLM84] Philippe Chassaing, Gérard Letac, and Marianne Mora. Brocot sequences and random walks in SL(2, IR). Lecture Notes in Mathematics, 1064:36–48, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073632/.

Clote:1980:WPR

[Clo80] Peter Clote. Weak partition relations, finite games, and independence results in Peano arithmetic. Lecture Notes in Mathematics, 834:92–107, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090161/.

Clote:1981:ABT

[Clo81a] Peter Clote. Anti-basis theorems and their relation to independence results in Peano arithmetic. Lecture Notes in Mathematics, 890:115–133, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095660/.

Clote:1981:NDM

[Clo81b] Peter Clote. A note on decidable model theory. Lecture Notes in Mathematics, 890:134–142, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095661/.

Carr:1980:SNL

[CM80a] J. Carr and M. Z. M. Malhardeen. Stability of nonconservative linear systems. Lecture Notes in Mathematics, 799:45–68, 1980.

CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089308/.

Coifman:1980:FAM

[CM80b] R. R. Coifman and Y. Meyer. Fourier analysis of multilinear convolutions, Calderón’s theorem, and analysis on Lipschitz curves. Lecture Notes in Mathematics, 779:104–122, 1980. CODEN LN-MAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087669/.

Colombeau:1981:FLC

[CM81] Jean-François Colombeau and Reinhold Meise. C^{∞}-functions on locally convex and on bornological vector spaces. Lecture Notes in Mathematics, 843:195–216, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089275/.

Clarkson:1982:CFS

[CM82a] P. A. Clarkson and J. B. McLeod. A connection formula for the second Painlevé transcendent. Lecture Notes in Mathematics, 964:135–142, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064994/.

Cowling:1982:ETC

[CM82b] M. G. Cowling and A. M. Mantero. Examples of twisted convolution operators. Lecture Notes in Mathematics, 908:210–216, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093290/.

Canada:1983:PSN

[CMA83] A. Cañada and P. Martínez-Amores. Periodic solutions of neutral functional differential equations. Lecture Notes in Mathematics, 1017:115–121, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103242/.

Cohn:1980:EVS

[CMH80] P. M. Cohn and M. Mahdavi-Hezavehi. Extensions of valuations on skew fields. Lecture Notes in Mathematics, 825:28–41, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089119/.

Chow:1983:POI

[CMPY83] Shui-Nee Chow, John Mallet-Paret, and James A. Yorke. A periodic orbit index which is a bifurcation invariant. Lecture Notes in Mathematics, 1007:109–131, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061414/.

Chou:1980:ISP

[CMS80] C. S. Chou, P.-A. Meyer, and C. Stricker. Sur les intégrales stochastiques de processus prévisibles non bornés. (French) []. Lecture Notes in Mathematics, 784:128–139, 1980. CODEN LN-MAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089478/.

Coifman:1983:NEF

[CMS83] R. R. Coifman, Y. Meyer, and E. M. Stein. Un nouveal espace fonctionnel adapté à l’étude des opérateurs définis par des intégrales singulières. (French) []. Lecture Notes in Mathematics, 992:1–15, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069149/.

Coifman:1984:SIB

[CMT^{+}84] R. R. Coifman, Yves Meyer, G. C. Tumarkin, Benjamin Muckenhoupt, Peter W. Jones, S. V. Kisliakov, I.È. Verbitsky, N. Ya. Krupnik, Stephen Semmes, Richard Rochberg, John Garnett, Albert Baernstein, Donald Sarason, J. M. Anderson, Sun-Yung A. Chang, et al. Singular integrals, BMO, H^{p}. Lecture Notes in Mathematics, 1043:308–352, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072189/.

Cecchi:1980:CTI

[CMZ80] M. Cecchi, M. Marini, and P. L. Zezza. A compactness theorem for integral operators and applications. Lecture Notes in Mathematics, 799:119–125, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089311/.

Chen:1982:SSI

[CN82] B. Chen and A. Noyola. A study of the stability of the interface between two inmiscible viscous fluids. Lecture Notes in Mathematics, 909:125–131, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092966/.

Cohen:1984:CLS

[CN84a] F. R. Cohen and J. A. Neisendorfer. A construction of p-local H-spaces. Lecture Notes in Mathematics, 1051:351–359, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075576/.

Cwikel:1984:CRC

[CN84b] Michael Cwikel and Per Nilsson. The coincidence of real and complex interpolation methods for couples of weighted Banach lattices. Lecture Notes in Mathematics, 1070:54–65, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099093/.

Chern:1981:RRM

[CO81] S.-S. Chern and R. Osserman. Remarks on the Riemannian metric of a minimal submanifold. Lecture Notes in Mathematics, 894: 49–90, 1981. CODEN LNMAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096224/.

Coates:1981:WMW

[Coa81] John Coates. The work of Mazur and Wiles on cyclotomic fields. Lecture Notes in Mathematics, 901:220–242, 1981. CODEN LN-MAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097200/.

Coban:1984:SVM

[Čob84] M. M. Čoban. Set-valued mappings and extension of continuous mappings. Lecture Notes in Mathematics, 1060:11–16, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099916/.

Coeure:1982:APP

[Coe82] Gérad Coeuré. Allocution prononcée par monsieur le Professeur Gérard Coeuré de l’Université de Lille I. En l’honneur du Professeur Pierre Lelong. (French) []. Lecture Notes in Mathematics, 919:187–189, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097050/.

Cohn:1981:NPE

[Coh81] Harvey Cohn. The next Pellian equation. Lecture Notes in Mathematics, 899:221–230, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096463/.

Cohn:1982:TMF

[Coh82] P. M. Cohn. Torsion modules and the factorization of matrices. Lecture Notes in Mathematics, 951:1–11, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067320/.

Cohn:1983:RCF

[Coh83] Harvey Cohn. Remarks on the cyclotomic Fricke groups. Lecture Notes in Mathematics, 971:15–23, 1983. CODEN LN-MAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067069/.

Collatz:1980:MFB

[Col80a] L. Collatz. Monotonicity and free boundary value problems. Lecture Notes in Mathematics, 773:33–45, 1980. CODEN LN-MAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094162/.

Collomb:1980:ERP

[Col80b] Gérard Collomb. Estimation de la regression par la méthode des k points les plus proches avec noyau: quelques propriétés de convergence ponctuelle. (French) []. Lecture Notes in Mathematics, 821: 159–175, 1980. CODEN LNMAA2. ISBN 3-540-10239-6 (print), 3-540-38318-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097428/.

Colombeau:1980:HLC

[Col80c] J.-F. Colombeau. Holomorphy in locally convex spaces and operators on the Fock spaces. Lecture Notes in Mathematics, 822: 46–60, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097748/.

Colton:1980:CRS

[Col80d] David Colton. Continuation and reflection of solutions to parabolic partial differential equations. Lecture Notes in Mathematics, 827:54–82, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091374/.

Collatz:1981:RBP

[Col81a] L. Collatz. Remark on bifurcation problems with several parameters. Lecture Notes in Mathematics, 846:82–87, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089827/.

Collatz:1981:EFG

[Col81b] Prof. Dr. L. Collatz. Einschliessungssätze für Fixpunkte. (German) []. Lecture Notes in Mathematics, 878:108–116, 1981. CO-DEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090678/.

Collino:1982:FGFa

[Col82a] Alberto Collino. The fundamental group of the Fano surface, I. Lecture Notes in Mathematics, 947:209–218, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093589/.

Collino:1982:FGFb

[Col82b] Alberto Collino. The fundamental group of the Fano surface, II. Lecture Notes in Mathematics, 947:219–220, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093590.

Colton:1982:ISP

[Col82c] David Colton. The inverse scattering problem for acoustic waves. Lecture Notes in Mathematics, 964:143–161, 1982. CODEN LN-MAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064995/.

Collatz:1983:ITS

[Col83] Prof. Dr. Lothar Collatz. Inclusion theorems for solutions of differential equations with aid of pointwise or vector monotonicity. Lecture Notes in Mathematics, 1032:93–109, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076794/.

Coleman:1984:PTM

[Col84] Bernard D. Coleman. A phenomenological theory of the mechanics of cold drawing. Lecture Notes in Mathematics, 1063: 76–142, 1984. CODEN LNMAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072154/.

Combet:1982:CPC

[Com82a] Edmond Combet. Cas des points critiques non-dégénérés. (French) []. Lecture Notes in Mathematics, 937:2–10, 1982. CO-DEN LNMAA2. ISBN 3-540-11566-8 (print), 3-540-39272-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096005/.

Combet:1982:CSI

[Com82b] Edmond Combet. Cas des singularités isolées. (French) []. Lecture Notes in Mathematics, 937:23–50, 1982. CODEN LN-MAA2. ISBN 3-540-11566-8 (print), 3-540-39272-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096007/.

Combet:1982:CPF

[Com82c] Edmond Combet. Changements de phases. (French) []. Lecture Notes in Mathematics, 937:58–75, 1982. CODEN LN-MAA2. ISBN 3-540-11566-8 (print), 3-540-39272-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096009/.

Combet:1982:DAI

[Com82d] Edmond Combet. Développement asymptotique des intégrales exponentielles. (French) []. Lecture Notes in Mathematics, 937: 11–22, 1982. CODEN LNMAA2. ISBN 3-540-11566-8 (print), 3-540-39272-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096006/.

Combet:1982:IED

[Com82e] Edmond Combet. Intégrales Exponentielles: Développements Asymptotiques, Propriétés Lagrangiennes, volume 937 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11566-8 (print), 3-540-39272-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 114 pp. LCCN QA3 .L28 no. 937. URL http:/

/link.springer.com/book/10.1007/BFb0096004;   http://
www.springerlink.com/content/978-3-540-39272-9.

Combet:1982:PNC

[Com82f] Edmond Combet. Phases naturelles en calcul des variations. (French) []. Lecture Notes in Mathematics, 937:89–111, 1982. CODEN LNMAA2. ISBN 3-540-11566-8 (print), 3-540-39272-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096011/.

Combet:1982:PNE

[Com82g] Edmond Combet. Préliminaires: Notes sur les exposants critiques. (French) []. Lecture Notes in Mathematics, 937:52–57, 1982. CODEN LNMAA2. ISBN 3-540-11566-8 (print), 3-540-39272-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096008/.

Combet:1982:VLA

[Com82h] Edmond Combet. Variétés lagrangiennes associées àux systèmes hamiltoniens. (French) []. Lecture Notes in Mathematics, 937: 76–88, 1982. CODEN LNMAA2. ISBN 3-540-11566-8 (print), 3-540-39272-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096010/.

Comer:1983:EPP

[Com83] Stephen D. Comer. Extension of polygroups by polygroups and their representations using color schemes. Lecture Notes in Mathematics, 1004:91–103, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063431/.

Conley:1980:QSP

[Con80a] C. Conley. A qualitative singular perturbation theorem. Lecture Notes in Mathematics, 819:65–89, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086980/.

Conlon:1980:TCS

[Con80b] Joseph G. Conlon. On a theorem of Conley and Smoller. Lecture Notes in Mathematics, 819:90–95, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086981/.

Constantinescu:1980:DMT

[Con80c] Corneliu Constantinescu. Duality in Measure Theory, volume 796 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09989-1 (print), 3-540-39275-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 197 pp. LCCN QA3 .L28 no. 796. URL http:

//link.springer.com/book/10.1007/BFb0089049;   http:/
/www.springerlink.com/content/978-3-540-39275-0.

Constantinescu:1980:DSM

[Con80d] Corneliu Constantinescu. Duals of spaces of measures. Lecture Notes in Mathematics, 796:45–113, 1980. CODEN LN-MAA2. ISBN 3-540-09989-1 (print), 3-540-39275-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089053/.

Constantinescu:1980:I

[Con80e] Corneliu Constantinescu. Introduction. Lecture Notes in Mathematics, 796:1–4, 1980. CODEN LNMAA2. ISBN 3-540-09989-1 (print), 3-540-39275-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089050/.

Constantinescu:1980:P

[Con80f] Corneliu Constantinescu. Preliminaries. Lecture Notes in Mathematics, 796:5–23, 1980. CODEN LNMAA2. ISBN 3-540-09989-1 (print), 3-540-39275-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089051/.

Constantinescu:1980:R

[Con80g] Corneliu Constantinescu. Representations. Lecture Notes in Mathematics, 796:24–44, 1980. CODEN LNMAA2. ISBN 3-540-09989-1 (print), 3-540-39275-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089052/.

Constantinescu:1980:VM

[Con80h] Corneliu Constantinescu. Vector measures. Lecture Notes in Mathematics, 796:114–190, 1980. CODEN LNMAA2. ISBN 3-540-09989-1 (print), 3-540-39275-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089054/.

Connes:1981:FAO

[Con81a] Alain Connes. Feuilletages et algèbres d’opérateurs. (French) []. Lecture Notes in Mathematics, 842:139–155, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089932/.

Conrad:1981:RCL

[Con81b] Paul Conrad. K-radical classes of lattice ordered groups. Lecture Notes in Mathematics, 848:186–207, 1981. CODEN LN-MAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090566/.

Conte:1982:TWH

[Con82a] A. Conte. On threefolds whose hyperplane sections are Enriques surfaces. Lecture Notes in Mathematics, 947:221–228, 1982. CO-DEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093591/.

Conte:1982:ATP

[Con82b] Alberto Conte, editor. Algebraic Threefolds: Proceedings of the 2nd1981 Session of the Centro Internazionale Matematico Estivo (C.I.M.E.), Held at Varenna, Italy, June 15–23, 1981, volume 947 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 947. URL http:/

/link.springer.com/book/10.1007/BFb0093583;   http://
www.springerlink.com/content/978-3-540-39342-9.

Conte:1983:TEA

[Con83] Alberto Conte. Two examples of algebraic threefolds whose hyperplane sections are Enriques surfaces. Lecture Notes in Mathematics, 997:124–130, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061639/.

Constantinescu:1984:ASC

[Con84] Adrian Constantinescu. On the algebraization of some complex schemes. Lecture Notes in Mathematics, 1056:112–131, 1984. CO-DEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071771/.

Cooke:1984:SNA

[Coo84] K. L. Cooke. Stability of non-autonomous delay differential equations by Liapunov functionals. Lecture Notes in Mathematics, 1076:41–52, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072764/.

Cordoba:1980:MFP

[Cór80a] A. Córdoba. Maximal functions: a problem of A. Zygmund. Lecture Notes in Mathematics, 779:154–161, 1980. CODEN LN-MAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087671/.

Cordoba:1980:M

[Cór80b] A. Córdoba. Multipliers of F (L^{P}). Lecture Notes in Mathematics, 779:162–177, 1980. CODEN LNMAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087672/.

Corones:1980:SHC

[Cor80c] James Corones. Some heuristic comments on solitons, integrability conditions and Lie groups. Lecture Notes in Mathematics, 810:23–42, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089972/.

Coram:1981:AFG

[Cor81a] Donald S. Coram. Approximate fibrations — a geometric perspective. Lecture Notes in Mathematics, 870:37–47, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089706/.

Cordellier:1981:GIR

[Cor81b] Florent Cordellier. Sur une généralisation de l’interpolation rationnelle. (French) []. Lecture Notes in Mathematics, 888:124– 136, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095581/.

Corrigan:1982:MS

[Cor82] E. F. Corrigan. Monopole solitons. Lecture Notes in Mathematics, 926:160–178, 1982. CODEN LNMAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092661/.

Cornea:1983:RDB

[Cor83a] Florin Ovidiu Cornea. On the Riesz decomposition of bi-excessive measures on locally compact groups. Lecture Notes in Mathematics, 1014:188–196, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072078/.

Corner:1983:EVDa

[Cor83b] A. L. S. Corner. On the existence of very decomposable Abelian groups. Lecture Notes in Mathematics, 1006:354–357, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21560-9_18/.

Corner:1983:EVDb

[Cor83c] A. L. S. Corner. On the existence of very decomposable abelian groups. Lecture Notes in Mathematics, 1006:354–357, 1983. CO-DEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103713/.

Cornish:1983:CPG

[Cor83d] William H. Cornish. A combinatorial problem and the generalized cosh. Lecture Notes in Mathematics, 1036:147–153, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071516/.

Cordellier:1984:UIH

[Cor84a] Florent Cordellier. Utilisation de l’invariance homographique dans les algorithmes de Losange. (French) []. Lecture Notes in Mathematics, 1071:62–94, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099610/.

Cornea:1984:CRB

[Cor84b] Aurel Cornea. Continuity of reduites and balayaged functions. Lecture Notes in Mathematics, 1096:173–182, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100111/.

Corwin:1984:MCN

[Cor84c] Lawrence Corwin. Matrix coefficients of nilpotent Lie groups. Lecture Notes in Mathematics, 1077:1–11, 1984. CODEN LN-MAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072334/.

Corwin:1984:PPN

[Cor84d] Lawrence Corwin. Primary projections on nilmanifolds. Lecture Notes in Mathematics, 1077:12–22, 1984. CODEN LN-MAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072335/.

Corwin:1984:SLI

[Cor84e] Lawrence Corwin. Solvability of left invariant differential operators on nilpotent Lie groups. Lecture Notes in Mathematics, 1077: 23–59, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072336/.

Costa:1982:SNP

[Cos82a] Douglas L. Costa. Semi-normality and projective modules. Lecture Notes in Mathematics, 924:400–412, 1982. CODEN LN-MAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092943/.

Coste:1982:ESA

[Cos82b] Michel Coste. Ensembles semi-algébriques. Lecture Notes in Mathematics, 959:109–138, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062252/.

Cossart:1984:DD

[Cos84] Vincent Cossart. Desingularization in dimension 2. Lecture Notes in Mathematics, 1101:79–98, 1984. CODEN LNMAA2. ISBN 3-540-13904-4 (print), 3-540-39125-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072261/.

Couchot:1981:MAL

[Cou81] François Couchot. Les modules artiniens et leurs enveloppes quasi-injectives. (French) []. Lecture Notes in Mathematics, 867: 380–395, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090394/.

Cowling:1983:PBS

[Cow83] Michael G. Cowling. Pointwise behavpour of solutions to Schrödinger equations. Lecture Notes in Mathematics, 992:83– 90, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069152/.

Cox:1982:PSA

[Cox82] M. G. Cox. Practical spline approximation. Lecture Notes in Mathematics, 965:79–112, 1982. CODEN LNMAA2. ISBN 3-540-11967-1 (print), 3-540-39558-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063201/.

Colombeau:1981:FBT

[CP81] J.-F. Colombeau and B. Perrot. The Fourier–Borel transform in infinitely many dimensions and applications. Lecture Notes in Mathematics, 843:163–186, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089273/.

Cullinan:1982:DSD

[CP82] M. P. Cullinan and M. J. D. Powell. Data smoothing by divided differences. Lecture Notes in Mathematics, 912:26–37, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093146/.

Caccetta:1983:CCN

[CP83a] Louis Caccetta and Norman J. Pullman. On clique covering numbers of cubic graphs. Lecture Notes in Mathematics, 1036:121– 127, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071513/.

Choksi:1983:ABC

[CP83b] J. R. Choksi and V. S. Prasad. Approximation and Baire category theorems in ergodic theory. Lecture Notes in Mathematics, 1033:94–113, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099849/.

Carstens:1984:ACC

[CP84a] Hans-Georg Carstens and Peter Päppinghaus. Abstract construction of counterexamples in recursive graph theory. Lecture Notes in Mathematics, 1104:39–62, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099478/.

Cwikel:1984:ISA

[CP84b] Michael Cwikel and Jaak Peetre, editors. Interpolation Spaces and Allied Topics in Analysis: Proceedings of the Conference held in Lund, Sweden, August 29–September 1, 1983, volume 1070 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1070. URL http:/

/link.springer.com/book/10.1007/BFb0099087;   http://
www.springerlink.com/content/978-3-540-38913-2.

Colombeau:1981:SDE

[CPD81] J.-F. Colombeau, B. Perrot, and T. A. W. Dwyer III. On the solvability of differential equations of infinite order in non-metrizable spaces. Lecture Notes in Mathematics, 843:187–194, 1981. CO-DEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089274/.

Chen:1981:SHA

[CQ81] C. C. Chen and N. F. Quimpo. On strongly Hamiltonian abelian group graphs. Lecture Notes in Mathematics, 884:23–34, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091805/.

Chen:1983:HCG

[CQ83] C. C. Chen and N. Quimpo. Hamiltonian Cayley graphs of order pq. Lecture Notes in Mathematics, 1036:1–5, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071505/.

Ciarlet:1980:EDM

[CR80a] Philippe G. Ciarlet and Patrick Rabier. Établissement du modèle de von Kármán à partir d’un modèle non linéaire d’élasticité tridimensionnelle. (French) []. Lecture Notes in Mathematics, 826: 1–66, 1980. CODEN LNMAA2. ISBN 3-540-10248-5 (print), 3-540-38336-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091529/.

Ciarlet:1980:EBP

[CR80b] Philippe G. Ciarlet and Patrick Rabier. Étude de la bifurcation perturbée dans les Équations de von Kármán. (French) []. Lecture Notes in Mathematics, 826:120–152, 1980. CODEN LNMAA2. ISBN 3-540-10248-5 (print), 3-540-38336-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091531/.

Ciarlet:1980:EEK

[CR80c] Philippe G. Ciarlet and Patrick Rabier. Étude des équations de von Kármán: Résultats d’existence: Introduction à la bifurcation. (French) []. Lecture Notes in Mathematics, 826:67–119, 1980. CODEN LNMAA2. ISBN 3-540-10248-5 (print), 3-540-38336-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091530/.

Ciarlet:1980:EK

[CR80d] Philippe G. Ciarlet and Patrick Rabier. Les Équations de von Kármán, volume 826 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10248-5 (print), 3-540-38336-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 181 pp. LCCN QA3 .L28 no. 826. URL http:

//link.springer.com/book/10.1007/BFb0091528;   http:/
/www.springerlink.com/content/978-3-540-38336-9.

Coven:1980:PEM

[CR80e] Ethan M. Coven and William L. Reddy. Positively expansive maps of compact manifolds. Lecture Notes in Mathematics, 819: 96–110, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086982/.

Cecil:1981:TSE

[CR81] Thomas E. Cecil and Patrick J. Ryan. Tight spherical embeddings. Lecture Notes in Mathematics, 838:94–104, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088845/.

Coste-Roy:1982:FSS

[CR82] Marie-Françoise Coste-Roy. Faisceau structural sur le spèctre réel et fonctions de Nash. (French) []. Lecture Notes in Mathematics, 959:406–432, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062267/.

Christensen:1983:NDP

[CR83] Jens Peter Reus Christensen and Paul Ressel. Norm-dependent positive definite functions on B-spaces. Lecture Notes in Mathematics, 990:47–53, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064262/.

Campbell:1984:SGO

[CR84a] Colin M. Campbell and Edmund F. Robertson. On the simple groups of order less than 10^{5}. Lecture Notes in Mathematics, 1098:15–20, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099657/.

Crossley:1984:URE

[CR84b] J. N. Crossley and J. B. Remmel. Undecidability and recursive equivalence II. Lecture Notes in Mathematics, 1104:79–100, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099480/.

Craciunas:1983:HQC

[Cră83] Silviu Crăciunaş. Homeomorphismes quasi-conformes sur des espaces localement convexes. (French) []. Lecture Notes in Mathematics, 1013:50–67, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066517/.

Cressie:1984:MS

[Cre84] N. A. C. Cressie. Modelling sets. Lecture Notes in Mathematics, 1091:138–149, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098807/.

Combe:1983:JPA

[CRSSC83] Ph. Combe, R. Rodriguez, M. Sirugue, and M. Sirugue-Collin. Jump processes and applications to the trigonometric interaction. Lecture Notes in Mathematics, 1031:16–30, 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072108/.

Combe:1984:SRT

[CRSSC84] Ph. Combe, R. Rodriguez, M. Sirugue, and M. Sirugue-Collin. Stochastic representation of thermal functionals. Lecture Notes in Mathematics, 1055:86–92, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071712/.

Coifman:1980:CMI

[CRW^{+}80] R. R. Coifman, R. Rochberg, G. Weiss, M. Cwikel, and Y. Sagher. The complex method for interpolation of operators acting on families of Banach spaces. Lecture Notes in Mathematics, 779:123– 153, 1980. CODEN LNMAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087670/.

Cameron:1980:SBA

[CS80a] Robert Horton Cameron and David Arne Storvick. Some Banach algebras of analytic Feynman integrable functionals. Lecture Notes in Mathematics, 798:18–67, 1980. CODEN LN-MAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097256/.

Chicone:1980:IHI

[CS80b] Carmen Chicone and R. C. Swanson. Infinitesimal hyperbolicity implies hyperbolicity. Lecture Notes in Mathematics, 819: 50–64, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086979/.

Choi:1981:CPU

[CS81a] Bong Dae Choi and Louis Sucheston. Continuous parameter uniform amarts. Lecture Notes in Mathematics, 860:85–98, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090609/.

Cooper:1981:UMC

[CS81b] J. B. Cooper and W. Schachermayer. Uniform measures and CoSaks spaces. Lecture Notes in Mathematics, 843:217–246, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089276/.

Carrell:1982:AHA

[CS82a] James B. Carrell and Andrew John Sommese. Almost homogeneous C∗ actions on compact complex surfaces. Lecture Notes in Mathematics, 956:29–33, 1982. CODEN LNMAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101507/.

Carrell:1982:GTH

[CS82b] James B. Carrell and Andrew John Sommese. A generalization of a theorem of Horrocks. Lecture Notes in Mathematics, 956: 23–28, 1982. CODEN LNMAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101506/.

Clausen:1982:PSR

[CS82c] Michael Clausen and Friedrich Stötzer. Pictures and skew (reverse) plane partitions. Lecture Notes in Mathematics, 969:100– 114, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062989/.

Cotlar:1982:MTF

[CS82d] Mischa Cotlar and Cora Sadosky. Majorized Toeplitz forms and weighted inequalities with general norms. Lecture Notes in Mathematics, 908:139–168, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093285/.

Capozzi:1983:NCA

[CS83a] A. Capozzi and A. Salvatore. A note on a class of autonomous Hamiltonian systems with strong resonance at infinity. Lecture Notes in Mathematics, 1017:132–139, 1983. CODEN LN-MAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103244/.

Cotlar:1983:VVI

[CS83b] Mischa Cotlar and Cora Sadosky. Vector valued inequalities of Marcinkiewicz–Zygmund and Grothendieck type for Toeplitz forms. Lecture Notes in Mathematics, 992:278–308, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069164/.

Chui:1984:DRA

[CS84a] Charles K. Chui and Xie-Chang Shen. Degree of rational approximation in digital filter realization. Lecture Notes in Mathematics, 1105:189–209, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072412/.

Connett:1984:RAC

[CS84b] W. C. Connett and A. L. Schwartz. A remark about Calderón’s upper s method of interpolation. Lecture Notes in Mathematics, 1070:48–53, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099092/.

Csorgo:1981:TGF

[Csö81] Miklós Csörgő. On a test for goodness-of-fit based on the empirical probability measure of Foutz and testing for exponentiality. Lecture Notes in Mathematics, 861:25–34, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097309/.

Colliot-Thelene:1982:VDN

[CT82] J.-L. Colliot-Thélène. Variantes du Nullstellensatz réel et anneaux formellement réels. (French) []. Lecture Notes in Mathematics, 959:98–108, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062251/.

Colliot-Thelene:1982:GAR

[CTCMR82] Jean-Louis Colliot-Thélène, Michel Coste, Louis Mahé, and Marie-Françoise Roy, editors. Géométrie Algébrique Réelle et Formes Quadratiques: Journées S.M.F., Université de Rennes 1, Mai 1981, volume 959 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 959. URL http:/

/link.springer.com/book/10.1007/BFb0062247;   http://
www.springerlink.com/content/978-3-540-39548-5.

Cuntz:1983:GHB

[Cun83] Joachim Cuntz. Generalized homomorphisms between C∗- algebras and KK-theory. Lecture Notes in Mathematics, 1031: 31–45, 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072109/.

Curry:1980:AFC

[Cur80a] J. H. Curry. An algorithm for finding closed orbits. Lecture Notes in Mathematics, 819:111–120, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086983/.

Curtis:1980:HRF

[Cur80b] Charles W. Curtis. Homology representations of finite groups. Lecture Notes in Mathematics, 832:177–194, 1980. CODEN LN-MAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088463/.

Cusick:1981:BDA

[Cus81] T. W. Cusick. Best Diophantine approximations for ternary linear forms, II. Lecture Notes in Mathematics, 899:231–238, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096464/.

Cushman:1983:PSA

[Cus83] Richard Cushman. Periodic solutions of the autonomous Josephson equation. Lecture Notes in Mathematics, 985:319–326, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062373/.

Cusick:1984:LBR

[Cus84] T. W. Cusick. Lower bounds for regulators. Lecture Notes in Mathematics, 1068:63–73, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099441/.

Cuyt:1981:NCA

[Cuy81] Annie A. M. Cuyt. Numerical comparison of abstract Padéapproximants and abstract rational approximants with other generalizations of the classical Padé-approximant. Lecture Notes in Mathematics, 888:137–157, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095582/.

Cuyt:1984:APA

[Cuy84a] Annie Cuyt. Abstract Padé approximants in operator theory. Lecture Notes in Mathematics, 1065:1–58, 1984. CODEN LN-MAA2. ISBN 3-540-13342-9 (print), 3-540-38878-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099707/.

Cuyt:1984:MMP

[Cuy84b] Annie Cuyt. The mechanism of the multivariate Padé process. Lecture Notes in Mathematics, 1071:95–103, 1984. CODEN LN-MAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099611/.

Cuyt:1984:MPA

[Cuy84c] Annie Cuyt. Multivariate Padé approximants. Lecture Notes in Mathematics, 1065:59–93, 1984. CODEN LNMAA2. ISBN 3-540-13342-9 (print), 3-540-38878-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099708/.

Cuyt:1984:PAO

[Cuy84d] Annie Cuyt. Padé Approximants for Operators: Theory and Applications, volume 1065 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13342-9 (print), 3-540-38878-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 138 pp. LCCN QA3 .L28 no. 1065.

URL http://link.springer.com/book/10.1007/BFb0099706;
 http://www.springerlink.com/content/978-3-540-38878-
4.

Cuyt:1984:SNO

[Cuy84e] Annie Cuyt. The solution of nonlinear operator equations. Lecture Notes in Mathematics, 1065:94–132, 1984. CODEN LN-MAA2. ISBN 3-540-13342-9 (print), 3-540-38878-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099709/.

Carmona:1981:NCH

[CV81] Jacques Carmona and Michèle Vergne, editors. Non Commutative Harmonic Analysis and Lie Groups: Actes du Colloque d’Analyse Harmonique Non Commutative, 16 au 20 juin 1980 Marseille–Luminy, volume 880 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 880. URL http:

//link.springer.com/book/10.1007/BFb0090400;   http:/
/www.springerlink.com/content/978-3-540-38783-1.

Caenepeel:1982:CPG

[CV82] S. Caenepeel and F. Van Oystaeyen. Crossed products over graded local rings. Lecture Notes in Mathematics, 917:25–42, 1982. CO-DEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092226/.

Carmona:1983:NCH

[CV83] Jaques Carmona and Michèle Vergne, editors. Non Commutative Harmonic Analysis and Lie Groups: Proceedings of the International Conference Held in Marseille Luminy, 21–26 June, 1982, volume 1020 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1020. URL http:/

/link.springer.com/book/10.1007/BFb0071493;   http://
www.springerlink.com/content/978-3-540-38700-8.

Cherlin:1984:CPA

[CV84a] G. L. Cherlin and H. Volger. Convexity properties and algebraic closure operators. Lecture Notes in Mathematics, 1103: 113–146, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099384/.

Chiantini:1984:SPS

[CV84b] Luca Chiantini and Paolo Valabrega. Some properties of subcanonical curves. Lecture Notes in Mathematics, 1092:152–160, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099361/.

Cullum:1980:CEE

[CW80] Jane Cullum and Ralph A. Willoughby. Computing eigenvectors (and eigenvalues) of large, symmetric matrices using Lanczos tridiagonalization. Lecture Notes in Mathematics, 773:46– 63, 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094163/.

Chao:1982:MTH

[CW82] Jia-Arng Chao and Wojbor A. Woyczyński, editors. Martingale Theory in Harmonic Analysis and Banach Spaces: Proceedings of the NSF-CBMS Conference Held at the Cleveland State University, Cleveland, Ohio, July 13–17, 1981, volume 939 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 939.

URL http://link.springer.com/book/10.1007/BFb0096252;
 http://www.springerlink.com/content/978-3-540-39284-

Cecconi:1983:MTO

[CZ83a] Jaurés P. Cecconi and Tullio Zolezzi, editors. Mathematical Theories of Optimization: Proceedings of the International Conference Held in S. Margherita Ligure (Genova) November 30– December 4, 1981, volume 979 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 979. URL http:

//link.springer.com/book/10.1007/BFb0066246;   http:/
/www.springerlink.com/content/978-3-540-39473-0.

Conley:1983:ITP

[CZ83b] C. Conley and E. Zehnder. An index theory for periodic solutions of a Hamiltonian system. Lecture Notes in Mathematics, 1007:132–145, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061415/.

Czedli:1983:CCS

[Czé83] Gábor Czédli. A characterization for congruence semi-distributivity. Lecture Notes in Mathematics, 1004:104–110, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063432/.

Czelakowski:1984:RFB

[Cze84] Janusz Czelakowski. Remarks on finitely based logics. Lecture Notes in Mathematics, 1103:147–168, 1984. CODEN LN-MAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099385/.

Czyz:1981:GBT

[Czy81] J. Czyż. On graded bundles and their geometry. Lecture Notes in Mathematics, 838:110–125, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088847/.

DaPrato:1984:ADE

[Da 84] G. Da Prato. Abstract differential equations and extrapolation spaces. Lecture Notes in Mathematics, 1076:53–61, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072765/.

DAntoni:1984:SIL

[D’A84] Claudio D’Antoni. Statistical independence of local algebras. Lecture Notes in Mathematics, 1055:93–102, 1984. CODEN LN-MAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071713/.

Dacorogna:1982:A

[Dac82a] Bernard Dacorogna. Applications. Lecture Notes in Mathematics, 922:59–73, 1982. CODEN LNMAA2. ISBN 3-540-11488-2 (print), 3-540-39171-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096147/.

Dacorogna:1982:CC

[Dac82b] Bernard Dacorogna. Compensated compactness. Lecture Notes in Mathematics, 922:7–58, 1982. CODEN LNMAA2. ISBN 3-540-11488-2 (print), 3-540-39171-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096146/.

Dacorogna:1982:DRP

[Dac82c] Bernard Dacorogna. Dual and relaxed problems. Lecture Notes in Mathematics, 922:74–99, 1982. CODEN LNMAA2. ISBN 3-540-11488-2 (print), 3-540-39171-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096148/.

Dacorogna:1982:I

[Dac82d] Bernard Dacorogna. Introduction. Lecture Notes in Mathematics, 922:1–6, 1982. CODEN LNMAA2. ISBN 3-540-11488-2 (print), 3-540-39171-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096145/.

Dacorogna:1982:WCW

[Dac82e] Bernard Dacorogna. Weak Continuity and Weak Lower Semicontinuity of Non-Linear Functionals, volume 922 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CO-DEN LNMAA2. ISBN 3-540-11488-2 (print), 3-540-39171-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 120 pp. LCCN QA3 .L28 no. 922. URL http:/

/link.springer.com/book/10.1007/BFb0096144;   http://
www.springerlink.com/content/978-3-540-39171-5. See er-

ratum in [LM88].

Dade:1980:ARM

[Dad80] Everett C. Dade. Algebraically rigid modules. Lecture Notes in Mathematics, 832:195–215, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088464/.

Daffer:1981:SSW

[Daf81] Peter Z. Daffer. Some strong and weak laws of large numbers for weighted sums in D[0, 1]. Lecture Notes in Mathematics, 860: 99–106, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090610/.

Dafermos:1983:SED

[Daf83] C. M. Dafermos. Stabilizing effects of dissipation. Lecture Notes in Mathematics, 1017:140–147, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103245/.

Dahlquist:1982:CGE

[Dah82] Germund Dahlquist. On the control of the global error in stiff initial value problems. Lecture Notes in Mathematics, 912:38– 49, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093147/.

deAcosta:1983:ILI

[dAKL83] Alejandro de Acosta, James Kuelbs, and Michel Ledoux. An inequality for the law of the iterated logarithm. Lecture Notes in Mathematics, 990:1–29, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064260/.

Dales:1983:CAR

[Dal83] H. G. Dales. Convolution algebras on the real line. Lecture Notes in Mathematics, 975:180–209, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064550/.

Doebner:1982:DGM

[DAP82] Heinz-Dietrich Doebner, Stig I. Andersson, and Herbert Rainer Petry, editors. Differential Geometric Methods in Mathematical Physics: Clausthal 1980 Proceedings of an International Conference Held at the Technical University of Clausthal, FRG, July 23–25, 1980, volume 905 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 905. URL http:

//link.springer.com/book/10.1007/BFb0092421;   http:/
/www.springerlink.com/content/978-3-540-39002-2.

Darling:1982:MMD

[Dar82] R. W. R. Darling. Martingales in manifolds — definition examples, and behaviour under maps. Lecture Notes in Mathematics, 921:217–236, 1982. CODEN LNMAA2. ISBN 3-540-11486-6 (print), 3-540-39167-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092652/.

Das:1981:CPS

[Das81] Prabir Das. Characterization of potentially self-complementary, self-converse degree-pair sequences for digraphs. Lecture Notes in Mathematics, 885:212–226, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092265/.

Dateyama:1983:IMH

[Dat83] Masahito Dateyama. Invariant measures for homeomorphisms with almost weak specification. Lecture Notes in Mathematics, 1021:93–96, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072906/.

Dauns:1981:GSR

[Dau81] John Dauns. Generalized semigroup rings. Lecture Notes in Mathematics, 848:235–254, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090570/.

Dauns:1982:SUM

[Dau82] John Dauns. Sums of uniform modules. Lecture Notes in Mathematics, 951:68–87, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067325/.

Davenport:1981:EMC

[Dav81a] James H. Davenport. Effective mathematics — the computer algebra viewpoint. Lecture Notes in Mathematics, 873:31–43, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090725/.

Davies:1981:TDA

[Dav81b] Laurie Davies. A theorem of Deny with applications to characterization problems. Lecture Notes in Mathematics, 861:35–41, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097310/.

Davis:1984:CGA

[Dav84] Michael W. Davis. Coxeter groups and aspherical manifolds. Lecture Notes in Mathematics, 1051:197–221, 1984. CODEN LN-MAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075568/.

Dawson:1981:CFS

[Daw81] J. E. Dawson. A construction for a family of sets and its application to matroids. Lecture Notes in Mathematics, 884:136– 147, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091815/.

Dawson:1982:MBO

[Daw82] Jeremy E. Dawson. Matroid bases, opposite families and some related algorithms. Lecture Notes in Mathematics, 952:225– 238, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061981/.

Dawson:1984:CSR

[Daw84] Jeremy E. Dawson. A collection of sets related to the Tutte polynomial of a matroid. Lecture Notes in Mathematics, 1073: 193–204, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073117/.

Dayton:1981:SIC

[Day81] Barry H. Dayton. Seminormality implies the Chinese remainder theorem. Lecture Notes in Mathematics, 854:124–126, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089518/.

Day:1983:GAM

[Day83] Alan Day. Geometrical applications in modular lattices. Lecture Notes in Mathematics, 1004:111–141, 1983. CODEN LN-MAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063433/.

deBoor:1982:TMA

[dB82] C. de Boor. Topics in multivariate approximation theory. Lecture Notes in Mathematics, 965:39–78, 1982. CODEN LN-MAA2. ISBN 3-540-11967-1 (print), 3-540-39558-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063200/.

deBartolomeis:1983:ACC

[dB83] Paolo de Bartolomeis. Sur l’analyticité complexe de certaines applications harmoniques. (French) []. Lecture Notes in Mathematics, 1028:27–40, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071676/.

deBoor:1984:AOS

[dB84a] C. de Boor. Approximation order from smooth bivariate pp functions. Lecture Notes in Mathematics, 1066:44–49, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099517/.

deBruin:1984:SCR

[dB84b] Marcel G. de Bruin. Some convergence results in simultaneous rational approximation to the set of hypergeometric functions {_{1}F_{1}(1; c_{i}; z)}_{i=1}^{n}. Lecture Notes in Mathematics, 1071:12– 33, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099607/.

Castelan:1980:LFM

[dBC80] Walter de Bona Castelan. A Liapunov functional for a matrix retarded difference-differential equation with several delay. Lecture Notes in Mathematics, 799:82–118, 1980. CODEN LN-MAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089310/.

deBartolomeis:1980:TPF

[dBT80] Paolo de Bartolomeis and Giuseppe Tomassini. Traces of pluriharmonic functions. Lecture Notes in Mathematics, 798:10–17, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097255/.

deBruin:1981:PAA

[dBvR81] M. G. de Bruin and H. van Rossum, editors. Padé Approximation and its Applications Amsterdam 1980: Proceedings of a Conference Held in Amsterdam, The Netherlands, October 29–31, 1980, volume 888 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 888. URL http:/

/link.springer.com/book/10.1007/BFb0095573;   http://
www.springerlink.com/content/978-3-540-38606-3.

doCarmo:1981:SMS

[dC81] Manfredo P. do Carmo. Stability of minimal submanifolds. Lecture Notes in Mathematics, 838:129–139, 1981. CODEN LN-MAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088849/.

deCarvalho:1983:SRA

[dCR83] Sonia de Carvalho and Robert Roussarie. Some remarks about homoclinic points of second order differential equations. Lecture Notes in Mathematics, 1007:88–95, 1983. CODEN LN-MAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061411/.

DeGiorgi:1983:CCV

[DD83] Ennio De Giorgi and Gianni Dal Maso. Γ-convergence and calculus of variations. Lecture Notes in Mathematics, 979:121– 143, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066252/.

deLaPradelle:1982:RIF

[de 82a] Arnaud de La Pradelle. Sur la représentation intégrale des fonctions pseudo-excessives. (French) []. Lecture Notes in Mathematics, 906:321–328, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093275/.

DelaTorre:1982:WET

[De 82b] Alberto De la Torre. Weights in ergodic theory. Lecture Notes in Mathematics, 908:128–138, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093284/.

DeLauney:1983:GHM

[De 83] Warwick De Launey. Generalised Hadamard matrices whose rows and columns form a group. Lecture Notes in Mathematics, 1036: 154–176, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071517/.

Deheuvels:1980:NPT

[Deh80] Paul Deheuvels. Non parametric tests of independence. Lecture Notes in Mathematics, 821:95–107, 1980. CODEN LN-MAA2. ISBN 3-540-10239-6 (print), 3-540-38318-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097426/.

Deheuvels:1981:MTI

[Deh81] Paul Deheuvels. Multivariate tests of independence. Lecture Notes in Mathematics, 861:42–50, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097311/.

Deligne:1980:SGC

[Del80a] P. Deligne. Sommes de Gauss cubiques et revêtements de SL(2), d’après S. J. Patterson. (French) []. Lecture Notes in Mathematics, 770:244–277, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096246/.

Delon:1980:HFE

[Del80b] Françoise Delon. Hensel fields in equal characteristic p > 0. Lecture Notes in Mathematics, 834:108–116, 1980. CODEN LN-MAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090162/.

Delahaye:1981:CAE

[Del81a] J.-P. Delahaye. Choix automatique entre suites de paramètres dans l’extrapolation de Richardson. (French) []. Lecture Notes in Mathematics, 888:158–172, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095583/.

Deligne:1981:HCAa

[Del81b] P. Deligne. Hodge cycles on Abelian varieties. Lecture Notes in Mathematics, 900:9–100, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38955-2_3/.

Deligne:1981:HCAb

[Del81c] P. Deligne. Hodge cycles on Abelian varieties. Lecture Notes in Mathematics, 900:418, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38955-2_12.

Deligne:1981:MGTa

[Del81d] P. Deligne. Motifs et groupes de Taniyama. (French) []. Lecture Notes in Mathematics, 900:261–279, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38955-2_6/.

Deligne:1981:MGTb

[Del81e] P. Deligne. Motifs et groupes de Taniyama. (French) []. Lecture Notes in Mathematics, 900:421, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38955-2_15.

Deligne:1981:GFD

[Del81f] Pierre Deligne. Le groupe fondamental du complément d’une courbe plane n’ayant que des points doubles ordinaires est abélien [d’après W. Fulton]. (French) []. Lecture Notes in Mathematics, 842:1–10, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089924/.

Dellacherie:1981:CAS

[Del81g] Claude Dellacherie. Capacities and analytic sets. Lecture Notes in Mathematics, 839:1–31, 1981. CODEN LNMAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090234/.

Delorme:1981:HCH

[Del81h] Patrick Delorme. Harish-Chandra homomorphisms and minimal K-types of real semisimple Lie groups. Lecture Notes in Mathematics, 880:69–73, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090404/.

Dellacherie:1982:AEP

[Del82a] C. Dellacherie. Appendice à l’exposé précedent. (French) []. Lecture Notes in Mathematics, 920:29–40, 1982. CODEN LN-MAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092768/.

Delves:1982:CMI

[Del82b] L. M. Delves. Chebyshev methods for integral and differential equations. Lecture Notes in Mathematics, 912:50–63, 1982. CO-DEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093148/.

Delahaye:1984:OFS

[Del84a] J.-P. Delahaye. Opérations sur des familles de suites et accélérabilité. (French) []. Lecture Notes in Mathematics, 1071: 104–116, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099612/.

Dellacherie:1984:SNE

[Del84b] Claude Dellacherie. Les sous-noyaux élémentaires. (French) []. Lecture Notes in Mathematics, 1096:183–222, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100112/.

Demazure:1980:BCD

[Dem80a] M. Demazure. A, b, c, d, e, f, etc. Lecture Notes in Mathematics, 777:221–227, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085885/.

Demazure:1980:SDPa

[Dem80b] M. Demazure. Surfaces de Del Pezzo — I. (French) []. Lecture Notes in Mathematics, 777:21–22, 1980. CODEN LN-MAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085875.

Demazure:1980:SDPb

[Dem80c] M. Demazure. Surfaces de Del Pezzo — II — Éclater n points dans P^{2}. (French) [Del Pezzo surfaces — II — Exploding n points in P^{2}]. Lecture Notes in Mathematics, 777:23–35, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085876/.

Demazure:1980:SDPc

[Dem80d] M. Demazure. Surfaces de Del Pezzo — III — positions presque générales. (French) []. Lecture Notes in Mathematics, 777:36– 49, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085877/.

Demazure:1980:SDPd

[Dem80e] M. Demazure. Surfaces de Del Pezzo — IV — systèmes anticanoniques. (French) []. Lecture Notes in Mathematics, 777:50–60, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085878/.

Demazure:1980:SDPe

[Dem80f] M. Demazure. Surfaces de Del Pezzo — V — modèles anticanoniques. (French) []. Lecture Notes in Mathematics, 777:61–69, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085879/.

Demazure:1981:CEP

[Dem81] Michel Demazure. Caractérisations de l’espace projectif (conjectures de Hartshorne et de Frankel) d’après Shigefumi Mori [6]. (French) []. Lecture Notes in Mathematics, 842:11–19, 1981. CO-DEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089925/.

Demailly:1982:RED

[Dem82a] J.-P. Demailly. Relations entre les différentes notions de fibrés et de courants positifs. (French) []. Lecture Notes in Mathematics, 919:56–76, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097043/.

Demailly:1982:SHM

[Dem82b] J.-P. Demailly. Scindage holomorphe d’un morphisme de fibrés vectoriels semi-positifs avec estimations L^{2}. (French) [Holomorphic splitting of a morphism of semi-positive vector sheaves with L^{2} estimates]. Lecture Notes in Mathematics, 919:77–107, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097044/.

Demailly:1983:CFS

[Dem83a] Jean-Pierre Demailly. Constructibilité des faisceaux de solutions des systèmes différentiels holonômes. (French) []. Lecture Notes in Mathematics, 1028:83–95, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071679/.

Demmel:1983:CNE

[Dem83b] James Demmel. The condition number of equivalence transformations that block diagonalize matrix pencils. Lecture Notes in Mathematics, 973:2–16, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062091/.

Demailly:1984:PSC

[Dem84] Jean-Pierre Demailly. Sur la propagation des singularités des courants positifs fermés. (French) []. Lecture Notes in Mathematics, 1094:53–64, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099155/.

Dennis:1982:AEPa

[Den82a] R. Keith Dennis, editor. Algebraic K-Theory: Proceedings of a Conference Held at Oberwolfach, June 1980 Part I, volume 966 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 966. URL http:/

/link.springer.com/book/10.1007/BFb0062163;   http://
www.springerlink.com/content/978-3-540-39553-9.

Dennis:1982:AEPb

[Den82b] R. Keith Dennis, editor. Algebraic K-Theory: Proceedings of a Conference Held at Oberwolfach, June 1980 Part II, volume 967 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 967. URL http:/

/link.springer.com/book/10.1007/BFb0061896;   http://
www.springerlink.com/content/978-3-540-39556-0.

Derdzinski:1981:CRM

[Der81] Andrzej Derdziński. Compact Riemannian manifolds with harmonic curvature and non-parallel Ricci tensor. Lecture Notes in Mathematics, 838:126–128, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088848/.

Derriennic:1984:CSR

[Der84] Y. Derriennic. Une condition suffisante de recurrence pour des chaines de Markov sur la droite. (French) []. Lecture Notes in Mathematics, 1064:49–55, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073633/.

Deser:1980:H

[Des80a] S. Deser. Hypergravities. Lecture Notes in Mathematics, 836: 256–262, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089740/.

Deser:1980:WDS

[Des80b] S. Deser. What does supergravity teach us about gravity? Lecture Notes in Mathematics, 775:49–68, 1980. CODEN LN-MAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092020/.

Deser:1982:MLD

[Des82] S. Deser. Massless limits and dual field theories. Lecture Notes in Mathematics, 905:3–11, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092423/.

Deshouillers:1984:QMC

[Des84] Jean-Marc Deshouillers. Sur quelques moyennes des coéfficients de Fourier de formes modulaires. (French) []. Lecture Notes in Mathematics, 1068:74–79, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099442/.

Dahlquist:1982:NMS

[DESS82] Germund Dahlquist, Lennart Edsberg, Gunilla Sköllermo, and Gustaf Söderlind. Are the numerical methods and software satisfactory for chemical kinetics? Lecture Notes in Mathematics, 968: 149–164, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064886/.

Dettweiler:1982:PMB

[Det82] Egbert Dettweiler. Poisson measures on Banach lattices. Lecture Notes in Mathematics, 928:16–45, 1982. CODEN LN-MAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093217/.

Dettweiler:1983:BSV

[Det83] E. Dettweiler. Banach space valued processes with independent increments and stochastic integration. Lecture Notes in Mathematics, 990:54–83, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064263/.

Dettweiler:1984:SIE

[Det84] E. Dettweiler. Stochastic integral equations and diffusions on Banach spaces. Lecture Notes in Mathematics, 1080:9–45, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099783/.

Devaney:1980:LTM

[Dev80] Robert L. Devaney. Linked twist mappings are almost Anosov. Lecture Notes in Mathematics, 819:121–145, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086984/.

Devlin:1981:MLC

[Dev81] Keith J. Devlin. Morass-like constructions of ℵ_{2}-trees in L. Lecture Notes in Mathematics, 872:1–36, 1981. CODEN LN-MAA2. ISBN 3-540-10849-1 (print), 3-540-38757-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098617/.

DeVore:1984:FHB

[DeV84] R. DeVore. The K functional for (h_{1}, bmo). Lecture Notes in Mathematics, 1070:66–79, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099094/.

deFigueiredo:1982:PSS

[dF82a] Djairo Guedes de Figueiredo. Positive solutions of semilinear elliptic problems. Lecture Notes in Mathematics, 957:34–87, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066233/.

deFrancia:1982:WNI

[dF82b] José L. Rubio de Francia. Weighted norm inequalities and vector valued inequalities. Lecture Notes in Mathematics, 908:86– 101, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093282/.

DeMeyer:1982:BGS

[DF82c] F. R. DeMeyer and T. J. Ford. On the Brauer group of surfaces and subrings of k[x, y]. Lecture Notes in Mathematics, 917: 211–221, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092236/.

deFalco:1984:PNC

[dF84] Diego de Falco. On the problem of non configurational observables in stochastic mechanics. Lecture Notes in Mathematics, 1055:103–112, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071714/.

deFigueiredo:1982:DEP

[dFH82] Djairo Guedes de Figueiredo and Chaim Samuel Hönig, editors. Differential Equations: Proceedings of the 1st Latin American School of Differential Equations, Held at São Paulo, Brazil, June 29–July 17, 1981, volume 957 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 957. URL http:

//link.springer.com/book/10.1007/BFb0066230;   http:/
/www.springerlink.com/content/978-3-540-39539-3.

Dellacherie:1982:ICF

[DFM82] C. Dellacherie, D. Feyel, and G. Mokobodzki. Intégrales de capacités fortement sous-additives. (French) []. Lecture Notes in Mathematics, 920:8–28, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092767/.

DeCaen:1980:PBM

[DG80a] D. De Caen and D. A. Gregory. Prime Boolean matrices. Lecture Notes in Mathematics, 829:76–82, 1980. CODEN LN-MAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088902/.

Debiard:1980:MCO

[DG80b] Amédée Debiard and Bernard Gaveau. Méthodes de contrôle optimal en analyse complexe. (French) []. Lecture Notes in Mathematics, 798:109–140, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097258/.

Dlab:1980:RTP

[DG80c] Vlastimil Dlab and Peter Gabriel, editors. Representation Theory I: Proceedings of the Workshop on the Present Trends in Representation Theory, Ottawa, Carleton University, August 13–18, 1979, volume 831 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10263-9 (print), 3-540-38385-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 831. URL http:/

/link.springer.com/book/10.1007/BFb0089777;   http://
www.springerlink.com/content/978-3-540-38385-7.

Dlab:1980:RTI

[DG80d] Vlastimil Dlab and Peter Gabriel, editors. Representation Theory II: Proceedings of the Second International Conference on Representations of Algebras Ottawa, Carleton University, August 13–25, 1979, volume 832 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 832. URL http:

//link.springer.com/book/10.1007/BFb0088456;   http:/
/www.springerlink.com/content/978-3-540-38387-1.

Demailly:1983:MSC

[DG83] Jean-Pierre Demailly and Bernard Gaveau. Majoration statistique de la courbure d’une variété analytique. (French) []. Lecture Notes in Mathematics, 1028:96–124, 1983. CODEN LN-MAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071680/.

deGiorgi:1984:DCM

[dG84a] Ennio de Giorgi. On a definition of Γ-convergence of measures. Lecture Notes in Mathematics, 1091:150–159, 1984. CODEN LN-MAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098808/.

DelCentina:1984:CRE

[DG84b] Andrea Del Centina and Alessandro Gimigliano. Curves on rational and elliptic normal cones which are set theoretically complete intersection. Lecture Notes in Mathematics, 1092:241–252, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099367/.

Drisch:1984:SLH

[DG84c] Thomas Drisch and Léonard Gallardo. Stable laws on the Heisenberg groups. Lecture Notes in Mathematics, 1064:56–79, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073634/.

deGuzman:1982:GSP

[dGdlV82] Miguel de Guzmán and Agustín de la Villa. A general system of polar coordinates with applications. Lecture Notes in Mathematics, 945:226–233, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096678/.

Desch:1984:WPA

[DGZ84] G. W. Desch, R. Grimmer, and M. Zeman. Wave propagation for abstract integrodifferential equations. Lecture Notes in Mathematics, 1076:62–70, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072766/.

Diekmann:1980:HMJ

[DH80] O. Diekmann and D. Hilhorst. How many jumps? variational characterization of the limit solution of a singular perturbation problem. Lecture Notes in Mathematics, 810:159–180, 1980. CO-DEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089980/.

Deconinck:1982:MFE

[DH82] Herman Deconinck and Charles Hirsch. A multigrid finite element method for the transonic potential equation. Lecture Notes in Mathematics, 960:387–409, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069935/.

Dragan:1983:UAE

[DH83] V. Drăgan and A. Halanay. Uniform asymptotic expresions for the fundamental matrix of singularly perturbed linear systems and applications. Lecture Notes in Mathematics, 985:215–245, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062369/.

Djrbashyan:1984:UMN

[DHH^{+}84] M. M. Djrbashyan, V. P. Havin, S. V. Hruščëv, B. Jöricke, N. G. Makarov, E. M. Dyn’kin, V. I. Matsaev, L. de Branges, J. A. Siddiqi, A. L. Vol’berg, and V. V. Napalkov. Uniqueness, moments, normality. Lecture Notes in Mathematics, 1043:520–562, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072192/.

deHaan:1981:LLT

[dHR81] L. de Haan and S. I. Resnick. Local limit theorem for sample extremes. Lecture Notes in Mathematics, 861:51–53, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097312/.

Deligne:1981:COC

[DI81a] P. Deligne and L. Illusie. Cristaux ordinaires et coordonnées canoniques. (French) []. Lecture Notes in Mathematics, 868:80– 137, 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090647/.

Deligne:1981:RSC

[DI81b] P. Deligne and Luc Illusie. Relèvement des surfaces K_{3} en caractéristique nulle. (French) []. Lecture Notes in Mathematics, 868: 58–79, 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090646/.

Dennis:1982:HHS

[DI82] R. Keith Dennis and Kiyoshi Igusa. Hochschild homology and the second obstruction for pseudoisotopy. Lecture Notes in Mathematics, 966:7–58, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062165/.

DePascale:1983:PSG

[DI83] E. De Pascale and R. Iannacci. Periodic solutions of generalized Liénard equations with delay. Lecture Notes in Mathematics, 1017:148–156, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103246/.

Diamond:1984:AGF

[Dia84] Jack Diamond. p-adic gamma functions and their applications. Lecture Notes in Mathematics, 1052:168–175, 1984. CODEN LN-MAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071542/.

Dickmann:1980:PRC

[Dic80a] M. A. Dickmann. On polynomials over real closed rings. Lecture Notes in Mathematics, 834:117–135, 1980. CODEN LN-MAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090163/.

Dicks:1980:CDO

[Dic80b] Warren Dicks. Cohomological dimension one. Lecture Notes in Mathematics, 790:101–120, 1980. CODEN LNMAA2. ISBN 3-540-09974-3 (print), 3-540-39210-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088146/.

Dicks:1980:D

[Dic80c] Warren Dicks. Decompositions. Lecture Notes in Mathematics, 790:55–100, 1980. CODEN LNMAA2. ISBN 3-540-09974-3 (print), 3-540-39210-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088145/.

Dicks:1980:FG

[Dic80d] Warren Dicks. Fundamental groups. Lecture Notes in Mathematics, 790:35–53, 1980. CODEN LNMAA2. ISBN 3-540-09974-3 (print), 3-540-39210-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088144/.

Dicks:1980:GAG

[Dic80e] Warren Dicks. Groups acting on graphs. Lecture Notes in Mathematics, 790:1–34, 1980. CODEN LNMAA2. ISBN 3-540-09974-3 (print), 3-540-39210-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088143/.

Dicks:1980:GTP

[Dic80f] Warren Dicks. Groups, Trees and Projective Modules, volume 790 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09974-3 (print), 3-540-39210-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 127 pp. LCCN QA3 .L28 no. 790. URL http:

//link.springer.com/book/10.1007/BFb0088140;   http:/
/www.springerlink.com/content/978-3-540-39210-1.

Diers:1982:CRP

[Die82a] Yves Diers. Un critère de représentabilité par sections continues de faisceaux. (French) []. Lecture Notes in Mathematics, 962: 51–61, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066885/.

Diestler:1982:DCI

[Die82b] Dennis J. Diestler. The discretization of continuous infinite sets of coupled ordinary linear differential equations: Application to the collision-induced dissociation of a diatomic molecule by an atom. Lecture Notes in Mathematics, 968:40–52, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064880/.

Dies:1983:CMP

[Die83a] Jacques-Édouard Dies. Chaînes de Markov sur les permutations. (French) [Markov chains on permutations], volume 1010 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 226 pp. LCCN QA3 .L28 no. 1010. URL http:/

/link.springer.com/book/10.1007/BFb0067150;   http://
www.springerlink.com/content/978-3-540-38663-6.

Dies:1983:CLS

[Die83b] Jacques-Edouard Dies. Classification des librairies et des structures mixtes. (French) []. Lecture Notes in Mathematics, 1010: 150–169, 1983. CODEN LNMAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067159/.

Dies:1983:LLS

[Die83c] Jacques-Edouard Dies. Librairies, librairies stationnaires. (French) [Bookstores, stationary]. Lecture Notes in Mathematics, 1010:19–36, 1983. CODEN LNMAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067152/.

Dies:1983:MSF

[Die83d] Jacques-Edouard Dies. Mesures stationnaires. (French) [Stationary measures]. Lecture Notes in Mathematics, 1010:37–60, 1983. CODEN LNMAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067153/.

Dies:1983:OPT

[Die83e] Jacques-Edouard Dies. Optimalité de la police de transposition. (French) []. Lecture Notes in Mathematics, 1010:171–207, 1983. CODEN LNMAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067160/.

Dies:1983:RPLa

[Die83f] Jacques-Edouard Dies. Récurrence positive des librairies (e, T_{ω}^{∞}, p) et (e, M_{ω}^{∞}, p). (French) []. Lecture Notes in Mathematics, 1010:61–71, 1983. CODEN LNMAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067154/.

Dies:1983:RPLb

[Die83g] Jacques-Edouard Dies. Recurrence positive des librairies mixtes. Lecture Notes in Mathematics, 1010:128–149, 1983. CODEN LNMAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067158/.

Dies:1983:SPF

[Die83h] Jacques-Edouard Dies. Structures de permutation. (French) []. Lecture Notes in Mathematics, 1010:1–18, 1983. CODEN LN-MAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067151/.

Dies:1983:SRS

[Die83i] Jacques-Edouard Dies. Structures recurrences, structures transientes. Lecture Notes in Mathematics, 1010:101–127, 1983. CO-DEN LNMAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067157/.

Dies:1983:TL

[Die83j] Jacques-Edouard Dies. Transience des librairies (e, T_{ω}^{∞}, p). Lecture Notes in Mathematics, 1010:72–84, 1983. CODEN LN-MAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067155/.

Dies:1983:VMF

[Die83k] Jacques-Edouard Dies. Variantes mixtes finies des librairies de Tsetlin transience des librairies (e, M_{ω}^{∞}, p). (French) []. Lecture Notes in Mathematics, 1010:85–100, 1983. CODEN LN-MAA2. ISBN 3-540-12669-4 (print), 3-540-38663-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067156/.

Dimiev:1983:PLF

[Dim83] Stancho Dimiev. Propriétés locales des fonctions presqueholomorphes. (French) []. Lecture Notes in Mathematics, 1039: 102–117, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073360/.

Dineen:1981:HGC

[Din81] Seán Dineen. Holomorphic germs on compact subsets of locally convex spaces. Lecture Notes in Mathematics, 843:247–263, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089277/.

Dinculeanu:1982:WCC

[Din82] Nicolae Dinculeanu. Weak compactness criteria in function spaces over a locally compact group. Lecture Notes in Mathematics, 945: 213–225, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096677/.

Dittmar:1980:NEQ

[Dit80] Bodo Dittmar. Ein neuer existenzbeweis für quasikonforme Abbildungen mit vorgegebener komplexer Dilatation. (German) []. Lecture Notes in Mathematics, 798:148–154, 1980. CODEN LN-MAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097260/.

Dixon:1981:CTC

[Dix81] Martyn Dixon. A conjugacy theorem for Carter subgroups in groups with min-p for all p. Lecture Notes in Mathematics, 848: 161–168, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090563/.

Dembinski:1980:SBS

[DJ80] V. Dembinski and K. Janßen. Standard balayage spaces and standard Markov processes. Lecture Notes in Mathematics, 787:84– 105, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086327/.

Donder:1981:SAC

[DJK81] D. Donder, R. B. Jensen, and B. J. Koppelberg. Some applications of the core model. Lecture Notes in Mathematics, 872:55–97, 1981. CODEN LNMAA2. ISBN 3-540-10849-1 (print), 3-540-38757-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098619/.

Djokovic:1981:CCC

[Djo81] Dragomir Ž. Djoković. Closures of conjugacy classes in classical real linear Lie groups. Lecture Notes in Mathematics, 848:

63–83, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090557/.

Dennis-Jones:1984:SHD

[DJW84] E. C. Dennis-Jones and S. S. Wainer. Subrecursive hierarchies via direct limits. Lecture Notes in Mathematics, 1104:117–128, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099482/.

Dimovski:1980:ITD

[DK80a] Ivan Hristov Dimovski and Virjinia Stoineva Kirjakova. On an integral transformation due to N. Obrechkoff. Lecture Notes in Mathematics, 798:141–147, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097259/.

Draxl:1980:SSB

[DK80b] P. Draxl and M. Kneser, editors. SK_{1} von Schiefkörpern: Seminar Bielefeld–Göttingen, 1976, volume 778 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 778. URL http:

//link.springer.com/book/10.1007/BFb0095915;   http:/
/www.springerlink.com/content/978-3-540-38599-8.

Draxl:1980:VGF

[DK80c] Peter Draxl and Martin Kneser. Vorwort. (German) [Foreword]. Lecture Notes in Mathematics, 778:3–6, 1980. CODEN LN-MAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095916/.

Dahlberg:1982:NAE

[DK82a] Björn E. J. Dahlberg and Carlos E. Kenig. A note on the almost everywhere behavior of solutions to the Schrödinger equation. Lecture Notes in Mathematics, 908:205–209, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093289/.

Delfs:1982:TSW

[DK82b] Hans Delfs and Manfred Knebusch. Zur Theorie der semialgebraischen Wege und Intervalle über einem reell abgeschlossenen Körper. (German) []. Lecture Notes in Mathematics, 959:299– 323, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062260/.

Dress:1982:CSE

[DK82c] Andreas W. M. Dress and Aderemi O. Kuku. A convenient setting for equivariant higher algebraic K-theory. Lecture Notes in Mathematics, 966:59–68, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062166/.

DiBlasio:1984:RAE

[DKS84] G. Di Blasio, K. Kunisch, and E. Sinestrari. Retarded abstract equations in Hilbert spaces. Lecture Notes in Mathematics, 1076: 71–77, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072767/.

Diaz:1984:ODF

[DKT84] Alejandro R. Díaz, Noboru Kikuchi, and J. E. Taylor. Optimal design formulations for finite element grid adaptation. Lecture Notes in Mathematics, 1086:56–76, 1984. CODEN LNMAA2. ISBN 3-540-13871-4 (print), 3-540-39061-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073069/.

Dolgachev:1981:FGC

[DL81] Igor Dolgachev and Anatoly Libgober. On the fundamental group of the complement to a discriminant variety. Lecture Notes in Mathematics, 862:1–25, 1981. CODEN LNMAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090888/.

Davis:1982:HMP

[DL82] Burgess Davis and John L. Lewis. The harmonic measure of porous membranes in R^{3}. Lecture Notes in Mathematics, 939: 29–35, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096257/.

Denenberg:1984:LSC

[DL84] Larry Denenberg and Harry R. Lewis. Logical syntax and computational complexity. Lecture Notes in Mathematics, 1104:101– 115, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099481/.

Dlab:1980:STL

[Dla80] Vlastimil Dlab. Structure des treillis linéaires libres. (French) []. Lecture Notes in Mathematics, 795:10–34, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090112/.

Dlab:1983:RRT

[Dla83] Vlastimil Dlab. The regular representations of the tame hereditary algebras. Lecture Notes in Mathematics, 1029:120–133, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098929/.

delaPradelle:1980:DRR

[dlP80] A. de la Pradelle. Deux remarques sur la représentation intégrale des fonctions excessives ou surharmoniques. (French) []. Lecture Notes in Mathematics, 814:229–239, 1980. CODEN LN-MAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094156/.

delaPradelle:1982:CME

[dlP82] Arnaud de la Pradelle. Sur le caractère métrique d’un espace harmonique de Brelot. (French) []. Lecture Notes in Mathematics, 906:308–320, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093274/.

Dugue:1981:AMP

[DLR81] Daniel Dugué, Eugene Lukacs, and Vijay K. Rohatgi, editors. Analytical Methods in Probability Theory: Proceedings of the Conference Held at Oberwolfach, Germany, June 9–14, 1980, volume 861 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 861. URL http:/

/link.springer.com/book/10.1007/BFb0097305;   http://
www.springerlink.com/content/978-3-540-36785-7.

deMonvel:1980:NVP

[dM80a] Louis Boutet de Monvel. Nombre de valeurs propres d’un opérateur elliptique et polynôme de Hilbert–Samuel [d’après V. Guillemin]. (French) []. Lecture Notes in Mathematics, 770:120– 131, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096239/.

Dodd:1980:BT

[DM80b] R. K. Dodd and H. C. Morris. Bäcklund transformations. Lecture Notes in Mathematics, 810:63–94, 1980. CODEN LN-MAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089974/.

Dodd:1980:GBT

[DM80c] R. K. Dodd and H. C. Morris. Generalised Bäcklund transformations for integrable evolution equations associated with N th order scattering problems. Lecture Notes in Mathematics, 810: 95–98, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089975/.

deMedrano:1981:TAM

[dM81a] Santiago López de Medrano. Topological aspects of matrix problems. Lecture Notes in Mathematics, 903:196–210, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092993/.

Deligne:1981:TCa

[DM81b] P. Deligne and J. S. Milne. Tannakian categories. Lecture Notes in Mathematics, 900:101–228, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38955-2_4/.

Deligne:1981:TCb

[DM81c] P. Deligne and J. S. Milne. Tannakian categories. Lecture Notes in Mathematics, 900:419, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38955-2_13.

Desolneux-Moulis:1981:OPS

[DM81d] Nicole Desolneux-Moulis. Orbites périodiques des systèmes Hamiltoniens autonomes [d’après Clarke, Ekeland–Lasry, Moser, Rabinowitz, Weinstein]. (French) []. Lecture Notes in Mathematics, 842:156–173, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089933/.

Diamond:1981:CEE

[DM81e] Harold G. Diamond and Kevin S. McCurley. Constructive elementary estimates for M(x). Lecture Notes in Mathematics, 899: 239–253, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096465/.

Dubois:1982:HKP

[DM82] Jacques Dubois and Pedro Morales. On the Hukuhara–Kneser property for some Cauchy problems in locally convex topological vector spaces. Lecture Notes in Mathematics, 964:162–170, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064996/.

deMoura:1983:LUN

[dM83] Carlos A. de Moura. A linear uncoupling numerical scheme for the nonlinear coupled thermoelastodynamics equations. Lecture Notes in Mathematics, 1005:204–211, 1983. CODEN LN-MAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112535/.

Davis:1984:CIC

[DM84] E. Davis and P. Maroscia. Complete intersections in P^{2}: Cayley– Bacharach characterizations. Lecture Notes in Mathematics, 1092: 253–269, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099368/.

Gonzalez:1983:RMV

[dMG83] J. L. de María González. Sur la régularité d’une mesure à valeurs dans un semigroupe. (French) []. Lecture Notes in Mathematics, 1033:230–239, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099859/.

Deligne:1981:FM

[DMOyS81a] Pierre Deligne, James S. Milne, Arthur Ogus, and Kuang yen Shih. Front matter. Lecture Notes in Mathematics, 900:i– v, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-38955-2/1.pdf.

Deligne:1981:GI

[DMOyS81b] Pierre Deligne, James S. Milne, Arthur Ogus, and Kuang yen Shih. General introduction. Lecture Notes in Mathematics, 900: 1–7, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38955-2_1/.

Deligne:1981:HCM

[DMOyS81c] Pierre Deligne, James S. Milne, Arthur Ogus, and Kuang yen Shih. Hodge Cycles, Motives, and Shimura Varieties, volume 900 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 423 (est.) pp. LCCN QA3 .L28 no. 900. URL http://link.

springer.com/book/10.1007/978-3-540-38955-2;   http:/
/www.springerlink.com/content/978-3-540-38955-2.

Deligne:1981:NC

[DMOyS81d] Pierre Deligne, James S. Milne, Arthur Ogus, and Kuang yen Shih. Notations and conventions. Lecture Notes in Mathematics, 900:8, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38955-2_2.

Dinh:1983:ASN

[DMPG83] Q. V. Dinh, B. Mantel, J. Periaux, and R. Glowinski. Approximate solution of the Navier–Stokes equations for incompressible viscous fluids, related domain decomposition methods. Lecture Notes in Mathematics, 1005:46–86, 1983. CODEN LN-MAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112525/.

deMelo:1981:CDM

[dMPvS81] W. de Melo, J. Palis, and S. J. van Strien. Characterising diffeomorphisms with modulus of stability one. Lecture Notes in Mathematics, 898:266–285, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091919/.

Degiovanni:1984:CPE

[DMT84] Marco Degiovanni, Antonio Marino, and Mario Tosques. Critical points and evolution equations. Lecture Notes in Mathematics, 1091:184–192, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098811/.

deOliveira:1980:NSS

[dO80a] M. C. de Oliveira. Non-singular structural stable flows on threedimensional manifolds. Lecture Notes in Mathematics, 799:322– 333, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089322/.

Deville:1980:SCS

[DO80b] Michel Deville and Steven A. Orszag. Spectral calculation of the stability of the circular Couette flow. Lecture Notes in Mathematics, 771:159–176, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086905/.

Dur:1982:IAE

[DO82] Arne Dür and Ulrich Oberst. Incidence algebras, exponential formulas and unipotent groups. Lecture Notes in Mathematics, 969: 133–166, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062991/.

Doitchinov:1984:SSS

[Doi84] D. Doitchinov. Supertopological spaces and a special class of extensions of topological spaces. Lecture Notes in Mathematics, 1060:17–25, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099917/.

Dolgachev:1982:WPV

[Dol82] Igor Dolgachev. Weighted projective varieties. Lecture Notes in Mathematics, 956:34–71, 1982. CODEN LNMAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101508/.

Dolbeault:1983:HCG

[Dol83a] Pierre Dolbeault. On holomorphic chains with given boundary in P^{n}(C). Lecture Notes in Mathematics, 1039:118–129, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073361/.

Dolgachev:1983:AGP

[Dol83b] I. Dolgachev, editor. Algebraic Geometry: Proceedings of the Third Midwest Algebraic Geometry Conference held at the University of Michigan, Ann Arbor, USA, November 14–15, 1981, volume 1008 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1008. URL http:/

/link.springer.com/book/10.1007/BFb0065695;   http://
www.springerlink.com/content/978-3-540-40971-7.

Domar:1980:SS

[Dom80] Yngve Domar. On spectral synthesis in R^{n}, n ≥ 2. Lecture Notes in Mathematics, 779:46–72, 1980. CODEN LNMAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087667/.

Domar:1983:BAR

[Dom83a] Y. Domar. A Banach algebra related to the disk algebra. Lecture Notes in Mathematics, 975:309–311, 1983. CODEN LN-MAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064563/.

Domar:1983:BTI

[Dom83b] Y. Domar. Bilaterally translation-invariant subspaces of weighted L^{p}(R). Lecture Notes in Mathematics, 975:210–213, 1983. CO-DEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064551/.

Domar:1983:STI

[Dom83c] Yngve Domar. A solution of the translation-invariant subspace problem for weighted L^{p} on R, R or R^{+}. Lecture Notes in Mathematics, 975:214–226, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064552/.

Donder:1981:CM

[Don81] Hans-Dieter Donder. Coarse morasses in L. Lecture Notes in Mathematics, 872:37–54, 1981. CODEN LNMAA2. ISBN 3-540-10849-1 (print), 3-540-38757-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098618/.

Donner:1982:BSF

[Don82a] Klaus Donner. Bisublinear and subbilinear functionals. Lecture Notes in Mathematics, 904:30–67, 1982. CODEN LN-MAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094136/.

Donner:1982:CEV

[Don82b] Klaus Donner. Cone embeddings for vector lattices. Lecture Notes in Mathematics, 904:2–11, 1982. CODEN LNMAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094134/.

Donner:1982:CVL

[Don82c] Klaus Donner. Convergence to vector lattice homomorphisms and essential sets. Lecture Notes in Mathematics, 904:162–173, 1982. CODEN LNMAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094141/.

Donner:1982:EVP

[Don82d] Klaus Donner. Extension of L^{1}-valued positive operators. Lecture Notes in Mathematics, 904:68–83, 1982. CODEN LN-MAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094137/.

Donner:1982:EPOb

[Don82e] Klaus Donner. Extension of Positive Operators and Korovkin Theorems, volume 904 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 182 pp. LCCN QA3 .L28 no. 904. URL http:

//link.springer.com/book/10.1007/BFb0094133;   http:/
/www.springerlink.com/content/978-3-540-38969-9.

Donner:1982:EPOa

[Don82f] Klaus Donner. Extension of positive operators in L^{p}-spaces. Lecture Notes in Mathematics, 904:84–104, 1982. CODEN LN-MAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094138/.

Donner:1982:KCE

[Don82g] Klaus Donner. The Korovkin closure for equicontinuous nets of positive operators. Lecture Notes in Mathematics, 904:105– 126, 1982. CODEN LNMAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094139/.

Donner:1982:KTI

[Don82h] Klaus Donner. Korovkin theorems for the identity mapping on classical Banach lattices. Lecture Notes in Mathematics, 904:127– 161, 1982. CODEN LNMAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094140/.

Donner:1982:VVH

[Don82i] Klaus Donner. A vector-valued Hahn–Banach theorem. Lecture Notes in Mathematics, 904:12–29, 1982. CODEN LN-MAA2. ISBN 3-540-11183-2 (print), 3-540-38969-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094135/.

Dora:1981:QRS

[Dor81] J. Della Dora. Quelques résultats sur la structure des tables de Padé–Hermite. (French) []. Lecture Notes in Mathematics, 888: 173–184, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095584/.

Douglas:1982:SMD

[Dou82] Jim Douglas, Jr. Simulation of miscible displacement in porous media by a modified method of characteristic procedure. Lecture Notes in Mathematics, 912:64–70, 1982. CODEN LN-MAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093149/.

Doyen:1981:LSS

[Doy81] Jean Doyen. Linear spaces and Steiner systems. Lecture Notes in Mathematics, 893:30–42, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091011/.

Dozo:1981:ACP

[Doz81a] E. Lami Dozo. Asymptotic centers in particular spaces. Lecture Notes in Mathematics, 886:199–207, 1981. CODEN LN-MAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092185/.

Dozzi:1981:DIT

[Doz81b] Markus Dozzi. On the decomposition and integration of twoparameter stochastic processes. Lecture Notes in Mathematics, 863:162–171, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091099/.

Damiano:1982:SRF

[DP82a] Robert F. Damiano and Zoltan Papp. Stable rings with finite global dimension. Lecture Notes in Mathematics, 951:47–67, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067324/.

Doebner:1982:TGN

[DP82b] Heinz-Dietrich Doebner and Tchavdar D. Palev, editors. Twistor Geometry and Non-Linear Systems: Review Lectures given at the 4th Bulgarian Summer School on Mathematical Problems of Quantum Field Theory, Held at Primorsko, Bulgaria, September 1980, volume 970 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 970. URL http:/

/link.springer.com/book/10.1007/BFb0066021;   http://
www.springerlink.com/content/978-3-540-39418-1.

DeConcini:1983:CSV

[DP83a] C. De Concini and C. Procesi. Complete symmetric varieties. Lecture Notes in Mathematics, 996:1–44, 1983. CO-DEN LNMAA2. ISBN 3-540-12319-9 (print), 3-540-40043-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063234/;
 http://link.springer.com/content/pdf/bfm:978-3-540-40043-
1/1.pdf.

Dobrushin:1983:CUG

[DP83b] R. L. Dobrushin and E. A. Pecherski. A criterion of the uniqueness of Gibbsian fields in the non-compact case. Lecture Notes in Mathematics, 1021:97–110, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072907/.

Demazure:1980:SSS

[DPT80] Michel Demazure, Henry Charles Pinkham, and Bernard Teissier, editors. Séminaire sur les Singularités des Surfaces: Centre de Mathématiques de l’École Polytechnique, Palaiseau 1976– 1977, volume 777 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 777. URL http:/

/link.springer.com/book/10.1007/BFb0085872;   http://
www.springerlink.com/content/978-3-540-38586-8.

Deuber:1982:CPT

[DPV82] W. Deuber, H. J. Prömel, and B. Voigt. A canonical partition theorem for chains in regular trees. Lecture Notes in Mathematics, 969:115–132, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062990/.

Dlab:1980:PAM

[DR80a] Vlastimil Dlab and Claus Michael Ringel. The preprojective algebra of a modulated graph. Lecture Notes in Mathematics, 832: 216–231, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088465/.

Dovermann:1980:ESS

[DR80b] Karl Heinz Dovermann and Melvin Rothenberg. An equivariant surgery sequence and equivariant diffeomorphism and homeomorphism classification (a survey). Lecture Notes in Mathematics, 788:257–280, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099249/.

Dayton:1981:SUP

[DR81] Barry H. Dayton and Leslie G. Roberts. Seminormality of unions of planes. Lecture Notes in Mathematics, 854:93–123, 1981. CO-DEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089517/.

DiMasi:1982:AMN

[DR82a] Giovanni B. Di Masi and Wolfgang J. Runggaldier. On approximation methods for nonlinear filtering. Lecture Notes in Mathematics, 972:249–259, 1982. CODEN LNMAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064864/.

Dubois:1982:SSS

[DR82b] D. W. Dubois and Tomás Recio. Subordinate structure sheaves. Lecture Notes in Mathematics, 959:324–342, 1982. CODEN LN-MAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062261/.

Draxl:1980:VDB

[Dra80a] P. Draxl. SK_{1} Bei Vollständig Diskret Bewerteten Körpern. (German) []. Lecture Notes in Mathematics, 778:55–66, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095923/.

Draxl:1980:ZVD

[Dra80b] P. Draxl. SK_{1} Bei Zweifach Vollständig Diskret Bewerteten Körpern. (German) []. Lecture Notes in Mathematics, 778:67– 75, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095924/.

Draxl:1980:SBR

[Dra80c] P. Draxl. SK_{1} von Schiefkorpern über Beliebig–Rangig Henselsch Bewerteten Körpern. (German) []. Lecture Notes in Mathematics, 778:94–100, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095926/.

Draxl:1980:AEB

[Dra80d] P. Draxl. Anhang: Ergänzende Bemerkungen Zum 8. Vortrag. (German) []. Lecture Notes in Mathematics, 778:76–93, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095925/.

Draxl:1980:EVK

[Dra80e] P. Draxl. Eigenschaften von SK_{1} und ‘Vernünftige’ Körper. (German) [Properties of SK_{1} and ‘reasonable’ bodies]. Lecture Notes in Mathematics, 778:47–54, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095922/.

Draxl:1980:LDD

[Dra80f] Peter Draxl. Eine Liftung der Dieudonné-Determinante und Anwendungen die Multiplikative Gruppe eines Schiefkörpers Betreffend. (German) []. Lecture Notes in Mathematics, 778:101– 116, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095927/.

Draux:1981:AET

[Dra81] André Draux. Approximants of exponential type general orthogonal polynomials. Lecture Notes in Mathematics, 888:185– 196, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095585/.

Dragan:1983:AES

[Dră83a] V. Drăgan. Asymptotic expansions for singularly perturbed differential matrix Riccati equations with applications to linearquadratic optimization problems. Lecture Notes in Mathematics, 985:370–399, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062377/.

Draux:1983:APD

[Dra83b] André Draux. Approximants de Padé en deux points. (French) []. Lecture Notes in Mathematics, 974:439–552, 1983. CODEN LNMAA2. ISBN 3-540-11984-1 (print), 3-540-39450-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066476/.

Draux:1983:ASF

[Dra83c] André Draux. Approximants des series de fonctions. (French) []. Lecture Notes in Mathematics, 974:553–600, 1983. CODEN LNMAA2. ISBN 3-540-11984-1 (print), 3-540-39450-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066477/.

Draux:1983:FLP

[Dra83d] André Draux. Fonctionnelles linéaires particulières. (French) []. Lecture Notes in Mathematics, 974:178–274, 1983. CODEN LN-MAA2. ISBN 3-540-11984-1 (print), 3-540-39450-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066473/.

Draux:1983:POFb

[Dra83e] André Draux. Polynômes Orthogonaux Formels —: Applications. (French) [Formal orthogonal polynomials] — applications, volume 974 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-11984-1 (print), 3-540-39450-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 625 pp. LCCN QA3 .L28 no. 974. URL http:

//link.springer.com/book/10.1007/BFb0066470;   http:/
/www.springerlink.com/content/978-3-540-39450-1.

Draux:1983:POFa

[Dra83f] André Draux. Polynômes orthogonaux. (French) [Orthogonal polynomials]. Lecture Notes in Mathematics, 974:1–96, 1983. CODEN LNMAA2. ISBN 3-540-11984-1 (print), 3-540-39450-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066471/.

Draux:1983:QGF

[Dra83g] André Draux. Quadratures de Gauss. (French) []. Lecture Notes in Mathematics, 974:405–438, 1983. CODEN LNMAA2. ISBN 3-540-11984-1 (print), 3-540-39450-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066475/.

Draux:1983:RTA

[Dra83h] André Draux. Relations tous azimuts. (French) [All-azimuth relations]. Lecture Notes in Mathematics, 974:275–404, 1983. CODEN LNMAA2. ISBN 3-540-11984-1 (print), 3-540-39450-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066474/.

Draux:1983:SAP

[Dra83i] André Draux. Systèmes adjacents de polynômes orthogonaux. (French) []. Lecture Notes in Mathematics, 974:97–177, 1983. CODEN LNMAA2. ISBN 3-540-11984-1 (print), 3-540-39450-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066472/.

Draux:1984:PAN

[Dra84] André Draux. The Padé approximants in a non-commutative algebra and their applications. Lecture Notes in Mathematics, 1071:117–131, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099613/.

Dress:1983:GES

[Dre83] Andreas Dress. Global equilibria and steady states of discrete networks, according to classical thermodynamics. Lecture Notes in Mathematics, 1031:46–61, 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072110/.

Drisch:1982:FET

[Dri82] T. Drisch. First elements of a theory of quantum mechanical limit distributions. Lecture Notes in Mathematics, 928:46–59, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093218/.

Drozd:1980:TWM

[Dro80] Ju. A. Drozd. Tame and wild matrix problems. Lecture Notes in Mathematics, 832:242–258, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088467/.

Dumortier:1981:GDP

[DRR81a] Freddy Dumortier, Paulo R. Rodrigues, and Robert Roussarie. Germs of Diffeomorphisms in the Plane, volume 902 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11177-8 (print), 3-540-38958-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 197 pp. LCCN QA3 .L28 no. 902. URL http:/

/link.springer.com/book/10.1007/BFb0095812;   http://
www.springerlink.com/content/978-3-540-38958-3.

Dumortier:1981:IDF

[DRR81b] Freddy Dumortier, Paulo R. Rodrigues, and Robert Roussarie. Introduction, definitions, formal study and statement of the results. Lecture Notes in Mathematics, 902:7–47, 1981. CODEN LNMAA2. ISBN 3-540-11177-8 (print), 3-540-38958-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095814/.

Dumortier:1981:PR

[DRR81c] Freddy Dumortier, Paulo R. Rodrigues, and Robert Roussarie. Proof of the C^{∞} results. Lecture Notes in Mathematics, 902: 126–172, 1981. CODEN LNMAA2. ISBN 3-540-11177-8 (print), 3-540-38958-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095817/.

Dumortier:1981:PTR

[DRR81d] Freddy Dumortier, Paulo R. Rodrigues, and Robert Roussarie. Proof of the topological results. Lecture Notes in Mathematics, 902:173–192, 1981. CODEN LNMAA2. ISBN 3-540-11177-8 (print), 3-540-38958-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095818/.

Dumortier:1981:STT

[DRR81e] Freddy Dumortier, Paulo R. Rodrigues, and Robert Roussarie. Stability of type I- and type II-singularities. Lecture Notes in Mathematics, 902:48–82, 1981. CODEN LNMAA2. ISBN 3-540-11177-8 (print), 3-540-38958-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095815/.

Dumortier:1981:STI

[DRR81f] Freddy Dumortier, Paulo R. Rodrigues, and Robert Roussarie. Stability of type III-singularities. Lecture Notes in Mathematics, 902:83–125, 1981. CODEN LNMAA2. ISBN 3-540-11177-8 (print), 3-540-38958-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095816/.

Dumortier:1981:SSM

[DRR81g] Freddy Dumortier, Paulo R. Rodrigues, and Robert Roussarie. Summary, some motivation and acknowledgments. Lecture Notes in Mathematics, 902:1–6, 1981. CODEN LNMAA2. ISBN 3-540-11177-8 (print), 3-540-38958-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095813/.

Dowbor:1980:HAR

[DRS80] P. Dowbor, C. M. Ringel, and D. Simson. Hereditary Artinian rings of finite representation type. Lecture Notes in Mathematics, 832:232–241, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088466/.

Drury:1983:RNI

[Dru83] S. W. Drury. Remarks on von Neumann’s inequality. Lecture Notes in Mathematics, 995:14–32, 1983. CODEN LN-MAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061886/.

Demailly:1980:REN

[DS80] J.-P. Demailly and H. Skoda. Relations entre les notions de positivités de P. A. Griffiths et de S. Nakano pour les fibrés vectoriels. (French) []. Lecture Notes in Mathematics, 822:304–309, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097764/.

deSalas:1982:VTB

[dS82] Juan B. Sancho de Salas. A vanishing theorem for birational morphisms. Lecture Notes in Mathematics, 961:275–284, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071288/.

Drozda:1983:JTI

[DSS83] Wiesława Drozda, Anna Szynal, and Jan Szynal. The Jenkins’ type inequality for Bazilevič functions. Lecture Notes in Mathematics, 1039:130–141, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073362/.

Dervieux:1980:FEM

[DT80] A. Dervieux and F. Thomasset. A finite element method for the simulation of a Rayleigh–Taylor instability. Lecture Notes in Mathematics, 771:145–158, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086904/.

Diestel:1983:PVM

[DU83] Joe Diestel and J. Jerry Uhl, Jr. Progress in vector measures — 1977–83. Lecture Notes in Mathematics, 1033:144–192, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099853/.

Dubois:1982:OLL

[Dub82a] R.-M. Dubois. Opératurs locaux localement dissipatifs à valeurs dans un espace de Hilbert. (French) []. Lecture Notes in Mathematics, 906:88–102, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093262/.

Dubois:1982:PMA

[Dub82b] R.-M. Dubois. Problèmes mixtes abstraits et principe du maximum parabol Ioue. (French) []. Lecture Notes in Mathematics, 906:103–113, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093263/.

Dubois:1984:SPD

[Dub84] Roger-Marie Dubois. Systèmes paraboliques et dissipativité pour différentes normes. (French) []. Lecture Notes in Mathematics, 1061:92–108, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099019/.

Dudley:1981:SRR

[Dud81] R. M. Dudley. Some recent results on empirical processes. Lecture Notes in Mathematics, 860:107–123, 1981. CODEN LN-MAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090611/.

Dudley:1984:CEP

[Dud84] R. M. Dudley. A course on empirical processes. Lecture Notes in Mathematics, 1097:1–142, 1984. CODEN LNMAA2. ISBN 3-540-13897-8 (print), 3-540-39109-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099432/.

Duflo:1981:CGL

[Duf81] Michel Duflo. Caractères des groupes de Lie résolubles. (French) []. Lecture Notes in Mathematics, 842:257–272, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089939/.

Duff:1982:FMT

[Duf82a] Iain S. Duff. Full matrix techniques in sparse Gaussian elimination. Lecture Notes in Mathematics, 912:71–84, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093150/.

Duff:1982:SMS

[Duf82b] Iain S. Duff. Sparse matrix software for elliptic PDE’s. Lecture Notes in Mathematics, 960:410–426, 1982. CODEN LN-MAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069936/.

Duff:1983:MAS

[Duf83] Iain Duff. A multifrontal approach for solving sparse linear equations. Lecture Notes in Mathematics, 1005:87–98, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112526/.

Duflo:1984:PFA

[Duf84] Michel Duflo. On the Plancherel formula for almost algebraic real Lie groups. Lecture Notes in Mathematics, 1077:101–165, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072338/.

Duistermaat:1982:AES

[Dui82] J. J. Duistermaat. Asymptotics of elementary spherical functions. Lecture Notes in Mathematics, 905:100–107, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092430/.

Duistermaat:1984:BPS

[Dui84] J. J. Duistermaat. Bifurcations of periodic solutions near equilibrium points of Hamiltonian systems. Lecture Notes in Mathematics, 1057:57–105, 1984. CODEN LNMAA2. ISBN 3-540-12931-6 (print), 3-540-38812-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098594/.

Dumcke:1984:MLM

[Düm84] R. Dümcke. Markovian limits of multi time correlation functions for open quantum systems. Lecture Notes in Mathematics, 1055: 113–118, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071715/.

Dung:1982:ESA

[Dun82] Nguyen Xuan Dung. Essential self-adjointness and selfadjointness for Generalized Schrödinger Operators. Lecture Notes in Mathematics, 964:171–178, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064997/.

Dupeyrat:1980:ESM

[Dup80] M. Dupeyrat. Experimental study of the mechanism of a new hydrodynamical instability observed at some interfaces between immiscible liquids. Lecture Notes in Mathematics, 782:122– 129, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090430/.

Duret:1980:CFA

[Dur80] Jean-Louis Duret. Les corps faiblement algébriquement clos non séparablement clos ont la propriété d’indépendance. (French) []. Lecture Notes in Mathematics, 834:136–162, 1980. CODEN LN-MAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090164/.

Duval:1980:PDS

[Duv80] C. Duval. On the prequantum description of spinning particles in an external gauge field. Lecture Notes in Mathematics, 836: 49–66, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089726/.

Duval:1983:EDA

[Duv83a] Anne Duval. Équations aux différences algébriques: solutions méromorphes dans C-R^{−}. système fondamental de solutions holomorphes dans un demi-plan. (French) []. Lecture Notes in Mathematics, 1015:102–135, 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071351/.

Duval:1983:EAI

[Duv83b] Anne Duval. Étude asymptotique d’une intégrale analogue à la fonction ‘Γ modifiée’. (French) []. Lecture Notes in Mathematics, 1015:50–63, 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071349/.

Duval:1983:SIE

[Duv83c] Anne Duval. Solutions irregulîeres d’équations aux différences polynômiales. (French) []. Lecture Notes in Mathematics, 1015: 64–101, 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071350/.

deVerdiere:1981:MSP

[dV81] Y. Colin de Verdière. La matrice de scattering pour l’opérateur de Schrödinger sur la droite réelle. (French) []. Lecture Notes in Mathematics, 842:246–256, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089938/.

daVeiga:1984:FDL

[dV84] H. Beirão da Veiga, editor. Fluid Dynamics: Lectures given at the 3rd 1982 Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Varenna, Italy, August 22–September 1, 1982, volume 1047 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12893-X (print), 3-540-38773-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1047. URL http:/

/link.springer.com/book/10.1007/BFb0072325;   http://
www.springerlink.com/content/978-3-540-38773-2.

Dayton:1981:SST

[DW81] Barry H. Dayton and Charles A. Weibel. A spectral sequence for the K-theory of affine glued schemes. Lecture Notes in Mathematics, 854:24–92, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089516/.

Dwilewicz:1983:DCR

[Dwi83] Roman Dwilewicz. Division of Cauchy–Riemann functions on hypersurfaces. Lecture Notes in Mathematics, 1039:142–167, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073363/.

Dydak:1981:LCQ

[Dyd81] Jerzy Dydak. Local n-connectivity of quotient spaces and onepoint compactifications. Lecture Notes in Mathematics, 870:48– 72, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089707/.

Dynkin:1981:RPA

[Dyn81] E. M. Dyn’kin. The rate of polynomial approximation in the complex domain. Lecture Notes in Mathematics, 864:90–142, 1981. CODEN LNMAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096997/.

Dugas:1981:IDS

[DZH81] M. Dugas and B. Zimmermann-Huisgen. Iterated direct sums and products of modules. Lecture Notes in Mathematics, 874: 179–193, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090534/.

Darling:1983:TCM

[DZM83] R. W. Darling, W. A. Zheng, and P.-A. Meyer. Le théorème de convergence des martingales dans les variétés riemanniennes. (French) []. Lecture Notes in Mathematics, 986:187–193, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068314/.

Eades:1980:CD

[Ead80] Peter Eades. Circulant (v, k, μ) designs. Lecture Notes in Mathematics, 829:83–93, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088903/.

Eades:1981:ROT

[Ead81] Peter Eades. Regularity and optimality for trees. Lecture Notes in Mathematics, 884:148–154, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091816/.

Eastham:1982:ATC

[Eas82a] M. S. P. Eastham. Asymptotic theory for a critical class of fourthorder differential equations. Lecture Notes in Mathematics, 964: 179–186, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064998/.

Eastwood:1982:TTP

[Eas82b] M. G. Eastwood. Twistor theory (the Penrose transform). Lecture Notes in Mathematics, 950:1–11, 1982. CODEN LNMAA2. ISBN 3-540-11596-X (print), 3-540-39366-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061875/.

Eastwood:1982:TDL

[Eas82c] Michael G. Eastwood. The twistor description of linear fields. Lecture Notes in Mathematics, 926:190–198, 1982. CODEN LN-MAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092663/.

Eastham:1983:LGA

[Eas83] M. S. P. Eastham. The Liouville–Green asymptotic theory for second-order differential equations: a new approach and some extensions. Lecture Notes in Mathematics, 1032:110–122, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076795/.

Eberlein:1983:CFB

[Ebe83] Ernst Eberlein. On the covariance function of Banach space valued very weak Bernoulli processes. Lecture Notes in Mathematics, 990:84–91, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064264/.

Eccles:1980:MPC

[Ecc80] Peter John Eccles. Multiple points of codimension one immersions. Lecture Notes in Mathematics, 788:23–38, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099236/.

Edgar:1980:LJS

[Edg80] G. A. Edgar. A long James space. Lecture Notes in Mathematics, 794:31–37, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088209/.

Edgar:1982:PCS

[Edg82] G. A. Edgar. On pointwise-compact sets of measurable functions. Lecture Notes in Mathematics, 945:24–28, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096661/.

Edgar:1983:TIR

[Edg83] G. A. Edgar. Two integral representations. Lecture Notes in Mathematics, 1033:193–198, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099854/.

Eckhaus:1982:TAS

[EdJ82] W. Eckhaus and E. M. de Jager, editors. Theory and Applications of Singular Perturbations: Proceedings of a Conference Held in Oberwolfach, August 16–22, 1981, volume 942 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 942.

URL http://link.springer.com/book/10.1007/BFb0094735;
 http://www.springerlink.com/content/978-3-540-39332-
0.

Edmunds:1982:ENN

[Edm82] D. E. Edmunds. Entropy numbers, s-numbers and eigenvalues. Lecture Notes in Mathematics, 964:213–231, 1982. CODEN LN-MAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065000/.

Edwards:1980:CID

[Edw80] Robert D. Edwards. Characterizing infinite dimensional manifolds topologically [after Henryk Toruńczyk]. Lecture Notes in Mathematics, 770:278–302, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096247/.

Eells:1980:EHH

[Eel80] James Eells. Examples of harmonic and holomorphic maps. Lecture Notes in Mathematics, 798:155–160, 1980. CODEN LN-MAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097261/.

Eells:1982:CAP

[Eel82] James Eells, editor. Complex Analysis: Proceedings of the Summer School Held at the International Centre for Theoretical Physics Trieste July 5–30, 1980, volume 950 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11596-X (print), 3-540-39366-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 950.

URL http://link.springer.com/book/10.1007/BFb0061874;
 http://www.springerlink.com/content/978-3-540-39366-
5.

Erven:1981:CCL

[EF81a] Joachim Erven and Bernd-Jürgen Falkowski. Continuous cohomology of Lie groups and Lie algebras. Lecture Notes in Mathematics, 877:1–10, 1981. CODEN LNMAA2. ISBN 3-540-10864-5 (print), 3-540-38780-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097354/.

Erven:1981:CTP

[EF81b] Joachim Erven and Bernd-Jürgen Falkowski. Continuous tensor products, infinitely divisible and factorizable representations. Lecture Notes in Mathematics, 877:11–31, 1981. CODEN LN-MAA2. ISBN 3-540-10864-5 (print), 3-540-38780-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097355/.

Erven:1981:FOCb

[EF81c] Joachim Erven and Bernd-Jürgen Falkowski. First order cohomology for SL(2; R) and SL(2; C). Lecture Notes in Mathematics, 877:48–77, 1981. CODEN LNMAA2. ISBN 3-540-10864-5 (print), 3-540-38780-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097357/.

Erven:1981:FOCa

[EF81d] Joachim Erven and Bernd-Jürgen Falkowski. First order cohomology groups for certain semi-direct products. Lecture Notes in Mathematics, 877:32–47, 1981. CODEN LNMAA2. ISBN 3-540-10864-5 (print), 3-540-38780-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097356/.

Erven:1981:FRS

[EF81e] Joachim Erven and Bernd-Jürgen Falkowski. Further results on semi-simple Lie groups. Lecture Notes in Mathematics, 877:78– 103, 1981. CODEN LNMAA2. ISBN 3-540-10864-5 (print), 3-540-38780-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097358/.

Erven:1981:GID

[EF81f] Joachim Erven and Bernd-Jürgen Falkowski. ’genuine’ infinitely divisible representations. Lecture Notes in Mathematics, 877:104– 120, 1981. CODEN LNMAA2. ISBN 3-540-10864-5 (print), 3-540-38780-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097359/.

Erven:1981:LOC

[EF81g] Joachim Erven and Bernd-Jürgen Falkowski. Low Order Cohomology and Applications, volume 877 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10864-5 (print), 3-540-38780-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 126 pp. LCCN QA3 .L28 no. 877.

URL http://link.springer.com/book/10.1007/BFb0097353;
 http://www.springerlink.com/content/978-3-540-38780-
0.

Erven:1982:CCI

[EF82] Joachim Erven and Bernd-Jürgen Falkowski. Continuous cohomology, infinitely divisible positive definite functions and continuous tensor products for SU(1, 1). Lecture Notes in Mathematics, 928:76–89, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093220/.

Eades:1982:EHI

[EFG82] Peter Eades, L. Foulds, and J. Giffin. An efficient heuristic for identifying a maximum weight planar subgraph. Lecture Notes in Mathematics, 952:239–251, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061982/.

Efroymson:1982:ETN

[Efr82] Gustave A. Efroymson. The extension theorem for Nash functions. Lecture Notes in Mathematics, 959:343–357, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062262/.

Eastham:1981:ATD

[EG81] M. S. P. Eastham and C. G. M. Grudniewicz. Asymptotic theory and deficiency indices for fourth and higher order self-adjoint equations: a simplified approach. Lecture Notes in Mathematics, 846:88–99, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089828/.

Eggleton:1980:PRI

[EH80] R. B. Eggleton and D. A. Holton. Pseudographic realizations of an infinitary degree sequence. Lecture Notes in Mathematics, 829:94–109, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088904/.

Eggleton:1981:SMR

[EH81a] R. B. Eggleton and D. A. Holton. Simple and multigraphic realizations of degree sequences. Lecture Notes in Mathematics, 884: 155–172, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091817/.

Eklof:1981:REG

[EH81b] Paul C. Eklof and Martin Huber. On the p-ranks of ext(a, g), assuming ch. Lecture Notes in Mathematics, 874:93–108, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090528/.

Eisenbud:1983:BNT

[EH83] D. Eisenbud and J. Harris. On the Brill–Noether theorem. Lecture Notes in Mathematics, 997:131–137, 1983. CODEN LN-MAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061640/.

Erickson:1984:GRT

[EH84] Martin Erickson and Frank Harary. Generalized Ramsey theory XV: Achievement and avoidance games for bipartite graphs. Lecture Notes in Mathematics, 1073:212–216, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073119/.

Epstein:1981:HSD

[EHK81] Richard L. Epstein, Richard Haas, and Richard L. Kramer. Hierarchies of sets and degrees below 0′. Lecture Notes in Mathematics, 859:32–48, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090937/.

Ennis:1983:FAS

[EHP83] Christopher Ennis, Morris W. Hirsch, and Charles Pugh. Foliations that are not approximable by smoother ones. Lecture Notes in Mathematics, 1007:146–176, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061416/.

Ehrhard:1983:PSE

[Ehr83] Antoine Ehrhard. Un principe de symétrisation dans les espaces de Gauss. (French) []. Lecture Notes in Mathematics, 990:92– 101, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064265/.

Ehrhard:1984:ISL

[Ehr84] Antoine Ehrhard. Sur l’inégalité de Sobolev logarithmique de gross. (French) []. Lecture Notes in Mathematics, 1059:194– 196, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100044/.

Eisenbud:1981:RNB

[Eis81] David Eisenbud. Report on normal bundles of curves in P_{3}. Lecture Notes in Mathematics, 867:141–147, 1981. CODEN LN-MAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090384/.

Eisenstat:1983:SOG

[Eis83] Stanley C. Eisenstat. Some observations on the generalized conjugate gradient method. Lecture Notes in Mathematics, 1005: 99–107, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112527/.

Ekeland:1983:DVM

[Eke83] I. Ekeland. Dual variational methods in non-convex optimization and differential equations. Lecture Notes in Mathematics, 979:108–120, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066251/.

Eklof:1983:STSa

[Ekl83a] Paul C. Eklof. Set theory and structure theorems. Lecture Notes in Mathematics, 1006:275–284, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21560-9_13/.

Eklof:1983:STSb

[Ekl83b] Paul C. Eklof. Set theory and structure theorems. Lecture Notes in Mathematics, 1006:275–284, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103708/.

ElMir:1980:FPE

[El 80] Hassine El Mir. Fonctions plurisousharmoniques et ensembles polaires. (French) []. Lecture Notes in Mathematics, 822:61–76, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097749/.

ElKaroui:1981:APD

[El 81] N. El Karoui. Les aspects probabilistes du contrôle stochastique. (French) []. Lecture Notes in Mathematics, 876:73–238, 1981. CODEN LNMAA2. ISBN 3-540-10860-2 (print), 3-540-38774-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097499/.

ElKaroui:1984:TDP

[El 84] Nicole El Karoui. Théorie du potentiel et contrôle stochastique. (French) []. Lecture Notes in Mathematics, 1096:261–279, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100114/.

Everitt:1983:ODE

[EL83] W. N. Everitt and R. T. Lewis, editors. Ordinary Differential Equations and Operators: A Tribute to F.V. Atkinson Proceedings of a Symposium held at Dundee, Scotland March–July 1982, volume 1032 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1032. URL http:/

/link.springer.com/book/10.1007/BFb0076788;   http://
www.springerlink.com/content/978-3-540-38689-6.

Elbert:1982:ONT

[Elb82] Á. Elbert. Oscillation and nonoscillation theorems for some nonlinear ordinary differential equations. Lecture Notes in Mathematics, 964:187–212, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064999/.

Elencwajg:1982:BGC

[Ele82] G. Elencwajg. The Brauer groups in complex geometry. Lecture Notes in Mathematics, 917:222–230, 1982. CODEN LN-MAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092237/.

Elie:1982:TDP

[Éli82] Laure Élie. Sur le théorème de dichotomie pour les marches aléatoires sur les espaces homogènes. (French) []. Lecture Notes in Mathematics, 928:60–75, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093219/.

Eliahou:1984:IDC

[Eli84a] Shalom Eliahou. Idéaux de définition des courbes monomiales. (French) []. Lecture Notes in Mathematics, 1092:229–240, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099366/.

Elie:1984:NPA

[Éli84b] Laure Élie. Noyaux potentiels associés aux marches aléatoires sur les espaces homogènes quelques exemples clefs dont le groupe affine. (French) []. Lecture Notes in Mathematics, 1096:223– 260, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100113/.

Ellers:1980:GRC

[Ell80] E. W. Ellers. Generators and relations for classical groups. Lecture Notes in Mathematics, 792:40–45, 1980. CODEN LN-MAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088659/.

Elliott:1981:SID

[Ell81] Robert J. Elliott. Stochastic integration and discontinuous martingales. Lecture Notes in Mathematics, 851:72–84, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088723/.

Ellingham:1982:CCC

[Ell82] M. N. Ellingham. Constructing certain cubic graphs. Lecture Notes in Mathematics, 952:252–274, 1982. CODEN LN-MAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061983/.

Ellingham:1983:ACL

[Ell83] M. N. Ellingham. The asymptotic connectivity of labelled coloured regular bipartite graphs. Lecture Notes in Mathematics, 1036:177–188, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071518/.

Elmroth:1984:SHB

[Elm84] Tony Elmroth. The space-homogeneous Boltzmann equation for molecular forces of infinite range. Lecture Notes in Mathematics, 1048:192–201, 1984. CODEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071881/.

Elworthy:1981:SMD

[Elw81] K. David Elworthy. Stochastic methods and differential geometry. Lecture Notes in Mathematics, 901:95–110, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097192/.

Evans:1983:NSA

[ELZ83] W. D. Evans, R. T. Lewis, and A. Zettl. Non-self-adjoint operators and their essential spectra. Lecture Notes in Mathematics, 1032:123–160, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076796/.

Eklof:1983:ERSa

[EM83a] Paul C. Eklof and Alan H. Mekler. On endomorphism rings of ω_{1}- separable primary groups. Lecture Notes in Mathematics, 1006: 320–339, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21560-9_16/.

Eklof:1983:ERSb

[EM83b] Paul C. Eklof and Alan H. Mekler. On endomorphism rings of ω_{1}- separable primary groups. Lecture Notes in Mathematics, 1006: 320–339, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103711/.

Emch:1981:SMP

[Emc81] Gérard G. Emch. Some mathematical problems in nonequilibrium statistical mechanics. Lecture Notes in Mathematics, 843:264–295, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089278/.

Emery:1980:CPV

[Éme80a] Michel Émery. Compensation de processus V.F. non localement intégrables. (French) []. Lecture Notes in Mathematics, 784:152– 160, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089481/.

Emery:1980:EDS

[Éme80b] Michel Émery. Équations différentielles stochastiques: La méthode de Métivier et Pellaumail. (French) []. Lecture Notes in Mathematics, 784:118–124, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089476/.

Emery:1980:MQE

[Éme80c] Michel Émery. Metrisabilité de quelques espaces de processus aléatoires. (French) []. Lecture Notes in Mathematics, 784:140– 147, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089479/.

Emery:1981:TGJ

[Éme81] Michel Émery. Le théorème de Garnett-Jones, d’après Varopoulos. (French) []. Lecture Notes in Mathematics, 850:278–284, 1981.

CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088375/.

Emery:1982:MEM

[Éme82] Michel Émery. En marge de l’exposé de Meyer “Géométrie différentielle stochastique”. (French) []. Lecture Notes in Mathematics, 921:208–216, 1982. CODEN LNMAA2. ISBN 3-540-11486-6 (print), 3-540-39167-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092651/.

Emery:1983:NEP

[Éme83] Michel Émery. Note sur l’exposé precedent. (French) []. Lecture Notes in Mathematics, 986:185–186, 1983. CODEN LN-MAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068313.

Edgar:1982:COS

[EMS82] G. A. Edgar, A. Millet, and L. Sucheston. On compactness and optimality of stopping times. Lecture Notes in Mathematics, 939: 36–61, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096258/.

Everitt:1983:CAS

[EN83] W. N. Everitt and F. Neuman. A concept of adjointness and symmetry of differential expressions based on the generalised Lagrange identity and Green’s formula. Lecture Notes in Mathematics, 1032:161–169, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076797/.

Endell:1984:AUG

[End84] R. Endell. Zur algebraischen Unabhängigkeit gewisser Werte der Exponentialfunktion. (German) []. Lecture Notes in Mathematics, 1068:80–85, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099443/.

England:1982:SHI

[Eng82] Roland England. Some hybrid implicit stiffly stable methods for ordinary differential equations. Lecture Notes in Mathematics, 909:147–158, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092969/.

Engber:1983:NGT

[Eng83a] Michael Engber. On the Noether gap theorem. Lecture Notes in Mathematics, 971:24–30, 1983. CODEN LNMAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067070/.

Engler:1983:SPI

[Eng83b] Hans Engler. On some parabolic integro-differential equations: Existence and asymptotics of solutions. Lecture Notes in Mathematics, 1017:161–167, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103248/.

Engl:1984:BSN

[Eng84] Heinz W. Engl. Behaviour of solutions of nonlinear alternative problems under perturbations of the linear part with rank change. Lecture Notes in Mathematics, 1107:82–101, 1984. CO-DEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101494/.

Enright:1982:DEM

[Enr82] W. H. Enright. Developing effective multistep methods for the numerical solution of systems of second order initial value problems. Lecture Notes in Mathematics, 909:159–165, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092970/.

Enright:1983:URT

[Enr83] Thomas J. Enright. Unitary representations for two real forms of a semisimple Lie algebra; a theory of comparison. Lecture Notes in Mathematics, 1024:1–29, 1983. CODEN LNMAA2. ISBN 3-540-12725-9 (print), 3-540-38709-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071430/.

Enss:1983:CTB

[Ens83] Volker Enss. Completeness of three body quantum scattering. Lecture Notes in Mathematics, 1031:62–88, 1983. CODEN LN-MAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072111/.

Enright:1981:PCK

[EP81] T. J. Enright and R. Parthasarathy. A proof of a conjecture of Kashiware and Vergne. Lecture Notes in Mathematics, 880: 74–90, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090405/.

Edjvet:1984:CLG

[EP84] M. Edjvet and Stephen J. Pride. The concept of ‘largeness’ in group theory II. Lecture Notes in Mathematics, 1098:29–54, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099659/.

Erbe:1983:ONP

[Erb83] Lynn H. Erbe. Oscillation and nonoscillation properties for second order nonlinear differential equations. Lecture Notes in Mathematics, 1017:168–176, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103249/.

Erdos:1980:SCP

[Erd80] P. Erdős. Some combinational problems in geometry. Lecture Notes in Mathematics, 792:46–53, 1980. CODEN LN-MAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088660/.

Erdos:1981:SPR

[Erd81a] P. Erdős. Some problems and results on additive and multiplicative number theory. Lecture Notes in Mathematics, 899: 171–182, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096460/.

Erdos:1981:SNP

[Erd81b] Paul Erdős. Some new problems and results in graph theory and other branches of combinatorial mathematics. Lecture Notes in Mathematics, 885:9–17, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092251/.

Erdos:1982:SNP

[Erd82] Paul Erdős. Some new problems and results in number theory. Lecture Notes in Mathematics, 938:50–74, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097173/.

Ericsson:1983:IAS

[Eri83] Thomas Ericsson. Implementation and applications of the spectral transformation Lanczos algorithm. Lecture Notes in Mathematics, 973:177–188, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062101/.

Ericksen:1984:TVO

[Eri84a] J. L. Ericksen. A thermodynamic view of order parameters for liquid crystals. Lecture Notes in Mathematics, 1063:27–36, 1984. CODEN LNMAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072150/.

Ericksen:1984:OPP

[Eri84b] Jerald L. Ericksen, editor. Orienting Polymers: Proceedings of a Workshop held at the IMA, University of Minnesota, Minneapolis March 21–26, 1983, volume 1063 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1063. URL http:/

/link.springer.com/book/10.1007/BFb0072148;   http://
www.springerlink.com/content/978-3-540-38869-2.

Erkama:1983:DBD

[Erk83] Timo Erkama. On domains of bounded dilatation. Lecture Notes in Mathematics, 1013:68–75, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066518/.

Ermine:1980:IDS

[Erm80a] J.-L. Ermine. Images directes à supports propres dans le cas d’un morphisme fortement q-concave. (French) []. Lecture Notes in Mathematics, 807:66–101, 1980. CODEN LNMAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090875/.

Ermine:1980:CER

[Erm80b] Jean-Louis Ermine. Courants ergodiques et répartition geométrique. (French) []. Lecture Notes in Mathematics, 807:180–198, 1980. CODEN LNMAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090880/.

Erne:1981:CIE

[Ern81a] Marcel Erné. A completion- invariant extension of the concept of continuous lattices. Lecture Notes in Mathematics, 871:45– 60, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089903/.

Erne:1981:SCS

[Ern81b] Marcel Erné. Scott convergence and Scott topology in partially ordered sets II. Lecture Notes in Mathematics, 871:61–96, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089904/.

Ershov:1980:SST

[Ers80] M. P. Ershov. Some selection theorems for partitions of sets without topology. Lecture Notes in Mathematics, 794:105–111, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088217/.

Ershov:1982:GIT

[Ers82] M. P. Ershov. On a generalization of the Ionescu Tulcea construction of a measure by transition kernels. Lecture Notes in Mathematics, 945:29–33, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096662/.

Everitt:1981:OPD

[ES81] W. N. Everitt and B. D. Sleeman, editors. Ordinary and Partial Differential Equations: Proceedings of the Sixth Conference Held at Dundee, Scotland, March 31–April 4, 1980, volume 846 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 846. URL http:/

/link.springer.com/book/10.1007/BFb0089819;   http://
www.springerlink.com/content/978-3-540-38538-7.

Everitt:1982:OPD

[ES82] W. N. Everitt and B. D. Sleeman, editors. Ordinary and Partial Differential Equations: Proceedings of the Seventh Conference Held at Dundee, Scotland, March 29–April 2, 1982, volume 964 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 964. URL http:/

/link.springer.com/book/10.1007/BFb0064984;   http://
www.springerlink.com/content/978-3-540-39561-4.

Edelson:1983:ASA

[ES83] Allan L. Edelson and Jerry D. Schuur. Asymptotic and strong asymptotic equivalence to polynomials for solutions of nonlinear differential equations. Lecture Notes in Mathematics, 1017: 157–160, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103247/.

Esnault:1981:CVD

[Esn81] Hélène Esnault. Classification des variétés de dimension 3 et plus [d’après T. Fujita, S. Iitaka, Y. Kawamata, K. Ueno, E. Viehweg]. (French) []. Lecture Notes in Mathematics, 901:111–131, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097193/.

Estabrook:1980:DGT

[Est80] Frank B. Estabrook. Differential geometry as a tool for applied mathematicians. Lecture Notes in Mathematics, 810:1–22, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089971/.

Esterle:1983:ECC

[Est83a] J. Esterle. Elements for a classification of commutative radical Banach algebras. Lecture Notes in Mathematics, 975:4–65, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064546.

Esterle:1983:QRC

[Est83b] J. Esterle. Quasimultipliers, representations of H^{∞}, and the closed ideal problem for communtative Banach algebras. Lecture Notes in Mathematics, 975:66–162, 1983. CODEN LN-MAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064548/.

Esterle:1984:PMF

[Est84] Jean Esterle. Problème de Michael et fonctions entières de plusieurs variables complexes. (French) []. Lecture Notes in Mathematics, 1094:65–83, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099156/.

Edrei:1983:FGZ

[ESV83a] Albert Edrei, Edward B. Saff, and Richard S. Varga. £-functions of genus zero are admissible in the sense of Hayman. Lecture Notes in Mathematics, 1002:91–92, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070491.

Edrei:1983:CPS

[ESV83b] Albert Edrei, Edward B. Saff, and Richard S. Varga. The circular portion of the Szegő curve (proof of theorem 3). Lecture Notes in Mathematics, 1002:77–79, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070487/.

Edrei:1983:CE

[ESV83c] Albert Edrei, Edward B. Saff, and Richard S. Varga. Crude estimates for |U_{m}(w)| and |Q_{m}(w)|. Lecture Notes in Mathematics, 1002:63–69, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070483/.

Edrei:1983:DE

[ESV83d] Albert Edrei, Edward B. Saff, and Richard S. Varga. A differential equation. Lecture Notes in Mathematics, 1002:53–60, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070480/.

Edrei:1983:DON

[ESV83e] Albert Edrei, Edward B. Saff, and Richard S. Varga. Discussion of our numerical results. Lecture Notes in Mathematics, 1002: 26–40, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070475/.

Edrei:1983:E

[ESV83f] Albert Edrei, Edward B. Saff, and Richard S. Varga. Estimates for G_{m}(w) and Q_{m}(w). Lecture Notes in Mathematics, 1002: 49–52, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070479/.

Edrei:1983:ENC

[ESV83g] Albert Edrei, Edward B. Saff, and Richard S. Varga. Estimates for J_{m}(w) near the circumference |w| = 1. Lecture Notes in Mathematics, 1002:60–62, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070481/.

Edrei:1983:EUS

[ESV83h] Albert Edrei, Edward B. Saff, and Richard S. Varga. Existence and uniqueness of the Szegő curve. Lecture Notes in Mathematics, 1002:62–63, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070482.

Edrei:1983:FAF

[ESV83i] Albert Edrei, Edward B. Saff, and Richard S. Varga. The functions U_{m}(w), Q_{m}(w), G_{m}(w) associated with £-functions of genus zero. Lecture Notes in Mathematics, 1002:92–94, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070492/.

Edrei:1983:I

[ESV83j] Albert Edrei, Edward B. Saff, and Richard S. Varga. Introduction. Lecture Notes in Mathematics, 1002:1–7, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070473/.

Edrei:1983:NC

[ESV83k] Albert Edrei, Edward B. Saff, and Richard S. Varga. Notational conventions. Lecture Notes in Mathematics, 1002:43–44, 1983.

CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070477.

Edrei:1983:OM

[ESV83l] Albert Edrei, Edward B. Saff, and Richard S. Varga. Outline of the method. Lecture Notes in Mathematics, 1002:41–43, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070476/.

Edrei:1983:PTb

[ESV83m] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 1. Lecture Notes in Mathematics, 1002:70–72, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070485/.

Edrei:1983:PTc

[ESV83n] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 2. Lecture Notes in Mathematics, 1002:72–77, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070486/.

Edrei:1983:PTd

[ESV83o] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 4. Lecture Notes in Mathematics, 1002:80–82, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070488/.

Edrei:1983:PTa

[ESV83p] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 5. Lecture Notes in Mathematics, 1002:70, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070484.

Edrei:1983:PTe

[ESV83q] Albert Edrei, Edward B. Saff, and Richard S. Varga. Proof of theorem 6. Lecture Notes in Mathematics, 1002:82–87, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070489/.

Edrei:1983:PFP

[ESV83r] Albert Edrei, Edward B. Saff, and Richard S. Varga. Properties of £-functions; proof of assertion I of theorem 7. Lecture Notes in Mathematics, 1002:87–91, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070490/.

Edrei:1983:PML

[ESV83s] Albert Edrei, Edward B. Saff, and Richard S. Varga. Properties of the Mittag-Leffler function of order 1 < λ < ∞. Lecture Notes in Mathematics, 1002:44–49, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070478/.

Edrei:1983:SOR

[ESV83t] Albert Edrei, Edward B. Saff, and Richard S. Varga. Statements of our results. Lecture Notes in Mathematics, 1002:7–25, 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070474/.

Edrei:1983:ZSP

[ESV83u] Albert Edrei, Edward B. Saff, and Richard S. Varga. Zeros of Sections of Power Series, volume 1002 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12318-0 (print), 3-540-40040-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 115 pp. LCCN QA3 .L28 no. 1002.

URL http://link.springer.com/book/10.1007/BFb0070472;
 http://www.springerlink.com/content/978-3-540-40040-
0.

Etayo:1984:DTB

[Eta84] J. J. Etayo. Derivations in the tangent bundle. Lecture Notes in Mathematics, 1045:43–52, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072164/.

Esnault:1982:RCF

[EV82] Hélène Esnault and Eckart Viehweg. Revétements cycliques. (French) [Cyclic coatings]. Lecture Notes in Mathematics, 947: 241–250, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093593/.

Evans:1981:SSO

[Eva81] W. D. Evans. On the spectra of Schrödinger operators with a complex potential. Lecture Notes in Mathematics, 846:100–108, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089829/.

Evans:1983:PIM

[Eva83] D. J. Evans. Preconditioned iterative methods for the generalized eigenvalue problem. Lecture Notes in Mathematics, 973:189–194, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062102/.

Everitt:1980:LPS

[Eve80a] W. N. Everitt. Legendre polynomials and singular differential operators. Lecture Notes in Mathematics, 827:83–106, 1980. CO-DEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091375/.

Everitt:1980:OPD

[Eve80b] W. N. Everitt, editor. Ordinary and Partial Differential Equations: Proceedings of the Fifth Conference Held at Dundee, Scotland, March 29–31, 1978, volume 827 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 827. URL http:

//link.springer.com/book/10.1007/BFb0091371;   http:/
/www.springerlink.com/content/978-3-540-38346-8.

Everitt:1983:QII

[EW83] W. N. Everitt and S. D. Wray. On quadratic integral inequalities associated with second-order symmetric differential expressions. Lecture Notes in Mathematics, 1032:170–223, 1983. CODEN LN-MAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076798/.

Ewald:1980:ALS

[Ewa80] Günter Ewald. Über die algorithmische Lösung des Steinitzproblems einer inneren Kennzeichnung polytopaler Sphären. (German) []. Lecture Notes in Mathematics, 792:54–58, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088661/.

Fadlalla:1983:BVP

[Fad83] Adib A. Fadlalla. On a boundary value problem in pseudoconvex domains. Lecture Notes in Mathematics, 1039:168–176, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073364/.

Faith:1982:EMP

[Fai82a] Carl Faith. Embedding modules in projectives: a report on a problem. Lecture Notes in Mathematics, 951:21–40, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067322/.

Faith:1982:SSI

[Fai82b] Carl Faith. Subrings of self-injective and FPF rings. Lecture Notes in Mathematics, 951:12–20, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067321/.

Faltings:1983:CLS

[Fal83a] Gerd Faltings. On the cohomology of locally symmetric Hermitian spaces. Lecture Notes in Mathematics, 1029:55–98, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098927/.

Faltings:1983:PAI

[Fal83b] Prof. Dr. Gerd Faltings. Properties of Arakelov’s intersection product. Lecture Notes in Mathematics, 997:138–145, 1983. CO-DEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061641/.

Falkowski:1984:ALK

[Fal84] B.-J. Falkowski. An analogue of the Lévy–Khintchin formula on SL(2; C). Lecture Notes in Mathematics, 1064:80–85, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073635/.

Farah:1982:TUA

[Far82] J. L. Farah. Towards a unified approach to data smoothing. Lecture Notes in Mathematics, 909:59–72, 1982. CODEN LN-MAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092960/.

Farber:1984:KSH

[Far84] M. Š. Farber. Knots and stable homotopy. Lecture Notes in Mathematics, 1060:140–150, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099930/.

Fasel:1980:NSC

[Fas80] Hermann F. Fasel. Numerical solution of the complete Navier– Stokes equations for the simulation of unsteady flows. Lecture Notes in Mathematics, 771:177–195, 1980. CODEN LN-MAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086906/.

Fedorchuk:1984:HWT

[Fed84] V. V. Fedorchuk. On hypermaps, which are trivial bundles. Lecture Notes in Mathematics, 1060:26–36, 1984. CODEN LN-MAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099918/.

Feferman:1984:BCC

[Fef84] Solomon Feferman. Between constructive and classical mathematics. Lecture Notes in Mathematics, 1104:143–162, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099484/.

Feinsilver:1982:CRB

[Fei82] P. Feinsilver. Canonical representation of the Bernoulli process. Lecture Notes in Mathematics, 928:90–95, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093221/.

Feinsilver:1984:BSS

[Fei84] Philip Feinsilver. Bernoulli systems in several variables. Lecture Notes in Mathematics, 1064:86–98, 1984. CODEN LN-MAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073636/.

Fekete:1983:HCC

[Fek83] Otto Fekete. The Hardy class of certain functions with integral representations. Lecture Notes in Mathematics, 1013:233– 238, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066532/.

Felgner:1980:HTA

[Fel80] Ulrich Felgner. Horn-theories of abelian groups. Lecture Notes in Mathematics, 834:163–173, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090165/.

Ferrara:1980:SGT

[Fer80] S. Ferrara. Supergravity as the gauge theory of supersymmetry. Lecture Notes in Mathematics, 836:246–255, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089739/.

Fernique:1981:TKR

[Fer81a] X. Fernique. Sur le théorème de Kantorovitch–Rubinstein dans les espaces polonais. (French) []. Lecture Notes in Mathematics, 850:6–10, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088357/.

Fernique:1981:MMM

[Fer81b] Xavier Fernique. Mesures majorantes, majoration et continuité de fonctions aléatoires, exemples de construction. (French) []. Lecture Notes in Mathematics, 860:124–137, 1981. CODEN LN-MAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090612/.

Fernique:1981:LCI

[Fer81c] Xavier Fernique. Sur les lois de certaines intégrales associées à des mouvements Browniens. (French) []. Lecture Notes in Mathematics, 850:1–5, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088356/.

Ferry:1981:FDC

[Fer81d] Steve Ferry. Finitely dominated compacta need not have finite type. Lecture Notes in Mathematics, 870:1–5, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089703/.

Ferry:1981:SHA

[Fer81e] Steve Ferry. A simple-homotopy approach to the finiteness obstruction. Lecture Notes in Mathematics, 870:73–81, 1981. CO-DEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089708/.

Fernique:1983:RFA

[Fer83a] X. Fernique. Régularité de fonctions aléatoires non gaussiennes. (French) []. Lecture Notes in Mathematics, 976:2–74, 1983. CO-DEN LNMAA2. ISBN 3-540-11987-6 (print), 3-540-39458-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067985/.

Ferrand:1983:ICL

[Fer83b] Mme Jacqueline Ferrand. Un invariant conforme Lie aux geodesiques conformes. (French) []. Lecture Notes in Mathematics, 1013: 76–86, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066519/.

Ferreyra:1984:REA

[Fer84] Guillermo S. Ferreyra. The robust equation approach to multidimensional stochastic nonlinear filtering. Lecture Notes in Mathematics, 1080:46–60, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099784/.

Feyel:1980:EFP

[Fey80a] D. Feyel. Espaces fonctionnels de processus. (French) []. Lecture Notes in Mathematics, 814:80–84, 1980. CODEN LN-MAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094148/.

Feyel:1980:ICP

[Fey80b] D. Feyel. Inégalités de convexité pour espaces de Banach adaptés. (French) []. Lecture Notes in Mathematics, 814:85–89, 1980. CO-DEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094149/.

Feyel:1980:GMP

[Fey80c] D. Feyel. Sur le gradient mutuel des potentiels. (French) []. Lecture Notes in Mathematics, 814:90–96, 1980. CODEN LN-MAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094150/.

Feyel:1982:RRC

[Fey82a] D. Feyel. Remarques sur un résultat de Choquet. (French) []. Lecture Notes in Mathematics, 906:114–117, 1982. CODEN LN-MAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093264/.

Feyel:1982:ECN

[Fey82b] D. Feyel. Sur les extensions cylindriques de Noyaux. (French) []. Lecture Notes in Mathematics, 906:118–125, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093265/.

Feyel:1984:PAP

[Fey84a] D. Feyel. Potentiels adjoints et potentiels directs. (French) []. Lecture Notes in Mathematics, 1061:109–121, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099020/.

Feyel:1984:RCP

[Fey84b] D. Feyel. Remarques sur les cônes de potentiels. (French) []. Lecture Notes in Mathematics, 1061:122–133, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099021/.

Feyel:1984:QAT

[Fey84c] Denis Feyel. Quelques applications d’un théorème de Moschovakis à la théorie du potentiel. (French) []. Lecture Notes in Mathematics, 1096:280–289, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100115/.

Fadell:1981:FPTb

[FF81] Edward Fadell and Gilles Fournier, editors. Fixed Point Theory: Proceedings of a Conference Held at Sherbrooke, Québec, Canada, June 2–21, 1980, volume 886 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 886. URL http:

//link.springer.com/book/10.1007/BFb0092171;   http:/
/www.springerlink.com/content/978-3-540-38600-1.

Freese:1983:UAL

[FG83] Ralph S. Freese and Octavio C. Garcia, editors. Universal Algebra and Lattice Theory: Proceedings of the Fourth International Conference Held at Puebla, Mexico, 1982, volume 1004 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1004.

URL http://link.springer.com/book/10.1007/BFb0063426;
 http://www.springerlink.com/content/978-3-540-40954-
0.

Frigerio:1984:SMD

[FG84] Alberto Frigerio and Vittorio Gorini. On stationary Markov dilations of quantum dynamical semigroups (some remarks inspired by the workshop). Lecture Notes in Mathematics, 1055: 119–125, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071716/.

Fadell:1981:FPTa

[FH81] E. Fadell and S. Husseini. A fixed point theory for fiber-preserving maps. Lecture Notes in Mathematics, 886:49–72, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092176/.

Farrell:1984:NCC

[FH84] F. T. Farrell and W. C. Hsiang. On Novikov’s conjecture for cocompact discrete subgroups of a Lie group. Lecture Notes in Mathematics, 1051:38–48, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075563/.

Fiedorowicz:1982:EAT

[FHM82] Z. Fiedorowicz, H. Hauschild, and J. P. May. Equivariant algebraic K-theory. Lecture Notes in Mathematics, 967:23–80, 1982.

CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061898/.

Fuji-Hara:1980:ADR

[FHV80] R. Fuji-Hara and S. A. Vanstone. On automorphisms of doubly resolvable designs. Lecture Notes in Mathematics, 829:29–36, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088897/.

Fathi:1983:PPS

[FHY83] A. Fathi, M.-R. Herman, and J.-C. Yoccoz. A proof of Pesin’s stable manifold theorem. Lecture Notes in Mathematics, 1007: 177–215, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061417/.

Fichera:1980:SDP

[Fic80] Gaetano Fichera. Singularities of 3-dimensional potential functions at the vertices and at the edges of the boundary. Lecture Notes in Mathematics, 827:107–114, 1980. CODEN LN-MAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091376/.

Fiedler:1983:GHB

[Fie83a] Bernold Fiedler. Global Hopf bifurcation in porous catalysts. Lecture Notes in Mathematics, 1017:177–184, 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103250/.

Fiedler:1983:CPS

[Fie83b] Miroslav Fiedler. Combinatorial properties of sign-patterns in some classes of matrices. Lecture Notes in Mathematics, 1018: 28–32, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071608/.

Filter:1984:IAM

[Fil84] Wolfgang Filter. Isolated and antiisolated measures. Lecture Notes in Mathematics, 1089:1–10, 1984. CODEN LN-MAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072596/.

Finucan:1982:SDG

[Fin82] H. M. Finucan. Some decompositions of generalised Catalan numbers. Lecture Notes in Mathematics, 952:275–293, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061984/.

Fisher:1981:SIR

[Fis81] Joe W. Fisher. Semiprime ideals in rings with finite group actions revisited. Lecture Notes in Mathematics, 867:365–379, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090393/.

Fitzgibbon:1984:VPF

[Fit84] W. E. Fitzgibbon. A variation of parameters formula for Burgers system. Lecture Notes in Mathematics, 1076:78–85, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072768/.

Flensted-Jensen:1981:FJE

[FJ81] Mogens Flensted-Jensen. K-finite joint eigenfunctions of U(g)^{K} on a non-Riemannian semisimple symmetric space G/H. Lecture Notes in Mathematics, 880:91–101, 1981. CODEN LN-MAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090406/.

Flensted-Jensen:1984:HAS

[FJ84] Mogens Flensted-Jensen. Harmonic analysis on semisimple symmetric spaces a method of duality. Lecture Notes in Mathematics, 1077:166–209, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072339/.

Faierman:1981:NME

[FK81] M. Faierman and I. Knowles. Non-modal eigenfunction expansions. Lecture Notes in Mathematics, 846:109–118, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089830/.

Fink:1983:NTD

[FK83a] A. M. Fink and Takaŝi Kusano. Nonoscillation theorems for differential equations with general deviating arguments. Lecture Notes in Mathematics, 1032:224–239, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076799/.

Fulton:1983:SAO

[FK83b] Charles T. Fulton and Allan M. Krall. Self-adjoint 4-th order boundary value problem in the limit- 4 case. Lecture Notes in Mathematics, 1032:240–256, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076800/.

Fulton:1983:AEC

[FKM83] William Fulton, Steven Kleiman, and Robert MacPherson. About the enumeration of contacts. Lecture Notes in Mathematics, 997: 156–196, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061643/.

Ferus:1981:GDG

[FKSW81] Dirk Ferus, Wolfgang Kühnel, Udo Simon, and Bernd Wegner, editors. Global Differential Geometry and Global Analysis: Proceedings of the Colloquium Held at the Technical University of Berlin, November 21–24, 1979, volume 838 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 838. URL http:

//link.springer.com/book/10.1007/BFb0088835;   http:/
/www.springerlink.com/content/978-3-540-38419-9.

Fulton:1981:CAA

[FL81] William Fulton and Robert Lazarsfeld. Connectivity and its applications in algebraic geometry. Lecture Notes in Mathematics, 862:26–92, 1981. CODEN LNMAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090889/.

Fiorentini:1984:TTG

[FL84a] Mario Fiorentini and Alexandru T. Lascu. Two theorems of G. Gherardelli on curves simple intersection of three surfaces. Lecture Notes in Mathematics, 1056:132–141, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071772/.

Froberg:1984:CA

[FL84b] R. Fröberg and D. Laksov. Compressed algebras. Lecture Notes in Mathematics, 1092:121–151, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099360/.

Fleckinger:1981:ADE

[Fle81] J. Fleckinger. Asymptotic distribution of eigenvalues of elliptic operators on unbounded domains. Lecture Notes in Mathematics, 846:119–128, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089831/.

Fleckinger:1982:SVN

[Fle82a] J. Fleckinger. On the singular values of non-self-adjoint operators of Schrödinger type. Lecture Notes in Mathematics, 964: 251–262, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065002/.

Fletcher:1982:SOC

[Fle82b] R. Fletcher. Second order corrections for non-differentiable optimization. Lecture Notes in Mathematics, 912:85–114, 1982. CO-DEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093151/.

Fleury:1982:ANC

[Fle82c] Patrick J. Fleury, editor. Advances in Non-Commutative Ring Theory: Proceedings of the Twelfth George H. Hudson Symposium Held at Plattsburgh, USA, April 23–25, 1981, volume 951 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 951. URL http:/

/link.springer.com/book/10.1007/BFb0067319;   http://
www.springerlink.com/content/978-3-540-39371-9.

Flicker:1982:CS

[Fli82a] Yuval Z. Flicker. The continuous spectrum. Lecture Notes in Mathematics, 927:129–152, 1982. CODEN LNMAA2. ISBN 3-540-11500-5 (print), 3-540-39202-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094276/.

Flicker:1982:C

[Fli82b] Yuval Z. Flicker. The correspondence. Lecture Notes in Mathematics, 927:183–199, 1982. CODEN LNMAA2. ISBN 3-540-11500-5 (print), 3-540-39202-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094278/.

Flicker:1982:ET

[Fli82c] Yuval Z. Flicker. Equality of traces. Lecture Notes in Mathematics, 927:153–182, 1982. CODEN LNMAA2. ISBN 3-540-11500-5 (print), 3-540-39202-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094277/.

Flicker:1982:LT

[Fli82d] Yuval Z. Flicker. Local theory. Lecture Notes in Mathematics, 927:1–48, 1982. CODEN LNMAA2. ISBN 3-540-11500-5 (print), 3-540-39202-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094273/.

Flicker:1982:TF

[Fli82e] Yuval Z. Flicker. The trace formula. Lecture Notes in Mathematics, 927:49–99, 1982. CODEN LNMAA2. ISBN 3-540-11500-5 (print), 3-540-39202-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094274/.

Flicker:1982:TFB

[Fli82f] Yuval Z. Flicker. The Trace Formula and Base Change for GL(3), volume 927 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11500-5 (print), 3-540-39202-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 204 pp. LCCN QA3 .L28 no. 927. URL http:

//link.springer.com/book/10.1007/BFb0094272;   http:/
/www.springerlink.com/content/978-3-540-39202-6.

Flicker:1982:TTF

[Fli82g] Yuval Z. Flicker. The twisted trace formula. Lecture Notes in Mathematics, 927:100–128, 1982. CODEN LNMAA2. ISBN 3-540-11500-5 (print), 3-540-39202-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094275/.

Fujiwara:1981:OEC

[FLM81] Hidenori Fujiwara, Gérard Lion, and Bernard Magneron. Opérateurs d’entrelacement & calcul d’obstruction sur des groupes de Lie résolubles. (French) []. Lecture Notes in Mathematics, 880:102–137, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090407/.

Floret:1980:AST

[Flo80a] Klaus Floret. Applications of the sup-theorem. Lecture Notes in Mathematics, 801:82–97, 1980. CODEN LNMAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091491/.

Floret:1980:BAT

[Flo80b] Klaus Floret. Best approximations and the theorem of R. C. James. Lecture Notes in Mathematics, 801:57–66, 1980. CODEN LNMAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091489/.

Floret:1980:BSW

[Flo80c] Klaus Floret. Bounding sets in the weak topology. Lecture Notes in Mathematics, 801:21–27, 1980. CODEN LNMAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091486/.

Floret:1980:CCS

[Flo80d] Klaus Floret. Countably compact sets and the theorem of Eberlein–Grothendieck. Lecture Notes in Mathematics, 801:7– 20, 1980. CODEN LNMAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091485/.

Floret:1980:PWC

[Flo80e] Klaus Floret. Pointwise and weak compactness in spaces of continuous functions. Lecture Notes in Mathematics, 801:45–56, 1980. CODEN LNMAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091488/.

Floret:1980:PTR

[Flo80f] Klaus Floret. Proof of the theorem of R. C. James. Lecture Notes in Mathematics, 801:67–81, 1980. CODEN LNMAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091490/.

Floret:1980:SCA

[Flo80g] Klaus Floret. Sequential compactness and angelic spaces. Lecture Notes in Mathematics, 801:28–44, 1980. CODEN LN-MAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091487/.

Floret:1980:SFL

[Flo80h] Klaus Floret. Some fundamentals of locally convex spaces. Lecture Notes in Mathematics, 801:1–6, 1980. CODEN LNMAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091484/.

Floret:1980:TRR

[Flo80i] Klaus Floret. The topology related to Rainwater’s theorem. Lecture Notes in Mathematics, 801:98–116, 1980. CODEN LN-MAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091492/.

Floret:1980:WCS

[Flo80j] Klaus Floret, editor. Weakly Compact Sets: Lectures Held at S.U.N.Y., Buffalo, in Spring 1978, volume 801 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09991-3 (print), 3-540-39283-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 801.

URL http://link.springer.com/book/10.1007/BFb0091483;
 http://www.springerlink.com/content/978-3-540-39283-
5.

Fox:1983:TNV

[FM83] Geoffrey Fox and Pedro Morales. Théorèmes de Nikodým et de Vitali–Hahn–Saks pour les mesures à valeurs dans un semigroupe uniforme. (French) []. Lecture Notes in Mathematics, 1033:199– 208, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099855/.

Faddeev:1984:TGA

[FM84] Ludwig D. Faddeev and Arkadii A. Mal’cev, editors. Topology: General and Algebraic Topology, and Applications Proceedings of the International Topological Conference held in Leningrad, August 23–27, 1982, volume 1060 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1060. URL http:/

/link.springer.com/book/10.1007/BFb0099913;   http://
www.springerlink.com/content/978-3-540-38863-0.

Fernique:1983:EEP

[FMSW83] X. Fernique, P. W. Millar, D. W. Stroock, and M. Weber, editors. École d’Été de Probabilités de Saint-Flour XI — 1981. (French) [], volume 976 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-11987-6 (print), 3-540-39458-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 976. URL http:

//link.springer.com/book/10.1007/BFb0067984;   http:/
/www.springerlink.com/content/978-3-540-39458-7.

Foulser:1982:ITP

[FMW82] David Foulser, Geoffrey Mason, and Michael Walker. On irreducible translation planes in odd characteristic, I. Lecture Notes in Mathematics, 969:167–175, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062992/.

Fife:1982:SPA

[FN82] Paul C. Fife and Basil Nicolaenko. The singular perturbation approach to flame theory with chain and competing reactions. Lecture Notes in Mathematics, 964:232–250, 1982. CODEN LN-MAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065001/.

Flaherty:1982:SPB

[FO82] Joseph E. Flaherty and Robert E. O’Malley, Jr. Singularly perturbed boundary value problems for nonlinear systems, including a challenging problem for a nonlinear beam. Lecture Notes in Mathematics, 942:170–191, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094747/.

Foias:1980:SFD

[Foi80] C. Foiaş. A survey on the functional dynamical system generated by the Navier–Stokes equations. Lecture Notes in Mathematics, 771:196–202, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086907/.

Follmer:1981:CIS

[Föl81a] H. Föllmer. Calcul d’Itô sans probabilités. (French) [Itô’s calculations without probabilities]. Lecture Notes in Mathematics, 850: 143–150, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088364/.

Follmer:1981:DP

[Föl81b] H. Föllmer. Dirichlet processes. Lecture Notes in Mathematics, 851:476–478, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088738/.

Fomenko:1984:ASC

[Fom84a] A. T. Fomenko. Algebraic structure of certain integrable Hamiltonian systems. Lecture Notes in Mathematics, 1108:103–127, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099554/.

Fomenko:1984:CPE

[Fom84b] A. T. Fomenko. On certain properties of extremals in variational problems. Lecture Notes in Mathematics, 1108:209–217, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099559/.

Fontaine:1983:CRC

[Fon83] Jean-Marc Fontaine. Cohomologie de de Rham, cohomologie cristalline et représentations p-adiques. (French) []. Lecture Notes in Mathematics, 1016:86–108, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099959/.

Font:1984:MTP

[Fon84] Josep M. Font. Monadicity in topological pseudo-Boolean algebras. Lecture Notes in Mathematics, 1103:169–192, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099386/.

Foreman:1983:MSI

[For83a] Matthew Foreman. More saturated ideals. Lecture Notes in Mathematics, 1019:1–27, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071691/.

Forger:1983:NSM

[For83b] Michael Forger. Nonlinear sigma models on symmetric spaces. Lecture Notes in Mathematics, 1037:38–80, 1983. CODEN LN-MAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073169/.

Forst:1984:SD

[For84a] Gunnar Forst. Self-decomposability on R and Z. Lecture Notes in Mathematics, 1064:99–115, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073637/.

Forster:1984:CIA

[For84b] Otto Forster. Complete intersections in affine algebraic varieties and Stein spaces. Lecture Notes in Mathematics, 1092: 1–28, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/BFb0099355/;   http://link.springer.com/content/
pdf/bfm:978-3-540-39089-3/1.pdf.

Fossum:1980:TFC

[Fos80] Robert Fossum. Tensor functors of complexes (d’après Nielsen). Lecture Notes in Mathematics, 795:291–318, 1980. CODEN LN-MAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090124/.

Fossum:1981:ITR

[Fos81] Robert M. Fossum. Invariant theory, representation theory, commutative algebra — ménage à trois. Lecture Notes in Mathematics, 867:1–37, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090378/.

Fossum:1982:DR

[Fos82] Robert M. Fossum. Decompositions revisited. Lecture Notes in Mathematics, 924:260–295, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092936/.

Foulds:1981:CLI

[Fou81a] L. R. Foulds. Critical link identification in a network. Lecture Notes in Mathematics, 884:173–186, 1981. CODEN LN-MAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091818/.

Fournier:1981:SAF

[Fou81b] Gilles Fournier. A simplical approach to the fixed point index. Lecture Notes in Mathematics, 886:73–102, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092177/.

Foxby:1980:HDC

[Fox80] Hans-Bjørn Foxby. Homological dimensions of complexes of modules. Lecture Notes in Mathematics, 795:360–368, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090128/.

Friedlander:1981:ECR

[FP81] Eric M. Friedlander and Brian Parshall. Etale cohomology of reductive groups. Lecture Notes in Mathematics, 854:127–140, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089519/.

Flondor:1983:SRM

[FP83] Paul Flondor and Eugen Pascu. Some results on mixed manifolds. Lecture Notes in Mathematics, 1014:17–26, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072067/.

Foulds:1980:DAN

[FR80] L. R. Foulds and R. W. Robinson. Determining the asymptotic number of phylogenetic trees. Lecture Notes in Mathematics, 829: 110–126, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088905/.

Foulds:1981:EBP

[FR81] L. R. Foulds and R. W. Robinson. Enumeration of binary phylogenetic trees. Lecture Notes in Mathematics, 884:187–202, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091819/.

Fink:1983:SAT

[FR83] James P. Fink and Werner C. Rheinboldt. Some analytic techniques for parametrized nonlinear equations and their discretizations. Lecture Notes in Mathematics, 1005:108–119, 1983. CO-DEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112528/.

Franks:1980:SDH

[Fra80] John M. Franks. Symbolic dynamics, homology, and knots. Lecture Notes in Mathematics, 819:146–157, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086985/.

Franzen:1981:ACF

[Fra81] Berthold Franzen. Algebraic compactness of filter quotients. Lecture Notes in Mathematics, 874:228–241, 1981. CODEN LN-MAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090537/.

Francoise:1983:PGT

[Fra83a] Jean-Pierre Françoise. Propriétés de généricité des transformations canoniques. (French) []. Lecture Notes in Mathematics, 1007:

216–260, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061418/.

Frank:1983:DPP

[Fra83b] András Frank. Disjoint paths in the plane. Lecture Notes in Mathematics, 1018:33–37, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071609/.

Frank:1983:ACF

[Fra83c] Evelyn Frank. Applications of continued fractions in function theory. Lecture Notes in Mathematics, 1013:239–251, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066533/.

Frank:1983:FCR

[Fra83d] O. Frank. Faces and components of random lattices. Lecture Notes in Mathematics, 1018:38–45, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071610/.

Frankiewicz:1984:SRE

[Fra84] R. Frankiewicz. Some remarks on embeddings of Boolean algebras. Lecture Notes in Mathematics, 1089:64–68, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072601/.

Frei:1982:KES

[Fre82a] Armin Frei. Kan extensions and systems of imprimitivity. Lecture Notes in Mathematics, 962:62–68, 1982. CODEN LN-MAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066886/.

Frenkel:1982:RAL

[Fre82b] I. B. Frenkel. Representations of affine Lie algebras, Hecke modular forms and Korteweg–de Vries type equations. Lecture Notes in Mathematics, 933:71–110, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093354/.

Freese:1983:SIA

[Fre83] Ralph Freese. Subdirectly irreducible algebras in modular varieties. Lecture Notes in Mathematics, 1004:142–152, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063434/.

Fried:1980:EVH

[Fri80] David Fried. Efficiency vs. hyperbolicity on tori. Lecture Notes in Mathematics, 819:175–189, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086987/.

Friedman:1981:NRD

[Fri81] D. Friedman. Natural α-RE degrees. Lecture Notes in Mathematics, 859:63–66, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090939/.

Friedrich:1982:GIY

[Fri82] Thomas Friedrich. A geometric introduction to Yang–Millsequations. Lecture Notes in Mathematics, 926:74–84, 1982. CO-DEN LNMAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092657/.

Fried:1983:PPT

[Fri83] David Fried. Periodic points and twisted coefficients. Lecture Notes in Mathematics, 1007:261–293, 1983. CODEN LN-MAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061419/.

Frohlich:1981:HC

[Frö81] A. Fröhlich. The Hermitian classgroup. Lecture Notes in Mathematics, 882:191–206, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092493/.

FroeseFischer:1982:ASS

[Fro82a] Charlotte Froese Fischer. Approximate solution of Schrödinger’s equation for atoms. Lecture Notes in Mathematics, 968:71– 81, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064882/.

Frolicher:1982:SS

[Frö82b] Alfred Frölicher. Smooth structures. Lecture Notes in Mathematics, 962:69–81, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066887/.

Frolik:1982:MFU

[Fro82c] Z. Frolík. Measure-fine uniform spaces II. Lecture Notes in Mathematics, 945:34–41, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096663/.

Frolik:1984:PUS

[Fro84] Z. Frolík. Paracompactness of uniform spaces in the spirit of the Tamano’s theorem. Lecture Notes in Mathematics, 1060:37– 44, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099919/.

Frunza:1983:QSM

[Fru83] Monica Frunză. On the quasiconformality of some mappings in normed spaces. Lecture Notes in Mathematics, 1013:87–94, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066520/.

Feferman:1981:IDS

[FS81a] S. Feferman and W. Sieg. Inductive definitions and subsystems of analysis. Lecture Notes in Mathematics, 897:16–77, 1981. CODEN LNMAA2. ISBN 3-540-11170-0 (print), 3-540-38649-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091895/.

Feferman:1981:PTE

[FS81b] Solomon Feferman and Wilfried Sieg. Proof theoretic equivalences between classical and constructive theories for analysis. Lecture Notes in Mathematics, 897:78–142, 1981. CODEN LN-MAA2. ISBN 3-540-11170-0 (print), 3-540-38649-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091896/.

Fein:1981:BGR

[FS81c] B. Fein and M. Schacher. Brauer groups of rational function fields over global fields. Lecture Notes in Mathematics, 844:46–74, 1981. CODEN LNMAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090477/.

Fejer:1981:PCT

[FS81d] Peter A. Fejer and Robert I. Soare. The plus-cupping theorem for the recursively enumerable degrees. Lecture Notes in Mathematics, 859:49–62, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090938/.

Friedlander:1981:AEE

[FS81e] Eric M. Friedlander and Michael R. Stein, editors. Algebraic K-Theory Evanston 1980: Proceedings of the Conference Held at Northwestern University Evanston, March 24–27, 1980, volume 854 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 854. URL http:/

/link.springer.com/book/10.1007/BFb0089514;   http://
www.springerlink.com/content/978-3-540-38646-9.

Fainberg:1983:SMP

[FS83] E. A. Faĭnberg and I. M. Sonin. Stationary and Markov policies in countable state dynamic programming. Lecture Notes in Mathematics, 1021:111–129, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072908/.

Farkas:1984:SFR

[FS84] E. J. Farkas and M. E. Szabo. A star-finite relational semantics for parallel programs. Lecture Notes in Mathematics, 1104: 129–142, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099483/.

Falkner:1982:TAR

[FSY82] N. Falkner, C. Stricker, and M. Yor. Temps d’arrêt riches et applications. (French) []. Lecture Notes in Mathematics, 920:213– 218, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092784/.

Figa-Talamanca:1980:SPD

[FT80] Alessandro Figà-Talamanca. Singular positive definite functions. Lecture Notes in Mathematics, 781:22–29, 1980. CODEN LN-MAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097644/.

Fuchs:1981:RBW

[Fuc81a] L. Fuchs. Reinhold Baer’s work on abelian groups. Lecture Notes in Mathematics, 874:XV–XXI, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090519.

Fuchssteiner:1981:GHN

[Fuc81b] Benno Fuchssteiner. Generalized Hewitt–Nachbin spaces arising in state-space completions. Lecture Notes in Mathematics, 843: 296–318, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089279/.

Fuglede:1982:FBF

[Fug82] Bent Fuglede. Fonctions BLD et fonctions finement surharmoniques. (French) []. Lecture Notes in Mathematics, 906:126– 157, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093266/.

Fukushima:1981:RLM

[Fuk81] M. Fukushima. On a representation of local martingale additive functionals of symmetric diffusions. Lecture Notes in Mathematics, 851:110–118, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088725/.

Fukushima:1982:FAM

[Fuk82a] M. Fukushima, editor. Functional Analysis in Markov Processes: Proceedings of the International Workshop Held at Katata, Japan, August 21–26, 1981 and of the International Conference Held at Kyoto, Japan, August 27–29, 1981, volume 923 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 923.

URL http://link.springer.com/book/10.1007/BFb0093034;
 http://www.springerlink.com/content/978-3-540-39155-
5.

Fukushima:1982:ACM

[Fuk82b] M. Fukushima. On absolute continuity of multidimensional symmetrizable diffusions. Lecture Notes in Mathematics, 923:146– 176, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093040/.

Fukushima:1983:CMI

[Fuk83] Masatoshi Fukushima. Capacitary maximal inequalities and an ergodic theorem. Lecture Notes in Mathematics, 1021:130– 136, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072909/.

Fukushima:1984:DFW

[Fuk84] Masatoshi Fukushima. A Dirichlet form on the Wiener space and properties on Brownian motion. Lecture Notes in Mathematics, 1096:290–300, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100116/.

Fulton:1983:NC

[Ful83] William Fulton. On nodal curves. Lecture Notes in Mathematics, 997:146–155, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061642/.

Fumy:1981:LDM

[Fum81] Walter Fumy. The large with design — materialized. Lecture Notes in Mathematics, 893:189–194, 1981. CODEN LN-MAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091019/.

Furstenberg:1981:RCE

[Fur81] Harry Furstenberg. Rigidity and cocycles for ergodic actions of semi-simple Lie groups. Lecture Notes in Mathematics, 842:273– 292, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089940/.

Franks:1980:AAF

[FW80a] John Franks and Bob Williams. Anomalous Anosov flows. Lecture Notes in Mathematics, 819:158–174, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086986/.

Frostman:1980:RPM

[FW80b] Otto Frostman and Hans Wallin. Riesz potentials on manifolds. Lecture Notes in Mathematics, 787:106–120, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086328/.

Foerster:1982:MSS

[FW82a] Hartmut Foerster and Kristian Witsch. Multigrid software for the solution of elliptic problems on rectangular domains: MGOO (release 1). Lecture Notes in Mathematics, 960:427–460, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069937/.

Frank:1982:CSP

[FW82b] L. S. Frank and W. D. Wendt. Coercive singular perturbations: Reduction and convergence. Lecture Notes in Mathematics, 942: 54–64, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094739/.

Franks:1981:KSD

[FY81] John Franks and Lai-Sang Young. A C^{2} Kupka–Smale diffeomorphism of the disk with no sources or sinks. Lecture Notes in Mathematics, 898:90–98, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091909/.

Flum:1980:P

[FZ80a] Jörg Flum and Martin Ziegler. Part I. Lecture Notes in Mathematics, 769:1–77, 1980. CODEN LNMAA2. ISBN 3-540-09732-5 (print), 3-540-38544-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097007/.

Flum:1980:PI

[FZ80b] Jörg Flum and Martin Ziegler. Part II. Lecture Notes in Mathematics, 769:78–147, 1980. CODEN LNMAA2. ISBN 3-540-09732-5 (print), 3-540-38544-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097008/.

Flum:1980:TMT

[FZ80c] Jörg Flum and Martin Ziegler. Topological Model Theory, volume 769 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 0-387-09732-5, 3-540-09732-5 (print), 3-540-38544-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 149 pp. LCCN QA3 .L28 no. 769. URL http:/

/link.springer.com/book/10.1007/BFb0097006;   http://
www.springerlink.com/content/978-3-540-38544-8.

Gabriel:1980:ARS

[Gab80] Peter Gabriel. Auslander-Reiten sequences and representationfinite algebras. Lecture Notes in Mathematics, 831:1–71, 1980. CODEN LNMAA2. ISBN 3-540-10263-9 (print), 3-540-38385-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089778/.

Gabber:1981:STA

[Gab81a] Ofer Gabber. Some theorems on Azumaya algebras. Lecture Notes in Mathematics, 844:129–209, 1981. CODEN LNMAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090480/.

Gabriel:1981:UCR

[Gab81b] P. Gabriel. The universal cover of a representation-finite algebra. Lecture Notes in Mathematics, 903:68–105, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092986/.

Gabriel:1981:AAI

[Gab81c] Pierre Gabriel. Algèbres auto-injectives de représentation finie (d’après Christine Riedtmann). (French) []. Lecture Notes in Mathematics, 842:20–39, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089926/.

Gafurov:1983:ERC

[Gaf83] M. U. Gafurov. On the estimate of the rate of convergence in the law of iterated logarithm. Lecture Notes in Mathematics, 1021:137–144, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072910/.

Galvao:1980:CSP

[Gal80] Carlos A. P. Galvão. Classical 1/2 spin particles interacting with gravitational fields: a supersymmetric model. Lecture Notes in Mathematics, 775:69–77, 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092021/.

Gallot:1981:MVR

[Gal81] Sylvestre Gallot. Minorations sur le λ_{1} des variétés riemanniennes. (French) []. Lecture Notes in Mathematics, 901:132–148, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097194/.

Gali:1982:OCS

[Gal82a] I. M. Gali. Optimal control of systems governed by elliptic operator of infinite order. Lecture Notes in Mathematics, 964:263– 271, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065003/.

Gallardo:1982:CMB

[Gal82b] Léonard Gallardo. Capacités, mouvement Brownien et problèmen de l’epine de Lebesgue sur les groupes de Lie nilpotents. (French) []. Lecture Notes in Mathematics, 928:96–120, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093222/.

Galligo:1982:ITG

[Gal82c] André Galligo. Invariants topologiques de germes d’applications stables et finies. (French) []. Lecture Notes in Mathematics, 961: 84–88, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071276/.

Galcuk:1983:PUS

[Gal83a] I. I. Galčuk. On the pathwise uniqueness of solutions of stochastic integral equations driven by martingales. Lecture Notes in Mathematics, 1021:145–152, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072911/.

Gallo:1983:DHC

[Gal83b] Daniel Gallo. A 3-dimensional hyperbolic collar lemma. Lecture Notes in Mathematics, 971:31–35, 1983. CODEN LN-MAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067071/.

Gallardo:1989:RAG

[Gal89] Léonard Gallardo. Rectification à l’article de L. Gallardo et O. Gebuhrer “Lois de probabilités infiniment divisibles sur les hypergroupes commutatifs discrets dénombrables (Probability Measures on Groups VII Lecture Notes 1064). (French) []. Lecture Notes in Mathematics, 1379:86–89, 1989. CODEN LNMAA2. ISBN 3-540-51401-5 (print), 3-540-46206-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087847/. See [GG84].

Gamelin:1980:JMS

[Gam80] T. W. Gamelin. Jensen measures, subharmonicity, and the conjugation operation. Lecture Notes in Mathematics, 781:30–42, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097645/.

Gangopadhyay:1981:CFB

[Gan81] T. Gangopadhyay. Characterization of forcibly bipartite selfcomplementary bipartitioned sequences. Lecture Notes in Mathematics, 885:237–260, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092267/.

Garnett:1980:TCB

[Gar80] John B. Garnett. Two constructions in BMO. Lecture Notes in Mathematics, 781:43–50, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097646/.

Garcia:1982:PSC

[Gar82a] Pedro L. García. Phase space of the coupled vectorial Klein– Gordon–Maxwell equations. Lecture Notes in Mathematics, 905: 133–149, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092433/.

Gardiner:1982:CDT

[Gar82b] A. Gardiner. Classifying distance-transitive graphs. Lecture Notes in Mathematics, 952:67–88, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061973/.

Gardner:1982:RBM

[Gar82c] R. J. Gardner. The regularity of Borel measures. Lecture Notes in Mathematics, 945:42–100, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096664/.

Garling:1982:SBS

[Gar82d] D. J. H. Garling. Stable Banach spaces, random measures and Orlicz function spaces. Lecture Notes in Mathematics, 928:121– 175, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093223/.

Gardiner:1983:TKM

[Gar83a] Frederick P. Gardiner. The Teichmüller–Kobayashi metric for infinite dimensional complex Teichmüller spaces. Lecture Notes in Mathematics, 971:48–67, 1983. CODEN LNMAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067073/.

Garnett:1983:SPF

[Gar83b] Lucy Garnett. Statistical properties of foliations. Lecture Notes in Mathematics, 1007:294–299, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061420/.

Gattazzo:1984:CVV

[Gat84] Remo Gattazzo. In characteristic p = 2 the Veronese variety V^{m} ⊂ P^{m(m+3)/2} and each of its generic projection is set-theoretic complete intersection. Lecture Notes in Mathematics, 1092:221– 228, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099365/.

Gauthier:1981:FPT

[Gau81] Gilles Gauthier. Fixed point theorems for approximative ANR’s. Lecture Notes in Mathematics, 886:103–115, 1981. CODEN LN-MAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092178/.

Gauthier:1983:CAU

[Gau83] Paul M. Gauthier. Carleman approximation on unbounded sets by harmonic functions with Newtonian singularities. Lecture Notes in Mathematics, 1039:177–181, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073365/.

Gaveau:1982:ICP

[Gav82] Bernard Gaveau. Intégrales de courbure et potentiels sur les hypersurfaces analytiques de C^{n}. (French) []. Lecture Notes in Mathematics, 919:108–122, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097045/.

Gaveau:1983:VFF

[Gav83] Bernard Gaveau. Valeurs frontières des fonctions harmoniques ou holomorphes et de leurs dérivées. (French) []. Lecture Notes in Mathematics, 1039:182–219, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073366/.

Gay:1980:DFA

[Gay80] Roger Gay. Division des fonctionnelles analytiques: applications aux fonctions entières de type exponentiel moyennes- périodiques. (French) []. Lecture Notes in Mathematics, 822:77–89, 1980. CO-DEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097750/.

Gay:1984:TRA

[Gay84] David M. Gay. A trust-region approach to linearly constrained optimization. Lecture Notes in Mathematics, 1066:72–105, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099519/.

Germain-Bonne:1984:CSA

[GB84] Bernard Germain-Bonne. Conditions suffisantes d’acceleration de la convergence. (French) []. Lecture Notes in Mathematics, 1071: 132–143, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099614/.

Gil:1982:NNF

[GBC82] Jose M. Amillo Gil, John A. Burns, and Eugene M. Cliff. Nonlinear neutral functional differential equations in product spaces. Lecture Notes in Mathematics, 964:118–134, 1982. CODEN LN-MAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064993/.

Gear:1982:ADT

[Gea82] C. W. Gear. Automatic detection and treatment of oscillatory and/or stiff ordinary differential equations. Lecture Notes in Mathematics, 968:190–206, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064888/.

Gebuhrer:1983:FHV

[Geb83] Marc-Olivier Gebuhrer. Sur les fonctions holomorphes à valeurs dans l’espace des martingales locales. (French) []. Lecture Notes in Mathematics, 986:123–124, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068307.

Gekeler:1984:AIB

[Gek84a] Eckart Gekeler. Approximation of initial boundary value problems. Lecture Notes in Mathematics, 1044:142–156, 1984. CODEN LNMAA2. ISBN 3-540-12880-8 (print), 3-540-38763-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071444/.

Gekeler:1984:DMM

[Gek84b] Eckart Gekeler. Direct multistep multiderivative methods for differential systems of second order. Lecture Notes in Mathematics, 1044:34–71, 1984. CODEN LNMAA2. ISBN 3-540-12880-8 (print), 3-540-38763-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071440/.

Gekeler:1984:DMS

[Gek84c] Eckart Gekeler. Discretization Methods for Stable Initial Value Problems, volume 1044 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12880-8 (print), 3-540-38763-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 201 pp. LCCN QA3 .L28 no. 1044.

URL http://link.springer.com/book/10.1007/BFb0071438;
 http://www.springerlink.com/content/978-3-540-38763-
3.

Gekeler:1984:LMMb

[Gek84d] Eckart Gekeler. Linear multistep methods and nonlinear differential systems of first order. Lecture Notes in Mathematics, 1044: 88–113, 1984. CODEN LNMAA2. ISBN 3-540-12880-8 (print), 3-540-38763-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071442/.

Gekeler:1984:LMMa

[Gek84e] Eckart Gekeler. Linear multistep methods and problems with leading matrix A(t) = a(t)A. Lecture Notes in Mathematics, 1044: 72–87, 1984. CODEN LNMAA2. ISBN 3-540-12880-8 (print), 3-540-38763-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071441/.

Gekeler:1984:MMM

[Gek84f] Eckart Gekeler. Multistep multiderivative methods for differential systems of first order. Lecture Notes in Mathematics, 1044: 1–33, 1984. CODEN LNMAA2. ISBN 3-540-12880-8 (print), 3-540-38763-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071439/.

Gekeler:1984:RKM

[Gek84g] Eckart Gekeler. Runge–Kutta methods for differential systems of first order. Lecture Notes in Mathematics, 1044:114–141, 1984. CODEN LNMAA2. ISBN 3-540-12880-8 (print), 3-540-38763-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071443/.

Gelfgren:1981:MPA

[Gel81a] J. K. Gelfgren. Multipoint Padé approximants converging to functions of Stieltjes’ type. Lecture Notes in Mathematics, 888:197– 207, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095586/.

Gelfond:1981:CTV

[Gel81b] Michael Gelfond. A class of theorems with valid constructive counterparts. Lecture Notes in Mathematics, 873:314–320, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090741/.

Gela:1983:AGM

[Gel83] Jarosław Gela. On the automorphism groups of mapping graphs. Lecture Notes in Mathematics, 1018:46–54, 1983. CODEN LN-MAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071611/.

Gentili:1983:DCC

[Gen83] Graziano Gentili. Distances on convex cones. Lecture Notes in Mathematics, 1022:1–31, 1983. CODEN LNMAA2. ISBN 3-540-12719-4 (print), 3-540-38702-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071600/.

Geoghegan:1981:FPF

[Geo81a] Ross Geoghegan. Fixed points in finitely dominated compacta: the geometric meaning of a conjecture of H. Bass. Lecture Notes in Mathematics, 870:6–22, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089704/.

Georg:1981:NID

[Geo81b] K. Georg. Numerical integration of the Davidenko equation. Lecture Notes in Mathematics, 878:128–161, 1981. CODEN LN-MAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090680/.

Georg:1981:NSU

[Geo81c] K. Georg. A numerically stable update for simplicial algorithms. Lecture Notes in Mathematics, 878:117–127, 1981. CODEN LN-MAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090679/.

Gerard:1980:MFS

[Gér80] Raymond Gérard. Meromorphic forms solutions of completely integrable Pfaffian systems with regular singularities. Lecture Notes in Mathematics, 810:99–117, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089976/.

Gerardin:1981:FAA

[Gér81a] Paul Gérardin. Formes automorphes associées àux cycles géodésiques des surfaces de Riemann hyperboliques. (French) [].

Lecture Notes in Mathematics, 901:23–35, 1981. CODEN LN-MAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097187/.

Gerardin:1981:IGL

[Gér81b] Paul Gérardin. Immeubles des groupes linéaires généraux. (French) []. Lecture Notes in Mathematics, 880:138–178, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090408/.

Gerl:1984:CFM

[Ger84] Peter Gerl. Continued fraction methods for random walks on N and on trees. Lecture Notes in Mathematics, 1064:131–146, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073639/.

Gutkowski:1984:KSI

[GF84] D. Gutkowski and M. V. Valdes Franco. Kolmogorovian statistical invariants for the Aspect–Rapisarda experiment. Lecture Notes in Mathematics, 1055:146–152, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071719/.

Gear:1982:AMH

[GG82] C. W. Gear and K. A. Gallivan. Automatic methods for highly oscillatory ordinary differential equations. Lecture Notes in Mathematics, 912:115–124, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093152/.

Gallardo:1984:LPI

[GG84] Léonard Gallardo and Olivier Gebuhrer. Lois de probabilité infiniment divisibles sur les hypergroupes commutatifs, discrets, denombrables. (French) []. Lecture Notes in Mathematics, 1064: 116–130, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/

BFb0073638/. See correction [Gal89].

Glass:1981:ETAb

[GGHJG81] A. M. W. Glass, Yuri Gurevich, W. Charles Holland, and Michèle Jambu-Giraudet. Elementary theory of automorphism groups of doubly homogeneous chains. Lecture Notes in Mathematics, 859: 67–82, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090940/.

Gourgeon:1982:CEF

[GH82] H. Gourgeon and J. P. Hennart. A class of exponentially fitted piecewise continuous methods for initial value problems. Lecture Notes in Mathematics, 909:200–207, 1982. CODEN LN-MAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092974/.

Gastineau-Hills:1983:KPSa

[GH83] H. M. Gastineau-Hills. Kronecker products of systems of orthogonal designs. Lecture Notes in Mathematics, 1036:189–205, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071519/.

Grifone:1984:CSC

[GH84] J. Grifone and E. Hassan. Connexions singulières et classe de Maslov. (French) []. Lecture Notes in Mathematics, 1045:71– 83, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072167/.

Gherardelli:1983:ITP

[Ghe83] Francesco Gherardelli, editor. Invariant Theory: Proceedings of the 1st 1982 Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) Held at Montecatini, Italy, June 10–18, 1982, volume 996 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12319-9 (print), 3-540-40043-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 996. URL http:/

/link.springer.com/book/10.1007/BFb0063233;   http://
www.springerlink.com/content/978-3-540-40043-1.

Gastineau-Hills:1983:KPSb

[GHH83] H. M. Gastineau-Hills and Joseph Hammer. Kronecker products of systems of higher dimensional orthogonal designs. Lecture Notes in Mathematics, 1036:206–216, 1983. CODEN LN-MAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071520/.

Ghisa:1980:MHM

[Ghi80] Dorin Ghişa. The modulus and the hyperbolic measure. Lecture Notes in Mathematics, 798:161–164, 1980. CODEN LN-MAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097262/.

Ghione:1983:PDT

[Ghi83] Franco Ghione. Un problème du type Brill–Noether pour les fibres vectoriels. (French) []. Lecture Notes in Mathematics, 997: 197–209, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061644/.

Grunewald:1980:GEE

[GHMN80] Fritz J. Grunewald, George Havas, J. L. Mennicke, and M. F. Newman. Groups of exponent eight. Lecture Notes in Mathematics, 806:49–188, 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091269/.

Ghoussoub:1982:MIF

[Gho82] N. A. Ghoussoub. Martingales of increasing functions. Lecture Notes in Mathematics, 939:62–72, 1982. CODEN LN-MAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096259/.

Gine:1983:LTR

[GHZ83] Evarist Giné, Marjorie G. Hahn, and Joel Zinn. Limit theorems for random sets: an application of probability in Banach space results. Lecture Notes in Mathematics, 990:112–135, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064267/.

Gierz:1982:SP

[Gie82a] Gerhard Gierz. C(X)-submodules of Γ(p). Lecture Notes in Mathematics, 955:80–85, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068873/.

Gierz:1982:ATS

[Gie82b] Gerhard Gierz. An adequate M-theory for ω-spaces. Lecture Notes in Mathematics, 955:154–158, 1982. CODEN LN-MAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068879/.

Gierz:1982:ADS

[Gie82c] Gerhard Gierz. An alternative description of spaces of sections: Function modules. Lecture Notes in Mathematics, 955: 44–59, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068870/.

Gierz:1982:BD

[Gie82d] Gerhard Gierz. Basic definitions. Lecture Notes in Mathematics, 955:8–21, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068866/.

Gierz:1982:BO

[Gie82e] Gerhard Gierz. Bundles of operators. Lecture Notes in Mathematics, 955:112–135, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068876/.

Gierz:1982:BTV

[Gie82f] Gerhard Gierz. Bundles of Topological Vector Spaces and Their Duality, volume 955 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 296 pp. LCCN QA3 .L28 no. 955. URL http:

//link.springer.com/book/10.1007/BFb0068863;   http:/
/www.springerlink.com/content/978-3-540-39437-2.

Gierz:1982:BLP

[Gie82g] Gerhard Gierz. Bundles with locally paracompact base spaces. Lecture Notes in Mathematics, 955:28–38, 1982. CODEN LN-MAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068868/.

Gierz:1982:CUB

[Gie82h] Gerhard Gierz. The closure of the ‘unit ball’ of a bundle and separation axioms. Lecture Notes in Mathematics, 955:183–199, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068881/.

Gierz:1982:D

[Gie82i] Gerhard Gierz. Duality. Lecture Notes in Mathematics, 955:159– 182, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068880/.

Gierz:1982:ECL

[Gie82j] Gerhard Gierz. Excursion: Continuous lattices and bundles. Lecture Notes in Mathematics, 955:136–143, 1982. CODEN LN-MAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068877/.

Gierz:1982:FBB

[Gie82k] Gerhard Gierz. Full bundles and bundles with completely regular base space. Lecture Notes in Mathematics, 955:22–27, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068867/.

Gierz:1982:I

[Gie82l] Gerhard Gierz. Introduction. Lecture Notes in Mathematics, 955: 1–6, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068864/.

Gierz:1982:LLI

[Gie82m] Gerhard Gierz. Local linear independence. Lecture Notes in Mathematics, 955:202–208, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068883/.

Gierz:1982:LTB

[Gie82n] Gerhard Gierz. Locally trivial bundles: a definition. Lecture Notes in Mathematics, 955:200–201, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068882.

Gierz:1982:SB

[Gie82o] Gerhard Gierz. M-structure and bundles. Lecture Notes in Mathematics, 955:144–153, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068878/.

Gierz:1982:MBB

[Gie82p] Gerhard Gierz. Morphisms between bundles. Lecture Notes in Mathematics, 955:95–111, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068875/.

Gierz:1982:NR

[Gie82q] Gerhard Gierz. Notational remarks. Lecture Notes in Mathematics, 955:7, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068865.

Gierz:1982:QBM

[Gie82r] Gerhard Gierz. Quotients of bundles and C(X)-modules. Lecture Notes in Mathematics, 955:86–94, 1982. CODEN LN-MAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068874/.

Gierz:1982:SAA

[Gie82s] Gerhard Gierz. Some algebraic aspects of ω-spaces. Lecture Notes in Mathematics, 955:60–61, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068871.

Gierz:1982:SWT

[Gie82t] Gerhard Gierz. Stone–Weierstraß theorems for bundles. Lecture Notes in Mathematics, 955:39–43, 1982. CODEN LN-MAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068869/.

Gierz:1982:TDS

[Gie82u] Gerhard Gierz. A third description of spaces of sections: C(X)- convex modules. Lecture Notes in Mathematics, 955:62–79, 1982. CODEN LNMAA2. ISBN 3-540-11610-9 (print), 3-540-39437-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068872/.

Gieseker:1983:CSS

[Gie83a] D. Gieseker. A construction of special space curves. Lecture Notes in Mathematics, 1008:51–60, 1983. CODEN LNMAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065698/.

Gieseker:1983:GIT

[Gie83b] D. Gieseker. Geometric invariant theory and applications to moduli problems. Lecture Notes in Mathematics, 996:45–73, 1983. CODEN LNMAA2. ISBN 3-540-12319-9 (print), 3-540-40043-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063235/.

Gilbert:1981:GHF

[Gil81a] R. P. Gilbert. Generalized hyperanalytic function theory. Lecture Notes in Mathematics, 846:129–147, 1981. CODEN LN-MAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089832/.

Gill:1981:GTR

[Gil81b] M. K. Gill. A graph theoretical recurrence formula for computing the characteristic polynomial of a matrix. Lecture Notes in Mathematics, 885:261–265, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092268/.

Gill:1981:NCA

[Gil81c] Mukhtiar Kaur Gill. A note concerning Acharya’s conjecture on a spectral measure of structural balance in a social system. Lecture Notes in Mathematics, 885:266–271, 1981. CODEN LN-MAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092269/.

Gillet:1981:CTS

[Gil81d] Henri Gillet. Comparison of K-theory spectral sequences, with applications. Lecture Notes in Mathematics, 854:141–167, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089520/.

Gill:1982:CAC

[Gil82a] John Gill. Convergence acceleration for continued fractions K(a_{n}/1) with lim a_{n} = 0. Lecture Notes in Mathematics, 932: 67–70, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093305/.

Gill:1982:TEA

[Gil82b] John Gill. Truncation error analysis for continued fractions K(a_{n}/1) where \sqrt{|a_{n}|} + \sqrt{|a_{n−1}|} < 1. Lecture Notes in Mathematics, 932:71–73, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093306/.

Gilligan:1983:HRH

[Gil83] Bruce Gilligan. Holomorphic reductions of homogeneous spaces. Lecture Notes in Mathematics, 1014:27–36, 1983. CODEN LN-MAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072068/.

Gine:1980:CDA

[Gin80] Evarist Giné. Corrections to domains of attraction in Banach spaces by Evarist Gine. Lecture Notes in Mathematics, 784: 17, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089465.

Gine:1981:CLT

[Gin81] Evarist Giné. Central limit theorems in Banach spaces: a survey. Lecture Notes in Mathematics, 860:138–152, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090613/.

Gindikin:1982:IGT

[Gin82] S. G. Gindikin. Integral geometry and twistors. Lecture Notes in Mathematics, 970:2–42, 1982. CODEN LNMAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066022/.

Gine:1983:CDP

[Gin83] Evarist Giné. A counterexample on domains of partial attraction in Banach spaces. Lecture Notes in Mathematics, 990:102– 111, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064266/.

Giovannitti:1983:EBGa

[Gio83a] Anthony Giovannitti. Extensions of Butler groups. Lecture Notes in Mathematics, 1006:164–170, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21560-9_5/.

Giovannitti:1983:EBGb

[Gio83b] Anthony Giovannitti. Extensions of Butler groups. Lecture Notes in Mathematics, 1006:164–170, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103700/.

Giraud:1981:SHBb

[Gir81a] J. Giraud. Surfaces de Hilbert–Blumenthal II. (French) []. Lecture Notes in Mathematics, 868:19–34, 1981. CODEN LN-MAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090644/.

Giraud:1981:SHBa

[Gir81b] J. Giraud. Surfaces d’Hilbert–Blumenthal. (French) [Hilbert– Blumenthal surfaces]. Lecture Notes in Mathematics, 868:1– 18, 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090643/.

Giraud:1981:SHBc

[Gir81c] J. Giraud. Surfaces d’Hilbert–Blumenthal III. (French) []. Lecture Notes in Mathematics, 868:35–57, 1981. CODEN LN-MAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090645/.

Giraud:1981:SAS

[GIR81d] Jean Giraud, Luc Illusie, and Michel Raynaud, editors. Surfaces Algébriques: Séminaire de Géométrie Algébrique d’Orsay 1976–78. (French) [], volume 868 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 868. URL http:

//link.springer.com/book/10.1007/BFb0090642;   http:/
/www.springerlink.com/content/978-3-540-38742-8.

Giry:1982:CAP

[Gir82] Michèle Giry. A categorical approach to probability theory. Lecture Notes in Mathematics, 915:68–85, 1982. CODEN LN-MAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092872/.

Giraud:1983:CJF

[Gir83a] J. Giraud. Condition de Jung four les revêtements radiciels de hauteur 1. (French) []. Lecture Notes in Mathematics, 1016:313– 333, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099969/.

Girko:1983:STN

[Gir83b] V. L. Girko. Spectral theory of nonself-adjoint random matrices. Lecture Notes in Mathematics, 1021:153–156, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072912/.

Giraud:1984:DLD

[Gir84a] Jean Giraud. Desingularization in low dimension. Lecture Notes in Mathematics, 1101:51–78, 1984. CODEN LNMAA2. ISBN 3-540-13904-4 (print), 3-540-39125-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072260/.

Girbau:1984:SED

[Gir84b] J. Girbau. Some examples of deformations of transversely holomorphic foliations. Lecture Notes in Mathematics, 1045:53– 62, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072165/.

Gisin:1984:MID

[Gis84] N. Gisin. A model of irreversible deterministic quantum dynamics. Lecture Notes in Mathematics, 1055:126–133, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071717/.

Gruson:1981:DCR

[GJ81] L. Gruson and C. U. Jensen. Dimensions cohomologiques reliées aux foncteurs \underline{lim (i)}. (French) []. Lecture Notes in Mathematics, 867:234–294, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090389/.

Gierz:1981:CIC

[GK81] Gerhard Gierz and Klaus Keimel. Continuous ideal completions and compactifications. Lecture Notes in Mathematics, 871:97– 124, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089905/.

Gilewicz:1981:PAI

[GL81] J. Gilewicz and E. Leopold. Padé approximant inequalities for the functions of the class S. Lecture Notes in Mathematics, 888: 208–219, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095587/.

Gaveau:1982:FHP

[GL82a] B. Gaveau and G. Laville. Fonctions holomorphes et particule chargée dans un champ magnétique uniforme. (French) []. Lecture Notes in Mathematics, 919:123–130, 1982. CODEN LN-MAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097046/.

Gaveau:1982:IDV

[GŁ82b] Bernard Gaveau and Julian Ławrynowicz. Intégrale de Dirichlet sur une variété complexe I. (French) []. Lecture Notes in Mathematics, 919:131–166, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097047/.

Gomez:1982:TMS

[GL82c] Susana Gomez and A. V. Levy. The tunnelling method for solving the constrained global optimization problem with several nonconnected feasible regions. Lecture Notes in Mathematics, 909: 34–47, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092958/.

Gutierrez:1982:HTB

[GL82d] J. A. Gutiérrez and H. E. Lacey. On the Hilbert transform for Banach space valued functions. Lecture Notes in Mathematics, 939: 73–80, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096260/.

Geymonat:1984:LTO

[GL84a] G. Geymonat and P. Leyland. The linear transport operator of fluid dynamics. Lecture Notes in Mathematics, 1047:127– 137, 1984. CODEN LNMAA2. ISBN 3-540-12893-X (print), 3-540-38773-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072328/.

Goldstein:1984:SCM

[GŁ84b] S. Goldstein and A. Łuczak. Sample continuity moduli theorem in von Neumann algebras. Lecture Notes in Mathematics, 1080: 61–68, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099785/.

Glass:1981:ETAa

[Gla81] A. M. W. Glass. Elementary types of automorphisms of linearly ordered sets — a survey. Lecture Notes in Mathematics, 848: 218–229, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090568/.

Gresho:1980:STD

[GLCS80] Philip M. Gresho, Robert L. Lee, Stevens T. Chan, and Robert L. Sani. Solution of the time-dependent incompressible Navier– Stokes and Boussinesq equations using the Galerkin finite element method. Lecture Notes in Mathematics, 771:203–222, 1980.

CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086908/.

Gliklikh:1984:RPT

[Gli84] Yu. E. Gliklikh. Riemannian parallel translation in non-linear mechanics. Lecture Notes in Mathematics, 1108:128–151, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099555/.

Gobel:1983:AGT

[GLM83] Rüdiger Göbel, Lee Lady, and Adolf Mader, editors. Abelian Group Theory: Proceedings of the Conference held at the University of Hawaii, Honolulu, USA, December 28, 1982–January 4, 1983, volume 1006 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1006. URL http://link.

springer.com/book/10.1007/978-3-662-21560-9;   http:/
/link.springer.com/book/10.1007/BFb0103695;   http://
www.springerlink.com/content/978-3-540-40051-6;   http:
//www.springerlink.com/content/978-3-662-21560-9.

Glonti:1983:TTT

[Glo83] O. A. Glonti. Transmission of ‘television type’ signals through a feedback channel. Lecture Notes in Mathematics, 1021:157– 166, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072913/.

Graham:1981:MLP

[GLP81] R. L. Graham, Wen-Ching Winnie Li, and J. L. Paul. Monochromatic lines in partitions of Z^{n}. Lecture Notes in Mathematics, 884:35–48, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091806/.

Gluck:1980:DBG

[Glu80] Herman Gluck. Dynamical behavior of geodesic fields. Lecture Notes in Mathematics, 819:190–215, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086988/.

Glynn:1983:TNS

[Gly83] David G. Glynn. Two new sequences of ovals in finite Desarguesian planes of even order. Lecture Notes in Mathematics, 1036: 217–229, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071521/.

Godsil:1980:GRN

[GM80a] C. D. Godsil and B. D. McKay. Graphs with regular neighbourhoods. Lecture Notes in Mathematics, 829:127–140, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088906/.

Grunewald:1980:FPG

[GM80b] F. J. Grunewald and J. Mennicke. Finiteness proofs for groups of exponent 8. Lecture Notes in Mathematics, 806:189–210, 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091270/.

Gumowski:1980:CSO

[GM80c] Igor Gumowski and Christian Mira. Conclusion and some open problems. Lecture Notes in Mathematics, 809:234–259, 1980. CO-DEN LNMAA2. ISBN 3-540-10017-2 (print), 3-540-38155-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089142/.

Gumowski:1980:ISP

[GM80d] Igor Gumowski and Christian Mira. Introduction and statement of the problem. Lecture Notes in Mathematics, 809:1–22, 1980. CODEN LNMAA2. ISBN 3-540-10017-2 (print), 3-540-38155-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089136/.

Gumowski:1980:RDD

[GM80e] Igor Gumowski and Christian Mira. Recurrences and Discrete Dynamic Systems, volume 809 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10017-2 (print), 3-540-38155-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 272 pp. LCCN QA3 .L28 no. 809.

URL http://link.springer.com/book/10.1007/BFb0089135;
 http://www.springerlink.com/content/978-3-540-38155-
6.

Gumowski:1980:SPF

[GM80f] Igor Gumowski and Christian Mira. Some properties of first order recurrences. Lecture Notes in Mathematics, 809:23–60, 1980. CODEN LNMAA2. ISBN 3-540-10017-2 (print), 3-540-38155-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089137/.

Gumowski:1980:SPS

[GM80g] Igor Gumowski and Christian Mira. Some properties of second order recurrences. Lecture Notes in Mathematics, 809:61–103, 1980. CODEN LNMAA2. ISBN 3-540-10017-2 (print), 3-540-38155-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089138/.

Gumowski:1980:SAC

[GM80h] Igor Gumowski and Christian Mira. Stochasticity in almost conservative recurrences. Lecture Notes in Mathematics, 809:184– 211, 1980. CODEN LNMAA2. ISBN 3-540-10017-2 (print), 3-540-38155-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089140/.

Gumowski:1980:SCR

[GM80i] Igor Gumowski and Christian Mira. Stochasticity in conservative recurrences. Lecture Notes in Mathematics, 809:104–183, 1980. CODEN LNMAA2. ISBN 3-540-10017-2 (print), 3-540-38155-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089139/.

Gumowski:1980:SSN

[GM80j] Igor Gumowski and Christian Mira. Stochasticity in strongly non-conservative recurrences. Lecture Notes in Mathematics, 809: 212–233, 1980. CODEN LNMAA2. ISBN 3-540-10017-2 (print), 3-540-38155-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089141/.

Graves-Morris:1981:ERR

[GM81] P. R. Graves-Morris. Efficient reliable rational interpolation. Lecture Notes in Mathematics, 888:28–63, 1981. CODEN LN-MAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095575/.

Gelfand:1982:VMS

[GM82a] Sergei Gelfand and Robert MacPherson. Verma modules and Schubert cells: a dictionary. Lecture Notes in Mathematics, 924: 1–50, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092924/.

Giusti:1982:SISa

[GM82b] M. Giusti and M. Merle. Singularités isolées et sections planes de variétés determinantielles. (French) []. Lecture Notes in Mathematics, 961:89–102, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071277/.

Giusti:1982:SISb

[GM82c] M. Giusti and M. Merle. Singularités isolées et sections planes de variétés determinantielles. (French) []. Lecture Notes in Mathematics, 961:103–118, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071278/.

Goresky:1982:TCA

[GM82d] M. Goresky and R. MacPherson. On the topology of complex algebraic maps. Lecture Notes in Mathematics, 961:119–129, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071279/.

Graf:1982:DMR

[GM82e] S. Graf and G. Mägerl. Disintegration of a measure with respect to a correspondence. Lecture Notes in Mathematics, 945: 167–169, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096671/.

Grigelionis:1982:WCR

[GM82f] B. Grigelionis and R. Mikulevičius. On weak convergence to random processes with boundary conditions. Lecture Notes in Mathematics, 972:260–275, 1982. CODEN LNMAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064865/.

Grasman:1983:SPE

[GM83a] J. Grasman and B. J. Matkowsky. Singular perturbations of epidemic models involving a threshold. Lecture Notes in Mathematics, 985:400–412, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062378/.

Grigelionis:1983:CWC

[GM83b] B. Grigelionis and R. Mikulevičius. On contiguity and weak convergence of probability measures. Lecture Notes in Mathematics, 1021:177–194, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072915/.

Germano:1984:PCS

[GM84a] G. Germano and S. Mazzanti. Partial closures and semantics of while: Towards an iteration-based theory of data types. Lecture Notes in Mathematics, 1104:163–174, 1984. CODEN LN-MAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099485/.

Gibert:1984:RSC

[GM84b] S. Gibert and A. Mukherjea. Results in semigroups in the context of non-homogeneous Markov chains: Tail idempotents and their structure for infinite-dimensional nonnegative matrices. Lecture Notes in Mathematics, 1064:147–160, 1984. CODEN LN-MAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073640/.

Graves-Morris:1984:GIV

[GMJ84] P. R. Graves-Morris and C. D. Jenkins. Generalised inverse vector valued rational interpolation. Lecture Notes in Mathematics, 1071:144–156, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099615/.

Graves-Morris:1984:MTV

[GMS84] P. R. Graves-Morris and E. B. Saff. A de Montessus theorem for vector valued rational interpolants. Lecture Notes in Mathematics, 1105:227–242, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072414/.

Graves-Morris:1984:RAI

[GMSV84] Peter Russell Graves-Morris, Edward B. Saff, and Richard S. Varga, editors. Rational Approximation and Interpolation: Proceedings of the United Kingdom–United States Conference held at Tampa, Florida, December 12–16, 1983, volume 1105 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LN-MAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1105; QA3 .L471 no.1105; QA3.L471. URL http:/

/link.springer.com/book/10.1007/BFb0072395;   http://
www.springerlink.com/content/978-3-540-39113-5.

Graf:1984:RH

[GMW84] S. Graf, R. D. Mauldin, and S. C. Williams. Random homeomorphisms. Lecture Notes in Mathematics, 1089:23–46, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072598/.

Gotay:1980:GCA

[GN80] Mark J. Gotay and James M. Nester. Generalized constraint algorithm and special presymplectic manifolds. Lecture Notes in Mathematics, 775:78–104, 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092022/.

Gammel:1982:NGJ

[GN82] J. L. Gammel and J. Nuttall. Note on generalized Jacobi polynomials. Lecture Notes in Mathematics, 925:258–270, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093514/.

Gnedenko:1983:LTR

[Gne83a] B. V. Gnedenko. On limit theorems for a random number of random variables. Lecture Notes in Mathematics, 1021:167– 176, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072914/.

Gnedenko:1983:SST

[Gne83b] B. V. Gnedenko. On some stability theorems. Lecture Notes in Mathematics, 982:24–31, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082058/.

Golub:1982:CTS

[GO82] Gene H. Golub and Michael L. Overton. Convergence of a twostage Richardson iterative procedure for solving systems of linear equations. Lecture Notes in Mathematics, 912:125–139, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093153/.

Gobel:1983:ERAa

[Göb83a] Rüdiger Göbel. Endomorphism rings of abelian groups. Lecture Notes in Mathematics, 1006:340–353, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21560-9_17/.

Gobel:1983:ERAb

[Göb83b] Rüdiger Göbel. Endomorphism rings of abelian groups. Lecture Notes in Mathematics, 1006:340–353, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103712/.

Godini:1983:BAI

[God83] G. Godini. Best approximation and intersections of balls. Lecture Notes in Mathematics, 991:44–54, 1983. CODEN LN-MAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061557/.

Gokturk:1983:RVQ

[Gök83] Zerrin Göktürk. Remarks on p-valent quasimeromorphic functions. Lecture Notes in Mathematics, 1013:95–99, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066521/.

Goldschlager:1980:AAC

[Gol80] Leslie M. Goldschlager. An approximation algorithm for computing the permanent. Lecture Notes in Mathematics, 829:141– 147, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088907/.

Goldie:1981:RRN

[Gol81a] A. W. Goldie. The reduced rank in Noetherian rings. Lecture Notes in Mathematics, 867:396–405, 1981. CODEN LN-MAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090395/.

Goldstein:1981:GDH

[Gol81b] Larry Joel Goldstein. On the generalized density hypothesis, I. Lecture Notes in Mathematics, 899:107–128, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096455/.

Goldstein:1982:AEE

[Gol82] Jerome A. Goldstein. Autocorrelation, equipartition of energy, and random evolutions. Lecture Notes in Mathematics, 928:176– 182, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093224/.

Goldstein:1983:SFN

[Gol83a] Norman Goldstein. Springer fibres with non-ample normal bundles. Lecture Notes in Mathematics, 1008:61–64, 1983. CODEN LNMAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065699/.

Golzani:1983:LSC

[Gol83b] Leonardo Golzani. Lipschitz spaces on compact rank one symmetric spaces. Lecture Notes in Mathematics, 992:139–160, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069156/.

Gonek:1981:ZHZ

[Gon81] Steven M. Gonek. The zeros of Hurwitz’s zeta-function on σ = 1/2. Lecture Notes in Mathematics, 899:129–140, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096456/.

Goodearl:1980:ESS

[Goo80] K. R. Goodearl. Extensions of simple by simple unit-regular rings. Lecture Notes in Mathematics, 825:42–58, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089120/.

Goodman:1981:RBP

[Goo81a] Nicolas D. Goodman. Reflections on Bishop’s philosophy of mathematics. Lecture Notes in Mathematics, 873:135–145, 1981. CO-DEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090732/.

Goodman:1981:GRS

[Goo81b] Victor Goodman. Growth rates for sums of i.i.d. Hilbert space valued random variables. Lecture Notes in Mathematics, 860:153– 175, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090614/.

Good:1983:ACF

[Goo83a] Anton Good. Analytic continuations and functional equations. Lecture Notes in Mathematics, 1040:86–97, 1983. CODEN LN-MAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073083/.

Good:1983:CSIa

[Goo83b] Anton Good. Computation of some integrals I. Lecture Notes in Mathematics, 1040:47–72, 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073080/.

Good:1983:CSIb

[Goo83c] Anton Good. Computation of some integrals II. Lecture Notes in Mathematics, 1040:80–86, 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073082/.

Good:1983:D

[Goo83d] Anton Good. Decompositions of G. Lecture Notes in Mathematics, 1040:17–26, 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073077/.

Good:1983:FCK

[Goo83e] Anton Good. Fourier coefficients and Kloosterman sums. Lecture Notes in Mathematics, 1040:39–46, 1983. CODEN LN-MAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073079/.

Good:1983:IRE

[Goo83f] Anton Good. Integral representations of eigenfunctions. Lecture Notes in Mathematics, 1040:26–39, 1983. CODEN LN-MAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073078/.

Good:1983:I

[Goo83g] Anton Good. Introduction. Lecture Notes in Mathematics, 1040: 1–13, 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073075/.

Good:1983:LAS

[Goo83h] Anton Good. Local Analysis of Selberg’s Trace Formula, volume 1040 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 128 pp. LCCN QA3 .L28 no. 1040. URL http:

//link.springer.com/book/10.1007/BFb0073074;   http:/
/www.springerlink.com/content/978-3-540-38698-8.

Good:1983:PST

[Goo83i] Anton Good. Poincaré series and their Fourier series expansions. Lecture Notes in Mathematics, 1040:73–80, 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073081/.

Good:1983:P

[Goo83j] Anton Good. Preliminaries. Lecture Notes in Mathematics, 1040: 14–17, 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073076/.

Good:1983:SFF

[Goo83k] Anton Good. Sum formulae (first form). Lecture Notes in Mathematics, 1040:98–112, 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073084/.

Good:1983:SFS

[Goo83l] Anton Good. Sum formulae (second form). Lecture Notes in Mathematics, 1040:113–122, 1983. CODEN LNMAA2. ISBN 3-540-12713-5 (print), 3-540-38698-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073085/.

Goodman:1983:DSA

[Goo83m] Sue Goodman. Dehn surgery on Anosov flows. Lecture Notes in Mathematics, 1007:300–307, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061421/.

Good:1984:AAM

[Goo84] A. Good. Analytical and arithmetical methods in the theory of Fuchsian groups. Lecture Notes in Mathematics, 1068:86– 103, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099444/.

Gorniewicz:1981:LCT

[Gór81] Lech Górniewicz. On the Lefschetz coincidence theorem. Lecture Notes in Mathematics, 886:116–139, 1981. CODEN LN-MAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092179/.

Gossez:1984:POS

[Gos84] J.-P. Gossez. On a property of Orlicz–Sobolev spaces. Lecture Notes in Mathematics, 1107:102–105, 1984. CODEN LN-MAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101495/.

Guyon:1981:IES

[GP81] X. Guyon and B. Prum. Identification et estimation de semimartingales representables par rapport à un brownien à un indice double. (French) []. Lecture Notes in Mathematics, 863:211– 232, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091102/.

Greub:1982:SS

[GP82] W. Greub and H.-R. Petry. Spinor structures. Lecture Notes in Mathematics, 905:170–185, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092436/.

Gallo:1983:KGR

[GP83a] Daniel M. Gallo and R. Michael Porter, editors. Kleinian Groups and Related Topics: Proceedings of the Workshop Held at Oaxtepec, Mexico, August 10–14, 1981, volume 971 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 971.

URL http://link.springer.com/book/10.1007/BFb0067066;
 http://www.springerlink.com/content/978-3-540-39426-
6.

Gallo:1983:PSO

[GP83b] Daniel M. Gallo and R. Michael Porter. Projective structures on open surfaces. Lecture Notes in Mathematics, 971:36–47, 1983. CODEN LNMAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067072/.

Gear:1983:DAS

[GP83c] C. W. Gear and L. R. Petzold. Differential/algebraic systems and matrix pencils. Lecture Notes in Mathematics, 973:75–89, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062095/.

Gear:1983:SIO

[GP83d] C. W. Gear and L. R. Petzold. Singular implicit ordinary differential equations and constraints. Lecture Notes in Mathematics, 1005:120–127, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112529/.

Gruson:1983:PCG

[GP83e] Laurent Gruson and Christian Peskine. Postulation des courbes gauches. (French) []. Lecture Notes in Mathematics, 997:218– 227, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061646/.

Gardner:1984:CIS

[GP84] R. J. Gardner and W. F. Pfeffer. Conditions that imply a space is Radon. Lecture Notes in Mathematics, 1089:11–22, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072597/.

Garcia:1980:Pa

[GPR80a] P. L. García and A. Pérez-Rendón. Preface. Lecture Notes in Mathematics, 836:265, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089741.

Garcia:1980:Pb

[GPR80b] P. L. García and A. Pérez-Rendón. Preface. Lecture Notes in Mathematics, 836:266–268, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089742/.

Garcia:1980:DGM

[GPRS80] P. L. García, A. Pérez-Rendón, and J. M. Souriau, editors. Differential Geometrical Methods in Mathematical Physics: Proceedings of the Conferences Held at Aix-en-Provence, September 3–7, 1979 and Salamanca, September 10–14, 1979, volume 836 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 836.

URL http://link.springer.com/book/10.1007/BFb0089722;
 http://www.springerlink.com/content/978-3-540-38405-

Gelbart:1983:AFF

[GPS83] Stephen Gelbart and Ilya Piatetski-Shapiro. Automorphic forms and L-functions for the unitary group. Lecture Notes in Mathematics, 1041:141–184, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073147/.

Goursaud:1982:TVK

[GPV82] J.-M. Goursaud, J.-L. Pascaud, and J. Valette. Sur les travaux de V. K. Kharchenko. (French) []. Lecture Notes in Mathematics, 924:322–355, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092939/.

Gong:1982:IMP

[GQ82] Guang Gong and Minping Qian. The invariant measures, probability flux and circulations of one-dimensional Markov processes. Lecture Notes in Mathematics, 923:188–198, 1982. CODEN LN-MAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093042/.

Gerard:1983:RCH

[GR83a] R. Gérard and J.-P. Ramis. Résidu d’une connexion holomorphe. (French) []. Lecture Notes in Mathematics, 1015:243–306, 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071354/.

Gerard:1983:EDS

[GR83b] Raymond Gérard and Jean-Pierre Ramis. Equations différentielles et systèmes de Pfaff dans le champ complexe — II: Séminaire, volume 1015 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 411 pp. LCCN QA3 .L28 no. 1015. URL http:

//link.springer.com/book/10.1007/BFb0071347;   http:/
/www.springerlink.com/content/978-3-540-38674-2.

Graf:1980:MWS

[Gra80] Siegfried Graf. Measurable weak selections. Lecture Notes in Mathematics, 794:117–140, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088219/.

Gray:1981:GWT

[Gra81a] Alfred Gray. A generalization of Weyl’s tube formula. Lecture Notes in Mathematics, 838:140–144, 1981. CODEN LN-MAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088850/.

Grayson:1981:DC

[Gra81b] Daniel R. Grayson. Dilogarithm computations for K_{3}. Lecture Notes in Mathematics, 854:168–178, 1981. CODEN LN-MAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089521/.

Granger:1982:SSH

[Gra82a] Jean-Michel Granger. Singularités des schémas de Hilbert ponctuels. (French) []. Lecture Notes in Mathematics, 961:130– 140, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071280/.

Gray:1982:VCP

[Gra82b] Alfred Gray. Vector cross products, harmonic maps and the Cauchy Riemann equations. Lecture Notes in Mathematics, 949: 57–74, 1982. CODEN LNMAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069756/.

Gray:1982:ATT

[Gra82c] John W. Gray. Asscoli’s theorem for topological categories. Lecture Notes in Mathematics, 915:86–104, 1982. CODEN LN-MAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092873/.

Gray:1982:EAS

[Gra82d] John W. Gray. Enriched algebras, spectra and homotopy limits. Lecture Notes in Mathematics, 962:82–99, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066888/.

Grayson:1982:FGG

[Gra82e] Daniel R. Grayson. Finite generation of K- groups of a curve over a finite field. Lecture Notes in Mathematics, 966:69–90, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062167/.

Grabiner:1983:WCA

[Gra83a] Sandy Grabiner. Weighted convolution algebras as analogues of Banach algebras of power series. Lecture Notes in Mathematics, 975:282–289, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064559/.

Graham:1983:DS

[Gra83b] George E. Graham. Differentiable semigroups. Lecture Notes in Mathematics, 998:57–127, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062028/.

Grasman:1983:AMM

[Gra83c] J. Grasman. Asymptotic methods in mathematical biology. Lecture Notes in Mathematics, 985:76–95, 1983. CODEN LN-MAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062363/.

Grasman:1983:IPB

[Gra83d] J. Grasman. Introduction to the paper by A. S. Bakaj. Lecture Notes in Mathematics, 985:96–98, 1983. CODEN LN-MAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062364/.

Graffi:1984:CLC

[Gra84a] D. Graffi. On the contributions of Lamberto Cesari to applied mathematics. Lecture Notes in Mathematics, 1107:1–18, 1984. CODEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101490/.

Gray:1984:DOP

[Gra84b] Brayton Gray. Desuspension at an odd prime. Lecture Notes in Mathematics, 1051:360–370, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075577/.

Green:1980:RPR

[Gre80a] Edward L. Green. Remarks on projective resolutions. Lecture Notes in Mathematics, 832:259–279, 1980. CODEN LN-MAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088468/.

Green:1980:CLM

[Gre80b] James A. Green. The Carter–Lusztig modules V_{λ,K}. Lecture Notes in Mathematics, 830:65–79, 1980. CODEN LNMAA2. ISBN 3-540-10258-2 (print), 3-540-38379-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092301/.

Green:1980:I

[Gre80c] James A. Green. Introduction. Lecture Notes in Mathematics, 830:1–17, 1980. CODEN LNMAA2. ISBN 3-540-10258-2 (print), 3-540-38379-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092297/.

Green:1980:M

[Gre80d] James A. Green. The modules D_{λ}, K. Lecture Notes in Mathematics, 830:50–64, 1980. CODEN LNMAA2. ISBN 3-540-10258-2 (print), 3-540-38379-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092300/.

Green:1980:PR

[Gre80e] James A. Green. Polynomial Representations of GL_{n}, volume 830 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10258-2 (print), 3-540-38379-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 118 pp. LCCN QA3 .L28 no. 830. URL http:

//link.springer.com/book/10.1007/BFb0092296;   http:/
/www.springerlink.com/content/978-3-540-38379-6.

Green:1980:PRS

[Gre80f] James A. Green. Polynomial representations of GL_{n}(K): The Schur algebra. Lecture Notes in Mathematics, 830:18–35, 1980. CODEN LNMAA2. ISBN 3-540-10258-2 (print), 3-540-38379-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092298/.

Green:1980:RTS

[Gre80g] James A. Green. Representation theory of the symmetric group. Lecture Notes in Mathematics, 830:80–112, 1980. CODEN LN-MAA2. ISBN 3-540-10258-2 (print), 3-540-38379-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092302/.

Green:1980:WC

[Gre80h] James A. Green. Weights and characters. Lecture Notes in Mathematics, 830:36–49, 1980. CODEN LNMAA2. ISBN 3-540-10258-2 (print), 3-540-38379-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092299/.

Green:1981:GGA

[Gre81a] Edward L. Green. Group-graded algebras and the zero relation problem. Lecture Notes in Mathematics, 903:106–115, 1981. CO-DEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092987/.

Green:1981:PR

[Gre81b] James A. Green. Polynomial representations of GL_{n}. Lecture Notes in Mathematics, 848:124–140, 1981. CODEN LN-MAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090560/.

Greenleaf:1981:LCS

[Gre81c] Newcomb Greenleaf. Liberal constructive set theory. Lecture Notes in Mathematics, 873:213–240, 1981. CODEN LN-MAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090736/.

Greenlee:1982:SPN

[Gre82a] W. M. Greenlee. Singular perturbation of non-self-adjoint elliptic eigenvalue problems. Lecture Notes in Mathematics, 942: 43–53, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094738/.

Greub:1982:CCA

[Gre82b] W. H. Greub. The complex Cayley algebra and the Lorentz group. Lecture Notes in Mathematics, 905:186–195, 1982. CO-DEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092437/.

Greuel:1982:DCF

[Gre82c] Gert-Martin Greuel. On deformation of curves and a formula of Deligne. Lecture Notes in Mathematics, 961:141–168, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071281/.

Greve:1982:GCM

[Gre82d] Georg Greve. General construction of monoidal closed structures in topological, uniform and nearness spaces. Lecture Notes in Mathematics, 962:100–114, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066889/.

Greve:1982:RSM

[Gre82e] Georg Greve. Rigid spaces and monoidal closedness. Lecture Notes in Mathematics, 915:105–111, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092874/.

Greim:1983:ISS

[Gre83] Peter Greim. Isometries and L^{p}-structure of separably valued Bochner L^{p}-spaces. Lecture Notes in Mathematics, 1033:209– 218, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099856/.

Greiner:1984:TPF

[Gre84] G. Greiner. A typical Perron–Frobenius theorem with applications to an age-dependent population equation. Lecture Notes in Mathematics, 1076:86–100, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072769/.

Grigorjan:1980:OGN

[Gri80] Levon Derenikovič Grigorjan. On the order of growth for the norm of the holomorphic component of a meromorphic function. Lecture Notes in Mathematics, 798:165–168, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097263/.

Grigelionis:1982:SNL

[Gri82] B. Grigelionis. Stochastic non linear filtering equations and semimartingales. Lecture Notes in Mathematics, 972:63–99, 1982. CODEN LNMAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064860/.

Griffiths:1984:NAP

[Gri84] David F. Griffiths, editor. Numerical Analysis: Proceedings of the 10th Biennial Conference held at Dundee, Scotland, June 28–July 1, 1983, volume 1066 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1066; QA3.L28 no.1066, QA 297 D915n 1983.

URL http://link.springer.com/book/10.1007/BFb0099514;
 http://www.springerlink.com/content/978-3-540-38881-
4.

Gross:1980:A

[Gro80a] Benedict H. Gross. Acknowledgements. Lecture Notes in Mathematics, 776:2, 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096756.

Gross:1980:AEC

[Gro80b] Benedict H. Gross. Arithmetic on Elliptic Curves with Complex Multiplication, volume 776 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 95 pp. LCCN QA3 .L28 no. 776.

URL http://link.springer.com/book/10.1007/BFb0096754;
 http://www.springerlink.com/content/978-3-540-38575-
2.

Gross:1980:C

[Gro80c] Benedict H. Gross. A classification. Lecture Notes in Mathematics, 776:23–33, 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096759/.

Gross:1980:GA

[Gro80d] Benedict H. Gross. Global arithmetic. Lecture Notes in Mathematics, 776:45–66, 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096761/.

Gross:1980:I

[Gro80e] Benedict H. Gross. Introduction. Lecture Notes in Mathematics, 776:1, 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096755.

Gross:1980:LA

[Gro80f] Benedict H. Gross. Local arithmetic. Lecture Notes in Mathematics, 776:34–44, 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096760/.

Gross:1980:NC

[Gro80g] Benedict H. Gross. Notation and conventions. Lecture Notes in Mathematics, 776:3, 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096757.

Gross:1980:CP

[Gro80h] Benedict H. Gross. The Q-curve a(p). Lecture Notes in Mathematics, 776:67–86, 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096762/.

Gross:1980:TCM

[Gro80i] Benedict H. Gross. The theory of complex multiplication. Lecture Notes in Mathematics, 776:4–22, 1980. CODEN LNMAA2. ISBN 3-540-09743-0 (print), 3-540-38575-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096758/.

Gromov:1981:HMA

[Gro81a] Michael Gromov. Hyperbolic manifolds according to Thurston and Jørgensen. Lecture Notes in Mathematics, 842:40–53, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089927/.

Grosswald:1981:OSL

[Gro81b] Emil Grosswald. An orthonormal system and its Lebesgue constants. Lecture Notes in Mathematics, 899:1–9, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096451/.

Gross:1982:TSM

[Gro82] Leonard Gross. Thermodynamics, statistical mechanics and random fields. Lecture Notes in Mathematics, 929:101–204, 1982. CODEN LNMAA2. ISBN 3-540-11547-1 (print), 3-540-39220-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095619/.

Gronbaek:1983:WDC

[Grø83] Niels Grønbæk. Weighted discrete convolution algebras. Lecture Notes in Mathematics, 975:295–300, 1983. CODEN LN-MAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064561/.

Gruman:1980:GGE

[Gru80] Lawrence Gruman. La géométrie globale des ensembles analytiques dans C^{n}. (French) []. Lecture Notes in Mathematics, 822: 90–99, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097751/.

Grusa:1982:AVP

[Gru82a] Karl-Ulrich Grusa. Anwendung der verallgemeinerten partiellen Integration auf die Bilinearform B(g, s) = (λg, λs). (German) []. Lecture Notes in Mathematics, 916:46–53, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093897/.

Grusa:1982:ASG

[Gru82b] Karl-Ulrich Grusa. Anwendungsspektrum der Splines. (German) []. Lecture Notes in Mathematics, 916:3–11, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093889/.

Grusa:1982:SIRa

[Gru82c] Karl-Ulrich Grusa. Der L_{g}-Spline, der im Randgebiet harmonisch und im Rechteck biharmonisch ist. (German) []. Lecture Notes in Mathematics, 916:63–64, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093902.

Grusa:1982:SIRb

[Gru82d] Karl-Ulrich Grusa. Der L_{g}-Spline, der im Randgebiet holomorph und im Rechteck harmonisch ist. (German) []. Lecture Notes in Mathematics, 916:93–100, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093907/.

Grusa:1982:KRG

[Gru82e] Karl-Ulrich Grusa. Der Kern der Randoperatoren. (German) []. Lecture Notes in Mathematics, 916:34–38, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093894/.

Grusa:1982:SVG

[Gru82f] Karl-Ulrich Grusa. Die L_{g}-Splines und die Variationsrechnung. (German) []. Lecture Notes in Mathematics, 916:58–59, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093899.

Grusa:1982:AG

[Gru82g] Karl-Ulrich Grusa. Die Anwendungen. (German) []. Lecture Notes in Mathematics, 916:62, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093901.

Grusa:1982:CG

[Gru82h] Karl-Ulrich Grusa. Die Charakterisierungssätze. (German) []. Lecture Notes in Mathematics, 916:15–27, 1982. CODEN LN-MAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093891/.

Grusa:1982:ELR

[Gru82i] Karl-Ulrich Grusa. Die explizite Lösung des Randgebietes. (German) []. Lecture Notes in Mathematics, 916:79–80, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093904.

Grusa:1982:ELI

[Gru82j] Karl-Ulrich Grusa. Die explizite Lösung im Rechteck. (German) []. Lecture Notes in Mathematics, 916:92, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093906.

Grusa:1982:KIL

[Gru82k] Karl-Ulrich Grusa. Die Konstruktion der iterativen Lösung im Randgebiet. (German) []. Lecture Notes in Mathematics, 916: 65–78, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093903/.

Grusa:1982:VFP

[Gru82l] Karl-Ulrich Grusa. Die verallgemeinerte Form der partiellen Integration. (German) []. Lecture Notes in Mathematics, 916:38– 43, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093895/.

Grusa:1982:EG

[Gru82m] Karl-Ulrich Grusa. Einleitung. (German) []. Lecture Notes in Mathematics, 916:1–3, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093888/.

Grusa:1982:KLI

[Gru82n] Karl-Ulrich Grusa. Konstruktion der Lösung im Rechteck. (German) []. Lecture Notes in Mathematics, 916:80–91, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093905/.

Grusa:1982:LCG

[Gru82o] Karl-Ulrich Grusa. Lemmata zu den Charakterisierungssätzen. (German) []. Lecture Notes in Mathematics, 916:53–57, 1982.

CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093898/.

Grusa:1982:NGN

[Gru82p] Karl-Ulrich Grusa. Nomenklatur. (German) [Nomenclature]. Lecture Notes in Mathematics, 916:12–14, 1982. CODEN LN-MAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093890/.

Grusa:1982:SSB

[Gru82q] Karl-Ulrich Grusa. Spezielle L_{g}-Splines und die Blendingfunktionen. (German) []. Lecture Notes in Mathematics, 916:60–61, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093900.

Grusa:1982:TLG

[Gru82r] Karl-Ulrich Grusa. Technische Lemmata. (German) [Technical lemmas]. Lecture Notes in Mathematics, 916:43–46, 1982. CO-DEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093896/.

Grusa:1982:UBG

[Gru82s] Karl-Ulrich Grusa. Untersuchung der Bilinearform. (German) []. Lecture Notes in Mathematics, 916:28–31, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093892/.

Grusa:1982:VSG

[Gru82t] Karl-Ulrich Grusa. Verallgemeinerter Spektralsatz. (German) [Generalized spectral theorem]. Lecture Notes in Mathematics, 916:31–33, 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093893/.

Grusa:1982:ZIS

[Gru82u] Karl-Ulrich Grusa. Zweidimensionale, interpolierende L_{g}-Splines und ihre Anwendungen. (German) [Two-dimensional interpolating L_{g}-splines and their applications], volume 916 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11213-8 (print), 3-540-39049-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 238 pp. LCCN QA3 .L28 no. 916. URL http:/

/link.springer.com/book/10.1007/BFb0093887;   http://
www.springerlink.com/content/978-3-540-39049-7.

Gruman:1983:EEP

[Gru83] Lawrence Gruman. Ensembles exceptionnels pour les applications holomorphes dans C^{n}. (French) []. Lecture Notes in Mathematics, 1028:125–162, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071681/.

Gruman:1984:SDE

[Gru84] Lawrence Gruman. Solutions of difference equations with nonconstant coefficients. Lecture Notes in Mathematics, 1094:84– 138, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099157/.

Gobel:1981:GTS

[GRW81] R. Göbel, S. V. Richkov, and B. Wald. A general theory of slender groups and Fuchs-44-groups. Lecture Notes in Mathematics, 874:194–201, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090535/.

Grzegorek:1982:SFS

[Grz82] E. Grzegorek. Symmetric σ-fields of sets and universal null sets. Lecture Notes in Mathematics, 945:101–109, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096665/.

Grzegorek:1984:RSB

[Grz84] E. Grzegorek. Remarks on some Borel structures. Lecture Notes in Mathematics, 1089:69–74, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072602/.

Getoor:1982:SRE

[GS82a] R. K. Getoor and M. J. Sharpe. Some remarks on energy and duality. Lecture Notes in Mathematics, 923:177–187, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093041/.

Goyal:1982:CPS

[GS82b] Vinod B. Goyal and Philip W. Schaefer. Comparison principles for some fourth order elliptic problems. Lecture Notes in Mathematics, 964:272–279, 1982. CODEN LNMAA2. ISBN 3-540-11968-X (print), 3-540-39561-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065004/.

Guillemin:1982:MR

[GS82c] Victor Guillemin and Shlomo Sternberg. Moments and reductions. Lecture Notes in Mathematics, 905:52–65, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092427/.

Gundy:1982:PIR

[GS82d] Richard F. Gundy and Martin L. Silverstein. On a probabilistic interpretation for the Riesz transforms. Lecture Notes in Mathematics, 923:199–203, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093043/.

Sun:1983:PAG

[gS83a] Ji guang Sun. Perturbation analysis for the generalized eigenvalue and the generalized singular value problem. Lecture Notes in Mathematics, 973:221–244, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062105/.

Guillemin:1983:FRT

[GS83b] Victor Guillemin and Shlomo Sternberg. The Frobenius reciprocity theorem from a symplectic point of view. Lecture Notes in Mathematics, 1037:242–256, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073175/.

Gut:1983:AB

[GS83c] Allan Gut and Klaus D. Schmidt. Amarts — a bibliography. Lecture Notes in Mathematics, 1042:237–258, 1983. CODEN LNMAA2. ISBN 3-540-12867-0 (print), 3-540-38754-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073703/.

Gut:1983:ASF

[GS83d] Allan Gut and Klaus D. Schmidt. Amarts and Set Function Processes, volume 1042 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12867-0 (print), 3-540-38754-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 258 (est.) pp. LCCN QA3 .L28 no. 1042. URL http:/

/link.springer.com/book/10.1007/BFb0073700;   http://
www.springerlink.com/content/978-3-540-38754-1.

Gutierrez:1983:ATI

[GS83e] C. Gutiérrez and J. Sotomayor. An approximation theorem for immersions with stable configurations of lines of principal curvature. Lecture Notes in Mathematics, 1007:332–368, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061423/.

Greco:1984:CIL

[GS84] Silvio Greco and Rosario Strano, editors. Complete Intersections: Lectures given at the 1st 1983 Session of the Centro Internationale Matematico Estivo (C.I.M.E.) held at Acireale (Catania), Italy, June 13–21, 1983, volume 1092 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1092. URL http:/

/link.springer.com/book/10.1007/BFb0099354;   http://
www.springerlink.com/content/978-3-540-39089-3.

Godet-Thobie:1980:SRA

[GT80a] C. Godet-Thobie. Some results about multimeasures and their selectors. Lecture Notes in Mathematics, 794:112–116, 1980. CO-DEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088218/.

Gorgadze:1980:GOS

[GT80b] Z. G. Gorgadze and V. I. Tarieladze. On geometry of Orlicz spaces. Lecture Notes in Mathematics, 828:47–51, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097394/.

Grabiner:1983:CUS

[GT83] Sandy Grabiner and Marc P. Thomas. A class of unicellular shifts which contains non-strictly cyclic shifts. Lecture Notes in Mathematics, 975:273–276, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064556/.

Griewank:1984:NEP

[GT84] A. Griewank and Ph. L. Toint. Numerical experiments with partially separable optimization problems. Lecture Notes in Mathematics, 1066:203–220, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099526/.

Guckenheimer:1980:GTE

[Guc80] John Guckenheimer. The growth of topological entropy for one dimensional maps. Lecture Notes in Mathematics, 819:216– 223, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086989/.

Guckenheimer:1981:CTB

[Guc81] John Guckenheimer. On a codimension two bifurcation. Lecture Notes in Mathematics, 898:99–142, 1981. CODEN LN-MAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091910/.

Guessous:1980:PSE

[Gue80] Hassan Guessous. Prolongement surharmonioue dans un espace harmonique notion de flux. (French) []. Lecture Notes in Mathematics, 787:121–143, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086329/.

Guerra:1984:PQM

[Gue84] Francesco Guerra. Probability and quantum mechanics the conceptual foundations of stochastic mechanics. Lecture Notes in Mathematics, 1055:134–145, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071718/.

Guichardet:1980:RCS

[Gui80a] Alain Guichardet. Représentations de G^{X} (G compact) selon Verchik–Gelfand–Graiev et Ismagilov. (French) []. Lecture Notes in Mathematics, 770:303–311, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096248/.

Guillope:1980:RPV

[Gui80b] C. Guillopé. Remarques sur un problème de valeurs propres non linéaires faisant intervenir des fonctions non différentiables. (French) []. Lecture Notes in Mathematics, 782:130–159, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090431/.

Guimaraes:1980:CEB

[Gui80c] Luiz Carlos Guimarães. Contact equivalence and bifurcation theory. Lecture Notes in Mathematics, 799:140–151, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089313/.

Guivarch:1980:QPA

[Gui80d] Y. Guivarc’h. Quelques propriétés asymptotiques des produits de matrices aléatoires. (French) []. Lecture Notes in Mathematics, 774:177–250, 1980. CODEN LNMAA2. ISBN 3-540-09741-4 (print), 3-540-38567-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089624/.

Guichardet:1981:GRG

[Gui81] A. Guichardet. Sur les groupes EXT^{n} des représentations des groupes de Lie résolubles. (French) []. Lecture Notes in Mathematics, 880:179–196, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090409/.

Guichardet:1982:ILG

[Gui82a] Alain Guichardet. Introduction to Lie groups and their representations. Lecture Notes in Mathematics, 948:131–163, 1982. CODEN LNMAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069845/.

Guin:1982:GAF

[Gui82b] Daniel Guin. Sur le groupe K_{3} d’un anneau. (French) []. Lecture Notes in Mathematics, 967:81–100, 1982. CODEN LN-MAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061899/.

Guillemonat:1983:RSS

[Gui83] A. Guillemonat. Représentations sphériques singulières. (French) [Singular spherical representations]. Lecture Notes in Mathematics, 1020:35–72, 1983. CODEN LNMAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071496/.

Guivarch:1984:ECP

[Gui84a] Y. Guivarc’h. Exposants caractéristiques des produits de matrices aléatoires en dépendance Markovienne. (French) []. Lecture Notes in Mathematics, 1064:161–181, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073641/.

Guivarch:1984:ATL

[Gui84b] Yues Guivarc’h. Application d’un théorème limite local à la transience et à la récurrence de marches de Markov. (French) []. Lecture Notes in Mathematics, 1096:301–332, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100117/.

Gulisashvili:1983:EPI

[Gul83] A. Gulisashvili. Estimates for the Pettis integral in interpolation spaces with some applications. Lecture Notes in Mathematics, 991:55–78, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061558/.

Gundy:1980:IPM

[Gun80a] R. F. Gundy. Inégalités pour martingales à un et deux indices: L’espace H^{p}. (French) []. Lecture Notes in Mathematics, 774: 251–334, 1980. CODEN LNMAA2. ISBN 3-540-09741-4 (print), 3-540-38567-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089625/.

Gundy:1980:MFC

[Gun80b] R. F. Gundy. Maximal function characterization of H^{p} for the bidisc. Lecture Notes in Mathematics, 781:51–58, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097647/.

Gundy:1984:TLI

[Gun84] R. F. Gundy. Temps locaux et l’intégrale d’aire de Lusin. (French) []. Lecture Notes in Mathematics, 1059:77–81, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100033/.

Gupta:1984:EGT

[Gup84] Narain Gupta. Extending groups via tree automorphisms. Lecture Notes in Mathematics, 1098:55–57, 1984. CODEN LN-MAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099660/.

Gurarie:1980:HAB

[Gur80] David Gurarie. Harmonic analysis based on crossed product algebras and motion groups. Lecture Notes in Mathematics, 781: 59–75, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097648/.

Gurevich:1984:TLT

[Gur84] Yuri Gurevich. Toward logic tailored for computational complexity. Lecture Notes in Mathematics, 1104:175–216, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099486/.

Gustafson:1981:RHA

[Gus81] William H. Gustafson. Remarks on the history and applications of integral representations. Lecture Notes in Mathematics, 882: 1–36, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092486/.

Gustafson:1982:HAT

[Gus82] William H. Gustafson. The history of algebras and their representations. Lecture Notes in Mathematics, 944:1–28, 1982. CODEN LNMAA2. ISBN 3-540-11577-3 (print), 3-540-39313-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094056/.

Gustafsson:1983:SAI

[Gus83] Bertil Gustafsson. Stability analysis for initial-boundary value problems. Lecture Notes in Mathematics, 1005:128–141, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112530/.

Gut:1983:ITA

[Gut83a] Allan Gut. An introduction to the theory of asymptotic martingales. Lecture Notes in Mathematics, 1042:1–49, 1983. CODEN LNMAA2. ISBN 3-540-12867-0 (print), 3-540-38754-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073701/.

Gutierrez:1983:SCF

[Gut83b] Carlos Gutiérrez. Smoothability of Cherry flows on two-manifolds. Lecture Notes in Mathematics, 1007:308–331, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061422/.

Guzinski:1981:ACP

[Guz81] W. Guziński. Acceleration of convergence of power iterative process. Lecture Notes in Mathematics, 888:220–227, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095588/.

Greco:1983:CRS

[GV83] Silvio Greco and Angelo Vistoli. On the construction of rational surfaces with assigned singularities. Lecture Notes in Mathematics, 997:210–217, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061645/.

Gonzalo:1984:CEG

[GV84] J. Gonzalo and F. Varela. Sur certaines expressions globales d’une forme de contact. (French) []. Lecture Notes in Mathematics, 1045:63–70, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072166/.

Gerritzen:1980:ARA

[GvdP80a] Lothar Gerritzen and Marius van der Put. Analytic reductions of algebraic curves. Lecture Notes in Mathematics, 817:154– 180, 1980. CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089962/.

Gerritzen:1980:AMC

[GvdP80b] Lothar Gerritzen and Marius van der Put. Automorphisms of Mumford curves. Lecture Notes in Mathematics, 817:213– 232, 1980. CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089964/.

Gerritzen:1980:CDG

[GvdP80c] Lothar Gerritzen and Marius van der Put. The curve of a discontinuous group and its Jacobian variety. Lecture Notes in Mathematics, 817:233–256, 1980. CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089965/.

Gerritzen:1980:DG

[GvdP80d] Lothar Gerritzen and Marius van der Put. Discontinuous groups. Lecture Notes in Mathematics, 817:1–36, 1980. CODEN LN-MAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089958/.

Gerritzen:1980:EGQ

[GvdP80e] Lothar Gerritzen and Marius van der Put. Examples: Groups of quaternions; Whittaker groups. Lecture Notes in Mathematics, 817:257–286, 1980. CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089966/.

Gerritzen:1980:GMC

[GvdP80f] Lothar Gerritzen and Marius van der Put. The geometry of Mumford curves. Lecture Notes in Mathematics, 817:70–124, 1980.

CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089960/.

Gerritzen:1980:G

[GvdP80g] Lothar Gerritzen and Marius van der Put. The group PSL(2, F_{q}[t]). Lecture Notes in Mathematics, 817:287–309, 1980. CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089967/.

Gerritzen:1980:JV

[GvdP80h] Lothar Gerritzen and Marius van der Put. Jacobian varieties. Lecture Notes in Mathematics, 817:181–212, 1980. CODEN LN-MAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089963/.

Gerritzen:1980:MCA

[GvdP80i] Lothar Gerritzen and Marius van der Put. Mumford curves via automorphic forms. Lecture Notes in Mathematics, 817:37–69, 1980. CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089959/.

Gerritzen:1980:SGM

[GvdP80j] Lothar Gerritzen and Marius van der Put. Schottky Groups and Mumford Curves, volume 817 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 317 pp. LCCN QA3 .L28 no. 817.

URL http://link.springer.com/book/10.1007/BFb0089957;
 http://www.springerlink.com/content/978-3-540-38304-
8.

Gerritzen:1980:TSC

[GvdP80k] Lothar Gerritzen and Marius van der Put. Totally split curves and universal coverings. Lecture Notes in Mathematics, 817:125– 153, 1980. CODEN LNMAA2. ISBN 3-540-10229-9 (print), 3-540-38304-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089961/.

Gobel:1981:AGT

[GW81] Rüdiger Göbel and Elbert Walker, editors. Abelian Group Theory: Proceedings of the Oberwolfach Conference, January 12–17, 1981, volume 874 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 874. URL http:/

/link.springer.com/book/10.1007/BFb0090518;   http://
www.springerlink.com/content/978-3-540-38767-1.

Guin-Walery:1981:OET

[GWL81] Dominique Guin-Waléry and Jean-Louis Loday. Obstruction à l’excision en K-théorie algébrique. (French) []. Lecture Notes in Mathematics, 854:179–216, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089522/.

Gobel:1981:GIV

[GWW81] Rüdiger Göbel, Burkhard Wald, and Petra Westphal. Groups of integer-valuated functions. Lecture Notes in Mathematics, 874: 161–178, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090533/.

Gustafson:1982:CSD

[GY82] Karl E. Gustafson and David P. Young. Computation of solenoidal (divergence-free) vector fields. Lecture Notes in Mathematics, 968: 95–113, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064884/.

Habets:1980:SPE

[Hab80a] P. Habets. Singular perturbations of elliptic boundary value problems. Lecture Notes in Mathematics, 827:115–130, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091377/.

Haboush:1980:CDO

[Hab80b] W. J. Haboush. Central differential operators on split semisimple groups over fields of positive characteristic. Lecture Notes in Mathematics, 795:35–85, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090113/.

Hackbusch:1982:MGC

[Hac82a] W. Hackbusch. Multi-grid convergence theory. Lecture Notes in Mathematics, 960:177–219, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069929/.

Hackbusch:1982:MGS

[Hac82b] W. Hackbusch. Multi-grid solution of continuation problems. Lecture Notes in Mathematics, 953:20–45, 1982. CODEN LN-MAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069372/.

Hackbusch:1982:MGI

[Hac82c] W. Hackbusch. On multi-grid iterations with defect correction. Lecture Notes in Mathematics, 960:461–473, 1982. CODEN LN-MAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069938/.

Hadzic:1982:FPT

[Had82] Olga Hadžić. Fixed point theorems in not necessarily locally convex topological vector spaces. Lecture Notes in Mathematics, 948: 118–130, 1982. CODEN LNMAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069844/.

Hafner:1981:RSF

[Haf81] James Lee Hafner. On the representation of the summatory functions of a class of arithmetical functions. Lecture Notes in Mathematics, 899:148–165, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096458/.

Hagis:1981:SLP

[Hag81] Peter Hagis, Jr. On the second largest prime divisor of an odd perfect number. Lecture Notes in Mathematics, 899:254–263, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096466/.

Hajek:1984:NNP

[Háj84] Petr Hájek. On a new notion of partial conservativity. Lecture Notes in Mathematics, 1104:217–232, 1984. CODEN LN-MAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099487/.

Halder:1980:RPG

[Hal80a] Heinz-Richard Halder. Regular permutation geometries. Lecture Notes in Mathematics, 792:59–60, 1980. CODEN LN-MAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088662.

Hale:1980:SRR

[Hal80b] Jack K. Hale. Some recent results on dissipative processes. Lecture Notes in Mathematics, 799:152–172, 1980. CODEN LN-MAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089314/.

Hale:1981:SBP

[Hal81] J. K. Hale. Stability and bifurcation in a parabolic equation. Lecture Notes in Mathematics, 898:143–153, 1981. CODEN LN-MAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091911/.

Hale:1983:IDD

[Hal83] Jack K. Hale. Infinite dimensional dynamical systems. Lecture Notes in Mathematics, 1007:379–400, 1983. CODEN LN-MAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061425/.

Hale:1984:IDB

[Hal84] Jack K. Hale. Introduction to dynamic bifurcation. Lecture Notes in Mathematics, 1057:106–151, 1984. CODEN LNMAA2. ISBN 3-540-12931-6 (print), 3-540-38812-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098595/.

Hambleton:1982:PSO

[Ham82] Ian Hambleton. Projective surgery obstructions on closed manifolds. Lecture Notes in Mathematics, 967:101–131, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061900/.

Hanlon:1980:IAG

[Han80] Phil Hanlon. The incidence algebra of a group reduced partially ordered set. Lecture Notes in Mathematics, 829:148–156, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088908/.

Hanazawa:1981:VKA

[Han81] Masazumi Hanazawa. Various kinds of Aronszajn tree with no subtree of a different kind. Lecture Notes in Mathematics, 891: 1–21, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090975/.

Hansell:1984:MSR

[Han84] Roger W. Hansell. A measurable selection and representation theorem in non-separable spaces. Lecture Notes in Mathematics, 1089:86–94, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072605/.

Happel:1981:RIS

[Hap81] Dieter Happel. Relative invariants and subgeneric orbits of quivers of finite and tame type. Lecture Notes in Mathematics, 903:

116–124, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092988/.

Harel:1980:CLP

[Har80a] Michel Harel. Convergence en loi pour la topologie de Skorohod du processus empirique multidimensionnel normalisé tronqué et semi-corrigé (Étude au voisinage de la frontière inférieure de [0, 1]^{1+k}). (French) []. Lecture Notes in Mathematics, 821:46– 85, 1980. CODEN LNMAA2. ISBN 3-540-10239-6 (print), 3-540-38318-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097424/.

Hart:1980:IMS

[Har80b] R. Hart. Invertible 2 × 2 matrices over skew polynomial rings. Lecture Notes in Mathematics, 825:59–62, 1980. CODEN LN-MAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089121/.

Hartman:1980:DOR

[Har80c] Alan Hartman. Doubly and orthogonally resolvable quadruple systems. Lecture Notes in Mathematics, 829:157–164, 1980. CO-DEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088909/.

Haraux:1981:APQ

[Har81a] Alain Haraux. Almost-periodic quasi-autonomous dissipative systems in a Hilbert space. Lecture Notes in Mathematics, 841:295– 309, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089619/.

Haraux:1981:ADS

[Har81b] Alain Haraux. Autonomous dissipative systems. Lecture Notes in Mathematics, 841:200–240, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089615/.

Haraux:1981:BTQ

[Har81c] Alain Haraux. Boundedness of trajectories for quasi-autonomous dissipative systems. Lecture Notes in Mathematics, 841:284– 294, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089618/.

Haraux:1981:GRQ

[Har81d] Alain Haraux. General results for quasi-autonomous periodic systems. Lecture Notes in Mathematics, 841:241–265, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089616/.

Haraux:1981:GLT

[Har81e] Alain Haraux. Generalities and local theory. Lecture Notes in Mathematics, 841:1–18, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089607/.

Haraux:1981:GEP

[Har81f] Alain Haraux. The global existence problem. Lecture Notes in Mathematics, 841:19–38, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089608/.

Haraux:1981:LCH

[Har81g] Alain Haraux. The linear case: Hilbertian theory and applications. Lecture Notes in Mathematics, 841:148–163, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089612/.

Haraux:1981:MAB

[Har81h] Alain Haraux. More on asymptotic behavior for solutions of the nonlinear dissipative forced wave equation. Lecture Notes in Mathematics, 841:266–283, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089617/.

Haraux:1981:NEE

[Har81i] Alain Haraux. Nonlinear Evolution Equations — Global Behavior of Solutions, volume 841 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 313 pp. LCCN QA3 .L28 no. 841.

URL http://link.springer.com/book/10.1007/BFb0089606;
 http://www.springerlink.com/content/978-3-540-38534-
9.

Haraux:1981:SWE

[Har81j] Alain Haraux. Schrödinger and wave equations with a logarithmic nonlinearity. Lecture Notes in Mathematics, 841:118–147, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089611/.

Haraux:1981:SES

[Har81k] Alain Haraux. Smoothing effect for some nonlinear evolution equations. Lecture Notes in Mathematics, 841:96–117, 1981. CO-DEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089610/.

Haraux:1981:SNM

[Har81l] Alain Haraux. Some nonlinear monotone cases. Lecture Notes in Mathematics, 841:164–183, 1981. CODEN LNMAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089613/.

Haraux:1981:SNN

[Har81m] Alain Haraux. Some nonlinear, non monotone cases. Lecture Notes in Mathematics, 841:184–199, 1981. CODEN LN-MAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089614/.

Haraux:1981:TMO

[Har81n] Alain Haraux. Theory of monotone operators and applications. Lecture Notes in Mathematics, 841:39–95, 1981. CODEN LN-MAA2. ISBN 3-540-10563-8 (print), 3-540-38534-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089609/.

Harris:1981:SSG

[Har81o] B. J. Harris. Some spectral gap results. Lecture Notes in Mathematics, 846:148–157, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089833/.

Harrison:1981:WIG

[Har81p] J. Harrison. Wandering intervals. (German) []. Lecture Notes in Mathematics, 898:154–163, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091912/.

Hartman:1981:SDP

[Har81q] A. Hartman. A singular direct product for quadruple systems. Lecture Notes in Mathematics, 884:211–220, 1981. CODEN LN-MAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091821/.

Harary:1983:AGT

[Har83] Frank Harary. An achievement game on a toroidal board. Lecture Notes in Mathematics, 1018:55–59, 1983. CODEN LN-MAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071612/.

Harary:1984:GC

[Har84] Frank Harary. General connectivity. Lecture Notes in Mathematics, 1073:83–92, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073107/.

Hasumi:1983:CT

[Has83a] Morisuke Hasumi. Cauchy theorems. Lecture Notes in Mathematics, 1027:144–159, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071454/.

Hasumi:1983:CSP

[Has83b] Morisuke Hasumi. Characterization of surfaces of Parreau– Widom type. Lecture Notes in Mathematics, 1027:179–207, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071456/.

Hasumi:1983:CPR

[Has83c] Morisuke Hasumi. Classification of plane regions. Lecture Notes in Mathematics, 1027:234–261, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071458/.

Hasumi:1983:ESP

[Has83d] Morisuke Hasumi. Examples of surfaces of Parreau–Widom type. Lecture Notes in Mathematics, 1027:208–233, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071457/.

Hasumi:1983:GL

[Has83e] Morisuke Hasumi. Green lines. Lecture Notes in Mathematics, 1027:119–143, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071453/.

Hasumi:1983:HC

[Has83f] Morisuke Hasumi. Hardy classes. Lecture Notes in Mathematics, 1027:64–82, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071451/.

Hasumi:1983:HCI

[Has83g] Morisuke Hasumi. Hardy Classes on Infinitely Connected Riemann Surfaces, volume 1027 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 280 pp. LCCN QA3 .L28 no. 1027.

URL http://link.springer.com/book/10.1007/BFb0071447;
 http://www.springerlink.com/content/978-3-540-38719-
0.

Hasumi:1983:MC

[Has83h] Morisuke Hasumi. Martin compactification. Lecture Notes in Mathematics, 1027:39–63, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071450/.

Hasumi:1983:MAF

[Has83i] Morisuke Hasumi. Multiplicative analytic functions. Lecture Notes in Mathematics, 1027:23–38, 1983. CODEN LN-MAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071449/.

Hasumi:1983:RSP

[Has83j] Morisuke Hasumi. Riemann surfaces of Parreau–Widom type. Lecture Notes in Mathematics, 1027:83–118, 1983. CODEN LN-MAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071452/.

Hasumi:1983:SIS

[Has83k] Morisuke Hasumi. Shift-invariant subspaces. Lecture Notes in Mathematics, 1027:160–178, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071455/.

Hasumi:1983:TRS

[Has83l] Morisuke Hasumi. Theory of Riemann surfaces: a quick review. Lecture Notes in Mathematics, 1027:1–22, 1983. CODEN LNMAA2. ISBN 3-540-12729-1 (print), 3-540-38719-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071448/.

Hauschild:1980:AKA

[Hau80a] H. Hauschild. Äquivariante konfigurationsräume und Abbildungsräume. (German) []. Lecture Notes in Mathematics, 788: 281–315, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099250/.

Hausmann:1980:CTD

[Hau80b] Jean-Claude Hausmann. h-Cobordisms and twisted doubles. Lecture Notes in Mathematics, 788:39–48, 1980. CODEN LN-MAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099237/.

Hausen:1981:CDP

[Hau81] Jutta Hausen. A cardinal-determined projectivity condition for abelian groups and modules. Lecture Notes in Mathematics, 874: 109–113, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090529/.

Hausen:1983:AGWa

[Hau83a] Jutta Hausen. Abelian groups which are uniserial as modules over their endomorphism rings. Lecture Notes in Mathematics, 1006:204–208, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21560-9_8/.

Hausen:1983:AGWb

[Hau83b] Jutta Hausen. Abelian groups which are uniserial as modules over their endomorphism rings. Lecture Notes in Mathematics, 1006: 204–208, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103703/.

Havel:1983:EGU

[Hav83] Ivan Havel. Embedding graphs in undirected and directed cubes. Lecture Notes in Mathematics, 1018:60–68, 1983. CODEN LN-MAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071613/.

Hawkes:1984:SGA

[Haw84] John Hawkes. Some geometric aspects of potential theory. Lecture Notes in Mathematics, 1095:130–154, 1984. CODEN LN-MAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099126/.

Hayashi:1981:STT

[Hay81] Susumu Hayashi. On set theories in toposes. Lecture Notes in Mathematics, 891:23–29, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090976/.

Hayes:1982:CKL

[Hay82] J. G. Hayes. Curved knot lines and surfaces with ruled segments. Lecture Notes in Mathematics, 912:140–156, 1982. CODEN LN-MAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093154/.

Hayman:1983:VDF

[Hay83] W. K. Hayman. Value distribution of functions regular in the unit disk. Lecture Notes in Mathematics, 981:13–43, 1983. CODEN LNMAA2. ISBN 3-540-12003-3 (print), 3-540-39480-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066382/.

Hazewinkel:1981:LAV

[Haz81a] Michiel Hazewinkel. On Lie algebras of vectorfields, Lie algebras of differential operators and (nonlinear) filtering. Lecture Notes in Mathematics, 894:91–106, 1981. CODEN LNMAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096225/.

Hazewinkel:1981:RSG

[Haz81b] Michiel Hazewinkel. On representations of the symmetric groups, nilpotent matrices, systems, vectorbundles and Schubert cells.

Lecture Notes in Mathematics, 903:145–167, 1981. CODEN LN-MAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092990/.

Hazod:1982:SPL

[Haz82] Wilfried Hazod. Stable probabilities on locally compact groups. Lecture Notes in Mathematics, 928:183–211, 1982. CODEN LN-MAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093225/.

Hazewinkel:1983:LIR

[Haz83] Michiel Hazewinkel. Lectures on invariants, representations and Lie algebras in systems and control theory. Lecture Notes in Mathematics, 1029:1–36, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098925/.

Hazod:1984:RSS

[Haz84a] W. Hazod. Remarks on [semi-] stable probabilities. Lecture Notes in Mathematics, 1064:182–203, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073642/.

Hazod:1984:SSP

[Haz84b] W. Hazod. Stable and semistable probabilities on groups and on vector spaces. Lecture Notes in Mathematics, 1080:69–89, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099786/.

Heath-Brown:1984:CFV

[HB84] D. R. Heath-Brown. Cubic forms in 10 variables. Lecture Notes in Mathematics, 1068:104–108, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099445/.

Halley:1983:PRS

[HD83] J. W. Halley and C. Dasgupta. Percolation and related systems in equilibrium statistical mechanics. Lecture Notes in Mathematics, 1035:260–282, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073263/.

Huber-Dyson:1984:HCF

[HD84] Verena Huber-Dyson. Hnn-constructing finite groups. Lecture Notes in Mathematics, 1098:58–62, 1984. CODEN LN-MAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099661/.

Hebisch:1983:RQF

[Heb83] U. Hebisch. On right quotient filters of semigroups. Lecture Notes in Mathematics, 998:261–269, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062035/.

Hedberg:1980:SSS

[Hed80] Lars Inge Hedberg. Spectral synthesis and stability in Sobolev spaces. Lecture Notes in Mathematics, 779:73–103, 1980. CODEN LNMAA2. ISBN 3-540-09748-1 (print), 3-540-38602-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087668/.

Heinkel:1980:DEU

[Hei80] B. Heinkel. Deux exemples d’utilisation de mesures majorantes. (French) []. Lecture Notes in Mathematics, 784:1–16, 1980. CO-DEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089464/.

Heinkel:1983:MML

[Hei83] Bernard Heinkel. Majorizing measures and limit theorems for c_{o}-valued random variables. Lecture Notes in Mathematics, 990: 136–149, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064268/.

Heinkel:1984:NDS

[Hei84] Bernard Heinkel. The non i.i.d. strong law of large numbers in 2-uniformly smooth Banach spaces. Lecture Notes in Mathematics, 1080:90–118, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099787/.

Hejhal:1983:DTF

[Hej83a] Dennis A. Hejhal. Development of the trace formula (version A). Lecture Notes in Mathematics, 1001:1–232, 1983. CODEN LNMAA2. ISBN 3-540-12323-7 (print), 3-540-40914-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061303/.

Hejhal:1983:PSS

[Hej83b] Dennis A. Hejhal. Poincaré series and the spectral decomposition of L_{2}(ΓH, χ). Lecture Notes in Mathematics, 1001:233–266, 1983. CODEN LNMAA2. ISBN 3-540-12323-7 (print), 3-540-40914-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061304/.

Hejhal:1983:STF

[Hej83c] Dennis A. Hejhal. The Selberg Trace Formula for PSL(2, R): Volume 2, volume 1001 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12323-7 (print), 3-540-40914-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 806 pp. LCCN QA3 .L28 no. 1001. URL http:/

/link.springer.com/book/10.1007/BFb0061302;   http://
www.springerlink.com/content/978-3-540-40914-4.

Hejhal:1983:SA

[Hej83d] Dennis A. Hejhal. Selected applications. Lecture Notes in Mathematics, 1001:429–506, 1983. CODEN LNMAA2. ISBN 3-540-12323-7 (print), 3-540-40914-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061307/.

Hejhal:1983:SE

[Hej83e] Dennis A. Hejhal. Some examples. Lecture Notes in Mathematics, 1001:507–600, 1983. CODEN LNMAA2. ISBN 3-540-12323-7 (print), 3-540-40914-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061308/.

Hejhal:1983:VBS

[Hej83f] Dennis A. Hejhal. Version B of the Selberg trace formula. Lecture Notes in Mathematics, 1001:267–330, 1983. CODEN LN-MAA2. ISBN 3-540-12323-7 (print), 3-540-40914-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061305/.

Hejhal:1983:VCS

[Hej83g] Dennis A. Hejhal. Version C of the Selberg trace formula. Lecture Notes in Mathematics, 1001:331–428, 1983. CODEN LN-MAA2. ISBN 3-540-12323-7 (print), 3-540-40914-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061306/.

Hejtmanek:1984:TDL

[Hej84] J. Hejtmanek. Time-dependent linear transport theory. Lecture Notes in Mathematics, 1048:1–59, 1984. CODEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071877/;   http://link.
springer.com/content/pdf/bfm:978-3-540-38777-0/1.pdf.

Helffer:1980:PSP

[Hel80a] Bernard Helffer. Propagation des singularités pour des problèmes aux limites [d’après R. B. Melrose, J. Sjöstrand]. (French) []. Lecture Notes in Mathematics, 770:1–20, 1980. CODEN LN-MAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096232/.

Heller:1980:STM

[Hel80b] Alex Heller. Some topological methods in abstract group theory. Lecture Notes in Mathematics, 795:86–97, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090114/.

Helgason:1981:XRT

[Hel81] Sigurdur Helgason. The X-ray transform on a symmetric space. Lecture Notes in Mathematics, 838:145–148, 1981. CODEN LN-MAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088851/.

Helms:1983:HSM

[Hel83] L. L. Helms. Hyperfinite spin models. Lecture Notes in Mathematics, 983:15–26, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065333/.

Helffer:1984:PDE

[Hel84a] B. Helffer. Partial differential equations on nilpotent groups. Lecture Notes in Mathematics, 1077:210–253, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072340/.

Helgason:1984:WEH

[Hel84b] Sigurdur Helgason. Wave equations on homogeneous spaces. Lecture Notes in Mathematics, 1077:254–287, 1984. CODEN LN-MAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072341/.

Hemker:1982:AMD

[Hem82a] P. W. Hemker. An accurate method without directional bias for the numerical solution of a 2-D elliptic singular perturbation problem. Lecture Notes in Mathematics, 942:192–206, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094748/.

Hemker:1982:MDC

[Hem82b] P. W. Hemker. Mixed defect correction iteration for the accurate solution of the convection diffusion equation. Lecture Notes in Mathematics, 960:485–501, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069940/.

Hemker:1983:NAS

[Hem83] P. W. Hemker. Numerical aspects of singular perturbation problems. Lecture Notes in Mathematics, 985:267–287, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062371/.

Hemker:1984:MMP

[Hem84] P. W. Hemker. Multigrid methods for problems with a small parameter in the highest derivative. Lecture Notes in Mathematics, 1066:106–121, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099520/.

Henaut:1980:CCT

[Hén80a] Alain Hénaut. Cycles et cône tangent de Zariski. (French) []. Lecture Notes in Mathematics, 807:145–166, 1980. CODEN LN-MAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090878/.

Hendy:1980:EMP

[Hen80b] M. D. Hendy. Establishing the minimality of phylogenetic trees from protein sequences. Lecture Notes in Mathematics, 829:165– 171, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088910/.

Henkin:1980:ARC

[Hen80c] Gennadiĭ Markovič Henkin. Analytic representation for crfunctions on submanifolds of codimension 2 in cents^{n}. Lecture Notes in Mathematics, 798:169–191, 1980. CODEN LN-MAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097264/.

Hennequin:1980:EEP

[Hen80d] P. L. Hennequin. École d’Été de Probabilités de Saint-Flour VIII — 1978. (French) [], volume 774 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09741-4 (print), 3-540-38567-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 334 pp. LCCN QA3 .L28 no. 774.

URL http://link.springer.com/book/10.1007/BFb0089622;
 http://www.springerlink.com/content/978-3-540-38567-
7.

Hennequin:1981:EEP

[Hen81a] P. L. Hennequin. École d’Été de Probabilités de Saint-Flour IX — 1979. (French) [], volume 876 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10860-2 (print), 3-540-38774-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 280 pp. LCCN QA3 .L28 no. 876.

URL http://link.springer.com/book/10.1007/BFb0097497;
 http://www.springerlink.com/content/978-3-540-38774-
9.

Henry:1981:DSL

[Hen81b] Daniel Henry. Dynamical systems and Liapunov stability. Lecture Notes in Mathematics, 840:82–97, 1981. CODEN LN-MAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089652/.

Henry:1981:ENP

[Hen81c] Daniel Henry. Examples of nonlinear parabolic equations in physical, biological and engineering problems. Lecture Notes in Mathematics, 840:41–46, 1981. CODEN LNMAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089650/.

Henry:1981:EUC

[Hen81d] Daniel Henry. Existence, uniqueness and continuous dependence. Lecture Notes in Mathematics, 840:47–81, 1981. CODEN LN-MAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089651/.

Henry:1981:GTS

[Hen81e] Daniel Henry. Geometric Theory of Semilinear Parabolic Equations, volume 840 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 348 pp. LCCN QA3 .L28 no. 840. URL http:

//link.springer.com/book/10.1007/BFb0089647;   http:/
/www.springerlink.com/content/978-3-540-38528-8.

Henry:1981:I

[Hen81f] Daniel Henry. Introduction. Lecture Notes in Mathematics, 840: 1–2, 1981. CODEN LNMAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089648.

Henry:1981:IMN

[Hen81g] Daniel Henry. Invariant manifolds near an equilibrium point. Lecture Notes in Mathematics, 840:143–187, 1981. CODEN LN-MAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089654/.

Henry:1981:LNE

[Hen81h] Daniel Henry. Linear nonautonomous equations. Lecture Notes in Mathematics, 840:188–246, 1981. CODEN LNMAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089655/.

Henry:1981:NPS

[Hen81i] Daniel Henry. Neighborhood of a periodic solution. Lecture Notes in Mathematics, 840:247–274, 1981. CODEN LNMAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089656/.

Henry:1981:NEP

[Hen81j] Daniel Henry. Neighborhood of an equilibrium point. Lecture Notes in Mathematics, 840:98–142, 1981. CODEN LN-MAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089653/.

Henry:1981:NIM

[Hen81k] Daniel Henry. The neighborhood of an invariant manifold. Lecture Notes in Mathematics, 840:275–313, 1981. CODEN LN-MAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089657/.

Henry:1981:P

[Hen81l] Daniel Henry. Preliminaries. Lecture Notes in Mathematics, 840: 3–40, 1981. CODEN LNMAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089649/.

Henry:1981:TE

[Hen81m] Daniel Henry. Two examples. Lecture Notes in Mathematics, 840: 314–330, 1981. CODEN LNMAA2. ISBN 3-540-10557-3 (print), 3-540-38528-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089658/.

Hennart:1982:NAP

[Hen82a] J. P. Hennart, editor. Numerical Analysis: Proceedings of the Third IIMAS Workshop Held at Cocoyoc, Mexico, January [19– 23,] 1981, volume 909 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 909. URL http:/

/link.springer.com/book/10.1007/BFb0092954;   http://
www.springerlink.com/content/978-3-540-38986-6.

Hennart:1982:TFE

[Hen82b] J. P. Hennart. Topics in finite element discretization of parabolic evolution problems. Lecture Notes in Mathematics, 909:185– 199, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092973/.

Hennequin:1982:EEP

[Hen82c] P. L. Hennequin, editor. École d’Été de Probabilités de Saint-Flour X — 1980. (French) [], volume 929 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11547-1 (print), 3-540-39220-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 929.

URL http://link.springer.com/book/10.1007/BFb0095617;
 http://www.springerlink.com/content/978-3-540-39220-
0.

Henry:1982:HRS

[Hen82d] Daniel B. Henry. How to remember the Sobolev inequalities. Lecture Notes in Mathematics, 957:97–109, 1982. CODEN LN-MAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066235/.

Hennefeld:1983:IRS

[Hen83] Julien Hennefeld. M-ideals, related spaces, and some approximation properties. Lecture Notes in Mathematics, 991:96–102, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061561/.

Hennequin:1984:EEP

[Hen84a] P. L. Hennequin, editor. École d’Été de Probabilités de Saint-Flour XII — 1982. (French) [], volume 1097 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13897-8 (print), 3-540-39109-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1097.

URL http://link.springer.com/book/10.1007/BFb0099431;
 http://www.springerlink.com/content/978-3-540-39109-
8.

Hensz:1984:SET

[Hen84b] E. Hensz. On some ergodic theorems for von Neumann algebras. Lecture Notes in Mathematics, 1080:119–123, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099788/.

Hermann:1980:GVF

[Her80a] Robert Hermann. A geometric variational formalism for the theory of nonlinear waves. Lecture Notes in Mathematics, 775:145– 161, 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092025/.

Hermanns:1980:CGE

[Her80b] Franz-Josef Hermanns. On certain groups of exponent eight generated by three involutions. Lecture Notes in Mathematics, 806: 231–245, 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091272/.

Herzer:1980:CKS

[Her80c] Armin Herzer. On characterisations of kinematic spaces by parallelisms. Lecture Notes in Mathematics, 792:61–67, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088663/.

Herb:1981:FIP

[Her81] Rebecca A. Herb. Fourier inversion and the Plancherel theorem. Lecture Notes in Mathematics, 880:197–210, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090410/.

Hergert:1982:ECV

[Her82a] Ferdinand Hergert. The equivalence classes of the Vasil’ev codes of length 15. Lecture Notes in Mathematics, 969:176–186, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062993/.

Hermann:1982:FSC

[Her82b] Robert Hermann. Fiber spaces, connections and Yang–Mills fields. Lecture Notes in Mathematics, 926:1–73, 1982. CODEN LNMAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092656/.

Hermans:1982:WPS

[Her82c] A. J. Hermans. Wave pattern of a ship sailing at low speed. Lecture Notes in Mathematics, 942:281–294, 1982. CODEN LN-MAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094753/.

Herrlich:1982:UCC

[Her82d] Horst Herrlich. Universal completions of concrete categories. Lecture Notes in Mathematics, 915:127–135, 1982. CODEN LN-MAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092876/.

Herb:1983:PTS

[Her83a] Rebecca Herb. The Plancherel theorem for semisimple Lie groups without compact Cartan subgroups. Lecture Notes in Mathematics, 1020:73–79, 1983. CODEN LNMAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071497/.

Hertle:1983:ABG

[Her83b] Alexander Hertle. On the asymptotic behaviour of Gaussian spherical integrals. Lecture Notes in Mathematics, 990:221– 234, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064275/.

Hernandez:1984:RBT

[Her84a] Eugenio Hernández. A relation between two interpolation methods. Lecture Notes in Mathematics, 1070:80–91, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099095/.

Herz:1984:PKM

[Her84b] Carl Herz. The Poisson kernel for _{sl} (3, R). Lecture Notes in Mathematics, 1096:333–346, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100118/.

Hess:1980:CSM

[Hes80] Harald Hess. Connections on symplectic manifolds and geometric quantization. Lecture Notes in Mathematics, 836:153–166, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089731/.

Hesselink:1981:CUA

[Hes81] Wim H. Hesselink. Concentration under actions of algebraic groups. Lecture Notes in Mathematics, 867:55–89, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090380/.

Hess:1983:EEM

[Hes83] H.-U. Hess. On Etcheberry’s extended Milutin lemma. Lecture Notes in Mathematics, 991:103–110, 1983. CODEN LN-MAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061562/.

Hess:1984:PSS

[Hes84] P. Hess. On positive solutions of semilinear periodic-parabolic problems. Lecture Notes in Mathematics, 1076:101–114, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072770/.

Heywood:1980:AFP

[Hey80a] John G. Heywood. Auxiliary flux and pressure conditions for Navier–Stokes problems. Lecture Notes in Mathematics, 771:223– 234, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086909/.

Heywood:1980:CSN

[Hey80b] John G. Heywood. Classical solutions of the Navier–Stokes equations. Lecture Notes in Mathematics, 771:235–248, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086910/.

Heyer:1981:AMM

[Hey81] Herbert Heyer. An application of the method of moments to the central limit theorem on hyperbolic spaces. Lecture Notes in Mathematics, 861:65–73, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097315/.

Heyer:1982:PMG

[Hey82] Herbert Heyer, editor. Probability Measures on Groups: Proceedings of the Sixth Conference Held at Oberwolfach, Germany, June 28–July 4, 1981, volume 928 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 0-387-11501-3 (paperback), 3-540-11501-3 (paperback), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 928; QA3 .L28 no. 928 QA273.A1. URL http:

//link.springer.com/book/10.1007/BFb0093215;   http:/
/www.springerlink.com/content/978-3-540-39206-4.

Heyer:1984:PMG

[Hey84] Herbert Heyer, editor. Probability Measures on Groups VII: Proceedings of a Conference held in Oberwolfach, 24–30 April 1983, volume 1064 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 0-387-13341-0, 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1064; QA3 .L28 no. 1064 QA273.A1. URL http:/

/link.springer.com/book/10.1007/BFb0073628;   http://
www.springerlink.com/content/978-3-540-38874-6.

Hahn-Goldberg:1982:RTI

[HG82a] Saul Hahn-Goldberg. A regularity theorem for inverse bounded and accretive operators in abstract Hilbert space. Lecture Notes in Mathematics, 957:88–96, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066234/.

Hans-Gill:1982:SRR

[HG82b] R. J. Hans-Gill. Some recent results on positive values of nonhomogeneous indefinite quadratic forms. Lecture Notes in Mathematics, 938:75–81, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097174/.

Hahn:1981:PTP

[HH81a] Marjorie G. Hahn and Peter Hahn. The pointwise translation problem for the Radon transform in Banach spaces. Lecture Notes in Mathematics, 860:176–186, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090615/.

Hastings:1981:SHI

[HH81b] Harold M. Hastings and Alex Heller. Splitting homotopy idempotents. Lecture Notes in Mathematics, 870:23–36, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089705/.

Hartshorne:1982:DPG

[HH82] Robin Hartshorne and André Hirschowitz. Droites en position générale dans l’espace projectif. (French) []. Lecture Notes in Mathematics, 961:169–188, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071282/.

Hambleton:1984:AMP

[HH84] Ian Hambleton and Jean-Claude Hausmann. Acyclic maps and Poincaré spaces. Lecture Notes in Mathematics, 1051:222–245, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075569/.

Hou:1984:SOE

[HHB84] Jean W. Hou, Edward J. Haug, and Robert L. Benedict. Shape optimization of elastic bars in torsion. Lecture Notes in Mathematics, 1086:31–55, 1984. CODEN LNMAA2. ISBN 3-540-13871-4 (print), 3-540-39061-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073068/.

Heinrich:1983:EEP

[HHM83] S. Heinrich, C. W. Henson, and L. C. Moore, Jr. Elementary equivalence of L_{1}-preduals. Lecture Notes in Mathematics, 991: 79–90, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061559/.

Havin:1984:LCA

[HHN84] Victor P. Havin, Sergei V. Hruščëv, and Nikolai K. Nikol’skii. Linear and Complex Analysis Problem Book: 199 Research Problems, volume 1043 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 720 pp. LCCN QA3 .L28 no. 1043. URL http:/

/link.springer.com/book/10.1007/BFb0072183;   http://
www.springerlink.com/content/978-3-540-38758-9.

Husch:1981:ECS

[HI81a] L. S. Husch and I. Ivanšić. Embedding compacta up to shape. Lecture Notes in Mathematics, 870:119–134, 1981. CODEN LN-MAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089712/.

Husch:1981:SC

[HI81b] L. S. Husch and I. Ivanšić. On shape concordances. Lecture Notes in Mathematics, 870:135–149, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089713/.

Hasminskii:1983:AEP

[HI83] R. Z. Hasminskii and I. A. Ibragimov. On asymptotic efficiency in the presence of an infinitedimensional nuisance parameter. Lecture Notes in Mathematics, 1021:195–229, 1983. CODEN LN-MAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072916/.

Hiai:1984:SLL

[Hia84] Fumio Hiai. Strong laws of large numbers for multivalued random variables. Lecture Notes in Mathematics, 1091:160–172, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098809/.

Higuchi:1982:ERA

[Hig82a] Isao Higuchi. Existence de resolvantes associées à un noyau verifiant le principe de domination. (French) []. Lecture Notes in Mathematics, 906:203–224, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093268/.

Higuchi:1982:RPC

[Hig82b] Isao Higuchi. Régularité et propriété de convergence dominée des potentiels d’un noyau-fonction non-symétrique. (French) []. Lecture Notes in Mathematics, 906:158–202, 1982. CODEN LN-MAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093267/.

Hillman:1981:AIL

[Hil81a] Jonathan A. Hillman. Alexander Ideals of Links, volume 895 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 178 pp. LCCN QA3 .L28 no. 895. URL http:/

/link.springer.com/book/10.1007/BFb0091682;   http://
www.springerlink.com/content/978-3-540-38643-8.

Hillman:1981:CES

[Hil81b] Jonathan A. Hillman. The Crowell exact sequence. Lecture Notes in Mathematics, 895:40–53, 1981. CODEN LNMAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091687/.

Hillman:1981:DIM

[Hil81c] Jonathan A. Hillman. Determinatal invariants of modules. Lecture Notes in Mathematics, 895:27–39, 1981. CODEN LN-MAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091686/.

Hillman:1981:LLG

[Hil81d] Jonathan A. Hillman. Links and link groups. Lecture Notes in Mathematics, 895:2–15, 1981. CODEN LNMAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091684/.

Hillman:1981:LBP

[Hil81e] Jonathan A. Hillman. Localizing the Blanchfield pairing. Lecture Notes in Mathematics, 895:116–139, 1981. CODEN LN-MAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091692/.

Hillman:1981:LP

[Hil81f] Jonathan A. Hillman. Longitudes and principality. Lecture Notes in Mathematics, 895:66–81, 1981. CODEN LNMAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091689/.

Hillman:1981:MSS

[Hil81g] Jonathan A. Hillman. Monorientable spanning surfaces. Lecture Notes in Mathematics, 895:140–152, 1981. CODEN LN-MAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091693/.

Hillman:1981:P

[Hil81h] Jonathan A. Hillman. Preliminaries. Lecture Notes in Mathematics, 895:1, 1981. CODEN LNMAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091683.

Hillman:1981:RAI

[Hil81i] Jonathan A. Hillman. Reduced Alexander ideals. Lecture Notes in Mathematics, 895:100–115, 1981. CODEN LNMAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091691/.

Hillman:1981:RL

[Hil81j] Jonathan A. Hillman. Ribbon links. Lecture Notes in Mathematics, 895:16–26, 1981. CODEN LNMAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091685/.

Hillman:1981:S

[Hil81k] Jonathan A. Hillman. Sublinks. Lecture Notes in Mathematics, 895:82–99, 1981. CODEN LNMAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091690/.

Hillman:1981:VAI

[Hil81l] Jonathan A. Hillman. The vanishing of Alexander ideals. Lecture Notes in Mathematics, 895:54–65, 1981. CODEN LN-MAA2. ISBN 3-540-11168-9 (print), 3-540-38643-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091688/.

Hilhorst:1982:PFB

[Hil82a] D. Hilhorst. A perturbed free boundary problem arising in the physics of ionized gases. Lecture Notes in Mathematics, 942:309– 317, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094755/.

Hiller:1982:ALA

[Hil82b] Howard Hiller. Affine Lie algebras and algebraic K-theory. Lecture Notes in Mathematics, 966:91–107, 1982. CODEN LN-MAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062168/.

Hilton:1982:RNG

[Hil82c] Peter Hilton. Relative nilpotent groups. Lecture Notes in Mathematics, 915:136–147, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092877/.

Hilton:1983:LCM

[Hil83] Peter Hilton. Localization of crossed-modules. Lecture Notes in Mathematics, 1029:311–324, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098937/.

Hildebrandt:1984:SRM

[Hil84] S. Hildebrandt. Some results on minimal surfaces with free boundaries. Lecture Notes in Mathematics, 1107:115–134, 1984. CO-DEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101497/.

Hinze:1982:NID

[Hin82] Juergen Hinze, editor. Numerical Integration of Differential Equations and Large Linear Systems: Proceedings of two Workshops Held at the University of Bielefeld Spring 1980, volume 968 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 968. URL http:/

/link.springer.com/book/10.1007/BFb0064877;   http://
www.springerlink.com/content/978-3-540-39374-0.

Hinman:1984:FAS

[Hin84] Peter G. Hinman. Finitely approximable sets. Lecture Notes in Mathematics, 1104:233–258, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099488/.

Hirsch:1980:PCD

[Hir80] F. Hirsch. Principe complet du maximum et principe complet du maximum relatif. (French) []. Lecture Notes in Mathematics, 787:144–158, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086330/.

Hirschfeld:1983:WCF

[Hir83] J. W. P. Hirschfeld. The Weil conjectures in finite geometry. Lecture Notes in Mathematics, 1036:6–23, 1983. CODEN LN-MAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071506/.

Hirsch:1984:ALT

[Hir84a] Francis Hirsch. Aspects linéaires de la théorie du potentiel dans les travaux de Jacques Deny. (French) []. Lecture Notes in Mathematics, 1096:22–33, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100105/.

Hirsch:1984:GES

[Hir84b] Francis Hirsch. Générateurs étendus et subordination au sens de Bochner. (French) []. Lecture Notes in Mathematics, 1061: 134–156, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099022/.

Hitchin:1982:CME

[Hit82] N. J. Hitchin. Complex manifolds and Einstein’s equations. Lecture Notes in Mathematics, 970:73–99, 1982. CODEN LN-MAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066025/.

Havin:1981:CUT

[HJ81] V. P. Havin and B. Jöricke. On a class of uniqueness theorems for convolutions. Lecture Notes in Mathematics, 864:143–170, 1981. CODEN LNMAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096998/.

Hardie:1982:PNO

[HJ82a] K. A. Hardie and A. V. Jansen. The Puppe and Nomura operators in the category of homotopy pairs. Lecture Notes in Mathematics, 915:112–126, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092875/.

Hoffmann-Jorgensen:1982:PGB

[HJ82b] J. Hoffmann-Jørgensen. Probability and geometry of Banach spaces. Lecture Notes in Mathematics, 948:164–229, 1982. CO-DEN LNMAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069846/.

Hsiang:1982:NHG

[HJ82c] W. C. Hsiang and B. Jahren. A note on the homotopy groups of the diffeomorphism groups of spherical space forms. Lecture Notes in Mathematics, 967:132–145, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061901/.

Herb:1984:LGR

[HJLR84] Rebecca Herb, Raymond Johnson, Ronald Lipsman, and Jonathan Rosenberg, editors. Lie Group Representations III: Proceedings of the Special Year held at the University of Maryland, College Park 1982–1983, volume 1077 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1077. URL http:/

/link.springer.com/book/10.1007/BFb0072333;   http://
www.springerlink.com/content/978-3-540-38936-1.

Hofmann:1983:RDA

[HJW83] Karl Heinrich Hofmann, Helmut Jürgensen, and Hanns Joachim Weinert, editors. Recent Developments in the Algebraic, Analytical, and Topological Theory of Semigroups: Proceedings of a Conference Held at Oberwolfach, Germany, May 24–30, 1981, volume 998 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 998. URL http:/

/link.springer.com/book/10.1007/BFb0062023;   http://
www.springerlink.com/content/978-3-540-40051-6.

Hahn:1980:GDA

[HK80a] Marjorie G. Hahn and Michael J. Klass. The generalized domain of attraction of spherically symmetric stable laws on R^{d}. Lecture Notes in Mathematics, 828:52–81, 1980. CODEN LN-MAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097395/.

Hurder:1980:HIF

[HK80b] S. Hurder and F. W. Kamber. Homotopy invariants of foliations. Lecture Notes in Mathematics, 788:49–61, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099238/.

Hahn:1981:SGD

[HK81] Marjorie G. Hahn and Michael J. Klass. A survey of generalized domains of attraction and operator norming methods. Lecture Notes in Mathematics, 860:187–218, 1981. CODEN LN-MAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090616/.

Hedberg:1984:AC

[HKD^{+}84] Lars Inge Hedberg, Thomas Kriete, Harry Dym, Arne Stray, V. I. Belyi, André Boivin, Paul M. Gauthier, J. Brennan, Thomas Bagby, A. A. Gonchar, Hans Wallin, Lee A. Rubel, G. M. Henkin, Donald E. Marshall, W. K. Hayman, et al. Approximation and capacities. Lecture Notes in Mathematics, 1043:431–519, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072191/.

Herb:1983:LGRb

[HKLR83] Rebecca Herb, Stephen Kudla, Ronald Lipsman, and Jonathan Rosenberg, editors. Lie Group Representations II: Proceedings of the Special Year held at the University of Maryland, College Park 1982–1983, volume 1041 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1041. URL http:/

/link.springer.com/book/10.1007/BFb0073143;   http://
www.springerlink.com/content/978-3-540-38699-5.

Henderson:1983:SPC

[HKP83] W. Henderson, R. W. Kennington, and C. E. M. Pearce. Stochastic processes and combinatoric identities. Lecture Notes in Mathematics, 1036:230–243, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071522/.

Haebich:1981:MME

[HL81] W. Haebich and J.-L. Lassez. Minimisatin of multiple entry finite automata. Lecture Notes in Mathematics, 884:203–210, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091820/.

Haefliger:1983:CCI

[HL83a] A. Haefliger and Banghe Li. Currents on a circle invariant by a Fuchsian group. Lecture Notes in Mathematics, 1007:369– 378, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061424/.

Hofmann:1983:FLS

[HL83b] K. H. Hofmann and J. D. Lawson. Foundations of Lie semigroups. Lecture Notes in Mathematics, 998:128–201, 1983. CO-DEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062029/.

Herb:1983:LGRa

[HLR83] Rebecca Herb, Ronald Lipsman, and Jonathan Rosenberg, editors. Lie Group Representations I: Proceedings of the Special Year held at the University of Maryland, College Park, 1982– 1983, volume 1024 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12725-9 (print), 3-540-38709-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1024. URL http:/

/link.springer.com/book/10.1007/BFb0071429;   http://
www.springerlink.com/content/978-3-540-38709-1.

Heyvaerts:1980:SFN

[HLSW80] J. Heyvaerts, J.-M. Lasry, M. Schatzman, and P. Witomski. Solar flares: a non linear eigenvalue problem in an unbounded domain. Lecture Notes in Mathematics, 782:160–191, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090432/.

Hirsch:1980:STD

[HM80] Francis Hirsch and Gabriel Mokobodzki. Séminaire de Théorie du Potentiel Paris, No. 5, volume 814 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 239 pp. LCCN QA3 .L28 no. 814.

URL http://link.springer.com/book/10.1007/BFb0094144;
 http://www.springerlink.com/content/978-3-540-38189-
1.

Hofmann:1981:LCC

[HM81] Karl H. Hofmann and Michael W. Mislove. Local compactness and continuous lattices. Lecture Notes in Mathematics, 871:209– 248, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089908/.

Henry:1982:LET

[HM82a] J.-P. G. Henry and M. Merle. Limites d’espaces tangents et transversalité de variétés polaires. (French) []. Lecture Notes in Mathematics, 961:189–199, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071283/.

Hirsch:1982:STD

[HM82b] Francis Hirsch and Gabriel Mokobodzki. Séminaire de Théorie du Potentiel Paris, No. 6, volume 906 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 328 pp. LCCN QA3 .L28 no. 906.

URL http://link.springer.com/book/10.1007/BFb0093258;
 http://www.springerlink.com/content/978-3-540-38971-
2.

Houdebine:1982:SCC

[HM82c] Jean Houdebine and Louis Mahé. Séparation des composantes connexes réelles dans le cas des variétés projectives. (French) []. Lecture Notes in Mathematics, 959:358–370, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062263/.

Hsiao:1982:SPT

[HM82d] George C. Hsiao and Richard C. MacCamy. Singular perturbations for the two-dimensional viscous flow problem. Lecture Notes in Mathematics, 942:229–244, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094750/.

Hammersley:1983:MFC

[HM83a] J. M. Hammersley and G. Mazzarino. Markov fields, correlated percolation, and the Ising model. Lecture Notes in Mathematics, 1035:201–245, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073261/.

Heinrich:1983:SOP

[HM83b] S. Heinrich and P. Mankiewicz. Some open problems in the nonlinear classification of Banach spaces. Lecture Notes in Mathematics, 991:91–95, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061560/.

Henson:1983:NAT

[HM83c] C. Ward Henson and L. C. Moore, Jr. Nonstandard analysis and the theory of Banach spaces. Lecture Notes in Mathematics, 983:27–112, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065334/.

Hirsch:1984:STD

[HM84a] Francis Hirsch and Gabriel Mokobodzki. Séminaire de Théorie du Potentiel Paris, No. 7, volume 1061 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 281 pp. LCCN QA3 .L28 no. 1061.

URL http://link.springer.com/book/10.1007/BFb0099013;
 http://www.springerlink.com/content/978-3-540-38866-
1.

Hognas:1984:LAV

[HM84b] Göran Högnäs and Arunava Mukherjea. On the limit of the average of the values of a function at random points. Lecture Notes in Mathematics, 1064:204–218, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073643/.

Holton:1982:CPT

[HMP82] D. A. Holton, B. D. Mc Kay, and M. D. Plummer. A corollary to Perfect’s theorem. Lecture Notes in Mathematics, 952: 294–299, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061985/.

Henkin:1981:CSAa

[HMT81a] L. Henkin, J. D. Monk, and A. Tarski. Cylindric set algebras and related structures. Lecture Notes in Mathematics, 883:1– 129, 1981. CODEN LNMAA2. ISBN 3-540-10881-5 (print), 3-540-38790-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095613/.

Henkin:1981:CSAb

[HMT^{+}81b] Leon Henkin, J. Donald Monk, Alfred Tarski, Hajnalka Andréka, and István Németi. Cylindric Set Algebras, volume 883 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981.

CODEN LNMAA2. ISBN 3-540-10881-5 (print), 3-540-38790-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 323 pp. LCCN QA3 .L28 no. 883. URL http:/

/link.springer.com/book/10.1007/BFb0095612;   http://
www.springerlink.com/content/978-3-540-38790-9.

Hartzman:1980:SNI

[HN80a] C. S. Hartzman and D. R. Naugler. Separatrices, non-isolated invariant sets and the Seifert conjecture. Lecture Notes in Mathematics, 819:224–245, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086990/.

Havas:1980:ACQ

[HN80b] George Havas and M. F. Newman. Application of computers to questions like those of burnside. Lecture Notes in Mathematics, 806:211–230, 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091271/.

Havin:1981:CAS

[HN81a] Victor P. Havin and Nikolai K. Nikol’skii, editors. Complex Analysis and Spectral Theory: Seminar, Leningrad 1979/ 80, volume 864 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 864. URL http:/

/link.springer.com/book/10.1007/BFb0096995;   http://
www.springerlink.com/content/978-3-540-38626-1.

Hirose:1981:RSS

[HN81b] Ken Hirose and Fujio Nakayasu. A representation for spector second order classes in computation theories on two types. Lecture Notes in Mathematics, 891:31–47, 1981. CODEN LN-MAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090977/.

Hong:1982:SDI

[HN82] S. S. Hong and L. D. Nel. Spectral dualities involving mixed structures. Lecture Notes in Mathematics, 915:198–204, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092880/.

Hughes:1983:MPD

[HN83] B. D. Hughes and B. W. Ninham, editors. The Mathematics and Physics of Disordered Media: Percolation, Random Walk, Modeling, and Simulation: Proceedings of a Workshop held at the IMA, University of Minnesota, Minneapolis February 13–19, 1983, volume 1035 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1035. URL http:/

/link.springer.com/book/10.1007/BFb0073254;   http://
www.springerlink.com/content/978-3-540-38693-3.

Hruscev:1981:UBE

[HNP81] S. V. Hruščev, N. K. Nikol’skii, and B. S. Pavlov. Unconditional bases of exponentials and of reproducing kernels. Lecture Notes in Mathematics, 864:214–335, 1981. CODEN LNMAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097000/.

Howes:1980:SPS

[HO80] F. A. Howes and R. E. O’Malley, Jr. Singular perturbations of semilinear second order systems. Lecture Notes in Mathematics, 827:131–150, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091378/.

Herrmann:1982:BEN

[HO82] M. Herrmann and U. Orbanz. Between equimultiplicity and normal flatness. Lecture Notes in Mathematics, 961:200–232, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071284/.

Hochster:1981:DIA

[Hoc81] Melvin Hochster. The dimension of an intersection in an ambient hypersurface. Lecture Notes in Mathematics, 862:93–106, 1981. CODEN LNMAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090890/.

Hodges:1984:FEF

[Hod84] Wilfrid Hodges. Finite extensions of finite groups. Lecture Notes in Mathematics, 1103:193–206, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099387/.

Hof:1981:VHB

[Hof81a] Hans-Christoph Im Hof. Visibility, horocycles, and the Bruhat decomposition. Lecture Notes in Mathematics, 838:149–153, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088852/.

Hoffmann:1981:FFR

[Hof81b] K.-H. Hoffmann. Fixpunktprinzipien und Freie Randwertaufgaben. (German) []. Lecture Notes in Mathematics, 878:162– 181, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090681/.

Hoffmann:1981:CPP

[Hof81c] Rudolf-E. Hoffmann. Continuous posets, prime spectra of completely distributive complete lattices, and Hausdorff compactifications. Lecture Notes in Mathematics, 871:159–208, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089907/.

Hoffmann:1981:PSS

[Hof81d] Rudolf-E. Hoffmann. Projective Sober spaces. Lecture Notes in Mathematics, 871:125–158, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089906/.

Hoffmann:1982:FCI

[Hof82a] Rudolf-E. Hoffmann. Factorization of cones II, with applications to weak Hausdorff spaces. Lecture Notes in Mathematics, 915: 148–170, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092878/.

Hoffmann:1982:MTC

[Hof82b]  Rudolf-E. Hoffmann.  Minimal topological completion of

KBan_{1} →_{K} Vec. Lecture Notes in Mathematics, 962:123–132, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066891/.

Hofmann:1982:EFC

[Hof82c] Karl Heinrich Hofmann. An essay on free compact groups. Lecture Notes in Mathematics, 915:171–197, 1982. CODEN LN-MAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092879/.

Hofmann:1983:RTT

[Hof83] Karl Heinrich Hofmann. Recent trends in the theory of semigroups in analysis and topology. Lecture Notes in Mathematics, 998:1–7, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062024/.

Hohlov:1983:SEC

[Hoh83] Yu. S. Hohlov. On stability estimation of certain characterization of the exponential distribution. Lecture Notes in Mathematics, 982:32–37, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082059/.

Holmann:1980:SHF

[Hol80] Harald Holmann. On the stability of holomorphic foliations. Lecture Notes in Mathematics, 798:192–202, 1980. CODEN LN-MAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097265/.

Holmes:1981:STP

[Hol81] Philip Holmes. Space- and time-periodic perturbations of the sine-Gordon equation. Lecture Notes in Mathematics, 898:164– 191, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091913/.

Holland:1983:SVL

[Hol83a] W. Charles Holland. A survey of varieties of lattice ordered groups. Lecture Notes in Mathematics, 1004:153–158, 1983. CO-DEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063435/.

Holton:1983:CG

[Hol83b] D. A. Holton. Cycles in graphs. Lecture Notes in Mathematics, 1036:24–48, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071507/.

Holevo:1984:CMI

[Hol84] A. S. Holevo. Covariant measurements and imprimitivity systems. Lecture Notes in Mathematics, 1055:153–172, 1984. CO-DEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071720/.

Honig:1980:VSI

[Hön80] Chaim Samuel Hönig. Volterra Stieltjes-integral equations. Lecture Notes in Mathematics, 799:173–216, 1980. CODEN LN-MAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089315/.

Honig:1982:AEL

[Hön82] Chaim Samuel Hönig. The adjoint equation of a linear Volterra Stieltjes-integral equation with a linear constraint. Lecture Notes in Mathematics, 957:110–125, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066236/.

Hoobler:1982:W

[Hoo82] Raymond T. Hoobler. When is Br(X) = Br′(X)? Lecture Notes in Mathematics, 917:231–244, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092238/.

Hoppner:1982:FWD

[Höp82] Michael Höppner. On the freeness of Whitehead-diagrams. Lecture Notes in Mathematics, 962:133–137, 1982. CODEN LN-MAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066892/.

Hopf:1983:BM

[Hop83a] Heinz Hopf. Back matter. Lecture Notes in Mathematics, 1000: 185–187, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-662-21563-0/1.pdf.

Hopf:1983:CSC

[Hop83b] Heinz Hopf. Closed surfaces with constant Gauss curvature (Hilbert’s method) — generalizations and problems — general remarks on Weingarten surfaces. Lecture Notes in Mathematics, 1000:123–135, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21563-0_10/.

Hopf:1983:CTO

[Hop83c] Heinz Hopf. The congruence theorem for ovaloids. Lecture Notes in Mathematics, 1000:163–173, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21563-0_13/.

Hopf:1983:DGL

[Hop83d] Heinz Hopf, editor. Differential Geometry in the Large: Seminar Lectures New York University 1946 and Stanford University 1956, volume 1000 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1000. URL http://link.

springer.com/book/10.1007/978-3-662-21563-0;   http:/
/www.springerlink.com/content/978-3-662-21563-0.

Hopf:1983:DGS

[Hop83e] Heinz Hopf. Differential geometry of surfaces in the small [summary sketch]. Lecture Notes in Mathematics, 1000:82–99, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21563-0_6/.

Hopf:1983:ECA

[Hop83f] Heinz Hopf. The elementary concept of area and volume. Lecture Notes in Mathematics, 1000:58–75, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21563-0_4/.

Hopf:1983:ECR

[Hop83g] Heinz Hopf. The Euler characteristic and related topics. Lecture Notes in Mathematics, 1000:3–29, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21563-0_1/.

Hopf:1983:FMa

[Hop83h] Heinz Hopf. Front matter. Lecture Notes in Mathematics, 1000: 1, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-662-21563-0/1/1.pdf.

Hopf:1983:FMb

[Hop83i] Heinz Hopf. Front matter. Lecture Notes in Mathematics, 1000: 77–80, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-662-21563-0/2/1.pdf.

Hopf:1983:FMc

[Hop83j] Heinz Hopf. Front matter. Lecture Notes in Mathematics, 1000: N2–VII, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-662-21563-0/1.pdf.

Hopf:1983:GCS

[Hop83k] Heinz Hopf. General closed surfaces of genus O with constant mean curvature — generalizations. Lecture Notes in Mathematics, 1000:136–146, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21563-0_11/.

Hopf:1983:HCO

[Hop83l] Heinz Hopf. Hadamard’s characterization of the ovaloids. Lecture Notes in Mathematics, 1000:119–122, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21563-0_9/.

Hopf:1983:I

[Hop83m] Heinz Hopf. Introduction. Lecture Notes in Mathematics, 1000: 81, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21563-0_5.

Hopf:1983:IIR

[Hop83n] Heinz Hopf. The isoperimetric inequality and related inequalities. Lecture Notes in Mathematics, 1000:47–57, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21563-0_
3/.

Hopf:1983:STE

[Hop83o] Heinz Hopf. Selected topics in elementary differential geometry. Lecture Notes in Mathematics, 1000:30–46, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21563-0_
2/.

Hopf:1983:SCS

[Hop83p] Heinz Hopf. Simple closed surfaces (of arbitrary genus) with constant mean curvature — generalizations. Lecture Notes in Mathematics, 1000:147–162, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21563-0_12/.

Hopf:1983:SSC

[Hop83q] Heinz Hopf. Singularities of surfaces with constant negative Gauss curvature. Lecture Notes in Mathematics, 1000:174–184, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21563-0_14/.

Hopf:1983:SGR

[Hop83r] Heinz Hopf. Some general remarks on closed surfaces in differential geometry. Lecture Notes in Mathematics, 1000:100– 106, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21563-0_7/.

Hopf:1983:TCC

[Hop83s] Heinz Hopf. The total curvature (curvatura integra) of a closed surface with Riemannian metric and Poincaré’s theorem on the singularities of fields of line elements. Lecture Notes in Mathematics, 1000:107–118, 1983. CODEN LNMAA2. ISBN 3-540-12004-1 (print), 3-662-21563-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21563-0_8/.

Horvathy:1980:CAW

[Hor80] Péter A. Horváthy. Classical action, the Wu–Yang phase factor and prequantization. Lecture Notes in Mathematics, 836: 67–90, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089727/.

Hornung:1982:AMN

[Hor82a] Ulrich Hornung. Adi-methods for nonlinear variational inequalities of evolution. Lecture Notes in Mathematics, 953:138–148, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069379/.

Horvathy:1982:PPI

[Hor82b] P. A. Horváthy. Prequantisation from path integral viewpoint. Lecture Notes in Mathematics, 905:197–206, 1982. CODEN LN-MAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092438/.

Hoskins:1982:FBM

[Hos82] J. A. Hoskins. Factoring binary matrices: a weaver’s approach. Lecture Notes in Mathematics, 952:300–326, 1982. CODEN LN-MAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061986/.

Howard:1981:COR

[How81a] W. A. Howard. Computability of ordinal recursion of type level two. Lecture Notes in Mathematics, 873:87–104, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090729/.

Howe:1981:AFL

[How81b] Roger Howe. Automorphic forms of low rank. Lecture Notes in Mathematics, 880:211–248, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090411/.

Howes:1982:ASS

[How82] F. A. Howes. The asymptotic solution of singularly perturbed Dirichlet problems with applications to the study of incompressible flows at high Reynolds number. Lecture Notes in Mathematics, 942:245–257, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094751/.

Hajek:1980:TOC

[HP80] Petr Hájek and Pavel Pudlák. Two orderings of the class of all countable models of Peano arithmetic. Lecture Notes in Mathematics, 834:174–185, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090166/.

Hughes:1983:RPR

[HP83] Barry D. Hughes and Stephen Prager. Random processes and random systems: an introduction. Lecture Notes in Mathematics, 1035:1–108, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073255/.

Hilton:1984:AEN

[HP84a] Peter J. Hilton and Markus Pfenniger. Allgemeine Eigenschaften nilpotenter Gruppen. (German) []. Lecture Notes in Mathematics, 1053:21–40, 1984. CODEN LNMAA2. ISBN 3-540-12910-3 (print), 3-540-38791-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071371/.

Hilton:1984:EEN

[HP84b] Peter J. Hilton and Markus Pfenniger. Endlich erzeugte nilpotente Gruppen. (German) [Finitely-generated nilpotent groups]. Lecture Notes in Mathematics, 1053:40–61, 1984. CODEN LN-MAA2. ISBN 3-540-12910-3 (print), 3-540-38791-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071372/.

Hilton:1984:ENG

[HP84c] Peter J. Hilton and Markus Pfenniger. Endliche und nilpotente Gruppen. (German) []. Lecture Notes in Mathematics, 1053:1– 20, 1984. CODEN LNMAA2. ISBN 3-540-12910-3 (print), 3-540-38791-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071370/.

Hilton:1984:LAG

[HP84d] Peter J. Hilton and Markus Pfenniger. Lokalisierung abelscher Gruppen. (German) [Localization of Abelian groups]. Lecture Notes in Mathematics, 1053:62–87, 1984. CODEN LN-MAA2. ISBN 3-540-12910-3 (print), 3-540-38791-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071373/.

Hilton:1984:LEZ

[HP84e] Peter J. Hilton and Markus Pfenniger. Lokalisierung einfach zusammenhängender CW-Komplexe. (German) []. Lecture Notes in Mathematics, 1053:148–163, 1984. CODEN LNMAA2. ISBN 3-540-12910-3 (print), 3-540-38791-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071375/.

Hilton:1984:LNC

[HP84f] Peter J. Hilton and Markus Pfenniger. Lokalisierung nilpotenter CW-Komplexe. (German) [Localization of nilpotent CW complexes]. Lecture Notes in Mathematics, 1053:164–219, 1984. CO-DEN LNMAA2. ISBN 3-540-12910-3 (print), 3-540-38791-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071376/.

Hilton:1984:LNG

[HP84g] Peter J. Hilton and Markus Pfenniger. Lokalisierung nilpotenter Gruppen. (German) [Localization of nilpotent groups]. Lecture Notes in Mathematics, 1053:88–147, 1984. CODEN LN-MAA2. ISBN 3-540-12910-3 (print), 3-540-38791-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071374/.

Hilton:1984:NGN

[HP84h] Peter J. Hilton and Markus Pfenniger, editors. Nilpotente Gruppen und nilpotente Räume: Nachdiplomvorlesung gehalten am Mathematik-Departement ETH Zürich 1981/82, volume 1053 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12910-3 (print), 3-540-38791-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1053. URL http:/

/link.springer.com/book/10.1007/BFb0071369;   http://
www.springerlink.com/content/978-3-540-38791-6.

Hudson:1984:CQD

[HP84i] R. L. Hudson and K. R. Parthasarathy. Construction of quantum diffusions. Lecture Notes in Mathematics, 1055:173–198, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071721/.

Hudson:1984:GWO

[HP84j] R. L. Hudson and K. R. Parthasarathy. Generalised Weyl operators. Lecture Notes in Mathematics, 1095:45–50, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099121/.

Happel:1980:VCD

[HPR80] Dieter Happel, Udo Preiser, and Claus Michael Ringel. Vinberg’s characterization of Dynkin diagrams using subadditive functions with application to DTr-periodic modules. Lecture Notes in Mathematics, 832:280–294, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088469/.

Hohlov:1984:HSA

[HPR84] A. V. Hohlov, A. V. Pazhitnov, and Ju. B. Rudjak. On the homotopical structure and applications of Morava’s extraordinary K-theories. Lecture Notes in Mathematics, 1060:151–163, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099931/.

Happel:1981:CTA

[HR81] Dieter Happel and Claus Michael Ringel. Construction of tilted algebras. Lecture Notes in Mathematics, 903:125–144, 1981. CO-DEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092989/.

Hedstrom:1982:AGM

[HR82] G. W. Hedstrom and G. H. Rodrique. Adaptive-grid methods for time-dependent partial differential equations. Lecture Notes in Mathematics, 960:474–484, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069939/.

Howe:1984:SMC

[HRW84] Roger Howe, Gail Ratcliff, and Norman Wildberger. Symbol mappings for certain nilpotent groups. Lecture Notes in Mathematics, 1077:288–320, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072342/.

Holley:1981:GOU

[HS81] R. Holley and D. Stroock. Generalized Ornstein–Uhlenbeck processes as limits of interacting systems. Lecture Notes in Mathematics, 851:152–168, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088727/.

Hering:1982:NPP

[HS82] Christoph Hering and Hans-Jörg Schaeffer. On the new projective planes of R. Figueroa. Lecture Notes in Mathematics, 969: 187–190, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062994/.

Hess:1984:AAE

[HS84] P. Hess and S. Senn. Another approach to elliptic eigenvalue problems with respect to indefinite weight functions. Lecture Notes in Mathematics, 1107:106–114, 1984. CODEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101496/.

Hsu:1980:CNC

[Hsu80] D. Frank Hsu. Cyclic Neofields and Combinatorial Designs, volume 824 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 230 pp. LCCN QA3 .L28 no. 824. URL http:

//link.springer.com/book/10.1007/BFb0089021.

Howell:1981:MZW

[HT81] Joseph O. Howell and Robert L. Taylor. Marcinkiewicz–Zygmund weak laws of large numbers for unconditional random elements in Banach spaces. Lecture Notes in Mathematics, 860:219–230, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090617/.

Hackbusch:1982:MMP

[HT82a] W. Hackbusch and U. Trottenberg, editors. Multigrid Methods: Proceedings of the Conference Held at Köln–Porz, November 23–27, 1981, volume 960 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 960. URL http:/

/link.springer.com/book/10.1007/BFb0069927;   http://
www.springerlink.com/content/978-3-540-39544-7.

Higgins:1982:FGH

[HT82b] P. J. Higgins and J. Taylor. The fundamental groupoid and the homotopy crossed complex of an orbit space. Lecture Notes in Mathematics, 962:115–122, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066890/.

Howell:1981:SLF

[HTW81] J. Howell, R. L. Taylor, and W. A. Woyczyński. Stability of linear forms in independent random variables in Banach spaces. Lecture Notes in Mathematics, 860:231–245, 1981. CODEN LN-MAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090618/.

Hambleton:1984:IMB

[HTW84] Ian Hambleton, Laurence Taylor, and Bruce Williams. An introduction to maps between surgery obstruction groups. Lecture Notes in Mathematics, 1051:49–127, 1984. CODEN LN-MAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075564/.

Hiriart-Urruty:1983:AFO

[HU83] J.-B. Hiriart-Urruty. The approximate first-order and secondorder directional derivatives for a convex function. Lecture Notes in Mathematics, 979:144–177, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066253/.

Huber:1983:MSTa

[Hub83a] Martin Huber. Methods of set theory and the abundance of separable Abelian p-groups. Lecture Notes in Mathematics, 1006: 304–319, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21560-9_15/.

Huber:1983:MSTb

[Hub83b] Martin Huber. Methods of set theory and the abundance of separable abelian p-groups. Lecture Notes in Mathematics, 1006: 304–319, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103710/.

Hueber:1980:BHP

[Hue80] Hermann Hueber. On boundary Harnack principles and singularities of extremal harmonic functions. Lecture Notes in Mathematics, 787:159–162, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086331/.

Huebschmann:1982:SEI

[Hue82] Johannes Huebschmann. Stem extensions of the infinite general linear group and large Steinberg groups. Lecture Notes in Mathematics, 966:108–111, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062169/.

Hughes:1981:D

[Hug81] D. R. Hughes. On designs. Lecture Notes in Mathematics, 893: 43–67, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091012/.

Hulanicki:1980:CCS

[Hul80] A. Hulanicki. A class of convolution semi-groups of measures on a Lie group. Lecture Notes in Mathematics, 828:82–101, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097396/.

Hull:1982:PCE

[Hul82] T. E. Hull. Precision control, exception handling and a choice of numerical algorithms. Lecture Notes in Mathematics, 912:169– 178, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093156/.

Hulek:1983:PGE

[Hul83] Klaus Hulek. Projective geometry of elliptic curves. Lecture Notes in Mathematics, 997:228–266, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061647/.

Humphreys:1980:GLS

[Hum80a] James E. Humphreys. GL_{n} and SL_{n} (over R). Lecture Notes in Mathematics, 789:28–64, 1980. CODEN LNMAA2. ISBN 3-540-09972-7 (print), 3-540-39198-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094571/.

Humphreys:1980:AAG

[Hum80b] James E. Humphreys. GL_{n} and SL_{n} (p-adic and adelic groups). Lecture Notes in Mathematics, 789:64–83, 1980. CODEN LN-MAA2. ISBN 3-540-09972-7 (print), 3-540-39198-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094572/.

Humphreys:1980:AGa

[Hum80c] James E. Humphreys. The additive group. Lecture Notes in Mathematics, 789:11–21, 1980. CODEN LNMAA2. ISBN 3-540-09972-7 (print), 3-540-39198-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094569/.

Humphreys:1980:AGb

[Hum80d] James E. Humphreys. Arithmetic Groups, volume 789 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09972-7 (print), 3-540-39198-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 158 pp. LCCN QA3 .L28 no. 789. URL http:/

/link.springer.com/book/10.1007/BFb0094567;   http://
www.springerlink.com/content/978-3-540-39198-2.

Humphreys:1980:CSP

[Hum80e] James E. Humphreys. The congruence subgroup problem. Lecture Notes in Mathematics, 789:84–148, 1980. CODEN LN-MAA2. ISBN 3-540-09972-7 (print), 3-540-39198-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094573/.

Humphreys:1980:HWM

[Hum80f] James E. Humphreys. Highest weight modules for semisimple Lie algebras. Lecture Notes in Mathematics, 831:72–103, 1980. CODEN LNMAA2. ISBN 3-540-10263-9 (print), 3-540-38385-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089779/.

Humphreys:1980:LCG

[Hum80g] James E. Humphreys. Locally compact groups and fields. Lecture Notes in Mathematics, 789:1–11, 1980. CODEN LNMAA2. ISBN 3-540-09972-7 (print), 3-540-39198-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094568/.

Humphreys:1980:MG

[Hum80h] James E. Humphreys. The multiplicative group. Lecture Notes in Mathematics, 789:21–28, 1980. CODEN LNMAA2. ISBN 3-540-09972-7 (print), 3-540-39198-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094570/.

Huneke:1983:IL

[Hun83] C. Huneke. The invariants of liaison. Lecture Notes in Mathematics, 1008:65–74, 1983. CODEN LNMAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065700/.

Hurlimann:1981:GBA

[Hür81] W. Hürlimann. Sur le groupe de Brauer d’un anneau de polynômes en caractéristique p et la théorie des invariants. (French) []. Lecture Notes in Mathematics, 844:229–274, 1981. CODEN LNMAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090482/.

Hurley:1982:NCL

[Hur82a] James F. Hurley. A note on the centers of Lie algebras of classical type. Lecture Notes in Mathematics, 933:111–116, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093355/.

Hurlimann:1982:BGD

[Hür82b] Werner Hürlimann. Brauer group and Diophantine geometry: a cohomological approach. Lecture Notes in Mathematics, 917: 43–65, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092227/.

Hurrelbrink:1982:TTR

[Hur82c] Jürgen Hurrelbrink. K_{2}(o) for two totally real fields of degree three and four. Lecture Notes in Mathematics, 966:112–114, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062170/.

Hurd:1983:NAR

[Hur83] Albert Emerson Hurd. Nonstandard Analysis — Recent Developments, volume 983 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 213 pp. LCCN QA3 .L28 no. 983. URL http:

//link.springer.com/book/10.1007/BFb0065331;   http:/
/www.springerlink.com/content/978-3-540-39602-4.

Husemoller:1981:DEL

[Hus81a] Dale Husemoller. La décomposition des espaces des lacets et la torsion impaire des groupes d’homotopie [d’après F. Cohen, J. C. Moore, J. Neisendorfer, et P. Selick]. (French) [The decomposition of loop spaces and the odd torsion of homotopy groups]. Lecture Notes in Mathematics, 842:54–72, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089928/.

Husseini:1981:CM

[Hus81b] S. Y. Husseini. Cobordisms of maps. Lecture Notes in Mathematics, 886:140–153, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092180/.

Husek:1982:ACT

[Huš82a] M. Hušek. Applications of category theory to uniform structures. Lecture Notes in Mathematics, 962:138–144, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066893/.

Husek:1982:COS

[Huš82b] M. Hušek. Categories of orderable spaces. Lecture Notes in Mathematics, 915:205–215, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092881/.

Hall:1981:SCP

[HV81a] R. R. Hall and I. Vincze. On a simultaneous characterization of the Poisson law and the gamma distribution. Lecture Notes in Mathematics, 861:54–59, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097313/.

Hruscev:1981:FIS

[HV81b] S. V. Hruščëv and S. A. Vinogradov. Free interpolation in the space of uniformly convergent Taylor series. Lecture Notes in Mathematics, 864:171–213, 1981. CODEN LNMAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096999/.

Hajek:1981:SPM

[HW81a] Bruce Hajek and Eugene Wong. Set-parametered martingales and multiple stochastic integration. Lecture Notes in Mathematics, 851:119–151, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088726/.

Hofmann:1981:SF

[HW81b] Karl H. Hofmann and Fred Watkins. The spectrum as a functor. Lecture Notes in Mathematics, 871:249–263, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089909/.

Huber:1981:HBC

[HW81c] Martin Huber and R. B. Warfield, Jr. Homomorphisms between cartesian powers of an abelian group. Lecture Notes in Mathematics, 874:202–227, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090536/.

Hairer:1982:CNL

[HW82] E. Hairer and G. Wanner. Characterization of non-linearly stable implicit Runge–Kutta methods. Lecture Notes in Mathematics, 968:207–219, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064889/.

Hwang:1981:CSM

[Hwa81] J. S. Hwang. Complete stable marriages and systems of I-m preferences. Lecture Notes in Mathematics, 884:49–63, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091807/.

Hyland:1981:FSC

[Hyl81] J. M. E. Hyland. Function spaces in the category of locales. Lecture Notes in Mathematics, 871:264–281, 1981. CODEN LN-MAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089910/.

He:1983:CSC

[HYZ83] S. W. He, J. A. Yan, and W. A. Zheng. Sur la convergence des semimartingales continues dans R^{n} et des martingales dans une variété. (French) []. Lecture Notes in Mathematics, 986:179– 184, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068312/.

Harper:1984:SSM

[HZ84a] J. Harper and A. Zabrodsky. H-spaces and self-maps. Lecture Notes in Mathematics, 1051:371–373, 1984. CODEN LN-MAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075578/.

He:1984:RCM

[HZ84b] S. W. He and W. A. Zheng. Remarques sur la convergence des martingales dans les variétés. (French) []. Lecture Notes in Mathematics, 1059:174–178, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100042/.

Amari:1983:DGS

[iA83] Shun ichi Amari. Differential geometry of statistical inference. Lecture Notes in Mathematics, 1021:26–40, 1983. CODEN LN-MAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072900/.

Iooss:1981:SCB

[IACT81] G. Iooss, A. Arneodo, P. Coullet, and C. Tresser. Simple computation of bifurcating invariant circles for mappings. Lecture Notes in Mathematics, 898:192–211, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091914/.

Igusa:1982:ATR

[Igu82] Kiyoshi Igusa. On the algebraic K-theory of a_{∞}-ring spaces. Lecture Notes in Mathematics, 967:146–194, 1982. CODEN LN-MAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061902/.

Iitaka:1983:CTL

[Iit83] Shigeru Iitaka. Characterization of two lines on a projective plane. Lecture Notes in Mathematics, 1016:432–448, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099974/.

Inglot:1984:LLL

[IJ84] Tadeusz Inglot and Teresa Jurlewicz. Log log law for Gaussian random variables in Orlicz spaces. Lecture Notes in Mathematics, 1080:124–129, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099789/.

Iwaniec:1980:SDE

[IK80] Tadeusz Iwaniec and Ryszard Kopiecki. Stability in the differential equations for quasiregular mappings. Lecture Notes in Mathematics, 798:203–214, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097266/.

Ilamed:1980:RVP

[Ila80] Yehiel Ilamed. On realizations of vector products by polynomials which are identities for matrix rings. Lecture Notes in Mathematics, 792:68–72, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088664/.

Ilamed:1982:CPA

[Ila82] Yehiel Ilamed. On central polynomials and algebraic algebras. Lecture Notes in Mathematics, 951:88–93, 1982. CODEN LN-MAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067326/.

Iliadis:1984:RTS

[Ili84] S. D. Iliadis. On rim-type of spaces. Lecture Notes in Mathematics, 1060:45–54, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099920/.

Illusie:1983:FDK

[Ill83] L. Illusie. Finiteness, duality, and Künneth theorems in the cohomology of the de Rham–Witt complex. Lecture Notes in Mathematics, 1016:20–72, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099957/.

Imaoka:1983:SRF

[Ima83] Teruo Imaoka. Some remarks on fundamental regular *- semigroups. Lecture Notes in Mathematics, 998:270–280, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062036/.

Immink:1984:AAD

[Imm84a] Geertrui K. Immink. Asymptotics of Analytic Difference Equations, volume 1085 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13867-6 (print), 3-540-39060-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 134 pp. LCCN QA3 .L28 no. 1085. URL http:

//link.springer.com/book/10.1007/BFb0100174;   http:/
/www.springerlink.com/content/978-3-540-39060-2.

Immink:1984:EPR

[Imm84b] Geertrui K. Immink. Existence proofs for right inverses of difference operators. Lecture Notes in Mathematics, 1085:21–75, 1984. CODEN LNMAA2. ISBN 3-540-13867-6 (print), 3-540-39060-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100176/.

Immink:1984:LDE

[Imm84c] Geertrui K. Immink. Linear difference equations. Lecture Notes in Mathematics, 1085:1–20, 1984. CODEN LNMAA2. ISBN 3-540-13867-6 (print), 3-540-39060-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100175/.

Immink:1984:NDE

[Imm84d] Geertrui K. Immink. Nonlinear difference equations. Lecture Notes in Mathematics, 1085:76–125, 1984. CODEN LN-MAA2. ISBN 3-540-13867-6 (print), 3-540-39060-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100177/.

Inglot:1980:CTS

[Ing80] T. Inglot. Convergence of two-sample empirical processes. Lecture Notes in Mathematics, 828:102–107, 1980. CODEN LN-MAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097397/.

Iochum:1984:CAJ

[Ioc84a] Bruno Iochum. Cônes associés aux J. B. W. algèbres. (French) []. Lecture Notes in Mathematics, 1049:175–185, 1984. CODEN LNMAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071366/.

Iochum:1984:CAA

[Ioc84b] Bruno Iochum. Cônes autopolaires et algèbres de Jordan, volume 1049 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 247 pp. LCCN QA3 .L28 no. 1049. URL http:/

/link.springer.com/book/10.1007/BFb0071358;   http://
www.springerlink.com/content/978-3-540-38779-4.

Iochum:1984:CAFb

[Ioc84c] Bruno Iochum. Cônes autopolaires facialement homogènes. Lecture Notes in Mathematics, 1049:49–71, 1984. CODEN LN-MAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071361/.

Iochum:1984:CAFa

[Ioc84d] Bruno Iochum. Cônes autopolaires. (French) []. Lecture Notes in Mathematics, 1049:12–48, 1984. CODEN LNMAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071360/.

Iochum:1984:COF

[Ioc84e] Bruno Iochum. Cônes orientables. (French) [Orientable cones]. Lecture Notes in Mathematics, 1049:165–174, 1984. CODEN LN-MAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071365/.

Iochum:1984:IFI

[Ioc84f] Bruno Iochum. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 1049:1–11, 1984. CODEN LN-MAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071359/.

Iochum:1984:LJC

[Ioc84g] Bruno Iochum. L’algèbre de Jordan d’un cône autopolaire facialement homogène. (French) []. Lecture Notes in Mathematics, 1049:72–94, 1984. CODEN LNMAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071362/.

Iochum:1984:TJB

[Ioc84h] Bruno Iochum. Traces sur les J. B. algèbres. (French) []. Lecture Notes in Mathematics, 1049:122–164, 1984. CODEN LN-MAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071364/.

Iochum:1984:TCA

[Ioc84i] Bruno Iochum. Traces sur un cône autopolaire. (French) []. Lecture Notes in Mathematics, 1049:95–121, 1984. CODEN LN-MAA2. ISBN 3-540-12901-4 (print), 3-540-38779-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071363/.

Ioffe:1980:RTM

[Iof80] A. D. Ioffe. Representation theorems for measurable multifunctions. Lecture Notes in Mathematics, 794:141–152, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088220/.

Ionescu:1984:EPV

[Ion84] Paltin Ionescu. Embedded projective varieties of small invariants. Lecture Notes in Mathematics, 1056:142–186, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071773/.

Iooss:1980:BIT

[Ioo80] G. Iooss. Bifurcation of invariant tori in R^{3}. Lecture Notes in Mathematics, 782:192–200, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090433/.

Iooss:1984:BTT

[Ioo84] Gérard Iooss. Bifurcation and transition to turbulence in hydrodynamics. Lecture Notes in Mathematics, 1057:152–201, 1984. CODEN LNMAA2. ISBN 3-540-12931-6 (print), 3-540-38812-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098596/.

Iozzi:1983:SFS

[IP83a] Alessandra Iozzi and Massimo A. Picardello. Spherical functions on symmetric graphs. Lecture Notes in Mathematics, 992:344– 386, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069168/.

Ito:1983:PTM

[IP83b] Kiyosi Itô and Jurii V. Prokhorov, editors. Probability Theory and Mathematical Statistics: Proceedings of the Fourth USSR–Japan Symposium, held at Tbilisi, USSR, August 23–29, 1982, volume 1021 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 0-387-12718-6, 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1021; QA3 .L28 no. 1021 QA273.A1.

URL http://link.springer.com/book/10.1007/BFb0072896;
 http://www.springerlink.com/content/978-3-540-38701-
5.

Ibrahim:1981:RLD

[IS81a] R. Ibrahim and B. D. Sleeman. A regular left-definite eigenvalue problem with eigenvalue parameter in the boundary conditions. Lecture Notes in Mathematics, 846:158–167, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089834/.

Irwin:1981:NCS

[IS81b] John Irwin and Tom Snabb. A new class of subgroups of P_{χo}Z. Lecture Notes in Mathematics, 874:154–160, 1981. CODEN LN-MAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090532/.

Ischebeck:1982:ROD

[Isc82] Friedrich Ischebeck. On real one-dimensional cycles. Lecture Notes in Mathematics, 959:371–380, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062264/.

Iserles:1981:GOS

[Ise81] Arieh Iserles. Generalized order star theory. Lecture Notes in Mathematics, 888:228–238, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095589/.

Iserles:1984:OSS

[Ise84a] A. Iserles. Order stars and the structure of Padé tableaux. Lecture Notes in Mathematics, 1071:166–175, 1984. CODEN LN-MAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099617/.

Iserles:1984:OSC

[Ise84b] Arieh Iserles. Order stars, contractivity and a Pick-type theorem. Lecture Notes in Mathematics, 1105:117–124, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072404/.

Ishida:1983:HES

[Ish83] Masa-Nori Ishida. Hirzebruch’s examples of surfaces of general type with c_{1}^{2} = 3c_{2}. Lecture Notes in Mathematics, 1016:412– 431, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099973/.

Isles:1981:RNS

[Isl81] David Isles. Remarks on the notion of standard non-isomorphic natural number series. Lecture Notes in Mathematics, 873:111– 134, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090731/.

Ito:1980:PEE

[Itô80] Masayuki Itô. Positive eigen elements for an infinitesimal generator of a diffusion semi-group and their integral representations. Lecture Notes in Mathematics, 787:163–184, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086332/.

Ito:1982:DCC

[Itô82] Masayuki Itô. Divisible convex cones constituted by Hunt convolution kernels. Lecture Notes in Mathematics, 923:204–226, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093044/.

Ito:1984:NCT

[Itô84a] Masayuki Itô. Les noyaux de convolution de type logarithmique. (French) []. Lecture Notes in Mathematics, 1096:347–392, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100119/.

Ito:1984:LNC

[Itô84b] Masayuki Itô. L’existence des noyaux de convolution N sur R^{+} vérifiant Nϵ(T V) et N ∉ B. (French) []. Lecture Notes in Mathematics, 1061:220–229, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099026/.

Ito:1984:LDN

[Itô84c] Masayuki Itô. L’unicité de la décomposition des noyaux de convolution de Hunt. (French) []. Lecture Notes in Mathematics, 1061:192–210, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099024/.

Ito:1984:PCC

[Itô84d] Masayuki Itô. Sur le problème de Choquet concernant la TVinégalité. (French) []. Lecture Notes in Mathematics, 1061:211– 219, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099025/.

Ito:1984:DNC

[Itô84e] Masayuki Itô. Sur une decomposition des noyaux de convolution de Hunt. (French) []. Lecture Notes in Mathematics, 1061: 157–191, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099023/.

Ize:1980:RNI

[IV80] A. F. Izé and A. Ventura. Relationship in the neighbourhood of infinity and asymptotic equivalence of neutral functional differential equations. Lecture Notes in Mathematics, 799:217–251, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089316/.

Ivascu:1983:ECD

[Iva83] D. Ivaşcu. The exponent of convergence of a discontinuous Möbius group. Lecture Notes in Mathematics, 1013:100–112, 1983. CO-DEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066522/.

Ivanov:1984:BS

[Iva84] A. A. Ivanov. Bitopological spaces. Lecture Notes in Mathematics, 1060:55–66, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099921/.

Ivnitsky:1983:AER

[Ivn83] V. A. Ivnitsky. Accuracy estimation of the results of complex systems simulation with vector output and several types of randomnesses. Lecture Notes in Mathematics, 982:38–59, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082060/.

Ivrii:1984:AMP

[Ivr84a] Victor Ivriĭ. The asymptotics for miscellaneous problems. Lecture Notes in Mathematics, 1100:177–226, 1984. CODEN LN-MAA2. ISBN 3-540-13361-5 (print), 3-540-38911-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072207/.

Ivrii:1984:ASO

[Ivr84b] Victor Ivriĭ. The asymptotics for second-order operators. Lecture Notes in Mathematics, 1100:1–176, 1984. CODEN LN-MAA2. ISBN 3-540-13361-5 (print), 3-540-38911-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072206/.

Ivrii:1984:PSA

[Ivr84c] Victor Ivriĭ. Precise Spectral Asymptotics for Elliptic Operators Acting in Fiberings over Manifolds with Boundary, volume 1100 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13361-5 (print), 3-540-38911-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 238 pp. LCCN QA3 .L28 no. 1100. URL http:

//link.springer.com/book/10.1007/BFb0072205;   http:/
/www.springerlink.com/content/978-3-540-38911-8.

Ion:1982:ZMW

[IvW82] P. D. F. Ion and W. von Waldenfels. Zeitgeordnete Momente des Weissen klassischen und Des Weissen Quantenrauschens. (German) []. Lecture Notes in Mathematics, 928:212–235, 1982. CO-DEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093226/.

Iwanaga:1980:TEA

[IW80] Yasuo Iwanaga and Takayoshi Wakamatsu. Trivial extension of Artin algebras. Lecture Notes in Mathematics, 832:295–301, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088470/.

Iwaniec:1983:CRD

[Iwa83] Tadeusz Iwaniec. On Cauchy–Riemann derivatives in several real variables. Lecture Notes in Mathematics, 1039:220–244, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073367/.

Ize:1980:FDE

[Izé80] Antonio Fernandes Izé, editor. Functional Differential Equations and Bifurcation: Proceedings of a Conference Held at São Carlos, Brazil, July 2–7, 1979, volume 799 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 799. URL http:

//link.springer.com/book/10.1007/BFb0089305;   http:/
/www.springerlink.com/content/978-3-540-39251-4.

Ize:1982:FPI

[Izé82a] A. F. Izé. On a fixed point index method for the analysis of the asymptotic behavior and boundary value problems of process and semidynamical systems. Lecture Notes in Mathematics, 957:126–144, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066237/.

Ize:1982:IBT

[Ize82b] Jorge Ize. Introduction to bifurcation theory. Lecture Notes in Mathematics, 957:145–202, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066238/.

Jaco:1980:EGJ

[Jac80a] William Jaco. On ends of groups and Johannson’s deformation theorem. Lecture Notes in Mathematics, 788:62–87, 1980. CO-DEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099239/.

Jacod:1980:ISP

[Jac80b] Jean Jacod. Intégrales stochastiques par rapport à une semimartingale vectorielle et changements de filtration. (French) []. Lecture Notes in Mathematics, 784:161–172, 1980. CODEN LN-MAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089482/.

Jacob:1981:CSP

[Jac81] Pierre Jacob. Convergences stochastiques des processus ponctuels composes à signe. (French) []. Lecture Notes in Mathematics, 861: 74–82, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097316/.

Jacobsen:1982:MCA

[Jac82a] Lisa Jacobsen. A method for convergence acceleration of continued fractions K(a_{n}/1). Lecture Notes in Mathematics, 932: 74–86, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093307/.

Jacobsen:1982:SPS

[Jac82b] Lisa Jacobsen. Some periodic sequences of circular convergence regions. Lecture Notes in Mathematics, 932:87–98, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093308/.

Jacka:1983:LTI

[Jac83a] S. D. Jacka. A local time inequality for martingales. Lecture Notes in Mathematics, 986:106–116, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068304/.

Jacquet:1983:RS

[Jac83b] Hervé Jacquet. On the residual spectrum of GL(n). Lecture Notes in Mathematics, 1041:185–208, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073148/.

Jacod:1984:GSP

[Jac84] Jean Jacod. Une généralisation des semimartingales: Les processus admettant un processus à accroissements indépendants tangent. (French) []. Lecture Notes in Mathematics, 1059:91– 118, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100035/.

Jadczyk:1980:CSC

[Jad80] A. Z. Jadczyk. Conformal structures and connections. Lecture Notes in Mathematics, 836:202–210, 1980. CODEN LN-MAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089736/.

Jager:1984:NTN

[Jag84] Hendrik Jager, editor. Number Theory Noordwijkerhout 1983: Proceedings of the Journées Arithmétiques held at Noordwijkerhout, The Netherlands July 11–15, 1983, volume 1068 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1068.

URL http://link.springer.com/book/10.1007/BFb0099436;
 http://www.springerlink.com/content/978-3-540-38906-
4.

Jajte:1980:DMH

[Jaj80] R. Jajte. V -decomposable measures on Hilbert spaces. Lecture Notes in Mathematics, 828:108–127, 1980. CODEN LN-MAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097398/.

Jajte:1984:FRA

[Jaj84a] R. Jajte. A few remarks on the almost uniform ergodic theorems in von Neumann algebras. Lecture Notes in Mathematics, 1080:130–143, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099790/.

Jajte:1984:NCS

[Jaj84b] R. Jajte. Non-commutative subadditive ergodic theorems for semifinite von Neumann algebras. Lecture Notes in Mathematics, 1064:219–234, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073644/.

Jakobczak:1980:DPS

[Jak80a] Piotr Jakóbczak. Decomposition property of A^{k}(D) in strictly pseudoconvex domains. Lecture Notes in Mathematics, 798:215– 223, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097267/.

Jakobson:1980:CIM

[Jak80b] M. V. Jakobson. Construction of invariant measures absolutely continuous with respect to dx for some maps of the interval. Lecture Notes in Mathematics, 819:246–257, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086991/.

Jakobsen:1982:GTA

[Jak82] H. P. Jakobsen. Group theoretical aspects of the chronometric theory. Lecture Notes in Mathematics, 905:165–169, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092435/.

Jakobczak:1983:DTB

[Jak83] Piotr Jakóbczak. The decomposition theorems in the bidisc. Lecture Notes in Mathematics, 1039:245–256, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073368/.

James:1983:EPU

[Jam83] Robert C. James. Extreme points and an unusual Banach space. Lecture Notes in Mathematics, 991:111–123, 1983. CODEN LN-MAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061563/.

James:1984:TCD

[Jam84] R. D. James. Theory for the cold-drawing of polymers. Lecture Notes in Mathematics, 1063:143–161, 1984. CODEN LN-MAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072155/.

Janssen:1982:SZO

[Jan82] Arnold Janssen. Some zero-one laws for semistable and selfdecomposable measures on locally convex spaces. Lecture Notes in Mathematics, 928:236–246, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093227/.

Janssen:1983:AQD

[Jan83a] K. Janßen. Applications of quasi Dirichlet bounded harmonic functions. Lecture Notes in Mathematics, 1014:204–219, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072080/.

Janssen:1983:SCB

[Jan83b] K. Janßen. Standard H-cones and balayage spaces. Lecture Notes in Mathematics, 1014:197–203, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072079/.

Jannsen:1984:SGG

[Jan84] Uwe Jannsen. On the structure of Galois groups as Galois modules. Lecture Notes in Mathematics, 1068:109–126, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099446/.

Jaworowski:1981:FPM

[Jaw81] Jan Jaworowski. Fibre preserving maps of sphere bundles into vector space bundles. Lecture Notes in Mathematics, 886:154– 162, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092181/.

Jaworski:1983:CRB

[Jaw83] Jerzy Jaworski. On the connectedness of a random bipartite mapping. Lecture Notes in Mathematics, 1018:69–74, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071614/.

Joseph:1981:LBP

[JB81] D. D. Joseph and K. Burns. Lectures on bifurcation from periodic orbits. Lecture Notes in Mathematics, 898:1–12, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091904/.

Jeffrey:1980:FFN

[Jef80] Alan Jeffrey. Far fields, nonlinear evolution equations, the Bäcklund transformation and inverse scattering. Lecture Notes in Mathematics, 810:118–134, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089977/.

Jeffrey:1981:STN

[Jef81] Alan Jeffrey. Some topics in nonlinear wave propagation. Lecture Notes in Mathematics, 846:168–184, 1981. CODEN LN-MAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089835/.

Jennane:1980:GCF

[Jen80] B. Jennane. Groupes de cohomologie d’un fibré holomorphe à base et à fibre de Stein. (French) []. Lecture Notes in Mathematics, 822:100–108, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097752/.

Jennings:1983:BSE

[Jen83] Alan Jennings. On bounds for symmetric eigenvalue problems. Lecture Notes in Mathematics, 973:195–205, 1983. CODEN LN-MAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062103/.

Jerome:1980:HLA

[Jer80] Joseph W. Jerome. Horizontal line analysis of the multidimensional porous medium equation: Existence, rate of convergence and maximum principles. Lecture Notes in Mathematics, 773: 64–82, 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094164/.

Jeulin:1980:A

[Jeu80a] Thierry Jeulin. Applications. Lecture Notes in Mathematics, 833: 89–131, 1980. CODEN LNMAA2. ISBN 3-540-10265-5 (print), 3-540-38390-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093545/.

Jeulin:1980:GAV

[Jeu80b] Thierry Jeulin. Grossissement à l’aide de variables honnêtes. (French) []. Lecture Notes in Mathematics, 833:73–88, 1980. CO-DEN LNMAA2. ISBN 3-540-10265-5 (print), 3-540-38390-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093544/.

Jeulin:1980:GIF

[Jeu80c] Thierry Jeulin. Grossissement initial. (French) [Initial magnification]. Lecture Notes in Mathematics, 833:26–60, 1980. CODEN LNMAA2. ISBN 3-540-10265-5 (print), 3-540-38390-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093542/.

Jeulin:1980:GPF

[Jeu80d] Thierry Jeulin. Grossissement progressif. (French) [Progressive magnification]. Lecture Notes in Mathematics, 833:61–72, 1980. CODEN LNMAA2. ISBN 3-540-10265-5 (print), 3-540-38390-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093543/.

Jeulin:1980:PFP

[Jeu80e] Thierry Jeulin. Préliminaires. (French) [Preliminaries]. Lecture Notes in Mathematics, 833:1–11, 1980. CODEN LNMAA2. ISBN 3-540-10265-5 (print), 3-540-38390-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093540/.

Jeulin:1980:RGH

[Jeu80f] Thierry Jeulin. Résultats généraux sur l’hypothèse H′. (French) []. Lecture Notes in Mathematics, 833:12–25, 1980. CODEN LNMAA2. ISBN 3-540-10265-5 (print), 3-540-38390-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093541/.

Jeulin:1980:SMG

[Jeu80g] Thierry Jeulin. Semi-Martingales et Grossissement d’une Filtration. (French) [Semi-Martingales and enlargment of a filtration], volume 833 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10265-5 (print), 3-540-38390-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 142 pp. LCCN QA3 .L28 no. 833. URL http:

//link.springer.com/book/10.1007/BFb0093539;   http:/
/www.springerlink.com/content/978-3-540-38390-1.

Jezek:1983:JIE

[Jež83] Jaroslav Ježek. On join-indecomposable equational theories. Lecture Notes in Mathematics, 1004:159–165, 1983. CODEN LN-MAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063436/.

Jambu-Giraudet:1981:IAG

[JG81] Michèle Jambu-Giraudet. Interprétations d’arithmétiques dans des groupes et des treillis. (French) []. Lecture Notes in Mathematics, 890:143–153, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095662/.

Jiang:1981:FPC

[Jia81] Boju Jiang. Fixed point classes from a differential viewpoint. Lecture Notes in Mathematics, 886:163–170, 1981. CODEN LN-MAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092182/.

Jongen:1980:CNM

[JJT80] H. Th. Jongen, P. Jonker, and F. Twilt. The continuous Newtonmethod for meromorphic functions. Lecture Notes in Mathematics, 810:181–239, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089981/.

Jezek:1983:IFC

[JK83] Jaroslav Ježek and Tomáš Kepka. Ideal-free CIM-groupoids and open convex sets. Lecture Notes in Mathematics, 1004:166– 175, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063437/.

Jensen:1980:MTR

[JL80] C. U. Jensen and H. Lenzing. Model theory and representations of algebras. Lecture Notes in Mathematics, 832:302–310, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088471/.

Jacod:1981:WSS

[JM81a] Jean Jacod and Jean Mémin. Weak and strong solutions of stochastic differential equations: Existence and stability. Lecture Notes in Mathematics, 851:169–212, 1981. CODEN LN-MAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088728/.

Jain:1981:GMC

[JM81b] Naresh C. Jain and Ditlev Monrad. Gaussian measures in certain function spaces. Lecture Notes in Mathematics, 860:246– 256, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090619/.

Jorsboe:1982:CHT

[JM82a] Ole Groth Jørsboe and Leif Mejlbro. The Carleson–Hunt Theorem on Fourier Series, volume 911 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11198-0 (print), 3-540-39006-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 122 pp. LCCN QA3 .L28 no. 911.

URL http://link.springer.com/book/10.1007/BFb0094072;
 http://www.springerlink.com/content/978-3-540-39006-
0.

Jorsboe:1982:C

[JM82b] Ole Groth Jørsboe and Leif Mejlbro. Chapter I. Lecture Notes in Mathematics, 911:3–21, 1982. CODEN LNMAA2. ISBN 3-540-11198-0 (print), 3-540-39006-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094073/.

Jorsboe:1982:CIa

[JM82c] Ole Groth Jørsboe and Leif Mejlbro. Chapter II. Lecture Notes in Mathematics, 911:23–43, 1982. CODEN LNMAA2. ISBN 3-540-11198-0 (print), 3-540-39006-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094074/.

Jorsboe:1982:CIb

[JM82d] Ole Groth Jørsboe and Leif Mejlbro. Chapter III. Lecture Notes in Mathematics, 911:45–68, 1982. CODEN LNMAA2. ISBN 3-540-11198-0 (print), 3-540-39006-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094075/.

Jorsboe:1982:CIc

[JM82e] Ole Groth Jørsboe and Leif Mejlbro. Chapter IV. Lecture Notes in Mathematics, 911:69–121, 1982. CODEN LNMAA2. ISBN 3-540-11198-0 (print), 3-540-39006-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094076/.

Jackson:1983:PWO

[JM83] Steve Jackson and Donald A. Martin. Pointclasses and wellordered unions. Lecture Notes in Mathematics, 1019:56–66, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071693/.

Jacobsen:1984:CLP

[JM84] Lisa Jacobsen and Arne Magnus. On the convergence of limit periodic continued fractions K(a_{n}/1), where a_{1} → −1/4. Lecture Notes in Mathematics, 1105:243–248, 1984. CODEN LN-MAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072415/.

Jeltsch:1982:LBA

[JN82] Rolf Jeltsch and Olavi Nevanlinna. Lower bounds for the accuracy of linear multistep methods. Lecture Notes in Mathematics, 968:280–291, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064895/.

Jockusch:1981:TEC

[Joc81] Carl G. Jockusch, Jr. Three easy constructions of recursively enumerable sets. Lecture Notes in Mathematics, 859:83–91, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090941/.

Joffe:1983:NAN

[Jof83] Alexander D. Joffe. New applications of nonsmooth analysis to nonsmooth optimization. Lecture Notes in Mathematics, 979:178– 201, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066254/.

Johnson:1981:TRC

[Joh81a] K. W. Johnson. Transversals, S-rings and centraliser rings of groups. Lecture Notes in Mathematics, 848:169–177, 1981. CO-DEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090564/.

Johnstone:1981:IT

[Joh81b] Peter T. Johnstone. Injective toposes. Lecture Notes in Mathematics, 871:284–297, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089912/.

Johnstone:1981:SAS

[Joh81c] Peter T. Johnstone. Scott is not always sober. Lecture Notes in Mathematics, 871:282–283, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089911.

Johnstone:1982:FTG

[Joh82] Peter T. Johnstone. Factorization theorems for geometric morphisms, II. Lecture Notes in Mathematics, 915:216–233, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092882/.

John:1983:TPN

[Joh83a] K. John. Tensor products and nuclearity. Lecture Notes in Mathematics, 991:124–129, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061564/.

Johnson:1983:ACM

[Joh83b] R. Johnson. Application of Carleson measures to partial differential equations and Fourier multiplier problems. Lecture Notes in Mathematics, 992:16–72, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069150/.

Johnson:1984:ABG

[Joh84] D. L. Johnson. Analogues of the braid group. Lecture Notes in Mathematics, 1098:63–68, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099662/.

Jondrup:1980:HPA

[Jøn80] S. Jøndrup. Hereditary P. I. algebras. Lecture Notes in Mathematics, 825:63–75, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089122/.

Jones:1981:OOT

[Jon81] Gary D. Jones. An ordering of oscillation types for y^{(n)} + py = 0. Lecture Notes in Mathematics, 846:185–192, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089836/.

Jones:1982:FPT

[Jon82] Marsha Finkel Jones. Flatness and f-projectivity of torsion-free modules and injective modules. Lecture Notes in Mathematics, 951:94–116, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067327/.

Jones:1984:IB

[Jon84] Peter W. Jones. On interpolation between H^{1} and H^{∞}. Lecture Notes in Mathematics, 1070:143–151, 1984. CODEN LN-MAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099098/.

Joseph:1980:DRB

[Jos80] Daniel D. Joseph. Direct and repeated bifurcation into turbulence. Lecture Notes in Mathematics, 771:249–271, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086911/.

Joseph:1981:KPG

[Jos81] A. Joseph. Kostant’s problem and Goldie rank. Lecture Notes in Mathematics, 880:249–266, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090412/.

Joshi:1982:OFI

[Jos82] V. S. Joshi. Order-free integers (modm). Lecture Notes in Mathematics, 938:93–100, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097176/.

Joseph:1983:CFC

[Jos83a] A. Joseph. Completion functors in the Q category. Lecture Notes in Mathematics, 1020:80–106, 1983. CODEN LNMAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071498/.

Joseph:1983:CPI

[Jos83b] A. Joseph. On the classification of primitive ideals in the enveloping algebra of a semisimple Lie algebra. Lecture Notes in Mathematics, 1024:30–76, 1983. CODEN LNMAA2. ISBN 3-540-12725-9 (print), 3-540-38709-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071431/.

Jost:1984:PEA

[Jos84a] Jürgen Jost. C^{1,α}-a priori estimates for arbitrary domains. Nonvariational existence proofs. Lecture Notes in Mathematics, 1062: 90–98, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0121266/.

Jost:1984:PE

[Jos84b] Jürgen Jost. A-priori C^{1,α}-estimates. Lecture Notes in Mathematics, 1062:65–77, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0121263/.

Jost:1984:PEB

[Jos84c] Jürgen Jost. A-priori estimates from below for the functional determinant of harmonic diffeomorphisms. Lecture Notes in Mathematics, 1062:77–86, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0121264/.

Jost:1984:AHM

[Jos84d] Jürgen Jost. Applications of harmonic maps between surfaces. Lecture Notes in Mathematics, 1062:112–126, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0121269/.

Jost:1984:CM

[Jos84e] Jürgen Jost. Conformal mappings. Lecture Notes in Mathematics, 1062:18–38, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0121260/.

Jost:1984:EHDb

[Jos84f] Jürgen Jost. The existence of harmonic diffeomorphisms between surfaces. Lecture Notes in Mathematics, 1062:106–112, 1984. CO-DEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0121268/.

Jost:1984:EHDa

[Jos84g] Jürgen Jost. The existence of harmonic diffeomorphisms which solve a Dirichlet problem. Lecture Notes in Mathematics, 1062: 86–90, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0121265/.

Jost:1984:ETH

[Jos84h] Jürgen Jost. Existence theorems for harmonic maps between surfaces. Lecture Notes in Mathematics, 1062:38–53, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0121261/.

Jost:1984:GC

[Jos84i] Jürgen Jost. Geometric considerations. Lecture Notes in Mathematics, 1062:12–18, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0121259/.

Jost:1984:HCP

[Jos84j] Jürgen Jost. Harmonic coordinates. C^{2,α}-a-priori estimates for harmonic maps. Lecture Notes in Mathematics, 1062:98–105, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0121267/.

Jost:1984:HMB

[Jos84k] Jürgen Jost. Harmonic Maps Between Surfaces (with a Special Chapter on Conformal Mappings), volume 1062 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 133 pp. LCCN QA3 .L28 no. 1062. URL http:/

/link.springer.com/book/10.1007/BFb0100160;   http://
www.springerlink.com/content/978-3-540-38868-5.

Jost:1984:I

[Jos84l] Jürgen Jost. Introduction. Lecture Notes in Mathematics, 1062: 1–12, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100161/.

Jost:1984:UT

[Jos84m] Jürgen Jost. Uniqueness theorems. Lecture Notes in Mathematics, 1062:54–65, 1984. CODEN LNMAA2. ISBN 3-540-13339-9 (print), 3-540-38868-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0121262/.

Joussen:1980:CAO

[Jou80] J. Joussen. On the construction of Archimedean orders of a free plane. Lecture Notes in Mathematics, 792:73–78, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088665/.

Journe:1983:W

[Jou83a] Jean-Lin Journé. A_{p} weights. Lecture Notes in Mathematics, 994: 16–28, 1983. CODEN LNMAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061461/.

Journe:1983:BMO

[Jou83b] Jean-Lin Journé. BMO(R^{n}, dx). Lecture Notes in Mathematics, 994:29–45, 1983. CODEN LNMAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061462/.

Journe:1983:CZOa

[Jou83c] Jean-Lin Journé. Calderón–Zygmund operators. Lecture Notes in Mathematics, 994:46–63, 1983. CODEN LNMAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061463/.

Journe:1983:CZOc

[Jou83d] Jean-Lin Journé. Calderón–Zygmund operators and Littlewood– Paley theory. Lecture Notes in Mathematics, 994:78–92, 1983. CODEN LNMAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061465/.

Journe:1983:CZOb

[Jou83e] Jean-Lin Journé. Calderón–Zygmund operators and pseudodifferential operators. Lecture Notes in Mathematics, 994:64– 77, 1983. CODEN LNMAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061464/.

Journe:1983:CZO

[Jou83f]  Jean-Lin Journé.  Calderón–Zygmund Operators, Pseudo-

Differential Operators and the Cauchy Integral of Calderón, volume 994 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 128 pp. LCCN QA3 .L28 no. 994. URL http:

//link.springer.com/book/10.1007/BFb0061458;   http:/
/www.springerlink.com/content/978-3-540-40033-2.

Journe:1983:CKL

[Jou83g] Jean-Lin Journé. The Cauchy kernel on Lipschitz curves. Lecture Notes in Mathematics, 994:93–109, 1983. CODEN LN-MAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061466/.

Journe:1983:HLM

[Jou83h] Jean-Lin Journé. The Hardy–Littlewood maximal operator. Lecture Notes in Mathematics, 994:7–15, 1983. CODEN LN-MAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061460/.

Journe:1983:P

[Jou83i] Jean-Lin Journé. Preliminaries on L_{B}^{P}(R^{n}, dμ). Lecture Notes in Mathematics, 994:1–6, 1983. CODEN LNMAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061459/.

Journe:1983:STG

[Jou83j] Jean-Lin Journé. Some techniques to generate new Calderón– Zygmund operators. Lecture Notes in Mathematics, 994:110– 122, 1983. CODEN LNMAA2. ISBN 3-540-12313-X (print), 3-540-40033-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061467/.

Jovanovic:1984:SCP

[Jov84] A. Jovanović. Some combinatorial properties of measures. Lecture Notes in Mathematics, 1089:75–78, 1984. CODEN LN-MAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072603/.

Jensen:1981:STM

[JP81] Ronald Björn Jensen and Alexander Prestel, editors. Set Theory and Model Theory: Proceedings of an Informal Symposium Held at Bonn, June 1–3, 1979, volume 872 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10849-1 (print), 3-540-38757-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 872. URL http:

//link.springer.com/book/10.1007/BFb0098616;   http:/
/www.springerlink.com/content/978-3-540-38757-2.

Janson:1984:HI

[JP84a] Svante Janson and Jaak Peetre. Harmonic interpolation. Lecture Notes in Mathematics, 1070:92–124, 1984. CODEN LN-MAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099096/.

Janson:1984:HOC

[JP84b] Svante Janson and Jaak Peetre. Higher order commutators of singular integral operators. Lecture Notes in Mathematics, 1070: 125–142, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099097/.

Jurgensen:1981:SPC

[JPW81] Helmut Jürgensen, Mario Petrich, and Hanns Joachim Weinert, editors. Semigroups: Proceedings of a Conference Held at Oberwolfach, Germany December 16–21, 1978, volume 855 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 855.

URL http://link.springer.com/book/10.1007/BFb0089803;
 http://www.springerlink.com/content/978-3-540-38656-
8.

Jones:1982:UTC

[JR82] William B. Jones and Walter M. Reid. Uniform twin-convergence regions for continued fractions K(a_{n}/1). Lecture Notes in Mathematics, 932:106–128, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093310/.

James:1980:EHT

[JS80] I. M. James and G. B. Segal. On equivariant homotopy theory. Lecture Notes in Mathematics, 788:316–330, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099251/.

Jiang:1981:FPS

[JS81] Boju Jiang and Helga Schirmer. Fixed point sets of continuous selfmaps on polyhedra. Lecture Notes in Mathematics, 886: 171–177, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092183/.

Jagannathan:1982:NTP

[JS82a] R. Jagannathan and T. S. Santhanam. On a number of theoretical problems involved in the study of the physics of spinsystems. Lecture Notes in Mathematics, 938:82–92, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097175/.

Jodeit:1982:ESR

[JS82b] Max Jodeit, Jr. and Richard K. Shaw. Extension of Stein’s L logL result to the closed interval. Lecture Notes in Mathematics, 908: 300–310, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093297/.

Jones:1982:DFC

[JS82c] William B. Jones and Allan Steinhardt. Digital filters and continued fractions. Lecture Notes in Mathematics, 932:129–151, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093311/.

Jones:1984:ASF

[JS84] William B. Jones and Allan Steinhardt. Applications of Schur fractions to digital filtering and signal processing. Lecture Notes in Mathematics, 1105:210–226, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072413/.

Jones:1982:SCF

[JT82] William B. Jones and W. J. Thron. Survey of continued fraction methods of solving moment problems and related topics. Lecture Notes in Mathematics, 932:4–37, 1982. CO-DEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093302/;
 http://link.springer.com/chapter/10.1007/BFb0093303/.

Jones:1982:ATC

[JTW82] William B. Jones, W. J. Thron, and Haakon Waadeland, editors. Analytic Theory of Continued Fractions: Proceedings of a Seminar–Workshop held at Loen, Norway, 1981, volume 932 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 932. URL http:/

/link.springer.com/book/10.1007/BFb0093301;   http://
www.springerlink.com/content/978-3-540-39276-7.

Janson:1983:ECM

[JTW83] Svante Janson, Mitchell Taibleson, and Guido Weiss. Elementary characterizations of the Morrey–Campanato spaces. Lecture Notes in Mathematics, 992:101–114, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069154/.

Junek:1983:LOS

[Jun83] H. Junek. Linear operators in (F)-spaces. Lecture Notes in Mathematics, 991:130–136, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061565/.

Junnila:1984:SZD

[Jun84] Heikki J. K. Junnila. On strongly zero-dimensional F_{σ}-metrizable stratifiable spaces. Lecture Notes in Mathematics, 1060:67–75, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099922/.

Jurek:1980:SPM

[Jur80] Zbigniew J. Jurek. On stability of probability measures in Euclidean spaces. Lecture Notes in Mathematics, 828:128–145, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097399/.

Jurek:1981:CTS

[Jur81] Zbigniew J. Jurek. Convergence of types, selfdecomposability and stability of measures on linear spaces. Lecture Notes in Mathematics, 860:257–267, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090620/.

Jurchescu:1983:EMF

[Jur83a] M. Jurchescu. Espaces mixtes. (French) []. Lecture Notes in Mathematics, 1014:37–57, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072069/.

Jurgensen:1983:TDV

[Jür83b] H. Jürgensen. Total disjunktive verallgemeinerte Bruck–Reilly– Erweiterungen von Halbgruppen und formale Sprachen. (German) [Total generalized disjunctive Bruck–Reilly extensions of semigroups and formal languages]. Lecture Notes in Mathematics, 998: 281–309, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062037/.

Jungnickel:1982:CTP

[JV82] Dieter Jungnickel and Klaus Vedder, editors. Combinatorial Theory: Proceedings of a Conference Held at Schloss Rauischholzhausen, May 6–9, 1982, volume 969 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 969.

URL http://link.springer.com/book/10.1007/BFb0062982;
 http://www.springerlink.com/content/978-3-540-39380-

Janssen:1984:FST

[JvdS84a] A. J. E. M. Janssen and P. van der Steen. The fundamental stations and their connections. Lecture Notes in Mathematics, 1078: 34–65, 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098944/.

Janssen:1984:FDT

[JvdS84b] A. J. E. M. Janssen and P. van der Steen. Further development of the theory of integration. Lecture Notes in Mathematics, 1078: 66–94, 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098945/.

Janssen:1984:GI

[JvdS84c] A. J. E. M. Janssen and P. van der Steen. General introduction. Lecture Notes in Mathematics, 1078:1–4, 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098941/.

Janssen:1984:ILC

[JvdS84d] A. J. E. M. Janssen and P. van der Steen. Integration on locally compact Hausdorff spaces. Lecture Notes in Mathematics, 1078:124–145, 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098947/.

Janssen:1984:IMS

[JvdS84e] A. J. E. M. Janssen and P. van der Steen. Integration on measure spaces. Lecture Notes in Mathematics, 1078:95–123, 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098946/.

Janssen:1984:IPT

[JvdS84f] A. J. E. M. Janssen and P. van der Steen. Integration on the product of two spaces. Lecture Notes in Mathematics, 1078:185– 216, 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098949/.

Janssen:1984:IT

[JvdS84g] A. J. E. M. Janssen and P. van der Steen. Integration Theory, volume 1078 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 224 pp. LCCN QA3 .L28 no. 1078. URL http:

//link.springer.com/book/10.1007/BFb0098940;   http:/
/www.springerlink.com/content/978-3-540-38937-8.

Janssen:1984:N

[JvdS84h] A. J. E. M. Janssen and P. van der Steen. Notation. Lecture Notes in Mathematics, 1078:5, 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098942.

Janssen:1984:P

[JvdS84i] A. J. E. M. Janssen and P. van der Steen. Preliminaries. Lecture Notes in Mathematics, 1078:6–33, 1984. CODEN LN-MAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098943/.

Janssen:1984:RNT

[JvdS84j] A. J. E. M. Janssen and P. van der Steen. The Radon–Nikodým theorem and its classical origin. Lecture Notes in Mathematics, 1078:146–184, 1984. CODEN LNMAA2. ISBN 3-540-13386-0 (print), 3-540-38937-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098948/.

Jacobsen:1982:SUF

[JW82] Lisa Jacobsen and Haakon Waadeland. Some useful formulas involving tails of continued fractions. Lecture Notes in Mathematics, 932:99–105, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093309/.

Jurgensen:1983:PPA

[JW83] Helmut Jürgensen and Hanns Joachim Weinert. Preface to the papers on the algebraic and combinatorial theories of semigroups, delivered at the Conference on semigroups in Oberwolfach 1981, May 24 through May 30. Lecture Notes in Mathematics, 998: 252–254, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062033/.

Jacobsen:1984:MCF

[JW84] Lisa Jacobsen and Haakon Waadeland. Modification of continued fractions. Lecture Notes in Mathematics, 1071:176–196, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099618/.

Jeulin:1981:DCF

[JY81] T. Jeulin and M. Yor. Sur les distributions de certaines fonctionnelles du mouvement brownien. (French) []. Lecture Notes in Mathematics, 850:210–226, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088370/.

Kac:1980:SRR

[Kač80] Victor G. Kač. Some remarks on representations of quivers and infinite root systems. Lecture Notes in Mathematics, 832:311– 327, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088472/.

Kac:1981:SLG

[Kac81] V. G. Kac. Simple Lie groups and the Legendre symbol. Lecture Notes in Mathematics, 848:110–123, 1981. CODEN LN-MAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090559/.

Kac:1982:SPI

[Kac82] Victor G. Kac. Some problems on infinite dimensional Lie algebras and their representations. Lecture Notes in Mathematics, 933:117–126, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093356/.

Kac:1983:RSR

[Kac83] Victor G. Kac. Root systems, representations of quivers and invariant theory. Lecture Notes in Mathematics, 996:74–108, 1983. CODEN LNMAA2. ISBN 3-540-12319-9 (print), 3-540-40043-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063236/.

Kaagstrom:1983:CKC

[Kåg83] Bo Kågström. On computing the Kronecker canonical form of regular (A − λB)-pencils. Lecture Notes in Mathematics, 973:30– 57, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062093/.

Kahane:1980:TPH

[Kah80] J.-P. Kahane. Sur le treizième problème de Hilbert, le théorème de superposition de Kolmogorov et les sommes algébriques d’arcs croissants. (French) []. Lecture Notes in Mathematics, 781:76– 101, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097649/.

Kahn:1982:SPI

[Kah82] Peter J. Kahn. Steenrod’s problem and k-invariants of certain classifying spaces. Lecture Notes in Mathematics, 967:195– 214, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061903/.

Kahane:1984:DCD

[Kah84] Jean-Pierre Kahane. Dimension capacitaire et dimension de Hausdorff. (French) []. Lecture Notes in Mathematics, 1096:393– 400, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100120/.

Kaiser:1980:HGT

[Kai80a] Gerald Kaiser. Holomorphic gauge theory. Lecture Notes in Mathematics, 775:122–144, 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092024/.

Kaiser:1980:RQT

[Kai80b] Gerald Kaiser. Relativistic quantum theory in complex spacetime. Lecture Notes in Mathematics, 836:171–176, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089733/.

Kaijser:1983:SMP

[Kai83] Sten Kaijser. A simple-minded proof of the Pisier–Grothendieck inequality. Lecture Notes in Mathematics, 995:33–55, 1983. CO-DEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061887/.

Kakuda:1981:PIT

[Kak81] Yuzuru Kakuda. Precipitousness of the ideal of thin sets on a measurable cardinal. Lecture Notes in Mathematics, 891:49–56, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090978/.

Kalla:1980:OFI

[Kal80] Shyam Lal Kalla. Operators of fractional integration. Lecture Notes in Mathematics, 798:258–280, 1980. CODEN LN-MAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097270/.

Kalka:1982:HMK

[Kal82] M. Kalka. Harmonic maps in Kähler geometry and deformation theory. Lecture Notes in Mathematics, 949:75–86, 1982. CODEN LNMAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069757/.

Kalashnikov:1983:CMF

[Kal83] V. V. Kalashnikov. A complete metric in the function space D[0, ∞) and its application. Lecture Notes in Mathematics, 982: 60–76, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082061/.

Kamin:1982:ESP

[Kam82] S. Kamin. On elliptic singular perturbation problems with several turning points. Lecture Notes in Mathematics, 942:145–149, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094745/.

Kanemaki:1980:HMA

[Kan80a] Shôji Kanemaki. Hermitian manifolds with almost product structures. Lecture Notes in Mathematics, 798:281–291, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097271/.

Kaniel:1980:AHE

[Kan80b] Shmuel Kaniel. Approximation of the hydrodynamic equations by a transport process. Lecture Notes in Mathematics, 771:272– 286, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086912/.

Kanda:1982:NPS

[Kan82] Mamoru Kanda. Notes on polar sets for Lévy processes on the line. Lecture Notes in Mathematics, 923:227–234, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093045/.

Kantorovitz:1983:CO

[Kan83a] Shmuel Kantorovitz. Convolution operators in L^{p}. Lecture Notes in Mathematics, 1012:49–58, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064296/.

Kantorovitz:1983:E

[Kan83b] Shmuel Kantorovitz. Examples. Lecture Notes in Mathematics, 1012:8–14, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064291/.

Kantorovitz:1983:F

[Kan83c] Shmuel Kantorovitz. The family S +ζV. Lecture Notes in Mathematics, 1012:38–48, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064295/.

Kantorovitz:1983:FU

[Kan83d] Shmuel Kantorovitz. The family S + ζV, S unbounded. Lecture Notes in Mathematics, 1012:82–98, 1983. CODEN LN-MAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064300/.

Kantorovitz:1983:FR

[Kan83e] Shmuel Kantorovitz. First reduction. Lecture Notes in Mathematics, 1012:15–19, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064292/.

Kantorovitz:1983:I

[Kan83f] Shmuel Kantorovitz. Introduction. Lecture Notes in Mathematics, 1012:1–5, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064289/.

Kantorovitz:1983:LA

[Kan83g] Shmuel Kantorovitz. Local analysis. Lecture Notes in Mathematics, 1012:146–170, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064303/.

Kantorovitz:1983:OC

[Kan83h] Shmuel Kantorovitz. Operational calculus. Lecture Notes in Mathematics, 1012:6–7, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064290.

Kantorovitz:1983:SR

[Kan83i] Shmuel Kantorovitz. Second reduction. Lecture Notes in Mathematics, 1012:20–24, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064293/.

Kantorovitz:1983:S

[Kan83j] Shmuel Kantorovitz. Similarity. Lecture Notes in Mathematics, 1012:65–72, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064298/.

Kantorovitz:1983:SC

[Kan83k] Shmuel Kantorovitz. Similarity (continued). Lecture Notes in Mathematics, 1012:99–122, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064301/.

Kantorovitz:1983:SO

[Kan83l] Shmuel Kantorovitz. Singular C^{n}-operators. Lecture Notes in Mathematics, 1012:123–145, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064302/.

Kantorovitz:1983:SRS

[Kan83m] Shmuel Kantorovitz. Some regular semigroups. Lecture Notes in Mathematics, 1012:59–64, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064297/.

Kantorovitz:1983:SA

[Kan83n] Shmuel Kantorovitz. Spectral analysis. Lecture Notes in Mathematics, 1012:73–81, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064299/.

Kantorovitz:1983:STB

[Kan83o] Shmuel Kantorovitz. Spectral Theory of Banach Space Operators: Ck-classification, abstract Volterra operators, similarity, spectrality, local spectral analysis, volume 1012 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 179 pp. LCCN QA3 .L28 no. 1012.

URL http://link.springer.com/book/10.1007/BFb0064288;
 http://www.springerlink.com/content/978-3-540-38666-
7.

Kantorovitz:1983:VE

[Kan83p] Shmuel Kantorovitz. Volterra elements. Lecture Notes in Mathematics, 1012:25–37, 1983. CODEN LNMAA2. ISBN 3-540-12673-2 (print), 3-540-38666-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064294/.

Kaniuth:1984:RLC

[Kan84a] Eberhard Kaniuth. ∗-regularity of locally compact groups. Lecture Notes in Mathematics, 1064:235–240, 1984. CODEN LN-MAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073645/.

Kannan:1984:RMN

[Kan84b] R. Kannan. Relaxation methods in nonlinear problems. Lecture Notes in Mathematics, 1107:135–153, 1984. CODEN LN-MAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101498/.

Kano:1984:GFG

[Kan84c] Mikio Kano. Graph factors with given properties. Lecture Notes in Mathematics, 1073:161–168, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073114/.

Kaplansky:1982:SSL

[Kap82] Irving Kaplansky. Some simple Lie algebras of characteristic 2. Lecture Notes in Mathematics, 933:127–129, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093357/.

Karcher:1980:RMG

[Kar80a] Hermann Karcher. Report on M. Gromov’s almost flat manifolds. Lecture Notes in Mathematics, 770:21–35, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096233/.

Karzel:1980:RS

[Kar80b] Helmut Karzel. Rectangular spaces. Lecture Notes in Mathematics, 792:79–91, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088666/.

Karim:1981:FPS

[Kar81] Rahmi Ibrahim Ibrahim Abdel Karim. Fundamental properties of the solutions of quasilinear differential equations of the fifth order. Lecture Notes in Mathematics, 846:193–202, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089837/.

Karanikas:1983:PPP

[Kar83] C. Karanikas. Pathological properties and S.I.P. measures on metrizable groups. Lecture Notes in Mathematics, 992:323– 332, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069166/.

Kassel:1981:TRI

[Kas81] Christian Kassel. K-théorie relative d’un idéal bilatère de carré nul: étude homologique en basse dimension. (French) []. Lecture Notes in Mathematics, 854:249–261, 1981. CODEN LN-MAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089524/.

Kaspar:1982:OMC

[Kas82a] Bernhard Kaspar. Overrelaxation in monotonically convergent iteration methods. Lecture Notes in Mathematics, 953:80–87, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069375/.

Kassel:1982:GTP

[Kas82b] Christian Kassel. Le groupe K_{3}(Z[∈]) n’a pas de p-torsion pour p ≠ 2 et 3. (French) []. Lecture Notes in Mathematics, 966:115– 121, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062171/.

Kashiwara:1983:VCS

[Kas83] M. Kashiwara. Vanishing cycle sheaves and holonomic systems of differential equations. Lecture Notes in Mathematics, 1016: 134–142, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099962/.

Kasue:1984:ESP

[Kas84] Atsushi Kasue. Estimates for solutions of Poisson equations and their applications to submanifolds. Lecture Notes in Mathematics, 1090:1–14, 1984. CODEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101562/.

Kato:1980:SFD

[Kat80a] Junji Kato. Stability in functional differential equations. Lecture Notes in Mathematics, 799:252–262, 1980. CODEN LN-MAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089317/.

Katok:1980:EAT

[Kat80b] Svetlana R. Katok. The estimation from above for the topological entropy of a diffeomorphism. Lecture Notes in Mathematics, 819:258–264, 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086992/.

Katz:1981:STL

[Kat81] N. Katz. Serre-tate local moduli. Lecture Notes in Mathematics, 868:138–202, 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090648/.

Kato:1982:GCC

[Kat82] Kazuya Kato. Galois cohomology of complete discrete valuation fields. Lecture Notes in Mathematics, 967:215–238, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061904/.

Kato:1983:CFT

[Kat83] Kazuya Kato. Class field theory and algebraic K-theory. Lecture Notes in Mathematics, 1016:109–126, 1983. CODEN LN-MAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099960/.

Kauffman:1980:WPP

[Kau80] Louis H. Kauffman. Weaving patterns and polynomials. Lecture Notes in Mathematics, 788:88–97, 1980. CODEN LN-MAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099240/.

Kaufmann:1981:EEE

[Kau81a] Matt Kaufmann. On existence of Σ_{n} end extensions. Lecture Notes in Mathematics, 859:92–103, 1981. CODEN LN-MAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090942/.

Kaup:1981:TCC

[Kau81b] Ludger Kaup. On the topology of compact complex surfaces. Lecture Notes in Mathematics, 843:319–340, 1981. CODEN LN-MAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089280/.

Kaup:1981:JAH

[Kau81c] Wilhelm Kaup. Jordan algebras and holomorphy. Lecture Notes in Mathematics, 843:341–365, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089281/.

Kauffman:1983:FFE

[Kau83] Robert M. Kauffman. Factorization and the Friedrichs extension for ordinary differential operators. Lecture Notes in Mathematics, 1032:257–266, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076801/.

Kawai:1981:ASN

[Kaw81] Toru Kawai. Axiom systems of nonstandard set theory. Lecture Notes in Mathematics, 891:57–65, 1981. CODEN LN-MAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090979/.

Kawohl:1982:SPA

[Kaw82] B. Kawohl. A singular perturbation approach to nonlinear elliptic boundary value problems. Lecture Notes in Mathematics, 942: 65–71, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094740/.

Kazhdan:1983:L

[Kaz83] David Kazhdan. On lifting. Lecture Notes in Mathematics, 1041: 209–249, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073149/.

Krause:1980:FDS

[KB80] Egon Krause and Fritz Bartels. Finite-difference solutions of the Navier–Stokes equations for axially symmetric flows in spherical gaps. Lecture Notes in Mathematics, 771:313–322, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086915/.

Krein:1984:MP

[KDB^{+}84] M. G. Krein, P. L. Duren, A. Baernstein, Josef Král, A. B. Gulisashvili, E. a. Gorin, E. A. Gorin, V. Ya. Lin, V. Ya. Lin, and M. G. Zaidenberg. Miscellaneous problems. Lecture Notes in Mathematics, 1043:631–663, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072196/.

Kearfott:1981:DFA

[Kea81] R. B. Kearfott. A derivative-free arc continuation method and a bifurcation technique. Lecture Notes in Mathematics, 878:182– 198, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090682/.

Keating:1982:WGD

[Kea82] M. E. Keating. Whitehead groups of dihedral 2-groups. Lecture Notes in Mathematics, 966:122–127, 1982. CODEN LN-MAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062172/.

Kearfott:1983:CMP

[Kea83] Ralph Baker Kearfott. Continuation methods and parametrized nonlinear least squares: Techniques and experiments. Lecture Notes in Mathematics, 1005:142–150, 1983. CODEN LN-MAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112531/.

Kechris:1981:HTP

[Kec81a] Alexander S. Kechris. Homogeneous trees and projective scales. Lecture Notes in Mathematics, 839:33–73, 1981. CODEN LN-MAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090235/.

Kechris:1981:NWD

[Kec81b] Alexander S. Kechris. A note on Wadge degrees. Lecture Notes in Mathematics, 839:165–168, 1981. CODEN LNMAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090240/.

Kechris:1981:SCS

[Kec81c] Alexander S. Kechris. Souslin cardinals, κ-Souslin sets and the scale property in the hyperprojective hierarchy. Lecture Notes in Mathematics, 839:127–146, 1981. CODEN LNMAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090238/.

Keedwell:1983:SGG

[Kee83] A. D. Keedwell. Sequenceable groups, generalized complete mappings, neofields and block designs. Lecture Notes in Mathematics, 1036:49–71, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071508/.

Kellerer:1980:BSP

[Kel80] Hans G. Kellerer. Baire sets in product spaces. Lecture Notes in Mathematics, 794:38–44, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088210/.

Kempf:1983:ETC

[Kem83] George R. Kempf. The elementary theory of correspondences. Lecture Notes in Mathematics, 971:68–78, 1983. CODEN LN-MAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067074/.

Kendall:1981:BMN

[Ken81] W. S. Kendall. Brownian motion, negative curvature, and harmonic maps. Lecture Notes in Mathematics, 851:479–491, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088739/.

Kendall:1984:BMS

[Ken84a] Wilfrid S. Kendall. Brownian motion on a surface of negative curvature. Lecture Notes in Mathematics, 1059:70–76, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100032/.

Kenmotsu:1984:DGS

[Ken84b] K. Kenmotsu, editor. Differential Geometry of Submanifolds: Proceedings of the Conference held at Kyoto, January 23–25, 1984, volume 1090 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1090. URL http:/

/link.springer.com/book/10.1007/BFb0101561;   http://
www.springerlink.com/content/978-3-540-39065-7.

Kettle:1982:CNB

[Ket82a] St. John G. Kettle. A class of natural bijections between Catalan families. Lecture Notes in Mathematics, 952:327–348, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061987/.

Kettler:1982:ACR

[Ket82b] Rob Kettler. Analysis and comparison of relaxation schemes in robust multigrid and preconditioned conjugate gradient methods. Lecture Notes in Mathematics, 960:502–534, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069941/.

Kettle:1983:CES

[Ket83a] S. G. Kettle. Classifying and enumerating some freely generated families of objects. Lecture Notes in Mathematics, 1036:

275–304, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071524/.

Kettle:1983:FES

[Ket83b] S. G. Kettle. Families enumerated by the Schröder–Etherington sequence and a renewal array it generates. Lecture Notes in Mathematics, 1036:244–274, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071523/.

Khaleelulla:1982:CTV

[Kha82a] S. M. Khaleelulla. Counterexamples in Topological Vector Spaces, volume 936 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxi + 179 pp. LCCN QA3 .L28 no. 936. URL http:/

/link.springer.com/book/10.1007/BFb0097678;   http://
www.springerlink.com/content/978-3-540-39268-2.

Khaleelulla:1982:HP

[Kha82b] S. M. Khaleelulla. Hereditary properties. Lecture Notes in Mathematics, 936:103–109, 1982. CODEN LNMAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097685/.

Khaleelulla:1982:LCS

[Kha82c] S. M. Khaleelulla. Locally convex spaces. Lecture Notes in Mathematics, 936:18–26, 1982. CODEN LNMAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097681/.

Khaleelulla:1982:OTV

[Kha82d] S. M. Khaleelulla. Ordered topological vector spaces. Lecture Notes in Mathematics, 936:77–102, 1982. CODEN LN-MAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097684/.

Khaleelulla:1982:PVS

[Kha82e] S. M. Khaleelulla. Prerequisites vector spaces and ordered vector spaces. Lecture Notes in Mathematics, 936:1–8, 1982. CODEN LNMAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097679/.

Khaleelulla:1982:SCL

[Kha82f] S. M. Khaleelulla. Special classes of locally convex spaces. Lecture Notes in Mathematics, 936:27–64, 1982. CODEN LN-MAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097682/.

Khaleelulla:1982:SCT

[Kha82g] S. M. Khaleelulla. Special classes of topological vector spaces. Lecture Notes in Mathematics, 936:65–76, 1982. CODEN LN-MAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097683/.

Khaleelulla:1982:TA

[Kha82h] S. M. Khaleelulla. Topological algebras. Lecture Notes in Mathematics, 936:137–157, 1982. CODEN LNMAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097687/.

Khaleelulla:1982:TB

[Kha82i] S. M. Khaleelulla. Topological bases. Lecture Notes in Mathematics, 936:110–136, 1982. CODEN LNMAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097686/.

Khaleelulla:1982:TVS

[Kha82j] S. M. Khaleelulla. Topological vector spaces. Lecture Notes in Mathematics, 936:9–17, 1982. CODEN LNMAA2. ISBN 3-540-11565-X (print), 3-540-39268-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097680/.

Komkov:1984:UGM

[KI84] Vadim Komkov and Carlton Irwin. Uniqueness for gradient methods in engineering optimization. Lecture Notes in Mathematics, 1086:93–118, 1984. CODEN LNMAA2. ISBN 3-540-13871-4 (print), 3-540-39061-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073071/.

Kifer:1980:PVB

[Kif80] Yuri Kifer. A probabilistic version of Bowen–Ruelle’s volume lemma. Lecture Notes in Mathematics, 819:291–299, 1980. CO-DEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086994/.

Kijowski:1980:NSS

[Kij80] Jerzy Kijowski. A new symplectic structure of field theory. Lecture Notes in Mathematics, 836:186–201, 1980. CODEN LN-MAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089735/.

Kimura:1981:CSR

[Kim81] Tatsuo Kimura. On the construction of some relative invariants for CL(n) (n = 6, 7, 8) by the decomposition of the Young diagrams. Lecture Notes in Mathematics, 867:38–54, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090379/.

King:1981:CPD

[Kin81] Donald R. King. Character polynomials of discrete series representations. Lecture Notes in Mathematics, 880:267–302, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090413/.

Kinzl:1982:CPP

[Kin82] Franz Kinzl. Convolution powers of probability measures on locally compact semigroups. Lecture Notes in Mathematics, 928: 247–257, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093228/.

Kinderlehrer:1984:E

[Kin84a] David Kinderlehrer. Erratum. Lecture Notes in Mathematics, 1063:167, 1984. CODEN LNMAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072157.

Kinderlehrer:1984:MA

[Kin84b] David Kinderlehrer. A mathematical afterthought. Lecture Notes in Mathematics, 1063:162–166, 1984. CODEN LNMAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072156/.

Kinzl:1984:AEL

[Kin84c] Franz Kinzl. Asymptotic equidistribution on locally compact semigroups. Lecture Notes in Mathematics, 1064:241–259, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073646/.

Kirk:1981:LNM

[Kir81] W. A. Kirk. Locally nonexpansive mappings in Banach spaces. Lecture Notes in Mathematics, 886:178–198, 1981. CODEN LN-MAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092184/.

Kirillov:1982:IDL

[Kir82] A. A. Kirillov. Infinite dimensional Lie groups; their orbits, invariants and representations. The geometry of moments. Lecture Notes in Mathematics, 970:101–123, 1982. CODEN LN-MAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066026/.

Kirchgassner:1984:WWC

[Kir84] K. Kirchgässner. Waves in weakly-coupled parabolic systems. Lecture Notes in Mathematics, 1107:154–180, 1984. CODEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101499/.

Kiselman:1981:HRS

[Kis81a] Christer O. Kiselman. How to recognize supports from the growth of functional transforms in real and complex analysis. Lecture Notes in Mathematics, 843:366–372, 1981. CODEN LN-MAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089282/.

Kisliakov:1981:WNA

[Kis81b] S. V. Kisliakov. What is needed for a O-absolutely summing operator to be nuclear? Lecture Notes in Mathematics, 864: 336–364, 1981. CODEN LNMAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097001/.

Kiselman:1982:SDN

[Kis82] Christer O. Kiselman. Stabilité du nombre de Lelong par restriction à une sous-variété. (French) []. Lecture Notes in Mathematics, 919:324–336, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097055/.

Kiselman:1983:GCP

[Kis83] Christer O. Kiselman. The growth of compositions of a plurisubharmonic function with entire mappings. Lecture Notes in Mathematics, 1039:257–263, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073369/.

Kiselman:1984:DOM

[Kis84a] Christer O. Kiselman. Sur la définition de l’opérateur de Monge– Ampère complexe. (French) []. Lecture Notes in Mathematics, 1094:139–150, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099158/.

Kisynski:1984:FIW

[Kis84b] Jan Kisyński. On a formula of N. Ikeda and S. Watanabe concerning the Lévy kernel. Lecture Notes in Mathematics, 1064: 260–279, 1984. CODEN LNMAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073647/.

Kitagawa:1984:CHS

[Kit84] Yoshihisa Kitagawa. Compact homogeneous submanifolds with parallel mean curvature. Lecture Notes in Mathematics, 1090: 93–98, 1984. CODEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101569/.

Kernevez:1980:PFW

[KJTB80] J.-P. Kernevez, G. Joly, D. Thomas, and B. Bunow. Pattern formation and wave propagation in the s-a system. Lecture Notes in Mathematics, 782:201–221, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090434/.

Kuribayashi:1983:PCF

[KK83] Izumi Kuribayashi and Akikazu Kuribayashi. On parameters of a canonical form of a curve of genus three. Lecture Notes in Mathematics, 1013:132–159, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066525/.

Kakosyan:1984:CDM

[KKM84a] Ashot Vazrikievich Kakosyan, Leo Borisovich Klebanov, and Joseph Aleksandrovich Melamed. Characterization of Distributions by the Method of Intensively Monotone Operators, volume 1088 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13857-9 (print), 3-540-39050-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 175 pp. LCCN QA3 .L28 no. 1088. URL http:

//link.springer.com/book/10.1007/BFb0101656;   http:/
/www.springerlink.com/content/978-3-540-39050-3.

Kakosyan:1984:CPAb

[KKM84b] Ashot Vazrikievich Kakosyan, Leo Borisovich Klebanov, and Joseph Aleksandrovich Melamed. Characterization problems associated with non-linear statistics and problems of reconstruction of distributions. Lecture Notes in Mathematics, 1088:71– 100, 1984. CODEN LNMAA2. ISBN 3-540-13857-9 (print), 3-540-39050-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101659/.

Kakosyan:1984:CPAa

[KKM84c] Ashot Vazrikievich Kakosyan, Leo Borisovich Klebanov, and Joseph Aleksandrovich Melamed. Characterization problems associated with properties of linear statistics. Lecture Notes in Mathematics, 1088:23–70, 1984. CODEN LNMAA2. ISBN 3-540-13857-9 (print), 3-540-39050-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101658/.

Kakosyan:1984:CMD

[KKM84d] Ashot Vazrikievich Kakosyan, Leo Borisovich Klebanov, and Joseph Aleksandrovich Melamed. Characterizations of multivariate distributions. Lecture Notes in Mathematics, 1088:133– 160, 1984. CODEN LNMAA2. ISBN 3-540-13857-9 (print), 3-540-39050-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101661/.

Kakosyan:1984:IMO

[KKM84e] Ashot Vazrikievich Kakosyan, Leo Borisovich Klebanov, and Joseph Aleksandrovich Melamed. Intensively monotone operators and their properties. Lecture Notes in Mathematics, 1088: 1–22, 1984. CODEN LNMAA2. ISBN 3-540-13857-9 (print), 3-540-39050-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101657/.

Kakosyan:1984:SPC

[KKM84f] Ashot Vazrikievich Kakosyan, Leo Borisovich Klebanov, and Joseph Aleksandrovich Melamed. On some problems of characterization of distributions associated with mathematical theory of reliability. Lecture Notes in Mathematics, 1088:101–132, 1984. CODEN LNMAA2. ISBN 3-540-13857-9 (print), 3-540-39050-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101660/.

Kechris:1981:ADS

[KKMW81] Alexander S. Kechris, Eugene M. Kleinberg, Yiannis N. Moschovakis, and W. Hugh Woodin. The axiom of determinacy, strong partition properties and nonsingular measures. Lecture Notes in Mathematics, 839:75–99, 1981. CODEN LN-MAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090236/.

Knill:1982:HMP

[KKS82] Ronald J. Knill, Morris Kalka, and Howard C. J. Sealey, editors. Harmonic Maps: Proceedings of the N.S.F.-C.B.M.S. Regional Conference, Held at Tulane University, New Orleans December 15–19, 1980, volume 949 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 949. URL http:

//link.springer.com/book/10.1007/BFb0069752;   http:/
/www.springerlink.com/content/978-3-540-39360-3.

Klarsfeld:1981:PAR

[Kla81] S. Klarsfeld. Padé approximants and related methods for computing boundary values on cuts. Lecture Notes in Mathematics, 888:255–262, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095591/.

Klares:1983:CTU

[Kla83a] Bernard Klarès. Classification topologique des n-uples de champs de vecteurs holomorphes commutatifs sur P_{n+1}(C). (French) []. Lecture Notes in Mathematics, 1015:307–321, 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071355/.

Klares:1983:DLI

[Kla83b] Bernard Klarès. Déformations localement iso-irrégulières de connexions linéaires complètement intégrables sur P_{m}(C). (French)

[]. Lecture Notes in Mathematics, 1015:322–411, 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071356/.

Kleiner:1981:IMC

[Kle81] Mark Kleiner. Induced modules and comodules and representations of BOCS’s and DGC’s. Lecture Notes in Mathematics, 903: 168–185, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092991/.

Kleiman:1982:BM

[Kle82a] Steven L. Kleiman. Back matter. Lecture Notes in Mathematics, 947:??, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-39342-9/1.pdf.

Kleiman:1982:PFM

[Kle82b] Steven L. Kleiman. Plane forms and multiple-point formulas. Lecture Notes in Mathematics, 947:287–310, 1982. CODEN LN-MAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093596/.

Kleiman:1984:ACS

[Kle84a] Steven L. Kleiman. About the conormal scheme. Lecture Notes in Mathematics, 1092:161–197, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099362/.

Klein:1984:ACO

[Kle84b] Abel Klein. The analytic continuation of a Osterwalder–Schrader positive representation of the Euclidean group to a representation of the Poincaré group. Lecture Notes in Mathematics, 1055: 199–205, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071722/.

Klein:1984:ACB

[Kle84c] Abel Klein. Appendix: a connection between quantum systems and stochastic processes. Lecture Notes in Mathematics, 1055: 206–209, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071723/.

Kletzing:1984:AWG

[Kle84d] Dennis Kletzing. Application to the Weyl groups of exceptional type. Lecture Notes in Mathematics, 1084:220–268, 1984. CODEN LNMAA2. ISBN 3-540-13865-X (print), 3-540-39056-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103433/.

Kletzing:1984:CG

[Kle84e] Dennis Kletzing. Constructions of Q-groups. Lecture Notes in Mathematics, 1084:58–104, 1984. CODEN LNMAA2. ISBN 3-540-13865-X (print), 3-540-39056-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103430/.

Kletzing:1984:GSR

[Kle84f] Dennis Kletzing. General structural results. Lecture Notes in Mathematics, 1084:7–57, 1984. CODEN LNMAA2. ISBN 3-540-13865-X (print), 3-540-39056-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103429/.

Kletzing:1984:I

[Kle84g] Dennis Kletzing. Introduction. Lecture Notes in Mathematics, 1084:1–4, 1984. CODEN LNMAA2. ISBN 3-540-13865-X (print), 3-540-39056-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103427/.

Kletzing:1984:LC

[Kle84h] Dennis Kletzing. Local characters. Lecture Notes in Mathematics, 1084:105–173, 1984. CODEN LNMAA2. ISBN 3-540-13865-X (print), 3-540-39056-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103431/.

Kletzing:1984:N

[Kle84i] Dennis Kletzing. Notation. Lecture Notes in Mathematics, 1084: 5–6, 1984. CODEN LNMAA2. ISBN 3-540-13865-X (print), 3-540-39056-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103428.

Kletzing:1984:RRG

[Kle84j] Dennis Kletzing. Rational representations of Q-groups. Lecture Notes in Mathematics, 1084:174–219, 1984. CODEN LN-MAA2. ISBN 3-540-13865-X (print), 3-540-39056-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103432/.

Kletzing:1984:SRG

[Kle84k] Dennis Kletzing. Structure and Representations of Q-groups, volume 1084 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13865-X (print), 3-540-39056-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 290 pp. LCCN QA3 .L28 no. 1084. URL http:/

/link.springer.com/book/10.1007/BFb0103426;   http://
www.springerlink.com/content/978-3-540-39056-5.

Kalina:1980:SBI

[KŁLS80] Jerzy Kalina, Julian Ławrynowicz, Ewa Ligocka, and Maciej Skwarczyński. On some biholomorphic invariants in the analysis on manifolds. Lecture Notes in Mathematics, 798:224–249, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097268/.

Korezlioglu:1981:PMD

[KLM81] H. Korezlioglu, P. Lefort, and G. Mazziotto. Une propriété markovienne et diffusions associées. (French) []. Lecture Notes in Mathematics, 863:245–274, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091104/.

Kadec:1984:EMS

[KLR^{+}84] M. I. Kadec, James Langley, Lee A. Rubel, A. A. Gol’Dberg, A.È. Erëmenko, I. V. Ostrovskii, A. A. Gol’Dberg, A.È. Erëmenko, I. V. Ostrovskii, B. Ya. Levin, I. V. Ostrovskii, B. Ya. Levin, S. Ya. Havinson, B. Ya. Levin, V. S. Azarin, et al. Entire, meromorphic and subharmonic functions. Lecture Notes in Mathematics, 1043:589–618, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072194/.

Kaiser:1980:GMM

[KM80] Gerald Kaiser and Jerrold E. Marsden, editors. Geometric Methods in Mathematical Physics: Proceedings of an NSF–CBMS Conference Held at the University of Lowell, Massachusetts, March 19–23, 1979, volume 775 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 775. URL http:

//link.springer.com/book/10.1007/BFb0092017;   http:/
/www.springerlink.com/content/978-3-540-38571-4.

Kaufmann:1982:SWD

[KM82] Susana Kaufmann and Antonio Montalvo. Standing waves in diffusive reacting systems. Lecture Notes in Mathematics, 909: 116–124, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092965/.

Klebanov:1983:DPS

[KM83] L. B. Klebanov and J. A. Melamed. Discretization in the problems of stability of characterization of the exponential distribution. Lecture Notes in Mathematics, 982:110–122, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082063/.

Kechris:1981:CSP

[KMM81] Alexander S. Kechris, Donald A. Martin, and Yiannis N. Moschovakis, editors. Cabal Seminar 77–79: Proceedings, Caltech–UCLA Logic Seminar 1977–79, volume 839 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 839.

URL http://link.springer.com/book/10.1007/BFb0090233;
 http://www.springerlink.com/content/978-3-540-38422-
9.

Kechris:1983:CSP

[KMM83a] Alexander S. Kechris, Donald A. Martin, and Yiannis N. Moschovakis, editors. Cabal Seminar 79–81: Proceedings, Caltech–UCLA Logic Seminar 1979–81, volume 1019 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1019.

URL http://link.springer.com/book/10.1007/BFb0071690;
 http://www.springerlink.com/content/978-3-540-38680-
3.

Klebanov:1983:ELS

[KMM83b] L. B. Klebanov, G. M. Manija, and J. A. Melamed. On the estimation of location and scale parameters of stable laws. Lecture Notes in Mathematics, 982:77–109, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082062/.

Korezlioglu:1981:PAD

[KMS81] Hayri Korezlioglu, Gerald Mazziotto, and Jacques Szpirglas, editors. Processus Aléatoires à Deux Indices: Colloque E.N.S.T.– C.N.E.T., Paris 1980, volume 863 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 863. URL http:

//link.springer.com/book/10.1007/BFb0091089;   http:/
/www.springerlink.com/content/978-3-540-38718-3.

Kolzow:1982:MTO

[KMS82] D. Kölzow and D. Maharam-Stone, editors. Measure Theory Oberwolfach 1981: Proceedings of the Conference Held at Oberwolfach, Germany, June 21–27, 1981, volume 945 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 945.

URL http://link.springer.com/book/10.1007/BFb0096658;
 http://www.springerlink.com/content/978-3-540-39324-
5.

Kechris:1983:IT

[KMS83] Alexander S. Kechris, Donald A. Martin, and Robert M. Solovay. Introduction to Q-theory. Lecture Notes in Mathematics, 1019:199–281, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071702/.

Kolzow:1984:MTO

[KMS84] D. Kölzow and D. Maharam-Stone, editors. Measure Theory Oberwolfach 1983: Proceedings of the Conference held at Oberwolfach, June 26–July 2, 1983, volume 1089 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1089. URL http:/

/link.springer.com/book/10.1007/BFb0072595;   http://
www.springerlink.com/content/978-3-540-39069-5.

Keynes:1980:EE

[KN80a] H. B. Keynes and D. Newton. Ergodicity in (G, σ)-extensions. Lecture Notes in Mathematics, 819:265–290, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086993/.

Koschorke:1980:TSS

[KN80b] Ulrich Koschorke and Walter D. Neumann, editors. Topology Symposium Siegen 1979: Proceedings of a Symposium Held at the University of Siegen, June 14–19, 1979, volume 788 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 788.

URL http://link.springer.com/book/10.1007/BFb0099234;
 http://www.springerlink.com/content/978-3-540-39187-

Kolb:1982:RMC

[KN82] Gerhard Kolb and Wilhelm Niethammer. Relaxation methods for the computation of the spectral norm. Lecture Notes in Mathematics, 953:149–158, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069380/.

Kim:1984:GKP

[KN84] Ann Chi Kim and Bernhard H. Neumann, editors. Groups — Korea 1983: Proceedings of a Conference on Combinatorial Group Theory, held at Kyoungju, Korea, August 26–31, 1983, volume 1098 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1098. URL http:/

/link.springer.com/book/10.1007/BFb0099655;   http://
www.springerlink.com/content/978-3-540-39102-9.

Knapp:1983:MTF

[Kna83a] A. W. Knapp. Minimal K-type formula. Lecture Notes in Mathematics, 1020:107–118, 1983. CODEN LNMAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071499/.

Knapp:1983:URB

[Kna83b] A. W. Knapp. Unitary representations and basic cases. Lecture Notes in Mathematics, 1024:77–98, 1983. CODEN LN-MAA2. ISBN 3-540-12725-9 (print), 3-540-38709-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071432/.

Knauer:1983:CMP

[Kna83c] Ulrich Knauer. Characterization of monoids by properties of finitely generated right acts and their right ideals. Lecture Notes in Mathematics, 998:310–332, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062038/.

Kneser:1980:SVD

[Kne80] M. Kneser. Schiefkörper Über Vollständig Diskret Bewerteten Körpern. (German) []. Lecture Notes in Mathematics, 778:43– 46, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095921/.

Knightly:1980:SDP

[Kni80] George H. Knightly. Some decay properties of solutions of the Navier–Stokes equations. Lecture Notes in Mathematics, 771: 287–298, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086913/.

Knight:1983:TPP

[Kni83] Frank B. Knight. A transformation from prediction to past of an L^{2}-stochastic process. Lecture Notes in Mathematics, 986:1– 14, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068295/.

Knight:1984:RT

[Kni84] F. B. Knight. On the ray topology. Lecture Notes in Mathematics, 1059:56–69, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100031/.

Knobloch:1980:HON

[Kno80] H.-W. Knobloch. Higher order necessary conditions in optimal control theory. Lecture Notes in Mathematics, 827:151–164, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091379/.

Knopp:1981:ANT

[Kno81] Marvin I. Knopp, editor. Analytic Number Theory: Proceedings of a Conference Held at Temple University, Philadelphia, May 12–15, 1980, volume 899 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 899. URL http:

//link.springer.com/book/10.1007/BFb0096450;   http:/
/www.springerlink.com/content/978-3-540-38953-8.

Knowles:1983:EPR

[Kno83] Ian Knowles. Eigenvalue problems and the Riemann zeta function II. Lecture Notes in Mathematics, 1032:267–297, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076802/.

Knus:1982:QM

[Knu82] M.-A. Knus. Quaternionic modules over P^{2}(R). Lecture Notes in Mathematics, 917:245–259, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092239/.

Kervaire:1981:GBS

[KO81a] Michel Kervaire and Manuel Ojanguren, editors. Groupe de Brauer: Séminaire, Les Plans-sur-Bex, Suisse 1980, volume 844 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 844. URL http:/

/link.springer.com/book/10.1007/BFb0090474;   http://
www.springerlink.com/content/978-3-540-38531-8.

Knus:1981:CEG

[KO81b] M.-A. Knus and M. Ojanguren. Cohomologie Étale et groupe de Brauer. (French) []. Lecture Notes in Mathematics, 844:210– 228, 1981. CODEN LNMAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090481/.

Kincaid:1983:IS

[KO83] David R. Kincaid and Thomas C. Oppe. Itpack on supercomputers. Lecture Notes in Mathematics, 1005:151–161, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112532/.

Kobayashi:1981:HC

[Kob81] Shoshichi Kobayashi. On holomorphic connections. Lecture Notes in Mathematics, 838:154–164, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088853/.

Kocay:1982:SNM

[Koc82] W. L. Kocay. Some new methods in reconstruction theory. Lecture Notes in Mathematics, 952:89–114, 1982. CODEN LN-MAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061974/.

Kohler:1981:DCC

[Köh81] Egmont Köhler. k-Difference-cycles and the construction of cyclic t-designs. Lecture Notes in Mathematics, 893:195–203, 1981. CO-DEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091020/.

Kohler:1982:QSA

[Köh82] Egmont Köhler. Quadruple systems over Z_{p} admitting the affine group. Lecture Notes in Mathematics, 969:212–228, 1982. CO-DEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062996/.

Kohno:1983:MPC

[Koh83] Mitsuhiko Kohno. A multi-point connection problem. Lecture Notes in Mathematics, 1015:136–171, 1983. CODEN LN-MAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071352/.

Koiso:1984:SGM

[Koi84] Miyuki Koiso. The stability and the Gauss map of minimal surfaces in R^{3}. Lecture Notes in Mathematics, 1090:77–92, 1984. CODEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101568/.

Kojima:1981:IVD

[Koj81] M. Kojima. An introduction to variable dimension algorithms for solving systems of equations. Lecture Notes in Mathematics, 878:199–237, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090683/.

Kolzow:1980:MTO

[Köl80] Dietrich Kölzow, editor. Measure Theory Oberwolfach 1979: Proceedings of the Conference Held at Oberwolfach, Germany, July 1–7, 1979, volume 794 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 794. URL http:/

/link.springer.com/book/10.1007/BFb0088206;   http://
www.springerlink.com/content/978-3-540-39221-7.

Kolaitis:1981:MTC

[Kol81a] Phokion G. Kolaitis. Model theoretic characterizations in generalized recursion theory. Lecture Notes in Mathematics, 859: 104–119, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090943/.

Kolar:1981:FPC

[Kol81b] Ivan Kolář. Fiber parallelism and connections. Lecture Notes in Mathematics, 838:165–173, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088854/.

Kolster:1982:EDP

[Kol82a] Manfred Kolster. Even dimensional projective surgery groups of finite groups. Lecture Notes in Mathematics, 967:239–279, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061905/.

Kolster:1982:IS

[Kol82b] Manfred Kolster. On injective stability for K_{2}. Lecture Notes in Mathematics, 966:128–168, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062173/.

Kolmogorov:1983:LFP

[Kol83] A. N. Kolmogorov. On logical foundations of probability theory. Lecture Notes in Mathematics, 1021:1–5, 1983. CODEN LNMAA2. ISBN 3-540-12718-6 (print), 3-540-38701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072897/.

Kolsrud:1984:LSF

[Kol84] Torbjörn Kolsrud. Localisation of superharmonic functions in not necessarily local Dirichlet spaces. Lecture Notes in Mathematics, 1096:401–411, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100121/.

Komatu:1980:OPF

[Kom80] Yûsaku Komatu. A one-parameter family of operators defined on analytic functions in a circle. Lecture Notes in Mathematics, 798:292–300, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097272/.

Komkov:1984:SAS

[Kom84a] Vadim Komkov. Sensitivity analysis in some engineering applications. Lecture Notes in Mathematics, 1086:1–30, 1984. CODEN LNMAA2. ISBN 3-540-13871-4 (print), 3-540-39061-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073067/.

Komkov:1984:SFA

[Kom84b] Vadim Komkov, editor. Sensitivity of Functionals with Applications to Engineering Sciences: Proceedings of a Special Session of the American Mathematical Society Spring Meeting held in New York City May 1983, volume 1086 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13871-4 (print), 3-540-39061-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1086. URL http:/

/link.springer.com/book/10.1007/BFb0073066;   http://
www.springerlink.com/content/978-3-540-39061-9.

Konarski:1982:PEA

[Kon82a] Jerzy Konarski. A pathological example of an action of k∗. Lecture Notes in Mathematics, 956:72–78, 1982. CODEN LN-MAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101509/.

Konarski:1982:PPO

[Kon82b] Jerzy Konarski. Properties of projective orbits of actions of affine algebraic groups. Lecture Notes in Mathematics, 956:79–91, 1982. CODEN LNMAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101510/.

Koornwinder:1980:PDS

[Koo80] Tom H. Koornwinder. A precise definition of separation of variables. Lecture Notes in Mathematics, 810:240–263, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089982/.

Koppelberg:1981:LSI

[Kop81] Sabine Koppelberg. A lattice structure on the isomorphism types of complete Boolean algebras. Lecture Notes in Mathematics, 872: 98–126, 1981. CODEN LNMAA2. ISBN 3-540-10849-1 (print), 3-540-38757-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098620/.

Kori:1980:PAB

[Kor80a] Tosiaki Kori. Problèmes au bord sur un espace harmonique. (French) []. Lecture Notes in Mathematics, 787:185–190, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086333/.

Korner:1980:IMM

[Kör80b] T. W. Körner. Ivašev Mustov in many dimensions. Lecture Notes in Mathematics, 781:102–120, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097650/.

Koras:1982:LRG

[Kor82] Mariusz Koras. Linearization of reductive group actions. Lecture Notes in Mathematics, 956:92–98, 1982. CODEN LN-MAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101511/.

Korner:1983:BPS

[Kör83] T. W. Körner. The behavior of power series on their circle of convergence. Lecture Notes in Mathematics, 995:56–94, 1983. CODEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061888/.

Koschorke:1980:VFM

[Kos80a] Ulrich Koschorke. Vector fields on (4q + 2)-manifolds. Lecture Notes in Mathematics, 788:98–108, 1980. CODEN LN-MAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099241/.

Kosniowski:1980:SFC

[Kos80b] Czes Kosniowski. Some formulae and conjectures associated with circle actions. Lecture Notes in Mathematics, 788:331–339, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099252/.

Koschorke:1981:BC

[Kos81a] Ulrich Koschorke. Bordism classification. Lecture Notes in Mathematics, 847:55–137, 1981. CODEN LNMAA2. ISBN 3-540-10572-7 (print), 3-540-38546-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090443/.

Koschorke:1981:EHC

[Kos81b] Ulrich Koschorke. Existence and homotopy classification. Lecture Notes in Mathematics, 847:17–54, 1981. CODEN LN-MAA2. ISBN 3-540-10572-7 (print), 3-540-38546-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090442/.

Koschorke:1981:F

[Kos81c] Ulrich Koschorke. Framefields. Lecture Notes in Mathematics, 847:138–298, 1981. CODEN LNMAA2. ISBN 3-540-10572-7 (print), 3-540-38546-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090444/.

Koschorke:1981:I

[Kos81d] Ulrich Koschorke. Introduction. Lecture Notes in Mathematics, 847:1–16, 1981. CODEN LNMAA2. ISBN 3-540-10572-7 (print), 3-540-38546-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090441/.

Koschorke:1981:VFO

[Kos81e] Ulrich Koschorke. Vector Fields and Other Vector Bundle Morphisms — a Singularity Approach, volume 847 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10572-7 (print), 3-540-38546-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 304 pp. LCCN QA3 .L28 no. 847. URL http:/

/link.springer.com/book/10.1007/BFb0090440;   http://
www.springerlink.com/content/978-3-540-38546-2.

Koshitani:1981:SRM

[Kos81f] Shigeo Koshitani. Some remarks on the modular group algebras. Lecture Notes in Mathematics, 903:186–195, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092992/.

Kostant:1982:PCG

[Kos82] Bertram Kostant. Poisson commutativity and the generalized periodic Toda lattice. Lecture Notes in Mathematics, 905:12– 28, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092424/.

Kosciuk:1983:ESM

[Kos83a] S. A. Kosciuk. Existence of solutions to the martingale problem in R^{d} for degenerate discontinuous coefficients. Lecture Notes in Mathematics, 983:116–119, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065336/.

Kosciuk:1983:SSP

[Kos83b] S. A. Kosciuk. Stochastic solutions to partial differential equations. Lecture Notes in Mathematics, 983:113–115, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065335/.

Koslinski:1983:CTV

[Koś83c] Antoni Kośliński. A characterization of a tree by its vertex — disconnections. Lecture Notes in Mathematics, 1018:111–113, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071619/.

Kotelenez:1981:LBH

[Kot81] P. Kotelenez. Local behaviour of Hilbert space valued stochastic integrals and the continuity of mild solutions of stochastic evolution equations. Lecture Notes in Mathematics, 851:492–496, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088740/.

Kovaceva:1983:RUM

[Kov83] Ralitza Krumova Kovačeva. The roots of unity and the mmeromorphic extension of functions. Lecture Notes in Mathematics, 1039:264–275, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073370/.

Kowalski:1980:CGA

[Kow80a] Oldřich Kowalski. The classification of generalized affine symmetric spaces in low dimensions. Lecture Notes in Mathematics, 805:149–157, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103332/.

Kowalski:1980:CGS

[Kow80b] Oldřich Kowalski. The classification of generalized symmetric Riemannian spaces in low dimensions. Lecture Notes in Mathematics, 805:131–148, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103331/.

Kowalski:1980:DM

[Kow80c] Oldřich Kowalski. Differentiable s-manifolds. Lecture Notes in Mathematics, 805:45–67, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103327/.

Kowalski:1980:DSG

[Kow80d] Oldřich Kowalski. Distinguished s-structures on generalized symmetric spaces. Lecture Notes in Mathematics, 805:111–130, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103330/.

Kowalski:1980:EGS

[Kow80e] Oldřich Kowalski. Existence of generalized symmetric spaces of solvable type. Lecture Notes in Mathematics, 805:158–161, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103333/.

Kowalski:1980:GPS

[Kow80f] Oldřich Kowalski. Generalized pointwise symmetric spaces. Lecture Notes in Mathematics, 805:164–168, 1980. CODEN LN-MAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103335/.

Kowalski:1980:GSR

[Kow80g] Oldřich Kowalski. Generalized symmetric Riemannian spaces. Lecture Notes in Mathematics, 805:1–26, 1980. CODEN LN-MAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103325/.

Kowalski:1980:GSS

[Kow80h] Oldřich Kowalski. Generalized Symmetric Spaces, volume 805 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 187 pp. LCCN QA3 .L28 no. 805. URL http:/

/link.springer.com/book/10.1007/BFb0103324;   http://
www.springerlink.com/content/978-3-540-39329-0.

Kowalski:1980:IGA

[Kow80i] Oldřich Kowalski. Irreducible generalized affine symmetric spaces. Lecture Notes in Mathematics, 805:161–163, 1980. CODEN LN-MAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103334/.

Kowalski:1980:LRM

[Kow80j] Oldřich Kowalski. Locally regular s-manifolds. Lecture Notes in Mathematics, 805:68–88, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103328/.

Kowalski:1980:NPS

[Kow80k] Oldřich Kowalski. Non-parallel s-structures on symmetric spaces. Lecture Notes in Mathematics, 805:168–171, 1980. CODEN LN-MAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103336/.

Kowalski:1980:OM

[Kow80l] Oldřich Kowalski. Operations with s-manifolds. Lecture Notes in Mathematics, 805:89–110, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103329/.

Kowalski:1980:RS

[Kow80m] Oldřich Kowalski. Reductive spaces. Lecture Notes in Mathematics, 805:27–44, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103326/.

Kowalski:1980:SAR

[Kow80n] Oldřich Kowalski. Some advanced results. Lecture Notes in Mathematics, 805:171–172, 1980. CODEN LNMAA2. ISBN 3-540-10002-4 (print), 3-540-39329-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103337.

Kowalewski:1981:ACP

[Kow81a] Christine Kowalewski. Accélération de la convergence pour certaines suites à convergence logarithmique. (French) []. Lecture Notes in Mathematics, 888:263–272, 1981. CODEN LN-MAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095592/.

Kowalski:1981:RMG

[Kow81b] Oldřich Kowalski. Riemannian manifolds the geodesic balls of which are near to the Euclidean balls by volume. Lecture Notes in Mathematics, 838:174–180, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088855/.

Kalina:1980:VMS

[KP80] Jerzy Kalina and Antoni Pierzchalski. A variation of the modulus of submanifold families. Lecture Notes in Mathematics, 798: 250–257, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097269/.

Kaijser:1982:CFI

[KP82] Sten Kaijser and Joan Wick Pelletier. A categorical framework for interpolation theory. Lecture Notes in Mathematics, 962:145– 152, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066894/.

Kartsatos:1984:SAE

[KP84a] A. G. Kartsatos and M. E. Parrott. A simplified approach to the existence and stability problem of a functional evolution equation in a general Banach space. Lecture Notes in Mathematics, 1076:115–122, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072771/.

Kranakis:1984:PHS

[KP84b] Evangelos Kranakis and Iain Phillips. Partitions and homogeneous sets for admissible ordinals. Lecture Notes in Mathematics, 1103:235–260, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099389/.

Kamps:1982:CTA

[KPT82] Klaus Heiner Kamps, Dieter Pumplün, and Walter Tholen, editors. Category Theory: Applications to Algebra, Logic and Topology Proceedings of the International Conference Held at Gummersbach, July 6–10, 1981, volume 962 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 962. URL http:

//link.springer.com/book/10.1007/BFb0066878;   http:/
/www.springerlink.com/content/978-3-540-39550-8.

Kaagstrom:1983:MPP

[KR83a] Bo Kågström and Axel Ruhe, editors. Matrix Pencils: Proceedings of a Conference Held at Pite Havsbad, Sweden, March 22–24, 1982, volume 973 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 973. URL http:

//link.springer.com/book/10.1007/BFb0062089;   http:/
/www.springerlink.com/content/978-3-540-39447-1.

Karonski:1983:NSB

[KR83b] Michał Karoński and Andrzej Ruciński. On the number of strictly balanced subgraphs of a random graph. Lecture Notes in Mathematics, 1018:79–83, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071616/.

Kraskiewicz:1983:QMQ

[KR83c] J. Kraśkiewicz and R. Raczka. Quantization of models of quantum field theory with solitons. Lecture Notes in Mathematics, 1037:257–285, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073176/.

Kral:1980:BVP

[Krá80a] Josef Král. Boundary value problems. Lecture Notes in Mathematics, 823:125–149, 1980. CODEN LNMAA2. ISBN 3-540-10227-2 (print), 3-540-38288-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091041/.

Kral:1980:CR

[Krá80b] Josef Král. Comments and references. Lecture Notes in Mathematics, 823:150–159, 1980. CODEN LNMAA2. ISBN 3-540-10227-2 (print), 3-540-38288-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091042/.

Kral:1980:CNO

[Krá80c] Josef Král. Contractivity of Neumann’s operator. Lecture Notes in Mathematics, 823:73–101, 1980. CODEN LNMAA2. ISBN 3-540-10227-2 (print), 3-540-38288-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091039/.

Kral:1980:DLP

[Krá80d] Josef Král. Double layer potentials. Lecture Notes in Mathematics, 823:28–72, 1980. CODEN LNMAA2. ISBN 3-540-10227-2 (print), 3-540-38288-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091038/.

Kral:1980:FRN

[Krá80e] Josef Král. Fredholm radius of the Neumann operator. Lecture Notes in Mathematics, 823:102–124, 1980. CODEN LN-MAA2. ISBN 3-540-10227-2 (print), 3-540-38288-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091040/.

Kral:1980:IOP

[Krá80f] Josef Král. Integral Operators in Potential Theory, volume 823 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10227-2 (print), 3-540-38288-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 171 pp. LCCN QA3 .L28 no. 823. URL http:/

/link.springer.com/book/10.1007/BFb0091035;   http://
www.springerlink.com/content/978-3-540-38288-1.

Kral:1980:IR

[Krá80g] Josef Král. Introductory remark. Lecture Notes in Mathematics, 823:1–5, 1980. CODEN LNMAA2. ISBN 3-540-10227-2 (print), 3-540-38288-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091036/.

Kral:1980:WND

[Krá80h] Josef Král. Weak normal derivatives of potentials. Lecture Notes in Mathematics, 823:6–27, 1980. CODEN LNMAA2. ISBN 3-540-10227-2 (print), 3-540-38288-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091037/.

Kramm:1980:DAS

[Kra80i] Bruno Kramm. (DFN)-analytic spaces, Stein algebras and a ‘universal’ holomorphic functional calculus. Lecture Notes in Mathematics, 822:109–128, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097753/.

Kraft:1982:GMR

[Kra82a] Hanspeter Kraft. Geometric methods in representation theory. Lecture Notes in Mathematics, 944:180–258, 1982. CODEN LN-MAA2. ISBN 3-540-11577-3 (print), 3-540-39313-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094059/.

Kraljevic:1982:RIB

[Kra82b] Hrvoje Kraljević. Rank and index in Banach algebras. Lecture Notes in Mathematics, 948:98–117, 1982. CODEN LN-MAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069843/.

Krasinski:1983:BIR

[Kra83] Tadeusz Krasiński. On biholomorphic invariants related to homology groups. Lecture Notes in Mathematics, 1039:276–284, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073371/.

Kree:1981:LDO

[Kré81a] Paul Krée. Linear differential operators on vector spaces. Lecture Notes in Mathematics, 843:373–404, 1981. CODEN LN-MAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089283/.

Kreith:1981:OPW

[Kre81b] Kurt Kreith. Oscillation properties of weakly nonlinear differential equations. Lecture Notes in Mathematics, 846:203–209, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089838/.

Kress:1981:BIE

[Kre81c] Rainer Kress. On boundary integral equation methods in stationary electromagnetic reflection. Lecture Notes in Mathematics, 846:210–226, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089839/.

Kreck:1984:SCM

[Kre84a] M. Kreck. Some closed 4-manifolds with exotic differentiable structure. Lecture Notes in Mathematics, 1051:246–262, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0075570/.

Kreck:1984:ASH

[Kre84b] Matthias Kreck. Addition and subtraction of handles. Lecture Notes in Mathematics, 1069:41–50, 1984. CODEN LN-MAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101014/.

Kreck:1984:AG

[Kre84c] Matthias Kreck. Application to SK-groups. Lecture Notes in Mathematics, 1069:95–100, 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101018/.

Kreck:1984:BGO

[Kre84d] Matthias Kreck. Bordism groups of orientation preserving diffeomorphisms. Lecture Notes in Mathematics, 1069:12–16, 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101009/.

Kreck:1984:BDR

[Kre84e] Matthias Kreck. Bordism of Diffeomorphisms and Related Topics, volume 1069 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 144 pp. LCCN QA3 .L28 no. 1069. URL http:/

/link.springer.com/book/10.1007/BFb0101007;   http://
www.springerlink.com/content/978-3-540-38912-5.

Kreck:1984:BDM

[Kre84f] Matthias Kreck. Bordism of diffeomorphisms on manifolds with additional normal structures like spin-, unitary structures or framings; orientation reversing diffeomorphisms and the unoriented case. Lecture Notes in Mathematics, 1069:72–94, 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101017/.

Kreck:1984:FWT

[Kre84g] Matthias Kreck. Fibrations over S^{1} within their bordism class and the computation of Δ_{∗}. Lecture Notes in Mathematics, 1069: 33–40, 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101013/.

Kreck:1984:I

[Kre84h] Matthias Kreck. Introduction. Lecture Notes in Mathematics, 1069:1–11, 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101008/.

Kreck:1984:ISD

[Kre84i] Matthias Kreck. The isometric structure of a diffeomorphism. Lecture Notes in Mathematics, 1069:22–26, 1984. CODEN LN-MAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101011/.

Kreck:1984:MTD

[Kre84j] Matthias Kreck. The mapping torus of a diffeomorphism. Lecture Notes in Mathematics, 1069:27–32, 1984. CODEN LN-MAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101012/.

Kreck:1984:MRR

[Kre84k] Matthias Kreck. Miscellaneous results: Ring structure, generators, relation to the inertia group. Lecture Notes in Mathematics, 1069:101–109, 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101019/.

Kreck:1984:PTE

[Kre84l] Matthias Kreck. Proof of theorem 5.5 in the even-dimensional case. Lecture Notes in Mathematics, 1069:54–71, 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101016/.

Kreck:1984:PTO

[Kre84m] Matthias Kreck. Proof of theorem 5.5 in the odd-dimensional case. Lecture Notes in Mathematics, 1069:51–53, 1984. CODEN LNMAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101015/.

Kreck:1984:RAE

[Kre84n] Matthias Kreck. Report about equivariant Witt groups. Lecture Notes in Mathematics, 1069:17–21, 1984. CODEN LN-MAA2. ISBN 3-540-13362-3 (print), 3-540-38912-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101010/.

Krishnamurthy:1981:FIM

[Kri81] E. V. Krishnamurthy. A form invariant multivariable polynomial representation of graphs. Lecture Notes in Mathematics, 885: 18–32, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092252/.

Krickeberg:1982:PPS

[Kri82] K. Krickeberg. Processus ponctuels en statistique. (French) []. Lecture Notes in Mathematics, 929:205–313, 1982. CODEN LN-MAA2. ISBN 3-540-11547-1 (print), 3-540-39220-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095620/.

Krolikowski:1983:BIR

[Kró83] Wiesław Królikowski. Biholomorphic invariants on relative homology groups. Lecture Notes in Mathematics, 1039:285–319, 1983.

CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073372/.

Krupski:1984:OIS

[Kru84a] Paweł Krupski. Open images of solenoids. Lecture Notes in Mathematics, 1060:76–83, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099923/.

Kruszynski:1984:EGT

[Kru84b] Paweł Kruszyński. Extensions of Gleason theorem. Lecture Notes in Mathematics, 1055:210–227, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071724/.

Kunzle:1980:ECF

[KS80a] H. P. Künzle and J. R. Savage. Equilibrium configurations of fluids in general relativity. Lecture Notes in Mathematics, 836: 211–220, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089737/.

Kupisch:1980:SAF

[KS80b] Herbert Kupisch and Eberhard Scherzler. Symmetric algebras of finite representation type. Lecture Notes in Mathematics, 832: 328–368, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088473/.

Karlsson:1981:SFD

[KS81a] J. Karlsson and E. B. Saff. Singularities of functions determined by the poles of Padé approximants. Lecture Notes in Mathematics, 888:239–254, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095590/.

Knapp:1981:SNI

[KS81b] A. W. Knapp and E. M. Stein. Some new intertwining operators for semisimple groups. Lecture Notes in Mathematics, 880: 303–336, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090414/.

Knapp:1982:SCI

[KS82] A. W. Knapp and B. Speh. Status of classification of irreducible unitary representations. Lecture Notes in Mathematics, 908:1– 38, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093278/.

Kambayashi:1983:ECA

[KS83a] T. Kambayashi and V. Srinivas. On étale coverings of the affine space. Lecture Notes in Mathematics, 1008:75–82, 1983. CODEN LNMAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065701/.

Klares:1983:SPA

[KS83b] Bernard Klarès and Charles Sadler. Systèmes de Pfaff et algèbres de Lie libres, étude d’une singularité polaire normale. (French) []. Lecture Notes in Mathematics, 1028:163–218, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071682/.

Knapp:1983:RBC

[KS83c] A. W. Knapp and B. Speh. The role of basic cases in classification: Theorems about unitary representations applicable to SU(N, 2). Lecture Notes in Mathematics, 1020:119–160, 1983. CODEN LNMAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071500/.

Knobloch:1983:EPI

[KS83d] Hans Wilhelm Knobloch and Klaus Schmitt, editors. Equadiff 82: Proceedings of the international conference held in Würzburg, FRG, August 23–28, 1982, volume 1017 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12686-4 (print), 3-540-38678-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1017. URL http:/

/link.springer.com/book/10.1007/BFb0103230;   http://
www.springerlink.com/content/978-3-540-38678-0.

Kreith:1983:BCH

[KS83e] K. Kreith and C. A. Swanson. Boundedness criteria for hyperbolic characteristic initial value problems. Lecture Notes in Mathematics, 1032:298–310, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076803/.

Krzyz:1983:CCF

[KS83f] Jan G. Krzyż and Anil K. Soni. Close-to-convex functions with quasiconformal extension. Lecture Notes in Mathematics, 1039: 320–327, 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073373/.

Kappel:1984:IDS

[KS84a] Franz Kappel and Wilhelm Schappacher, editors. Infinite-Dimensional Systems: Proceedings of the Conference on Operator Semigroups and Applications held in Retzhof (Styria), Austria, June 5–11, 1983, volume 1076 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1076. URL http:/

/link.springer.com/book/10.1007/BFb0072759;   http://
www.springerlink.com/content/978-3-540-38932-3.

Kelley:1984:MIN

[KS84b] J. L. Kelley and T. P. Srinivasan. Measure and integral — a new gambit. Lecture Notes in Mathematics, 1089:120–126, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072608/.

Koh:1984:CSC

[KS84c] K. M. Koh and N. Sauer. Concentric subgraphs, closed subsets and dense graphs. Lecture Notes in Mathematics, 1073:100– 118, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073109/.

Kravchenko:1980:NIU

[KSS80] V. I. Kravchenko, Yu. D. Shevelev, and V. V. Shchennikov. Numerical investigation of unsteady viscous incompressible flow about bodies for varying conditions of their motion. Lecture Notes in Mathematics, 771:323–359, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086916/.

Kechris:1981:ADP

[KSS81] Alexander S. Kechris, Robert M. Solovay, and John R. Steel. The axiom of determinacy and the prewellordering property. Lecture Notes in Mathematics, 839:101–125, 1981. CODEN LN-MAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090237/.

Kamber:1982:HF

[KT82a] Franz W. Kamber and Philippe Tondeur. Harmonic foliations. Lecture Notes in Mathematics, 949:87–121, 1982. CODEN LN-MAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069758/.

Kerber:1982:CSC

[KT82b] Adalbert Kerber and Karl-Josef Thürlings. Counting symmetry classes of functions by weight and automorphism group. Lecture Notes in Mathematics, 969:191–211, 1982. CODEN LN-MAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062995/.

Katona:1983:EPE

[KT83] G. O. H. Katona and T. G. Tarján. Extremal problems with excluded subgraphs in the n-cube. Lecture Notes in Mathematics, 1018:84–93, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071617/.

Kurke:1984:SEV

[KT84] Herbert Kurke and Horst Theel. Some examples of vector bundles on the flag variety F (1, 2). Lecture Notes in Mathematics, 1056:187–254, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071774/.

Kublanovskaya:1983:ASS

[Kub83] V. N. Kublanovskaya. An approach to solving the spectral problem of A − λB. Lecture Notes in Mathematics, 973:17–29, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062092/.

Kueker:1981:EEM

[Kue81a] David W. Kueker. L_{∞_{ω_{1}}} -elementarily equivalent models of power ω_{1}. Lecture Notes in Mathematics, 859:120–131, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090944/.

Kuelbs:1981:LIL

[Kue81b] J. Kuelbs. The law of the iterated logarithm for Banach space valued random variables. Lecture Notes in Mathematics, 860: 268–278, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090621/.

Kuhlbrandt:1980:STP

[Küh80] Harold Kühlbrandt. On sharply 2-transitive permutation sets. Lecture Notes in Mathematics, 792:92–96, 1980. CODEN LN-MAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088667/.

Kuhnau:1983:BEE

[Küh83a] Reiner Kühnau. Bemerkungen zu Extremalproblemen bei in einem speziellen Mittel quasikonformen Abbildungen. (German) []. Lecture Notes in Mathematics, 1013:125–131, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066524/.

Kuhnau:1983:EIM

[Küh83b] Reiner Kühnau. Über Extremalprobleme bei im Mittel quasikonformen Abbildungen. (German) []. Lecture Notes in Mathematics, 1013:113–124, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066523/.

Kumar:1981:SRC

[Kum81] Shawpawn Das Kumar. Set-reconstruction of chain sizes in a class of finite topologies. Lecture Notes in Mathematics, 885: 227–236, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092266/.

Kummerer:1984:EMD

[Küm84a] Burkhard Kümmerer. Examples of Markov dilations over the 2 × 2 matrices. Lecture Notes in Mathematics, 1055:228–244, 1984. CODEN LNMAA2. ISBN 3-540-12915-4 (print), 3-540-38798-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071725/.

Kumpera:1984:CSE

[Kum84b] A. Kumpera. Sur la cohomologie des systèmes d’équations différentielles et des pseudogroupes de Lie. (French) []. Lecture Notes in Mathematics, 1045:84–94, 1984. CODEN LN-MAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072168/.

Kunita:1981:DSS

[Kun81a] Hiroshi Kunita. On the decomposition of solutions of stochastic differential equations. Lecture Notes in Mathematics, 851:

213–255, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088729/.

Kunita:1981:SEI

[Kun81b] Hiroshi Kunita. Some extensions of Itô’s formula. Lecture Notes in Mathematics, 850:118–141, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088362/.

Kunze:1981:HGG

[Kun81c] Michael Kunze. Die Halbgruppe der Grössenklassen eines geordneten lokalen Rings. (German) []. Lecture Notes in Mathematics, 855:19–25, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089805/.

Kunita:1982:SPD

[Kun82] Hiroshi Kunita. Stochastic partial differential equations connected with non-linear filtering. Lecture Notes in Mathematics, 972:100– 169, 1982. CODEN LNMAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064861/.

Kunita:1984:SDE

[Kun84] H. Kunita. Stochastic differential equations and stochastic flows of diffeomorphisms. Lecture Notes in Mathematics, 1097:143– 303, 1984. CODEN LNMAA2. ISBN 3-540-13897-8 (print), 3-540-39109-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099433/.

Kuo:1980:FWT

[Kuo80] Hui-Hsiung Kuo. Fourier–Wiener transform on Brownian functionals. Lecture Notes in Mathematics, 828:146–161, 1980. CO-DEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097400/.

Kupershmidt:1980:GJB

[Kup80] B. A. Kupershmidt. Geometry of jet bundles and the structure of Lagrangian and Hamiltonian formalisms. Lecture Notes in Mathematics, 775:162–218, 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092026/.

Kurepa:1982:SCF

[Kur82] Svetozar Kurepa. Semigroups and cosine functions. Lecture Notes in Mathematics, 948:47–72, 1982. CODEN LNMAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069841/.

Kuratowski:1983:PSC

[Kur83a] Kasimierz Kuratowski. On the problem of skew curves in topology [1]. Lecture Notes in Mathematics, 1018:1–13, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071605/.

Kursten:1983:LDU

[Kür83b] K.-D. Kürsten. Local duality of ultraproducts of Banach lattices. Lecture Notes in Mathematics, 991:137–142, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061566/.

Kurtz:1984:AWC

[Kur84] Thomas G. Kurtz. Approaches to weak convergence. Lecture Notes in Mathematics, 1091:173–183, 1984. CODEN LN-MAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098810/.

Kusuoka:1982:AFW

[Kus82] Shigeo Kusuoka. Analytic functionals of Wiener process and absolute continuity. Lecture Notes in Mathematics, 923:1–46, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093035/.

Kvaratskhelia:1980:UCR

[Kva80] V. V. Kvaratskhelia. On unconditional convergence of random series in Banach spaces. Lecture Notes in Mathematics, 828:162– 166, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097401/.

Kotani:1982:KST

[KW82] S. Kotani and S. Watanabe. Krein’s spectral theory of strings and generalized diffusion processes. Lecture Notes in Mathematics, 923:235–259, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093046/.

Kreitz:1984:UAC

[KW84] Christoph Kreitz and Klaus Weihrauch. A unified approach to constructive and recursive analysis. Lecture Notes in Mathematics, 1104:259–278, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099489/.

Kwasnik:1983:K

[Kwa83] Maria Kwaśnik. On the (k; l)-kernels. Lecture Notes in Mathematics, 1018:114–121, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071620/.

Kwong:1983:MRI

[Kwo83] Man Kam Kwong. Matrix Riccati inequality and oscillation of second order differential systems. Lecture Notes in Mathematics, 1032:311–327, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076804/.

Kaijser:1984:ITD

[KWP84] Sten Kaijser and Joan Wick-Pelletier. Interpolation theory and duality. Lecture Notes in Mathematics, 1070:152–168, 1984. CO-DEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099099/.

Koplik:1983:PCF

[KWW83] Joel Koplik, David Wilkinson, and Jorge F. Willemsen. Percolation and capillary fluid displacement. Lecture Notes in Mathematics, 1035:169–183, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073259/.

Kovenya:1980:IDS

[KY80] V. M. Kovenya and N. N. Yanenko. The implicit difference schemes for numerical solving the Navier–Stokes equations. Lecture Notes in Mathematics, 771:299–312, 1980. CODEN LN-MAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086914/.

Koh:1984:GTS

[KY84] Khee Meng Koh and Hian Poh Yap, editors. Graph Theory Singapore 1983: Proceedings of the First Southeast Asian Graph Theory Colloquium, held in Singapore May 10–28, 1983, volume 1073 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1073. URL http:/

/link.springer.com/book/10.1007/BFb0073099;   http://
www.springerlink.com/content/978-3-540-38924-8.

Kalashnikov:1983:SPS

[KZ83a] V. V. Kalashnikov and V. M. Zolotarev, editors. Stability Problems for Stochastic Models: Proceedings of the 6th International Seminar Held in Moscow, USSR, April 1982, volume 982 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 982. URL http:/

/link.springer.com/book/10.1007/BFb0082055;   http://
www.springerlink.com/content/978-3-540-39598-0.

Kwong:1983:NAS

[KZ83b] Man Kam Kwong and A. Zettl. A new approach to second order linear oscillation theory. Lecture Notes in Mathematics, 1032: 328–345, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076805/.

Lacher:1981:GTM

[Lac81] R. C. Lacher. Generalized three-manifolds. Lecture Notes in Mathematics, 870:82–92, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089709/.

Lachance:1984:BMI

[Lac84a] Michael A. Lachance. Bernstein and Markov inequalities for constrained polynomials. Lecture Notes in Mathematics, 1105:125– 135, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072405/.

Lacroix:1984:RSO

[Lac84b] Jean Lacroix. The random Schrödinger operator in a strip. Lecture Notes in Mathematics, 1064:280–297, 1984. CODEN LN-MAA2. ISBN 3-540-13341-0 (print), 3-540-38874-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073648/.

Lady:1981:RBH

[Lad81] E. L. Lady. Relations between hom, ext, and tensor product for certain categories of modules over Dedekind domains. Lecture Notes in Mathematics, 874:53–61, 1981. CODEN LN-MAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090523/.

Lady:1983:SSRa

[Lad83a] E. L. Lady. A seminar on splitting rings for torsion free modules over Dedekind domains. Lecture Notes in Mathematics, 1006:1–48, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21560-9_1/;   http://link.springer.
com/content/pdf/bfm:978-3-662-21560-9/1.pdf.

Lady:1983:SSRb

[Lad83b] E. L. Lady. A seminar on splitting rings for torsion free modules over Dedekind domains. Lecture Notes in Mathematics, 1006:1– 48, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103696/.

Lagarias:1981:CRA

[Lag81] J. C. Lagarias. A complement to Rident’s p-adic generalization of the Thue–Siegel–Roth theorem. Lecture Notes in Mathematics, 899:264–275, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096467/.

Lajos:1981:ICC

[Laj81] S. Lajos. (m, n)-Ideal characterizations of certain classes of semigroups. Lecture Notes in Mathematics, 855:26–32, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089806/.

Lambek:1982:TMC

[Lam82] J. Lambek. Toposes are monadic over categories. Lecture Notes in Mathematics, 962:153–166, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066895/.

Lam:1984:UAS

[Lam84] S. P. Lam. Unstable algebras over the Steenrod algebra. Lecture Notes in Mathematics, 1051:374–392, 1984. CODEN LN-MAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075579/.

Lanckau:1980:GVO

[Lan80a] Eberhard Lanckau. General Vekua operators. Lecture Notes in Mathematics, 798:301–311, 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097273/.

Landrock:1980:SRL

[Lan80b] Peter Landrock. Some remarks on Loewy lengths of projective modules. Lecture Notes in Mathematics, 832:369–381, 1980. CO-DEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088474/.

Langevin:1980:SCP

[Lan80c] R. Langevin. Singularités complexes, points critiques et intégrales de courbure. (French) []. Lecture Notes in Mathematics, 822:129– 143, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097754/.

Lannes:1980:FEP

[Lan80d] Jean Lannes. Un faux espace projectif réel de dimension 4 [d’après Sylvain E. Cappell et Julius L. Shaneson]. (French) []. Lecture Notes in Mathematics, 770:36–56, 1980. CODEN LN-MAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096234/.

Lander:1981:CBT

[Lan81a] E. S. Lander. Characterization of biplanes by their automorphism groups. Lecture Notes in Mathematics, 893:204–218, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091021/.

Lanford:1981:STI

[Lan81b] Oscar E. Lanford III. Smooth transformations of intervals. Lecture Notes in Mathematics, 901:36–54, 1981. CODEN LN-MAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097188/.

Lancaster:1982:GHM

[Lan82a] Peter Lancaster. Generalized Hermitian matrices: a new frontier for numerical analysis? Lecture Notes in Mathematics, 912: 179–189, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093157/.

Lange:1982:FEA

[Lan82b] L. J. Lange. δ-fraction expansions of analytic functions. Lecture Notes in Mathematics, 932:152–175, 1982. CODEN LN-MAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093312/.

Langer:1982:SFD

[Lan82c] Heinz Langer. Spectral functions of definitizable operators in Krein spaces. Lecture Notes in Mathematics, 948:1–46, 1982. CODEN LNMAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069840/.

Langevin:1984:EGI

[Lan84a] R. Langevin. Énergies et géométrie intégrale. (French) []. Lecture Notes in Mathematics, 1045:95–103, 1984. CODEN LN-MAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072169/.

Lannes:1984:IBF

[Lan84b] J. Lannes. Sur les immersions de Boy. (French) []. Lecture Notes in Mathematics, 1051:263–270, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0075571/.

Lascoux:1980:PKR

[Las80a] A. Lascoux. Produit de Kronecker des représentations du groupe symétrique. (French) []. Lecture Notes in Mathematics, 795:319– 329, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090125/.

Lassoued:1980:EDA

[Las80b] Mme L. Lassoued. Espaces de Dirichlet et axiomatique Brelot. (French) []. Lecture Notes in Mathematics, 814:117–152, 1980. CODEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094152/.

Lascar:1981:SFS

[Las81a] Bernard Lascar. Solutions faibles et solutions fortes du problème ¯∂u = f où f est une fonction à croissance polynômiale sur un espace de Hilbert. (French) []. Lecture Notes in Mathematics, 843: 405–436, 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089284/.

Lascar:1981:PMPb

[Las81b] Richard Lascar. Paramétrices microlocales de problèmes aux limites pour une classe d’équations pseudo-différentielles à caractéristiques de multiplicité variable. (French) []. Lecture Notes in Mathematics, 856:106–221, 1981. CODEN LNMAA2. ISBN 3-540-10702-9 (print), 3-540-38658-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089186/.

Lascar:1981:PMPa

[Las81c] Richard Lascar. Paramétrices microlocales pour un problème de Cauchy hyperbolique à caractéristiques de multiplicité variable. (French) []. Lecture Notes in Mathematics, 856:27–105, 1981. CODEN LNMAA2. ISBN 3-540-10702-9 (print), 3-540-38658-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089185/.

Lascar:1981:PSS

[Las81d] Richard Lascar. Propagation des Singularités des Solutions d’équations Pseudo–Différentielles à Caractéristiques de Multiplicités Variables. (French) [Propagation of singularities of solutions of pseudo-differential equations with features of variable multiplicities], volume 856 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10702-9 (print), 3-540-38658-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 237 pp. LCCN QA3 .L28 no. 856. URL http:/

/link.springer.com/book/10.1007/BFb0089183;   http://
www.springerlink.com/content/978-3-540-38658-2.

Lascar:1981:PSP

[Las81e] Richard Lascar. Propagation des singularités pour des opérateurs pseudo-différentiels à caractéristiques de multiplicité variable. (French) []. Lecture Notes in Mathematics, 856:1–26, 1981. CO-DEN LNMAA2. ISBN 3-540-10702-9 (print), 3-540-38658-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089184/.

Lascar:1981:REP

[Las81f] Richard Lascar. Une relation entre la propagation des singularités et la propagation du support pour des opérateurs hyperboliques. (French) []. Lecture Notes in Mathematics, 856:222–236, 1981. CODEN LNMAA2. ISBN 3-540-10702-9 (print), 3-540-38658-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089187/.

Lasiecka:1984:AAD

[Las84] I. Lasiecka. Approximations of analytic and differentiable semigroups — rate of convergence with nonsmooth initial conditions. Lecture Notes in Mathematics, 1076:123–138, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072772/.

Laub:1980:CKS

[Lau80a] Jesper Laub. Convolution kernels satisfying the domination principle. Lecture Notes in Mathematics, 814:153–165, 1980. CODEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094153/.

Laub:1980:RSP

[Lau80b] Jesper Laub. A remark on the singular part of a convolution kernel satisfying the domination principle. Lecture Notes in Mathematics, 787:191–193, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086334/.

Laugwitz:1983:CGF

[Lau83a] Detlef Laugwitz. Θ-calculus as a generalization of field extension an alternative approach to nonstandard analysis. Lecture Notes in Mathematics, 983:120–133, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065337/.

Laumon:1983:CDM

[Lau83b] G. Laumon. Sur la catégorie dérivées des D-modules filtrés. (French) []. Lecture Notes in Mathematics, 1016:151–237, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099964/.

Laumon:1983:VCB

[Lau83c] G. Laumon. Vanishing cycles over a base of dimension ≥ 1. Lecture Notes in Mathematics, 1016:143–150, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099963/.

Laursen:1983:ISR

[Lau83d] K. B. Laursen. Ideal structure in radical sequence algebras. Lecture Notes in Mathematics, 975:248–257, 1983. CODEN LN-MAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064554/.

Lawrynowicz:1980:AFK

[Ław80] Julian Ławrynowicz, editor. Analytic Functions Kozubnik 1979: Proceedings of a Conference Held in Kozubnik, Poland, April 19–25, 1979, volume 798 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09985-9 (print), 3-540-39247-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 798. URL http:/

/link.springer.com/book/10.1007/BFb0097253;   http://
www.springerlink.com/content/978-3-540-39247-7.

Lawniczak:1982:GMO

[Ław82] Anna T. Ławniczak. Gaussian measures on Orlicz spaces and abstract Wiener spaces. Lecture Notes in Mathematics, 939:81– 97, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096261/.

Lawrynowicz:1983:AFB

[Ław83] Julian Ławrynowicz, editor. Analytic Functions Błażejewko 1982: Proceedings of a Conference held in Błażejewko, Poland, August 19–27, 1982, volume 1039 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12712-7 (print), 3-540-38697-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1039. URL http:/

/link.springer.com/book/10.1007/BFb0073353;   http://
www.springerlink.com/content/978-3-540-38697-1.

Lazarsfeld:1984:SAT

[Laz84] Robert Lazarsfeld. Some applications of the theory of positive vector bundles. Lecture Notes in Mathematics, 1092:29–61, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099356/.

Lagha-Benabdallah:1984:LEF

[LB84] A. Lagha-Benabdallah. Limites des équations d’un fluide compressible lorsque la compressibilité tend vers zéro. (French) []. Lecture Notes in Mathematics, 1047:139–165, 1984. CODEN LNMAA2. ISBN 3-540-12893-X (print), 3-540-38773-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072329/.

Levy-Bruhl-Laperriere:1981:SDR

[LBL81] Anne Lévy-Bruhl-Laperrière. Spectre du de Rham–Hodge sur l’espace projectif quaternionique. (French) []. Lecture Notes in Mathematics, 867:98–129, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090382/.

Lesieur:1980:ITS

[LC80] Marcel Lesieur and Jean-Pierre Chollet. Introduction aux théories statistiques de la turbulence pleinement développée. (French) [].

Lecture Notes in Mathematics, 782:101–121, 1980. CODEN LN-MAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090429/.

Lelong:1983:SAP

[LDS83] Pierre Lelong, Pierre Dolbeault, and Henri Skoda, editors. Séminaire d’Analyse P.Lelong–P.Dolbeault–H. Skoda: Années 1981/1983. (French) [], volume 1028 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1028. URL http:/

/link.springer.com/book/10.1007/BFb0071674;   http://
www.springerlink.com/content/978-3-540-38724-4.

LeJan:1982:DMS

[Le 82a] Yves Le Jan. Dual Markovian semigroups and processes. Lecture Notes in Mathematics, 923:47–75, 1982. CODEN LN-MAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093036/.

LePage:1982:TLP

[Le 82b] Émile Le Page. Théorèmes limitès pour les produits de matrices aléatoires. (French) []. Lecture Notes in Mathematics, 928: 258–303, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093229/.

LeGall:1983:ADT

[Le 83a] J.-F. Le Gall. Applications du temps local aux équations différentielles stochastiques unidimensionnelles. (French) []. Lecture Notes in Mathematics, 986:15–31, 1983. CODEN LN-MAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068296/.

LePotier:1983:FVS

[Le 83b] J. Le Potier. Fibrés vectoriels sur les surfaces K3. (French) []. Lecture Notes in Mathematics, 1028:225–238, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071684/.

LeGall:1984:ODS

[Le 84a] J.-F. Le Gall. One-dimensional stochastic differential equations involving the local times of the unknown process. Lecture Notes in Mathematics, 1095:51–82, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099122/.

LeJan:1984:FCL

[Le 84b] Yves Le Jan. Fonctions “cad-lag” sur les trajectoires d’un processus de Ray. (French) []. Lecture Notes in Mathematics, 1096: 412–418, 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100122/.

Leandre:1983:ETF

[Léa83] R. Léandre. Un exemple en théorie des flots stochastiques. (French) []. Lecture Notes in Mathematics, 986:158–161, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068310/.

Ledoux:1981:IBP

[Led81a] Michel Ledoux. Inégalités de Burkholder pour martingales indexées par N×N. (French) []. Lecture Notes in Mathematics, 863: 122–127, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091096/.

Ledoux:1981:LDL

[Led81b] Michel Ledoux. La loi du logarithme itéré bornée dans les espaces de Banach. (French) []. Lecture Notes in Mathematics, 850: 11–37, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088358/.

Ledoux:1983:TLC

[Led83] Michel Ledoux. Sur les théorèmes limités dans certains espaces de Banach lisses. (French) [On some limit theorems in smooth Banach spaces]. Lecture Notes in Mathematics, 990:150–169, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064269/.

Ledoux:1984:RCL

[Led84a] M. Ledoux. A remark on the central limit theorem in Banach spaces. Lecture Notes in Mathematics, 1080:144–151, 1984. CO-DEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099791/.

Ledrappier:1984:QPE

[Led84b] F. Ledrappier. Quelques propriétés des exposants caractéristiques. (French) []. Lecture Notes in Mathematics, 1097:305–396, 1984. CODEN LNMAA2. ISBN 3-540-13897-8 (print), 3-540-39109-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099434/.

Legrand:1982:FP

[Leg82a] André Legrand. B-fibres principaux. Lecture Notes in Mathematics, 941:17–42, 1982. CODEN LNMAA2. ISBN 3-540-11575-7 (print), 3-540-39303-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094694/.

Legrand:1982:DSS

[Leg82b] André Legrand. Différentbielle de la suite spectrale de Shih. (French) []. Lecture Notes in Mathematics, 941:109–127, 1982. CODEN LNMAA2. ISBN 3-540-11575-7 (print), 3-540-39303-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094697/.

Legrand:1982:ESA

[Leg82c] André Legrand. Ensembles simpliciaux au-dessus de B. (French) []. Lecture Notes in Mathematics, 941:1–16, 1982. CODEN LN-MAA2. ISBN 3-540-11575-7 (print), 3-540-39303-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094693/.

Legrand:1982:FGF

[Leg82d] André Legrand. Fibres en groupes de fibre de type K(π, n). (French) []. Lecture Notes in Mathematics, 941:43–83, 1982. CO-DEN LNMAA2. ISBN 3-540-11575-7 (print), 3-540-39303-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094695/.

Legrand:1982:HESa

[Leg82e] André Legrand. Homotopie de l’espace des sections d’un fibre en groupes. (French) []. Lecture Notes in Mathematics, 941:84– 108, 1982. CODEN LNMAA2. ISBN 3-540-11575-7 (print), 3-540-39303-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094696/.

Legrand:1982:HESb

[Leg82f] André Legrand. Homotopie des Espaces de Sections. (French) [Homotopy of section spaces], volume 941 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11575-7 (print), 3-540-39303-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 132 pp. LCCN QA3 .L28 no. 941. URL http:/

/link.springer.com/book/10.1007/BFb0094692;   http://
www.springerlink.com/content/978-3-540-39303-0.

Lehel:1983:CTH

[Leh83a] J. Lehel. A covering theorem for hypergraphs. Lecture Notes in Mathematics, 1018:122–124, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071621/.

Lehto:1983:RNM

[Leh83b] Olli Lehto. On Rolf Nevanlinna’s mathematical work and on his role, together with Simion Stoilow, as a promotor of Romanian– Finnish mathematical relations. Lecture Notes in Mathematics, 1013:1–7, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066514/.

Leichtweiss:1981:BCE

[Lei81] K. Leichtweiß. E. B. Christoffels Einfluss auf die Geometrie. (German) []. Lecture Notes in Mathematics, 838:1–11, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088836/.

Lelong:1980:PCC

[Lel80] Pierre Lelong. Potentiels canoniques et comparaison de deux méthodes pour la résolution du ∂ ¯∂ à croissance. (French) []. Lecture Notes in Mathematics, 822:144–168, 1980. CODEN LN-MAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097755/.

Lelong:1982:CDN

[Lel82a] Pierre Lelong. Calcul du nombre densité ν (x,f) et lemme de Schwarz pour les fonctions plurisousharmoniques dans un espace vectoriel topologique. (French) []. Lecture Notes in Mathematics, 919:167–176, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097048/.

LeLong:1982:RPL

[LeL82b] Pierre LeLong. Réponse de Pierre Lelong. (French) []. Lecture Notes in Mathematics, 919:190–191, 1982. CODEN LN-MAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097051.

Lelong:1983:DAO

[Lel83] Pierre Lelong. Discontinuité et annulation de l’opérateur de Monge–Ampère complexe. (French) []. Lecture Notes in Mathematics, 1028:219–224, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071683/.

Lembcke:1980:MET

[Lem80] Jörn Lembcke. On a measure extension theorem of Bierlein. Lecture Notes in Mathematics, 794:45–48, 1980. CODEN LN-MAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088211/.

Lemaire:1981:ALE

[Lem81a] Jean-Michel Lemaire. Anneaux locaux et espaces de lacets à séries de Poincaré irrationnelles [d’après Anick, Roos, etc. …]. (French) []. Lecture Notes in Mathematics, 901:149–156, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097195/.

Lemaire:1981:EAH

[Lem81b] Luc Lemaire. Existence des applications harmoniques et courbure des variétés. (French) []. Lecture Notes in Mathematics, 842: 174–195, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089934/.

Lembcke:1982:SPM

[Lem82] Jörn Lembcke. On simultaneous preimage measures on Hausdorff spaces. Lecture Notes in Mathematics, 945:110–115, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096666/.

Lenglart:1980:AEP

[Len80a] E. Lenglart. Appendice à l’exposé précédent: Inégalités de semi martingales. (French) []. Lecture Notes in Mathematics, 784:49– 52, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089468/.

Lenglart:1980:IMP

[Len80b] E. Lenglart. Sur l’inégalité de Metivier–Pellaumail. (French) []. Lecture Notes in Mathematics, 784:125–127, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089477/.

Lenstra:1981:PTA

[Len81] H. W. Lenstra, Jr. Primality testing algorithms [after Adleman, Rumely and Williams]. Lecture Notes in Mathematics, 901:243– 257, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097201/.

Lenski:1984:EQT

[Len84] Wolfgang Lenski. Elimination of quantifiers for the theory of Archimedean ordered divisible groups in a logic with Ramsey quantifiers. Lecture Notes in Mathematics, 1103:261–280, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099390/.

Leptin:1980:BLG

[Lep80] Horst Leptin. Bemerkungen über Linksideale in Gruppenalgebren. (German) []. Lecture Notes in Mathematics, 781:121– 141, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097651/.

LePage:1981:MID

[LeP81] Raoul LePage. Multidimensional infinitely divisible variables and processes Part II. Lecture Notes in Mathematics, 860:279– 284, 1981. CODEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090622/.

Lepowsky:1982:ALA

[Lep82] J. Lepowsky. Affine Lie algebras and combinatorial identities. Lecture Notes in Mathematics, 933:130–156, 1982. CODEN LN-MAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093358/.

Lerman:1981:RLO

[Ler81] Manuel Lerman. On recursive linear orderings. Lecture Notes in Mathematics, 859:132–142, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090945/.

Lesieur:1980:APP

[Les80] Léonce Lesieur. Sur les anneaux premiers principaux à gauche. (French) []. Lecture Notes in Mathematics, 795:191–208, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090120/.

Lescot:1983:SBP

[Les83] Jack Lescot. La serie de bass d’un produit fibre d’anneaux locaux. (French) []. Lecture Notes in Mathematics, 1029:218–239, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098933/.

Leskow:1984:DVL

[Leś84] Jacek Leśkow. On different versions of the law of iterated logarithm for R^{∞} and 1_{p} valued Wiener process. Lecture Notes in Mathematics, 1080:152–161, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099792/.

Letac:1981:PCP

[Let81] Gérard Letac. Problèmes classiques de probabilité sur un couple de Gelfand. (French) []. Lecture Notes in Mathematics, 861: 93–120, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097318/.

Leutwiler:1980:HFB

[Leu80] Heinz Leutwiler. Harmonic functions of bounded mean oscillation and a generalization to vector lattices of continuous functions. Lecture Notes in Mathematics, 787:194–208, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086335/.

Leung:1982:SHM

[Leu82] Pui-Fai Leung. On the stability of harmonic maps. Lecture Notes in Mathematics, 949:122–129, 1982. CODEN LNMAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069759/.

Levasseur:1980:IPC

[Lev80a] Thierry Levasseur. Idéaux premiers et completion dans les algèbres enveloppantes d’algèbres de Lie nilpotentes. (French) []. Lecture Notes in Mathematics, 795:116–160, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090116/.

Levi:1980:PFR

[Lev80b] Mark Levi. Periodically forced relaxation oscillations. Lecture Notes in Mathematics, 819:300–317, 1980. CODEN LN-MAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086995/.

Levine:1980:FMb

[Lev80c] Jerome P. Levine. Δ_{i} of finite modules. Lecture Notes in Mathematics, 772:8–9, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091576.

Levine:1980:OT

[Lev80d] Jerome P. Levine. Π-only torsion. Lecture Notes in Mathematics, 772:21–23, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091581/.

Levine:1980:ASK

[Lev80e] Jerome P. Levine. Algebraic Structure of Knot Modules, volume 772 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 104 pp. LCCN QA3 .L28 no. 772. URL http:

//link.springer.com/book/10.1007/BFb0091572;   http:/
/www.springerlink.com/content/978-3-540-38555-4.

Levine:1980:CFD

[Lev80f] Jerome P. Levine. Classification fails in degree 4. Lecture Notes in Mathematics, 772:46–47, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091590.

Levine:1980:CPM

[Lev80g] Jerome P. Levine. Classification of Π-primary modules. Lecture Notes in Mathematics, 772:40–46, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091589/.

Levine:1980:CDP

[Lev80h] Jerome P. Levine. Classification of derived product structure. Lecture Notes in Mathematics, 772:16–18, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091578/.

Levine:1980:CEM

[Lev80i] Jerome P. Levine. Classification of elementary modules. Lecture Notes in Mathematics, 772:36–39, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091587/.

Levine:1980:CPS

[Lev80j] Jerome P. Levine. Classification of product structure. Lecture Notes in Mathematics, 772:53–59, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091592/.

Levine:1980:CP

[Lev80k] Jerome P. Levine. Completion of proof. Lecture Notes in Mathematics, 772:39–40, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091588.

Levine:1980:DES

[Lev80l] Jerome P. Levine. The derived exact sequences. Lecture Notes in Mathematics, 772:1–4, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091573/.

Levine:1980:FMa

[Lev80m] Jerome P. Levine. Finite modules. Lecture Notes in Mathematics, 772:4–6, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091574/.

Levine:1980:HEM

[Lev80n] Jerome P. Levine. Homogeneous and elementary modules. Lecture Notes in Mathematics, 772:32–34, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091585/.

Levine:1980:ICD

[Lev80o] Jerome P. Levine. Inductive construction of derived sequences. Lecture Notes in Mathematics, 772:24–26, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091583/.

Levine:1980:IRD

[Lev80p] Jerome P. Levine. Inductive recovery of derived sequences. Lecture Notes in Mathematics, 772:26–32, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091584/.

Levine:1980:PSP

[Lev80q] Jerome P. Levine. Product structure on Π-primary modules. Lecture Notes in Mathematics, 772:48–53, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091591/.

Levine:1980:PSF

[Lev80r] Jerome P. Levine. Product structure on finite modules. Lecture Notes in Mathematics, 772:9–16, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091577/.

Levine:1980:RI

[Lev80s] Jerome P. Levine. Rational invariants. Lecture Notes in Mathematics, 772:18–20, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091579/.

Levine:1980:REM

[Lev80t] Jerome P. Levine. Realization of elementary modules. Lecture Notes in Mathematics, 772:34–36, 1980. CODEN LN-MAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091586/.

Levine:1980:RFM

[Lev80u] Jerome P. Levine. Realization of finite modules. Lecture Notes in Mathematics, 772:6–8, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091575/.

Levine:1980:SRT

[Lev80v] Jerome P. Levine. Statement of realization theorem. Lecture Notes in Mathematics, 772:23–24, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091582.

Levine:1980:TFM

[Lev80w] Jerome P. Levine. Z-torsion-free modules. Lecture Notes in Mathematics, 772:20–21, 1980. CODEN LNMAA2. ISBN 3-540-09739-2 (print), 3-540-38555-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091580.

Levasseur:1981:AOD

[Lev81] Thierry Levasseur. Anneaux d’opérateurs différentiels. (French) []. Lecture Notes in Mathematics, 867:157–173, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090386/.

Levasseur:1982:DKA

[Lev82a] Thierry Levasseur. Sur la dimension de Krull de l’algèbre enveloppante d’une algèbre de Lie semi-simple. (French) []. Lecture Notes in Mathematics, 924:173–183, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092932/.

Levine:1982:DIT

[Lev82b] Marc Levine. Deformations of irregular threefolds. Lecture Notes in Mathematics, 947:269–286, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093595/.

Lewowicz:1982:SET

[Lew82] Jorge Lewowicz. Sobre estabilidad topologica. (Spanish) [On topological stability]. Lecture Notes in Mathematics, 957:203– 211, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066239/.

Lex:1983:SA

[Lex83] W. Lex. Simple acts. Lecture Notes in Mathematics, 998:339– 349, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062040/.

Lutz:1981:EPN

[LG81a] Robert Lutz and Michel Goze. Elementary practice of non standard analysis. Lecture Notes in Mathematics, 881:1–37, 1981.

CODEN LNMAA2. ISBN 3-540-10879-3 (print), 3-540-70808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093398/.

Lutz:1981:LFN

[LG81b] Robert Lutz and Michel Goze. Logical foundations of nonstandard analysis. Lecture Notes in Mathematics, 881:38–79, 1981. CODEN LNMAA2. ISBN 3-540-10879-3 (print), 3-540-70808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093399/.

Lutz:1981:NSA

[LG81c] Robert Lutz and Michel Goze. Non standard analysis as a tool in perturbation problems. Lecture Notes in Mathematics, 881: 126–252, 1981. CODEN LNMAA2. ISBN 3-540-10879-3 (print), 3-540-70808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093401/.

Lutz:1981:NAP

[LG81d] Robert Lutz and Michel Goze. Nonstandard Analysis: A Practical Guide with Applications, volume 881 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10879-3 (print), 3-540-70808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 261 pp. LCCN QA3 .L28 no. 881.

URL http://link.springer.com/book/10.1007/BFb0093397;
 http://www.springerlink.com/content/978-3-540-70808-
7.

Lutz:1981:SCT

[LG81e] Robert Lutz and Michel Goze. Some classical topics from a nonstandard point of view. Lecture Notes in Mathematics, 881:80– 125, 1981. CODEN LNMAA2. ISBN 3-540-10879-3 (print), 3-540-70808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093400/.

Liao:1984:PCL

[Lia84] Ming Liao. Path continuity and last exit distributions. Lecture Notes in Mathematics, 1059:119–126, 1984. CODEN LN-MAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100036/.

Lichnerowicz:1980:DQ

[Lic80a] André Lichnerowicz. Deformations and quantization. Lecture Notes in Mathematics, 775:105–121, 1980. CODEN LN-MAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092023/.

Lichnerowicz:1980:EED

[Lic80b] André Lichnerowicz. Existence et équivalence de déformations associatives associées à une variété symplectique. (French) []. Lecture Notes in Mathematics, 836:177–185, 1980. CODEN LN-MAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089734/.

Lichnerowicz:1984:GCA

[Lic84a] A. Lichnerowicz. Geometry and cohomologies associated with a contact manifold. Lecture Notes in Mathematics, 1045:104– 116, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072170/.

Lichtenbaum:1984:VZF

[Lic84b] S. Lichtenbaum. Values of zeta-functions at non-negative integers. Lecture Notes in Mathematics, 1068:127–138, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099447/.

Lieb:1980:LVC

[Lie80] I. Lieb. L’opérateur ¯∂ sur une variété q-concave. (French) []. Lecture Notes in Mathematics, 822:169–173, 1980. CODEN LN-MAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097756/.

Lieberman:1982:HVF

[Lie82] David I. Lieberman. Holomorphic vector fields and rationality. Lecture Notes in Mathematics, 956:99–117, 1982. CODEN LNMAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101512/.

Lin:1981:WLC

[Lin81] T. Y. Lin. Wu-like classes and generalized Peterson–Stein classes. Lecture Notes in Mathematics, 886:208–248, 1981. CODEN LN-MAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092186/.

Lindberg:1982:CDC

[Lin82a] Bengt Lindberg. Compact deferred correction formulas. Lecture Notes in Mathematics, 968:220–233, 1982. CODEN LN-MAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064890/.

Lindel:1982:PMP

[Lin82b] Hartmut Lindel. On projective modules over polynomial rings over regular rings. Lecture Notes in Mathematics, 966:169– 179, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062174/.

Lindstrom:1983:SIH

[Lin83] Tom L. Lindstrøm. Stochastic integration in hyperfinite dimensional linear spaces. Lecture Notes in Mathematics, 983:134– 161, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065338/.

Linde:1984:ESL

[Lin84a] W. Linde. Extensions of the Slepian lemma to p-stable measures. Lecture Notes in Mathematics, 1080:162–169, 1984. CO-DEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099793/.

Lindstrom:1984:FI

[Lin84b] Per Lindström. On faithful interpretability. Lecture Notes in Mathematics, 1104:279–288, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099490/.

Lions:1983:COS

[Lio83] J.-L. Lions. Contrôle optimal de systèmes à états multiples. (French) []. Lecture Notes in Mathematics, 979:202–215, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066255/.

Lipsman:1983:EGW

[Lip83] Ronald L. Lipsman. On the existence of a generalized Weil representation. Lecture Notes in Mathematics, 1020:161–178, 1983. CODEN LNMAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071501/.

Little:1982:IDP

[Lit82] Charles H. C. Little. An interesting decomposition of k_{4n,4n} into planar subgraphs. Lecture Notes in Mathematics, 952:353– 357, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061989/.

Liulevicius:1980:ETH

[Liu80] Arunas Liulevicius. Equivariant K-theory and homotopy rigidity. Lecture Notes in Mathematics, 788:340–358, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099253/.

Liu:1984:PTA

[Liu84] Shih-Chao Liu. A proof-theoretic approach to non standard analysis (continued). Lecture Notes in Mathematics, 1103:281– 296, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099391/.

Lejeune-Jalabert:1980:AAR

[LJ80] M. Lejeune-Jalabert. Arcs analytiques et resolution minimale des singularités des surfaces quasi homogènes. (French) []. Lecture Notes in Mathematics, 777:303–336, 1980. CODEN LN-MAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085890/.

Lejeune-Jalabert:1982:LRF

[LJ82] M. Lejeune-Jalabert. Liaison et residu. (French) []. Lecture Notes in Mathematics, 961:233–240, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071285/.

Ljapin:1981:POT

[Lja81] E. S. Ljapin. Partielle Operationen in der Theorie der Halbgruppen. (German) []. Lecture Notes in Mathematics, 855:33–48, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089807/.

Lawrynowicz:1983:BCT

[ŁK83a] Julian Ławrynowicz and Jan Krzyź. Basic concepts and theorems in the analytic theory of quasiconformal mappings. Lecture Notes in Mathematics, 978:3–52, 1983. CODEN LNMAA2. ISBN 3-540-11989-2 (print), 3-540-39464-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065378/.

Lawrynowicz:1983:NA

[ŁK83b] Julian Ławrynowicz and Jan Krzyź. Notation and abbreviations. Lecture Notes in Mathematics, 978:1–2, 1983. CODEN LNMAA2. ISBN 3-540-11989-2 (print), 3-540-39464-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065377.

Lawrynowicz:1983:PM

[ŁK83c] Julian Ławrynowicz and Jan Krzyź. The parametrical methods. Lecture Notes in Mathematics, 978:53–134, 1983. CODEN LNMAA2. ISBN 3-540-11989-2 (print), 3-540-39464-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065379/.

Lawrynowicz:1983:QMP

[ŁK83d] Julian Ławrynowicz and Jan Krzyź. Quasiconformal Mappings in the Plane: Parametrical Methods, volume 978 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-11989-2 (print), 3-540-39464-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 178 pp. LCCN QA3 .L28 no. 978. URL http:/

/link.springer.com/book/10.1007/BFb0065376;   http://
www.springerlink.com/content/978-3-540-39464-8.

Lawrynowicz:1983:RVM

[ŁK83e] Julian Ławrynowicz and Jan Krzyź. A review of variational methods and basic applications in electrical engineering. Lecture Notes in Mathematics, 978:135–162, 1983. CODEN LNMAA2. ISBN 3-540-11989-2 (print), 3-540-39464-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065380/.

Lehmer:1981:CNS

[LL81] D. H. Lehmer and Emma Lehmer. Cyclotomy for non-squarefree modul I. Lecture Notes in Mathematics, 899:276–300, 1981. CO-DEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096468/.

Lorentz:1984:MI

[LL84] G. G. Lorentz and R. A. Lorentz. Multivariate interpolation. Lecture Notes in Mathematics, 1105:136–144, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072406/.

Lloyd:1983:SAL

[Llo83] N. G. Lloyd. Small amplitude limit cycles of polynomial differential equations. Lecture Notes in Mathematics, 1032:346–357, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076806/.

Lenglart:1980:PUC

[LLP80] E. Lenglart, D. Lépingle, and M. Pratelli. Présentation unifiée de certaines inégalités de la théorie des martingales. (French)

[]. Lecture Notes in Mathematics, 784:26–48, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089467/.

Lambert:1980:NMP

[LM80] J. D. Lambert and R. J. Y. McLeod. Numerical methods for phase-plane problems in ordinary differential equations. Lecture Notes in Mathematics, 773:83–97, 1980. CODEN LN-MAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094165/.

Liukkonen:1983:MAL

[LM83] John R. Liukkonen and Michael Mislove. Measure algebras of locally compact semilattices. Lecture Notes in Mathematics, 998: 202–214, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062030/.

Lambert:1984:SWP

[LM84a] F. Lambert and M. Musette. Solitary waves, padéons and solitons. Lecture Notes in Mathematics, 1071:197–212, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099619/.

Lisica:1984:CPS

[LM84b] Ju. T. Lisica and S. Mardešić. Coherent prohomotopy and a strong shape category of topological spaces. Lecture Notes in Mathematics, 1060:164–173, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099932/.

Li:1988:MLN

[LM88] Caizhong Li and Libin Mou. A mistake in “Lecture Notes in Mathematics 922”. Journal of Chengdu University of Science and Technology, 6(??):79–82, 1988. CODEN CKDXDB. ISSN 0253-

Levy:1982:TGO

[LMGC82] A. V. Lévy, A. Montalvo, S. Gómez, and A. Calderon. Topics in global optimization. Lecture Notes in Mathematics, 909: 18–33, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092957/.

London:1982:MASa

[LMRW82a] R. R. London, H. P. McKean, L. C. G. Rogers, and David Williams. A martingale approach to some Wiener–Hopf problems, I. Lecture Notes in Mathematics, 920:41–67, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092769/.

London:1982:MASb

[LMRW82b] R. R. London, H. P. McKean, L. C. G. Rogers, and David Williams. A martingale approach to some Wiener–Hopf problems, II. Lecture Notes in Mathematics, 920:68–90, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092770/.

Linde:1980:SMA

[LMW80] W. Linde, V. Mandrekar, and A. Weron. p-stable measures and p-absolutely summing operators. Lecture Notes in Mathematics, 828:167–178, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097402/.

Liu:1983:AED

[LO83] Kam-Moon Liu and Eduardo L. Ortiz. Approximation of eigenvalues defined by ordinary differential equations with the tau method. Lecture Notes in Mathematics, 973:90–102, 1983. CO-DEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062096/.

Lagarias:1984:NAC

[LO84] J. C. Lagarias and A. M. Odlyzko. New algorithms for computing π(x). Lecture Notes in Mathematics, 1052:176–193, 1984. CODEN LNMAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071543/.

Loday:1981:BM

[Lod81] Jean-Louis Loday. On the boundary map K_{3}(Δ/I) → K_{2}(Δ, I). Lecture Notes in Mathematics, 854:262–268, 1981. CODEN LN-MAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089525/.

Loeb:1980:RBS

[Loe80] Peter A. Loeb. A regular boundary supporting representing measures of bounded functions in a Bauer harmonic space. Lecture Notes in Mathematics, 787:209–211, 1980. CODEN LN-MAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086336/.

Loffler:1980:HIS

[Löf80] Peter Löffler. Homotopielineare Involutionen auf Sphären. (German) []. Lecture Notes in Mathematics, 788:359–363, 1980. CO-DEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099254/.

Lohman:1983:ESD

[Loh83] Robert H. Lohman. On the existence of spheres and dual spheres without gap points. Lecture Notes in Mathematics, 991:143– 148, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061567/.

Longman:1981:DCA

[Lon81] I. M. Longman. Difficulties of convergence acceleration. Lecture Notes in Mathematics, 888:273–289, 1981. CODEN LN-MAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095593/.

Londen:1984:AER

[Lon84] S.-O. Londen. Asymptotic estimates for resolvents of some integral equations. Lecture Notes in Mathematics, 1076:139–146, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072773/.

Looijenga:1981:TTK

[Loo81a] Eduard Looijenga. A Torelli theorem for Kähler–Einstein K3 surfaces. Lecture Notes in Mathematics, 894:107–112, 1981. CODEN LNMAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096226/.

Lootgieter:1981:RMT

[Loo81b] J. C. Lootgieter. Sur un résultat de M. Talagrand. (French) []. Lecture Notes in Mathematics, 850:276–277, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088374.

Lopera:1984:NSF

[Lop84] J. F. T. Lopera. A note on semisimple flat homogeneous spaces. Lecture Notes in Mathematics, 1045:117–124, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072171/.

Lorenz:1981:PIG

[Lor81a] Martin Lorenz. Prime ideals in group algebras of polycyclic-by-finite groups: Vertices and sources. Lecture Notes in Mathematics, 867:406–420, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090396/.

Lorimer:1981:CFP

[Lor81b] Peter Lorimer. The construction of finite projective planes. Lecture Notes in Mathematics, 884:64–76, 1981. CODEN LN-MAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091808/.

Lorenz:1982:NBV

[Lor82] Priv. Doz. Dr. Jens Lorenz. Nonlinear boundary value problems with turning points and properties of difference schemes. Lecture Notes in Mathematics, 942:150–169, 1982. CODEN LN-MAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094746/.

Lorch:1983:PCC

[Lor83] Edgar R. Lorch. Problems in the classification of certain compact spaces. Lecture Notes in Mathematics, 991:149–155, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061568/.

Losert:1980:CMS

[Los80] V. Losert. A counterexample on measurable selections and strong lifting. Lecture Notes in Mathematics, 794:153–159, 1980. CO-DEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088221/.

Losert:1982:SLC

[Los82] Viktor Losert. Strong liftings for certain classes of compact spaces. Lecture Notes in Mathematics, 945:170–179, 1982. CO-DEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096672/.

Losert:1984:SRI

[Los84] Viktor Losert. Some remarks on invariant liftings. Lecture Notes in Mathematics, 1089:95–110, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072606/.

Louveau:1983:SRW

[Lou83a] A. Louveau. Some results in the Wadge hierarchy of Borel sets. Lecture Notes in Mathematics, 1019:28–55, 1983. CODEN LN-MAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071692/.

Louveau:1983:ECM

[Lou83b] Alain Louveau. Effective construction of measures. Lecture Notes in Mathematics, 1033:227–229, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099858/.

Loud:1984:SEB

[Lou84] W. S. Loud. Some examples of bifurcation. Lecture Notes in Mathematics, 1057:202–208, 1984. CODEN LNMAA2. ISBN 3-540-12931-6 (print), 3-540-38812-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098597/.

Loy:1983:UNP

[Loy83] Richard J. Loy. The uniqueness of norm problem in Banach algebras with finite dimensional radical. Lecture Notes in Mathematics, 975:317–327, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064565/.

Lion:1981:ERG

[LP81] Gérard Lion and Patrice Perrin. Extension des représentations de groupes unipotents p-adiques. Calculs d’obstructions. (French) []. Lecture Notes in Mathematics, 880:337–356, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090415/.

Lepowsky:1984:SMT

[LP84a] James Lepowsky and Mirko Primc. Standard modules for type one affine Lie algebras. Lecture Notes in Mathematics, 1052:194– 251, 1984. CODEN LNMAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071544/.

Lewis:1984:EES

[LP84b] J. T. Lewis and J. V. Pulé. The equivalence of ensembles in statistical mechanics. Lecture Notes in Mathematics, 1095:25–35, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099119/.

Fuglede:1980:APS

[lPBF80] Mr. le Prof Bent Fuglede. Asymptotic paths for subharmonic functions and polygonal connectedness of fine domains. Lecture Notes in Mathematics, 814:97–116, 1980. CODEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094151/.

Levin:1980:GPL

[LR80a] Frank Levin and Gerhard Rosenberger. Generalized power laws. Lecture Notes in Mathematics, 806:246–248, 1980. CODEN LN-MAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091273/.

Louie:1980:SSI

[LR80b] Donald Louie and Balram S. Rajput. Support and seminorm integrability theorems for r-semistable probability measures on LCTVS. Lecture Notes in Mathematics, 828:179–195, 1980. CO-DEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097403/.

Laha:1981:DPM

[LR81] R. G. Laha and V. K. Rohatgi. Decomposition of probability measures on locally compact abelian groups. Lecture Notes in Mathematics, 861:83–92, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097317/.

Laine:1983:VDT

[LR83a] Ilpo Laine and Seppo Rickman, editors. Value Distribution Theory: Proceedings of the Nordic Summer School in Mathematics Held at Joensuu, Finland June 1–12, 1981, volume 981 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12003-3 (print), 3-540-39480-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 981.

URL http://link.springer.com/book/10.1007/BFb0066380;

 http://www.springerlink.com/content/978-3-540-39480-
8.

Lazarsfeld:1983:LGC

[LR83b] Robert Lazarsfeld and Prabhakar Rao. Linkage of general curves of large degree. Lecture Notes in Mathematics, 997:267–289, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061648/.

Lentini:1983:PNH

[LR83c] M. Lentini and A. Reinoza. Piecewise nonlinear homotopies. Lecture Notes in Mathematics, 1005:162–169, 1983. CODEN LN-MAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112533/.

Loday-Richaud:1983:TIE

[LR83d] Michèle Loday-Richaud. Théorèmes d’indices dans les espaces de type Gevrey généralisé. (French) []. Lecture Notes in Mathematics, 1015:172–241, 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071353/.

Lelong:1980:SPL

[LS80] Pierre Lelong and Henri Skoda, editors. Séminaire Pierre Lelong– Henri Skoda (Analyse) Années 1978/79, volume 822 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 822.

URL http://link.springer.com/book/10.1007/BFb0097744;
 http://www.springerlink.com/content/978-3-540-38320-
8.

Lambek:1981:IPF

[LS81a] J. Lambek and P. J. Scott. Independence of premisses and the free topos. Lecture Notes in Mathematics, 873:191–207, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090734/.

Lovasz:1981:SCA

[LS81b] L. Lovász and A. Schriver. Some combinatorial applications of the new linear programming algorithm. Lecture Notes in Mathematics, 885:33–41, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092253/.

Lelong:1982:SPL

[LS82a] Pierre Lelong and Henri Skoda, editors. Séminaire Pierre Lelong–Henri Skoda (Analyse) Années 1980/81: Colloque de Wimereux, Mai 1981 “Les fonctions plurisousharmoniques en dimension finie ou infinie”, organisé en l’honneur de Pierre Lelong, volume 919 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 919. URL http:/

/link.springer.com/book/10.1007/BFb0097040;   http://
www.springerlink.com/content/978-3-540-39147-0.

Lutz:1982:ANA

[LS82b] Robert Lutz and Tewfik Sari. Applications of nonstandard analysis to boundary value problems in singular perturbation theory. Lecture Notes in Mathematics, 942:113–135, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094743/.

Langlands:1983:PVA

[LS83a] R. Langlands and D. Shelstad. On principal values on p-adic manifolds. Lecture Notes in Mathematics, 1041:250–279, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073150/.

Lascoux:1983:SFM

[LS83b] Alain Lascoux and Marcel-Paul Schützenberger. Symmetry and flag manifolds. Lecture Notes in Mathematics, 996:118–144, 1983. CODEN LNMAA2. ISBN 3-540-12319-9 (print), 3-540-40043-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063238/.

Lerman:1981:LYU

[LSS81] Manuel Lerman, James H. Schmerl, and Robert I. Soare, editors. Logic Year 1979–80: The University of Connecticut, USA, volume 859 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 859. URL http:/

/link.springer.com/book/10.1007/BFb0090934;   http://
www.springerlink.com/content/978-3-540-38673-5.

Looijenga:1981:GSU

[LST81] E. Looijenga, D. Siersma, and F. Takens, editors. Geometry Symposium Utrecht 1980: Proceedings of a Symposium Held at the University of Utrecht, The Netherlands, August 27–29, 1980, volume 894 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 894. URL http:/

/link.springer.com/book/10.1007/BFb0096221;   http://
www.springerlink.com/content/978-3-540-38641-4.

Lashof:1984:STF

[LT84a] Richard Lashof and Laurence Taylor. Smoothing theory and Freedman’s work on four manifolds. Lecture Notes in Mathematics, 1051:271–292, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075572/.

Laurent-Thiebaut:1984:TBV

[LT84b] Christine Laurent-Thiébaut. Théorème de Bochner sur une variété de Stein. (French) []. Lecture Notes in Mathematics, 1094: 151–161, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099159/.

Levenberg:1984:CC

[LT84c] Norman Levenberg and B. A. Taylor. Comparison of capacities in C^{n}. Lecture Notes in Mathematics, 1094:162–172, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099160/.

Lam:1982:FSS

[LTS82] C. Lam, L. Thiel, and S. Swiercz. A feasibility study of a search for ovals in a projective plane of order 10. Lecture Notes in Mathematics, 952:349–352, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061988/.

Luedeman:1981:CCP

[Lue81] John K. Luedeman. Completely cyclic projective S-systems. Lecture Notes in Mathematics, 855:49–54, 1981. CODEN LN-MAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089808/.

Luedeman:1983:TTS

[Lue83] John K. Luedeman. Torsion theories and semigroups of quotients. Lecture Notes in Mathematics, 998:350–373, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062041/.

Lukes:1980:SSP

[Luk80a] Jaroslav Lukeš. On the set of semiregular points. Lecture Notes in Mathematics, 787:212–218, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086337/.

Lukierski:1980:QSM

[Luk80b] Jerzy Lukierski. Quaternionic and supersymmetric σ- models. Lecture Notes in Mathematics, 836:221–245, 1980. CODEN LN-MAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089738/.

Lukacs:1981:CCT

[Luk81] Eugene Lukacs. Construction of characterization theorems. Lecture Notes in Mathematics, 861:121–127, 1981. CODEN LN-MAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097319/.

Lumer:1980:AOL

[Lum80a] G. Lumer. Approximation d’opérateurs locaux et de solutions d’équations d’évolution. (French) []. Lecture Notes in Mathematics, 814:166–185, 1980. CODEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094154/.

Lumer:1980:CLO

[Lum80b] G. Lumer. Connecting of local operators and evolution equations on networks. Lecture Notes in Mathematics, 787:219–234, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086338/.

Lumer:1984:EDG

[Lum84] G. Lumer. Équations de diffusion générales sur des réseaux infinis. (French) []. Lecture Notes in Mathematics, 1061:230–243, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099027/.

Luneburg:1981:EBS

[Lün81] Heinz Lüneburg. Ein Einfacher Beweis für den Satz von Zsigmondy über primitive Primeiler von A^{N} − 1. (German) []. Lecture Notes in Mathematics, 893:219–222, 1981. CODEN LN-MAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091022/.

Luneburg:1982:DN

[Lün82] Heinz Lüneburg. On Dedekind numbers. Lecture Notes in Mathematics, 969:251–257, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062998/.

Lusztig:1983:LCW

[Lus83] G. Lusztig. Left cells in Weyl groups. Lecture Notes in Mathematics, 1024:99–111, 1983. CODEN LNMAA2. ISBN 3-540-12725-9 (print), 3-540-38709-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071433/.

Lutz:1982:SCO

[Lut82] Dieter Lutz. Strongly continuous operator cosine functions. Lecture Notes in Mathematics, 948:73–97, 1982. CODEN LN-MAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069842/.

Lagomasino:1984:SRA

[LV84] G. López Lagomasino and V. V. Vavilov. Survey on recent advances in inverse problems of Padé approximation theory. Lecture Notes in Mathematics, 1105:11–26, 1984. CODEN LN-MAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072397/.

Libgober:1981:AGP

[LW81] Anatoly Libgober and Philip Wagreich, editors. Algebraic Geometry: Proceedings of the Midwest Algebraic Geometry Conference, University of Illinois at Chicago Circle, May 2–3, 1980, volume 862 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 862. URL http:/

/link.springer.com/book/10.1007/BFb0090887;   http://
www.springerlink.com/content/978-3-540-38720-6.

Lemaire:1982:IAC

[LW82] Luc Lemaire and John C. Wood. An introduction to analysis on complex manifolds. Lecture Notes in Mathematics, 950:12– 104, 1982. CODEN LNMAA2. ISBN 3-540-11596-X (print), 3-540-39366-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061876/.

LeGall:1983:EST

[LY83] J.-F. Le Gall and M. Yor. Sur l’équation stochastique de Tsirelson. (French) []. Lecture Notes in Mathematics, 986:81–88, 1983. CO-DEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068302/.

Lyndon:1984:SUC

[Lyn84] R. C. Lyndon. Some uses of coset graphs. Lecture Notes in Mathematics, 1098:69–74, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099663/.

Lyubeznik:1984:LCM

[Lyu84] Gennady Lyubeznik. On the local cohomology modules for ideals generated by monomials in an R-sequence. Lecture Notes in Mathematics, 1092:214–220, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099364/.

Lenz:1982:AOS

[LZ82] Hanfried Lenz and Herbert Zeitler. Arcs and ovals in Steiner triple systems. Lecture Notes in Mathematics, 969:229–250, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062997/.

Macintyre:1980:RQA

[Mac80] Angus Macintyre. Ramsey quantifiers in arithmetic. Lecture Notes in Mathematics, 834:186–210, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090167/.

Macdonald:1981:ALA

[Mac81a] I. G. Macdonald. Affine Lie algebras and modular forms. Lecture Notes in Mathematics, 901:258–276, 1981. CODEN LN-MAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097202/.

Macdonald:1981:SCR

[Mac81b] I. G. Macdonald. Some conjectures for root systems and finite Coxeter groups. Lecture Notes in Mathematics, 867:90–97, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090381/.

Machado:1981:FAH

[Mac81c] Silvio Machado, editor. Functional Analysis, Holomorphy, and Approximation Theory: Proceedings of the Seminário de Análise Funcional, Holomorfia e Teoria da Aproximação, Universidade Federal do Rio de Janeiro, Brazil, August 7–11, 1978, volume 843 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 843. URL http:/

/link.springer.com/book/10.1007/BFb0089265;   http://
www.springerlink.com/content/978-3-540-38529-5.

Macintyre:1981:CTF

[Mac81d] Angus Macintyre. The complexity of types in field theory. Lecture Notes in Mathematics, 859:143–156, 1981. CODEN LN-MAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090946/.

Macintyre:1981:LE

[Mac81e] Angus Macintyre. The laws of exponentiation. Lecture Notes in Mathematics, 890:185–197, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095664/.

Macki:1983:RGZ

[Mac83] Jack W. Macki. Regular growth and zero-tending solutions. Lecture Notes in Mathematics, 1032:358–374, 1983. CODEN LN-MAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076807/.

Maddox:1980:CMC

[Mad80a] Ivor J. Maddox. Characterization of matrix classes. Lecture Notes in Mathematics, 786:36–64, 1980. CODEN LNMAA2. ISBN 3-540-09764-3 (print), 3-540-38946-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088200/.

Maddox:1980:CT

[Mad80b] Ivor J. Maddox. Consistency theorems. Lecture Notes in Mathematics, 786:84–105, 1980. CODEN LNMAA2. ISBN 3-540-09764-3 (print), 3-540-38946-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088202/.

Maddox:1980:GKT

[Mad80c] Ivor J. Maddox. Generalized Köthe–Toeplitz duals. Lecture Notes in Mathematics, 786:19–35, 1980. CODEN LNMAA2. ISBN 3-540-09764-3 (print), 3-540-38946-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088199/.

Maddox:1980:IMO

[Mad80d] Ivor J. Maddox. Infinite Matrices of Operators, volume 786 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09764-3 (print), 3-540-38946-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 122 pp. LCCN QA3 .L28 no. 786. URL http:/

/link.springer.com/book/10.1007/BFb0088196;   http://
www.springerlink.com/content/978-3-540-38946-0.

Maddox:1980:I

[Mad80e] Ivor J. Maddox. Introduction. Lecture Notes in Mathematics, 786:1–2, 1980. CODEN LNMAA2. ISBN 3-540-09764-3 (print), 3-540-38946-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088197.

Maddox:1980:NT

[Mad80f] Ivor J. Maddox. Notation and terminology. Lecture Notes in Mathematics, 786:3–18, 1980. CODEN LNMAA2. ISBN 3-540-09764-3 (print), 3-540-38946-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088198/.

Maddox:1980:ONM

[Mad80g] Ivor J. Maddox. Operator Nörlund means. Lecture Notes in Mathematics, 786:106–116, 1980. CODEN LNMAA2. ISBN 3-540-09764-3 (print), 3-540-38946-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088203/.

Maddox:1980:TT

[Mad80h] Ivor J. Maddox. Tauberian theorems. Lecture Notes in Mathematics, 786:65–83, 1980. CODEN LNMAA2. ISBN 3-540-09764-3 (print), 3-540-38946-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088201/.

Maeda:1980:DIH

[Mae80a] Fumi-Yuki Maeda. Dirichlet Integrals on Harmonic Spaces, volume 803 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 180 pp. LCCN QA3 .L28 no. 803. URL http:/

/link.springer.com/book/10.1007/BFb0093376;   http://
www.springerlink.com/content/978-3-540-39301-6.

Maeda:1980:EFH

[Mae80b] Fumi-Yuki Maeda. Energy-finite harmonic functions and Green’s formula. Lecture Notes in Mathematics, 803:86–100, 1980. CO-DEN LNMAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093381/.

Maeda:1980:FC

[Mae80c] Fumi-Yuki Maeda. Functional completion. Lecture Notes in Mathematics, 803:134–154, 1980. CODEN LNMAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093383/.

Maeda:1980:GM

[Mae80d] Fumi-Yuki Maeda. Gradient measures. Lecture Notes in Mathematics, 803:33–55, 1980. CODEN LNMAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093379/.

Maeda:1980:HS

[Mae80e] Fumi-Yuki Maeda. Harmonic spaces. Lecture Notes in Mathematics, 803:1–15, 1980. CODEN LNMAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093377/.

Maeda:1980:RB

[Mae80f] Fumi-Yuki Maeda. Royden boundary. Lecture Notes in Mathematics, 803:154–169, 1980. CODEN LNMAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093384/.

Maeda:1980:SAH

[Mae80g] Fumi-Yuki Maeda. Self-adjoint harmonic spaces and Green potentials. Lecture Notes in Mathematics, 803:56–85, 1980. CODEN LNMAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093380/.

Maeda:1980:SDF

[Mae80h] Fumi-Yuki Maeda. Spaces of Dirichlet-finite and energy-finite harmonic functions. Lecture Notes in Mathematics, 803:101– 134, 1980. CODEN LNMAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093382/.

Maeda:1980:SFP

[Mae80i] Fumi-Yuki Maeda. Superharmonic functions and potentials. Lecture Notes in Mathematics, 803:16–32, 1980. CODEN LN-MAA2. ISBN 3-540-09995-6 (print), 3-540-39301-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093378/.

Maehara:1981:SFF

[Mae81] Shôji Maehara. Semi-formal finitist proof of the transfinite induction in an initial segment of Cantor’s second number class. Lecture Notes in Mathematics, 891:67–79, 1981. CODEN LN-MAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090980/.

Magnus:1980:TEB

[Mag80] Robert Magnus. Topological equivalence in bifurcation theory. Lecture Notes in Mathematics, 799:263–276, 1980. CODEN LN-MAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089318/.

Magurn:1981:UUW

[Mag81] Bruce A. Magurn. Uses of units in Whitehead groups. Lecture Notes in Mathematics, 882:261–268, 1981. CODEN LN-MAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092499/.

Magnus:1982:STP

[Mag82] Arne Magnus. On the structure of the two-point Padé table. Lecture Notes in Mathematics, 932:176–193, 1982. CODEN LN-MAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093313/.

Magiera:1983:EPR

[Mag83] Józef Magiera. Exterior partitions of a rectangle into rectangles and their graphs. Lecture Notes in Mathematics, 1018:125– 130, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071622/.

Magnus:1984:RAJ

[Mag84] Alphonse Magnus. Riccati acceleration of Jacobi continued fractions and Laguerre–Hahn orthogonal polynomials. Lecture Notes in Mathematics, 1071:213–230, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099620/.

Maharam:1984:PRM

[Mah84] Dorothy Maharam. On the planar representation of a measurable subfield. Lecture Notes in Mathematics, 1089:47–57, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072599/.

Maisonneuve:1981:LDT

[Mai81] B. Maisonneuve. On Lévy’s downcrossing theorem and various extensions. Lecture Notes in Mathematics, 850:191–205, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088368/.

Major:1981:HPC

[Maj81a] Péter Major. History of the problems. Comments. Lecture Notes in Mathematics, 849:105–121, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094045/.

Major:1981:MWIb

[Maj81b] Péter Major. Multiple Wiener–Itô Integrals: With Applications to Limit Theorems, volume 849 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 127 pp. LCCN QA3 .L28 no. 849.

URL http://link.springer.com/book/10.1007/BFb0094036;
 http://www.springerlink.com/content/978-3-540-38557-
8.

Major:1981:MWIa

[Maj81c] Péter Major. Multiple Wiener-Itô integrals. Lecture Notes in Mathematics, 849:22–37, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094040/.

Major:1981:NCL

[Maj81d] Péter Major. Non-central limit theorems. Lecture Notes in Mathematics, 849:80–105, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094044/.

Major:1981:LP

[Maj81e] Péter Major. On a limit problem. Lecture Notes in Mathematics, 849:1–6, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094037/.

Major:1981:OWI

[Maj81f] Péter Major. On the original Wiener–Itô integral. Lecture Notes in Mathematics, 849:73–79, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094043/.

Major:1981:PIF

[Maj81g] Péter Major. The proof of itô’s formula. The diagram formula and some of its consequences. Lecture Notes in Mathematics, 849: 37–55, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094041/.

Major:1981:RSM

[Maj81h] Péter Major. Random spectral measures. Lecture Notes in Mathematics, 849:13–22, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094039/.

Major:1981:SFC

[Maj81i] Péter Major. Subordinated fields, construction of self-similar fields. Lecture Notes in Mathematics, 849:55–73, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094042/.

Major:1981:WP

[Maj81j] Péter Major. Wick polynomials. Lecture Notes in Mathematics, 849:6–13, 1981. CODEN LNMAA2. ISBN 3-540-10575-1 (print), 3-540-38557-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094038/.

Majcher:1983:SRR

[Maj83] Zofia Majcher. On sequences representable by Γ^{−}-regular graphs. Lecture Notes in Mathematics, 1018:131–138, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071623/.

Majda:1984:SSE

[Maj84a] Andrew Majda. Smooth solutions for the equations of compressible and incompressible fluid flow. Lecture Notes in Mathematics, 1047:75–126, 1984. CODEN LNMAA2. ISBN 3-540-12893-X (print), 3-540-38773-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072326/;   http://link.springer.com/chapter/
10.1007/BFb0072327/;   http://link.springer.com/content/
pdf/bfm:978-3-540-38773-2/1.pdf.

Majima:1984:PLD

[Maj84b] Hideyuki Majima. ▽-Poincaré’s lemma and ▽-Derham cohomology theorem. Lecture Notes in Mathematics, 1075:134–151, 1984. CODEN LNMAA2. ISBN 3-540-13375-5 (print), 3-540-38931-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071555/.

Majima:1984:AAI

[Maj84c] Hideyuki Majima. Asymptotic Analysis for Integrable Connections with Irregular Singular Points, volume 1075 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13375-5 (print), 3-540-38931-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 159 pp. LCCN QA3 .L28 no. 1075. URL http:/

/link.springer.com/book/10.1007/BFb0071550;   http://
www.springerlink.com/content/978-3-540-38931-6.

Majima:1984:ADV

[Maj84d] Hideyuki Majima. Asymptotic developability and vanishing theorems in asymptotic analysis. Lecture Notes in Mathematics, 1075: 13–48, 1984. CODEN LNMAA2. ISBN 3-540-13375-5 (print), 3-540-38931-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071552/.

Majima:1984:ETA

[Maj84e] Hideyuki Majima. Existence theorems of asymptotic solutions and splitting lemmas. Lecture Notes in Mathematics, 1075:49– 110, 1984. CODEN LNMAA2. ISBN 3-540-13375-5 (print), 3-540-38931-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071553/.

Majima:1984:GI

[Maj84f] Hideyuki Majima. General introduction. Lecture Notes in Mathematics, 1075:1–12, 1984. CODEN LNMAA2. ISBN 3-540-13375-5 (print), 3-540-38931-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071551/.

Majima:1984:SPR

[Maj84g] Hideyuki Majima. Stokes phenomena and Riemann–Hilbert– Birkhoff problem. Lecture Notes in Mathematics, 1075:111– 133, 1984. CODEN LNMAA2. ISBN 3-540-13375-5 (print), 3-540-38931-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071554/.

Makagon:1980:REB

[Mak80a] A. Makagon. Remark on the extrapolation of Banach space valued stationary processes. Lecture Notes in Mathematics, 828: 196–207, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097404/.

Maki:1980:BPFa

[Mäk80b] Sirpa Mäki. Bergström’s product formula. Lecture Notes in Mathematics, 797:32–36, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088943/.

Maki:1980:BPFb

[Mäk80c] Sirpa Mäki. Bergström’s product formula in the case of real cyclic sextic fields. Lecture Notes in Mathematics, 797:37–51, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088944/.

Maki:1980:CN

[Mäk80d] Sirpa Mäki. The class number of K_{6}. Lecture Notes in Mathematics, 797:58–59, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088946.

Maki:1980:CP

[Mäk80e] Sirpa Mäki. The computer program. Lecture Notes in Mathematics, 797:62–67, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088948/.

Maki:1980:DUR

[Mäk80f] Sirpa Mäki. The Determination of Units in Real Cyclic Sextic Fields, volume 797 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 198 pp. LCCN QA3 .L28 no. 797. URL http:/

/link.springer.com/book/10.1007/BFb0088938;   http://
www.springerlink.com/content/978-3-540-39242-2.

Maki:1980:FC

[Mäk80g] Sirpa Mäki. Formulas for computing ξ_{A}. Lecture Notes in Mathematics, 797:52–57, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088945/.

Maki:1980:FS

[Mäk80h] Sirpa Mäki. The function M and the structure of U_{R}. Lecture Notes in Mathematics, 797:23–31, 1980. CODEN LN-MAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088942/.

Maki:1980:I

[Mäk80i] Sirpa Mäki. Introduction. Lecture Notes in Mathematics, 797: 1–5, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088939/.

Maki:1980:NR

[Mäk80j] Sirpa Mäki. Numerical results. Lecture Notes in Mathematics, 797:68–193, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088949/.

Maki:1980:RCC

[Mäk80k] Sirpa Mäki. Real cyclic cubic fields. Lecture Notes in Mathematics, 797:6–11, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088940/.

Maki:1980:RCS

[Mäk80l] Sirpa Mäki. Real cyclic sextic fields. Lecture Notes in Mathematics, 797:12–22, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088941/.

Maki:1980:SR

[Mäk80m] Sirpa Mäki. The signature rank of U_{6}. Lecture Notes in Mathematics, 797:60–61, 1980. CODEN LNMAA2. ISBN 3-540-09984-0 (print), 3-540-39242-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088947.

Makkai:1981:TT

[Mak81] M. Makkai. The topos of types. Lecture Notes in Mathematics, 859:157–201, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090947/.

Malliavin:1980:GTI

[Mal80a] Marie-Paule Malliavin. Grade et théorème d’intersection en algèbre non commutative. (French) []. Lecture Notes in Mathematics, 825:76–87, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089123/.

Malliavin:1980:SAP

[Mal80b] Marie-Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil et Marie-Paule Malliavin: Proceedings, Paris 1979 (32ème Année). (French) [], volume 795 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 795. URL http:

//link.springer.com/book/10.1007/BFb0090110;   http:/
/www.springerlink.com/content/978-3-540-39230-9.

Malliavin:1980:PLG

[Mal80c] Paul Malliavin. C^{∞} parametrix on Lie groups and two steps factorization on convolution algebras. Lecture Notes in Mathematics, 781:142–156, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097652/.

Malliavin:1981:SAP

[Mal81a] Marie-Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil et Marie-Paule Malliavin: Proceedings, Paris 1980 (33ème Année). (French) [], volume 867 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 867. URL http:

//link.springer.com/book/10.1007/BFb0090377;   http:/
/www.springerlink.com/content/978-3-540-38737-4.

Maltsiniotis:1981:TBN

[Mal81b] Georges Maltsiniotis. Le théorème de Brill–Noether. (French) []. Lecture Notes in Mathematics, 901:157–175, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097196/.

Malcolmson:1982:CUM

[Mal82a] Peter Malcolmson. Construction of universal matrix localizations. Lecture Notes in Mathematics, 951:117–131, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067328/.

Malliavin:1982:GTI

[Mal82b] Marie-Paule Malliavin. Grade et théorème d’intersection en algèbre non commutative. (French) []. Lecture Notes in Mathematics, 924:167–172, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092931/.

Malliavin:1982:SAP

[Mal82c] Marie-Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil et Marie-Paule Malliavin: Proceedings, Paris 1981 (34ème Année). (French) [], volume 924 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 924. URL http:

//link.springer.com/book/10.1007/BFb0092923;   http:/
/www.springerlink.com/content/978-3-540-39188-3.

Malliavin:1982:UPA

[Mal82d] Marie-Paule Malliavin. Ultra-produits d’algèbres de Lie. (French) []. Lecture Notes in Mathematics, 924:157–166, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092930/.

Malliavin:1983:SAP

[Mal83] Marie-Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil et Marie-Paule Malliavin: Proceedings, Paris 1982 (35ème Année). (French) [], volume 1029 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1029. URL http:/

/link.springer.com/book/10.1007/BFb0098924;   http://
www.springerlink.com/content/978-3-540-38686-5.

Maligranda:1984:FSS

[Mal84] Lech Maligranda. The K-functional for symmetric spaces. Lecture Notes in Mathematics, 1070:169–182, 1984. CODEN LN-MAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099100/.

Manders:1980:CCD

[Man80] Kenneth L. Manders. Computational complexity of decision problems in elementary number theory. Lecture Notes in Mathematics, 834:211–227, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090168/.

Mandrekar:1981:DAP

[Man81a] V. Mandrekar. Domain of attraction problem on Banach spaces: a survey. Lecture Notes in Mathematics, 860:285–290, 1981. CO-DEN LNMAA2. ISBN 3-540-10822-X (print), 3-540-38710-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090623/.

Mane:1981:DCI

[Mañ81b] Ricardo Mañé. On the dimension of the compact invariant sets of certain non-linear maps. Lecture Notes in Mathematics, 898: 230–242, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091916/.

Manning:1981:MTE

[Man81c] Anthony Manning. More topological entropy for geodesic flows. Lecture Notes in Mathematics, 898:243–249, 1981. CODEN LN-MAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091917/.

Manin:1982:GFC

[Man82] Yu. I. Manin. Gauge fields and cohomology of analytic sheaves. Lecture Notes in Mathematics, 970:43–52, 1982. CODEN LN-MAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066023/.

Mandrekar:1983:SRV

[Man83] V. Mandrekar. Some remarks on various definitions of Feynman integral. Lecture Notes in Mathematics, 990:170–177, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064270/.

Manders:1984:IMT

[Man84] Kenneth L. Manders. Interpretations and the model theory of the classical geometries. Lecture Notes in Mathematics, 1103: 297–330, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099392/.

Marmaridis:1980:RF

[Mar80a] Nikolaos Marmaridis. Reflection functors. Lecture Notes in Mathematics, 832:382–395, 1980. CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088475/.

Martini:1980:GAD

[Mar80b] Rodolfo Martini, editor. Geometrical Approaches to Differential Equations: Proceedings of the Fourth Scheveningen Conference on Differential Equations, The Netherlands August 26–31, 1979, volume 810 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 810. URL http:/

/link.springer.com/book/10.1007/BFb0089970;   http://
www.springerlink.com/content/978-3-540-38166-2.

Mardesic:1981:ILR

[Mar81a] Sibe Mardešić. Inverse limits and resolutions. Lecture Notes in Mathematics, 870:239–252, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089718/.

Margenstern:1981:TMR

[Mar81b] M. Margenštern. Le théorème de Matiyassévitch et résultats connexes. (French) []. Lecture Notes in Mathematics, 890:198– 241, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095665/.

Markowsky:1981:MGS

[Mar81c] George Markowsky. A motivation and generalization of Scott’s notion of a continuous lattice. Lecture Notes in Mathematics, 871: 298–307, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089913/.

Markowsky:1981:PCC

[Mar81d] George Markowsky. Propaedeutic to chain-complete posets with basis. Lecture Notes in Mathematics, 871:308–314, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089914/.

Markus:1981:CMT

[Mar81e] L. Markus. Controllability of multi-trajectories on Lie groups. Lecture Notes in Mathematics, 898:250–265, 1981. CODEN LN-MAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091918/.

Marmaridis:1981:SRR

[Mar81f] Nikolaos Marmaridis. Some remarks on reflection functors. Lecture Notes in Mathematics, 903:211–221, 1981. CODEN LN-MAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092994/.

Martelli:1981:SFO

[Mar81g] Mario Martelli. Semi-Fredholm operators and hyperbolic problems. Lecture Notes in Mathematics, 886:249–264, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092187/.

Martin:1981:MID

[Mar81h] Donald A. Martin. Π_{2}^{1} monotone inductive definitions. Lecture Notes in Mathematics, 839:215–233, 1981. CODEN LN-MAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090243/.

Martinet:1981:NCV

[Mar81i] Jean Martinet. Normalisation des champs de vecteurs holomorphes (d’après A.-D. Brjuno [2]). (French) []. Lecture Notes in Mathematics, 901:55–70, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097189/.

Marsden:1982:SSR

[Mar82a] Jerrold E. Marsden. Spaces of solutions of relativistic field theories with constraints. Lecture Notes in Mathematics, 905:29–43, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092425/.

Maruyama:1982:ETT

[Mar82b] Masaki Maruyama. Elementary transformations in the theory of algebraic vector bundles. Lecture Notes in Mathematics, 961: 241–266, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071286/.

Marcellini:1983:RBE

[Mar83a] Paolo Marcellini. A relation between existence of minima for nonconvex integrals and uniqueness for non-strictly-convex integrals of the calculus of variations. Lecture Notes in Mathematics, 979: 216–231, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066256/.

Maroscia:1983:SPR

[Mar83b] Paolo Maroscia. Some problems and results on finite sets of points in P^{n}. Lecture Notes in Mathematics, 997:290–314, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061649/.

Martens:1983:RFT

[Mar83c] Henrik H. Martens. Remarks on de Franchis’ theorem. Lecture Notes in Mathematics, 1013:160–163, 1983. CODEN LN-MAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066526/.

Martin:1983:LCO

[Mar83d] Donald A. Martin. The largest countable this, that, and the other. Lecture Notes in Mathematics, 1019:97–106, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071698/.

Martin:1983:RGQ

[Mar83e] Donald A. Martin. The real game quantifier propagates scales. Lecture Notes in Mathematics, 1019:157–171, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071700/.

Martio:1983:QMH

[Mar83f] O. Martio. Quasiconformal mappings and F -harmonic measure. Lecture Notes in Mathematics, 1013:164–170, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066527/.

Maruyama:1983:SCJ

[Mar83g] Masaki Maruyama. Singularities of the curve of jumping lines of a vector bundle of rank 2 on P^{2}. Lecture Notes in Mathematics, 1016:370–411, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099972/.

Marchioro:1984:VTE

[Mar84a] C. Marchioro. Vortex theory and Euler and Navier–Stokes evolution in two dimensions. Lecture Notes in Mathematics, 1047: 167–174, 1984. CODEN LNMAA2. ISBN 3-540-12893-X (print), 3-540-38773-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072330/.

Maritz:1984:SRM

[Mar84b] P. Maritz. Some remarks on measurable and semi-continuous multifunctions. Lecture Notes in Mathematics, 1089:111–119, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072607/.

Masuda:1980:RSN

[Mas80] Kyûya Masuda. On the regularity of solutions of the nonstationary Navier–Stokes equations. Lecture Notes in Mathematics, 771:360–370, 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086917/.

Massa:1981:MAD

[Mas81] Sylvio Massa. Multi-applications du type de Kannan. (French) []. Lecture Notes in Mathematics, 886:265–269, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092188/.

Masani:1982:ORF

[Mas82a] P. Masani. The outer regularization of finitely-additive measures over normal topological spaces. Lecture Notes in Mathematics, 945:116–144, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096667/.

Mastrangelo:1982:PDM

[Mas82b] M. Mastrangelo. Processus de diffusion multigroupe frontières fixes et variables. (French) []. Lecture Notes in Mathematics, 906: 225–251, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093269/.

Mastrangelo:1982:RSE

[Mas82c] M. Mastrangelo. Résolution des systèmes d’équations de diffusion par les intégrales stochastiques d’Itô. (French) []. Lecture Notes in Mathematics, 906:252–260, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093270/.

Maskit:1983:PWG

[Mas83a] Bernard Maskit. Panelled web groups. Lecture Notes in Mathematics, 971:79–108, 1983. CODEN LNMAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067075/.

Massaoud:1983:CIA

[Mas83b] Hédi Ben Massaoud. Courants intermédiaires associés à un Courant positif fermé. (French) []. Lecture Notes in Mathematics, 1028:41–68, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071677/.

Mascaro:1984:SR

[Mas84a] F. Mascaró. Some results on diff^{ω}(R^{n}). Lecture Notes in Mathematics, 1045:125–133, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072172/.

Mashimo:1984:SSM

[Mas84b] Katsuya Mashimo. On some stable minimal cones in R^{7}. Lecture Notes in Mathematics, 1090:107–115, 1984. CODEN LN-MAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101571/.

Mason:1984:SPA

[Mas84c] J. C. Mason. Some properties and applications of Chebyshev polynomial and rational approximation. Lecture Notes in Mathematics, 1105:27–48, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072398/.

Mason:1984:EFF

[Mas84d] R. C. Mason. Equations over function fields. Lecture Notes in Mathematics, 1068:149–157, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099449/.

Matsuda:1980:FOA

[Mat80] Michihiko Matsuda. First Order Algebraic Differential Equations: A Differential Algebraic Approach, volume 804 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 111 pp. LCCN

QA3 .L28 no. 804. URL http://link.springer.com/book/10.
1007/BFb0091495.

Matchett:1982:ESL

[Mat82] Andrew Matchett. Exact sequences for locally free class groups. Lecture Notes in Mathematics, 967:280–290, 1982. CODEN LN-MAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061906/.

Mauldin:1980:SST

[Mau80] R. Daniel Mauldin. Some selection theorems and problems. Lecture Notes in Mathematics, 794:160–165, 1980. CODEN LN-MAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088222/.

Maury:1982:TIG

[Mau82] Guy Maury. Un théorème de l’ideal à gauche principal dans certains anneaux. (French) []. Lecture Notes in Mathematics, 924: 184–196, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092933/.

Mayne:1980:UEP

[May80] David Q. Mayne. On the use of exact penalty functions to determine step length in optimization algorithms. Lecture Notes in Mathematics, 773:98–109, 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094166/.

Meis:1982:NCP

[MB82a] Theodor Meis and Wolfgang Baaske. Numerical computation of periodic solutions of a nonlinear wave equation. Lecture Notes in Mathematics, 953:159–171, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069381/.

Meyer:1982:VSD

[MB82b] P.-A. Meyer and J. M. Bismut. Variation des solutions d’une E.D.S.. (French) [Variation of the solutions of stochastic differential equation]. Lecture Notes in Mathematics, 921:151–164, 1982. CODEN LNMAA2. ISBN 3-540-11486-6 (print), 3-540-39167-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092649/.

McGill:1981:DPR

[Mc 81] P. Mc Gill. A direct proof of the Ray–Knight theorem. Lecture Notes in Mathematics, 850:206–209, 1981. CODEN LN-MAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088369/.

McAvaney:1980:CTP

[McA80] K. L. McAvaney. A conjecture on two-point deleted subgraphs of cartesian products. Lecture Notes in Mathematics, 829:172– 185, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088911/.

McAvaney:1981:CMV

[McA81] Kevin L. McAvaney, editor. Combinatorial Mathematics VIII: Proceedings of the Eighth Australian Conference on Combinatorial Mathematics Held at Deakin University, Geelong, Australia, August 25–29, 1980, volume 884 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 884. URL http:

//link.springer.com/book/10.1007/BFb0091801;   http:/
/www.springerlink.com/content/978-3-540-38792-3.

McAdam:1983:AIBb

[McA83a] Stephen McAdam. A^{∗}(I) − B^{∗}(I). Lecture Notes in Mathematics, 1023:8–11, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071577/.

McAdam:1983:AIBa

[McA83b] Stephen McAdam. A^{∗}(I) and B^{∗}(I). Lecture Notes in Mathematics, 1023:1–7, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071576/.

McAdam:1983:AI

[McA83c] Stephen McAdam. Ā^{∗}(I). Lecture Notes in Mathematics, 1023: 12–25, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071578/.

McAdam:1983:AG

[McA83d] Stephen McAdam. Asymptotic grade. Lecture Notes in Mathematics, 1023:55–60, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071582/.

McAdam:1983:APD

[McA83e] Stephen McAdam. Asymptotic Prime Divisors, volume 1023 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 118 pp. LCCN QA3 .L28 no. 1023. URL http:/

/link.springer.com/book/10.1007/BFb0071575;   http://
www.springerlink.com/content/978-3-540-38704-6.

McAdam:1983:AS

[McA83f] Stephen McAdam. Asymptotic sequences. Lecture Notes in Mathematics, 1023:32–41, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071580/.

McAdam:1983:ASI

[McA83g] Stephen McAdam. Asymptotic sequences over ideals. Lecture Notes in Mathematics, 1023:42–54, 1983. CODEN LN-MAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071581/.

McAdam:1983:CB

[McA83h] Stephen McAdam. A characterization of Ā^{∗} (i). Lecture Notes in Mathematics, 1023:26–31, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071579/.

McAdam:1983:CR

[McA83i] Stephen McAdam. Conforming relations. Lecture Notes in Mathematics, 1023:68–75, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071584/.

McAdam:1983:IT

[McA83j] Stephen McAdam. Ideal transforms. Lecture Notes in Mathematics, 1023:76–88, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071585/.

McAdam:1983:M

[McA83k] Stephen McAdam. Miscellaneous. Lecture Notes in Mathematics, 1023:89–109, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071586/.

McAdam:1983:W

[McA83l] Stephen McAdam. When A^{∗} = Ā^{∗}. Lecture Notes in Mathematics, 1023:61–67, 1983. CODEN LNMAA2. ISBN 3-540-12722-4 (print), 3-540-38704-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071583/.

McCrudden:1982:LTC

[McC82] M. McCrudden. Local tightness of convolution semigroups over locally compact groups. Lecture Notes in Mathematics, 928:304– 314, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093230/.

McDill:1982:FFH

[McD82] J. M. McDill. Factorization of functors having left adjoints. Lecture Notes in Mathematics, 915:234–246, 1982. CODEN LN-MAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092883/.

McGehee:1982:CLQ

[McG82] O. Carruth McGehee. A conjecture of Littlewood, a question of Lusin, and a principle of Fourier transform behavior. Lecture Notes in Mathematics, 908:39–47, 1982. CODEN LN-MAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093279/.

McGill:1984:TCB

[McG84] Paul McGill. Time changes of Brownian motion and the conditional excursion theorem. Lecture Notes in Mathematics, 1095:

83–99, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099123/.

McKenna:1980:SDN

[McK80] Kenneth McKenna. Some Diophantine Nullstellensätze. Lecture Notes in Mathematics, 834:228–247, 1980. CODEN LN-MAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090169/.

McKenzie:1983:FFL

[McK83] Ralph McKenzie. Finite forbidden lattices. Lecture Notes in Mathematics, 1004:176–205, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063438/.

McLeod:1982:SAG

[McL82] Robin J. Y. McLeod. Some applications of geometry in numerical analysis. Lecture Notes in Mathematics, 912:190–229, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093158/.

Martini:1982:GTG

[MdJ82] Rodolfo Martini and Eduardus M. de Jager, editors. Geometric Techniques in Gauge Theories: Proceedings of the Fifth Scheveningen Conference on Differential Equations, The Netherlands August 23–28, 1981, volume 926 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 926. URL http:

//link.springer.com/book/10.1007/BFb0092655;   http:/
/www.springerlink.com/content/978-3-540-39192-0.

Mdzinarishvili:1981:AST

[Mdz81] L. D. Mdzinarishvili. Application of the shape theory in the characterization of exact homology theories and the strong shape homotopic theory. Lecture Notes in Mathematics, 870:253–262, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089719/.

Mauldin:1980:RC

[ME80] R. Daniel Mauldin and Paul Erdős. Rotations of the circle. Lecture Notes in Mathematics, 794:53–56, 1980. CODEN LN-MAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088213/.

Mehta:1983:BM

[Meh83a] Vikram B. Mehta. Back matter. Lecture Notes in Mathematics, 996:154–163, 1983. CODEN LNMAA2. ISBN 3-540-12319-9 (print), 3-540-40043-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-40043-1/1.pdf.

Mehta:1983:SRT

[Meh83b] Vikram B. Mehta. On some restriction theorems for semistable bundles. Lecture Notes in Mathematics, 996:145–153, 1983. CO-DEN LNMAA2. ISBN 3-540-12319-9 (print), 3-540-40043-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063239/.

Meister:1980:SCI

[Mei80] Prof. Dr. Erhard Meister. Some classes of integral and integrodifferential equations of convolutional type. Lecture Notes in Mathematics, 827:182–228, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091381/.

Mekler:1981:SGM

[Mek81] Alan H. Mekler. ℵ_{1}-separable groups of mixed type. Lecture Notes in Mathematics, 874:114–126, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090530/.

Mekler:1983:PFAa

[Mek83a] Alan H. Mekler. Proper forcing and Abelian groups. Lecture Notes in Mathematics, 1006:285–303, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21560-9_14/.

Mekler:1983:PFAb

[Mek83b] Alan H. Mekler. Proper forcing and abelian groups. Lecture Notes in Mathematics, 1006:285–303, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103709/.

Mela:1980:CCA

[Mél80] Jean-François Méla. Le calcul sur les caractères de l’algèbre M(G) et le problème ‘μ ∗ L^{1} fermé?’ [d’après les travaux de B. Host et F. Parreau]. (French) []. Lecture Notes in Mathematics, 770: 151–169, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096241/.

Melnikov:1982:SCL

[Mel82] V. K. Mel’nikov. Symmetries and conservation laws of dynamical systems. Lecture Notes in Mathematics, 970:146–172, 1982. CODEN LNMAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066029/.

Meldrum:1983:RSE

[Mel83] John David Philip Meldrum. Regular semigroups of endomorphisms of groups. Lecture Notes in Mathematics, 998:374–384, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062042/.

Mennicke:1980:BGP

[Men80a] Jens L. Mennicke, editor. Burnside Groups: Proceedings of a Workshop Held at the University of Bielefeld, Germany June– July 1977, volume 806 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 806. URL http:/

/link.springer.com/book/10.1007/BFb0091265;   http://
www.springerlink.com/content/978-3-540-38120-4.

Menzala:1980:HTE

[Men80b] Gustavo Perla Menzala. On a Hartree type equation: Existence of regular solutions. Lecture Notes in Mathematics, 799: 277–288, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089319/.

Menzala:1982:SRW

[Men82] Gustavo Perla Menzala. Some remarks on a wave equation with a nonlocal interaction. Lecture Notes in Mathematics, 957:229– 236, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066241/.

Mennicken:1983:EPD

[Men83] Reinhard Mennicken. Eigenvalue problems depending nonlinearly on the parameter. Lecture Notes in Mathematics, 991:156– 181, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061569/.

Mennicke:1984:DG

[Men84] J. L. Mennicke. Discontinuous groups. Lecture Notes in Mathematics, 1098:75–80, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099664/.

Merle:1980:ACS

[Mer80a] M. Merle. Les anneaux coniques sont de Cohen–Macaulay d’après A. G. Kouchnirenko. (French) []. Lecture Notes in Mathematics, 777:295–302, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085889/.

Merle:1980:PNE

[Mer80b] M. Merle. Polyèdre de Newton, éventail et désingularisation d’après A. N. Varchenko ([1]). (French) []. Lecture Notes in Mathematics, 777:289–294, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085888/.

Merlier:1981:QPI

[Mer81] Thérèse Merlier. Quelques propriétés des idéaux d’un demigroupe. (French) []. Lecture Notes in Mathematics, 855:73– 81, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089810/.

Merucci:1984:AIF

[Mer84] C. Merucci. Applications of interpolation with a function parameter to Lorentz, Sobolev and Besov spaces. Lecture Notes in Mathematics, 1070:183–201, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099101/.

Metelli:1983:TRP

[Met83] Claudia Metelli. On type-related properties of torsion-free Abelian groups. Lecture Notes in Mathematics, 1006:253–267, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
662-21560-9_11/;   http://link.springer.com/chapter/10.
1007/BFb0103706/.

Meurman:1982:E

[Meu82] Arne Meurman. An embedding of PSL(2, 13) in C0. Lecture Notes in Mathematics, 933:157–165, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093359/.

Meyer:1980:RJG

[Mey80a] P.-A. Meyer. Les résultats de Jeulin sur le grossissement des tribus. (French) []. Lecture Notes in Mathematics, 784:173– 188, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089483/.

Meyer:1980:RSF

[Mey80b] P.-A. Meyer. Sur un résultat de L. Schwartz. (French) []. Lecture Notes in Mathematics, 784:102–103, 1980. CODEN LN-MAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089472.

Meyer:1980:NIS

[Mey80c] Yves Meyer. Les nouvelles intégrales singulières de Calderón. (French) []. Lecture Notes in Mathematics, 770:57–65, 1980. CO-DEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096235/.

Meyer:1981:DGF

[Mey81a] P.-A. Meyer. A differential geometric formalism for the Itô calculus. Lecture Notes in Mathematics, 851:256–270, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088730/.

Meyer:1981:FED

[Mey81b] P.-A. Meyer. Flot d’une équation différentielle stochastique (d’après Malliavin, Bismut, Kunita). (French) []. Lecture Notes in Mathematics, 850:103–117, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088361/.

Meyer:1981:GSS

[Mey81c] P.-A. Meyer. Géométrie stochastique sans larmes. (French) []. Lecture Notes in Mathematics, 850:44–102, 1981. CODEN LN-MAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088360/.

Meyer:1981:RTL

[Mey81d] P.-A. Meyer. Retour sur la théorie de Littlewood–Paley. (French) []. Lecture Notes in Mathematics, 850:151–166, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088365/.

Meyer:1981:TEP

[Mey81e] P.-A. Meyer. Théorie élémentaire des processus à deux indices. (French) []. Lecture Notes in Mathematics, 863:1–39, 1981. CO-DEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091090/.

Meyer:1981:QTP

[Mey81f] P.-A. Meyer. Une question de théorie des processus. (French) []. Lecture Notes in Mathematics, 850:142, 1981. CODEN LN-MAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088363.

Meyer:1981:RSE

[Mey81g] Yves Meyer. Régularité des solutions des équations aux dérivées partielles non linéaires. (French) []. Lecture Notes in Mathematics, 842:293–302, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089941/.

Meyer:1982:ARD

[Mey82a] P.-A. Meyer. Appendice: Un résultat de D. Williams. (French) []. Lecture Notes in Mathematics, 920:133, 1982. CODEN LN-MAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092773.

Meyer:1982:GDS

[Mey82b] P.-A. Meyer. Géométrie différentielle stochastique (bis). (French) []. Lecture Notes in Mathematics, 921:165–207, 1982. CODEN LNMAA2. ISBN 3-540-11486-6 (print), 3-540-39167-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092650/.

Meyer:1982:IEE

[Mey82c] P.-A. Meyer. Interpolation entre espaces d’Orlicz. (French) []. Lecture Notes in Mathematics, 920:153–158, 1982. CODEN LN-MAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092778/.

Meyer:1982:NPO

[Mey82d] P.-A. Meyer. Note sur les processus d’Ornstein–Uhlenbeck. (French) []. Lecture Notes in Mathematics, 920:95–132, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092772/.

Meyer:1982:ISF

[Mey82e] P.-A. Meyer. Sur une inégalité de Stein. (French) []. Lecture Notes in Mathematics, 920:151–152, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092777.

Meyer:1982:WRQ

[Mey82f] R. E. Meyer. Wave reflection and quasiresonance. Lecture Notes in Mathematics, 942:84–112, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094742/.

Meyer:1984:TRP

[Mey84] P.-A. Meyer. Transformations de Riesz pour les lois gaussiennes. (French) []. Lecture Notes in Mathematics, 1059:179–193, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100043/.

Mhaskar:1984:SFT

[Mha84] H. N. Mhaskar. On the smoothness of Fourier transforms. Lecture Notes in Mathematics, 1070:202–207, 1984. CODEN LN-MAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099102/.

Michel:1981:BSC

[Mic81a] Pascal Michel. Borne supérieure de la complexité de la théorie de N muni de la relation de divisibilité. (French) []. Lecture Notes in Mathematics, 890:242–250, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095666/.

Michler:1981:BMO

[Mic81b] Gerhard O. Michler. On blocks with multiplicity one. Lecture Notes in Mathematics, 903:242–256, 1981. CODEN LN-MAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092996/.

Michalak:1983:MTG

[Mic83] Danuta Michalak. On middle and total graphs with coarseness number equal 1. Lecture Notes in Mathematics, 1018:139– 150, 1983. CODEN LNMAA2. ISBN 3-540-12687-2 (print), 3-540-38679-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071624/.

Mika:1982:SSP

[Mik82a] Janusz Mika. Singular-singularly perturbed linear equations in Banach spaces. Lecture Notes in Mathematics, 942:72–83, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094741/.

Mikhailov:1982:RIM

[Mik82b] A. V. Mikhaĭlov. Relativistically invariant models of the field theory integrable by the inverse scattering method. Lecture Notes in Mathematics, 970:186–196, 1982. CODEN LNMAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066031/.

Mills:1980:TAU

[Mil80a] George Mills. A tree analysis of unprovable combinatorial statements. Lecture Notes in Mathematics, 834:248–311, 1980. CO-DEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090170/.

Milojevic:1980:ASS

[Mil80b] P. S. Milojević. Approximation- solvability of some nonlinear operator equations with applications. Lecture Notes in Mathematics, 799:289–316, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089320/.

Milgram:1981:OIS

[Mil81a] R. James Milgram. Odd index subgroups of units in cyclotomic fields and applications. Lecture Notes in Mathematics, 854:269– 298, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089526/.

Millar:1981:CEM

[Mil81b] Terrence Millar. Counter-examples via model completions. Lecture Notes in Mathematics, 859:215–229, 1981. CODEN LN-MAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090949/.

Miller:1981:HRE

[Mil81c] David P. Miller. High recursively enumerable degrees and the anti-cupping property. Lecture Notes in Mathematics, 859:230– 245, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090950/.

Millet:1981:CRS

[Mil81d] Annie Millet. Convergence and regularity of strong submartingales. Lecture Notes in Mathematics, 863:50–58, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091092/.

Milojevic:1982:SOE

[Mil82] P. S. Milojević. Solvability of operator equations involving nonlinear perturbations of Fredholm mappings of nonnegative index and applications. Lecture Notes in Mathematics, 957:212–228, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066240/.

Milev:1983:EID

[Mil83a] N. Milev. Équivalence indéfiniment différentiable et équivalence analytique réelle pour les germes d’ensembles analytiques. (French) []. Lecture Notes in Mathematics, 1014:58–67, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072070/.

Millar:1983:MPA

[Mil83b] P. Warwick Millar. The minimax principle in asymptotic statistical theory. Lecture Notes in Mathematics, 976:75–265, 1983. CODEN LNMAA2. ISBN 3-540-11987-6 (print), 3-540-39458-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067986/.

Milman:1984:RBF

[Mil84a] Mario Milman. Rearrangements of BMO functions and interpolation. Lecture Notes in Mathematics, 1070:208–212, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099103/.

Milner:1984:LMT

[Mil84b] E. C. Milner. Lectures on the marriage theorem of Aharoni, Nash-Williams and Shelah. Lecture Notes in Mathematics, 1073: 55–79, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073105/.

Mines:1981:CMN

[Min81] Ray Mines. Cotorsion modules over Noetherian hereditary rings. Lecture Notes in Mathematics, 874:242–250, 1981. CODEN LN-MAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090538/.

Mingarelli:1983:ADE

[Min83a] Angelo B. Mingarelli. Asymptotic distribution of the eigenvalues of non-definite Sturm–Liouville problems. Lecture Notes in Mathematics, 1032:375–383, 1983. CODEN LNMAA2. ISBN 3-540-12702-X (print), 3-540-38689-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076808/.

Mingarelli:1983:Ca

[Min83b] Angelo B. Mingarelli. Chapter 1. Lecture Notes in Mathematics, 989:1–27, 1983. CODEN LNMAA2. ISBN 3-540-12294-X (print), 3-540-39865-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070769/.

Mingarelli:1983:Cb

[Min83c] Angelo B. Mingarelli. Chapter 2. Lecture Notes in Mathematics, 989:28–117, 1983. CODEN LNMAA2. ISBN 3-540-12294-X (print), 3-540-39865-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070770/.

Mingarelli:1983:Cc

[Min83d] Angelo B. Mingarelli. Chapter 3. Lecture Notes in Mathematics, 989:118–196, 1983. CODEN LNMAA2. ISBN 3-540-12294-X (print), 3-540-39865-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070771/.

Mingarelli:1983:Cd

[Min83e] Angelo B. Mingarelli. Chapter 4. Lecture Notes in Mathematics, 989:197–224, 1983. CODEN LNMAA2. ISBN 3-540-12294-X (print), 3-540-39865-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070772/.

Mingarelli:1983:Ce

[Min83f] Angelo B. Mingarelli. Chapter 5. Lecture Notes in Mathematics, 989:225–255, 1983. CODEN LNMAA2. ISBN 3-540-12294-X (print), 3-540-39865-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070773/.

Mingarelli:1983:E

[Min83g] Angelo B. Mingarelli. Errata. Lecture Notes in Mathematics, 989:256, 1983. CODEN LNMAA2. ISBN 3-540-12294-X (print), 3-540-39865-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070774.

Mingarelli:1983:VSI

[Min83h] Angelo B. Mingarelli. Volterra–Stieltjes Integral Equations and Generalized Ordinary Differential Expressions, volume 989 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CO-DEN LNMAA2. ISBN 3-540-12294-X (print), 3-540-39865-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiiv + 318 pp. LCCN QA3 .L28 no. 989. URL http:/

/link.springer.com/book/10.1007/BFb0070768;   http://
www.springerlink.com/content/978-3-540-39865-3.

Mislin:1982:CCP

[Mis82] Guido Mislin. Classes caractéristiques pour les représentations de groupes discrets. (French) []. Lecture Notes in Mathematics, 924: 296–309, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092937/.

McKean:1984:PP

[MIS^{+}84a] H. P. McKean, I. A. Ibragimov, V. N. Solev, S. V. Hruščëv, V. V. Peller, V. F. Gaposhkin, A. M. Vershik, V. N. Sudakov, and N. A. Sapogov. Probabilistic problems. Lecture Notes in Mathematics, 1043:82–107, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072186/.

Mishchenko:1984:SEO

[Mis84b] A. S. Mishchenko. The structure of extension orbits of Lie algebras. Lecture Notes in Mathematics, 1108:24–30, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099550/.

Misiewicz:1984:SRE

[Mis84c] Jolanta Misiewicz. Some remarks on elliptically contoured measures. Lecture Notes in Mathematics, 1080:170–174, 1984. CO-DEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099794/.

Mittelmann:1981:NSC

[Mit81] H.-D. Mittelmann. On the numerical solution of contact problems. Lecture Notes in Mathematics, 878:258–274, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090685/.

Mittelmann:1982:FSN

[Mit82a] H.-D. Mittelmann. A fast solver for nonlinear eigenvalue problems. Lecture Notes in Mathematics, 953:46–67, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069373/.

Mittelmann:1982:MGM

[Mit82b] H.-D. Mittelmann. Multi-grid methods for simple bifurcation problems. Lecture Notes in Mathematics, 960:558–575, 1982. CO-DEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069944/.

Mitter:1982:LNF

[Mit82c] Sanjoy K. Mitter. Lectures on nonlinear filtering and stochastic control. Lecture Notes in Mathematics, 972:170–207, 1982. CODEN LNMAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064862/.

Mitter:1983:RPN

[Mit83] Sanjoy K. Mitter. Remarks on pathwise nonlinear filtering. Lecture Notes in Mathematics, 979:232–235, 1983. CODEN LN-MAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066257/.

Mityagin:1984:ITM

[Mit84] Boris Mityagin. An interpolation theorem for modular spaces. Lecture Notes in Mathematics, 1070:10–23, 1984. CODEN LN-MAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099089/.

Miyanishi:1981:NCA

[Miy81a] Masayoshi Miyanishi. Non-complete Algebraic Surfaces, volume 857 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10703-7 (print), 3-540-38660-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 244 pp. LCCN QA3 .L28 no. 857. URL http:/

/link.springer.com/book/10.1007/BFb0090459;   http://
www.springerlink.com/content/978-3-540-38660-5.

Miyanishi:1981:NCAa

[Miy81b] Masayoshi Miyanishi. Non-complete algebraic surfaces with logarithmic Kodaira dimension-∞. Lecture Notes in Mathematics, 857:1–129, 1981. CODEN LNMAA2. ISBN 3-540-10703-7 (print), 3-540-38660-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090460/.

Miyanishi:1981:NCAb

[Miy81c] Masayoshi Miyanishi. Non-complete algebraic surfaces with logarithmic Kodaira dimension 0 or 1. Lecture Notes in Mathematics, 857:130–191, 1981. CODEN LNMAA2. ISBN 3-540-10703-7 (print), 3-540-38660-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090461/.

Miyanishi:1981:NCAc

[Miy81d] Masayoshi Miyanishi. Non-complete algebraic surfaces with logarithmic Kodaira dimension 2. Lecture Notes in Mathematics, 857: 192–233, 1981. CODEN LNMAA2. ISBN 3-540-10703-7 (print), 3-540-38660-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090462/.

Miyatake:1981:LPF

[Miy81e] Tohru Miyatake. On the length of proofs in a formal system of recursive arithmetic. Lecture Notes in Mathematics, 891:81– 108, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090981/.

Miyahara:1983:FEI

[Miy83a] Yoshio Miyahara. Filtering equations for infinite dimensional nonlinear filtering problems. Lecture Notes in Mathematics, 1033: 240–246, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099860/.

Miyanishi:1983:ARA

[Miy83b] Masayoshi Miyanishi. On the affine-ruledness of algebraic varieties. Lecture Notes in Mathematics, 1016:449–489, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099975/.

Miyaoka:1983:RVH

[Miy83c] Yoichi Miyaoka. A remark on variation of the Hodge structure on curves. Lecture Notes in Mathematics, 1016:354–369, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099971/.

Miyaoka:1984:TED

[Miy84] Reiko Miyaoka. Taut embeddings and Dupin hypersurfaces. Lecture Notes in Mathematics, 1090:15–23, 1984. CODEN LN-MAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101563/.

Miwa:1982:IHQ

[MJ82] Tetsuji Miwa and Michio Jimbo. Introduction to holonomic quantum fields. Lecture Notes in Mathematics, 925:28–36, 1982. CO-DEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093497/.

Mihalevic:1983:SQS

[MK83] M. V. Mihalevič and L. B. Košlai. Some questions of stability theory of the stochastic economical models. Lecture Notes in Mathematics, 982:123–135, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082064/.

Meis:1982:AMM

[MLM82] Th. Meis, H. Lehmann, and H. Michael. Application of the multigrid method to a nonlinear indefinite problem. Lecture Notes in Mathematics, 960:545–557, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069943/.

Maitre:1982:CNC

[MM82a] J. F. Maitre and F. Musy. The contraction number of a class of two-level methods; an exact evaluation for some finite element subspaces and model problems. Lecture Notes in Mathematics, 960:535–544, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069942/.

Mitter:1982:NFS

[MM82b] Sanjoy K. Mitter and Antonio Moro, editors. Nonlinear Filtering and Stochastic Control: Proceedings of the 3rd 1981 Session of the Centro Internazionale Matematico Estivo (C.I.M.E.), Held at Cortona, July 1–10, 1981, volume 972 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 972. URL http:

//link.springer.com/book/10.1007/BFb0064858;   http:/
/www.springerlink.com/content/978-3-540-39431-0.

Mori:1983:UMS

[MM83] Shigefumi Mori and Shigeru Mukai. The uniruledness of the moduli space of curves of genus 11. Lecture Notes in Mathematics, 1016:334–353, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099970/.

Mignot:1980:VPR

[MMP80] F. Mignot, F. Murat, and J.-P. Puel. Variation d’un point de retournement par rapport au domaine. (French) []. Lecture Notes in Mathematics, 782:222–254, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090435/.

Mazziotto:1981:DPCa

[MMS81] G. Mazziotto, E. Merzbach, and J. Szpirglas. Discontinuités des processus croissants et martingales à variation intégrable. (French) []. Lecture Notes in Mathematics, 863:59–83, 1981. CO-DEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091093/.

Matos:1981:SHT

[MN81] Mário C. Matos and Leopoldo Nachbin. Silva-holomorphy types. Lecture Notes in Mathematics, 843:437–487, 1981. CODEN LN-MAA2. ISBN 3-540-10560-3 (print), 3-540-38529-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089285/.

Mukherjea:1982:CNS

[MN82] Arunava Mukherjea and Anastase Nakassis. Convergence of nonhomogeneous stochastic chains with countable states: an application to measures on semigroups. Lecture Notes in Mathematics, 928:315–336, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093231/.

Massabo:1981:SNE

[MNP81] I. Massabo, P. Nistri, and J. Pejsachowicz. On the solvability of nonlinear equations in Banach spaces. Lecture Notes in Mathematics, 886:270–299, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092189/.

Madsen:1984:ATA

[MO84] Ib H. Madsen and Robert A. Oliver, editors. Algebraic Topology Aarhus 1982: Proceedings of a conference held in Aarhus, Denmark, August 1–7, 1982, volume 1051 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1051. URL http:/

/link.springer.com/book/10.1007/BFb0075559;   http://
www.springerlink.com/content/978-3-540-38782-4.

Moet:1982:AAF

[Moe82] H. J. K. Moet. Asymptotic analysis of the free boundary in singularly perturbed elliptic variational inequalities. Lecture Notes in Mathematics, 942:2–18, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094736/.

Moishezon:1981:SBC

[Moi81] B. G. Moishezon. Stable branch curves and braid monodromies. Lecture Notes in Mathematics, 862:107–192, 1981. CODEN LN-MAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090891/.

Mok:1983:AET

[Mok83] Mgaiming Mok. Application of an extension theorem for closed positive currents to Kähler geometry. Lecture Notes in Mathematics, 1028:239–252, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071685/.

Moler:1982:DML

[Mol82a] Cleve Moler. Demonstration of a matrix laboratory. Lecture Notes in Mathematics, 909:84–98, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092962/.

Molino:1982:SPK

[Mol82b] Pierre Molino. Simple pseudopotentials for the KdV-equation. Lecture Notes in Mathematics, 926:206–219, 1982. CODEN LN-MAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092665/.

Molzon:1983:SEV

[Mol83] R. E. Molzon. Some examples in value distribution theory. Lecture Notes in Mathematics, 981:90–100, 1983. CODEN LN-MAA2. ISBN 3-540-12003-3 (print), 3-540-39480-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066384/.

Moncrief:1980:BSE

[Mon80a] Vincent Moncrief. The branching of solutions of Einstein’s equations. Lecture Notes in Mathematics, 775:30–48, 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092019/.

Montgomery:1980:CC

[Mon80b] Susan Montgomery. Chain conditions. Lecture Notes in Mathematics, 818:69–88, 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091567/.

Montgomery:1980:FRF

[Mon80c] Susan Montgomery. Fixed Rings of Finite Automorphism Groups of Associative Rings, volume 818 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 126 pp. LCCN QA3 .L28 no. 818.

URL http://link.springer.com/book/10.1007/BFb0091561;
 http://www.springerlink.com/content/978-3-540-38308-
6.

Montgomery:1980:IOA

[Mon80d] Susan Montgomery. Inner and outer automorphisms of semiprime rings. Lecture Notes in Mathematics, 818:38–54, 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091565/.

Montgomery:1980:PI

[Mon80e] Susan Montgomery. Polynomial identities. Lecture Notes in Mathematics, 818:89–106, 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091568/.

Montgomery:1980:P

[Mon80f] Susan Montgomery. Preliminaries. Lecture Notes in Mathematics, 818:1–5, 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091562/.

Montgomery:1980:RNT

[Mon80g] Susan Montgomery. Rings with no |G|-torsion: Existence of fixed elements, the prime radical and Jacobson radical of R^{G}. Lecture Notes in Mathematics, 818:6–18, 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091563/.

Montgomery:1980:RNN

[Mon80h] Susan Montgomery. Rings with no nilpotent elements: Existence of fixed elements and traces. Lecture Notes in Mathematics, 818: 55–68, 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091566/.

Montgomery:1980:SRI

[Mon80i] Susan Montgomery. Simple rings: Inner and outer automorphisms, simple fixed rings, R as an R^{G}-module. Lecture Notes in Mathematics, 818:19–37, 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091564/.

Montgomery:1980:SGR

[Mon80j] Susan Montgomery. Skew group rings and applications to modules. Lecture Notes in Mathematics, 818:107–118, 1980. CODEN LNMAA2. ISBN 3-540-10232-9 (print), 3-540-38308-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091569/.

Montgomery:1982:TFA

[Mon82] Susan Montgomery. Trace functions and affine fixed rings for groups acting on non-commutative rings. Lecture Notes in Mathematics, 924:356–374, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092940/.

Montesinos:1984:SII

[Mon84] A. Montesinos. Some integral invariants of plane fields on Riemannian manifolds. Lecture Notes in Mathematics, 1045:134– 141, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072173/.

Moreau:1980:LSP

[Mor80a] Jean-Charles Moreau. Lemmes de Schwarz en plusieurs variables et applications arithmétiques. (French) []. Lecture Notes in Mathematics, 822:174–190, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097757/.

Morgan:1980:BTD

[Mor80b] Elizabeth J. Morgan. Balanced ternary designs with block size three. Lecture Notes in Mathematics, 829:186–198, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088912/.

Morton:1980:PGM

[Mor80c] K. W. Morton. Petrov–Galerkin methods for non-self-adjoint problems. Lecture Notes in Mathematics, 773:110–118, 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094167/.

Morales:1981:CPP

[Mor81a] Pedro Morales. Contraction principle in pseudo-uniform spaces. Lecture Notes in Mathematics, 886:300–308, 1981. CODEN LN-MAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092190/.

Morgan:1981:AGF

[Mor81b] John W. Morgan. Actions de groupes finis sur S^{3}: La conjecture de P. A. Smith [d’après Thurston et Meeks–Yau]. (French) []. Lecture Notes in Mathematics, 901:277–289, 1981. CODEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097203/.

Morrison:1981:CCG

[Mor81c] Kent Morrison. The connected component group of an algebra. Lecture Notes in Mathematics, 903:257–262, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092997/.

Mori:1982:TWC

[Mor82a] Shigefumi Mori. Threefolds whose canonical bundles are not numerically effective. Lecture Notes in Mathematics, 947:155– 189, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093587/.

Morton:1982:FEM

[Mor82b] K. W. Morton. Finite element methods for non-self-adjoint problems. Lecture Notes in Mathematics, 965:113–148, 1982. CODEN LNMAA2. ISBN 3-540-11967-1 (print), 3-540-39558-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063202/.

Morales:1984:BUS

[Mor84a] Pedro Morales. Boundedness for uniform semigroup-valued set functions. Lecture Notes in Mathematics, 1089:153–164, 1984.

CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072611/.

Moreno:1984:SPE

[Mor84b] Carlos J. Moreno. Some problems of effectivity in arithmetic, geometry and analysis. Lecture Notes in Mathematics, 1052:252– 272, 1984. CODEN LNMAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071545/.

Moschovakis:1981:DDT

[Mos81a] Joan Rand Moschovakis. A disjunctive decomposition theorem for classical theories. Lecture Notes in Mathematics, 873:250– 259, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090738/.

Moschovakis:1981:GHM

[Mos81b] Yiannis N. Moschovakis. On the Grilliot–Harrington–MacQueen theorem. Lecture Notes in Mathematics, 859:246–267, 1981. CO-DEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090951/.

Moschovakis:1981:OGP

[Mos81c] Yiannis N. Moschovakis. Ordinal games and playful models. Lecture Notes in Mathematics, 839:169–201, 1981. CODEN LN-MAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090241/.

Moschovakis:1983:SCS

[Mos83] Yiannis N. Moschovakis. Scales on coinductive sets. Lecture Notes in Mathematics, 1019:77–85, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071696/.

Moschovakis:1984:ARF

[Mos84a] Yiannis N. Moschovakis. Abstract recursion as a foundation for the theory of algorithms. Lecture Notes in Mathematics, 1104: 289–364, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099491/.

Moscovici:1984:LFH

[Mos84b] Henri Moscovici. Lefschetz formulae for Hecke operators. Lecture Notes in Mathematics, 1077:321–358, 1984. CODEN LN-MAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072343/.

Motohashi:1981:HFD

[Mot81] Nobuyoshi Motohashi. Homogeneous formulas and definability theorems. Lecture Notes in Mathematics, 891:109–116, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090982/.

Malta:1981:FVF

[MP81] I. P. Malta and J. Palis. Families of vector fields with finite modulus of stability. Lecture Notes in Mathematics, 898:212–229, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091915/.

McCoy:1982:CES

[MP82] Barry M. McCoy and Jacques H. H. Perk. Continuous exponents of spin correlation functions of inhomogeneous layered Ising models. Lecture Notes in Mathematics, 925:12–27, 1982. CO-DEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093495/;
 http://link.springer.com/chapter/10.1007/BFb0093496/.

Miranda:1983:IES

[MP83a] Guillermo Miranda and Henry Power. Integral equation solution of Oseen’s flow with a free surface. Lecture Notes in Mathematics, 1005:170–203, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112534/.

Morrison:1983:GSR

[MP83b] Ian Morrison and Ulf Persson. The group of sections on a rational elliptic surface. Lecture Notes in Mathematics, 997:321–347, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061651/.

Martin:1984:DDI

[MP84a] Bernd Martin and Gerhard Pfister. Distinguished deformations of isolated singularities of plane curves. Lecture Notes in Mathematics, 1056:255–263, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071775/.

Mokobodzki:1984:TDP

[MP84b] Gabriel Mokobodzki and Didier Pinchon, editors. Théorie du Potentiel: Proceedings of the Colloque Jacques Deny held at Orsay, June 20–23, 1983, volume 1096 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13894-3 (print), 3-540-39106-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1096. URL http:/

/link.springer.com/book/10.1007/BFb0100102;   http://
www.springerlink.com/content/978-3-540-39106-7.

Maury:1980:OMA

[MR80] Guy Maury and Jacques Raynaud. Ordres Maximaux au Sens de K. Asano. (French) [Maximum orders in the sense of K. Asano], volume 808 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 192 pp. LCCN QA3 .L28 no. 808. URL http:

//link.springer.com/book/10.1007/BFb0090774.

Manaster:1981:SDP

[MR81a] A. B. Manaster and J. B. Remmel. Some decision problems for subtheories of two-dimensional partial orderings. Lecture Notes in Mathematics, 859:202–214, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090948/.

McAloon:1981:MKP

[MR81b] Kenneth Mc Aloon and Jean-Pierre Ressayre. Les méthodes de Kieby–Paris et la théorie des ensembles. (French) []. Lecture Notes in Mathematics, 890:154–184, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095663/.

Mines:1981:DD

[MR81c] Ray Mines and Fred Richman. Dedekind domains. Lecture Notes in Mathematics, 873:16–30, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090724/.

Mehta:1983:HBC

[MR83] V. B. Mehta and A. Ramanathan. Homogeneous bundles in characteristic p. Lecture Notes in Mathematics, 997:315–320, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061650/.

McCabe:1984:FSH

[MR84a] J. H. McCabe and A. Sri Ranga. Ĵ fractions and the strong Hamburger moment problem. Lecture Notes in Mathematics, 1071: 231–237, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099621/.

Muller:1984:MSP

[MR84b] Gert H. Müller and Michael M. Richter, editors. Models and Sets: Proceedings of the Logic Colloquium held in Aachen, July 18–23, 1983 Part I, volume 1103 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1103. URL http:/

/link.springer.com/book/10.1007/BFb0099377;   http://
www.springerlink.com/content/978-3-540-39115-9.

Mauceri:1983:HAP

[MRW83] Giancarlo Mauceri, Fulvio Ricci, and Guido Weiss, editors. Harmonic Analysis: Proceedings of a Conference Held in Cortona, Italy, July 1–9, 1982, volume 992 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 992. URL http:

//link.springer.com/book/10.1007/BFb0069148;   http:/
/www.springerlink.com/content/978-3-540-39885-1.

Maharam:1980:OOF

[MS80a] Dorothy Maharam and A. H. Stone. One-to-one functions and a problem on subfields. Lecture Notes in Mathematics, 794:49– 52, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088212/.

Mlak:1980:DOM

[MS80b] W. Mlak and F. H. Szafraniec. Dilations with operator multipliers. Lecture Notes in Mathematics, 828:208–214, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097405/.

Mardesic:1981:STG

[MS81a] Sibe Mardešić and Jack Segal, editors. Shape Theory and Geometric Topology: Proceedings of a Conference Held at the Inter-University Centre of Postgraduate Studies, Dubrovnik, Yugoslavia, January 19–30, 1981, volume 870 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 870.

URL http://link.springer.com/book/10.1007/BFb0089702;
 http://www.springerlink.com/content/978-3-540-38749-

Mazziotto:1981:DPCb

[MS81b] G. Mazziotto and J. Szpirglas. Sur les discontinuités d’un processus cad-lag à deux indices. (French) []. Lecture Notes in Mathematics, 863:84–90, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091094/.

Milne:1981:CSVa

[MS81c] J. S. Milne and K.-Y. Shih. Conjugates of Shimura varieties. Lecture Notes in Mathematics, 900:280–356, 1981. CODEN LN-MAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38955-2_
7/.

Milne:1981:CSVb

[MS81d] J. S. Milne and K.-Y. Shih. Conjugates of Shimura varieties. Lecture Notes in Mathematics, 900:422, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38955-2_16.

Milne:1981:LCTa

[MS81e] J. S. Milne and K.-Y. Shih. Langlands’s construction of the Taniyama Group. Lecture Notes in Mathematics, 900:229–260, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38955-2_5/.

Milne:1981:LCTb

[MS81f] J. S. Milne and K.-Y. Shih. Langlands’s construction of the Taniyama Group. Lecture Notes in Mathematics, 900:420, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38955-2_14.

MacDonald:1982:EMA

[MS82a] John MacDonald and Arthur Stone. Essentially monadic adjunctions. Lecture Notes in Mathematics, 962:167–174, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066896/.

Marczyk:1982:CDL

[MS82b] A. Marczyk and Z. Skupień. Characterizations of different line graphs of graphs. Lecture Notes in Mathematics, 952:358–364, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061990/.

Mauldin:1982:RM

[MS82c] R. Daniel Mauldin and A. H. Stone. Realization of maps. Lecture Notes in Mathematics, 945:145–149, 1982. CODEN LN-MAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096668/.

Mohan:1982:PTN

[MS82d] A. L. Mohan and D. Suryanarayana. Perfect totient numbers. Lecture Notes in Mathematics, 938:101–105, 1982. CODEN LN-MAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097177/.

Molzon:1982:CTC

[MS82e] Robert E. Molzon and Bernard Shiffman. Capacity, Tchebycheff constant, and transfinite hyperdiameter on complex projective space. Lecture Notes in Mathematics, 919:337–357, 1982. CODEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097056/.

Martin:1983:ES

[MS83a] Donald A. Martin and John R. Steel. The extent of scales in L(R). Lecture Notes in Mathematics, 1019:86–96, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071697/.

McNulty:1983:INB

[MS83b] George F. McNulty and Caroline R. Shallon. Inherently nonfinitely based finite algebras. Lecture Notes in Mathematics, 1004: 206–231, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063439/.

Migliorini:1983:DS

[MS83c] F. Migliorini and J. Szép. On Γ-decomposition of semigroups. Lecture Notes in Mathematics, 998:385–403, 1983. CODEN LN-MAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062043/.

Montroll:1983:WCD

[MS83d] Elliott W. Montroll and Michael F. Shlesinger. On the wedding of certain dynamical processes in disordered complex materials to the theory of stable (lévy) distribution functions. Lecture Notes in Mathematics, 1035:109–137, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073256/.

Malone:1983:VBD

[MSP83] George H. Malone, Suk Youn Suh, and Stephen Prager. Variational bounds on the diffusive and hydrodynamic permeabilities of randomly perforated sheets. Lecture Notes in Mathematics, 1035:370–390, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073268/.

Meixner:1980:F

[MSW80a] Josef Meixner, Friedrich W. Schäfke, and Gerhard Wolf. Foundations. Lecture Notes in Mathematics, 837:1–62, 1980. CODEN LNMAA2. ISBN 3-540-10282-5 (print), 3-540-38411-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096195/.

Meixner:1980:MF

[MSW80b] Josef Meixner, Friedrich W. Schäfke, and Gerhard Wolf. Mathieu functions. Lecture Notes in Mathematics, 837:63–101, 1980.

CODEN LNMAA2. ISBN 3-540-10282-5 (print), 3-540-38411-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096196/.

Meixner:1980:MFS

[MSW80c] Josef Meixner, Friedrich W. Schäfke, and Gerhard Wolf. Mathieu Functions and Spheroidal Functions and Their Mathematical Foundations: Further Studies, volume 837 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10282-5 (print), 3-540-38411-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 126 pp. LCCN QA3 .L28 no. 837. URL http:/

/link.springer.com/book/10.1007/BFb0096194;   http://
www.springerlink.com/content/978-3-540-38411-3.

Meixner:1980:SF

[MSW80d] Josef Meixner, Friedrich W. Schäfke, and Gerhard Wolf. Spheroidal functions. Lecture Notes in Mathematics, 837:102– 116, 1980. CODEN LNMAA2. ISBN 3-540-10282-5 (print), 3-540-38411-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096197/.

Murphy:1984:BA

[MSW^{+}84] G. J. Murphy, M. R. F. Smyth, T. T. West, Vlastimil Pták, N. J. Young, Jaroslav Zemánek, H. G. Dales, Gavin Brown, William Moran, Satoru Igari, Wiesław Żelazko, L. de Branges, T. W. Gamelin, Donald Sarason, J. Wermer, et al. Banach algebras. Lecture Notes in Mathematics, 1043:48–81, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072185/.

Merle:1980:CAA

[MT80a] M. Merle and B. Teissier. Conditions d’adjonction, d’après du val. (French) []. Lecture Notes in Mathematics, 777:229–245, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085886/.

Misfeld:1980:DNS

[MT80b] Jürgen Misfeld and Helga Tecklenburg. Dimension of nearaffine spaces. Lecture Notes in Mathematics, 792:97–109, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088668/.

MacDonald:1982:DMI

[MT82] John MacDonald and Walter Tholen. Decomposition of morphisms into infinitely many factors. Lecture Notes in Mathematics, 962:175–189, 1982. CODEN LNMAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066897/.

Marcja:1984:CBS

[MT84] Annalisa Marcja and Carlo Toffalori. On Cantor–Bendixson spectra containing (1, 1) — I. Lecture Notes in Mathematics, 1103: 331–350, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099393/.

Muller:1981:LSH

[MTT81] Gert H. Müller, Gaisi Takeuti, and Tosiyuki Tugué, editors. Logic Symposia Hakone 1979, 1980: Proceedings of Conferences Held in Hakone, Japan March 21–24, 1979 and February 4–7, 1980, volume 891 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 891. URL http:/

/link.springer.com/book/10.1007/BFb0090974;   http://
www.springerlink.com/content/978-3-540-38633-9.

Mueller:1981:GBC

[Mue81] Julia Mueller. Gaps between consecutive zeta zeros. Lecture Notes in Mathematics, 899:141–147, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096457/.

Mueller:1982:ETD

[Mue82] Carl Mueller. Exit times of diffusions. Lecture Notes in Mathematics, 939:98–105, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096262/.

Mueller:1984:TPA

[Mue84] J. Mueller. On Thue’s principle and its applications. Lecture Notes in Mathematics, 1068:158–166, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099450/.

Mumford:1983:KDS

[Mum83] David Mumford. On the Kodaira dimension of the Siegel modular variety. Lecture Notes in Mathematics, 997:348–375, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061652/.

Munkholm:1980:SMV

[Mun80] Hans J. Munkholm. Simplices of maximal volume in hyperbolic space, Gromov’s norm, and Gromov’s proof of Mostow’s rigidity theorem (following Thurston). Lecture Notes in Mathematics, 788:109–124, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099242/.

Mundici:1984:AMT

[Mun84] Daniele Mundici. Abstract model-theory and nets of C∗-algebras: Noncommutative interpolation and preservation properties. Lecture Notes in Mathematics, 1103:351–377, 1984. CODEN LN-MAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099394/.

Murre:1982:CFT

[Mur82a] J. P. Murre. Classification of Fano threefolds according to Fano and Iskovskih. Lecture Notes in Mathematics, 947:35–92, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093585/.

Murty:1982:SRR

[Mur82b] M. Ram Murty. Some ω-results for Ramanujan’s τ-function. Lecture Notes in Mathematics, 938:123–137, 1982. CODEN LN-MAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097180/.

Murawski:1984:CNT

[Mur84] Roman Murawski. A contribution to nonstandard teratology. Lecture Notes in Mathematics, 1103:379–388, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099395/.

Musial:1983:WRN

[Mus83a] Kazimierz Musiał. The weak Radon–Nikodým property in conjugate Banach spaces. Lecture Notes in Mathematics, 991:182– 187, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061570/.

Musielak:1983:CMS

[Mus83b] Julian Musielak. Countably modulared spaces. Lecture Notes in Mathematics, 1034:112–128, 1983. CODEN LNMAA2. ISBN 3-540-12706-2 (print), 3-540-38692-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072213/.

Musielak:1983:FMD

[Mus83c] Julian Musielak. Families of modulars depending on a parameter. Lecture Notes in Mathematics, 1034:129–150, 1983. CODEN LNMAA2. ISBN 3-540-12706-2 (print), 3-540-38692-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072214/.

Musielak:1983:MS

[Mus83d] Julian Musielak. Modular spaces. Lecture Notes in Mathematics, 1034:1–32, 1983. CODEN LNMAA2. ISBN 3-540-12706-2 (print), 3-540-38692-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072211/.

Musielak:1983:OS

[Mus83e] Julian Musielak. Orlicz spaces. Lecture Notes in Mathematics, 1034:33–111, 1983. CODEN LNMAA2. ISBN 3-540-12706-2 (print), 3-540-38692-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072212/.

Musielak:1983:OSM

[Mus83f] Julian Musielak. Orlicz Spaces and Modular Spaces, volume 1034 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12706-2 (print), 3-540-38692-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 222 pp. LCCN QA3 .L28 no. 1034. URL http:/

/link.springer.com/book/10.1007/BFb0072210;   http://
www.springerlink.com/content/978-3-540-38692-6.

Musielak:1983:SAM

[Mus83g] Julian Musielak. Some applications of modular spaces. Lecture Notes in Mathematics, 1034:151–163, 1983. CODEN LN-MAA2. ISBN 3-540-12706-2 (print), 3-540-38692-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072215/.

Mutzbauer:1983:TGa

[Mut83a] Otto Mutzbauer. Type graph. Lecture Notes in Mathematics, 1006:228–252, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21560-9_10/.

Mutzbauer:1983:TGb

[Mut83b] Otto Mutzbauer. Type graph. Lecture Notes in Mathematics, 1006:228–252, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103705/.

Martinez-Villa:1980:ASE

[MV80] Roberto Martínez-Villa. Algebras stably equivalent to lhereditary. Lecture Notes in Mathematics, 832:396–431, 1980.

CODEN LNMAA2. ISBN 3-540-10264-7 (print), 3-540-38387-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088476/.

Makarov:1981:MNS

[MV81a] N. G. Makarov and V. I. Vasjunin. A model for noncontractions and stability of the continuous spectrum. Lecture Notes in Mathematics, 864:365–412, 1981. CODEN LNMAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097002/.

Martinez-Villa:1981:ASE

[MV81b] Roberto Martínez-Villa. Algebras stably equivalent to factors of hereditary. Lecture Notes in Mathematics, 903:222–241, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092995/.

Mawhin:1980:RNP

[MW80] J. Mawhin and M. Willem. Range of nonlinear perturbations of linear operators with an infinite dimensional kernel. Lecture Notes in Mathematics, 827:165–181, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091380/.

Marki:1981:SCS

[MW81] L. Márki and R. Wiegandt. On semisimple classes of semigroups with zero. Lecture Notes in Mathematics, 855:55–72, 1981. CO-DEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089809/.

Marki:1982:RRC

[MW82] L. Márki and R. Wiegandt. Remarks on radicals in categories. Lecture Notes in Mathematics, 962:190–196, 1982. CODEN LN-MAA2. ISBN 3-540-11961-2 (print), 3-540-39550-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066898/.

Mawhin:1984:VMB

[MW84] J. Mawhin and M. Willem. Variational methods and boundary value problems for vector second order differential equations and applications to the pendulum equation. Lecture Notes in Mathematics, 1107:181–192, 1984. CODEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101500/.

Mantero:1983:UBR

[MZ83] Anna Maria Mantero and Anna Zappa. Uniformly bounded representations and L^{P} -convolution operators on a free group. Lecture Notes in Mathematics, 992:333–343, 1983. CODEN LN-MAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069167/.

Nacinovich:1982:CAC

[Nac82] Mauro Nacinovich. Complex analysis and complexes of differential operators. Lecture Notes in Mathematics, 950:105–195, 1982. CODEN LNMAA2. ISBN 3-540-11596-X (print), 3-540-39366-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061877/.

Nagano:1982:SHM

[Nag82] T. Nagano. Stability of harmonic maps between symmetric spaces. Lecture Notes in Mathematics, 949:130–137, 1982. CODEN LN-MAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069760/.

Naitoh:1984:SSC

[Nai84] Hiroo Naitoh. Symmetric submanifolds of compact symmetric spaces. Lecture Notes in Mathematics, 1090:116–128, 1984. CO-DEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101572/.

Najman:1984:RCS

[Naj84] B. Najman. The rate of convergence in singular perturbations of parabolic equations. Lecture Notes in Mathematics, 1076:147– 167, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072774/.

Namikawa:1980:AVC

[Nam80a] Yukihiko Namikawa. An application of the Voronoi compactification to the theory of moduli of polarized abelian varieties. Lecture Notes in Mathematics, 812:95–112, 1980. CODEN LN-MAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091060/.

Namikawa:1980:BCS

[Nam80b] Yukihiko Namikawa. Boundary components and the structure of parabolic subgroups. Lecture Notes in Mathematics, 812:15– 28, 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091055/.

Namikawa:1980:C

[Nam80c] Yukihiko Namikawa. The case of g = 1. Lecture Notes in Mathematics, 812:11–14, 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091054/.

Namikawa:1980:ERT

[Nam80d] Yukihiko Namikawa. Examples: reduction theory of positive quadratic forms. Lecture Notes in Mathematics, 812:85–94, 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091059/.

Namikawa:1980:MPM

[Nam80e] Yukihiko Namikawa. Main problem and main results. Lecture Notes in Mathematics, 812:7–10, 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091053/.

Namikawa:1980:RSD

[Nam80f] Yukihiko Namikawa. Realization as a Siegel domain of the third kind, and Satake compactification. Lecture Notes in Mathematics, 812:29–38, 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091056/.

Namikawa:1980:SUP

[Nam80g] Yukihiko Namikawa. The Siegel upperhalf plane and the symplectic group. Lecture Notes in Mathematics, 812:1–6, 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091052/.

Namikawa:1980:TTE

[Nam80h] Yukihiko Namikawa. Theory of torus embeddings. Lecture Notes in Mathematics, 812:39–57, 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091057/.

Namikawa:1980:TCD

[Nam80i] Yukihiko Namikawa. Toroidal compactification due to Mumford. Lecture Notes in Mathematics, 812:58–84, 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091058/.

Namikawa:1980:TCS

[Nam80j] Yukihiko Namikawa. Toroidal Compactification of Siegel Spaces, volume 812 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10021-0 (print), 3-540-38176-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 162 pp. LCCN QA3 .L28 no. 812. URL http:/

/link.springer.com/book/10.1007/BFb0091051;   http://
www.springerlink.com/content/978-3-540-38176-1.

Namba:1981:BVC

[Nam81] Kanji Namba. Boolean valued combinatorics. Lecture Notes in Mathematics, 891:117–153, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090983/.

Narasimhan:1982:DCS

[Nar82] M. S. Narasimhan. Deformations of complex structures and holomorphic vector bundles. Lecture Notes in Mathematics, 950:196– 209, 1982. CODEN LNMAA2. ISBN 3-540-11596-X (print), 3-540-39366-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061878/.

Narkiewicz:1984:AF

[Nar84a] Władysław Narkiewicz. Additive functions. Lecture Notes in Mathematics, 1087:51–61, 1984. CODEN LNMAA2. ISBN 3-540-13872-2 (print), 3-540-39063-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100184/.

Narkiewicz:1984:GR

[Nar84b] Władysław Narkiewicz. General results. Lecture Notes in Mathematics, 1087:1–11, 1984. CODEN LNMAA2. ISBN 3-540-13872-2 (print), 3-540-39063-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100181/.

Narkiewicz:1984:LRS

[Nar84c] Władysław Narkiewicz. Linear recurrent sequences. Lecture Notes in Mathematics, 1087:28–50, 1984. CODEN LNMAA2. ISBN 3-540-13872-2 (print), 3-540-39063-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100183/.

Narkiewicz:1984:MF

[Nar84d] Władysław Narkiewicz. Multiplicative functions. Lecture Notes in Mathematics, 1087:62–95, 1984. CODEN LNMAA2. ISBN 3-540-13872-2 (print), 3-540-39063-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100185/.

Narkiewicz:1984:PLF

[Nar84e] Władysław Narkiewicz. Polynomial-like functions. Lecture Notes in Mathematics, 1087:96–113, 1984. CODEN LNMAA2. ISBN 3-540-13872-2 (print), 3-540-39063-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100186/.

Narkiewicz:1984:PS

[Nar84f] Władysław Narkiewicz. Polynomial sequences. Lecture Notes in Mathematics, 1087:12–27, 1984. CODEN LNMAA2. ISBN 3-540-13872-2 (print), 3-540-39063-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100182/.

Narkiewicz:1984:UDS

[Nar84g] Władysław Narkiewicz. Uniform Distribution of Sequences of Integers in Residue Classes, volume 1087 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13872-2 (print), 3-540-39063-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 125 pp. LCCN QA3 .L28 no. 1087.

URL http://link.springer.com/book/10.1007/BFb0100180;
 http://www.springerlink.com/content/978-3-540-39063-
3.

Nastasescu:1980:THP

[Năs80] Constantin Năstăsescu. Théorème de Hopkins pour les catégories de Grothendieck. (French) []. Lecture Notes in Mathematics, 825: 88–93, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089124/.

Nathanson:1981:WPS

[Nat81] Melvyn B. Nathanson. Waring’s problem for sets of density zero. Lecture Notes in Mathematics, 899:301–310, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096469/.

Nathanson:1984:EOS

[Nat84] Melvyn B. Nathanson. The exact order of subsets of additive bases. Lecture Notes in Mathematics, 1052:273–277, 1984. CO-DEN LNMAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071546/.

Naveira:1984:SLL

[Nav84a] A. M. Naveira. A Schur-like lemma for the nk-manifolds of constant type. Lecture Notes in Mathematics, 1045:142–146, 1984.

CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072174/.

Naveira:1984:DGP

[Nav84b] Antonio M. Naveira, editor. Differential Geometry: Proceedings of the International Symposium held at Peñíscola, Spain, October 3–10, 1982, volume 1045 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1045. URL http:/

/link.springer.com/book/10.1007/BFb0072159;   http://
www.springerlink.com/content/978-3-540-38766-4.

Nazarova:1981:PR

[Naz81] L. A. Nazarova. Poset representations. Lecture Notes in Mathematics, 882:345–356, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092504/.

Nevanlinna:1980:WMF

[NC80] Rolf Nevanlinna and Cabiria Andreian Cazacu. Wertverteilung meromorpher Funktionen in ein- und mehrfach zusammenhängenden Gebieten. (German) [], volume 783 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09759-7 (print), 3-540-38640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 145 pp. LCCN QA3 .L28 no. 783.

URL http://link.springer.com/book/10.1007/BFb0099698;
 http://www.springerlink.com/content/978-3-540-38640-
7.

Nestoridis:1983:IAF

[ND83] V. Nestoridis and N. Danikas. Interval averages of H^{1}-functions and BMO norm of inner functions. Lecture Notes in Mathematics, 992:174–192, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069158/.

Neeman:1982:GTG

[Ne’82] Yuval Ne’eman. Gauge-theory ghosts and ghost-gauge theories. Lecture Notes in Mathematics, 905:241–259, 1982. CODEN LN-MAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092441/.

Nelson:1981:CA

[Nel81] Evelyn Nelson. Z-continuous algebras. Lecture Notes in Mathematics, 871:315–334, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089915/.

Nel:1982:EAC

[Nel82] L. D. Nel. Enriched algebraic categories with applications in functional analysis. Lecture Notes in Mathematics, 915:247–259, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092884/.

Netuka:1980:CDP

[Net80] Ivan Netuka. The classical Dirichlet problem and its generalizations. Lecture Notes in Mathematics, 787:235–266, 1980. CODEN LNMAA2. ISBN 3-540-09967-0 (print), 3-540-39183-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086339/.

Netuka:1982:RSG

[Net82a] Ivan Netuka. La représentation de la solution généralisée à l’aide des solutions classiques du problème de Dirichlet. (French) []. Lecture Notes in Mathematics, 906:261–268, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093271/.

Netuka:1982:LDP

[Net82b] Ivan Netuka. L’unicité du problème de Dirichlet généralisé pour un compact. (French) []. Lecture Notes in Mathematics, 906:269– 281, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093272/.

Neumann:1980:IPH

[Neu80] Walter D. Neumann. An invariant of plumbed homology spheres. Lecture Notes in Mathematics, 788:125–144, 1980. CODEN LN-MAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099243/.

Neumaier:1981:CER

[Neu81] A. Neumaier. On a class of edge-regular graphs. Lecture Notes in Mathematics, 893:223–230, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091023/.

Neumaier:1982:SBZ

[Neu82a] Arnold Neumaier. Simple bounds for zeros of systems of equations. Lecture Notes in Mathematics, 953:88–105, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069376/.

Neumaier:1982:RSQ

[Neu82b] Dr. A. Neumaier. Regular sets and quasi-symmetric 2-designs. Lecture Notes in Mathematics, 969:258–275, 1982. CODEN LN-MAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062999/.

Neumann:1984:CQ

[Neu84a] B. H. Neumann. Commutative quandles. Lecture Notes in Mathematics, 1098:81–86, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099665/.

Neunzert:1984:INB

[Neu84b] H. Neunzert. An introduction to the nonlinear Boltzmann–Vlasov equation. Lecture Notes in Mathematics, 1048:60–110, 1984. CO-DEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071878/.

Newman:1980:P

[New80] M. F. Newman. Problems. Lecture Notes in Mathematics, 806: 249–254, 1980. CODEN LNMAA2. ISBN 3-540-10006-7 (print), 3-540-38120-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091274/.

Newman:1981:SAP

[New81] D. J. Newman. Sequences without arithmetic progressions. Lecture Notes in Mathematics, 899:311–314, 1981. CODEN LN-MAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096470/.

Newman:1984:MGP

[New84] M. F. Newman. Metabelian groups of prime-power exponent. Lecture Notes in Mathematics, 1098:87–98, 1984. CODEN LN-MAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099666/.

Nicoletti:1980:GCA

[Nic80] G. Nicoletti. Generating cryptomorphic axiomatizations of matroids. Lecture Notes in Mathematics, 792:110–113, 1980. CO-DEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088669/.

Nickel:1982:ANF

[Nic82] Karl Nickel. Das Auflösungsverhalten von nichtlinearen Fixmengen-Systemen. (German) []. Lecture Notes in Mathematics, 953:106– 124, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069377/.

Niculescu:1983:OCO

[Nic83] Constantin Niculescu. Order σ-continuous operators on Banach lattices. Lecture Notes in Mathematics, 991:188–201, 1983. CO-DEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061571/.

Niemi:1980:CWC

[Nie80] Hannu Niemi. On the construction of Wold–Cramér decomposition for bivariate stationary processes. Lecture Notes in Mathematics, 828:215–222, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097406/.

Niemi:1984:GIM

[Nie84] Hannu Niemi. Grothendieck’s inequality and minimal orthogonally scattered dilations. Lecture Notes in Mathematics, 1080: 175–187, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099795/.

Nakajima:1984:CQS

[NiW84] Haruhisa Nakajima and Kei ichi Watanabe. The classification of quotient singularities which are complete intersections. Lecture Notes in Mathematics, 1092:102–120, 1984. CODEN LN-MAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099359/.

Nocedal:1983:NMS

[NO83] Jorge Nocedal and Michael L. Overton. Numerical methods for solving inverse eigenvalue problems. Lecture Notes in Mathematics, 1005:212–226, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112536/.

Nobelis:1981:FAL

[Nob81] Ph. Nobelis. Fonctions aléatoires lipschitziennes. (French) []. Lecture Notes in Mathematics, 850:38–43, 1981. CODEN LN-MAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088359/.

Nocedal:1982:SLN

[Noc82] Jorge Nocedal. Solving large nonlinear systems of equations arising in mechanics. Lecture Notes in Mathematics, 909:132– 141, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092967/.

Norguet:1980:FPV

[Nor80a] François Norguet, editor. Fonctions de Plusieurs Variables Complexes IV: Séminaire François Norguet Octobre 1977–Juin 1979, volume 807 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 807. URL http:/

/link.springer.com/book/10.1007/BFb0090872;   http://
www.springerlink.com/content/978-3-540-38146-4.

Normann:1980:TS

[Nor80b] Dag Normann. Ct(n) as a topological space. Lecture Notes in Mathematics, 811:49–79, 1980. CODEN LNMAA2. ISBN 3-540-10019-9 (print), 3-540-38170-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098603/.

Normann:1980:CVR

[Nor80c] Dag Normann. Computability vs recursion. Lecture Notes in Mathematics, 811:80–115, 1980. CODEN LNMAA2. ISBN 3-540-10019-9 (print), 3-540-38170-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098604/.

Normann:1980:CS

[Nor80d] Dag Normann. The computable structure on Ct(k). Lecture Notes in Mathematics, 811:116–144, 1980. CODEN LNMAA2. ISBN 3-540-10019-9 (print), 3-540-38170-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098605/.

Normann:1980:CF

[Nor80e] Dag Normann. The countable functionals. Lecture Notes in Mathematics, 811:23–48, 1980. CODEN LNMAA2. ISBN 3-540-10019-9 (print), 3-540-38170-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098602/.

Normann:1980:MTS

[Nor80f] Dag Normann. The maximal type structure. Lecture Notes in Mathematics, 811:1–22, 1980. CODEN LNMAA2. ISBN 3-540-10019-9 (print), 3-540-38170-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098601/.

Normann:1980:RCF

[Nor80g] Dag Normann. Recursion on the Countable Functionals, volume 811 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10019-9 (print), 3-540-38170-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 191 pp. LCCN QA3 .L28 no. 811. URL http:/

/link.springer.com/book/10.1007/BFb0098600;   http://
www.springerlink.com/content/978-3-540-38170-9.

Normann:1980:S

[Nor80h] Dag Normann. Sections. Lecture Notes in Mathematics, 811: 145–172, 1980. CODEN LNMAA2. ISBN 3-540-10019-9 (print), 3-540-38170-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098606/.

Normann:1980:SFR

[Nor80i] Dag Normann. Some further results and topics. Lecture Notes in Mathematics, 811:173–185, 1980. CODEN LNMAA2. ISBN 3-540-10019-9 (print), 3-540-38170-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098607/.

Norsett:1980:CPC

[Nør80j] Syvert P. Nørsett. Collocation and perturbed collocation methods. Lecture Notes in Mathematics, 773:119–132, 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094168/.

Nottrot:1982:I

[Not82a] Roel Nottrot. Introduction. Lecture Notes in Mathematics, 963: 1–19, 1982. CODEN LNMAA2. ISBN 3-540-11963-9 (print), 3-540-39551-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067451/.

Nottrot:1982:OPMa

[Not82b] Roel Nottrot. Optimal processes on manifolds. Lecture Notes in Mathematics, 963:20–29, 1982. CODEN LNMAA2. ISBN 3-540-11963-9 (print), 3-540-39551-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067452/.

Nottrot:1982:OPMb

[Not82c] Roel Nottrot. Optimal Processes on Manifolds: an Application of Stokes’ Theorem, volume 963 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11963-9 (print), 3-540-39551-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 124 pp. LCCN QA3 .L28 no. 963.

URL http://link.springer.com/book/10.1007/BFb0067450;
 http://www.springerlink.com/content/978-3-540-39551-
5.

Nottrot:1982:PDF

[Not82d] Roel Nottrot. Processes, described by first order partial differential equations. Lecture Notes in Mathematics, 963:41–53, 1982. CODEN LNMAA2. ISBN 3-540-11963-9 (print), 3-540-39551-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067454/.

Nottrot:1982:PDO

[Not82e] Roel Nottrot. Processes, described by ordinary differential equations. Lecture Notes in Mathematics, 963:30–40, 1982. CODEN LNMAA2. ISBN 3-540-11963-9 (print), 3-540-39551-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067453/.

Nottrot:1982:PDS

[Not82f] Roel Nottrot. Processes, described by second order partial differential equations. Lecture Notes in Mathematics, 963:54–119, 1982. CODEN LNMAA2. ISBN 3-540-11963-9 (print), 3-540-39551-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067455/.

Nowak:1982:UMM

[Now82] Zenon P. Nowak. Use of the multigrid method for Laplacian problems in three dimensions. Lecture Notes in Mathematics, 960: 576–598, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069945/.

Nguyen-Quang-Do:1984:FCM

[NQ– D84] T. Nguyẽn-Quang-– Dõ. Formations de classes et modules d’Iwasawa. (French) []. Lecture Notes in Mathematics, 1068: 167–185, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099451/.

Nitecki:1980:GTD

[NR80] Zbigniew Nitecki and Clark Robinson, editors. Global Theory of Dynamical Systems: Proceedings of an International Conference Held at Northwestern University, Evanston, Illinois, June 18–22, 1979, volume 819 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10236-1 (print), 3-540-38312-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 819. URL http:

//link.springer.com/book/10.1007/BFb0086975;   http:/
/www.springerlink.com/content/978-3-540-38312-3.

Nicolau:1984:CHF

[NR84] M. Nicolau and A. Reventós. Compact Hausdorff foliations. Lecture Notes in Mathematics, 1045:147–153, 1984. CODEN LN-MAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072175/.

Nowak:1981:SPD

[NS81a] Sławomir Nowak and Stanisław Spież. Some properties of deformation dimension. Lecture Notes in Mathematics, 870:93– 104, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089710/.

Nualart:1981:CIP

[NS81b] D. Nualart and M. Sanz. The conditional independence property in filtrations associated to stopping lines. Lecture Notes in Mathematics, 863:202–210, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091101/.

Nelson:1982:NSS

[NSE82] P. Nelson, S. Sagong, and I. T. Elder. Numerical solution of singular two-point boundary-value problems by invariant imbedding. Lecture Notes in Mathematics, 909:166–174, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092971/.

Nikolov:1982:STV

[NT82] P. A. Nikolov and I. T. Todorov. Space-time versus phase space approach to relativistic particle dynamics. Lecture Notes in Mathematics, 970:197–216, 1982. CODEN LNMAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066032/.

Nualart:1981:MVI

[Nua81] D. Nualart. Martingales à variation indépendante du chemin. (French) []. Lecture Notes in Mathematics, 863:128–148, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091097/.

Nussbaum:1981:ENP

[Nus81] Roger D. Nussbaum. Eigenvectors of nonlinear positive operators and the linear Krein–Rutman theorem. Lecture Notes in Mathematics, 886:309–330, 1981. CODEN LNMAA2. ISBN 3-540-11152-2 (print), 3-540-38600-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092191/.

Nuttall:1982:CPA

[Nut82] J. Nuttall. The convergence of Padé approximants and their generalizations. Lecture Notes in Mathematics, 925:246–257, 1982.

CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093513/.

Nauwelaerts:1982:AZC

[NV82] Erna Nauwelaerts and Jan Van Geel. Arithmetical Zariski central rings. Lecture Notes in Mathematics, 951:132–142, 1982. CODEN LNMAA2. ISBN 3-540-11597-8 (print), 3-540-39371-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067329/.

Nygaard:1983:SAV

[Nyg83] Niels O. Nygaard. On supersingular abelian varieties. Lecture Notes in Mathematics, 1008:83–101, 1983. CODEN LN-MAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065702/.

Oberschelp:1982:ALC

[Obe82] W. Oberschelp. Asymptotic 0–1 laws in combinatorics. Lecture Notes in Mathematics, 969:276–292, 1982. CODEN LN-MAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063000/.

Oberlin:1983:MPR

[Obe83] Daniel M. Oberlin. L^{p}–L^{q} mapping properties of the Radon transform. Lecture Notes in Mathematics, 995:95–102, 1983. CODEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061889/.

Oda:1983:HGS

[Oda83a] Tadao Oda. Hironaka group schemes and resolution of singularities. Lecture Notes in Mathematics, 1016:295–312, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099968/.

Oda:1983:TCH

[Oda83b] Takayuki Oda. Transcendental cycles on Hilbert modular surfaces. Lecture Notes in Mathematics, 1016:261–270, 1983. CO-DEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099966/.

Odifreddi:1981:TD

[Odi81] Piergiorgio Odifreddi. Trees and degrees. Lecture Notes in Mathematics, 839:235–271, 1981. CODEN LNMAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090244/.

Oesterle:1980:TFW

[Oes80] Joseph Oesterlé. Travaux de Ferrero et Washington sur le nombre de classes d’idéaux des corps cyclotomiques. (French) []. Lecture Notes in Mathematics, 770:170–182, 1980. CODEN LN-MAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096242/.

OFarrell:1983:QRA

[O’F83] A. G. O’Farrell. Qualitative rational approximation on plane compacta. Lecture Notes in Mathematics, 995:103–122, 1983. CODEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061890/.

Ogus:1981:A

[Ogu81a] Arthur Ogus. Addendum 1989. Lecture Notes in Mathematics, 900:415, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38955-2_9.

Ogus:1981:GI

[Ogu81b] Arthur Ogus. General introduction. Lecture Notes in Mathematics, 900:416, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38955-2_10.

Ogus:1981:HCCa

[Ogu81c] Arthur Ogus. Hodge cycles and crystalline cohomology. Lecture Notes in Mathematics, 900:357–414, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38955-2_8/.

Ogus:1981:HCCb

[Ogu81d] Arthur Ogus. Hodge cycles and crystalline cohomology. Lecture Notes in Mathematics, 900:423, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38955-2_17.

Ogus:1981:NC

[Ogu81e] Arthur Ogus. Notations and conventions. Lecture Notes in Mathematics, 900:417, 1981. CODEN LNMAA2. ISBN 3-540-11174-3 (print), 3-540-38955-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38955-2_11.

Ohnita:1984:FSM

[Ohn84] Yoshihiro Ohnita. The first standard minimal immersions of compact irreducible symmetric spaces. Lecture Notes in Mathematics, 1090:37–49, 1984. CODEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101565/.

Ohsawa:1984:HEC

[Ohs84] Takeo Ohsawa. Holonorphic embedding of compact s.p.c. manifolds into complex manifolds as real hypersurfaces. Lecture Notes in Mathematics, 1090:64–76, 1984. CODEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101567/.

Ojanguren:1982:URP

[Oja82] Manuel Ojanguren. Unités représentées par des formes quadratiques ou par des normes réduites. (French) []. Lecture Notes in Mathematics, 967:291–299, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061907/.

Okada:1984:TPV

[Oka84] Susumu Okada. A tensor product vector integral. Lecture Notes in Mathematics, 1089:127–145, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072609/.

Okoh:1981:CMT

[Oko81] Frank Okoh. Cotorsion modules over tame finite-dimensional hereditary algebras. Lecture Notes in Mathematics, 903:263– 269, 1981. CODEN LNMAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092998/.

Okuyama:1984:ACO

[Oku84a] Yasuo Okuyama. Absolute convergence of orthogonal series. Lecture Notes in Mathematics, 1067:1–8, 1984. CODEN LN-MAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099713/.

Okuyama:1984:ANSa

[Oku84b] Yasuo Okuyama. Absolute Nörlund summability almost everywhere of Fourier series. Lecture Notes in Mathematics, 1067:9– 23, 1984. CODEN LNMAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099714/.

Okuyama:1984:ANSb

[Oku84c] Yasuo Okuyama. Absolute Nörlund summability almost everywhere of orthogonal series. Lecture Notes in Mathematics, 1067: 24–36, 1984. CODEN LNMAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099715/.

Okuyama:1984:ANSd

[Oku84d] Yasuo Okuyama. Absolute Nörlund summability factors of conjugate series of Fourier series. Lecture Notes in Mathematics, 1067: 61–76, 1984. CODEN LNMAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099718/.

Okuyama:1984:ANSc

[Oku84e] Yasuo Okuyama. Absolute Nörlund summability factors of Fourier series. Lecture Notes in Mathematics, 1067:47–60, 1984. CODEN LNMAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099717/.

Okuyama:1984:ARS

[Oku84f] Yasuo Okuyama. Absolute Riesz summability almost everywhere of orthogonal series. Lecture Notes in Mathematics, 1067:37– 46, 1984. CODEN LNMAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099716/.

Okuyama:1984:ASF

[Oku84g] Yasuo Okuyama. Absolute Summability of Fourier Series and Orthogonal Series, volume 1067 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 118 pp. LCCN QA3 .L28 no. 1067.

URL http://link.springer.com/book/10.1007/BFb0099712;
 http://www.springerlink.com/content/978-3-540-38902-
6.

Okuyama:1984:LPAb

[Oku84h] Yasuo Okuyama. Local property of absolute Nörlund summability of Fourier series. Lecture Notes in Mathematics, 1067:91– 111, 1984. CODEN LNMAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099720/.

Okuyama:1984:LPAa

[Oku84i] Yasuo Okuyama. Local property of absolute Riesz summability of Fourier series. Lecture Notes in Mathematics, 1067:77–90, 1984. CODEN LNMAA2. ISBN 3-540-13355-0 (print), 3-540-38902-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099719/.

Olech:1983:BSD

[Ole83] Czesław Olech. Boundary solutions of differential inclusion. Lecture Notes in Mathematics, 979:236–239, 1983. CODEN LN-MAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066258/.

Olech:1984:DSC

[Ole84] Czesław Olech. Decomposability as a substitute for convexity. Lecture Notes in Mathematics, 1091:193–205, 1984. CODEN LN-MAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098812/.

Oliva:1980:LNE

[Oli80] W. M. Oliva. The Levin–Nohel equation on the torus. Lecture Notes in Mathematics, 799:317–321, 1980. CODEN LN-MAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089321/.

Oliver:1981:FGR

[Oli81] Robert Oliver. Sk_{1} for finite group rings: III. Lecture Notes in Mathematics, 854:299–337, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089527/.

Olsson:1981:IBI

[Ols81] Jörn B. Olsson. Inequalities for blocktheoretic invariants. Lecture Notes in Mathematics, 903:270–284, 1981. CODEN LN-MAA2. ISBN 3-540-11179-4 (print), 3-540-38963-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092999/.

OMalley:1982:SPA

[O’M82] Robert E. O’Malley, Jr. A singular perturbations approach to reduced-order modeling and decoupling for large scale linear systems. Lecture Notes in Mathematics, 968:246–255, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064892/.

OMalley:1983:SFD

[O’M83] Robert E. O’Malley, Jr. Slow/fast decoupling for linear boundary value problems. Lecture Notes in Mathematics, 985:248– 266, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062370/.

Ono:1981:UEM

[ON81] Hiroakira Ono and Akira Nakamura. Undecidability of extensions of the monadic first-order theory of successor and twodimensional finite automata. Lecture Notes in Mathematics, 891: 155–174, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090984/.

Onneweer:1982:GLS

[Onn82] C. W. Onneweer. Generalized Lipschitz spaces and Herz spaces on certain totally disconnected groups. Lecture Notes in Mathematics, 939:106–121, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096263/.

Odlyzko:1980:CI

[OR80] A. Odlyzko and B. Richmond. On the compositions of an integer. Lecture Notes in Mathematics, 829:199–210, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088913/.

Obretenov:1983:CBE

[OR83] A. Obretenov and S. Rachev. Characterizations of the bivariate exponential distribution and Marshall–Olkin distribution and stability. Lecture Notes in Mathematics, 982:136–150, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082065/.

Orbanz:1984:ERA

[Orb84] Ulrich Orbanz. Embedded resolution of algebraic surfaces after Abhyankar (characteristic 0). Lecture Notes in Mathematics, 1101:1–49, 1984. CODEN LNMAA2. ISBN 3-540-13904-4 (print), 3-540-39125-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072259/.

Orecchia:1982:CSO

[Ore82] Ferruccio Orecchia. The conductor of some one-dimensional rings and the computation of their K-theory groups. Lecture Notes in Mathematics, 966:180–196, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062175/.

Orecchia:1983:GHF

[Ore83] Ferruccio Orecchia. Generalized Hilbert functions of Cohen– Macaulay varieties. Lecture Notes in Mathematics, 997:376– 390, 1983. CODEN LNMAA2. ISBN 3-540-12320-2 (print), 3-540-40047-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061653/.

Ortoleva:1980:DPA

[Ort80] Peter Ortoleva. Dynamic Padé approximant and behavior singularities in nonlinear physico-chemical systems. Lecture Notes in Mathematics, 782:255–264, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090436/.

Orzech:1982:BGC

[Orz82] Morris Orzech. Brauer groups and class groups for a Krull domain. Lecture Notes in Mathematics, 917:66–90, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092228/.

Onuchic:1980:QPC

[OS80] Nelson Onuchic and Adalberto Spezamiglio. Qualitative properties of certain ordinary differential systems. Lecture Notes in Mathematics, 799:334–347, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089323/.

Orlik:1981:CRG

[OS81] Peter Orlik and Louis Solomon. Complexes for reflection groups. Lecture Notes in Mathematics, 862:193–207, 1981. CODEN LN-MAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090892/.

Osgood:1984:EBD

[Osg84] Charles F. Osgood. Effective bounds on the Diophantine approximation of algebraic functions, and Nevanlinna theory. Lecture Notes in Mathematics, 1052:278–285, 1984. CODEN LN-MAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071547/.

Oshima:1981:FAS

[Osh81] Toshio Oshima. Fourier analysis on semisimple symmetric spaces. Lecture Notes in Mathematics, 880:357–369, 1981. CODEN LN-MAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090416/.

Oshima:1982:PRS

[Ōsh82] Yōichi Ōshima. Potential of recurrent symmetric Markov processes and its associated Dirichlet spaces. Lecture Notes in Mathematics, 923:260–275, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093047/.

Osterholz:1980:TG

[Ost80] D. Osterholz. Der Trägheitsschiefkörper. (German) []. Lecture Notes in Mathematics, 778:33–42, 1980. CODEN LN-MAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095920/.

Ostrovskii:1983:GEC

[Ost83] I. V. Ostrovskiĭ. On the growth of entire characteristic functions. Lecture Notes in Mathematics, 982:151–155, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082066/.

Ott:1981:SRR

[Ott81] Udo Ott. Some remarks on representation theory in finite geometry. Lecture Notes in Mathematics, 893:68–110, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091013/.

Orzech:1982:SRB

[OV82] M. Orzech and A. Verschoren. Some remarks on Brauer groups of Krull domains. Lecture Notes in Mathematics, 917:91–95, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092229/.

Overholt:1982:CEV

[Ove82a] Marius Overholt. A class of element and value regions for continued fractions. Lecture Notes in Mathematics, 932:194–205, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093314/.

Overton:1982:FAE

[Ove82b] Michael L. Overton. A fast algorithm for the Euclidean distance location problem. Lecture Notes in Mathematics, 909:99– 105, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092963/.

Osborne:1983:STF

[OW83] M. Scott Osborne and Garth Warner. The Selberg trace formula IV: Inner product formulae (final considerations). Lecture Notes in Mathematics, 1024:112–263, 1983. CODEN LNMAA2. ISBN 3-540-12725-9 (print), 3-540-38709-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071434/.

Palis:1983:GDP

[Pal83] J. Palis, Jr., editor. Geometric Dynamics: Proceedings of the International Symposium held at the Instituto de Matématica Pura e Aplicada Rio de Janeiro, Brasil, July–August 1981, volume 1007 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1007. URL http:/

/link.springer.com/book/10.1007/BFb0061405;   http://
www.springerlink.com/content/978-3-540-40969-4.

Palczewski:1984:CPB

[Pal84] Andrzej Palczewski. The Cauchy problem for the Boltzmann equation. A survey of recent results. Lecture Notes in Mathematics, 1048:202–206, 1984. CODEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071882/.

Paneitz:1982:HSS

[Pan82] Stephen M. Paneitz. Hermitian structures on solution varieties of nonlinear relativistic wave equations. Lecture Notes in Mathematics, 905:108–118, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092431/.

Panaretos:1983:ECM

[Pan83a] J. Panaretos. An elementary characterization of the multinomial and the multivariate hypergeometric distributions. Lecture Notes in Mathematics, 982:156–164, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082067/.

Paneitz:1983:DPN

[Pan83b] Stephen M. Paneitz. Determination of a polarization by nonlinear scattering, and examples of the resulting quantization. Lecture Notes in Mathematics, 1037:286–300, 1983. CODEN LN-MAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073177/.

Papanicolaou:1983:DRW

[Pap83] George C. Papanicolaou. Diffusions and random walks in random media. Lecture Notes in Mathematics, 1035:391–399, 1983.

CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073269/.

Pap:1984:DGM

[Pap84a] Gyula Pap. Dependence of Gaussian measure on covariance in Hilbert space. Lecture Notes in Mathematics, 1080:188–194, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099796/.

Papangelou:1984:URD

[Pap84b] F. Papangelou. The uniqueness of regular DLR measures for certain one-dimensional spin systems. Lecture Notes in Mathematics, 1095:36–44, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099120/.

Paquet:1982:FDP

[Paq82] Luc Paquet. Formule de Duhamel et problème de la chaleur. (French) []. Lecture Notes in Mathematics, 906:282–307, 1982. CODEN LNMAA2. ISBN 3-540-11185-9 (print), 3-540-38971-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093273/.

Paris:1980:HCM

[Par80] J. B. Paris. A hierarchy of cuts in models of arithmetic. Lecture Notes in Mathematics, 834:312–337, 1980. CODEN LN-MAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090171/.

Paris:1981:SCR

[Par81] J. B. Paris. Some conservation results for fragments of arithmetic. Lecture Notes in Mathematics, 890:251–262, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095667/.

Pardon:1982:GCW

[Par82a] William Pardon. A ‘gersten conjecture’ for Witt groups. Lecture Notes in Mathematics, 967:300–315, 1982. CODEN LN-MAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061908/.

Pardoux:1982:ENL

[Par82b] É. Pardoux. Equations of non-linear filtering; and application to stochastic control with partial observation. Lecture Notes in Mathematics, 972:208–248, 1982. CODEN LNMAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064863/.

Parisi:1982:TIS

[Par82c] G. Parisi. Trace identities for the Schrödinger operator and the WKB method. Lecture Notes in Mathematics, 925:178–183, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093508/.

Park:1982:SBW

[Par82d] Chull Park. Stochastic barriers for the Wiener process and a mathematical model. Lecture Notes in Mathematics, 939:122– 130, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096264/.

Parter:1982:SFB

[Par82e] Seymour V. Parter. On the swirling flow between rotating coaxial disks: a survey. Lecture Notes in Mathematics, 942:258–280, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094752/.

Parthasarathy:1983:AAU

[Par83a] T. Parthasarathy. Assorted applications of univalence mapping results. Lecture Notes in Mathematics, 977:77–89, 1983. CODEN LNMAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065575/.

Parthasarathy:1983:FGU

[Par83b] T. Parthasarathy. Fundamental global univalence results of Gale– Nikaido–Inada. Lecture Notes in Mathematics, 977:17–27, 1983. CODEN LNMAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065569/.

Parthasarathy:1983:FGR

[Par83c] T. Parthasarathy. Further generalizations and remarks. Lecture Notes in Mathematics, 977:90–100, 1983. CODEN LN-MAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065576/.

Parthasarathy:1983:GHB

[Par83d] T. Parthasarathy. Global homeomorphisms between finite dimensional spaces. Lecture Notes in Mathematics, 977:28–40, 1983. CODEN LNMAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065570/.

Parthasarathy:1983:GUR

[Par83e] T. Parthasarathy. Global univalent results in R^{2} and R^{3}. Lecture Notes in Mathematics, 977:49–58, 1983. CODEN LN-MAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065572/.

Parthasarathy:1983:GUT

[Par83f] T. Parthasarathy. On Global Univalence Theorems, volume 977 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 106 pp. LCCN QA3 .L28 no. 977. URL http:/

/link.springer.com/book/10.1007/BFb0065566;   http://
www.springerlink.com/content/978-3-540-39462-4.

Parthasarathy:1983:GSA

[Par83g] T. Parthasarathy. On the global stability of an autonomous system on the plane. Lecture Notes in Mathematics, 977:59–67, 1983. CODEN LNMAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065573/.

Parthasarathy:1983:MM

[Par83h] T. Parthasarathy. p-matrices and n-matrices. Lecture Notes in Mathematics, 977:6–16, 1983. CODEN LNMAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065568/.

Parthasarathy:1983:PSP

[Par83i] T. Parthasarathy. Preliminaries and statement of the problem. Lecture Notes in Mathematics, 977:1–5, 1983. CODEN LNMAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065567/.

Parthasarathy:1983:SCV

[Par83j] T. Parthasarathy. Scarf’s conjecture and its validity. Lecture Notes in Mathematics, 977:41–48, 1983. CODEN LN-MAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065571/.

Parthasarathy:1983:UML

[Par83k] T. Parthasarathy. Univalence for mappings with Leontief type Jacobians. Lecture Notes in Mathematics, 977:68–76, 1983. CODEN LNMAA2. ISBN 3-540-11988-4 (print), 3-540-39462-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065574/.

Paszkiewicz:1980:RBO

[Pas80] A. Paszkiewicz. Representation of a bounded operator as a finite linear combination of projectors and some inequalities for a functional on B(H). Lecture Notes in Mathematics, 828:223– 233, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097407/.

Pasemann:1983:LNA

[Pas83] F. B. Pasemann. Linearized non-abelian gauge quantum field theories. Lecture Notes in Mathematics, 1037:81–114, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073170/.

Paulauskas:1980:RCC

[Pau80] V. Paulauskas. The rates of convergence in the central limit theorem in Banach spaces. Lecture Notes in Mathematics, 828:234– 243, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097408/.

Paulauskas:1983:DNG

[Pau83] Vygantas Paulauskas. On the density of the norm of Gaussian vector in Banach spaces. Lecture Notes in Mathematics, 990: 178–197, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064271/.

Pavel:1984:SPN

[Pav84] N. H. Pavel. Some problems on non-linear semigroups and the blow-up of integral solutions. Lecture Notes in Mathematics, 1076: 168–186, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072775/.

Pandit:1982:DSI

[PD82a] Sudhakar G. Pandit and Sadashiv G. Deo. Differential Systems Involving Impulses, volume 954 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11606-0 (print), 3-540-39428-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 102 pp. LCCN QA3 .L28 no. 954.

URL http://link.springer.com/book/10.1007/BFb0067476;
 http://www.springerlink.com/content/978-3-540-39428-
0.

Pandit:1982:EU

[PD82b] Sudhakar G. Pandit and Sadashiv G. Deo. Existence and uniqueness. Lecture Notes in Mathematics, 954:11–33, 1982. CODEN LNMAA2. ISBN 3-540-11606-0 (print), 3-540-39428-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067478/.

Pandit:1982:IS

[PD82c] Sudhakar G. Pandit and Sadashiv G. Deo. Impulsive systems. Lecture Notes in Mathematics, 954:56–77, 1982. CODEN LN-MAA2. ISBN 3-540-11606-0 (print), 3-540-39428-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067480/.

Pandit:1982:LSM

[PD82d] Sudhakar G. Pandit and Sadashiv G. Deo. Lyapunov’s second method. Lecture Notes in Mathematics, 954:78–97, 1982. CODEN LNMAA2. ISBN 3-540-11606-0 (print), 3-540-39428-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067481/.

Pandit:1982:P

[PD82e] Sudhakar G. Pandit and Sadashiv G. Deo. Preliminaries. Lecture Notes in Mathematics, 954:1–10, 1982. CODEN LNMAA2. ISBN 3-540-11606-0 (print), 3-540-39428-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067477/.

Pandit:1982:SAE

[PD82f] Sudhakar G. Pandit and Sadashiv G. Deo. Stability and asymptotic equivalence. Lecture Notes in Mathematics, 954:34–55, 1982. CODEN LNMAA2. ISBN 3-540-11606-0 (print), 3-540-39428-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067479/.

Puystjens:1980:MPI

[PdS80] R. Puystjens and H. de Smet. The Moore–Penrose inverse for matrices over skew polynomial rings. Lecture Notes in Mathematics, 825:94–103, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089125/.

Pedersen:1984:ICC

[Ped84a] Erik Kjær Pedersen. K_{−i}-invariants of chain complexes. Lecture Notes in Mathematics, 1060:174–186, 1984. CODEN LN-MAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099933/.

Pedersen:1984:SAN

[Ped84b] Pauli Pedersen. Sensitivity analysis for non-selfadjoint problems. Lecture Notes in Mathematics, 1086:119–130, 1984. CODEN LNMAA2. ISBN 3-540-13871-4 (print), 3-540-39061-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073072/.

Peetre:1984:TIS

[Pee84] Jaak Peetre. The theory of interpolation spaces — its origin, prospects for the future. Lecture Notes in Mathematics, 1070: 1–9, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099088/.

Penney:1984:HAU

[Pen84] Richard Penney. Harmonic analysis on unbounded homogeneous domains in C^{n}. Lecture Notes in Mathematics, 1077:359– 374, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072344/.

Percsy:1981:GUE

[Per81a] Nicolas Percsy. Géométries uniquely embeddable in projective spaces. Lecture Notes in Mathematics, 893:231–241, 1981. CO-DEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091024/.

Perrin:1981:RSI

[Per81b] Patrice Perrin. Representations de Schrödinger. Indice de Maslov et groupe metaplectique. Lecture Notes in Mathematics, 880: 370–407, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090417/.

Perjes:1982:ITP

[Per82a] Zoltán Perjés. Introduction to twistor particle theory. Lecture Notes in Mathematics, 970:53–72, 1982. CODEN LN-MAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066024/.

Perkins:1982:LTP

[Per82b] Edwin Perkins. Local time and pathwise uniqueness for stochastic differential equations. Lecture Notes in Mathematics, 920: 201–208, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092781/.

Perkins:1983:SIP

[Per83a] Edwin Perkins. Stochastic integrals and progressive measurability — an example. Lecture Notes in Mathematics, 986:67–71, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068299/.

Perkins:1983:SPN

[Per83b] Edwin Perkins. Stochastic processes and nonstandard analysis. Lecture Notes in Mathematics, 983:162–185, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065339/.

Persson:1984:DSI

[Per84] L.-E. Persson. Descriptions of some interpolation spaces in offdiagonal cases. Lecture Notes in Mathematics, 1070:213–231, 1984. CODEN LNMAA2. ISBN 3-540-13363-1 (print), 3-540-38913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099104/.

Petridis:1980:DVD

[Pet80] N. C. Petridis. Distance and volume decreasing theorems for a family of harmonic mappings of Riemannian manifolds. Lecture Notes in Mathematics, 781:157–170, 1980. CODEN LN-MAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097653/.

Peters:1981:CSC

[Pet81] H. Peters. Change of structure and chaos for solutions of ẋ(t) = −f(x(t − 1)). Lecture Notes in Mathematics, 878:325–350, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090687/.

Peterson:1982:ALA

[Pet82] Dale H. Peterson. Affine Lie algebras and theta-functions. Lecture Notes in Mathematics, 933:166–175, 1982. CODEN LN-MAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093360/.

Petrova:1983:SCU

[Pet83] L. L. Petrova. On the stability of characterizations of the unit distribution. Lecture Notes in Mathematics, 982:165–171, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082068/.

Dinghas:1980:PKA

[phAD80a] Dr. phil. habil. Alexander Dinghas. Das Problem der konformen Abbildung von universellen Überlagerungsflächen auf die Einheitskreisscheibe. Invariante partielle Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 783:46–67, 1980. CODEN LNMAA2. ISBN 3-540-09759-7 (print), 3-540-38640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099701/.

Dinghas:1980:BCF

[phAD80b] Dr. phil. habil. Alexander Dinghas. Der Begriff der charakteristischen Funktion. Nevanlinnas Charakterisierung rationaler Stellen. Der erste und der zweite Nevanlinnasche Hauptsatz der Wertverteilungstheorie. Ee Nevanlinnaschen Defektrelationen für den parabolischen und den hyperbolischen Fall. (German) []. Lecture Notes in Mathematics, 783:68–116, 1980. CODEN LN-MAA2. ISBN 3-540-09759-7 (print), 3-540-38640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099702/.

Dinghas:1980:GFE

[phAD80c] Dr. phil. habil. Alexander Dinghas. Die Greensche Funktion und die Evans–Selbergsche Kapazitätsfunktion.Eine allgemeine Formel von F. und R. Nevanlinna. (German) []. Lecture Notes in Mathematics, 783:20–46, 1980. CODEN LNMAA2. ISBN 3-540-09759-7 (print), 3-540-38640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099700/.

Dinghas:1980:HFM

[phAD80d] Dr. phil. habil. Alexander Dinghas. Harmonische Funktionen. Das Maximumprinzip für harmonische und stetige subharmonische Funktionen. Klassische Integralidentitäten. Die Grundformel der logarithmischen Methode. Anwendungen. (German) []. Lecture Notes in Mathematics, 783:1–20, 1980. CODEN LNMAA2. ISBN 3-540-09759-7 (print), 3-540-38640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099699/.

Dinghas:1980:WMF

[phAD80e] Dr. phil. habil. Alexander Dinghas. Wertverteilungsprobleme meromorpher Funktionen in mehrfach zusammenhängenden Gebieten. (German) []. Lecture Notes in Mathematics, 783:116– 131, 1980. CODEN LNMAA2. ISBN 3-540-09759-7 (print), 3-540-38640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099703/.

Pelczynski:1984:S

[PHM^{+}84] A. Pełczyński, V. P. Havin, G. J. Murphy, T. T. West, Yu. I. Lyubich, A. P. Calderón, Donald Sarason, A. A. Gol’dberg, V. S. Azarin, Walter Rudin, and E. M. Dyn’kin. Solutions. Lecture Notes in Mathematics, 1043:664–699, 1984. CODEN LN-MAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072197/.

Pichorides:1980:NES

[Pic80a] S. K. Pichorides. On the l’ norm of exponential sums. Lecture Notes in Mathematics, 781:171–176, 1980. CODEN LN-MAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097654/.

Pickert:1980:PPE

[Pic80b] Günter Pickert. Partial planes with exactly two complete parallel classes. Lecture Notes in Mathematics, 792:114–127, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088670/.

Pickert:1980:PFM

[Pic80c] Günter Pickert. A problem of free mobility. Lecture Notes in Mathematics, 792:128–131, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088671/.

Pierre:1980:RPP

[Pie80] Michel Pierre. Representant precis d’un potentiel parabolique. (French) []. Lecture Notes in Mathematics, 814:186–228, 1980. CODEN LNMAA2. ISBN 3-540-10025-3 (print), 3-540-38189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094155/.

Pierce:1983:TPB

[Pie83a] R. S. Pierce. Tensor products of Boolean algebras. Lecture Notes in Mathematics, 1004:232–239, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063440/.

Pierzchalski:1983:QDM

[Pie83b] A. Pierzchalski. On quasiconformal deformations on manifolds. Lecture Notes in Mathematics, 1013:171–181, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066528/.

Pillay:1981:CMA

[Pil81a] Anand Pillay. Cuts in models of arithmetic. Lecture Notes in Mathematics, 890:13–20, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095653/.

Pillay:1981:MPA

[Pil81b] Anand Pillay. Models of Peano arithmetic. Lecture Notes in Mathematics, 890:1–12, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095652/.

Pillay:1981:PPD

[Pil81c] Anand Pillay. Partition properties and definable types in Peano arithmetic. Lecture Notes in Mathematics, 890:263–269, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095668/.

Pinkham:1980:RSP

[Pin80a] H. Pinkham. Résolution simultanée de points doubles rationnels. (French) []. Lecture Notes in Mathematics, 777:179–203, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085883/.

Pinkham:1980:SKF

[Pin80b] H. Pinkham. Singularités de Klein — I. (French) []. Lecture Notes in Mathematics, 777:1–9, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085873/.

Pinkham:1980:SKI

[Pin80c] H. Pinkham. Singularités de Klein — II. (French) []. Lecture Notes in Mathematics, 777:10–20, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085874/.

Pinkham:1980:SRS

[Pin80d] H. Pinkham. Singularités rationnelles de surfaces. (French) []. Lecture Notes in Mathematics, 777:147–178, 1980. CODEN LN-MAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085882/.

Pinsky:1981:HSP

[Pin81] Mark A. Pinsky. Homogenization and stochastic parallel displacement. Lecture Notes in Mathematics, 851:271–284, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088731/.

Pintz:1984:RTP

[Pin84] J. Pintz. On the remainder term of the prime number formula and the zeros of Riemann’s zeta-function. Lecture Notes in Mathematics, 1068:186–197, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099452/.

Pisier:1980:NEB

[Pis80] Gilles Pisier. De nouveaux espaces de Banach sans la propriété d’approximation [d’après A. Szankowski]. (French) []. Lecture Notes in Mathematics, 770:312–327, 1980. CODEN LN-MAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096249/.

Pisier:1982:DBT

[Pis82] Gilles Pisier. On the duality between type and cotype. Lecture Notes in Mathematics, 939:131–144, 1982. CODEN LN-MAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096265/.

Pisier:1983:SAM

[Pis83] Gilles Pisier. Some applications of the metric entropy condition to harmonic analysis. Lecture Notes in Mathematics, 995:123– 154, 1983. CODEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061891/.

Pitman:1981:NMI

[Pit81] Jim Pitman. A note on L_{2} maximal inequalities. Lecture Notes in Mathematics, 850:251–258, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088372/.

Plesken:1981:ATO

[Ple81] Wilhelm Plesken. Applications of the theory of orders to crystallographic groups. Lecture Notes in Mathematics, 882:37–92, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092487/.

Pless:1982:UCC

[Ple82] Vera Pless. On the uses of contracted codes. Lecture Notes in Mathematics, 952:115–122, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061975/.

Plesken:1983:BM

[Ple83a] Wilhelm Plesken. Blocks of multiplicity 1. Lecture Notes in Mathematics, 1026:80–91, 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071563/.

Plesken:1983:BCD

[Ple83b] Wilhelm Plesken. Blocks with cyclic defect groups. Lecture Notes in Mathematics, 1026:131–145, 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071566/.

Plesken:1983:CFG

[Ple83c] Wilhelm Plesken. The conductor formula for graduated hulls of selfdual orders. Lecture Notes in Mathematics, 1026:43–58, 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071561/.

Plesken:1983:EGR

[Ple83d] Wilhelm Plesken. Examples of group rings. Lecture Notes in Mathematics, 1026:92–108, 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071564/.

Plesken:1983:GGO

[Ple83e] Wilhelm Plesken. Graduated and graduable orders. Lecture Notes in Mathematics, 1026:7–42, 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071560/.

Plesken:1983:GRF

[Ple83f] Wilhelm Plesken. Group Rings of Finite Groups Over p-adic Integers, volume 1026 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 151 pp. LCCN QA3 .L28 no. 1026. URL http:

//link.springer.com/book/10.1007/BFb0071558;   http:/
/www.springerlink.com/content/978-3-540-38717-6.

Plesken:1983:I

[Ple83g] Wilhelm Plesken. Introduction. Lecture Notes in Mathematics, 1026:1–6, 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071559/.

Plesken:1983:PBO

[Ple83h] Wilhelm Plesken. The principal 2-block of SL_{2}(q) for odd prime powers q. Lecture Notes in Mathematics, 1026:109–130, 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071565/.

Plesken:1983:SOD

[Ple83i] Wilhelm Plesken. Selfdual orders with decomposition numbers 0 and 1. Lecture Notes in Mathematics, 1026:59–79, 1983. CODEN LNMAA2. ISBN 3-540-12728-3 (print), 3-540-38717-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071562/.

Pagallo:1983:BQG

[PM83] Giulia Pagallo and Consuelo Maulino. A bipartite quotient graph model for unsymmetric matrices. Lecture Notes in Mathematics, 1005:227–239, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112537/.

Pardee:1983:MOS

[PMC83] W. J. Pardee, W. L. Morris, and B. N. Cox. Microscopic origins of stochastic crack growth. Lecture Notes in Mathematics, 1035:414–431, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073271/.

Poenaru:1980:TTD

[Poé80] Valentin Poénaru. Travaux de Thurston sur les difféomorphismes des surfaces et l’espace de Teichmüller. (French) []. Lecture Notes in Mathematics, 770:66–79, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096236/.

Poguntke:1980:SSM

[Pog80] Detlev Poguntke. Symmetry (or simple modules) of some Banach algebras. Lecture Notes in Mathematics, 781:177–193, 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097655/.

Pohl:1981:PLR

[Poh81a] William F. Pohl. The probability of linking of random closed curves. Lecture Notes in Mathematics, 894:113–126, 1981. CO-DEN LNMAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096227/.

Pohlers:1981:PTA

[Poh81b] W. Pohlers. Proof-theoretical analysis of ID_{ν} by the method of local predicativity. Lecture Notes in Mathematics, 897:261– 357, 1981. CODEN LNMAA2. ISBN 3-540-11170-0 (print), 3-540-38649-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091900/.

Pollett:1984:DAN

[Pol84] P. K. Pollett. Distributional approximations for networks of quasireversible queues. Lecture Notes in Mathematics, 1095:108– 129, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099125/.

Poornima:1983:SS

[Poo83] S. Poornima. On the Sobolev spaces W^{k,1}(R^{n}). Lecture Notes in Mathematics, 992:161–173, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069157/.

Popa:1983:NLS

[Pop83a] Eugen Popa. Natural localisation of a standard H-cone. Lecture Notes in Mathematics, 1014:220–227, 1983. CODEN LN-MAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072081/.

Popa:1983:PRI

[Pop83b] Nicolae Popa. The primariness of rearrangement invariant function p-spaces, 0 < p ≤ 1. Lecture Notes in Mathematics, 991: 202–214, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061572/.

Popescu:1984:ZUT

[Pop84] Dorin Popescu. On Zariski’s uniformization theorem. Lecture Notes in Mathematics, 1056:264–296, 1984. CODEN LN-MAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071776/.

Porter:1983:PMN

[Por83] R. Michael Porter. Prescribed monodromy on noncompact surfaces. Lecture Notes in Mathematics, 971:109–117, 1983. CODEN LNMAA2. ISBN 3-540-11975-2 (print), 3-540-39426-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067076/.

Potra:1982:CCN

[Pot82] F.-A. Potra. On the convergence of a class of Newton-like methods. Lecture Notes in Mathematics, 953:125–137, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069378/.

Powell:1984:NMC

[Pow84] M. J. D. Powell. Nonconvex minimization calculations and the conjugate gradient method. Lecture Notes in Mathematics, 1066: 122–141, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099521/.

Poyatos:1983:PSS

[Poy83] Francisco Poyatos. G-principal series of Stocks in an algebra. Lecture Notes in Mathematics, 1004:240–244, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063441/.

Perrin:1981:PGC

[PP81] D. Perrin and J.-F. Perrot. À propos des groupes dans certains monoïdes syntactiques. (French) []. Lecture Notes in Mathematics, 855:82–91, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089811/.

Pagallo:1982:SMS

[PP82a] G. Pagallo and V. Pereyra. Smooth monotone spline interpolation. Lecture Notes in Mathematics, 909:142–146, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092968/.

Pogrebkov:1982:STT

[PP82b] A. K. Pogrebkov and M. C. Polivanov. Some topics in the theory of singular solutions of nonlinear equations. Lecture Notes in Mathematics, 970:129–145, 1982. CODEN LNMAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066028/.

Pietsch:1983:BST

[PPS83] Albrecht Pietsch, Nicolae Popa, and Ivan Singer, editors. Banach Space Theory and its Applications: Proceedings of the First Romanian–GDR Seminar Held at Bucharest, Romania, August 31–September 6, 1981, volume 991 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 991. URL http:

//link.springer.com/book/10.1007/BFb0061552;   http:/
/www.springerlink.com/content/978-3-540-39877-6.

Petridis:1980:HAI

[PPV80] Nicholas Petridis, Stylianos K. Pichorides, and Nicolas Varopoulos, editors. Harmonic Analysis Iraklion 1978: Proceedings of a Conference Held at the University of Crete, Iraklion, Greece, July 1978, volume 781 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 781. URL http:/

/link.springer.com/book/10.1007/BFb0097640;   http://
www.springerlink.com/content/978-3-540-38632-2.

Pelletier:1982:SBS

[PR82] Joan Wick Pelletier and Robert Rosebrugh. On subcategories of Banach spaces in sheaves. Lecture Notes in Mathematics, 915: 260–274, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092885/.

Pereyra:1983:NMP

[PR83] Victor Pereyra and Alfonso Reinoza, editors. Numerical Methods: Proceedings of the International Workshop Held at Caracas, June 14–18, 1982, volume 1005 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1005; QA3 .L35 v.1005, QA 297 N86 1982. URL http:/

/link.springer.com/book/10.1007/BFb0112521;   http://
www.springerlink.com/content/978-3-540-40967-0.

Prestel:1984:FAFa

[PR84a] Alexander Prestel and Peter Roquette. Formally p-adic fields. Lecture Notes in Mathematics, 1050:92–121, 1984. CODEN LN-MAA2. ISBN 3-540-12890-5 (print), 3-540-38768-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071467/.

Prestel:1984:FAFb

[PR84b] Alexander Prestel and Peter Roquette. Formally p-adic Fields, volume 1050 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12890-5 (print), 3-540-38768-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 167 pp. LCCN QA3 .L28 no. 1050. URL http:

//link.springer.com/book/10.1007/BFb0071461;   http:/
/www.springerlink.com/content/978-3-540-38768-8.

Prestel:1984:FFA

[PR84c] Alexander Prestel and Peter Roquette. Function fields over padically closed fields. Lecture Notes in Mathematics, 1050:122– 152, 1984. CODEN LNMAA2. ISBN 3-540-12890-5 (print), 3-540-38768-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071468/.

Prestel:1984:GET

[PR84d] Alexander Prestel and Peter Roquette. The general embedding theorem. Lecture Notes in Mathematics, 1050:62–82, 1984. CO-DEN LNMAA2. ISBN 3-540-12890-5 (print), 3-540-38768-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071465/.

Prestel:1984:IM

[PR84e] Alexander Prestel and Peter Roquette. Introduction and motivation. Lecture Notes in Mathematics, 1050:1–11, 1984. CODEN LNMAA2. ISBN 3-540-12890-5 (print), 3-540-38768-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071462/.

Prestel:1984:MTA

[PR84f] Alexander Prestel and Peter Roquette. Model theory of p-adically closed fields. Lecture Notes in Mathematics, 1050:83–91, 1984. CODEN LNMAA2. ISBN 3-540-12890-5 (print), 3-540-38768-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071466/.

Prestel:1984:ACF

[PR84g] Alexander Prestel and Peter Roquette. p-adically closed fields. Lecture Notes in Mathematics, 1050:33–61, 1984. CODEN LN-MAA2. ISBN 3-540-12890-5 (print), 3-540-38768-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071464/.

Prestel:1984:V

[PR84h] Alexander Prestel and Peter Roquette. p-valuations. Lecture Notes in Mathematics, 1050:12–32, 1984. CODEN LN-MAA2. ISBN 3-540-12890-5 (print), 3-540-38768-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071463/.

Praeger:1980:SGC

[Pra80] Cheryl E. Praeger. Symmetric graphs and a characterization of the odd graphs. Lecture Notes in Mathematics, 829:211–219, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088914/.

Praeger:1982:WSG

[Pra82a] Cheryl E. Praeger. When are symmetric graphs characterised by their local properties? Lecture Notes in Mathematics, 952: 123–141, 1982. CODEN LNMAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061976/.

Prasad:1982:SHM

[Pra82b] V. S. Prasad. A survey of homeomorphic measures. Lecture Notes in Mathematics, 945:150–154, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096669/.

Pratelli:1983:MDT

[Pra83] Maurizio Pratelli. Majorations dans L^{p} du type Metivier– Pellaumail pour les semimartingales. (French) []. Lecture Notes in Mathematics, 986:125–131, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068308/.

Praeger:1984:SGC

[Pra84] Cheryl E. Praeger. Symmetric graphs and the classification of the finite simple groups. Lecture Notes in Mathematics, 1098: 99–110, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099667/.

Prestel:1981:PRC

[Pre81] Alexander Prestel. Pseudo real closed fields. Lecture Notes in Mathematics, 872:127–156, 1981. CODEN LNMAA2. ISBN 3-540-10849-1 (print), 3-540-38757-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098621/.

Preuss:1982:CDN

[Pre82] Gerhard Preuss. Connectednesses and disconnectednesses in Snear. Lecture Notes in Mathematics, 915:275–292, 1982. CODEN LNMAA2. ISBN 3-540-11211-1 (print), 3-540-39041-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092886/.

Preston:1983:AAV

[Pre83a] Chris Preston. An “almost all” version of theorem 5.1. Lecture Notes in Mathematics, 999:165–180, 1983. CODEN LN-MAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061758/.

Preston:1983:GRH

[Pre83b] Chris Preston. Getting rid of homtervals. Lecture Notes in Mathematics, 999:143–156, 1983. CODEN LNMAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061756/.

Preston:1983:I

[Pre83c] Chris Preston. Introduction. Lecture Notes in Mathematics, 999: 1–18, 1983. CODEN LNMAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061750/.

Preston:1983:IF

[Pre83d] Chris Preston. The iterates of functions in S. Lecture Notes in Mathematics, 999:82–108, 1983. CODEN LNMAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061754/.

Preston:1983:IMI

[Pre83e] Chris Preston. Iterates of Maps on an Interval, volume 999 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 205 pp. LCCN QA3 .L28 no. 999. URL http:/

/link.springer.com/book/10.1007/BFb0061749;   http://
www.springerlink.com/content/978-3-540-40055-4.

Preston:1983:KS

[Pre83f] Chris Preston. Kneading sequences. Lecture Notes in Mathematics, 999:157–164, 1983. CODEN LNMAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061757/.

Preston:1983:ODT

[Pre83g] Chris Preston. Occurrence of the different types of behaviour. Lecture Notes in Mathematics, 999:181–197, 1983. CODEN LN-MAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061759/.

Preston:1983:PMF

[Pre83h] Chris Preston. Piecewise monotone functions. Lecture Notes in Mathematics, 999:19–40, 1983. CODEN LNMAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061751/.

Preston:1983:PNS

[Pre83i] Chris Preston. Property R and negative Schwarzian derivatives. Lecture Notes in Mathematics, 999:60–81, 1983. CODEN LNMAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061753/.

Preston:1983:R

[Pre83j] Chris Preston. Reductions. Lecture Notes in Mathematics, 999: 109–142, 1983. CODEN LNMAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061755/.

Preston:1983:WBP

[Pre83k] Chris Preston. Well-behaved piecewise monotone functions. Lecture Notes in Mathematics, 999:41–59, 1983. CODEN LN-MAA2. ISBN 3-540-12322-9 (print), 3-540-40055-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061752/.

Prestel:1984:LFR

[Pre84a] Alexander Prestel. Lectures on Formally Real Fields, volume 1093 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 125 pp. LCCN QA3 .L28 no. 1093. URL http:/

/link.springer.com/book/10.1007/BFb0101548;   http://
www.springerlink.com/content/978-3-540-39093-0.

Prestel:1984:OSF

[Pre84b] Alexander Prestel. Orderings and semiorderings of fields. Lecture Notes in Mathematics, 1093:1–14, 1984. CODEN LN-MAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101549/.

Prestel:1984:QFFa

[Pre84c] Alexander Prestel. Quadratic forms over formally real fields. Lecture Notes in Mathematics, 1093:15–24, 1984. CODEN LN-MAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101550/.

Prestel:1984:QFFb

[Pre84d] Alexander Prestel. Quadratic forms over formally real fields. Lecture Notes in Mathematics, 1093:109–118, 1984. CODEN LNMAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101558/.

Prestel:1984:RAC

[Pre84e] Alexander Prestel. Real algebraic closures. Lecture Notes in Mathematics, 1093:25–33, 1984. CODEN LNMAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101551/.

Prestel:1984:RHF

[Pre84f] Alexander Prestel. Real Henselian fields. Lecture Notes in Mathematics, 1093:85–96, 1984. CODEN LNMAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101556/.

Prestel:1984:RP

[Pre84g] Alexander Prestel. Real places. Lecture Notes in Mathematics, 1093:70–84, 1984. CODEN LNMAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101555/.

Prestel:1984:SF

[Pre84h] Alexander Prestel. SAP-fields. Lecture Notes in Mathematics, 1093:97–108, 1984. CODEN LNMAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101557/.

Prestel:1984:SNM

[Pre84i] Alexander Prestel. Some notions from model theory. Lecture Notes in Mathematics, 1093:34–48, 1984. CODEN LN-MAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101552/.

Prestel:1984:SOS

[Pre84j] Alexander Prestel. The space of orderings and semiorderings. Lecture Notes in Mathematics, 1093:62–69, 1984. CODEN LN-MAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101554/.

Prestel:1984:TPR

[Pre84k] Alexander Prestel. The transfer-principle for real closed fields. Lecture Notes in Mathematics, 1093:49–61, 1984. CODEN LN-MAA2. ISBN 3-540-13885-4 (print), 3-540-39093-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101553/.

Priddy:1981:CC

[Pri81] Stewart Priddy. On a conjecture concerning K_{∗}(ZZ/p^{2}). Lecture Notes in Mathematics, 854:338–342, 1981. CODEN LN-MAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089528/.

Priouret:1982:RPP

[Pri82] Pierre Priouret. Remarques sur les petites perturbations de systèmes dynamiques. (French) []. Lecture Notes in Mathematics, 920:184–200, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092780/.

Price:1984:UFB

[Pri84] Gareth C. Price. The unique factorisation of Brownian products. Lecture Notes in Mathematics, 1095:166–168, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099128/.

Prochazka:1981:BST

[Pro81] Ladislav Procházka. ^{∞} p-basic subgroups of torsion free abelian groups. Lecture Notes in Mathematics, 874:127–153, 1981. CO-DEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090531/.

Price:1984:AI

[PRW84] Gareth C. Price, L. C. G. Rogers, and David Williams. BM(R^{3}) and its area integral ∫ β × dβ. Lecture Notes in Mathematics, 1095:155–165, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099127/.

Peitgen:1981:PSS

[PS81] H.-O. Peitgen and K. Schmitt. Positive and spurious solutions of nonlinear eigenvalue problems. Lecture Notes in Mathematics, 878:275–324, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090686/.

Piatetski-Shapiro:1983:WW

[PS83a] Ilya Piatetski-Shapiro. Work of Waldspurger. Lecture Notes in Mathematics, 1041:280–302, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073151/.

Protter:1983:EIL

[PS83b] Philip Protter and Alain-Sol Sznitman. An equation involving local time. Lecture Notes in Mathematics, 986:62–66, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068298/.

Pidkuiko:1984:HSD

[PS84a] S. I. Pidkuiko and A. M. Stepin. On Hamiltonian systems with dynamical symmetries. Lecture Notes in Mathematics, 1108:152– 172, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099556/.

Pourahmadi:1984:SLT

[PS84b] M. Pourahmadi and H. Salehi. On subordination and linear transformation of harmonizable and periodically correlated processes. Lecture Notes in Mathematics, 1080:195–213, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099797/.

Pritchard:1984:LQO

[PS84c] A. J. Pritchard and D. Salamon. The linear quadratic optimal control problem for infinite dimensional systems with unbounded input and output operators. Lecture Notes in Mathematics, 1076: 187–202, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072776/.

Preiss:1982:DMH

[PT82] David Preiss and Jaroslav Tišer. Differentiation of measures on Hilbert spaces. Lecture Notes in Mathematics, 945:194–207, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096675/.

Pullman:1983:CCG

[Pul83] Norman J. Pullman. Clique coverings of graphs — a survey. Lecture Notes in Mathematics, 1036:72–85, 1983. CODEN LN-MAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071509/.

Patwardhan:1981:ASB

[PV81] G. A. Patwardhan and M. N. Vartak. On the adjugate of a symmetrical balanced incomplete block design with λ = 1. Lecture Notes in Mathematics, 885:133–152, 1981. CODEN LN-MAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092258/.

Pierce:1983:RANa

[PV83a] R. S. Pierce and C. I. Vinsonhaler. Realizing algebraic number fields. Lecture Notes in Mathematics, 1006:49–96, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21560-9_2/.

Pierce:1983:RANb

[PV83b] R. S. Pierce and C. I. Vinsonhaler. Realizing algebraic number fields. Lecture Notes in Mathematics, 1006:49–96, 1983. CO-DEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103697/.

Potempa:1984:IDN

[PW84] Thom C. Potempa and Mary Fanett Wheeler. An implicit diffusive numerical procedure for a slightly compressible miscible displacement problem in porous media. Lecture Notes in Mathematics, 1066:239–263, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099528/.

Pacholski:1980:MTA

[PWW80] Leszek Pacholski, Jedrzej Wierzejewski, and Alec J. Wilkie, editors. Model Theory of Algebra and Arithmetic: Proceedings of the Conference on Applications of Logic to Algebra and Arithmetic Held at Karpacz, Poland, September 1–7, 1979, volume 834 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 834. URL http:/

/link.springer.com/book/10.1007/BFb0090155;   http://
www.springerlink.com/content/978-3-540-38393-2.

Pitman:1981:BPI

[PY81] Jim Pitman and Marc Yor. Bessel processes and infinitely divisible laws. Lecture Notes in Mathematics, 851:285–370, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088732/.

Pitman:1982:DPB

[PY82] Jim Pitman and Marc Yor. Sur une décomposition des ponts de Bessel. (French) []. Lecture Notes in Mathematics, 923:276– 285, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093048/.

Queyrut:1981:GGS

[Que81] Jacques Queyrut. S-groupes de Grothendieck et structure galoisienne des anneaux d’entiers. (French) []. Lecture Notes in Mathematics, 882:219–239, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092495/.

Queyrut:1984:SGF

[Que84] Jacques Queyrut. Structures galoisiennes. (French) [Galois structures]. Lecture Notes in Mathematics, 1068:198–209, 1984. CO-DEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099453/.

Quintas:1984:RSR

[Qui84] Louis V. Quintas. Random subgraphs of regular graphs. Lecture Notes in Mathematics, 1073:150–160, 1984. CODEN LN-MAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073113/.

Rabinowitz:1982:MPT

[Rab82] Paul H. Rabinowitz. The mountain pass theorem: Theme and variations. Lecture Notes in Mathematics, 957:237–271, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066242/.

Raby:1983:TZS

[Rab83] Gilles Raby. Théorème des zéros sous-analytiques et inégalités de łojasiewicz. (French) []. Lecture Notes in Mathematics, 1028: 253–265, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071686/.

Rabitz:1984:SMM

[Rab84] Herschel Rabitz. Sensitivity methods for mathematical modelling. Lecture Notes in Mathematics, 1086:77–92, 1984. CODEN LNMAA2. ISBN 3-540-13871-4 (print), 3-540-39061-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073070/.

Rachev:1983:MMR

[Rac83] S. T. Rachev. Minimal metrics in the real random variables space. Lecture Notes in Mathematics, 982:172–190, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082069/.

Rafel:1983:ACG

[Raf83] G. G. Rafel. Application of a combined Galerkin-averaging method. Lecture Notes in Mathematics, 985:349–369, 1983. CO-DEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062376/.

Ramachandra:1982:AMP

[Ram82a] K. Ramachandra. Addendum to my paper a brief summary of some results in the analytic theory of numbers. Lecture Notes in Mathematics, 938:119–122, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097179/.

Ramachandra:1982:BSS

[Ram82b] K. Ramachandra. A brief summary of some results in the analytic theory of numbers — II. Lecture Notes in Mathematics, 938: 106–118, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097178/.

Randell:1980:STZ

[Ran80] Richard Randell. Some topology of Zariski surfaces. Lecture Notes in Mathematics, 788:145–164, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099244/.

Range:1982:BRC

[Ran82] R. Michael Range. Boundary regularity for the Cauchy–Riemann complex. Lecture Notes in Mathematics, 919:177–186, 1982. CO-DEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097049/.

Rangaswamy:1983:SBSa

[Ran83a] K. M. Rangaswamy. On strongly balanced subgroups of separable torsion-free Abelian groups. Lecture Notes in Mathematics, 1006:268–274, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21560-9_12/.

Rangaswamy:1983:SBSb

[Ran83b] K. M. Rangaswamy. On strongly balanced subgroups of separable torsion-free abelian groups. Lecture Notes in Mathematics, 1006: 268–274, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103707/.

Raoult:1980:SNP

[Rao80] Jean-Pierre Raoult, editor. Statistique non Paramétrique Asymptotique: Actes des Journées Statistiques, Rouen, France, Juin 1979, volume 821 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10239-6 (print), 3-540-38318-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 821. URL http:/

/link.springer.com/book/10.1007/BFb0097421;   http://
www.springerlink.com/content/978-3-540-38318-5.

Rao:1981:CGT

[Rao81] Siddani Bhaskara Rao, editor. Combinatorics and Graph Theory: Proceedings of the Symposium Held at the Indian Statistical Institute, Calcutta, February 25–29, 1980, volume 885 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 885.

URL http://link.springer.com/book/10.1007/BFb0092249;
 http://www.springerlink.com/content/978-3-540-47037-
3.

Ratiu:1980:IT

[Raţ80] Tudor Raţiu. Involution theorems. Lecture Notes in Mathematics, 775:219–257, 1980. CODEN LNMAA2. ISBN 3-540-09742-2 (print), 3-540-38571-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092027/.

Rautmann:1980:AMN

[Rau80] Reimund Rautmann, editor. Approximation Methods for Navier– Stokes Problems: Proceedings of the Symposium Held by the International Union of Theoretical and Applied Mechanics (IU-TAM) at the University of Paderborn, Germany, September 9–15, 1979, volume 771 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09734-1 (print), 3-540-38550-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 771. URL http:/

/link.springer.com/book/10.1007/BFb0086897;   http://
www.springerlink.com/content/978-3-540-38550-9.

Raviart:1984:PAL

[Rav84] P.-A. Raviart. Particle approximation of linear hyperbolic equations of the first order. Lecture Notes in Mathematics, 1066:142– 158, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099522/.

Rawnsley:1980:NPP

[Raw80] J. H. Rawnsley. Non-positive polarizations and half-forms. Lecture Notes in Mathematics, 836:145–152, 1980. CODEN LN-MAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089730/.

Raynaud:1981:FVI

[Ray81] M. Raynaud. Fibrés vectoriels instables — applications aux surfaces (d’après Bogomolov). (French) []. Lecture Notes in Mathematics, 868:293–314, 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090653/.

Raynaud:1983:AMC

[Ray83] M. Raynaud. Around the Mordell conjecture for function fields and a conjecture of Serge Lang. Lecture Notes in Mathematics, 1016:1–19, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099956/.

Ray-Chaudhuri:1981:ATS

[RC81] Dijen K. Ray-Chaudhuri. Affine triple systems. Lecture Notes in Mathematics, 885:60–69, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092255/.

Rocha-Caridi:1982:RIH

[RC82] Alvany Rocha-Caridi. Resolutions of irreducible highest weight modules over infinite dimensional graded Lie algebras. Lecture Notes in Mathematics, 933:176–190, 1982. CODEN LN-MAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093361/.

Ruiz-Claeyssen:1980:AIA

[RCC80] Julio Ruíz-Claeyssen and Bernardo Cockburn. Applications of the integral averaging bifurcation method to retarded functional differential equations. Lecture Notes in Mathematics, 799:348– 367, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089324/.

Read:1981:SGG

[Rea81] Ronald C. Read. A survey of graph generation techniques. Lecture Notes in Mathematics, 884:77–89, 1981. CODEN LN-MAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091809/.

Rea:1983:FLO

[Rea83] Claudio Rea. Formes de Levi d’ordre supérieur et réduction des variables dans l’équation d’une hypersurface réelle. (French) []. Lecture Notes in Mathematics, 1014:68–77, 1983. CODEN LN-MAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072071/.

Redner:1983:DEP

[Red83a] Sidney Redner. Directionality effects in percolation. Lecture Notes in Mathematics, 1035:246–259, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073262/.

Redner:1983:RPC

[Red83b] Sidney Redner. Recent progress and current puzzles in percolation. Lecture Notes in Mathematics, 1035:184–200, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073260/.

Rehmann:1980:KG

[Reh80] U. Rehmann. Kommutatoren in GL_{n}(D). (German) []. Lecture Notes in Mathematics, 778:117–123, 1980. CODEN LN-MAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095928/.

Rehmann:1982:SCS

[Reh82] U. Rehmann. A survey of the congruence subgroup problem. Lecture Notes in Mathematics, 966:197–207, 1982. CODEN LN-MAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062176/.

Reid:1981:AGF

[Rei81] J. D. Reid. Abelian groups finitely generated over their endomorphism rings. Lecture Notes in Mathematics, 874:41–52, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090522/.

Reid:1982:PFE

[Rei82a] Walter M. Reid. Parameterizations and factorizations of element regions for continued fractions K(a_{n}/1). Lecture Notes in Mathematics, 932:206–224, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093315/.

Reinhardt:1982:AAF

[Rei82b] Hans-Jürgen Reinhardt. Analysis of adaptive finite element methods for −ϵU′′ + U′ = F based on a-posteriori error estimates. Lecture Notes in Mathematics, 942:207–227, 1982. CODEN LN-MAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094749/.

Reiten:1982:UAS

[Rei82c] Idun Reiten. The use of almost split sequences in the representation theory of Artin algebras. Lecture Notes in Mathematics, 944: 29–104, 1982. CODEN LNMAA2. ISBN 3-540-11577-3 (print), 3-540-39313-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094057/.

Reid:1983:AGCa

[Rei83a] J. D. Reid. Abelian groups cyclic over their endomorphism rings. Lecture Notes in Mathematics, 1006:190–203, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21560-9_
7/.

Reid:1983:AGCb

[Rei83b] J. D. Reid. Abelian groups cyclic over their endomorphism rings. Lecture Notes in Mathematics, 1006:190–203, 1983. CO-DEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103702/.

Reich:1984:DNS

[Rei84a] S. Reich. On the differentiability of nonlinear semigroups. Lecture Notes in Mathematics, 1076:203–208, 1984. CODEN LN-MAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072777/.

Reid:1984:MRC

[Rei84b] K. B. Reid. Monochromatic reachability, complementary cycles, and single arc reversals in tournaments. Lecture Notes in Mathematics, 1073:11–21, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073101/.

Remlinger:1982:GBD

[Rem82] Jörg Remlinger. Generalized block designs as approximations for optimal coverings. Lecture Notes in Mathematics, 969:293– 301, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063001/.

Renault:1980:AG

[Ren80a] Jean Renault. The C^{∗}-algebra of a groupoid. Lecture Notes in Mathematics, 793:47–120, 1980. CODEN LNMAA2. ISBN 3-540-09977-8 (print), 3-540-39218-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091075/.

Renault:1980:GAA

[Ren80b] Jean Renault. A Groupoid Approach to C^{∗}-Algebras, volume 793 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09977-8 (print), 3-540-39218-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 160 pp. LCCN QA3 .L28 no. 793. URL http:/

/link.springer.com/book/10.1007/BFb0091072;   http://
www.springerlink.com/content/978-3-540-39218-7.

Renault:1980:I

[Ren80c] Jean Renault. Introduction. Lecture Notes in Mathematics, 793: 1–4, 1980. CODEN LNMAA2. ISBN 3-540-09977-8 (print), 3-540-39218-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091073/.

Renault:1980:LCG

[Ren80d] Jean Renault. Locally compact groupoids. Lecture Notes in Mathematics, 793:5–46, 1980. CODEN LNMAA2. ISBN 3-540-09977-8 (print), 3-540-39218-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091074/.

Renault:1980:SE

[Ren80e] Jean Renault. Some examples. Lecture Notes in Mathematics, 793:121–147, 1980. CODEN LNMAA2. ISBN 3-540-09977-8 (print), 3-540-39218-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091076/.

Revesz:1981:LTI

[Rév81] P. Révész. Local time and invariance. Lecture Notes in Mathematics, 861:128–145, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097320/.

Reyn:1980:GLC

[Rey80] J. W. Reyn. Generation of limit cycles from separatrix polygons in the phase plane. Lecture Notes in Mathematics, 810: 264–289, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089983/.

Romero-Fuster:1984:GEG

[RF84] M. C. Romero-Fuster. Generic embeddings, Gauss maps and stratipications. Lecture Notes in Mathematics, 1045:164–181, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072177/.

Rhemtulla:1981:PRO

[Rhe81] A. H. Rhemtulla. Polycyclic right-ordered groups. Lecture Notes in Mathematics, 848:230–234, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090569/.

Richard:1981:SAS

[Ric81a] Denis Richard. De la structure additive à la saturation des modèles de Peano et à une classification des sous-langages de l’arithmétique. (French) []. Lecture Notes in Mathematics, 890: 270–296, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095669/.

Richardson:1981:ASC

[Ric81b] R. W. Richardson. An application of the Serre conjecture to semisimple algebraic groups. Lecture Notes in Mathematics, 848: 141–151, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090561/.

Richman:1981:CMP

[Ric81c] Fred Richman, editor. Constructive Mathematics: Proceedings of the New Mexico State University Conference Held at Las Cruces, New Mexico, August 11–15, 1980, volume 873 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 873.

URL http://link.springer.com/book/10.1007/BFb0090721;
 http://www.springerlink.com/content/978-3-540-38759-
6.

Richman:1981:SC

[Ric81d] Fred Richman. Seidenberg’s condition P. Lecture Notes in Mathematics, 873:1–11, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090722/.

Ricci:1982:CZK

[Ric82] Fulvio Ricci. Calderón–Zygmund kernels on nilpotent Lie groups. Lecture Notes in Mathematics, 908:217–227, 1982. CODEN LN-MAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093291/.

Rickman:1983:VDQ

[Ric83] Seppo Rickman. Value distribution of quasiregular mappings. Lecture Notes in Mathematics, 981:220–245, 1983. CODEN LN-MAA2. ISBN 3-540-12003-3 (print), 3-540-39480-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066386/.

Ringel:1980:RBT

[Rin80a] Claus Michael Ringel. Report on the Brauer–Thrall conjectures: Rojter’s theorem and the theorem of Nazarova and Rojter (on algorithms for solving vector-space problems. I). Lecture Notes in Mathematics, 831:104–136, 1980. CODEN LNMAA2. ISBN 3-540-10263-9 (print), 3-540-38385-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089780/.

Ringel:1980:TAA

[Rin80b] Claus Michael Ringel. Tame algebras (on algorithms for solving vector space problems. II). Lecture Notes in Mathematics, 831: 137–287, 1980. CODEN LNMAA2. ISBN 3-540-10263-9 (print), 3-540-38385-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089781/.

Ringel:1983:STS

[Rin83] Claus Michael Ringel. Separating tubular series. Lecture Notes in Mathematics, 1029:134–158, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098930/.

Ringel:1984:CSS

[Rin84a] Claus Michael Ringel. Construction of stable separating tubular families. Lecture Notes in Mathematics, 1099:113–166, 1984. CODEN LNMAA2. ISBN 3-540-13905-2 (print), 3-540-39127-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072873/.

Ringel:1984:DA

[Rin84b] Claus Michael Ringel. Directed algebras. Lecture Notes in Mathematics, 1099:324–361, 1984. CODEN LNMAA2. ISBN 3-540-13905-2 (print), 3-540-39127-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072876/.

Ringel:1984:IQF

[Rin84c] Claus Michael Ringel. Integral quadratic forms. Lecture Notes in Mathematics, 1099:1–40, 1984. CODEN LNMAA2. ISBN 3-540-13905-2 (print), 3-540-39127-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072871/.

Ringel:1984:QMC

[Rin84d] Claus Michael Ringel. Quivers, module categories, subspace categories (notation, results, some proofs). Lecture Notes in Mathematics, 1099:41–112, 1984. CODEN LNMAA2. ISBN 3-540-13905-2 (print), 3-540-39127-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072872/.

Ringel:1984:TAI

[Rin84e] Claus Michael Ringel. Tame Algebras and Integral Quadratic Forms, volume 1099 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13905-2 (print), 3-540-39127-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 376 pp. LCCN QA3 .L28 no. 1099. URL http:/

/link.springer.com/book/10.1007/BFb0072870;   http://
www.springerlink.com/content/978-3-540-39127-2.

Ringel:1984:TFT

[Rin84f] Claus Michael Ringel. Tilting functors and tubular extensions (notation, results, some proofs). Lecture Notes in Mathematics, 1099:167–267, 1984. CODEN LNMAA2. ISBN 3-540-13905-2 (print), 3-540-39127-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072874/.

Ringel:1984:TA

[Rin84g] Claus Michael Ringel. Tubular algebras. Lecture Notes in Mathematics, 1099:268–323, 1984. CODEN LNMAA2. ISBN 3-540-13905-2 (print), 3-540-39127-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072875/.

Risler:1982:HSA

[Ris82] J.-J. Risler. Sur l’homologie des surfaces algèbriques réelles. (French) []. Lecture Notes in Mathematics, 959:381–385, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062265/.

Ritter:1981:CGF

[Rit81] Jürgen Ritter. The class group à la Fröhlich. Lecture Notes in Mathematics, 882:174–190, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092492/.

Rivlin:1984:SUC

[Riv84] T. J. Rivlin. The strong uniqueness constant in complex approximation. Lecture Notes in Mathematics, 1105:145–149, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072407/.

Rizza:1984:NTF

[Riz84] G. B. Rizza. Nijenhuis tensor field and weakly Kàhler manifolds. Lecture Notes in Mathematics, 1045:154–163, 1984. CO-DEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072176/.

Robbin:1981:AKS

[Rob81a] J. W. Robbin. Algebraic Kupka–Smale theory. Lecture Notes in Mathematics, 898:286–301, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091920/.

Robinson:1981:DSF

[Rob81b] Clark Robinson. Differentiability of the stable foliation for the model Lorenz equations. Lecture Notes in Mathematics, 898:302– 315, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091921/.

Robinson:1981:RRF

[Rob81c] Derek J. S. Robinson. Recent results on finite complete groups. Lecture Notes in Mathematics, 848:178–185, 1981. CODEN LN-MAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090565/.

Roberts:1983:MSQ

[Rob83a] John E. Roberts. Mathematical structure in quantum field theory. Lecture Notes in Mathematics, 1031:89–96, 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072112/.

Robinson:1983:LSF

[Rob83b] Stephen M. Robinson. Local structure of feasible sets in nonlinear programming, Part I: Regularity. Lecture Notes in Mathematics, 1005:240–251, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112538/.

Robbiano:1984:FAF

[Rob84a] Lorenzo Robbiano. Factorial and almost factorial schemes in weighted projective spaces. Lecture Notes in Mathematics, 1092: 62–84, 1984. CODEN LNMAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099357/.

Robinson:1984:DPI

[Rob84b] Derek J. S. Robinson. Decision problems for infinite soluble groups. Lecture Notes in Mathematics, 1098:111–117, 1984. CO-DEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099668/.

Roczen:1984:SPC

[Roc84] M. Roczen. Some properties of the canonical resolutions of the 3dimensional singularities A_{n}, D_{n}, E_{n} over a field of characteristic ≠ 2. Lecture Notes in Mathematics, 1056:297–365, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071777/.

Rodier:1981:DSP

[Rod81] François Rodier. Décomposition de la série principale des groupes réductifs p-adiques. (French) []. Lecture Notes in Mathematics, 880:408–424, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090418/.

Rodding:1984:SLP

[Röd84] Dieter Rödding. Some logical problems connected with a modular decomposition theory of automata. Lecture Notes in Mathematics, 1104:365–388, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099492/.

Roggenkamp:1980:LTO

[Rog80] K. W. Roggenkamp. The lattice type of orders: a diagrammatic approach. I. Lecture Notes in Mathematics, 825:104–129, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089126/.

Rogers:1981:SIB

[Rog81a] L. C. G. Rogers. Stochastic integrals: Basic theory. Lecture Notes in Mathematics, 851:56–71, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088722/.

Rogers:1981:WCB

[Rog81b] L. C. G. Rogers. Williams’ characterisation of the Brownian excursion law: proof and applications. Lecture Notes in Mathematics, 850:227–250, 1981. CODEN LNMAA2. ISBN 3-540-10689-8 (print), 3-540-38610-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088371/.

Roggenkamp:1981:SIG

[Rog81c] K. W. Roggenkamp. Structure of integral group rings. Lecture Notes in Mathematics, 867:421–440, 1981. CODEN LN-MAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090397/.

Roggenkamp:1981:IRA

[Rog81d] Klaus W. Roggenkamp, editor. Integral Representations and Applications: Proceedings of a Conference held at Oberwolfach, Germany, June 22–28, 1980, volume 882 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 882. URL http:

//link.springer.com/book/10.1007/BFb0092485;   http:/
/www.springerlink.com/content/978-3-540-38789-3.

Rogers:1984:BLT

[Rog84a] L. C. G. Rogers. Brownian local times and branching processes. Lecture Notes in Mathematics, 1059:42–55, 1984. CODEN LN-MAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100030/.

Roggenkamp:1984:AII

[Rog84b] K. W. Roggenkamp. Automorphisms and isomorphisms of integral group rings of finite groups. Lecture Notes in Mathematics, 1098:118–135, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099669/.

Rohatgi:1981:RCC

[Roh81] Vijay K. Rohatgi. On the rate of convergence in the central limit theorem. Lecture Notes in Mathematics, 861:146–154, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097321/.

Rohl:1984:IIG

[Röh84] Frank Röhl. On induced isomorphisms of group rings. Lecture Notes in Mathematics, 1098:136–141, 1984. CODEN LN-MAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099670/.

Rojter:1980:MPR

[Roj80] A. V. Rojter. Matrix problems and representations of Bocs’s. Lecture Notes in Mathematics, 831:288–324, 1980. CODEN LN-MAA2. ISBN 3-540-10263-9 (print), 3-540-38385-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089782/.

Rolland:1982:ECG

[Rol82] Raymond Rolland. Étude des coupures dans les groupes et corps ordonnéas. (French) []. Lecture Notes in Mathematics, 959:386– 405, 1982. CODEN LNMAA2. ISBN 3-540-11959-0 (print), 3-540-39548-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062266/.

Romanowska:1983:AFP

[Rom83] Anna Romanowska. Algebras of functions from partially ordered sets into distributive lattices. Lecture Notes in Mathematics, 1004: 245–256, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063442/.

Roque:1982:ECA

[Roq82] J. Louis Roque. Étude d’une classe d’algèbres artiniennes locales non commutatives. (French) []. Lecture Notes in Mathematics, 924:241–259, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092935/.

Rosoff:1981:MED

[Ros81a] Jeffrey A. Rosoff. The monoid of effective divisor classes on a complex torus. Lecture Notes in Mathematics, 862:208–231, 1981. CODEN LNMAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090893/.

Rosset:1981:GRV

[Ros81b] Shmuel Rosset. The Goldie rank of virtually polycyclic groups. Lecture Notes in Mathematics, 844:35–45, 1981. CODEN LN-MAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090476/.

Rosinski:1982:CLT

[Ros82] Jan Rosiński. Central limit theorems for dependent random vectors in Banach spaces. Lecture Notes in Mathematics, 939:157– 180, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096267/.

Rosenberg:1983:GTP

[Ros83a] I. G. Rosenberg. Galois theory for partial algebras. Lecture Notes in Mathematics, 1004:257–272, 1983. CODEN LNMAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063443/.

Rosenthal:1983:SE

[Ros83b] H. P. Rosenthal. Sign-embeddings of L^{1}. Lecture Notes in Mathematics, 995:155–165, 1983. CODEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061892/.

Rost:1983:HGD

[Ros83c] H. Rost. Hydrodynamik gekoppelter diffusionen: Fluktuationen im gleichgewicht. Lecture Notes in Mathematics, 1031:97– 107, 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072113/.

Ros:1984:SGS

[Ros84a] A. Ros. Spectral geometry of submanifolds in the complex projective space. Lecture Notes in Mathematics, 1045:182–185, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072178/.

Roseau:1984:SRM

[Ros84b] M. Roseau. Stabilité de régime des machines tournantes et problèmes associés. (French) []. Lecture Notes in Mathematics, 1107:193–214, 1984. CODEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101501/.

Rosenberger:1984:DEU

[Ros84c] Gerhard Rosenberger. Über Darstellungen von Elementen und Untergruppen in freien Produkten. (German) []. Lecture Notes in Mathematics, 1098:142–160, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099671/.

Rossmann:1984:CCI

[Ros84d] W. Rossmann. Characters as contour integrals. Lecture Notes in Mathematics, 1077:375–388, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072345/.

Rost:1984:DSD

[Ros84e] Hermann Rost. Diffusion de sphères dures dans la droite rèelle: comportement macroscopique et èquilibre local. (French) []. Lecture Notes in Mathematics, 1059:127–143, 1984. CODEN LN-MAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100037/.

Rothe:1984:BND

[Rot84a] Franz Rothe. Basic notations and definitions. Lecture Notes in Mathematics, 1072:5–54, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099280/.

Rothe:1984:B

[Rot84b] Franz Rothe. The Brusselator. Lecture Notes in Mathematics, 1072:140–147, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099284/.

Rothe:1984:CR

[Rot84c] Franz Rothe. Chemical reactions. Lecture Notes in Mathematics, 1072:157–171, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099286/.

Rothe:1984:CTU

[Rot84d] Franz Rothe. Corollary of theorem 1 (uniqueness and maximality). Lecture Notes in Mathematics, 1072:54–103, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099281/.

Rothe:1984:FNS

[Rot84e] Franz Rothe. The FitzHugh–Nagumo system. Lecture Notes in Mathematics, 1072:148–156, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099285/.

Rothe:1984:GMM

[Rot84f] Franz Rothe. The Gierer-Meinhardt model. Lecture Notes in Mathematics, 1072:126–139, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099283/.

Rothe:1984:GSR

[Rot84g] Franz Rothe. Global Solutions of Reaction–Diffusion Systems, volume 1072 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 216 pp. LCCN QA3 .L28 no. 1072. URL http:

//link.springer.com/book/10.1007/BFb0099278;   http:/
/www.springerlink.com/content/978-3-540-38917-0.

Rothe:1984:I

[Rot84h] Franz Rothe. Introduction. Lecture Notes in Mathematics, 1072: 1–4, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099279/.

Rothe:1984:NRM

[Rot84i] Franz Rothe. A nuclear reactor model. Lecture Notes in Mathematics, 1072:172–187, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099287/.

Rothe:1984:RST

[Rot84j] Franz Rothe. Review of standard theorems. Lecture Notes in Mathematics, 1072:104–125, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099282/.

Rothe:1984:VLM

[Rot84k] Franz Rothe. The Volterra–Lotka model. Lecture Notes in Mathematics, 1072:188–210, 1984. CODEN LNMAA2. ISBN 3-540-13365-8 (print), 3-540-38917-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099288/.

Rothschild:1984:ASP

[Rot84l] Linda Preiss Rothschild. Analyticity of solutions of partial differential equations on nilpotent Lie groups. Lecture Notes in Mathematics, 1077:389–395, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072346/.

Rousseau:1981:IEV

[Rou81a] G. Rousseau. Instabilité dans les espaces vectoriels. (French) []. Lecture Notes in Mathematics, 868:263–276, 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090651/.

Rousseau:1981:IFV

[Rou81b] G. Rousseau. Instabilité dans les fibrés vectoriels (d’après Bogomolov). (French) []. Lecture Notes in Mathematics, 868:277– 292, 1981. CODEN LNMAA2. ISBN 3-540-10842-4 (print), 3-540-38742-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090652/.

Ray:1980:F

[RP80] Nigel Ray and Erik K. Pedersen. A fibration for DiffΣ^{n}. Lecture Notes in Mathematics, 788:165–171, 1980. CODEN LN-MAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099245/.

Raynaud-Pimienta:1984:LED

[RP84] Mme M. J. Raynaud-Pimienta. Lien entre (1) la différentiabilité en norme (2) la différentiabilité fine (3) les capacités Newtoniennes. (French) []. Lecture Notes in Mathematics, 1061:244– 270, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099028/.

Rychlik:1980:GAC

[RR80] E. Rychlik and Z. Rychlik. The generalized Anscombe condition and its applications in random limit theorems. Lecture Notes in Mathematics, 828:244–250, 1980. CODEN LNMAA2. ISBN 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097409/.

Rajput:1984:PSP

[RRM84] Balram S. Rajput and Kavi Rama-Murthy. Properties of semistable probability measures on R^{m}. Lecture Notes in Mathematics, 1080:214–232, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099798/.

Roberts:1981:SEG

[RS81] Joel Roberts and Robert Speiser. Schubert’s enumerative geometry of triangles from a modern viewpoint. Lecture Notes in Mathematics, 862:272–281, 1981. CODEN LNMAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090895/.

Rosinski:1982:PRM

[RS82a] Jan Rosiński and Jerzy Szulga. Product random measures and double stochastic integrals. Lecture Notes in Mathematics, 939: 181–199, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096268/.

Ruzsa:1982:IPD

[RS82b] Imre Z. Ruzsa and Gábor J. Székely. Irreducible and prime distributions. Lecture Notes in Mathematics, 928:354–361, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093233/.

Raynaud:1983:AGP

[RS83a] Michel Raynaud and Tetsuji Shioda, editors. Algebraic Geometry: Proceedings of the Japan–France Conference held at Tokyo and Kyoto, October 5–14, 1982, volume 1016 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1016.

URL http://link.springer.com/book/10.1007/BFb0099955;
 http://www.springerlink.com/content/978-3-540-38676-

Reich:1983:QMP

[RS83b] Edgar Reich and Kurt Strebel. Quasiconformal mappings of the punctured plane. Lecture Notes in Mathematics, 1013:182– 212, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066529/.

Richter:1983:TNA

[RS83c] M. M. Richter and M. E. Szabo. Towards a nonstandard analysis of programs. Lecture Notes in Mathematics, 983:186–203, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065340/.

Rychlik:1984:RCD

[RS84] Z. Rychlik and I. Szyszkowski. On the rate of convergence for distributions of integral type functionals. Lecture Notes in Mathematics, 1080:255–275, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099801/.

Robinson:1980:CMV

[RSW80] Robert W. Robinson, George W. Southern, and Walter D. Wallis, editors. Combinatorial Mathematics VII: Proceedings of the Seventh Australian Conference on Combinatorial Mathematics Held at the University of Newcastle, Australia, August 20–24, 1979, volume 829 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 829. URL http:/

/link.springer.com/book/10.1007/BFb0088895;   http://
www.springerlink.com/content/978-3-540-38376-5.

Riddle:1982:MFL

[RU82] Lawrence H. Riddle and J. J. Uhl, Jr. Martingales and the fine line between Asplund spaces and spaces not containing a copy of ℓ_{1}. Lecture Notes in Mathematics, 939:145–156, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096266/.

Ruckle:1982:ADS

[Ruc82] William H. Ruckle. Absolutely divergent series and Banach operator ideals. Lecture Notes in Mathematics, 939:200–210, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096269/.

Ruhe:1983:TSA

[Ruh83] Axel Ruhe. The two-sided Arnoldi algorithm for nonsymmetric eigenvalue problems. Lecture Notes in Mathematics, 973:104–120, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062097/.

Ruitenburg:1981:FE

[Rui81] W. Ruitenburg. Field extensions. Lecture Notes in Mathematics, 873:12–15, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090723/.

Ruiz:1983:ENP

[Ruí83] Luis Ruíz. Étude numérique d’un problème d’elasticité nonlinéaire. (French) []. Lecture Notes in Mathematics, 1005:266– 283, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112540/.

Ruppert:1983:SMR

[Rup83] Wolfgang Ruppert. On structural methods and results in the theory of compact semitopological semigroups. Lecture Notes in Mathematics, 998:215–238, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062031/.

Ruppert:1984:CSS

[Rup84a] Wolfgang Ruppert. Compact Semitopological Semigroups: an Intrinsic Theory, volume 1079 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13387-9 (print), 3-540-38938-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 260 pp. LCCN QA3 .L28 no. 1079.

URL http://link.springer.com/book/10.1007/BFb0073675;
 http://www.springerlink.com/content/978-3-540-38938-
5.

Ruppert:1984:EFC

[Rup84b] Wolfgang Ruppert. Elementary facts and constructions. Lecture Notes in Mathematics, 1079:8–46, 1984. CODEN LN-MAA2. ISBN 3-540-13387-9 (print), 3-540-38938-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073677/.

Ruppert:1984:I

[Rup84c] Wolfgang Ruppert. Introduction. Lecture Notes in Mathematics, 1079:1–7, 1984. CODEN LNMAA2. ISBN 3-540-13387-9 (print), 3-540-38938-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073676/.

Ruppert:1984:JC

[Rup84d] Wolfgang Ruppert. Joint continuity. Lecture Notes in Mathematics, 1079:47–100, 1984. CODEN LNMAA2. ISBN 3-540-13387-9 (print), 3-540-38938-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073678/.

Ruppert:1984:SCG

[Rup84e] Wolfgang Ruppert. Semigroup compactifications of groups. Lecture Notes in Mathematics, 1079:101–167, 1984. CODEN LN-MAA2. ISBN 3-540-13387-9 (print), 3-540-38938-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073679/.

Ruppert:1984:SLM

[Rup84f] Wolfgang Ruppert. Semigroups living in manifolds. Lecture Notes in Mathematics, 1079:168–231, 1984. CODEN LNMAA2. ISBN 3-540-13387-9 (print), 3-540-38938-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073680/.

Ruschendorf:1980:WCW

[Rüs80] Ludger Rüschendorf. Weak convergence of the weighted multiparameter empirical process. Lecture Notes in Mathematics, 821:

86–94, 1980. CODEN LNMAA2. ISBN 3-540-10239-6 (print), 3-540-38318-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097425/.

Russell:1982:DED

[Rus82a] Robert D. Russell. Difficulties in evaluating differential equation software. Lecture Notes in Mathematics, 909:175–184, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092972/.

Russell:1982:GCB

[Rus82b] Robert D. Russell. Global codes for BVODEs and their comparison. Lecture Notes in Mathematics, 968:256–268, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064893/.

Russek:1984:HEG

[Rus84a] Andrzej Russek. Hermite expansions of generalized Brownian functionals. Lecture Notes in Mathematics, 1080:233–243, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099799/.

Russell:1984:FFM

[Rus84b] Peter Russell. Factoring the Frobenius morphism of an algebraic surface. Lecture Notes in Mathematics, 1056:366–380, 1984. CODEN LNMAA2. ISBN 3-540-12930-8 (print), 3-540-38811-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071778/.

Ruymgaart:1980:UAA

[Ruy80] F. H. Ruymgaart. A unified approach to the asymptotic distribution theory of certain midrank statistics. Lecture Notes in Mathematics, 821:1–18, 1980. CODEN LNMAA2. ISBN 3-540-10239-6 (print), 3-540-38318-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097422/.

Ruzsa:1982:ICS

[Ruz82] Imre Z. Ruzsa. Infinite convolution and shift-convergence of measures on topological groups. Lecture Notes in Mathematics, 928: 337–353, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093232/.

Ruscheweyh:1984:MMN

[RV84] S. Ruscheweyh and R. S. Varga. On the minimum moduli of normalized polynomials. Lecture Notes in Mathematics, 1105: 150–159, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072408/.

Ricci:1982:HAP

[RW82] Fulvio Ricci and Guido Weiss, editors. Harmonic Analysis: Proceedings of a Conference Held at the University of Minnesota, Minneapolis, April 20–30, 1981, volume 908 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 908.

URL http://link.springer.com/book/10.1007/BFb0093277;
 http://www.springerlink.com/content/978-3-540-38973-
6.

Rockafellar:1983:DSP

[RW83] R. T. Rockafellar and Roger J.-B. Wets. A dual solution procedure for quadratic stochastic programs with simple recourse. Lecture Notes in Mathematics, 1005:252–265, 1983. CODEN LN-MAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112539/.

Rockafellar:1984:VSI

[RW84a] R. T. Rockafellar and Roger J.-B. Wets. Variational systems, an introduction. Lecture Notes in Mathematics, 1091:1–54, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098800/.

Rogers:1984:DEW

[RW84b] L. C. G. Rogers and David Williams. A differential equation in Wiener–Hopf theory. Lecture Notes in Mathematics, 1095:187– 199, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099130/.

Rooney:1984:MG

[RW84c] Joe Rooney and Robin J. Wilson. The mobility of a graph. Lecture Notes in Mathematics, 1073:135–149, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073112/.

Rand:1981:DST

[RY81] David Rand and Lai-Sang Young, editors. Dynamical Systems and Turbulence, Warwick 1980: Proceedings of a Symposium Held at the University of Warwick 1979/80, volume 898 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 898.

URL http://link.springer.com/book/10.1007/BFb0091903;
 http://www.springerlink.com/content/978-3-540-38945-
3.

Rychlik:1984:SCL

[Ryc84] Eugeniusz Rychlik. Some central limit theorems for randomly indexed sequences of random vectors. Lecture Notes in Mathematics, 1080:244–254, 1984. CODEN LNMAA2. ISBN 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099800/.

Rydel:1983:POQ

[Ryd83] A. Rydel. On Poisson output of queueing systems. Lecture Notes in Mathematics, 982:191–212, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082070/.

Saad:1983:PMS

[Saa83] Youcef Saad. Projection methods for solving large sparse eigenvalue problems. Lecture Notes in Mathematics, 973:121–144, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062098/.

Saad:1984:FGE

[Saa84a] Hedi Ben Saad. Fonction de Green sur un espace de Brelot. (French) []. Lecture Notes in Mathematics, 1061:40–53, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099016/.

Saad:1984:GNA

[Saa84b] Hedi Ben Saad. Généralisation des noyaux V_{h} et applications. (French) []. Lecture Notes in Mathematics, 1061:14–39, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099015/.

Sablonniere:1984:PTA

[Sab84] Paul Sablonnière. Padé-type approximants for multivariate series of functions. Lecture Notes in Mathematics, 1071:238–251, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099622/.

Sagraloff:1980:NSL

[Sag80] Boris Sagraloff. Normal solvability of linear partial differential operators in C^{∞}(ω). Lecture Notes in Mathematics, 810:290– 305, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089984/.

Sagher:1982:NIT

[Sag82] Y. Sagher. A new interpolation theorem. Lecture Notes in Mathematics, 908:189–198, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093287/.

Sahimi:1983:CET

[Sah83] Muhammad Sahimi. Critical exponents and thresholds for percolation and conduction. Lecture Notes in Mathematics, 1035: 314–346, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073266/.

Sakarovitch:1981:DMT

[Sak81] Jacques Sakarovitch. Description des monoïdes de type fini. (French) []. Lecture Notes in Mathematics, 855:92–115, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089812/.

Sakai:1982:A

[Sak82a] Makoto Sakai. Applications. Lecture Notes in Mathematics, 934: 100–125, 1982. CODEN LNMAA2. ISBN 3-540-11562-5 (print), 3-540-39257-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095896/.

Sakai:1982:CQD

[Sak82b] Makoto Sakai. Construction of quadrature domains. Lecture Notes in Mathematics, 934:4–47, 1982. CODEN LNMAA2. ISBN 3-540-11562-5 (print), 3-540-39257-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095894/.

Sakai:1982:E

[Sak82c] Makoto Sakai. Errata. Lecture Notes in Mathematics, 934:e1– e3, 1982. CODEN LNMAA2. ISBN 3-540-11562-5 (print), 3-540-39257-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095897/.

Sakai:1982:I

[Sak82d] Makoto Sakai. Introduction. Lecture Notes in Mathematics, 934: 1–3, 1982. CODEN LNMAA2. ISBN 3-540-11562-5 (print), 3-540-39257-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095893/.

Sakai:1982:PQD

[Sak82e] Makoto Sakai. Properties of quadrature domains. Lecture Notes in Mathematics, 934:48–99, 1982. CODEN LNMAA2. ISBN 3-540-11562-5 (print), 3-540-39257-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095895/.

Sakai:1982:QD

[Sak82f] Makoto Sakai. Quadrature Domains, volume 934 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11562-5 (print), 3-540-39257-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 133 pp. LCCN QA3 .L28 no. 934. URL http:/

/link.springer.com/book/10.1007/BFb0095892;   http://
www.springerlink.com/content/978-3-540-39257-6.

Salles:1980:DPS

[Sal80] Danielle Salles. Dualisation de la platitude systèmes projectifs de modules plats systèmes projectifs de modules injectifs. (French) []. Lecture Notes in Mathematics, 795:369–401, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090129/.

Sales:1982:DSG

[Sal82a] F. Puerta Sales. Déformations semiuniverselles et germes d’espaces analytiques C∗-équivariantes. (French) []. Lecture Notes in Mathematics, 961:267–274, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071287/.

Saltman:1982:GA

[Sal82b] David J. Saltman. Generic algebras. Lecture Notes in Mathematics, 917:96–117, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092230/.

Salamon:1983:TFD

[Sal83] Simon Salamon. Topics in four-dimensional Riemannian geometry. Lecture Notes in Mathematics, 1022:33–124, 1983. CODEN LNMAA2. ISBN 3-540-12719-4 (print), 3-540-38702-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071601/.

Salinetti:1984:MIS

[Sal84a] Gabriella Salinetti, editor. Multifunctions and Integrands: Stochastic Analysis, Approximation and Optimization Proceedings of a Conference held in Catania, Italy, June 7–16, 1983, volume 1091 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1091. URL http:/

/link.springer.com/book/10.1007/BFb0098799;   http://
www.springerlink.com/content/978-3-540-39083-1.

Salvadori:1984:BTA

[Sal84b] Luigi Salvadori, editor. Bifurcation Theory and Applications: Lectures given at the 2nd 1983 Session of the Centro Internationale Matematico Estivo (C.I.M.E.) held at Montecatini, Italy, June 24–July 2, 1983, volume 1057 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12931-6 (print), 3-540-38812-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1057. URL http:/

/link.springer.com/book/10.1007/BFb0098592;   http://
www.springerlink.com/content/978-3-540-38812-8.

Samaga:1980:UAM

[Sam80] Dr. H.-J. Samaga. A unified approach to Miquel’s theorem and its degenerations. Lecture Notes in Mathematics, 792:132–142, 1980. CODEN LNMAA2. ISBN 3-540-09976-X (print), 3-540-39214-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088672/.

Sampson:1982:FTE

[Sam82a] G. Sampson. Fourier transform estimates. Lecture Notes in Mathematics, 908:247–273, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093293/.

Sampson:1982:CHM

[Sam82b] J. H. Sampson. On a class of harmonic maps. Lecture Notes in Mathematics, 949:138–139, 1982. CODEN LNMAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069761.

Sandling:1981:GHT

[San81] Robert Sandling. Graham Higman’s thesis “Units in Group Rings”. Lecture Notes in Mathematics, 882:93–116, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092488/.

Sanders:1983:DJE

[San83] Jan A. Sanders. The (driven) Josephson equation: an exercise in asymptotics. Lecture Notes in Mathematics, 985:288–318, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062372/.

Sapronov:1984:BSS

[Sap84] Yu. I. Sapronov. Branching of solutions of smooth Fredholm equations. Lecture Notes in Mathematics, 1108:31–53, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099551/.

Sarvas:1983:STS

[Sar83] Jukka Sarvas. Stability theorem for spatial quasiconformal deformations. Lecture Notes in Mathematics, 1013:213–224, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066530/.

Sarnak:1984:ANT

[Sar84] Peter Sarnak. Additive number theory and Maass forms. Lecture Notes in Mathematics, 1052:286–309, 1984. CODEN LN-MAA2. ISBN 3-540-12909-X (print), 3-540-38788-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071548/.

Saut:1980:RNL

[Sau80] J.-C. Saut. Remarks on a non linear equation arising in population genetics. Lecture Notes in Mathematics, 782:265–282, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090437/.

Sawyer:1982:TWN

[Saw82] Eric T. Sawyer. Two weight norm inequalities for certain maximal and integral operators. Lecture Notes in Mathematics, 908: 102–127, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093283/.

Sazonov:1981:FDC

[Saz81a] Vjačeslav V. Sazonov. The finite dimensional case. Lecture Notes in Mathematics, 879:1–68, 1981. CODEN LNMAA2. ISBN 3-540-10863-7 (print), 3-540-38778-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096863/.

Sazonov:1981:IDC

[Saz81b] Vjačeslav V. Sazonov. The infinite dimensional case. Lecture Notes in Mathematics, 879:69–101, 1981. CODEN LN-MAA2. ISBN 3-540-10863-7 (print), 3-540-38778-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096864/.

Sazonov:1981:NAS

[Saz81c] Vjačeslav V. Sazonov. Normal Approximation: Some Recent Advances, volume 879 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 0-387-10863-7 (paperback), 3-540-10863-7 (paperback), 3-540-38778-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 104 pp. LCCN QA3 .L28 no. 879. URL http:/

/link.springer.com/book/10.1007/BFb0096862;   http://
www.springerlink.com/content/978-3-540-38778-7.

Stanton:1980:CSC

[SB80] R. G. Stanton and J. A. Bate. A computer search for B-coverings. Lecture Notes in Mathematics, 829:37–50, 1980. CODEN LN-MAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088898/.

Sbordone:1983:CMS

[Sbo83] Carlo Sbordone. On the compactness of minimizing sequences of variational problems. Lecture Notes in Mathematics, 979:240– 255, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066259/.

Schwartz:1981:CSMa

[SC81a] Laurent Schwartz and Paul R. Chernoff. Completely summing maps. Hilbert–Schmidt and nuclear maps. Lecture Notes in Mathematics, 852:9–15, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096726/.

Schwartz:1981:CSMb

[SC81b] Laurent Schwartz and Paul R. Chernoff. Completely summing maps: Six equivalent properties. p-radonifying maps. Lecture Notes in Mathematics, 852:18–25, 1981. CODEN LN-MAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096728/.

Schwartz:1981:EPS

[SC81c] Laurent Schwartz and Paul R. Chernoff. Equivalent properties, summing vs. factorization. Lecture Notes in Mathematics, 852: 61–67, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096737/.

Schwartz:1981:GSM

[SC81d] Laurent Schwartz and Paul R. Chernoff. Gauss-summing maps. Completion of Grothendieck factorization theorem. TLC and ILL. Lecture Notes in Mathematics, 852:78–85, 1981. CODEN LN-MAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096740/.

Schwartz:1981:GPB

[SC81e] Laurent Schwartz and Paul R. Chernoff. Geometry and Probability in Banach Spaces, volume 852 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 101 pp. LCCN QA3 .L28 no. 852.

URL http://link.springer.com/book/10.1007/BFb0096723;
 http://www.springerlink.com/content/978-3-540-38617-
9.

Schwartz:1981:KCP

[SC81f] Laurent Schwartz and Paul R. Chernoff. Kahane contraction principle. p-Gauss type the Gauss type interval is open. Lecture Notes in Mathematics, 852:51–55, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096735/.

Schwartz:1981:KIC

[SC81g] Laurent Schwartz and Paul R. Chernoff. Kahane inequality. The case of L^{p}Z-type. Lecture Notes in Mathematics, 852:46–51, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096734/.

Schwartz:1981:MTC

[SC81h] Laurent Schwartz and Paul R. Chernoff. Martingale type and cotype. Results of Pisier. Twelve properties equivalent to superreflexivity. Type for subspaces of L^{p} (Rosenthal theorem). Lecture Notes in Mathematics, 852:92–98, 1981. CODEN LN-MAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096742/.

Schwartz:1981:MCM

[SC81i] Laurent Schwartz and Paul R. Chernoff. More on cylindrical measures and stochastic processes. Lecture Notes in Mathematics, 852:42–45, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096733/.

Schwartz:1981:NEP

[SC81j] Laurent Schwartz and Paul R. Chernoff. Non-existence of (2 + ϵ)- Pietsch spaces, ultrapowers. Lecture Notes in Mathematics, 852: 67–72, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096738/.

Schwartz:1981:GL

[SC81k] Laurent Schwartz and Paul R. Chernoff. p-Gauss laws. Lecture Notes in Mathematics, 852:29–32, 1981. CODEN LN-MAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096730/.

Schwartz:1981:IM

[SC81l] Laurent Schwartz and Paul R. Chernoff. p-integral maps. Lecture Notes in Mathematics, 852:15–17, 1981. CODEN LN-MAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096727/.

Schwartz:1981:PSA

[SC81m] Laurent Schwartz and Paul R. Chernoff. p-Pietsch spaces. Application: Brownian motion. Lecture Notes in Mathematics, 852: 38–41, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096732/.

Schwartz:1981:PFT

[SC81n] Laurent Schwartz and Paul R. Chernoff. Pietsch factorization theorem. Lecture Notes in Mathematics, 852:5–9, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096725/.

Schwartz:1981:PIW

[SC81o] Laurent Schwartz and Paul R. Chernoff. The Pietsch interval. The weakest non-trivial superproperty. Cotypes, Rademacher vs. Gauss. Lecture Notes in Mathematics, 852:72–78, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096739/.

Schwartz:1981:PPC

[SC81p] Laurent Schwartz and Paul R. Chernoff. Proof of the Pietsch conjecture. Lecture Notes in Mathematics, 852:32–38, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096731/.

Schwartz:1981:FMT

[SC81q] Laurent Schwartz and Paul R. Chernoff. q-Factorization, Maurey’s theorem Grothendieck factorization theorem. Lecture Notes in Mathematics, 852:56–61, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096736/.

Schwartz:1981:RT

[SC81r] Laurent Schwartz and Paul R. Chernoff. Radonification theorem. Lecture Notes in Mathematics, 852:25–29, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096729/.

Schwartz:1981:SRS

[SC81s] Laurent Schwartz and Paul R. Chernoff. Super-reflexive spaces. Modulus of convexity, q-convexity ‘trees’ and Kelly–Chatteryji Theorem, Enflo theorem. Modulus of smoothness, p-smoothness. Properties equivalent to super-reflexivity. Lecture Notes in Mathematics, 852:85–92, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096741/.

Schwartz:1981:TCB

[SC81t] Laurent Schwartz and Paul R. Chernoff. Type and cotype for a Banach space p-summing maps. Lecture Notes in Mathematics, 852:1–5, 1981. CODEN LNMAA2. ISBN 3-540-10691-X (print), 3-540-38617-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096724/.

Chern:1983:SCM

[sC83] Shiing shen Chern. On surfaces of constant mean curvature in a three-dimensional space of constant curvature. Lecture Notes in Mathematics, 1007:104–108, 1983. CODEN LNMAA2. ISBN 3-540-12336-9 (print), 3-540-40969-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061413/.

Scedrov:1981:CIR

[Šče81a] Andrej Ščedrov. Consistency and independence results in intuitionistic set theory. Lecture Notes in Mathematics, 873:54– 86, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090727/.

Scedrov:1981:E

[Šče81b] Andrej Ščedrov. Errata. Lecture Notes in Mathematics, 873: 86, 1981. CODEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090728.

Scheringer:1980:PDD

[Sch80a] D. Scheringer. Präsentation der GL_{n} (D) und Dieudonné-Determinante. (German) []. Lecture Notes in Mathematics, 778: 9–20, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095918/.

Schmidt:1980:AAN

[Sch80b] Wolfgang M. Schmidt. Approximation by algebraic numbers. Lecture Notes in Mathematics, 785:251–288, 1980. CODEN LN-MAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38645-2_
8/.

Schmidt:1980:AIN

[Sch80c] Wolfgang M. Schmidt. Approximation to irrational numbers by rationals. Lecture Notes in Mathematics, 785:1–26, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38645-2_1/.

Schmidt:1980:BM

[Sch80d] Wolfgang M. Schmidt. Back matter. Lecture Notes in Mathematics, 785:289–299, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-38645-2/1.pdf.

Schmidt:1980:DA

[Sch80e] Wolfgang M. Schmidt. Diophantine Approximation, volume 785 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 299 pp. LCCN QA3 .L28 no. 785. URL http://link.

springer.com/book/10.1007/978-3-540-38645-2;   http:/
/www.springerlink.com/content/978-3-540-38645-2.

Schmidt:1980:FM

[Sch80f] Wolfgang M. Schmidt. Front matter. Lecture Notes in Mathematics, 785:I–XII, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-38645-2/1.pdf.

Schmidt:1980:GM

[Sch80g] Wolfgang M. Schmidt. Games and measures. Lecture Notes in Mathematics, 785:48–79, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38645-2_3/.

Schmidt:1980:IPP

[Sch80h] Wolfgang M. Schmidt. Integer points in parallelepipeds. Lecture Notes in Mathematics, 785:80–113, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38645-2_4/.

Schmidt:1980:NFE

[Sch80i] Wolfgang M. Schmidt. Norm form equations. Lecture Notes in Mathematics, 785:208–250, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38645-2_7/.

Schmidt:1980:RT

[Sch80j] Wolfgang M. Schmidt. Roth’s theorem. Lecture Notes in Mathematics, 785:114–150, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38645-2_5/.

Schmidt:1980:SA

[Sch80k] Wolfgang M. Schmidt. Simultaneous approximation. Lecture Notes in Mathematics, 785:27–47, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38645-2_2/.

Schmidt:1980:SAA

[Sch80l] Wolfgang M. Schmidt. Simultaneous approximation to algebraic numbers. Lecture Notes in Mathematics, 785:151–207, 1980. CODEN LNMAA2. ISBN 3-540-09762-7 (print), 3-540-38645-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38645-2_6/.

Schneider:1980:DEA

[Sch80m] Hans-Jürgen Schneider. Decomposable extensions of affine groups. Lecture Notes in Mathematics, 795:98–115, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090115/.

Schneider:1980:HVB

[Sch80n] Michael Schneider. Holomorphic vector bundles on P_{n}. Lecture Notes in Mathematics, 770:80–102, 1980. CODEN LN-MAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096237/.

Schwartz:1980:DMB

[Sch80o] Laurent Schwartz. Diffusion et mouvement brownien sur une variété sans bord. (French) []. Lecture Notes in Mathematics, 780:101–120, 1980. CODEN LNMAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096141/.

Schwartz:1980:LPA

[Sch80p] Laurent Schwartz. Localisation des processus attachés à une semimartingale vectorielle; Équivalences de semi-martingales vectorielles. (French) []. Lecture Notes in Mathematics, 780:14–28, 1980. CODEN LNMAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096136/.

Schwartz:1980:LSM

[Sch80q] Laurent Schwartz. Localisation des semi-martingales, et passage du local au global. (French) []. Lecture Notes in Mathematics, 780:7–13, 1980. CODEN LNMAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096135/.

Schwartz:1980:MCVa

[Sch80r] Laurent Schwartz. Martingales conformes à valeurs vectorielles et leurs localisations. (French) []. Lecture Notes in Mathematics, 780:29–37, 1980. CODEN LNMAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096137/.

Schwartz:1980:MCVb

[Sch80s] Laurent Schwartz. Martingales conformes sur une variété analytique complexe. (French) []. Lecture Notes in Mathematics, 822: 191–198, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097758/.

Schwartz:1980:MSM

[Sch80t] Laurent Schwartz. Martingales et semi-martingales conformes à valeurs dans des variétés analytiques complexes. (French) []. Lecture Notes in Mathematics, 780:38–55, 1980. CODEN LN-MAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096138/.

Schwartz:1980:SMVa

[Sch80u] Laurent Schwartz. Semi-martingales à valeurs dans une variété différentielle. (French) []. Lecture Notes in Mathematics, 780:1– 6, 1980. CODEN LNMAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096134/.

Schwartz:1980:SMVb

[Sch80v] Laurent Schwartz. Semi-Martingales sur des Variétés, et Martingales Conformes sur des Variétés Analytiques Complexes. (French) [Semi-Martingales on varieties and conforming martingales on complex analytical varieties], volume 780 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CO-DEN LNMAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 132 pp. LCCN QA3 .L28 no. 780. URL http:/

/link.springer.com/book/10.1007/BFb0096133;   http://
www.springerlink.com/content/978-3-540-38612-4.

Schwartz:1980:SESb

[Sch80w] Laurent Schwartz. Sous-espaces stables de martingales complexes. Sous-espaces stables et intégrales stochastiques associés à une semi-martingale conforme à valeurs dans une variété Canalytique. (French) []. Lecture Notes in Mathematics, 780:86– 100, 1980. CODEN LNMAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096140/.

Schwartz:1980:SESa

[Sch80x] Laurent Schwartz. Sous-espaces stables de martingales réeles. Sous-espaces stables et intégrales stochastiques associées à une semi-martingale à valeurs dans une variété. (French) []. Lecture Notes in Mathematics, 780:56–85, 1980. CODEN LN-MAA2. ISBN 3-540-09749-X (print), 3-540-38612-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096139/.

Schmerl:1981:RSR

[Sch81a] James H. Schmerl. Recursively saturated, rather classless models of Peano arithmetic. Lecture Notes in Mathematics, 859:268– 282, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090952/.

Schoeller:1981:HAL

[Sch81b] C. Schoeller. Homologie d’anneaux locaux de dimension d’immersion 3. (French) []. Lecture Notes in Mathematics, 867: 214–233, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090388/.

Schwarz:1981:CPL

[Sch81c] Friedhelm Schwarz. ’continuity’ properties in lattices of topological structures. Lecture Notes in Mathematics, 871:335–347, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089916/.

Schachermayer:1982:MCL

[Sch82a] Walter Schachermayer. Measurable and continuous linear functionals on spaces of uniformly continuous functions. Lecture Notes in Mathematics, 945:155–166, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096670/.

Schechtman:1982:LTI

[Sch82b] Gideon Schechtman. Lévy type inequality for a class of finite metric spaces. Lecture Notes in Mathematics, 939:211–215, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096270/.

Schenzel:1982:BTD

[Sch82c] Peter Schenzel. Bemerkungen zur Theorie Dualisierender Komplexe. (German) []. Lecture Notes in Mathematics, 907:28– 57, 1982. CODEN LNMAA2. ISBN 3-540-11187-5 (print), 3-540-38972-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094126/.

Schenzel:1982:DKI

[Sch82d] Peter Schenzel. Dualisierende Komplexe in der Iokalen Algebra und Buchsbaum- Ringe, volume 907 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11187-5 (print), 3-540-38972-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 161 pp. LCCN QA3 .L28 no. 907.

URL http://link.springer.com/book/10.1007/BFb0094123;
 http://www.springerlink.com/content/978-3-540-38972-
9.

Schenzel:1982:DKB

[Sch82e] Peter Schenzel. Dualisierender Komplex und Buchsbaum-Moduln. (German) []. Lecture Notes in Mathematics, 907:93– 112, 1982. CODEN LNMAA2. ISBN 3-540-11187-5 (print), 3-540-38972-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094128/.

Schenzel:1982:EGI

[Sch82f] Peter Schenzel. Enleitung. (German) [Introduction]. Lecture Notes in Mathematics, 907:1–9, 1982. CODEN LNMAA2. ISBN 3-540-11187-5 (print), 3-540-38972-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094124/.

Schenzel:1982:KBB

[Sch82g] Peter Schenzel. Konstruktion und Beispiele von Buchsbaum-Ringen. (German) []. Lecture Notes in Mathematics, 907:113– 131, 1982. CODEN LNMAA2. ISBN 3-540-11187-5 (print), 3-540-38972-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094129/.

Schenzel:1982:SKK

[Sch82h] Peter Schenzel. Simpliziale Komplexe und Kombinatorik. (German) []. Lecture Notes in Mathematics, 907:132–151, 1982. CO-DEN LNMAA2. ISBN 3-540-11187-5 (print), 3-540-38972-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094130/.

Schenzel:1982:VEB

[Sch82i] Peter Schenzel. Vorbereitende Ergebnisse und Bezeichnungen. (German) []. Lecture Notes in Mathematics, 907:10–27, 1982. CODEN LNMAA2. ISBN 3-540-11187-5 (print), 3-540-38972-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094125/.

Schenzel:1982:VLK

[Sch82j] Peter Schenzel. Zum Verschwinden Lokaler Kohomologiemoduln. (German) []. Lecture Notes in Mathematics, 907:58–92, 1982. CODEN LNMAA2. ISBN 3-540-11187-5 (print), 3-540-38972-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094127/.

Schue:1982:RLA

[Sch82k] John Schue. Representations of Lie p-algebras. Lecture Notes in Mathematics, 933:191–202, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093362/.

Schumaker:1982:OSS

[Sch82l] Larry L. Schumaker. Optimal spline solutions of systems of ordinary differential equations. Lecture Notes in Mathematics, 957: 272–283, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066243/.

Schwartz:1982:EFE

[Sch82m] Laurent Schwartz. Errata. (French) [Errata]. Lecture Notes in Mathematics, 921:149–150, 1982. CODEN LNMAA2. ISBN 3-540-11486-6 (print), 3-540-39167-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092648.

Schwartz:1982:GDD

[Sch82n] Laurent Schwartz. Géométrie différentielle du 2ème ordre, semimartingales et équations différentielles stochastiques sur une variété différentielle. (French) []. Lecture Notes in Mathematics, 921:1–148, 1982. CODEN LNMAA2. ISBN 3-540-11486-6 (print), 3-540-39167-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092647/.

Schwarz:1982:FRC

[Sch82o] A. S. Schwarz. A few remarks on the construction of solutions of non-linear equations. Lecture Notes in Mathematics, 970:124– 128, 1982. CODEN LNMAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066027/.

Schwarz:1982:SRQ

[Sch82p] H. R. Schwarz. Simultaneous Rayleigh-quotient iteration methods for large sparse generalized eigenvalue problems. Lecture Notes in Mathematics, 968:384–398, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

Schmets:1983:ACS

[Sch83a] Jean Schmets. Absolutely convex subsets of C(X; E). Lecture Notes in Mathematics, 1003:24–43, 1983. CODEN LN-MAA2. ISBN 3-540-12327-X (print), 3-540-40951-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061452/.

Schmets:1983:BLF

[Sch83b] Jean Schmets. Bounded linear functionals on C_{P} (X; E). Lecture Notes in Mathematics, 1003:110–112, 1983. CODEN LN-MAA2. ISBN 3-540-12327-X (print), 3-540-40951-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061455/.

Schmets:1983:CLC

[Sch83c] Jean Schmets. Characterization of locally convex properties of the C_{P} (X; E) spaces. Lecture Notes in Mathematics, 1003:70– 109, 1983. CODEN LNMAA2. ISBN 3-540-12327-X (print), 3-540-40951-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061454/.

Schmets:1983:SVVa

[Sch83d] Jean Schmets. Spaces of vector-valued continuous functions. Lecture Notes in Mathematics, 1003:1–23, 1983. CODEN LN-MAA2. ISBN 3-540-12327-X (print), 3-540-40951-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061451/.

Schmets:1983:SVVb

[Sch83e] Jean Schmets. Spaces of Vector-Valued Continuous Functions, volume 1003 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12327-X (print), 3-540-40951-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 117 pp. LCCN QA3 .L28 no. 1003. URL http:/

/link.springer.com/book/10.1007/BFb0061450;   http://
www.springerlink.com/content/978-3-540-40951-9.

Schmets:1983:SD

[Sch83f] Jean Schmets. Structure of the dual of C_{P} (X; E). Lecture Notes in Mathematics, 1003:44–69, 1983. CODEN LNMAA2. ISBN 3-540-12327-X (print), 3-540-40951-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061453/.

Schmidt:1983:AMT

[Sch83g] Klaus D. Schmidt. Amarts — a measure theoretic approach. Lecture Notes in Mathematics, 1042:51–236, 1983. CODEN LN-MAA2. ISBN 3-540-12867-0 (print), 3-540-38754-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073702/.

Schmidt:1983:JDV

[Sch83h] Klaus D. Schmidt. On the Jordan decomposition for vector measures. Lecture Notes in Mathematics, 990:198–203, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064272/.

Schreiber:1983:SAE

[Sch83i] Robert Schreiber. Systolic arrays for eigenvalues. Lecture Notes in Mathematics, 1005:284–296, 1983. CODEN LNMAA2. ISBN 3-540-12334-2 (print), 3-540-40967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0112541/.

Schwermer:1983:HDF

[Sch83j] Joachim Schwermer. Das Hochheben einer Differentialform auf e′(P) zu einer Familie von Formen auf Γ ⋂ P/X. (German) []. Lecture Notes in Mathematics, 988:53–62, 1983. CODEN LN-MAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070273/.

Schwermer:1983:EAD

[Sch83k] Joachim Schwermer. Die Eisensteinkohomologie arithmetisch definierter Untergruppen von SL_{3}Q. (German) []. Lecture Notes in Mathematics, 988:136–163, 1983. CODEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070279/.

Schwermer:1983:KRB

[Sch83l] Joachim Schwermer. Die Kohomologie des Randes der Borel– Serre-Kompaktifizierung. (German) []. Lecture Notes in Mathematics, 988:105–122, 1983. CODEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070277/.

Schwermer:1983:EG

[Sch83m] Joachim Schwermer. Einleitung. (German) []. Lecture Notes in Mathematics, 988:1–19, 1983. CODEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070269/.

Schwermer:1983:ESM

[Sch83n] Joachim Schwermer. Eisensteinkohomologieklassen, die Spitzenformen auf e′(P), P maximal parabolich, zugeordnet sind. (German) []. Lecture Notes in Mathematics, 988:123–135, 1983. CO-DEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070278/.

Schwermer:1983:EGE

[Sch83o] Joachim Schwermer. Eisensteinkohomologieklassen. (German) [Eisenstein cohomology classes]. Lecture Notes in Mathematics, 988:63–84, 1983. CODEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070274/.

Schwermer:1983:HES

[Sch83p] Joachim Schwermer. Holomorphieaussagen für Eisensteinreihen an speziellen Werten. (German) []. Lecture Notes in Mathematics, 988:85–93, 1983. CODEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070275/.

Schwermer:1983:KADa

[Sch83q] Joachim Schwermer. Kohomologie arithmetisch definierter Gruppen. (German) [Cohomology of arithmetically defined groups]. Lecture Notes in Mathematics, 988:24–43, 1983. CODEN LN-MAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070271/.

Schwermer:1983:KADb

[Sch83r] Joachim Schwermer. Kohomologie arithmetisch definierter Gruppen und Eisensteinreihen. (German) [Cohomology of arithmeticallydefined groups and Eisenstein series], volume 988 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CO-DEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 170 pp. LCCN QA3 .L28 no. 988. URL http:/

/link.springer.com/book/10.1007/BFb0070268;   http://
www.springerlink.com/content/978-3-540-39623-9.

Schwermer:1983:KSI

[Sch83s] Joachim Schwermer. Kohomologie einer Seite e′(P) im Rand von ΓX. (German) []. Lecture Notes in Mathematics, 988:44–52, 1983.

CODEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070272/.

Schwermer:1983:ULQ

[Sch83t] Joachim Schwermer. Nicht-Unitarisierbarkeit von Langlands-Quotienten und Holomorphieaussagen für Eisensteinreihen. (German) []. Lecture Notes in Mathematics, 988:94–103, 1983. CO-DEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070276/.

Schwermer:1983:TNG

[Sch83u] Joachim Schwermer. Terminologie und Notationen. (German) []. Lecture Notes in Mathematics, 988:21–22, 1983. CODEN LNMAA2. ISBN 3-540-12292-3 (print), 3-540-39623-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070270.

Schmerl:1984:DEF

[Sch84a] Ulf R. Schmerl. Diophantine equations in a fragment of number theory. Lecture Notes in Mathematics, 1104:389–398, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099493/.

Schmidt:1984:AVK

[Sch84b] Claus-Günther Schmidt. Abelsche Varietäten mit komplexer Multiplikation. (German) []. Lecture Notes in Mathematics, 1082: 24–39, 1984. CODEN LNMAA2. ISBN 3-540-13863-3 (print), 3-540-39052-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100167/.

Schmidt:1984:AAV

[Sch84c] Claus-Günther Schmidt. Arithmetik Abelscher Varietäten mit komplexer Multiplikation, volume 1082 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13863-3 (print), 3-540-39052-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 96 pp. LCCN QA3 .L28 no. 1082.

URL http://link.springer.com/book/10.1007/BFb0100164;

 http://www.springerlink.com/content/978-3-540-39052-
7.

Schmidt:1984:CVF

[Sch84d] Claus-Günther Schmidt. Die CM-Varietäten der Fermat– Jacobischen. (German) []. Lecture Notes in Mathematics, 1082: 78–88, 1984. CODEN LNMAA2. ISBN 3-540-13863-3 (print), 3-540-39052-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100171/.

Schmidt:1984:HAC

[Sch84e] Claus-Günther Schmidt. Die Halbsystem Abelscher CM-Typen. (German) []. Lecture Notes in Mathematics, 1082:39–44, 1984. CODEN LNMAA2. ISBN 3-540-13863-3 (print), 3-540-39052-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100168/.

Schmidt:1984:EG

[Sch84f] Claus-Günther Schmidt. Einleitung. (German) []. Lecture Notes in Mathematics, 1082:1–3, 1984. CODEN LNMAA2. ISBN 3-540-13863-3 (print), 3-540-39052-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100165/.

Schmidt:1984:GAK

[Sch84g] Claus-Günther Schmidt. Geometrische Annullator-Kriterien. (German) []. Lecture Notes in Mathematics, 1082:44–58, 1984. CODEN LNMAA2. ISBN 3-540-13863-3 (print), 3-540-39052-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100169/.

Schmidt:1984:GTG

[Sch84h] Claus-Günther Schmidt. Grössencharaktere vom Typ A_{o}. (German) []. Lecture Notes in Mathematics, 1082:4–23, 1984. CODEN LNMAA2. ISBN 3-540-13863-3 (print), 3-540-39052-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100166/.

Schmidt:1984:MUE

[Sch84i] Claus-Günther Schmidt. Modulikörper und unverzweigte Erweiterungen. (German) []. Lecture Notes in Mathematics, 1082: 58–78, 1984. CODEN LNMAA2. ISBN 3-540-13863-3 (print), 3-540-39052-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100170/.

Schmitt:1984:MSC

[Sch84j] Peter H. Schmitt. Model- and substructure complete theories of ordered Abelian groups. Lecture Notes in Mathematics, 1103: 389–418, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099396/.

Schneider:1984:RAT

[Sch84k] P. Schneider. Rigid-analytic L-transforms. Lecture Notes in Mathematics, 1068:216–230, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099455/.

Scott:1981:HAP

[Sco81] Leonard Scott. Hecke actions on Picard groups and class groups. Lecture Notes in Mathematics, 882:251–260, 1981. CODEN LN-MAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092498/.

Sealey:1982:HDS

[Sea82] H. C. J. Sealey. Harmonic diffeomorphisms of surfaces. Lecture Notes in Mathematics, 949:140–145, 1982. CODEN LN-MAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069762/.

Sebbar:1980:PSH

[Seb80a] A. Sebbar. Prolongement des solutions holomorphes de certains opérateurs différentiels d’ordre infini à coefficients constants. (French) []. Lecture Notes in Mathematics, 822:199–220, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097759/.

Seberry:1980:HDO

[Seb80b] Jennifer Seberry. Higher dimensional orthogonal designs and Hadamard matrices. Lecture Notes in Mathematics, 829:220– 223, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088915/.

Secrest:1982:OTU

[Sec82] Don Secrest. An overview of the techniques in use for solving the coupled equations of scattering theory. Lecture Notes in Mathematics, 968:1–21, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064878/.

Segal:1982:I

[Seg82a] I. E. Segal. Introduction. Lecture Notes in Mathematics, 905: 2, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092422.

Segal:1982:PTG

[Seg82b] I. E. Segal. Particle theory and global geometry. Lecture Notes in Mathematics, 905:151–164, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092434/.

Segal:1983:NWE

[Seg83] I. E. Segal. Nonlinear wave equations. Lecture Notes in Mathematics, 1037:115–141, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073171/.

Seidel:1981:GTD

[Sei81] J. J. Seidel. Graphs and two-distance sets. Lecture Notes in Mathematics, 884:90–98, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091810/.

Seldin:1981:CAC

[Sel81a] Jonathan P. Seldin. A constructive approach to classical mathematics. Lecture Notes in Mathematics, 873:105–110, 1981. CO-DEN LNMAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090730/.

Seligman:1981:RCC

[Sel81b] George B. Seligman. Representations of certain classical Lie algebras. Lecture Notes in Mathematics, 848:33–40, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090555/.

Seligman:1983:HEC

[Sel83] George B. Seligman. Higher even Clifford algebras. Lecture Notes in Mathematics, 1029:159–191, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098931/.

Semadeni:1982:BBL

[Sem82a] Zbigniew Semadeni. Bases of broken-line functions of a single variable. Lecture Notes in Mathematics, 918:21–43, 1982. CODEN LNMAA2. ISBN 3-540-11481-5 (print), 3-540-39143-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094631/.

Semadeni:1982:BPM

[Sem82b] Zbigniew Semadeni. Bases of pyramidal and multiaffine functions of several variables. Lecture Notes in Mathematics, 918: 44–85, 1982. CODEN LNMAA2. ISBN 3-540-11481-5 (print), 3-540-39143-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094632/.

Semadeni:1982:FPB

[Sem82c] Zbigniew Semadeni. Further properties of bases in C(X). Lecture Notes in Mathematics, 918:86–130, 1982. CODEN LN-MAA2. ISBN 3-540-11481-5 (print), 3-540-39143-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094633/.

Semadeni:1982:GPB

[Sem82d] Zbigniew Semadeni. General properties of bases in C(X). Lecture Notes in Mathematics, 918:1–20, 1982. CODEN LNMAA2. ISBN 3-540-11481-5 (print), 3-540-39143-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094630/.

Semadeni:1982:SBB

[Sem82e] Zbigniew Semadeni. Schauder Bases in Banach Spaces of Continuous Functions, volume 918 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11481-5 (print), 3-540-39143-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 136 pp. LCCN QA3 .L28 no. 918.

URL http://link.springer.com/book/10.1007/BFb0094629;
 http://www.springerlink.com/content/978-3-540-39143-
2.

Semmes:1983:ACA

[Sem83] Stephen Semmes. Another characterization of H^{p}, 0 < p < ∞, with an application to interpolation. Lecture Notes in Mathematics, 992:212–226, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069161/.

Senatov:1983:RBL

[Sen83a] V. V. Senatov. On a relation between Lévy–Prohorov metrics and ideal metrics. Lecture Notes in Mathematics, 982:213–216, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082071/.

Seneor:1983:SIR

[Sén83b] R. Sénéor. Superrenormalizable infra red theories. Lecture Notes in Mathematics, 1031:108–113, 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072114/.

Sentilles:1984:SMT

[Sen84] Dennis Sentilles. Some measure theoretic implications for the Pettis integral. Lecture Notes in Mathematics, 1089:165–170, 1984.

CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072612/.

Serre:1980:TSE

[Ser80] D. Serre. Triplets de solutions d’une équation aux derivées partielles elliptique non linéaire. (French) []. Lecture Notes in Mathematics, 782:283–296, 1980. CODEN LNMAA2. ISBN 3-540-09758-9 (print), 3-540-38637-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090438/.

Serrin:1984:AMS

[Ser84] J. Serrin. Applied mathematics and scientific thought. Lecture Notes in Mathematics, 1107:19–27, 1984. CODEN LN-MAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101491/.

Schaeffer:1981:GIS

[SH81a] David Schaeffer and John Hayden. General introduction to steady state bifurcation. Lecture Notes in Mathematics, 898:13–47, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091905/.

Semmes:1981:SCA

[SH81b] Stephen Semmes and Franklin Haimo. Some complete antinormal subgroups of infinite symmetric groups. Lecture Notes in Mathematics, 848:262–273, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090572/.

Streit:1983:WNA

[SH83] L. Streit and T. Hida. White noise analysis and its application to Feynman integral. Lecture Notes in Mathematics, 1033:219– 226, 1983. CODEN LNMAA2. ISBN 3-540-12703-8 (print), 3-540-38690-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099857/.

Schroeder-Heister:1984:GRQ

[SH84] Peter Schroeder-Heister. Generalized rules for quantifiers and the completeness of the intuitionistic operators &, ν, ⊃, λ, ∀, ϵ. Lecture Notes in Mathematics, 1104:399–426, 1984. CODEN LN-MAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099494/.

Sharpe:1980:LTS

[Sha80] M. J. Sharpe. Local times and singularities of continuous local martingales. Lecture Notes in Mathematics, 784:76–101, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089471/.

Shampine:1982:SOQ

[Sha82] L. F. Shampine. Solving ODEs in quasi steady state. Lecture Notes in Mathematics, 968:234–245, 1982. CODEN LN-MAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064891/.

Shampine:1984:GEE

[Sha84a] Lawrence F. Shampine. Global error estimation for stiff ODEs. Lecture Notes in Mathematics, 1066:159–168, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099523/.

Sharko:1984:MMF

[Sha84b] V. V. Sharko. Minimal Morse functions. Lecture Notes in Mathematics, 1108:218–234, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099560/.

Sheehan:1981:FRT

[She81a] John Sheehan. Finite Ramsey theory is hard. Lecture Notes in Mathematics, 884:99–106, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091811/.

Shelstad:1981:BCM

[She81b] D. Shelstad. Base change and a matching theorem for real groups. Lecture Notes in Mathematics, 880:425–482, 1981. CODEN LN-MAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090419/.

Sher:1981:CTS

[She81c] R. B. Sher. Complement theorems in shape theory. Lecture Notes in Mathematics, 870:150–168, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089714/.

Sherman:1981:TNS

[She81d] Clayton Sherman. K′-theory of Noetherian schemes. Lecture Notes in Mathematics, 854:343–371, 1981. CODEN LN-MAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089529/.

Shelah:1982:PAR

[She82a] Saharon Shelah. α-properness and not adding reals. Lecture Notes in Mathematics, 940:153–194, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21543-2_5/.

Shelah:1982:BM

[She82b] Saharon Shelah. Back matter. Lecture Notes in Mathematics, 940: 492–498, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-662-21543-2/1.pdf.

Shelah:1982:CCE

[She82c] Saharon Shelah. Changing cofinalities; equi-consistency results. Lecture Notes in Mathematics, 940:354–393, 1982. CODEN LN-MAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21543-2_

Shelah:1982:FBF

[She82d] Saharon Shelah. Forcing, basic facts. Lecture Notes in Mathematics, 940:1–38, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21543-2_1/.

Shelah:1982:FM

[She82e] Saharon Shelah. Front matter. Lecture Notes in Mathematics, 940:I–XXIX, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21543-2/1.pdf.

Shelah:1982:IF

[She82f] Saharon Shelah. Iteration of forcing. Lecture Notes in Mathematics, 940:39–72, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21543-2_2/.

Shelah:1982:NIF

[She82g] Saharon Shelah. Notes on improper forcing. Lecture Notes in Mathematics, 940:394–409, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21543-2_12/.

Shelah:1982:ONP

[She82h] Saharon Shelah. The October notes on proper forcing. Lecture Notes in Mathematics, 940:258–277, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21543-2_8/.

Shelah:1982:OCC

[She82i] Saharon Shelah. On oracle-c.c. and “P (ω) / finite has no nontrivial automorphism”. Lecture Notes in Mathematics, 940:114– 152, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21543-2_4/.

Shelah:1982:SPF

[She82j] Saharon Shelah. On semi-proper forcing. Lecture Notes in Mathematics, 940:304–353, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21543-2_10/.

Shelah:1982:WDP

[She82k] Saharon Shelah. On weak diamonds and the power of ext. Lecture Notes in Mathematics, 940:461–491, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21543-2_14/.

Shelah:1982:PPT

[She82l] Saharon Shelah. p-points and preservation theorems. Lecture Notes in Mathematics, 940:195–232, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21543-2_6/.

Shelah:1982:PFa

[She82m] Saharon Shelah. Proper forcing. Lecture Notes in Mathematics, 940:73–113, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21543-2_3/.

Shelah:1982:PFb

[She82n] Saharon Shelah. Proper Forcing, volume 940 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxix + 496 pp. LCCN QA3 .L28

no. 940. URL http://link.springer.com/book/10.1007/978-
3-662-21543-2;   http://www.springerlink.com/content/
978-3-662-21543-2.

Shelah:1982:SNP

[She82o] Saharon Shelah. The September notes on proper forcing. Lecture Notes in Mathematics, 940:233–257, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21543-2_7/.

Shelah:1982:SHD

[She82p] Saharon Shelah. Souslin hypothesis does not imply “Every Aronszajn Tree is Special”. Lecture Notes in Mathematics, 940:278– 303, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21543-2_9/.

Shelah:1982:SCL

[She82q] Saharon Shelah. Strong covering lemma and the G.C.H. Lecture Notes in Mathematics, 940:410–460, 1982. CODEN LNMAA2. ISBN 3-540-11593-5 (print), 3-662-21543-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21543-2_13/.

Sherman:1982:GRA

[She82r] Clayton Sherman. Group representations and algebraic K-theory. Lecture Notes in Mathematics, 966:208–243, 1982. CODEN LN-MAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062177/.

Shinoda:1981:SET

[Shi81a] Juichi Shinoda. Sections and envelopes of type 2 objects. Lecture Notes in Mathematics, 891:175–188, 1981. CODEN LN-MAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090985/.

Shirokov:1981:DMI

[Shi81b] N. A. Shirokov. Division and multiplication by inner functions in spaces of analytic functions smooth up to the boundary. Lecture Notes in Mathematics, 864:413–439, 1981. CODEN LN-MAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097003/.

Shiffman:1983:ICG

[Shi83a] Bernard Shiffman. Introduction to the Carlson–Griffiths equidistribution theory. Lecture Notes in Mathematics, 981:44–89, 1983.

CODEN LNMAA2. ISBN 3-540-12003-3 (print), 3-540-39480-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066383/.

Shimizu:1983:SLM

[Shi83b] R. Shimizu. On the stability of lack of memory characterization of the exponential distribution. Lecture Notes in Mathematics, 982:217–226, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082072/.

Shioda:1983:ACC

[Shi83c] Tetsuji Shioda. Algebraic cycles on a certain hypersurface. Lecture Notes in Mathematics, 1016:271–294, 1983. CODEN LN-MAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099967/.

Shore:1981:DUG

[Sho81] Richard A. Shore. The degrees of unsolvability: Global results. Lecture Notes in Mathematics, 859:283–301, 1981. CODEN LNMAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090953/.

Shortt:1984:BSS

[Sho84] R. M. Shortt. Big sets are strongly Blackwell. Lecture Notes in Mathematics, 1089:58–63, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072600/.

Sibuya:1980:CFP

[Sib80] Yasutaka Sibuya. Convergence of formal power series solutions of a system of nonlinear differential equations at an irregular singular point. Lecture Notes in Mathematics, 810:135–142, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089978/.

Siciak:1984:HNF

[Sic84] Józef Siciak. Highly noncontinuable functions on polynomially convex sets. Lecture Notes in Mathematics, 1094:173–178, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099161/.

Sidney:1983:EPD

[Sid83] S. J. Sidney. Exposed points for duals of separable Fréchet spaces. Lecture Notes in Mathematics, 995:166–168, 1983. CO-DEN LNMAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061893/.

Siebenmann:1980:VRI

[Sie80a] L. Siebenmann. On vanishing of the Rohlin invariant and nonfinitely amphicheiral homology 3-spheres. Lecture Notes in Mathematics, 788:172–222, 1980. CODEN LNMAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099246/.

Siebenmann:1980:ACD

[Sie80b] Laurent Siebenmann. Amorces de la chirurgie en dimension quatre: un S^{3} × R exotique [d’après Andrew H. Casson and Michael H. Freedman]. (French) []. Lecture Notes in Mathematics, 770: 183–207, 1980. CODEN LNMAA2. ISBN 3-540-09733-3 (print), 3-540-38547-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096243/.

Sieg:1981:IDC

[Sie81a] Wilfried Sieg. Inductive definitions, constructive ordinals, and normal derivations. Lecture Notes in Mathematics, 897:143– 187, 1981. CODEN LNMAA2. ISBN 3-540-11170-0 (print), 3-540-38649-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091897/.

Siegberg:1981:CMP

[Sie81b] H.-W. Siegberg. Chaotic mappings on S^{1} periods one, two, three imply chaos on S^{1}. Lecture Notes in Mathematics, 878:351– 370, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090688/.

Siebert:1982:CHP

[Sie82] Eberhard Siebert. Continuous hemigroups of probability measures on a Lie group. Lecture Notes in Mathematics, 928:362–402, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093234/.

Sigal:1983:BVRa

[Sig83a] Israel Michael Sigal. Boundary values of the resolvent. General single-channel case. Lecture Notes in Mathematics, 1011: 60–66, 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065398/.

Sigal:1983:BVRb

[Sig83b] Israel Michael Sigal. Boundary values of the resolvent. The general case. Lecture Notes in Mathematics, 1011:67–72, 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065399/.

Sigal:1983:EP

[Sig83c] Israel Michael Sigal. Exact parametrices. Lecture Notes in Mathematics, 1011:44–50, 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065396/.

Sigal:1983:IQS

[Sig83d] Israel Michael Sigal. Instability of quasibound-state and embedded eigenvalues. Lecture Notes in Mathematics, 1011:76–81, 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065401/.

Sigal:1983:I

[Sig83e] Israel Michael Sigal. Introduction. Lecture Notes in Mathematics, 1011:1–7, 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065392/.

Sigal:1983:NDA

[Sig83f] Israel Michael Sigal. Non-dilation-analytic potentials. Lecture Notes in Mathematics, 1011:73–75, 1983. CODEN LN-MAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065400/.

Sigal:1983:QSF

[Sig83g] Israel Michael Sigal. Quasibound states and the finiteness of the discrete spectrum. Lecture Notes in Mathematics, 1011:51– 59, 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065397/.

Sigal:1983:STM

[Sig83h] Israel Michael Sigal. Scattering Theory for Many-Body Quantum Mechanical Systems — Rigorous Results, volume 1011 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 132 pp. LCCN QA3 .L28 no. 1011. URL http:/

/link.springer.com/book/10.1007/BFb0065391;   http://
www.springerlink.com/content/978-3-540-38664-3.

Sigal:1983:STB

[Sig83i] Israel Michael Sigal. Scattering theory for n-body systems. Lecture Notes in Mathematics, 1011:31–43, 1983. CODEN LN-MAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065395/.

Sigal:1983:SD

[Sig83j] Israel Michael Sigal. Spectral decomposition. Lecture Notes in Mathematics, 1011:8–13, 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065393/.

Sigal:1983:TSS

[Sig83k] Israel Michael Sigal. Two-space scattering theory. Lecture Notes in Mathematics, 1011:14–30, 1983. CODEN LNMAA2. ISBN 3-540-12672-4 (print), 3-540-38664-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065394/.

Siganov:1983:SRA

[Šig83l] I. Šiganov. Several remarks on applications of one approach to studies of characterization problems of Pólya’s theorem type. Lecture Notes in Mathematics, 982:227–237, 1983. CODEN LN-MAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082073/.

Silvester:1982:SLR

[Sil82] J. R. Silvester. On the Gl_{n} of a semi-local ring. Lecture Notes in Mathematics, 966:244–260, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062178/.

Silva:1983:IUR

[Sil83] M. W. Baldoni Silva. L^{2} index and unitary representations. Lecture Notes in Mathematics, 1020:1–10, 1983. CODEN LN-MAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071494/.

Simms:1980:GAF

[Sim80] D. J. Simms. Geometric aspects of the Feynman integral. Lecture Notes in Mathematics, 836:167–170, 1980. CODEN LN-MAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089732/.

Schmidt:1982:AMG

[SJ82] Wolfgang Schmidt and Antony Jameson. Applications of multigrid methods for transonic flow calculations. Lecture Notes in Mathematics, 960:599–613, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069946/.

Spence:1984:NTN

[SJ84] A. Spence and A. Jepson. Numerical techniques for nonlinear multi-parameter problems. Lecture Notes in Mathematics, 1066: 169–185, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099524/.

Sjogren:1982:WSS

[Sjö82] Peter Sjögren. A weak spectral synthesis property for Hardy and Lipschitz spaces. Lecture Notes in Mathematics, 908:285– 296, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093295/.

Sjogren:1983:MFM

[Sjö83] Peter Sjögren. On the maximal function for the Mehler kernel. Lecture Notes in Mathematics, 992:73–82, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069151/.

Stanton:1980:IRT

[SK80] R. G. Stanton and J. G. Kalbfleisch. Introduction to Ramsey theory. Lecture Notes in Mathematics, 829:1–28, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088896/.

Skoda:1980:DAB

[Sko80a] H. Skoda. Diviseurs d’aire bornée dans la boule de C^{2}: réflexions sur un article de Bo Berndtsson. (French) []. Lecture Notes in Mathematics, 822:238–251, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097761/.

Skoda:1980:RPT

[Sko80b] H. Skoda. Remarques à propos des théorèmes d’annulation pour les fibrés semi-positifs. (French) []. Lecture Notes in Mathematics, 822:252–258, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097762/.

Skoda:1980:RSG

[Sko80c] Henri Skoda. Relèvement des sections globales dans les fibrés semi-positifs. (French) []. Lecture Notes in Mathematics, 822: 259–303, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097763/.

Skrypnik:1984:TMI

[Skr84] I. V. Skrypnik. Topological methods of investigation of general nonlinear elliptical boundary value problems. Lecture Notes in Mathematics, 1108:235–256, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099561/.

Slater:1984:SDC

[Sla84] Peter J. Slater. Some definitions of central structures. Lecture Notes in Mathematics, 1073:169–178, 1984. CODEN LN-MAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073115/.

Sleeman:1980:MPD

[Sle80] B. D. Sleeman. Multiparameter periodic differential equations. Lecture Notes in Mathematics, 827:229–250, 1980. CODEN LN-MAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091382/.

Slessenger:1984:SSC

[Sle84] P. H. Slessenger. On subsets of the Skolem class of exponential polynomials. Lecture Notes in Mathematics, 1104:427–436, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099495/.

Slodowy:1980:DT

[Slo80a] Peter Slodowy. Deformation theory. Lecture Notes in Mathematics, 815:4–16, 1980. CODEN LNMAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090296/.

Slodowy:1980:DSS

[Slo80b] Peter Slodowy. Deformations of simple singularities. Lecture Notes in Mathematics, 815:121–151, 1980. CODEN LN-MAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090302/.

Slodowy:1980:NES

[Slo80c] Peter Slodowy. Nilpotent elements in simple Lie algebras. Lecture Notes in Mathematics, 815:103–120, 1980. CODEN LN-MAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090301/.

Slodowy:1980:QAA

[Slo80d] Peter Slodowy. The quotient of the adjoint action. Lecture Notes in Mathematics, 815:17–42, 1980. CODEN LNMAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090297/.

Slodowy:1980:RGA

[Slo80e] Peter Slodowy. Regular group actions. Lecture Notes in Mathematics, 815:1–3, 1980. CODEN LNMAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090295/.

Slodowy:1980:RAQ

[Slo80f] Peter Slodowy. The resolution of the adjoint quotient. Lecture Notes in Mathematics, 815:43–59, 1980. CODEN LN-MAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090298/.

Slodowy:1980:SSb

[Slo80g] Peter Slodowy. Simple singularities. Lecture Notes in Mathematics, 815:70–102, 1980. CODEN LNMAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090300/.

Slodowy:1980:SSS

[Slo80h] Peter Slodowy. Simple Singularities and Simple Algebraic Groups, volume 815 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 175 pp. LCCN QA3 .L28 no. 815. URL http:/

/link.springer.com/book/10.1007/BFb0090294;   http://
www.springerlink.com/content/978-3-540-38191-4.

Slodowy:1980:SSa

[Slo80i] Peter Slodowy. Subregular singularities. Lecture Notes in Mathematics, 815:60–69, 1980. CODEN LNMAA2. ISBN 3-540-10026-1 (print), 3-540-38191-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090299/.

Slodowy:1982:CGD

[Slo82] Peter Slodowy. Chevalley groups over C((t)) and deformations of simply elliptic singularities. Lecture Notes in Mathematics, 961: 285–301, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071289/.

Slodowy:1983:PSK

[Slo83] P. Slodowy. Platonic solids, Kleinian singularities, and Lie groups. Lecture Notes in Mathematics, 1008:102–138, 1983. CODEN LNMAA2. ISBN 3-540-12337-7 (print), 3-540-40971-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065703/.

Stanton:1980:SNR

[SM80] R. G. Stanton and R. C. Mullin. Some new results on the covering numbers N(t, k, v). Lecture Notes in Mathematics, 829:51– 58, 1980. CODEN LNMAA2. ISBN 3-540-10254-X (print), 3-540-38376-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088899/.

Stanton:1981:SPD

[SM81] R. G. Stanton and R. C. Mullin. Some properties of H-designs. Lecture Notes in Mathematics, 884:1–7, 1981. CODEN LN-MAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091802/.

Schuss:1983:DSD

[SM83a] Z. Schuss and B. J. Matkowsky. Dynamical systems driven by small white noise: Asymptotic analysis and applications. Lecture Notes in Mathematics, 985:2–34, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062360/.

Schuss:1983:IPV

[SM83b] Z. Schuss and B. J. Matkowsky. Introduction to the paper by V. F. Butuzov and A. B. Vasil’eva. Lecture Notes in Mathematics, 985:35–74, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062361/.

Shlesinger:1983:FSP

[SM83c] Michael F. Shlesinger and Elliott W. Montroll. Fractal stochastic processes: Clusters and intermittancies. Lecture Notes in Mathematics, 1035:138–152, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073257/.

Smith:1981:TTA

[Smi81] Rick L. Smith. Two theorems on autostability in p-groups. Lecture Notes in Mathematics, 859:302–311, 1981. CODEN LN-MAA2. ISBN 3-540-10708-8 (print), 3-540-38673-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090954/.

Smith:1982:ARP

[Smi82] P. F. Smith. The Artin–Rees property. Lecture Notes in Mathematics, 924:197–240, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092934/.

Smith:1983:SCT

[Smi83] Brent Smith. A strong convergence theorem for H^{1}(T). Lecture Notes in Mathematics, 995:169–173, 1983. CODEN LN-MAA2. ISBN 3-540-12314-8 (print), 3-540-40036-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061894/.

Smyrnelis:1984:ADE

[Smy84] Emmanuel P. Smyrnélis. Axiome de domination dans les espaces biharmoniques. (French) []. Lecture Notes in Mathematics, 1061: 271–281, 1984. CODEN LNMAA2. ISBN 3-540-13338-0 (print), 3-540-38866-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099029/.

Snaith:1982:APT

[Sna82] Victor P. Snaith. Asymptotic phenomena in the K-theory of group rings. Lecture Notes in Mathematics, 966:261–270, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062179/.

Snaith:1984:UHL

[Sna84] Victor Snaith. Unitary K-homology and the Lichtenbaum– Quillen conjecture on the algebraic K-theory of schemes. Lecture Notes in Mathematics, 1051:128–155, 1984. CODEN LN-MAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075565/.

Sniatycki:1983:CQ

[Śni83a] Jedrzej Śniatycki. Constraints and quantization. Lecture Notes in Mathematics, 1037:301–334, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073178/.

Snider:1983:DRF

[Sni83b] Robert L. Snider. The division ring of fractions of a group ring. Lecture Notes in Mathematics, 1029:325–339, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098938/.

SoaresdeSouza:1982:DSA

[Soa82] Geraldo Soares de Souza. The dyadic special atom space. Lecture Notes in Mathematics, 908:297–299, 1982. CODEN LN-MAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093296/.

Soardi:1983:NLS

[Soa83] Paolo M. Soardi. On nonisotropic Lipschitz spaces. Lecture Notes in Mathematics, 992:115–138, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069155/.

Sommese:1981:HS

[Som81] Andrew John Sommese. Hyperplane sections. Lecture Notes in Mathematics, 862:232–271, 1981. CODEN LNMAA2. ISBN 3-540-10833-5 (print), 3-540-38720-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090894/.

Sommese:1982:ADF

[Som82a] Andrew John Sommese. Ample divisors on 3-folds. Lecture Notes in Mathematics, 947:229–240, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093592/.

Sommese:1982:SEA

[Som82b] Andrew John Sommese. Some examples of C∗ actions. Lecture Notes in Mathematics, 956:118–124, 1982. CODEN LN-MAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101513/.

Soria:1983:ICS

[Sor83] Fernando Soria. Integral characterization of a space generated by blocks. Lecture Notes in Mathematics, 992:201–211, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069160/.

Sot:1982:BDU

[Sot82a] Richard Sot. Brief discussion of unramified and étale homomorphisms. Lecture Notes in Mathematics, 935:45–53, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095803/.

Sot:1982:CSA

[Sot82b] Richard Sot. Characterizations of simple A-algebras in terms of étale homomorphisms; invariance of the property of being a simple algebra under composition and change of base. Lecture Notes in Mathematics, 935:89–102, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095807/.

Sot:1982:DSH

[Sot82c] Richard Sot. Descent of simple homomorphisms and removal of all Noetherian assumptions in chapter 7 and chapter 8. Lecture Notes in Mathematics, 935:103–116, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095808/.

Sot:1982:EAW

[Sot82d] Richard Sot. Every k-algebra a which is essentially of finite type over k and simple is a regular local ring. Lecture Notes in Mathematics, 935:35–44, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095802/.

Sot:1982:FI

[Sot82e] Richard Sot. Fitting ideals. Lecture Notes in Mathematics, 935: 57–72, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095805/.

Sot:1982:KD

[Sot82f] Richard Sot. The Kähler 1-differentials. Lecture Notes in Mathematics, 935:18–34, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095801/.

Sot:1982:PJC

[Sot82g] Richard Sot. Proof of the Jacobian criterion and some characterizations of simple k-algebras and A-algebras. Lecture Notes in Mathematics, 935:73–88, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095806/.

Sot:1982:SMA

[Sot82h] Richard Sot. Simple Morphisms in Algebraic Geometry, volume 935 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 146 pp. LCCN QA3 .L28 no. 935. URL http:

//link.springer.com/book/10.1007/BFb0095799;   http:/
/www.springerlink.com/content/978-3-540-39265-1.

Sot:1982:SMP

[Sot82i] Richard Sot. Simple morphisms of preschemes and translation of previous theorems into the language of preschemes. Lecture Notes in Mathematics, 935:117–127, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095809/.

Sot:1982:SCT

[Sot82j] Richard Sot. Some corollaries to theorem 3.5. Lecture Notes in Mathematics, 935:54–56, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095804/.

Sot:1982:ZTJ

[Sot82k] Richard Sot. The Zariski topology, the Jacobian criterion and examples of simple algebras over a field k. Lecture Notes in Mathematics, 935:1–17, 1982. CODEN LNMAA2. ISBN 3-540-11564-1 (print), 3-540-39265-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095800/.

Sotomayor:1982:SSS

[Sot82l] J. Sotomayor. Structurally stable second order differential equations. Lecture Notes in Mathematics, 957:284–301, 1982. CODEN LNMAA2. ISBN 3-540-11951-5 (print), 3-540-39539-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066244/.

Souriau:1980:GDF

[Sou80] J.-M. Souriau. Groupes différentiels. (French) []. Lecture Notes in Mathematics, 836:91–128, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089728/.

Soule:1981:RTD

[Sou81a] C. Soulé. Rational K-theory of the dual numbers of a ring of algebraic integers. Lecture Notes in Mathematics, 854:402–408, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089531/.

Soule:1981:HAR

[Sou81b] Christophe Soulé. On higher p-adic regulators. Lecture Notes in Mathematics, 854:372–401, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089530/.

Soule:1982:OET

[Sou82] C. Soulé. Operations on étale K-theory. Applications. Lecture Notes in Mathematics, 966:271–303, 1982. CODEN LN-MAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062180/.

Souriau:1983:MUC

[Sou83] Jean-Marie Souriau. Un modèle d’univers confronté aux observations. (French) []. Lecture Notes in Mathematics, 1031:114– 160, 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072115/.

Schulze-Pillot:1980:SDG

[SP80] R. Schulze-Pillot. Struktur von GL_{n} (D). (German) []. Lecture Notes in Mathematics, 778:7–8, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095917.

Sanchez-Palencia:1983:HMS

[SP83] E. Sánchez-Palencia. Homogenization method for the study of composite media. Lecture Notes in Mathematics, 985:192– 214, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062368/.

Spaltenstein:1982:CUS

[Spa82a] Nicolas Spaltenstein. Classes Unipotentes et Sous-groupes de Borel, volume 946 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11585-4 (print), 3-540-39335-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 259 pp. LCCN QA3 .L28 no. 946. URL http:

//link.springer.com/book/10.1007/BFb0096302;   http:/
/www.springerlink.com/content/978-3-540-39335-1.

Spaltenstein:1982:CUF

[Spa82b] Nicolas Spaltenstein. Classes unipotentes. (French) [Unipotent classes]. Lecture Notes in Mathematics, 946:8–41, 1982. CODEN LNMAA2. ISBN 3-540-11585-4 (print), 3-540-39335-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096304/.

Spaltenstein:1982:D

[Spa82c] Nicolas Spaltenstein. Dualité. Lecture Notes in Mathematics, 946: 178–231, 1982. CODEN LNMAA2. ISBN 3-540-11585-4 (print), 3-540-39335-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096306/.

Spaltenstein:1982:NRF

[Spa82d] Nicolas Spaltenstein. Notations et rappels. (French) []. Lecture Notes in Mathematics, 946:1–7, 1982. CODEN LNMAA2. ISBN 3-540-11585-4 (print), 3-540-39335-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096303/.

Spaltenstein:1982:PFV

[Spa82e] Nicolas Spaltenstein. Points fixes sur la variété des sous-groupes de Borel. (French) []. Lecture Notes in Mathematics, 946:42– 177, 1982. CODEN LNMAA2. ISBN 3-540-11585-4 (print), 3-540-39335-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096305/.

Spaltenstein:1982:T

[Spa82f] Nicolas Spaltenstein. Tables. Lecture Notes in Mathematics, 946: 232–250, 1982. CODEN LNMAA2. ISBN 3-540-11585-4 (print), 3-540-39335-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096307/.

Spallek:1983:PARa

[Spa83a] K. Spallek. Produktzerlegung und Äquivalenz von Raumkeimen I Der allgemeine Fall. (German) [Product separation and equivalence of space nuclei I. The general case]. Lecture Notes in Mathematics, 1014:78–100, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072072/.

Spallek:1983:PARb

[Spa83b] K. Spallek. Produktzerlegung und Äquivalenz von Raumkeimen II Der komplexe Fall. (German) []. Lecture Notes in Mathematics, 1014:101–111, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072073/.

Speh:1981:URC

[Spe81a] Birgit Speh. Unitary representations of SL(n, R) and the cohomology of congruence subgroups. Lecture Notes in Mathematics, 880:483–505, 1981. CODEN LNMAA2. ISBN 3-540-10872-6 (print), 3-540-38783-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090420/.

Spehner:1981:PSM

[Spe81b] J.-C. Spehner. Les présentations des sous-monoïdes de rang 3 d’un monoïde libre. (French) []. Lecture Notes in Mathematics, 855:116–155, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089813/.

Spingarn:1984:MAP

[Spi84] Jonathan E. Spingarn. Multifunctions associated with parameterized classes of constrained optimization problems. Lecture Notes in Mathematics, 1091:206–215, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098813/.

Spohn:1984:BHB

[Spo84] Herbert Spohn. Boltzmann hierarchy and Boltzmann equation. Lecture Notes in Mathematics, 1048:207–220, 1984. CODEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071883/.

Springer:1983:SPT

[Spr83] T. A. Springer. Séries de Poincarè dans la théorie des invariants. (French) []. Lecture Notes in Mathematics, 1029:37–54, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098926/.

Srinivasan:1981:BGL

[Sri81] Bhama Srinivasan. The blocks of the general linear group GL(n, q). Lecture Notes in Mathematics, 848:152–160, 1981. CO-DEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090562/.

Schafke:1980:CPL

[SS80a] R. Schäfke and D. Schmidt. Connection problems for linear ordinary differential equations in the complex domain. Lecture Notes in Mathematics, 810:306–317, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089985/.

Smorynski:1980:CEP

[SS80b] C. Smoryński and J. Stavi. Cofinal extension preserves recursive saturation. Lecture Notes in Mathematics, 834:338–345, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090172/.

Sageev:1981:WCS

[SS81a] G. Sageev and S. Shelah. Weak compactness and the structure of Ext(A, Z). Lecture Notes in Mathematics, 874:87–92, 1981. CODEN LNMAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090527/.

Shrikhande:1981:DAM

[SS81b] S. S. Shrikhande and N. M. Singhi. Designs, adjacency multigraphs and embeddings: a survey. Lecture Notes in Mathematics, 885:113–132, 1981. CODEN LNMAA2. ISBN 3-540-11151-4 (print), 3-540-47037-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092257/.

Sally:1983:FTO

[SS83] Paul J. Sally, Jr. and Joseph A. Shalika. The Fourier transform of orbital integrals on SL_{2} over a p-adic field. Lecture Notes in Mathematics, 1041:303–340, 1983. CODEN LNMAA2. ISBN 3-540-12715-1 (print), 3-540-38699-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073152/.

Sternin:1984:CCP

[SS84a] B. Yu. Sternin and V. E. Shatalov. Characteristic Cauchy problem on a complex-analytic manifold. Lecture Notes in Mathematics, 1108:54–76, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099552/.

Sternin:1984:CGL

[SS84b] B. Yu. Sternin and V. E. Shatalov. The contact geometry and linear differential equations. Lecture Notes in Mathematics, 1108: 257–277, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099562/.

Simon:1984:OT

[SSM^{+}84] Barry Simon, L. A. Sahnovich, H. P. Mckean, L. D. Faddeev, B. S. Pavlov, N. G. Makarov, M. S. Birman, T. Ya. Azizov, I. S. Iohvidov, L. de Branges, N. K. Nikol’skii, V. I. Vasyunin, S. N. Naboko, B. Szőkefalvi-Hagy, R. Teodorescu, et al. Operator theory. Lecture Notes in Mathematics, 1043:108–250, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072187/.

Solchenbach:1982:ESN

[SSTW82] Karl Solchenbach, Klaus Stüben, Ulrich Trottenberg, and Kristian Witsch. Efficient solution of a nonlinear heat conduction problem by use of fast elliptic reduction and multigrid methods. Lecture Notes in Mathematics, 968:114–148, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064885/.

Sauer:1983:EFA

[SSW83] N. Sauer, M. G. Stone, and R. H. Weedmark. Every finite algebra with congruence lattice M_{7} has principal congruences. Lecture Notes in Mathematics, 1004:273–292, 1983. CODEN LN-MAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063444/.

Stuben:1982:MMF

[ST82] Klaus Stüben and Ulrich Trottenberg. Multigrid methods: Fundamental algorithms, model problem analysis and applications. Lecture Notes in Mathematics, 960:1–176, 1982. CODEN LN-MAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069928/.

Shee:1984:GCR

[ST84] S. C. Shee and H. H. Teh. A graphical colour-representation of a group. Lecture Notes in Mathematics, 1073:80–82, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073106/.

Stanton:1981:FRC

[Sta81] R. G. Stanton. Further results on covering integers of the form 1+ k2^{N} by primes. Lecture Notes in Mathematics, 884:107–114, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091812/.

Stafford:1982:GME

[Sta82a] J. T. Stafford. Generating modules efficiently over noncommutative rings. Lecture Notes in Mathematics, 924:72–88, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092926/.

Stanton:1982:ONR

[Sta82b] R. G. Stanton. Old and new results on perfect coverings. Lecture Notes in Mathematics, 952:142–149, 1982. CODEN LN-MAA2. ISBN 3-540-11601-X (print), 3-540-39375-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061977/.

Staffans:1984:SGC

[Sta84] O. J. Staffans. Semigroups generated by a convolution equation. Lecture Notes in Mathematics, 1076:209–226, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072778/.

Stegeman:1980:SPS

[Ste80a] J. D. Stegeman. Some problems on spectral synthesis. Lecture Notes in Mathematics, 781:194–203, 1980. CODEN LN-MAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097656/.

Stetter:1980:MAN

[Ste80b] Hans J. Stetter. Modular analysis of numerical software. Lecture Notes in Mathematics, 773:133–145, 1980. CODEN LN-MAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094169/.

Steel:1981:CPP

[Ste81a] John R. Steel. Closure properties of pointclasses. Lecture Notes in Mathematics, 839:147–163, 1981. CODEN LNMAA2. ISBN 3-540-10288-4 (print), 3-540-38422-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090239/.

Stenger:1981:AUT

[Ste81b] F. Stenger. An algorithm for ultrasonic tomography based on inversion of the Helmholtz equation. Lecture Notes in Mathematics, 878:371–406, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090689/.

Steenbrink:1982:PGC

[Ste82a] Joseph Steenbrink. On the Picard group of certain smooth surfaces in weighted projective spaces. Lecture Notes in Mathematics, 961:302–313, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071290/.

Stetter:1982:GEE

[Ste82b] H. J. Stetter. Global error estimation in ordinary initial value problems. Lecture Notes in Mathematics, 968:269–279, 1982.

CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064894/.

Steel:1983:S

[Ste83a] John R. Steel. Scales in L(R). Lecture Notes in Mathematics, 1019:107–156, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071699/.

Steel:1983:SS

[Ste83b] John R. Steel. Scales on Σ_{1}^{1} sets. Lecture Notes in Mathematics, 1019:72–76, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071695/.

Stein:1983:FNS

[Ste83c] E. M. Stein. An H^{1} function with non-summable Fourier expansion. Lecture Notes in Mathematics, 992:193–200, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069159/.

Stell:1983:MDM

[Ste83d] George Stell. Models of disordered media: Some new results, including some new connections between composite-media, fluidstate, and random-flight theories. Lecture Notes in Mathematics, 1035:347–369, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073267/.

Stephen:1983:WDM

[Ste83e] Michael J. Stephen. Waves in disordered media. Lecture Notes in Mathematics, 1035:400–413, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073270/.

Stewart:1983:MCG

[Ste83f] G. W. Stewart. A method for computing the generalized singular value decomposition. Lecture Notes in Mathematics, 973:207–220, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062104/.

Stenger:1984:PSR

[Ste84a] Frank Stenger. Polynomial, sinc and rational function methods for approximating analytic functions. Lecture Notes in Mathematics, 1105:49–72, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072399/.

Stetter:1984:SDC

[Ste84b] Hans J. Stetter. Sequential defect correction for high-accuracy floating-point algorithms. Lecture Notes in Mathematics, 1066: 186–202, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099525/.

Stienstra:1981:TPR

[Sti81] Jan Stienstra. On K_{2} and K_{3} of truncated polynomial rings. Lecture Notes in Mathematics, 854:409–455, 1981. CODEN LN-MAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089532/.

Stinson:1983:RSS

[Sti83] D. R. Stinson. Room squares and subsquares. Lecture Notes in Mathematics, 1036:86–95, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071510/.

Stoica:1980:ALO

[Sto80a] Lucretiu Stoica. The addition of local operators on product spaces. Lecture Notes in Mathematics, 816:50–66, 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091611/.

Stoica:1980:CN

[Sto80b] Lucretiu Stoica. Chapter notation. Lecture Notes in Mathematics, 816:1, 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091607.

Stoica:1980:CV

[Sto80c] Lucretiu Stoica. Chapter V. Lecture Notes in Mathematics, 816: 67–75, 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091612/.

Stoica:1980:FR

[Sto80d] Lucretiu Stoica. Feller resolvents. Lecture Notes in Mathematics, 816:76–95, 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091613/.

Stoica:1980:LO

[Sto80e] Lucretiu Stoica. Local operators. Lecture Notes in Mathematics, 816:1–20, 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091608/.

Stoica:1980:LOM

[Sto80f] Lucretiu Stoica. Local Operators and Markov Processes, volume 816 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 104 pp. LCCN QA3 .L28 no. 816. URL http:/

/link.springer.com/book/10.1007/BFb0091606;   http://
www.springerlink.com/content/978-3-540-70644-1.

Stoica:1980:NPS

[Sto80g] Lucretiu Stoica. Note on the product of semigroups in Hilbert spaces. Lecture Notes in Mathematics, 816:96–99, 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091614/.

Stoica:1980:QS

[Sto80h] Lucretiu Stoica. Quasiharmonic spaces. Lecture Notes in Mathematics, 816:21–43, 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091609/.

Stoica:1980:TPG

[Sto80i] Lucretiu Stoica. Topological properties of global transition function. Lecture Notes in Mathematics, 816:44–49, 1980. CODEN LNMAA2. ISBN 3-540-10028-8 (print), 3-540-70644-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091610/.

Stoll:1982:IVD

[Sto82a] Wilhelm Stoll. Introduction to value distribution theory of meromorphic maps. Lecture Notes in Mathematics, 950:210–359, 1982. CODEN LNMAA2. ISBN 3-540-11596-X (print), 3-540-39366-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061879/.

Stone:1982:PTR

[Sto82b] Charles J. Stone. Potential theory for recurrent symmetric infinitely divisible processes. Lecture Notes in Mathematics, 928: 403–451, 1982. CODEN LNMAA2. ISBN 3-540-11501-3 (print), 3-540-39206-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093235/.

Stoica:1983:SMR

[Sto83a] Lucreţiu Stoica. On sub-Markov resolvents. The restriction to an open set and the Dirichlet problem. Lecture Notes in Mathematics, 1014:228–245, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072082/.

Stoll:1983:AWT

[Sto83b] Wilhelm Stoll. The Ahlfors–Weyl theory of meromorphic maps on parabolic manifolds. Lecture Notes in Mathematics, 981:101– 219, 1983. CODEN LNMAA2. ISBN 3-540-12003-3 (print), 3-540-39480-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066385/.

Stricker:1980:POS

[Str80a] C. Stricker. Projection optionnelle et semimartingales. (French) []. Lecture Notes in Mathematics, 784:112–115, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089474/.

Stricker:1980:PSF

[Str80b] C. Stricker. Prolongement des semimartingales. (French) [Extension of semimartingales]. Lecture Notes in Mathematics, 784: 104–111, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089473/.

Stralka:1981:FCL

[Str81a] Albert Stralka. Fundamental congruences on Lawson semilattices. Lecture Notes in Mathematics, 871:348–359, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089917/.

Strambach:1981:GL

[Str81b] Karl Strambach. Geometry and loops. Lecture Notes in Mathematics, 893:111–147, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091014/.

Streater:1981:EQM

[Str81c] R. F. Streater. Euclidean quantum mechanics and stochastic integrals. Lecture Notes in Mathematics, 851:371–393, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088733/.

Stroock:1981:MCA

[Str81d] Daniel W. Stroock. The Malliavin calculus and its applications. Lecture Notes in Mathematics, 851:394–432, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088734/.

Stricker:1982:ICF

[Str82a] C. Stricker. Les intervalles de constance de ⟨X, X⟩. (French) []. Lecture Notes in Mathematics, 920:219–220, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092785.

Stroock:1982:SMS

[Str82b] Daniel W. Stroock. On the spectrum of Markov semigroups and the existence of invariant measures. Lecture Notes in Mathematics, 923:286–307, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093049/.

Stroock:1983:SAS

[Str83a] Daniel W. Stroock. Some applications of stochastic calculus to partial differential equations. Lecture Notes in Mathematics, 976: 267–382, 1983. CODEN LNMAA2. ISBN 3-540-11987-6 (print), 3-540-39458-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067987/.

Stroyan:1983:IAM

[Str83b] K. D. Stroyan. Infinitesimal analysis of l^{∞} in its Mackey topology. Lecture Notes in Mathematics, 983:204–213, 1983. CODEN LNMAA2. ISBN 3-540-12279-6 (print), 3-540-39602-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065341/.

Stredulinsky:1984:A

[Str84a] Edward W. Stredulinsky. Applications. Lecture Notes in Mathematics, 1074:96–139, 1984. CODEN LNMAA2. ISBN 3-540-13370-4 (print), 3-540-38928-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101273/.

Stredulinsky:1984:BR

[Str84b] Edward W. Stredulinsky. Basic results. Lecture Notes in Mathematics, 1074:32–96, 1984. CODEN LNMAA2. ISBN 3-540-13370-4 (print), 3-540-38928-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101272/.

Stredulinsky:1984:I

[Str84c] Edward W. Stredulinsky. Introduction. Lecture Notes in Mathematics, 1074:1–2, 1984. CODEN LNMAA2. ISBN 3-540-13370-4 (print), 3-540-38928-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101269.

Stredulinsky:1984:N

[Str84d] Edward W. Stredulinsky. Notation. Lecture Notes in Mathematics, 1074:3, 1984. CODEN LNMAA2. ISBN 3-540-13370-4 (print), 3-540-38928-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101270.

Stredulinsky:1984:PA

[Str84e] Edward W. Stredulinsky. Preliminary analysis. Lecture Notes in Mathematics, 1074:3–32, 1984. CODEN LNMAA2. ISBN 3-540-13370-4 (print), 3-540-38928-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101271/.

Stredulinsky:1984:WID

[Str84f] Edward W. Stredulinsky. Weighted Inequalities and Degenerate Elliptic Partial Differential Equations, volume 1074 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13370-4 (print), 3-540-38928-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 143 pp. LCCN QA3 .L28 no. 1074. URL http:/

/link.springer.com/book/10.1007/BFb0101268;   http://
www.springerlink.com/content/978-3-540-38928-6.

Stricker:1984:ADC

[Str84g] C. Stricker. Approximation du crochet de certaines semimartingales continues. (French) []. Lecture Notes in Mathematics, 1059:

144–147, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100038/.

Stricker:1984:CSF

[Str84h] C. Stricker. Caractérisation des semimartingales. (French) []. Lecture Notes in Mathematics, 1059:148–153, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100039/.

Semenov-Tian-Shansky:1982:GTA

[STS82] M. A. Semenov-Tian-Shansky. Group-theoretical aspects of completely integrable systems. Lecture Notes in Mathematics, 970: 173–185, 1982. CODEN LNMAA2. ISBN 3-540-11972-8 (print), 3-540-39418-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066030/.

Stuhler:1980:NSA

[Stu80] U. Stuhler. Normen und Spuren in Assoziativen Algebren. (German) []. Lecture Notes in Mathematics, 778:21–32, 1980. CODEN LNMAA2. ISBN 3-540-09747-3 (print), 3-540-38599-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095919/.

Stuhler:1982:CRA

[Stu82] Ulrich Stuhler. On the cohomology of SL_{n} over rings of algebraic functions. Lecture Notes in Mathematics, 967:316–359, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061909/.

Succi:1983:PCF

[Suc83] Francesco Succi. Pseudovariétés complexes. (French) [Complex pseudovarieties]. Lecture Notes in Mathematics, 1014:112–121, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072074/.

Sullivan:1981:GPH

[Sul81a] Dennis Sullivan. Growth of positive harmonic functions and Kleinian group limit sets of zero planar measure and Hausdorf dimension two. In celebration of Nico Kuiper’s sixtieth birthday. Lecture Notes in Mathematics, 894:127–144, 1981. CODEN LNMAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096228/.

Sullivan:1981:TTG

[Sul81b] Dennis Sullivan. Travaux de Thurston sur les groupes quasifuchsiens et les variétés hyperboliques de dimension 3 fibrées sur s^{1}. (French) []. Lecture Notes in Mathematics, 842:196–214, 1981. CODEN LNMAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089935/.

Sundararaman:1982:CHT

[Sun82] D. Sundararaman. Compact Hausdorff transversally holomorphic foliations. Lecture Notes in Mathematics, 950:360–376, 1982. CODEN LNMAA2. ISBN 3-540-11596-X (print), 3-540-39366-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061880/.

Suslin:1982:MST

[Sus82a] A. A. Suslin. Mennicke symbols and their applications in the ktheory of fields. Lecture Notes in Mathematics, 966:334–356, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062182/.

Suslin:1982:SAT

[Sus82b] A. A. Suslin. Stability in algebraic K-theory. Lecture Notes in Mathematics, 966:304–333, 1982. CODEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062181/.

Suwa:1983:RCA

[Suw83] Noriyuki Suwa. De Rham cohomology of algebraic surfaces with q = −p_{a} in char. p. Lecture Notes in Mathematics, 1016:73– 85, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099958/.

Sivaramakrishnan:1982:CEC

[SV82] R. Sivaramakrishnan and B. K. Vijayan. On certain expoential and character sums. Lecture Notes in Mathematics, 938:138– 156, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097181/.

Sibony:1980:SRG

[SW80] Nessim Sibony and Pit-Mann Wong. Some results on global analytic sets. Lecture Notes in Mathematics, 822:221–237, 1980. CODEN LNMAA2. ISBN 3-540-10241-8 (print), 3-540-38320-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097760/.

Stanton:1981:CSN

[SW81] R. G. Stanton and H. C. Williams. Computation of some number-theoretic coverings. Lecture Notes in Mathematics, 884: 8–13, 1981. CODEN LNMAA2. ISBN 3-540-10883-1 (print), 3-540-38792-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091803/.

Strauss:1983:AES

[SW83] Walter Strauss and Stephen Wainger. An application of L^{p} estimates to scattering theory. Lecture Notes in Mathematics, 992: 91–100, 1983. CODEN LNMAA2. ISBN 3-540-12299-0 (print), 3-540-39885-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069153/.

Szynal:1984:PTV

[SW84] Dominik Szynal and Aleksander Weron, editors. Probability Theory on Vector Spaces III: Proceedings of a Conference held in Lublin, Poland, August 24–31, 1983, volume 1080 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LN-MAA2. ISBN 0-387-13388-7, 3-540-13388-7 (print), 3-540-38939-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1080; QA3 .L28 no. 1080 QA273.43. URL http:/

/link.springer.com/book/10.1007/BFb0099781;   http://
www.springerlink.com/content/978-3-540-38939-2.

Spreen:1984:EOT

[SY84] Dieter Spreen and Paul Young. Effective operators in a topological setting. Lecture Notes in Mathematics, 1104:437–451, 1984. CODEN LNMAA2. ISBN 3-540-13901-X (print), 3-540-39119-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099496/.

Salce:1981:PF

[SZ81a] Luigi Salce and Paolo Zanardo. On a paper of I. Fleischer. Lecture Notes in Mathematics, 874:76–86, 1981. CODEN LN-MAA2. ISBN 3-540-10855-6 (print), 3-540-38767-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090526/.

Sehgal:1981:SGR

[SZ81b] Sudarshan K. Sehgal and Hans Zassenhaus. On the supercentre of a group and its ring theoretic generalization. Lecture Notes in Mathematics, 882:117–144, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092489/.

Szendrei:1981:CST

[Sze81] Mária B. Szendrei. On closed sets of term functions on bands. Lecture Notes in Mathematics, 855:156–181, 1981. CODEN LN-MAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089814/.

Szeto:1982:SRA

[Sze82] George Szeto. Splitting rings for Azumaya quaternion algebras. Lecture Notes in Mathematics, 917:118–125, 1982. CODEN LN-MAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092231/.

Szpirglas:1980:PRR

[Szp80] A. Szpirglas. Platitude des revetements ramifies. Lecture Notes in Mathematics, 807:167–179, 1980. CODEN LNMAA2. ISBN 3-540-10015-6 (print), 3-540-38146-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090879/.

Szucs:1980:CMS

[Szű80] András Szűcs. Cobordism of maps with simplest singularities. Lecture Notes in Mathematics, 788:223–244, 1980. CODEN LN-MAA2. ISBN 3-540-09968-9 (print), 3-540-39187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099247/.

Szulga:1983:MT

[Szu83] A. Szulga. On the metrics of the type ζ. Lecture Notes in Mathematics, 982:238–243, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082074/.

Taft:1982:NSD

[Taf82] Earl J. Taft. Noncocommutative sequences of divided powers. Lecture Notes in Mathematics, 933:203–209, 1982. CODEN LN-MAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093363/.

Taha:1982:ASC

[Tah82] Faleh Taha. Algèbres simples centrales sur les corps ultra-produits de corps p-adiques. (French) []. Lecture Notes in Mathematics, 924:89–128, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092927/.

Takens:1980:MBN

[Tak80] Floris Takens. Moduli and bifurcations: Non-transversal intersections of invariant manifolds of vectorfields. Lecture Notes in Mathematics, 799:368–384, 1980. CODEN LNMAA2. ISBN 3-540-09986-7 (print), 3-540-39251-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089325/.

Takakuwa:1984:VAB

[Tak84] Shōichirō Takakuwa. A variational approach to the boundary value problem for hypersurfaces with prescribed mean curvature. Lecture Notes in Mathematics, 1090:50–63, 1984. CODEN LNMAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101566/.

Talagrand:1980:NEC

[Tal80] M. Talagrand. Non-existence de certaines sections mesurables et contre-exemples en théorie du relèvement. (French) []. Lecture Notes in Mathematics, 794:166–175, 1980. CODEN LN-MAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088223/.

Talagrand:1982:RFC

[Tal82a] Michel Talagrand. Sur les résultats de Feyel concernant les épaisseurs. (French) []. Lecture Notes in Mathematics, 920:1– 7, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092766/.

Talay:1982:HDS

[Tal82b] Denis Talay. How to discretize stochastic differential equations. Lecture Notes in Mathematics, 972:276–292, 1982. CODEN LN-MAA2. ISBN 3-540-11976-0 (print), 3-540-39431-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064866/.

Talhaoui:1983:PNS

[Tal83] A. Talhaoui. Le problème ¯∂ de Neumann et ses estimations souselliptiques. (French) []. Lecture Notes in Mathematics, 1028:266– 317, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071687/.

Tamari:1982:GTA

[Tam82a] Dov Tamari. A graphic theory of associativity and wordchain patterns. Lecture Notes in Mathematics, 969:302–320, 1982. CODEN LNMAA2. ISBN 3-540-11971-X (print), 3-540-39380-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063002/.

Tammi:1982:RS

[Tam82b] Olli Tammi. ℜ(a_{3} + λa_{2}) in S(b). Lecture Notes in Mathematics, 913:60–98, 1982. CODEN LNMAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094064/.

Tammi:1982:CSC

[Tam82c] Olli Tammi. Consequences for the second coefficient body in S_{R} (b). Lecture Notes in Mathematics, 913:155–165, 1982. CODEN LNMAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094069/.

Tammi:1982:DED

[Tam82d] Olli Tammi. Determination of extremal domains. Lecture Notes in Mathematics, 913:99–121, 1982. CODEN LNMAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094065/.

Tammi:1982:DCB

[Tam82e] Olli Tammi. Determination of the coefficient body (a_{2}, a_{3}) by aid of a perfect square method. Lecture Notes in Mathematics, 913:1–48, 1982. CODEN LNMAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094062/.

Tammi:1982:EPB

[Tam82f] Olli Tammi. Extremum Problems for Bounded Univalent Functions II, volume 913 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 168 pp. LCCN QA3 .L28 no. 913. URL http:

//link.springer.com/book/10.1007/BFb0094061;   http:/
/www.springerlink.com/content/978-3-540-39012-1.

Tammi:1982:GPI

[Tam82g] Olli Tammi. A generalized power-inequality for S_{R} (b)-functions. Lecture Notes in Mathematics, 913:144–154, 1982. CODEN LN-MAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094068/.

Tammi:1982:LC

[Tam82h] Olli Tammi. The linear combination a_{3} + λa_{2} in S_{R}(b). Lecture Notes in Mathematics, 913:49–59, 1982. CODEN LN-MAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094063/.

Tammi:1982:LIF

[Tam82i] Olli Tammi. A Löwner-identity for finding the power-inequality for a second coefficient body in S_{R}(b). Lecture Notes in Mathematics, 913:122–136, 1982. CODEN LNMAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094066/.

Tammi:1982:SDE

[Tam82j] Olli Tammi. Schiffer’s differential equation for functions 2: 3. Lecture Notes in Mathematics, 913:137–143, 1982. CODEN LN-MAA2. ISBN 3-540-11200-6 (print), 3-540-39012-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094067/.

Tang:1981:WRH

[Tan81a] A. Tang. Wadge reducibility and Hausdorff difference hierarchy in P ω. Lecture Notes in Mathematics, 871:360–371, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089918/.

Tannenbaum:1981:ART

[Tan81b] Allen Tannenbaum. Algebraic realization theory. Lecture Notes in Mathematics, 845:76–103, 1981. CODEN LNMAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093324/.

Tannenbaum:1981:BGU

[Tan81c] Allen Tannenbaum. The Brauer group and unirationality: an example of Artin–Mumford. Lecture Notes in Mathematics, 844: 103–128, 1981. CODEN LNMAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090479/.

Tannenbaum:1981:FSS

[Tan81d] Allen Tannenbaum. Feedback and stabilization of systems with parameter uncertainty. Lecture Notes in Mathematics, 845:119– 149, 1981. CODEN LNMAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093326/.

Tannenbaum:1981:GML

[Tan81e] Allen Tannenbaum. Global moduli of linear time-invariant dynamical systems. Lecture Notes in Mathematics, 845:48–64, 1981. CODEN LNMAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093322/.

Tannenbaum:1981:IST

[Tan81f] Allen Tannenbaum. Invariance and System Theory: Algebraic and Geometric Aspects, volume 845 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 161 pp. LCCN QA3 .L28 no. 845.

URL http://link.springer.com/book/10.1007/BFb0093318;
 http://www.springerlink.com/content/978-3-540-38536-
3.

Tannenbaum:1981:ITO

[Tan81g] Allen Tannenbaum. Invariant theory and orbit space problems. Lecture Notes in Mathematics, 845:36–47, 1981. CODEN LN-MAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093321/.

Tannenbaum:1981:LML

[Tan81h] Allen Tannenbaum. Local moduli of linear time-invariant dynamical systems. Lecture Notes in Mathematics, 845:65–75, 1981. CODEN LNMAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093323/.

Tannenbaum:1981:GRT

[Tan81i] Allen Tannenbaum. On the geometry of rational transfer functions. Lecture Notes in Mathematics, 845:104–118, 1981. CODEN LNMAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093325/.

Tannenbaum:1981:SBA

[Tan81j] Allen Tannenbaum. Some basic algebraic geometry. Lecture Notes in Mathematics, 845:1–17, 1981. CODEN LNMAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093319/.

Tannenbaum:1981:SBS

[Tan81k] Allen Tannenbaum. Some basic system theory. Lecture Notes in Mathematics, 845:18–35, 1981. CODEN LNMAA2. ISBN 3-540-10565-4 (print), 3-540-38536-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093320/.

Tanre:1983:AHF

[Tan83a] Daniel Tanré. Algèbre homotopique. (French) []. Lecture Notes in Mathematics, 1025:39–72, 1983. CODEN LNMAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071485/.

Tanre:1983:CWP

[Tan83b] Daniel Tanré. Crochets de Whitehead et produits de Massey d’ordre supérieur. (French) []. Lecture Notes in Mathematics, 1025:109–132, 1983. CODEN LNMAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071488/.

Tanre:1983:HRM

[Tan83c] Daniel Tanré. Homotopie Rationnelle: Modèles de Chen, Quillen, Sullivan, volume 1025 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 211 pp. LCCN QA3 .L28 no. 1025. URL http:

//link.springer.com/book/10.1007/BFb0071482;   http:/
/www.springerlink.com/content/978-3-540-38712-1.

Tanre:1983:IAF

[Tan83d] Daniel Tanré. Inventaire algébrique. (French) []. Lecture Notes in Mathematics, 1025:1–22, 1983. CODEN LNMAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071483/.

Tanre:1983:LEA

[Tan83e] Daniel Tanré. Liaison entre algèbres commutatives et algèbres de Lie différentielles graduées. (French) []. Lecture Notes in Mathematics, 1025:23–38, 1983. CODEN LNMAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071484/.

Tanre:1983:MCE

[Tan83f] Daniel Tanré. Modèle de Chen d’un espace. (French) []. Lecture Notes in Mathematics, 1025:100–108, 1983. CODEN LN-MAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071487/.

Tanre:1983:MQFb

[Tan83g] Daniel Tanré. Modèle de Quillen d’une fibration et classifiant. (French) []. Lecture Notes in Mathematics, 1025:157–197, 1983. CODEN LNMAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071490/.

Tanre:1983:MQFa

[Tan83h] Daniel Tanré. Modèle de Quillen d’une fibration et suite spectrale d’Eilenberg–Moore. (French) []. Lecture Notes in Mathematics, 1025:133–156, 1983. CODEN LNMAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071489/.

Tanre:1983:MQS

[Tan83i] Daniel Tanré. Modèles de Quillen et de Sullivan d’un espace. (French) []. Lecture Notes in Mathematics, 1025:73–99, 1983. CODEN LNMAA2. ISBN 3-540-12726-7 (print), 3-540-38712-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071486/.

Tirrell:1984:OMP

[TAP84] Matthew Tirrell, Douglas Adolf, and Stephen Prager. Orientation and motion at a polymer-polymer interface: Interdiffusion of fluorescent-labelled macromolecules. Lecture Notes in Mathematics, 1063:37–45, 1984. CODEN LNMAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072151/.

Tarski:1982:PIM

[Tar82] Jan Tarski. Path integrals over manifolds. Lecture Notes in Mathematics, 905:229–239, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092440/.

Tauvel:1980:ADP

[Tau80] Patrice Tauvel. Sur l’application de Dixmier pour les algèbres de Lie resolubles. (French) []. Lecture Notes in Mathematics, 795: 161–171, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090117/.

Taussky:1981:SFC

[Tau81] Olga Taussky. Some facts concerning integral representations of ideals in an algebraic number field. Lecture Notes in Mathematics, 882:145–158, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092490/.

Taylor:1981:FCL

[Tay81] M. J. Taylor. Fröhlich’s conjecture, logarithmic methods and Swan modules. Lecture Notes in Mathematics, 882:207–218, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092494/.

Taylor:1982:LDD

[Tay82] M. J. Taylor. A logarithmic description of determinants in group rings. Lecture Notes in Mathematics, 967:360–370, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061910/.

Taylor:1983:EEP

[Tay83] B. A. Taylor. An estimate for an extremal plurisubharmonic function on C^{n}. Lecture Notes in Mathematics, 1028:318–328, 1983. CODEN LNMAA2. ISBN 3-540-12731-3 (print), 3-540-38724-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071688/.

Tay:1984:RMG

[Tay84] Tiong-Seng Tay. Rigidity of multi-graphs II. Lecture Notes in Mathematics, 1073:129–134, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073111/.

Teissier:1980:RSF

[Tei80a] B. Teissier. Résolution simultanée — I — familles de courbes. (French) [Simultaneous resolution — I curve families]. Lecture Notes in Mathematics, 777:71–81, 1980. CODEN LN-MAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085880/.

Teissier:1980:RSI

[Tei80b] B. Teissier. Résolution simultanée — II — résolution simultanée et cycles évanescents. (French) []. Lecture Notes in Mathematics, 777:82–146, 1980. CODEN LNMAA2. ISBN 3-540-09746-5 (print), 3-540-38586-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085881/.

Teicher:1981:ACB

[Tei81a] Henry Teicher. Almost certain behavior of row sums of double arrays. Lecture Notes in Mathematics, 861:155–165, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097322/.

Teissier:1981:VTP

[Tei81b] Bernard Teissier. Variétés toriques et polytopes. (French) []. Lecture Notes in Mathematics, 901:71–84, 1981. CODEN LN-MAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097190/.

Teissier:1982:VPI

[Tei82] Bernard Teissier. Variétés polaires II. multiplicitès polaires, sections planes, et conditions de Whitney. (French) []. Lecture Notes in Mathematics, 961:314–491, 1982. CODEN LNMAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071291/.

Tenenbaum:1984:RRT

[Ten84] G. Tenenbaum. Un rapport sur de recents travaux en théorie analytique des nombres. (French) []. Lecture Notes in Mathematics, 1068:231–242, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099456/.

Testard:1983:AME

[Tes83] D. Testard. Almost Mathieu equation for small and big coupling constants. Lecture Notes in Mathematics, 1031:161–171, 1983. CODEN LNMAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072116/.

Tahara:1984:CGF

[TH84] Ken-Ichi Tahara and Akinori Hosomi. On the circle groups of finite nilpotent rings. Lecture Notes in Mathematics, 1098:161– 179, 1984. CODEN LNMAA2. ISBN 3-540-13890-0 (print), 3-540-39102-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099672/.

Thas:1983:GFP

[Tha83] J. A. Thas. Geometries in finite projective spaces: Recent results. Lecture Notes in Mathematics, 1036:96–110, 1983. CODEN LNMAA2. ISBN 3-540-12708-9 (print), 3-540-38694-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071511/.

Thierrin:1981:DSC

[Thi81] G. Thierrin. Decomposition of some classes of subsets in a semigroup. Lecture Notes in Mathematics, 855:182–188, 1981. CO-DEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089815/.

Thibault:1984:CMC

[Thi84] L. Thibault. Continuity of measurable convex multifunctions. Lecture Notes in Mathematics, 1091:216–224, 1984. CODEN LN-MAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098814/.

Thom:1981:PNS

[Tho81a] R. Thom. Sur le problème des normales à une sphère convexe, et l’approximation des applications “collapsantes”. (French) []. Lecture Notes in Mathematics, 894:145–153, 1981. CODEN LNMAA2. ISBN 3-540-11167-0 (print), 3-540-38641-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096229/.

Thomas:1981:IRT

[Tho81b] C. B. Thomas. Integral representations in the theory of finite cw-complexes. Lecture Notes in Mathematics, 882:269–286, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092500/.

Thomsen:1982:USM

[Tho82] Per Grove Thomsen. The use of sparse matrix techniques in ODEcodes. Lecture Notes in Mathematics, 968:301–309, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064897/.

Thomas:1983:ARA

[Tho83] Marc P. Thomas. Approximation in the radical algebra l^{1}(ω_{n}) when ω_{n} is star-shaped. Lecture Notes in Mathematics, 975:258– 272, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064555/.

Thomas:1984:IDI

[Tho84a] Erik G. F. Thomas. Invariant Daniell integrals. (German) []. Lecture Notes in Mathematics, 1089:146–152, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072610/.

Thomee:1984:M

[Tho84b] Vidar Thomée. The H^{1} and H^{−1} methods. Lecture Notes in Mathematics, 1054:187–204, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071802/.

Thomee:1984:CDSa

[Tho84c] Vidar Thomée. Completely discrete schemes for the homogeneous equation. Lecture Notes in Mathematics, 1054:92–105, 1984. CO-DEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071797/.

Thomee:1984:CDSb

[Tho84d] Vidar Thomée. Completely discrete schemes for the inhomogeneous equation. Lecture Notes in Mathematics, 1054:106– 125, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071798/.

Thomee:1984:GFE

[Tho84e] Vidar Thomée. Galerkin Finite Element Methods for Parabolic Problems, volume 1054 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 237 pp. LCCN QA3 .L28 no. 1054.

URL http://link.springer.com/book/10.1007/BFb0071790;
 http://www.springerlink.com/content/978-3-540-38793-
0.

Thomee:1984:MNE

[Tho84f] Vidar Thomée. Maximum-norm estimates. Lecture Notes in Mathematics, 1054:62–75, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071795/.

Thomee:1984:MLM

[Tho84g] Vidar Thomée. The method of lumped masses. Lecture Notes in Mathematics, 1054:166–186, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071801/.

Thomee:1984:MM

[Tho84h] Vidar Thomée. A mixed method. Lecture Notes in Mathematics, 1054:205–220, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071803/.

Thomee:1984:NNE

[Tho84i] Vidar Thomée. Negative norm estimates and superconvergence. Lecture Notes in Mathematics, 1054:76–91, 1984. CODEN LN-MAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071796/.

Thomee:1984:NP

[Tho84j] Vidar Thomée. A nonlinear problem. Lecture Notes in Mathematics, 1054:149–165, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071800/.

Thomee:1984:PEM

[Tho84k] Vidar Thomée. Parabolic equations with more general elliptic operators. Lecture Notes in Mathematics, 1054:49–61, 1984. CO-DEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071794/.

Thomee:1984:SMB

[Tho84l] Vidar Thomée. Semidiscrete methods based on more general approximations of the elliptic problem. Lecture Notes in Mathematics, 1054:17–32, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071792/.

Thomee:1984:SP

[Tho84m] Vidar Thomée. A singular problem. Lecture Notes in Mathematics, 1054:221–235, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071804/.

Thomee:1984:SNS

[Tho84n] Vidar Thomée. Smooth and non-smooth data error estimates for the homogeneous equation. Lecture Notes in Mathematics, 1054: 33–48, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071793/.

Thomee:1984:SGM

[Tho84o] Vidar Thomée. The standard Galerkin method. Lecture Notes in Mathematics, 1054:1–16, 1984. CODEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071791/.

Thomee:1984:TDD

[Tho84p] Vidar Thomée. Time discretization by the discontinuous Galerkin method. Lecture Notes in Mathematics, 1054:126–148, 1984. CO-DEN LNMAA2. ISBN 3-540-12911-1 (print), 3-540-38793-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071799/.

Tignol:1981:CIN

[Tig81] J.-P. Tignol. Corps à involution neutralisés par une extension abélienne Élémentaire. (French) []. Lecture Notes in Mathematics, 844:1–34, 1981. CODEN LNMAA2. ISBN 3-540-10562-X (print), 3-540-38531-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090475/.

Tignol:1982:DAD

[Tig82] J.-P. Tignol. Sur les decompositions des algèbres à division en produit tensoriel d’algèbres cycliques. (French) []. Lecture Notes in Mathematics, 917:126–145, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092232/.

Tits:1981:GCP

[Tit81] Jacques Tits. Groupes à croissance polynômiale. (French) []. Lecture Notes in Mathematics, 901:176–188, 1981. CODEN LN-MAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097197/.

Tugue:1981:IPE

[TN81] Tosiyuki Tugué and Hisao Nomoto. Independence of a problem in elementary analysis from set theory. Lecture Notes in Mathematics, 891:307–321, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090987/.

Taylor:1984:IBM

[TNV^{+}84] B. A. Taylor, N. K. Nikol’skii, S. A. Vinogradov, J. Bruna, N. A. Shirokov, A. F. Leontiev, Yu. A. Brudnyi, and Misha Zafran. Interpolation, bases, multiplers. Lecture Notes in Mathematics, 1043:563–588, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072193/.

Todd:1982:IPL

[Tod82] Michael J. Todd. An introduction to piecewise-linear homotopy algorithms for solving systems of equations. Lecture Notes in Mathematics, 965:149–202, 1982. CODEN LNMAA2. ISBN 3-540-11967-1 (print), 3-540-39558-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063203/.

Tognoli:1981:AAM

[Tog81] A. Tognoli. Algebraic approximation of manifolds and spaces. Lecture Notes in Mathematics, 842:73–94, 1981. CODEN LN-MAA2. ISBN 3-540-10292-2 (print), 3-540-38437-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089929/.

Tomisaki:1982:DFA

[Tom82] Matsuyo Tomisaki. Dirichlet forms associated with direct product diffusion processes. Lecture Notes in Mathematics, 923:76– 119, 1982. CODEN LNMAA2. ISBN 3-540-11484-X (print), 3-540-39155-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093037/.

Tonev:1983:CBA

[Ton83] Toma V. Tonev. Commutative Banach algebras and analytic functions of countable-many variables. Lecture Notes in Mathematics, 1014:122–128, 1983. CODEN LNMAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072075/.

Topsoe:1980:TTG

[Top80] Flemming Topsøe. Thin trees and geometrical criteria for Lebesgue nullsets. Lecture Notes in Mathematics, 794:57–78, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088214/.

Toppila:1983:INT

[Top83] Sakari Toppila. An introduction to Nevanlinna theory. Lecture Notes in Mathematics, 981:1–12, 1983. CODEN LNMAA2. ISBN 3-540-12003-3 (print), 3-540-39480-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066381/.

teRiele:1984:NVL

[tR84] Herman J. J. te Riele. New very large amicable pairs. Lecture Notes in Mathematics, 1068:210–215, 1984. CODEN LN-MAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099454/.

Trautman:1982:YMT

[Tra82a] Andrzej Trautman. Yang–Mills theory and gravitation: a comparison. Lecture Notes in Mathematics, 926:179–189, 1982. CODEN LNMAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092662/.

Trautmann:1982:OVY

[Tra82b] Gunther Trautmann. Olomorphic vectorbundles and Yang–Mills fields. Lecture Notes in Mathematics, 950:377–401, 1982. CODEN LNMAA2. ISBN 3-540-11596-X (print), 3-540-39366-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061881/.

Traynor:1984:MFT

[Tra84] Tim Traynor. Modular functions and their Fréchet–Nikodým topologies. Lecture Notes in Mathematics, 1089:171–180, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072613/.

Trefethen:1984:SBE

[Tre84] Lloyd N. Trefethen. Square blocks and equioscillation in the Padé, Walsh, and CF tables. Lecture Notes in Mathematics, 1105:170– 181, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072410/.

Triggiani:1984:BFR

[Tri84] R. Triggiani. A^{ϵ}-bounded, finite rank perturbations of s.c. group generators A: Counterexamples to generation and to another condition for well-posedness. Lecture Notes in Mathematics, 1076: 227–253, 1984. CODEN LNMAA2. ISBN 3-540-13376-3 (print), 3-540-38932-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072779/.

Trnkova:1984:APP

[Trn84] Věra Trnková. Arithmetical properties of the product of homeomorphism types of spaces. Lecture Notes in Mathematics, 1060: 84–94, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099924/.

Troesch:1982:EME

[Tro82a] Albert Troesch. Étude macroscopique de l’équation de van der Pol. (French) []. Lecture Notes in Mathematics, 942:136–144, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094744/.

Trotman:1982:RSS

[Tro82b] David Trotman. Regular stratifications and sufficiency of jets. Lecture Notes in Mathematics, 961:492–500, 1982. CODEN LN-MAA2. ISBN 3-540-11969-8 (print), 3-540-39367-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071292/.

Troallic:1983:SSP

[Tro83] Jean-Pierre Troallic. Semigroupes semitopologiques et presquepériodicité. (French) []. Lecture Notes in Mathematics, 998:239– 251, 1983. CODEN LNMAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062032/.

Takeuti:1981:HVU

[TT81] Gaisi Takeuti and Satoko Titani. Heyting valued universes of intuitionistic set theory. Lecture Notes in Mathematics, 891:189– 306, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090986/.

Tulczyjew:1980:ELR

[Tul80] W. M. Tulczyjew. The Euler–Lagrange resolution. Lecture Notes in Mathematics, 836:22–48, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089725/.

Turner:1982:TNA

[Tur82] Peter R. Turner, editor. Topics in Numerical Analysis: Proceedings of the S.E.R.C. Summer School, Lancaster, July 19– August 21, 1981, volume 965 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11967-1 (print), 3-540-39558-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 965. URL http:

//link.springer.com/book/10.1007/BFb0063198;   http:/
/www.springerlink.com/content/978-3-540-39558-4.

Tricerri:1984:SDA

[TV84] F. Tricerri and L. Vanhecke. Self-dual and anti-self-dual homogeneous structures. Lecture Notes in Mathematics, 1045:186– 194, 1984. CODEN LNMAA2. ISBN 3-540-12882-4 (print), 3-540-38766-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072179/.

Takahashi:1981:PFO

[TW81] Y. Takahashi and S. Watanabe. The probability functionals (Onsager–Machlup functions) of diffusion processes. Lecture Notes in Mathematics, 851:433–463, 1981. CODEN LN-MAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088735/.

Thron:1982:MCF

[TW82a] W. J. Thron and Haakon Waadeland. Modifications of continued fractions, a survey. Lecture Notes in Mathematics, 932:38– 66, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093304/.

Thron:1982:CTC

[TW82b] W. J. Thron and Haakon Waadeland. On a certain transformation of continued fractions. Lecture Notes in Mathematics, 932: 225–240, 1982. CODEN LNMAA2. ISBN 3-540-11567-6 (print), 3-540-39276-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093316/.

Truman:1984:SAA

[TW84] Aubrey Truman and David Williams, editors. Stochastic Analysis and Applications: Proceedings of the International Conference held in Swansea, April 11–15, 1983, volume 1095 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1095.

URL http://link.springer.com/book/10.1007/BFb0099117;
 http://www.springerlink.com/content/978-3-540-39103-
6.

Ueno:1982:BGA

[Uen82] Kenji Ueno. Bimeromorphic geometry of algebraic and analytic threefolds. Lecture Notes in Mathematics, 947:1–34, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/BFb0093584/;   http://link.springer.com/content/
pdf/bfm:978-3-540-39342-9/1.pdf.

Uesu:1981:ITT

[Ues81] Tadahiro Uesu. Intuitionistic theories and toposes. Lecture Notes in Mathematics, 891:323–358, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090988/.

Uhlenbeck:1982:EHM

[Uhl82] Karen K. Uhlenbeck. Equivariant harmonic maps into spheres. Lecture Notes in Mathematics, 949:146–158, 1982. CODEN LN-MAA2. ISBN 3-540-11595-1 (print), 3-540-39360-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069763/.

Ulyanov:1984:SAO

[Ul’84] V. M. Ul’yanov. The sequential absolute and the other analogs of the absolute. Lecture Notes in Mathematics, 1060:95–104, 1984. CODEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099925/.

Ullom:1981:RRI

[Ull81] Stephen V. Ullom. Ratios of rings of integers as Galios modules. Lecture Notes in Mathematics, 882:240–246, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092496/.

Ullom:1982:CAC

[Ull82] Stephen V. Ullom. Character action on the class group of Fröhlich. Lecture Notes in Mathematics, 967:371–375, 1982. CODEN LN-MAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061911/.

Ungar:1982:ESL

[Ung82] Tomas Ungar. Elementary systems for Lie algebra bundle actions. Lecture Notes in Mathematics, 905:66–89, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092428/.

vanAsselt:1982:MGM

[vA82] E. J. van Asselt. The multi grid method and artificial viscosity. Lecture Notes in Mathematics, 960:313–326, 1982. CODEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069931/.

Valadier:1984:SBB

[Val84a] M. Valadier. Some Bang–Bang theorems. Lecture Notes in Mathematics, 1091:225–234, 1984. CODEN LNMAA2. ISBN 3-540-13882-X (print), 3-540-39083-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098815/.

Valla:1984:STC

[Val84b] Giuseppe Valla. On set-theoretic complete intersections. Lecture Notes in Mathematics, 1092:85–101, 1984. CODEN LN-MAA2. ISBN 3-540-13884-6 (print), 3-540-39089-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099358/.

Valli:1984:FBP

[Val84c] Alberto Valli. Free boundary problems for compressible viscous fluids. Lecture Notes in Mathematics, 1047:175–187, 1984. CO-DEN LNMAA2. ISBN 3-540-12893-X (print), 3-540-38773-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072331/.

VanOystaeyen:1980:GAA

[Van80] F. Van Oystaeyen. Graded Azumaya algebras and Brauer groups. Lecture Notes in Mathematics, 825:158–171, 1980. CODEN LN-MAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089129/.

Vangipuram:1982:PCC

[Van82] S. Vangipuram. Partitions with congruence conditions and color restrictions. Lecture Notes in Mathematics, 938:157–169, 1982. CODEN LNMAA2. ISBN 3-540-11568-4 (print), 3-540-39279-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097182/.

VanDenEssen:1983:STF

[Van83a] Arno Van Den Essen. A study of the torsion-free part of the cokernel from the operator \frac{∂}{∂x_{n}} acting on a D_{n}-module. Lecture Notes in Mathematics, 1029:99–119, 1983. CODEN LNMAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098928/.

VanDooren:1983:RSD

[Van83b] Paul Van Dooren. Reducing subspaces: Definitions, properties and algorithms. Lecture Notes in Mathematics, 973:58–73, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062094/.

VanLoan:1983:GSA

[Van83c] Charles Van Loan. A generalized SVD analysis of some weighting methods for equality constrained least squares. Lecture Notes in Mathematics, 973:245–262, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062106/.

Vanderbauwhede:1984:SBE

[Van84] André Vanderbauwhede. Stability of bifurcating equilibria and the principle of reduced stability. Lecture Notes in Mathematics, 1057:209–223, 1984. CODEN LNMAA2. ISBN 3-540-12931-6 (print), 3-540-38812-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098598/.

Varadarajan:1984:APE

[Var84] V. S. Varadarajan. Asymptotic properties of eigenvalues and eigenfunctions of invariant differential operators on symmetric and locally symmetric spaces. Lecture Notes in Mathematics, 1077:396–436, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072347/.

Vaserstein:1981:NSR

[Vas81] L. N. Vaserstein. On the normal subgroups of GL_{n} over a ring. Lecture Notes in Mathematics, 854:456–465, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089533/.

vanderAa:1983:A

[vdA83] Els van der Aa. Appendix. Lecture Notes in Mathematics, 985: 184–191, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062367/.

vanderBurgh:1983:ATF

[vdB83] A. H. P. van der Burgh. An asymptotic theory for the free vibrations of an iced two-conductor bundled transmission line. Lecture Notes in Mathematics, 985:413–430, 1983. CODEN LN-MAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062379/.

vandenDries:1980:SMT

[vdD80] Lou van den Dries. Some model theory and number theory for models of weak systems of arithmetic. Lecture Notes in Mathematics, 834:346–362, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090173/.

vanderHouwen:1982:TIP

[vdH82] P. J. van der Houwen. On the time integration of parabolic differential equations. Lecture Notes in Mathematics, 912:157– 168, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093155/.

vanderHoeven:1984:CCS

[vdHM84] G. F. van der Hoeven and I. Moerdijk. Constructing choice sequences from lawless sequences of neighbourhood functions. Lecture Notes in Mathematics, 1103:207–234, 1984. CODEN LN-MAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099388/.

vanderKallen:1981:SDR

[vdK81] Wilberd van der Kallen. Stability for K_{2} of Dedekind rings of arithmetic type. Lecture Notes in Mathematics, 854:217– 248, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089523/.

vanderKallen:1982:DBW

[vdK82] Wilberd van der Kallen. SL_{3}(C[X]) does not have bounded word length. Lecture Notes in Mathematics, 966:357–361, 1982. CO-DEN LNMAA2. ISBN 3-540-11965-5 (print), 3-540-39553-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062183/.

vanderLinden:1984:ERI

[vdL84] F. J. van der Linden. Euclidean rings of integers of fourth degree fields. Lecture Notes in Mathematics, 1068:139–148, 1984. CODEN LNMAA2. ISBN 3-540-13356-9 (print), 3-540-38906-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099448/.

vanderLaan:1981:LRO

[vdLT81] G. van der Laan and A. J. J. Talman. Labelling rules and orientation: On Sperner’s lemma and Brouwer degree. Lecture Notes in Mathematics, 878:238–257, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090684/.

vanDulst:1983:FPN

[vDS83] D. van Dulst and Brailey Sims. Fixed points of nonexpansive mappings and Chebyshev centers in Banach spaces with norms of type (KK). Lecture Notes in Mathematics, 991:35–43, 1983. CODEN LNMAA2. ISBN 3-540-12298-2 (print), 3-540-39877-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061556/.

vonderTwer:1981:SRM

[vdT81] T. von der Twer. Some remarks on the mathematical incompleteness of Peano’s arithmetic found by Paris and Harrington. Lecture Notes in Mathematics, 872:157–174, 1981. CODEN LN-MAA2. ISBN 3-540-10849-1 (print), 3-540-38757-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098622/.

Veksler:1984:NZD

[Vek84] A. I. Veksler. Nettings in zero-dimensional compact spaces. Lecture Notes in Mathematics, 1060:105–114, 1984. CODEN LN-MAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099926/.

Venema:1981:EST

[Ven81] Gerard A. Venema. Embeddings in shape theory. Lecture Notes in Mathematics, 870:169–185, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089715/.

Verhulst:1980:PSC

[Ver80a] F. Verhulst. Periodic solutions of continuous self-gravitating systems. Lecture Notes in Mathematics, 810:318–339, 1980. CODEN LNMAA2. ISBN 3-540-10018-0 (print), 3-540-38166-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089986/.

Verschoren:1980:BPR

[Ver80b] A. Verschoren. Birationality of P.I. rings and non-commutative varieties. Lecture Notes in Mathematics, 825:172–197, 1980. CO-DEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089130/.

Verdier:1981:ALS

[Ver81a] Jean-Louis Verdier. Algèbres de Lie, systèmes hamiltoniens, courbes algébriques [d’après M. Adler et P. van Moerbeke]. (French) []. Lecture Notes in Mathematics, 901:85–94, 1981. CO-DEN LNMAA2. ISBN 3-540-11176-X (print), 3-540-38956-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097191/.

Verschoren:1981:PGA

[Ver81b] A. Verschoren. Pour une géométrie algébrique non-commutative. (French) []. Lecture Notes in Mathematics, 867:319–350, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090391/.

Verschoren:1982:CLB

[Ver82a] A. Verschoren. A check list on Brauer groups. Lecture Notes in Mathematics, 917:279–300, 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092241/.

Verschoren:1982:BGQ

[Ver82b] Alain Verschoren. The Brauer group of a quasi affine-scheme. Lecture Notes in Mathematics, 917:260–278, 1982. CODEN LN-MAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092240/.

Verdier:1983:GMF

[Ver83a] J.-L. Verdier. Géométrie microlocale. (French) []. Lecture Notes in Mathematics, 1016:127–133, 1983. CODEN LNMAA2. ISBN 3-540-12685-6 (print), 3-540-38676-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0099961/.

Verhulst:1983:AAI

[Ver83b] F. Verhulst. Asymptotic Analysis II — : Surveys and New Trends, volume 985 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 497 pp. LCCN QA3 .L28 no. 985. URL http:

//link.springer.com/book/10.1007/BFb0062359;   http:/
/www.springerlink.com/content/978-3-540-39612-3.

Verhulst:1983:AAH

[Ver83c] F. Verhulst. Asymptotic analysis of Hamiltonian systems. Lecture Notes in Mathematics, 985:137–183, 1983. CODEN LN-MAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062366/.

Verschoren:1983:RIC

[Ver83d] A. Verschoren. Relative invariants for commutative rings. Lecture Notes in Mathematics, 1029:240–266, 1983. CODEN LN-MAA2. ISBN 3-540-12699-6 (print), 3-540-38686-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098934/.

Verdera:1984:UAS

[Ver84a] Joan Verdera. Uniform approximation by solutions of higher order Cauchy–Riemann equations. Lecture Notes in Mathematics, 1094:179–184, 1984. CODEN LNMAA2. ISBN 3-540-13886-2 (print), 3-540-39096-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099162/.

Verona:1984:AS

[Ver84b] Andrei Verona. Abstract stratifications. Lecture Notes in Mathematics, 1102:2–9, 1984. CODEN LNMAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101674/.

Verona:1984:ASF

[Ver84c] Andrei Verona. Abstract stratifications (with faces). Lecture Notes in Mathematics, 1102:37–57, 1984. CODEN LN-MAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101678/.

Verona:1984:ATM

[Ver84d] Andrei Verona. Abstract Thom mappings. Lecture Notes in Mathematics, 1102:20–28, 1984. CODEN LNMAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101676/.

Verona:1984:CVF

[Ver84e] Andrei Verona. Controlled vector fields. Lecture Notes in Mathematics, 1102:9–20, 1984. CODEN LNMAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101675/.

Verona:1984:MF

[Ver84f] Andrei Verona. Manifolds with faces. Lecture Notes in Mathematics, 1102:29–37, 1984. CODEN LNMAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101677/.

Verona:1984:NC

[Ver84g] Andrei Verona. Notation and convention. Lecture Notes in Mathematics, 1102:1, 1984. CODEN LNMAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101673.

Verona:1984:SMS

[Ver84h] Andrei Verona. Stratified Mappings — Structure and Triangulability, volume 1102 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 160 pp. LCCN QA3 .L28 no. 1102. URL http:/

/link.springer.com/book/10.1007/BFb0101672;   http://
www.springerlink.com/content/978-3-540-39111-1.

Verona:1984:SAT

[Ver84i] Andrei Verona. The structure of abstract Thom mappings. Lecture Notes in Mathematics, 1102:57–120, 1984. CODEN LN-MAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101679/.

Verona:1984:TAS

[Ver84j] Andrei Verona. Triangulation of abstract stratifications. Lecture Notes in Mathematics, 1102:120–129, 1984. CODEN LN-MAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101680/.

Verona:1984:TNA

[Ver84k] Andrei Verona. Triangulation of nice abstract Thom mappings. Lecture Notes in Mathematics, 1102:129–141, 1984. CODEN LNMAA2. ISBN 3-540-13898-6 (print), 3-540-39111-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101681/.

Vershik:1984:CNC

[Ver84l] A. M. Vershik. Classical and non-classical dynamics with constraints. Lecture Notes in Mathematics, 1108:278–301, 1984. CO-DEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099563/.

Vesley:1981:IIC

[Ves81] Richard Vesley. An intuitionistic infinitesimal calculus. Lecture Notes in Mathematics, 873:208–212, 1981. CODEN LN-MAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090735/.

Vesentini:1983:GSL

[Ves83] Edoardo Vesentini, editor. Geometry Seminar “Luigi Bianchi”: Lectures given at the Scuola Normale Superiore, 1982, volume 1022 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNMAA2. ISBN 3-540-12719-4 (print), 3-540-38702-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1022. URL http:/

/link.springer.com/book/10.1007/BFb0071599;   http://
www.springerlink.com/content/978-3-540-38702-2.

vanHarten:1982:ASP

[vH82] A. van Harten. Applications of singular perturbation techniques to combustion theory. Lecture Notes in Mathematics, 942:295– 308, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094754/.

vanHarten:1983:IRP

[vH83a] A. van Harten. Introduction to the research papers. Lecture Notes in Mathematics, 985:348, 1983. CODEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062375.

vanHarten:1983:IFS

[vH83b] A. van Harten. On the interaction of flames and sound. Lecture Notes in Mathematics, 985:327–346, 1983. CODEN LN-MAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062374/.

vanHarn:1981:SDD

[vHSV81] K. van Harn, F. W. Steutel, and W. Vervaat. Self-decomposable discrete distributions and branching processes. Lecture Notes in Mathematics, 861:60–64, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097314/.

vanIseghem:1984:MRA

[vI84] Jeannette van Iseghem. Multipoints rational approximants. Lecture Notes in Mathematics, 1071:157–165, 1984. CODEN LN-MAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099616/.

Vidal:1980:DCL

[Vid80] Robert Vidal. Derivations d’un corps local à corps residuel de caractéristique nulle et algébriquement clos. (French) []. Lecture Notes in Mathematics, 795:209–224, 1980. CODEN LN-MAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090121/.

Vigneras:1980:AQCa

[Vig80a] Marie-France Vignéras. Algèbres de quaternions sur un corps. (French) []. Lecture Notes in Mathematics, 800:1–30, 1980. CO-DEN LNMAA2. ISBN 3-540-09983-2 (print), 3-540-39237-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091028/.

Vigneras:1980:AQCc

[Vig80b] Marie-France Vignéras. Algèbres de quaternions sur un corps global. (French) []. Lecture Notes in Mathematics, 800:57–102, 1980. CODEN LNMAA2. ISBN 3-540-09983-2 (print), 3-540-39237-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091030/.

Vigneras:1980:AQCb

[Vig80c] Marie-France Vignéras. Algèbres de quaternions sur un corps local. (French) []. Lecture Notes in Mathematics, 800:31–56, 1980. CODEN LNMAA2. ISBN 3-540-09983-2 (print), 3-540-39237-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091029/.

Vigneras:1980:AGA

[Vig80d] Marie-France Vignéras. Application aux groupes arithmétiques. (French) []. Lecture Notes in Mathematics, 800:103–137, 1980. CODEN LNMAA2. ISBN 3-540-09983-2 (print), 3-540-39237-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091031/.

Vigneras:1980:AAQ

[Vig80e] Marie-France Vignéras. Arithmétique des Algèbres de Quaternions, volume 800 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09983-2 (print), 3-540-39237-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 169 pp. LCCN QA3 .L28 no. 800. URL http:

//link.springer.com/book/10.1007/BFb0091027;   http:/
/www.springerlink.com/content/978-3-540-39237-8.

Vigneras:1980:AQQ

[Vig80f] Marie-France Vignéras. Arithmétique des quaternions, quand la condition d’Eichler n’est plus vérifiée. (French) []. Lecture Notes in Mathematics, 800:138–156, 1980. CODEN LNMAA2. ISBN 3-540-09983-2 (print), 3-540-39237-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091032/.

Vigue:1983:DBS

[Vig83] Jean-Pierre Vigué. Domaines bornés symétriques. (French) []. Lecture Notes in Mathematics, 1022:125–177, 1983. CODEN LNMAA2. ISBN 3-540-12719-4 (print), 3-540-38702-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071602/.

Vinogradov:1984:CND

[Vin84a] A. M. Vinogradov. Category of nonlinear differential equations. Lecture Notes in Mathematics, 1108:77–102, 1984. CODEN LN-MAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099553/.

Vinti:1984:NAO

[Vin84b] Calogero Vinti, editor. Nonlinear Analysis and Optimization: Proceedings of the International Conference held in Bologna, Italy, May 3–7, 1982, volume 1107 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13903-6 (print), 3-540-39123-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1107. URL http:/

/link.springer.com/book/10.1007/BFb0101489;   http://
www.springerlink.com/content/978-3-540-39123-4.

vanLeeuwen:1980:RCC

[vL80] L. C. A. van Leeuwen. Radicals and chain conditions. Lecture Notes in Mathematics, 825:153–157, 1980. CODEN LN-MAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089128/.

vanLeeuwen:1981:CFB

[vL81] L. C. A. van Leeuwen. Compositions of a finite band with respect to an idempotent. Lecture Notes in Mathematics, 855:189– 210, 1981. CODEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089816/.

vanLeeuwen:1983:RMR

[vL83a] L. C. A. van Leeuwen. Remarks on a monoid radical theory. Lecture Notes in Mathematics, 998:333–338, 1983. CODEN LN-MAA2. ISBN 3-540-12321-0 (print), 3-540-40051-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062039/.

Vaughan-Lee:1983:NPV

[VL83b] M. R. Vaughan-Lee. Nilpotence in permutable varieties. Lecture Notes in Mathematics, 1004:293–308, 1983. CODEN LN-MAA2. ISBN 3-540-12329-6 (print), 3-540-40954-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063445/.

Vladimirov:1982:SCV

[Vla82] V. S. Vladimirov. Several complex variables in mathematical physics. Lecture Notes in Mathematics, 919:358–386, 1982. CO-DEN LNMAA2. ISBN 3-540-11482-3 (print), 3-540-39147-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097057/.

Voss:1983:PFP

[VLA83] R. F. Voss, R. B. Laibowitz, and E. I. Alessandrini. Percolation and fractal properties of thin gold films. Lecture Notes in Mathematics, 1035:153–168, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073258/.

vanOystaeyen:1980:RTA

[vO80] F. van Oystaeyen, editor. Ring Theory Antwerp 1980: Proceedings, University of Antwerp U.I.A. Antwerp, Belgium, May 6–9, 1980, volume 825 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 825. URL http:/

/link.springer.com/book/10.1007/BFb0089114;   http://
www.springerlink.com/content/978-3-540-38334-5.

Vogan:1980:SID

[Vog80] David A. Vogan, Jr. The size of infinite dimensional representations. Lecture Notes in Mathematics, 795:172–178, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090118/.

Vogel:1982:LNC

[Vog82] Pierre Vogel. Localisation non commutative de formes quadratiques. (French) []. Lecture Notes in Mathematics, 967:376– 389, 1982. CODEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061912/.

Vogan:1983:UUD

[Vog83] David A. Vogan, Jr. Understanding the unitary dual. Lecture Notes in Mathematics, 1024:264–286, 1983. CODEN LN-MAA2. ISBN 3-540-12725-9 (print), 3-540-38709-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071435/.

Vogell:1984:CIA

[Vog84] Wolrad Vogell. The canonical involution on the algebraic Ktheory of spaces. Lecture Notes in Mathematics, 1051:156–172, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075566/.

Volberg:1981:TTF

[Vol81] A. L. Vol’berg. Thin and thick families of rational fractions. Lecture Notes in Mathematics, 864:440–480, 1981. CODEN LN-MAA2. ISBN 3-540-10695-2 (print), 3-540-38626-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097004/.

Volcic:1982:LDI

[Vol82a] Alessio Volčič. Liftings and Daniell integrals. Lecture Notes in Mathematics, 945:180–186, 1982. CODEN LNMAA2. ISBN 3-540-11580-3 (print), 3-540-39324-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096673/.

Volcic:1982:SCM

[Vol82b] Aljoša Volčič. On some counterexamples in measure theory. Lecture Notes in Mathematics, 948:235–239, 1982. CODEN LN-MAA2. ISBN 3-540-11594-3 (print), 3-540-39356-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069848/.

Volcic:1984:WPG

[Vol84] A. Volčič. Well-posedness of the Gardner–McMullen reconstruction problem. Lecture Notes in Mathematics, 1089:199–210, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072615/.

Voros:1982:ZFQ

[Vor82] A. Voros. Zeta functions of the quartic (and homogeneous anharmonic) oscillators. Lecture Notes in Mathematics, 925:184– 208, 1982. CODEN LNMAA2. ISBN 3-540-11483-1 (print), 3-540-39152-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093509/.

vanOystaeyen:1981:AV

[vOV81a] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Algebraic varieties. Lecture Notes in Mathematics, 887:227–269, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091737/.

vanOystaeyen:1981:BQ

[vOV81b] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Birationality and quasivarieties. Lecture Notes in Mathematics, 887: 334–351, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091741/.

vanOystaeyen:1981:CQS

[vOV81c] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Coherent and quasicoherent sheaves of modules over an algebraic k-variety. Lecture Notes in Mathematics, 887:270–294, 1981. CO-DEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091738/.

vanOystaeyen:1981:FTL

[vOV81d] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. The finishing touch on localization. Lecture Notes in Mathematics, 887:119–169, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091735/.

vanOystaeyen:1981:G

[vOV81e] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Generalities. Lecture Notes in Mathematics, 887:9–35, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091732/.

vanOystaeyen:1981:GR

[vOV81f] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Graded rings. Lecture Notes in Mathematics, 887:100–118, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091734/.

vanOystaeyen:1981:I

[vOV81g] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Introduction. Lecture Notes in Mathematics, 887:1–7, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091731/.

vanOystaeyen:1981:NCA

[vOV81h] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Non-Commutative Algebraic Geometry: An Introduction, volume 887 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 404 pp. LCCN QA3 .L28 no. 887. URL http:

//link.springer.com/book/10.1007/BFb0091730;   http:/
/www.springerlink.com/content/978-3-540-38601-8.

vanOystaeyen:1981:NCV

[vOV81i] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. A non-commutative version of the Riemann–Roch theorem for curves. Lecture Notes in Mathematics, 887:352–381, 1981. CO-DEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091742/.

vanOystaeyen:1981:PSE

[vOV81j] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Products, subvarieties etc …. Lecture Notes in Mathematics, 887: 295–325, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091739/.

vanOystaeyen:1981:RTR

[vOV81k] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Representation theory revisited. Lecture Notes in Mathematics, 887: 326–333, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091740/.

vanOystaeyen:1981:SNA

[vOV81l] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Some noncommutative algebra. Lecture Notes in Mathematics, 887: 36–99, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091733/.

vanOystaeyen:1981:SSS

[vOV81m] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Structure sheaves and schemes. Lecture Notes in Mathematics, 887: 170–226, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091736/.

vanOystaeyen:1981:WP

[vOV81n] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren. Work in progress, [91]. Lecture Notes in Mathematics, 887:382– 388, 1981. CODEN LNMAA2. ISBN 3-540-11153-0 (print), 3-540-38601-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091743/.

vanOystaeyen:1982:BGR

[vOV82] Freddy M. J. van Oystaeyen and Alain H. M. J. Verschoren, editors. Brauer Groups in Ring Theory and Algebraic Geometry: Proceedings, University of Antwerp U.I.A., Belgium, August 17–28, 1981, volume 917 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11216-2 (print), 3-540-39057-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 917. URL http:

//link.springer.com/book/10.1007/BFb0092224;   http:/
/www.springerlink.com/content/978-3-540-39057-2.

vanStrien:1981:BCH

[vS81a] Sebastian J. van Strien. On the bifurcations creating horseshoes. Lecture Notes in Mathematics, 898:316–351, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091922/.

vanStrien:1981:SCA

[vS81b] Sebastian J. van Strien. Saddle connections of arcs of diffeomorphisms: Moduli of stability. Lecture Notes in Mathematics, 898: 352–365, 1981. CODEN LNMAA2. ISBN 3-540-11171-9 (print), 3-540-38945-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091923/.

Vust:1980:FPT

[Vus80] Thierry Vust. Foncteurs polynômiaux et théorie des invariants. (French) []. Lecture Notes in Mathematics, 795:330–340, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090126/.

VanDeuren:1980:AGR

[VV80] J.-P. Van Deuren and F. Van Dystaeyen. Arithmetically graded rings. I. Lecture Notes in Mathematics, 825:130–152, 1980. CO-DEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089127/.

VanDeuren:1981:GRR

[VVV81] J.-P. Van Deuren, J. Van Geel, and F. Van Oystaeyen. Genus and a Riemann–Roch theorem for non-commutative function fields in one variable. Lecture Notes in Mathematics, 867:295–318, 1981. CODEN LNMAA2. ISBN 3-540-10841-6 (print), 3-540-38737-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090390/.

vonWeizsacker:1980:REM

[vW80] Heinrich von Weizsäcker. Remark on extremal measure extensions. Lecture Notes in Mathematics, 794:79–80, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088215.

Vinsonhaler:1983:PICa

[VW83a] C. I. Vinsonhaler and W. J. Wickless. Projective and injective classes of completely decomposable groups. Lecture Notes in Mathematics, 1006:144–163, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21560-9_4/.

Vinsonhaler:1983:PICb

[VW83b] C. I. Vinsonhaler and W. J. Wickless. Projective and injective classes of completely decomposable groups. Lecture Notes in Mathematics, 1006:144–163, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103699/.

Wachspress:1980:NST

[Wac80] E. L. Wachspress. The numerical solution of turbulent flow problems in general geometry. Lecture Notes in Mathematics, 773: 146–163, 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094170/.

Wachspress:1982:DPE

[Wac82] E. L. Wachspress. Discrete pressure equations in incompressible flow problems. Lecture Notes in Mathematics, 909:106–115, 1982. CODEN LNMAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092964/.

Wagner:1980:SMS

[Wag80] Daniel H. Wagner. Survey of measurable selection theorems: an update. Lecture Notes in Mathematics, 794:176–219, 1980. CO-DEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088224/.

Wagreich:1982:GFD

[Wag82] Philip Wagreich. The growth function of a discrete group. Lecture Notes in Mathematics, 956:125–144, 1982. CODEN LN-MAA2. ISBN 3-540-11946-9 (print), 3-540-39528-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101514/.

Wahlbin:1982:QEP

[Wah82] Lars B. Wahlbin. A quasioptimal estimate in piecewise polynomial Galerkin approximation of parabolic problems. Lecture Notes in Mathematics, 912:230–245, 1982. CODEN LNMAA2. ISBN 3-540-11199-9 (print), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093159/.

Walsh:1981:OIP

[Wal81a] John B. Walsh. Optional increasing paths. Lecture Notes in Mathematics, 863:172–201, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091100/.

Walsh:1981:DCD

[Wal81b] John J. Walsh. Dimension, cohomological dimension, and celllike mappings. Lecture Notes in Mathematics, 870:105–118, 1981. CODEN LNMAA2. ISBN 3-540-10846-7 (print), 3-540-38749-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089711/.

Waldhausen:1982:OAT

[Wal82a] Friedhelm Waldhausen. Operations in the algebraic K-theory of spaces. Lecture Notes in Mathematics, 967:390–409, 1982. CO-DEN LNMAA2. ISBN 3-540-11966-3 (print), 3-540-39556-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061913/.

Walsh:1982:NRS

[Wal82b] John B. Walsh. A non reversible semi-martingale. Lecture Notes in Mathematics, 920:212, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092783.

Wald:1983:PMPa

[Wal83a] Burkhard Wald. On κ-products modulo μ-products. Lecture Notes in Mathematics, 1006:362–370, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103715/.

Wald:1983:PMPb

[Wal83b] Burkhard Wald. On k-products modulo μ-products. Lecture Notes in Mathematics, 1006:362–370, 1983. CODEN LNMAA2. ISBN 3-540-12335-0 (print), 3-662-21560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21560-9_20/.

Wallach:1983:AEG

[Wal83c] Nolan R. Wallach. Asymptotic expansions of generalized matrix entries of representations of real reductive groups. Lecture Notes in Mathematics, 1024:287–369, 1983. CODEN LNMAA2. ISBN 3-540-12725-9 (print), 3-540-38709-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071436/.

Wallin:1983:DMP

[Wal83d] Hans Wallin. Divergence of multipoint Padé approximation. Lecture Notes in Mathematics, 1014:246–255, 1983. CODEN LN-MAA2. ISBN 3-540-12683-X (print), 3-540-38672-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072083/.

Waldhausen:1984:ATS

[Wal84] Friedhelm Waldhausen. Algebraic K-theory of spaces, localization, and the chromatic filtration of stable homotopy. Lecture Notes in Mathematics, 1051:173–195, 1984. CODEN LN-MAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0075567/.

Wang:1981:ELC

[Wan81] Y. H. Wang. Extensions of Lukacs’ characterization of the gamma distribution. Lecture Notes in Mathematics, 861:166–177, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097323/.

Ward:1982:TTG

[War82a] R. S. Ward. Twistor techniques in gauge theories. Lecture Notes in Mathematics, 926:199–205, 1982. CODEN LNMAA2. ISBN 3-540-11497-1 (print), 3-540-39192-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092664/.

Warshauer:1982:B

[War82b] Max L. Warshauer. The boundary. Lecture Notes in Mathematics, 914:133–183, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098267/.

Warshauer:1982:C

[War82c] Max L. Warshauer. Conventions. Lecture Notes in Mathematics, 914:11, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098262.

Warshauer:1982:DAO

[War82d] Max L. Warshauer. A detailed analysis of the octagon. Lecture Notes in Mathematics, 914:226–248, 1982. CODEN LN-MAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098270/.

Warshauer:1982:GB

[War82e] Max L. Warshauer. The global boundary. Lecture Notes in Mathematics, 914:207–225, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098269/.

Warshauer:1982:I

[War82f] Max L. Warshauer. Introduction. Lecture Notes in Mathematics, 914:1–10, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098261/.

Warshauer:1982:NMO

[War82g] Max L. Warshauer. Non-maximal orders. Lecture Notes in Mathematics, 914:184–206, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098268/.

Warshauer:1982:O

[War82h] Max L. Warshauer. The octagon over Z. Lecture Notes in Mathematics, 914:249–258, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098271/.

Warshauer:1982:PG

[War82i] Max L. Warshauer. Polynomials. (German) []. Lecture Notes in Mathematics, 914:70–78, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098265/.

Warshauer:1982:WGF

[War82j] Max L. Warshauer. Witt group of a field. Lecture Notes in Mathematics, 914:79–132, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098266/.

Warshauer:1982:WGD

[War82k] Max L. Warshauer. The Witt Group of Degree k Maps and Asymmetric Inner Product Spaces, volume 914 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 269 pp. LCCN QA3 .L28 no. 914. URL http:/

/link.springer.com/book/10.1007/BFb0098260;   http://
www.springerlink.com/content/978-3-540-39027-5.

Warshauer:1982:WI

[War82l] Max L. Warshauer. Witt invariants. Lecture Notes in Mathematics, 914:47–69, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098264/.

Warshauer:1982:WR

[War82m] Max L. Warshauer. The Witt ring. Lecture Notes in Mathematics, 914:12–46, 1982. CODEN LNMAA2. ISBN 3-540-11201-4 (print), 3-540-39027-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098263/.

Washington:1981:AF

[Was81] Lawrence C. Washington. p-adic L-functions at s = 0 and s = 1. Lecture Notes in Mathematics, 899:166–170, 1981. CODEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096459/.

Watson:1980:NAP

[Wat80] G. Alistair Watson, editor. Numerical Analysis: Proceedings of the 8th Biennial Conference Held at Dundee, Scotland, June 26–29, 1979, volume 773 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 773. URL http:/

/link.springer.com/book/10.1007/BFb0094158;   http://
www.springerlink.com/content/978-3-540-38562-2.

Watson:1982:NAP

[Wat82] George Alistair Watson, editor. Numerical Analysis: Proceedings of the 9th Biennial Conference, held at Dundee, Scotland, June 23–26, 1981, volume 912 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 0-387-11199-9 (softcover), 3-540-11199-9 (softcover), 3-540-39009-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 912; QA1 .L471; QA297 .D915n 1981. URL http://www.

springerlink.com/content/978-3-540-39009-1.

Watanabe:1984:ENF

[Wat84a] Kinji Watanabe. Sur les ensembles nodaux. (French) []. Lecture Notes in Mathematics, 1090:99–106, 1984. CODEN LN-MAA2. ISBN 3-540-13873-0 (print), 3-540-39065-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101570/.

Watson:1984:NST

[Wat84b] G. A. Watson. The numerical solution of total l_{p} approximation problems. Lecture Notes in Mathematics, 1066:221–238, 1984. CODEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099527/.

Werner:1984:PAA

[WB84] Helmut Werner and Hans Josef Bünger, editors. Padé Approximation and its Applications Bad Honnef 1983: Proceedings of a Conference held at Bad Honnef, Germany March 7–10, 1983, volume 1071 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1071. URL http:/

/link.springer.com/book/10.1007/BFb0099605;   http://
www.springerlink.com/content/978-3-540-38914-9.

Weber:1981:NAS

[Web81] H. Weber. On the numerical approximation of secondary bifurcation problems. Lecture Notes in Mathematics, 878:407–425, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090690/.

Weber:1983:AIF

[Web83a] M. Weber. Analyse infinitesimale de fonctions aléatoires. (French) []. Lecture Notes in Mathematics, 976:383–465, 1983. CODEN LNMAA2. ISBN 3-540-11987-6 (print), 3-540-39458-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067988/.

Weber:1983:PSS

[Web83b] Michel Weber. Polar sets of some Gaussian processes. Lecture Notes in Mathematics, 990:204–214, 1983. CODEN LN-MAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064273/.

Weber:1984:DJL

[Web84a] Claude Weber. La démonstration de J. Levine des théorèmes de A. Plans. (French) []. Lecture Notes in Mathematics, 1051: 315–330, 1984. CODEN LNMAA2. ISBN 3-540-12902-2 (print), 3-540-38782-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0075574/.

Weber:1984:GVV

[Web84b] Hans Weber. Group- and vector-valued S-bounded contents. Lecture Notes in Mathematics, 1089:181–198, 1984. CODEN LNMAA2. ISBN 3-540-13874-9 (print), 3-540-39069-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072614/.

Weck:1981:SCS

[Wec81] Sibylle Weck. Scott convergence and Scott topology in partially ordered sets I. Lecture Notes in Mathematics, 871:372–383, 1981. CODEN LNMAA2. ISBN 3-540-10848-3 (print), 3-540-38755-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089919/.

Wedin:1983:ABS

[Wed83] Per Åke Wedin. On angles between subspaces of a finite dimensional inner product space. Lecture Notes in Mathematics, 973:263–285, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062107/.

Weit:1980:SAL

[Wei80] Yitzhak Weit. On spectral analysis in locally compact groups. Lecture Notes in Mathematics, 781:204–213, 1980. CODEN LN-MAA2. ISBN 3-540-09756-2 (print), 3-540-38632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097657/.

Weibel:1981:MVS

[Wei81a] C. A. Weibel. Meyer Vietoris sequences and module structures on NK. Lecture Notes in Mathematics, 854:466–493, 1981. CODEN LNMAA2. ISBN 3-540-10698-7 (print), 3-540-38646-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089534/.

Weinert:1981:SRQ

[Wei81b] Hanns Joachim Weinert. On special right quotient filters of semigroups. Lecture Notes in Mathematics, 855:211–221, 1981. CO-DEN LNMAA2. ISBN 3-540-10701-0 (print), 3-540-38656-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089817/.

Weiland:1982:EAN

[Wei82a] C. Weiland. Erfahrungen bei der Anwendung numerischer Verfahren zur Lösung nichtlinearer hyperbolischer Differentialgleichungssysteme. (German) []. Lecture Notes in Mathematics, 953: 172–187, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069382/.

Weinstein:1982:SC

[Wei82b] Alan Weinstein. The symplectic “category”. Lecture Notes in Mathematics, 905:45–51, 1982. CODEN LNMAA2. ISBN 3-540-11197-2 (print), 3-540-39002-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092426/.

Weinryb:1983:EED

[Wei83] Sophie Weinryb. Étude d’une équation différentielle stochastique avec temps local. (French) []. Lecture Notes in Mathematics, 986: 72–77, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068300/.

Weispfenning:1984:QED

[Wei84] Volker Weispfenning. Quantifier elimination and decision procedures for valued fields. Lecture Notes in Mathematics, 1103: 419–472, 1984. CODEN LNMAA2. ISBN 3-540-13900-1 (print), 3-540-39115-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0099397/.

Wells:1983:TGR

[Wel83] R. O. Wells, Jr. The twistor-geometric representation of classical field theories. Lecture Notes in Mathematics, 1037:142–169, 1983. CODEN LNMAA2. ISBN 3-540-12710-0 (print), 3-540-38695-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073172/.

Weron:1980:PTV

[Wer80] A. Weron, editor. Probability Theory on Vector Spaces II: Proceedings, Błażejewko, Poland, September 17–23, 1979, volume 828 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 0-387-10253-1, 3-540-10253-1 (print), 3-540-38350-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 828; QA3 .L28 no. 828 QA273.43.

URL http://link.springer.com/book/10.1007/BFb0097389;
 http://www.springerlink.com/content/978-3-540-38350-
5.

Wermer:1981:ARD

[Wer81a] John Wermer. Applications of the Riesz decomposition. Lecture Notes in Mathematics, 408:129–137, 1981. CODEN LN-MAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069706/.

Wermer:1981:B

[Wer81b] John Wermer. Barriers. Lecture Notes in Mathematics, 408:100– 107, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069703/.

Wermer:1981:C

[Wer81c] John Wermer. Capacity. Lecture Notes in Mathematics, 408: 26–33, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069693/.

Wermer:1981:CC

[Wer81d] John Wermer. The cone condition. Lecture Notes in Mathematics, 408:60–65, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069698/.

Wermer:1981:Ea

[Wer81e] John Wermer. Electrostatics. Lecture Notes in Mathematics, 408: 1–10, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069690/.

Wermer:1981:Eb

[Wer81f] John Wermer. Energy. Lecture Notes in Mathematics, 408:34– 40, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069694/.

Wermer:1981:EED

[Wer81g] John Wermer. Existence of the equilibrium distribution. Lecture Notes in Mathematics, 408:41–49, 1981. CODEN LN-MAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069695/.

Wermer:1981:FS

[Wer81h] John Wermer. Fundamental solutions. Lecture Notes in Mathematics, 408:17–25, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069692/.

Wermer:1981:GF

[Wer81i] John Wermer. Green’s function. Lecture Notes in Mathematics, 408:74–83, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069700/.

Wermer:1981:KTb

[Wer81j] John Wermer. Kellogg’s theorem. Lecture Notes in Mathematics, 408:108–113, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069704/.

Wermer:1981:KTa

[Wer81k] John Wermer. The Kelvin transform. Lecture Notes in Mathematics, 408:84–90, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069701/.

Wermer:1981:MPP

[Wer81l] John Wermer. Maximum principle for potentials. Lecture Notes in Mathematics, 408:50–55, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069696/.

Wermer:1981:PM

[Wer81m] John Wermer. Perron’s method. Lecture Notes in Mathematics, 408:91–99, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069702/.

Wermer:1981:PE

[Wer81n] John Wermer. Poisson’s equation. Lecture Notes in Mathematics, 408:11–16, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069691/.

Wermer:1981:PT

[Wer81o] John Wermer. Potential Theory, volume 408 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CO-DEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 166 pp. LCCN QA3 .L28 no. 408. URL http:/

/link.springer.com/book/10.1007/BFb0069689;   http://
www.springerlink.com/content/978-3-540-38407-6.

Wermer:1981:RDT

[Wer81p] John Wermer. The Riesz decomposition theorem. Lecture Notes in Mathematics, 408:114–128, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069705/.

Wermer:1981:SBH

[Wer81q] John Wermer. Singularities of bounded harmonic functions. Lecture Notes in Mathematics, 408:66–73, 1981. CODEN LN-MAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069699/.

Wermer:1981:UED

[Wer81r] John Wermer. Uniqueness of the equilibrium distribution. Lecture Notes in Mathematics, 408:56–59, 1981. CODEN LN-MAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069697/.

Wermer:1981:WC

[Wer81s] John Wermer. Wiener’s criterion. Lecture Notes in Mathematics, 408:138–157, 1981. CODEN LNMAA2. ISBN 3-540-10276-0 (print), 3-540-38407-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069707/.

Werner:1981:NLS

[Wer81t] Helmut Werner. Non-linear splines, some applications to singular problems. Lecture Notes in Mathematics, 888:64–77, 1981. CODEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095576/.

Werner:1981:SIC

[Wer81u] W. Werner. Some improvements of classical iterative methods for the solution of nonlinear equations. Lecture Notes in Mathematics, 878:426–440, 1981. CODEN LNMAA2. ISBN 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090691/.

Werner:1982:SDP

[Wer82] Wilhelm Werner. On the simultaneous determination of polynomial roots. Lecture Notes in Mathematics, 953:188–202, 1982. CODEN LNMAA2. ISBN 3-540-11602-8 (print), 3-540-39379-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069383/.

Wesseling:1980:RCM

[Wes80] P. Wesseling. The rate of convergence of a multiple grid method. Lecture Notes in Mathematics, 773:164–184, 1980. CODEN LN-MAA2. ISBN 3-540-09740-6 (print), 3-540-38562-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094171/.

Wesseling:1982:REM

[Wes82] P. Wesseling. A robust and efficient multigrid method. Lecture Notes in Mathematics, 960:614–630, 1982. CODEN LN-MAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069947/.

Wets:1983:FLS

[Wet83] Roger J.-B. Wets. A formula for the level sets of epi-limits and some applications. Lecture Notes in Mathematics, 979:256– 268, 1983. CODEN LNMAA2. ISBN 3-540-11999-X (print), 3-540-39473-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066260/.

Wheeler:1980:EAM

[Whe80] Robert F. Wheeler. Extensions of a σ-additive measure to the projective cover. Lecture Notes in Mathematics, 794:81–104, 1980. CODEN LNMAA2. ISBN 3-540-09979-4 (print), 3-540-39221-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088216/.

Whittington:1983:BPP

[Whi83] S. G. Whittington. Branched polymers, polymer gels and percolation theory. Lecture Notes in Mathematics, 1035:283–299, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073264/.

Whiteman:1984:STD

[Whi84] J. R. Whiteman. Singularities in three-dimensional elliptic problems and their treatment with finite element methods. Lecture Notes in Mathematics, 1066:264–275, 1984. CODEN LN-MAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0099529/.

Wierman:1983:CPP

[Wie83] John C. Wierman. Critical probabilities in percolation models. Lecture Notes in Mathematics, 1035:300–313, 1983. CODEN LNMAA2. ISBN 3-540-12707-0 (print), 3-540-38693-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073265/.

Wilkie:1980:ACT

[Wil80a] A. J. Wilkie. Applications of complexity theory to Σ_{o}-definability problems in arithmetic. Lecture Notes in Mathematics, 834:363– 369, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090174/.

Wilmers:1980:MSM

[Wil80b] George Wilmers. Minimally saturated models. Lecture Notes in Mathematics, 834:370–380, 1980. CODEN LNMAA2. ISBN 3-540-10269-8 (print), 3-540-38393-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090175/.

Wilkie:1981:DOR

[Wil81a] A. J. Wilkie. On discretely ordered rings in which every definable ideal is principal. Lecture Notes in Mathematics, 890:297– 303, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095670/.

Williams:1981:SIP

[Wil81b] David Williams, editor. Stochastic Integrals: Proceedings of the LMS Durham Symposium, July 7–17, 1980, volume 851 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 851.

URL http://link.springer.com/book/10.1007/BFb0088719;
 http://www.springerlink.com/content/978-3-540-38613-
1.

Williams:1981:BBD

[Wil81c] David Williams. “to begin at the beginning: … ”. Lecture Notes in Mathematics, 851:1–55, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088721/.

Wilmers:1981:OCR

[Wil81d] George Wilmers. An observation concerning the relationship between finite and infinitary Σ_{1}^{1}. Lecture Notes in Mathematics, 890:304–306, 1981. CODEN LNMAA2. ISBN 3-540-11159-X (print), 3-540-38629-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095671/.

Wilson:1981:CPS

[Wil81e] Robert Lee Wilson. The classification problem for simple Lie algebras of characteristic p. Lecture Notes in Mathematics, 848: 13–32, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090554/.

Wilson:1981:NPK

[Wil81f] Robert Lee Wilson. A “new” proof of Kaplansky’s theorem on simple Lie algebras of rank one. Lecture Notes in Mathematics, 848:1–12, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090553/.

Wilson:1981:EMT

[Wil81g] S. M. J. Wilson. An extension of Miyata’s Theorem on the transfer map from the classgroup of a finite dihedral group to that of its cyclic maximal subgroup. Lecture Notes in Mathematics, 882: 247–250, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092497/.

Williams:1982:PTN

[Wil82a] David Williams. A ‘potential-theoretic’ note on the quadratic Wiener–Hopf equation for Q-matrices. Lecture Notes in Mathematics, 920:91–94, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092771/.

Wilson:1982:SQC

[Wil82b] P. M. H. Wilson. Some questions on the canonical ring of threefolds of general type. Lecture Notes in Mathematics, 947:251– 268, 1982. CODEN LNMAA2. ISBN 3-540-11587-0 (print), 3-540-39342-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093594/.

Wilson:1982:ELA

[Wil82c] Robert Lee Wilson. Eclidean Lie algebras are universal central extensions. Lecture Notes in Mathematics, 933:210–213, 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093364/.

Williams:1983:SCL

[Wil83a] Floyd L. Williams. Solution of a conjecture of Langlands. Lecture Notes in Mathematics, 1020:179–187, 1983. CODEN LN-MAA2. ISBN 3-540-12717-8 (print), 3-540-38700-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071502/.

Willis:1983:NPF

[Wil83b] G. A. Willis. The norms of powers of functions in the Volterra algebra. Lecture Notes in Mathematics, 975:280–281, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064558.

Williamson:1984:PAN

[Wil84] R. A. Williamson. Padé approximations in the numerical solution of hyperbolic differential equations. Lecture Notes in Mathematics, 1071:252–264, 1984. CODEN LNMAA2. ISBN 3-540-13364-X (print), 3-540-38914-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099623/.

Wimp:1980:USF

[Wim80] Jet Wimp. Uniform scale functions and the asymptotic expansion of integrals. Lecture Notes in Mathematics, 827:251–271, 1980. CODEN LNMAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091383/.

Winter:1981:CTS

[Win81] David J. Winter. A combinatorial theory of symmetry and applications to Lie algebras. Lecture Notes in Mathematics, 848: 41–62, 1981. CODEN LNMAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090556/.

Winter:1982:LAR

[Win82a] David Winter, editor. Lie Algebras and Related Topics: Proceedings of a Conference Held at New Brunswick, New Jersey, May 29–31, 1981, volume 933 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 933. URL http:

//link.springer.com/book/10.1007/BFb0093349;   http:/
/www.springerlink.com/content/978-3-540-39262-0.

Winter:1982:FJS

[Win82b] David J. Winter. The fitting and Jordan structure of affine semigroups. Lecture Notes in Mathematics, 933:214–232, 1982. CO-DEN LNMAA2. ISBN 3-540-11563-3 (print), 3-540-39262-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093365/.

Wissbrun:1984:ODL

[Wis84] Kurt F. Wissbrun. Orientation development in liquid crystal polymers. Lecture Notes in Mathematics, 1063:1–26, 1984. CODEN LNMAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072149/.

Wexler-Kreindler:1980:SPS

[WK80a] Elena Wexler-Kreindler. Skew power series rings and some homological properties of filtered rings. Lecture Notes in Mathematics, 825:198–209, 1980. CODEN LNMAA2. ISBN 3-540-10246-9 (print), 3-540-38334-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089131/.

Wexler-Kreindler:1980:DPM

[WK80b] Elena Wexler-Kreindler. Sur la dimension projective des modules filtres sur des anneaux filtres complets. (French) []. Lecture Notes in Mathematics, 795:225–250, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090122/.

Wold:1983:MCP

[WMW83] S. Wold, H. Martens, and H. Wold. The multivariate calibration problem in chemistry solved by the PLS method. Lecture Notes in Mathematics, 973:286–293, 1983. CODEN LNMAA2. ISBN 0-387-11983-3, 3-540-11983-3 (print), 3-540-39447-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062108/.

Wohlhauser:1983:RAG

[Woh83] Alfred Wohlhauser. Raumwinkelkonforme Abbildungen. (German) [Solid angle conformal mapping]. Lecture Notes in Mathematics, 1013:225–231, 1983. CODEN LNMAA2. ISBN 3-540-12682-1 (print), 3-540-38671-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066531/.

Wolf:1980:RRI

[Wol80] Joseph A. Wolf. Representations that remain irreducible on parabolic subgroups. Lecture Notes in Mathematics, 836:129– 144, 1980. CODEN LNMAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089729/.

Wolfe:1981:UID

[Wol81a] Stephen James Wolfe. On the unimodality of infinitely divisible distribution functions II. Lecture Notes in Mathematics, 861: 178–183, 1981. CODEN LNMAA2. ISBN 3-540-10823-8 (print), 3-540-36785-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097324/.

Wolff:1981:SPS

[Wol81b] Karl Erich Wolff. Strong point stable designs. Lecture Notes in Mathematics, 893:242–250, 1981. CODEN LNMAA2. ISBN 3-540-11166-2 (print), 3-540-38639-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091025/.

Wolff:1982:NIS

[Wol82a] Thomas H. Wolff. A note on interpolation spaces. Lecture Notes in Mathematics, 908:199–204, 1982. CODEN LNMAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093288/.

Wolniewicz:1982:NIL

[Wol82b] L. Wolniewicz. Numerical integration of linear inhomogeneous ordinary differential equations appearing in the nonadiabatic theory of small molecules. Lecture Notes in Mathematics, 968:82– 94, 1982. CODEN LNMAA2. ISBN 3-540-11970-1 (print), 3-540-39374-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064883/.

Wolinska:1983:PD

[Wol83] A. Wolińska. On a problem of Dugué. Lecture Notes in Mathematics, 982:244–253, 1983. CODEN LNMAA2. ISBN 3-540-12278-8 (print), 3-540-39598-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082075/.

Woodin:1983:AUS

[Woo83a] W. Hugh Woodin. Ad and the uniqueness of the supercompact measures on P ω_{1}(λ). Lecture Notes in Mathematics, 1019:67– 71, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071694/.

Woodin:1983:SCR

[Woo83b] W. Hugh Woodin. Some consistency results in ZFC using AD. Lecture Notes in Mathematics, 1019:172–198, 1983. CODEN LNMAA2. ISBN 3-540-12688-0 (print), 3-540-38680-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071701/.

Woyczynski:1982:ABM

[Woy82] Wojbor A. Woyczyński. Asymptotic behavior of martingales in Banach spaces II. Lecture Notes in Mathematics, 939:216– 225, 1982. CODEN LNMAA2. ISBN 3-540-11569-2 (print), 3-540-39284-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096271/.

Woyczynski:1983:SAB

[Woy83] Wojbor A. Woyczyński. Survey of asymptotic behavior of sums of independent random vectors and general martingales in Banach spaces. Lecture Notes in Mathematics, 990:215–220, 1983. CODEN LNMAA2. ISBN 3-540-12295-8 (print), 3-540-39870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064274/.

Waelbroeck:1984:SAS

[WSF^{+}84] L. Waelbroeck, Allen L. Shields, Richard Frankfurt, Boris Korenblum, Frank Forelli, T. Wolff, David L. Williams, J. Bruna, Yngve Domar, Marc Thomas, J.-P. Kahane, V. P. Gurarii, V. P. Palamodov, V. M. Trutnev, V. A. Tkachenko, et al. Spectral analysis and synthesis. Lecture Notes in Mathematics, 1043:353–430, 1984. CODEN LNMAA2. ISBN 3-540-12869-7 (print), 3-540-38758-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072190/.

Wang:1984:BNP

[WTL84] Jianfang Wang, Songlin Tian, and Jiuqiang Liu. The binding number of product graphs. Lecture Notes in Mathematics, 1073: 119–128, 1984. CODEN LNMAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073110/.

Wuytack:1981:CPA

[Wuy81] Luc Wuytack. On the conditioning of the Padé approximation problem. Lecture Notes in Mathematics, 888:78–89, 1981. CO-DEN LNMAA2. ISBN 3-540-11154-9 (print), 3-540-38606-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095577/.

Yammine:1980:LIS

[Yam80] Sleiman Yammine. Localisation des idéaux semi-premiers et extension des scalaires dans les algèbres noetheriennes sur un corps. (French) []. Lecture Notes in Mathematics, 795:251–290, 1980. CODEN LNMAA2. ISBN 3-540-09980-8 (print), 3-540-39230-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090123/.

Yammine:1982:IPA

[Yam82] Sleiman Yammine. Idéaux primitifs dans des algèbres universelles. (French) []. Lecture Notes in Mathematics, 924:148–156, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092929/.

Yan:1980:RPN

[Yan80] Jia-An Yan. Remarques sur l’I.S. de processus non bornés. (French) []. Lecture Notes in Mathematics, 784:148–151, 1980. CODEN LNMAA2. ISBN 3-540-09760-0 (print), 3-540-38642-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089480/.

Yan:1983:TKS

[Yan83a] J. A. Yan. Sur un théorème de Kazamaki–Sekiguchi. (French) []. Lecture Notes in Mathematics, 986:121–122, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068306.

Yan:1983:RSF

[Yan83b] Jia-An Yan. Une remarque sur les solutions faibles des équations différentielles stochastiques unidimensionnelles. (French) []. Lecture Notes in Mathematics, 986:78–80, 1983. CODEN LN-MAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068301/.

Yap:1984:CCG

[Yap84] H. P. Yap. Computational complexity of graph properties. Lecture Notes in Mathematics, 1073:35–54, 1984. CODEN LN-MAA2. ISBN 3-540-13368-2 (print), 3-540-38924-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073104/.

Yasugi:1981:HBT

[Yas81] Mariko Yasugi. The Hahn–Banach theorem and a restricted inductive definition. Lecture Notes in Mathematics, 891:359– 394, 1981. CODEN LNMAA2. ISBN 3-540-11161-1 (print), 3-540-38633-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090989/.

Chen:1981:CSF

[yCV81] Bang yen Chen and Leopold Verstraelen. Characterizations of space forms by hypersurfaces. Lecture Notes in Mathematics, 838:105–109, 1981. CODEN LNMAA2. ISBN 3-540-10285-X (print), 3-540-38419-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088846/.

Yor:1981:TDF

[Yor81] M. Yor. Sur la théorie du filtrage. (French) []. Lecture Notes in Mathematics, 876:239–280, 1981. CODEN LNMAA2. ISBN 3-540-10860-2 (print), 3-540-38774-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097500/.

Yor:1983:DBC

[Yor83] Marc Yor. Le drap brownien comme limite en loi de temps locaux linéaires. (French) []. Lecture Notes in Mathematics, 986: 89–105, 1983. CODEN LNMAA2. ISBN 3-540-12289-3 (print), 3-540-39614-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068303/.

Yor:1984:SRB

[Yor84] Marc Yor. On square-root boundaries for Bessel processes, and pole-seeking Brownian motion. Lecture Notes in Mathematics, 1095:100–107, 1984. CODEN LNMAA2. ISBN 3-540-13891-9 (print), 3-540-39103-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099124/.

Ytrehus:1984:NHS

[Ytr84] Tor Ytrehus. A nonlinear half-space problem in the kinetic theory of gases. Lecture Notes in Mathematics, 1048:221–243, 1984.

CODEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071884/.

Yessenin-Volpin:1981:AIF

[YV81] A. S. Yessenin-Volpin. About infinity, finiteness and finitization (in connection with the foundations of mathematics). Lecture Notes in Mathematics, 873:274–313, 1981. CODEN LN-MAA2. ISBN 3-540-10850-5 (print), 3-540-38759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090740/.

Zafran:1982:MOC

[Zaf82] Misha Zafran. Multipliers of L_{p} and the operational calculus. Lecture Notes in Mathematics, 908:228–246, 1982. CODEN LN-MAA2. ISBN 3-540-11188-3 (print), 3-540-38973-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093292/.

Zaharov:1984:TCH

[Zah84] V. K. Zaharov. Topological characterization of the hyperstonean cover. Lecture Notes in Mathematics, 1060:115–118, 1984. CO-DEN LNMAA2. ISBN 3-540-13337-2 (print), 3-540-38863-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0099927/.

Zakai:1981:SRI

[Zak81] Moshe Zakai. Some remarks on integration with respect to weak martingales. Lecture Notes in Mathematics, 863:149–161, 1981. CODEN LNMAA2. ISBN 3-540-10832-7 (print), 3-540-38718-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091098/.

Zayed:1982:CAR

[Zay82] Maher Zayed. Caractérisation des algèbres de représentation finie sur des corps algébriquement clos. (French) []. Lecture Notes in Mathematics, 924:129–147, 1982. CODEN LNMAA2. ISBN 3-540-11496-3 (print), 3-540-39188-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092928/.

Zapas:1984:NSP

[ZC84] L. J. Zapas and J. M. Crissman. The necking of semicrystalline polymers in tension. Lecture Notes in Mathematics, 1063:46– 63, 1984. CODEN LNMAA2. ISBN 3-540-13340-2 (print), 3-540-38869-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072152/.

Zehnder:1983:PSH

[Zeh83] E. Zehnder. Periodic solutions of Hamiltonian equations. Lecture Notes in Mathematics, 1031:172–213, 1983. CODEN LN-MAA2. ISBN 3-540-12705-4 (print), 3-540-38691-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072117/.

Zelevinsky:1981:FA

[Zel81a] Andrey V. Zelevinsky. First applications. Lecture Notes in Mathematics, 869:71–106, 1981. CODEN LNMAA2. ISBN 3-540-10824-6 (print), 3-540-38711-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090290/.

Zelevinsky:1981:I

[Zel81b] Andrey V. Zelevinsky. Introduction. Lecture Notes in Mathematics, 869:1–11, 1981. CODEN LNMAA2. ISBN 3-540-10824-6 (print), 3-540-38711-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090288/.

Zelevinsky:1981:RFC

[Zel81c] Andrey V. Zelevinsky. Representations of Finite Classical Groups: A Hopf Algebra Approach, volume 869 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10824-6 (print), 3-540-38711-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 184 pp. LCCN QA3 .L28 no. 869.

URL http://link.springer.com/book/10.1007/BFb0090287;
 http://www.springerlink.com/content/978-3-540-38711-
4.

Zelevinsky:1981:RGL

[Zel81d] Andrey V. Zelevinsky. Representations of general linear and affine groups over finite fields. Lecture Notes in Mathematics, 869:107– 148, 1981. CODEN LNMAA2. ISBN 3-540-10824-6 (print), 3-540-38711-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090291/.

Zelevinsky:1981:STP

[Zel81e] Andrey V. Zelevinsky. Structural theory of PSH-algebras. Lecture Notes in Mathematics, 869:12–70, 1981. CODEN LN-MAA2. ISBN 3-540-10824-6 (print), 3-540-38711-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090289/.

Zeleznikow:1981:NPO

[Zel81f] John Zeleznikow. The natural partial order on semirings. Lecture Notes in Mathematics, 848:255–261, 1981. CODEN LN-MAA2. ISBN 3-540-10573-5 (print), 3-540-38549-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090571/.

Zheng:1984:RMC

[Zhe84] Weian Zheng. Une remarque sur une même I.S. calculée dans deux filtrations. (French) []. Lecture Notes in Mathematics, 1059: 172–173, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100041.

Zieschang:1981:FGMb

[Zie81a] Heiner Zieschang. Finite Groups of Mapping Classes of Surfaces, volume 875 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 3-540-10857-2 (print), 3-540-38771-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 340 pp. LCCN QA3 .L28 no. 875. URL http:/

/link.springer.com/book/10.1007/BFb0090465;   http://
www.springerlink.com/content/978-3-540-38771-8.

Zieschang:1981:MRS

[Zie81b] Heiner Zieschang. Marked Riemann surfaces and Fuchsian groups and Fenchel’s realization theorem. Lecture Notes in Mathematics, 875:140–203, 1981. CODEN LNMAA2. ISBN 3-540-10857-2 (print), 3-540-38771-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090468/.

Zieschang:1981:NCF

[Zie81c] Heiner Zieschang. The Nielsen construction of finite groups of mapping classes. Lecture Notes in Mathematics, 875:204– 253, 1981. CODEN LNMAA2. ISBN 3-540-10857-2 (print), 3-540-38771-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090469/.

Zieschang:1981:FGMa

[Zie81d] Heiner Zieschang. On finite groups of mapping classes of 3manifolds: Some applications. Lecture Notes in Mathematics, 875: 296–329, 1981. CODEN LNMAA2. ISBN 3-540-10857-2 (print), 3-540-38771-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090471/.

Zieschang:1981:NRPb

[Zie81e] Heiner Zieschang. On the Nielsen realization problem. Lecture Notes in Mathematics, 875:254–295, 1981. CODEN LN-MAA2. ISBN 3-540-10857-2 (print), 3-540-38771-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090470/.

Zieschang:1981:NRPa

[Zie81f] Heiner Zieschang. On the Nielsen realization problem, preliminaries. Lecture Notes in Mathematics, 875:1–64, 1981. CODEN LNMAA2. ISBN 3-540-10857-2 (print), 3-540-38771-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090466/.

Zieschang:1981:TCG

[Zie81g] Heiner Zieschang. Tori, crystallographic groups and the realization of mapping classes. Lecture Notes in Mathematics, 875:65– 139, 1981. CODEN LNMAA2. ISBN 3-540-10857-2 (print), 3-540-38771-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090467/.

Zheng:1984:QRF

[ZM84a] W. A. Zheng and P.-A. Meyer. Quelques résultats de ^{≪}mécanique stochastique^{≫}. (French) []. Lecture Notes in Mathematics, 1059: 223–244, 1984. CODEN LNMAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100047/.

Zheng:1984:ISN

[ZM84b] Wei-An Zheng and P.-A. Meyer. Intégrales stochastiques non monotones. (French) [Nonmonotonic stochastic integrals]. Lecture Notes in Mathematics, 1059:154–171, 1984. CODEN LN-MAA2. ISBN 3-540-13332-1 (print), 3-540-38858-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100040/.

Zouakia:1983:TCE

[Zou83] F. Zouakia. The theory of Cohen elements. Lecture Notes in Mathematics, 975:163–178, 1983. CODEN LNMAA2. ISBN 3-540-11985-X (print), 3-540-39454-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064549/.

Zuckerman:1984:QPS

[Zuc84] Gregg J. Zuckerman. Quantum physics and semisimple symmetric spaces. Lecture Notes in Mathematics, 1077:437–454, 1984. CODEN LNMAA2. ISBN 3-540-13385-2 (print), 3-540-38936-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072348/.

Zieschang:1980:DCC

[ZVC80a] Heiner Zieschang, Elmar Vogt, and Hans-Dieter Coldewey. 2dimensional complexes and combinatorial presentations of groups. Lecture Notes in Mathematics, 835:19–63, 1980. CODEN LN-MAA2. ISBN 3-540-10024-5 (print), 3-540-69981-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089694/.

Zieschang:1980:APG

[ZVC80b] Heiner Zieschang, Elmar Vogt, and Hans-Dieter Coldewey. Automorphisms of planar groups. Lecture Notes in Mathematics, 835: 178–241, 1980. CODEN LNMAA2. ISBN 3-540-10024-5 (print), 3-540-69981-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089697/.

Zieschang:1980:FGG

[ZVC80c] Heiner Zieschang, Elmar Vogt, and Hans-Dieter Coldewey. Free groups and graphs. Lecture Notes in Mathematics, 835:1–18, 1980. CODEN LNMAA2. ISBN 3-540-10024-5 (print), 3-540-69981-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089693/.

Zieschang:1980:CAT

[ZVC80d] Heiner Zieschang, Elmar Vogt, and Hans-Dieter Coldewey. On the complex analytic theory of Riemann surfaces and planar discontinuous groups. Lecture Notes in Mathematics, 835:242–286, 1980. CODEN LNMAA2. ISBN 3-540-10024-5 (print), 3-540-69981-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089698/.

Zieschang:1980:TTS

[ZVC80e] Heiner Zieschang, Elmar Vogt, and Hans-Dieter Coldewey. On the topological theory of surfaces. Lecture Notes in Mathematics, 835:287–309, 1980. CODEN LNMAA2. ISBN 3-540-10024-5 (print), 3-540-69981-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089699/.

Zieschang:1980:PDG

[ZVC80f] Heiner Zieschang, Elmar Vogt, and Hans-Dieter Coldewey. Planar discontinuous groups. Lecture Notes in Mathematics, 835: 106–177, 1980. CODEN LNMAA2. ISBN 3-540-10024-5 (print), 3-540-69981-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089696/.

Zieschang:1980:S

[ZVC80g] Heiner Zieschang, Elmar Vogt, and Hans-Dieter Coldewey. Surfaces. Lecture Notes in Mathematics, 835:64–105, 1980. CODEN LNMAA2. ISBN 3-540-10024-5 (print), 3-540-69981-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089695/.

Zieschang:1980:SPD

[ZVC80h] Heiner Zieschang, Elmar Vogt, and Hans-Dieter Coldewey. Surfaces and Planar Discontinuous Groups: Revised and Expanded Translation Translated from the German by J. Stillwell, volume 835 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNMAA2. ISBN 3-540-10024-5 (print), 3-540-69981-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 334 pp. LCCN QA3 .L28 no. 835. URL http:/

/link.springer.com/book/10.1007/BFb0089692;   http://
www.springerlink.com/content/978-3-540-69981-1.

Zvyagin:1984:TGC

[Zvy84] V. G. Zvyagin. On the theory of generalized condensing perturbations of continuous mappings. Lecture Notes in Mathematics, 1108:173–193, 1984. CODEN LNMAA2. ISBN 3-540-13910-9 (print), 3-540-39132-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0099557/.

Zweifel:1984:BEP

[Zwe84] Paul F. Zweifel. The Boltzmann equation and its properties. Lecture Notes in Mathematics, 1048:111–175, 1984. CODEN LNMAA2. ISBN 3-540-12899-9 (print), 3-540-38777-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071879/.